A Complete Bibliography of Publications in Bulletin canadien de mathématiques = Canadian Mathematical Bulletin

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
     WWW URL: https://www.math.utah.edu/~beebe/

30 November 2023 Version 2.39

Title word cross-reference

# [1443]. (−1) [2560]. (−1)r = r [1114]. (0, , 1) [873, 1307, 508, 197]. (0, , p/4) [1937]. (0, 1) [1328]. (0, 2) [1463]. (0, ∞) [3113]. (0, N) [4252]. (1, , 1) [1074]. (1, 1) [3797, 3796, 3220]. (1 − x)^{α}(1 + x)^{η} + Mδ(x + 1) + Nδ(x − 1) [2027]. (1 − x^{n})/(1 − x) [347]. SL(2, R) [3275]. (2s + 1) [2258]. (3, 1) [3394]. (4n + 3) [4535]. (a) [2437]. (α, β) [3698]. (α, η) [2567]. (L^{1})^{∗} [166]. (C, k) [1558]. (f(x))^{2} − f(x^{2}) = h(x) [58]. (G) [1097]. (G, , σ) [822]. (g, f, n) [3784]. (G, Z(G)) [4299]. (γ, r) [164]. (k + j) [3930]. (k, r) [1512]. (Z/a[rtimes]Z/b) × SL_{2} (F_{p}) [3578]. (mHM) [2005]. (mod, p) [656]. (modp) [394, 1937]. (modr) [1647]. (n, n, ∞; k) [3935]. (N − 1) [1900]. (p, q) = 1 [2180]. (PU) [1379]. (r, λ) [1529]. (sinhx)/cosx [740]. (Sp(n), O(V)) [3551]. (t, m, s) [3131]. (Z_{2})^{k} [2885]. ∗ [4085, 1379, 3687, 3566, 2930]. −D [594]. −Δ + V [3513]. −Δ_{p} − Δ_{q} [4319]. 0 [2796, 1271, 2389, 1885, 4098, 1611, 2445, 1554].

{0, 1} [3758]. 1 [4491, 728, 1271, 4390, 1432, 2445, 3243]. {1, 1 + s, 1 + s + t} [2980]. 1, ⋯, n [1875]. 1/(r + s − 1) [1158]. 10 [155, 209]. 11 [3529]. 14 [494]. 15 [3579]. 16 [1868]. 2 [4034, 4346, 3167, 3169, 2997, 3222, 2008, 1213, 2817, 2658, 1700, 4134, 2453, 4086, 3290, 982, 4517, 3892, 2843, 2403, 3693, 3511, 2248, 2522, 2433, 3857, 1223, 4342, 1935, 2901, 2494, 1575, 2159, 3760, 3389, 1339, 1389, 3699, 1932, 4130, 4057]. 2^{ℵ_{α}} [1439]. 2^{h} [3449]. 2^{h}c [3449]. 2d [4114]. \binom{2n}{n} [221]. \binom{2n}{n} = \binom{2a}{a}\binom{2b}{b} [161]. 2p [4198]. 2S^{2} [1624]. 3 [2771, 3409, 3864, 2819, 4139, 764, 2265, 1196, 2646, 4659, 2987, 4412, 2847, 3818, 3006, 2959, 2159, 2032, 4452, 4303]. 3 + 1 [3919]. 3, , 6_{b,,c} [603]. 33 [4392]. 3 × 3 [4291]. 4 [4475, 2652, 3604, 3591, 2321, 1223, 3419, 1253, 3553, 4152, 2886, 3119, 1740, 3055]. 4, , 4_{b,,c} [603]. 4m + 1 [4446]. 5 [4424, 3288, 3055]. {5, 3, 3} [211]. 5_{2} [4116]. 6 [2469, 2569, 3804]. {6, 6} [118]. 7 [1994]. 9 [709, 259, 172]. = [3066]. [(d/dt) − A]x = 0 [1859]. [−1, 1] [996]. [1/2, 1) [3598]. [a, , b] [1636]. [nθ] [1608]. [S, α_{n}] [1407]. [T, α_{n}] [1407]. ^{∗} [3370, 3615, 2948]. ^{∗} [2004, 3102]. _{0} [3361]. _{p}F_{p}(x) [3958]. A [2784, 2945, 641]. A(Ω) [1789]. A + A [1056]. A − A [1056]. A/I [3836]. A = XY Z [1137]. A^{+} [2600]. A_{∞}^{+}(g) [2982]. A^{∗} [968]. A^{p} [3019]. A_{4} [4140]. A_{n} [1362]. AF [3337]. AH [4409]. ℵ_{0} [1335]. ℵ_{α} [1439, 730]. α [4414, 2685, 3639, 3746, 2902]. α + η + 1 [2567]. AN_{2} ∧ N_{2}A [763]. A × B [574]. Aut(μ) [3931]. ax + b [4423]. ax^{2} + by^{2} ≡ c(modk) [605]. B [1305, 678, 2717, 3481, 4143]. B = ZY X [1137]. B^{∗} [1226]. B_{2} [3311]. B_{4} [2431]. B_{p_{1},p_{2}} [4218]. ¯∂ [4816]. R^{3} [3084]. β_{1} [2255]. βX − X [993]. ⋂ R(T − λ) [2497]. BMO [3100]. BP_{∗}((BZ/p^{k})^{n}) [2730]. bv_{0} [2665]. C [2131, 1821]. C(K) [3064]. C(K) [2732]. C(n) [3265]. C(N)oC(N(−1)) [2652]. C(X) [4176, 1751, 999, 668]. C∗ [4064, 4220, 4302]. c > 1 [3449]. C[0, , 1] [1270]. C[0, 1] [2948]. C^{∗} [4350, 4496, 4426, 4089, 4774, 4487, 4077, 4187, 2593, 3863, 3747, 3832, 3771, 3197, 2955, 3014, 2929, 3836, 3357, 2516, 3174, 3583, 2439, 2679, 2628, 2359, 3436, 3601, 2650, 3364, 2869]. C^{∗}(C_{λ}^{∗}(G), C_{h}^{∗}o(G)) [2485]. C_{λ}^{∗}(G) ⊗_{max} C_{h}^{∗}o(G) [2485]. C^{1} [3592]. C^{2} [3221, 3246]. C^{∗} [4701, 4782, 1557, 3186, 1884, 1955, 1943, 3374, 3461, 2040, 3268, 956, 908, 3671, 2985]. C^{∗}(X) [1299, 1023]. C^{∞} [569, 888, 824]. C^{N} [3844, 3476, 3286, 2557, 2824]. c_{0} [2619, 885, 2732, 395, 3077]. c_{0}(Γ) [2076]. C_{i} [3500]. c_{o} [3956]. C_{p} [3206]. C_{p}(X) [2676, 3673]. C′_{∧} [4364]. B [1288]. B(X) [3413]. B_{1}(Ω) [2020]. D_{p−1}^{p} [4703]. J_{0} [2986]. ↕^{r} [1442]. L_{c}(E, , F) [1580]. T_{0} [534]. CAT(0) [3810]. cb [3889]. CP^{1} [3505]. CR [3390]. D [3009, 2384, 4006, 2404, 3741]. D(2, 1; α) [3265]. D(m, n) [3265]. D + M [1245]. d^{2}y^{1}/dt^{2} = f(y, y, t) [3516]. d^{2}y^{2}/dt^{2} = g(y, y, t) [3516]. D^{H} [4507]. D_{4} [4499]. ẍ + f(s)x = 0 [407]. Delta

[2585, 3688, 466]. Δ(x, N) = ∑_{n<xN(n,N)=1} 1 − xϕ(N) [3068].

Δu + Ku^{n+2n−2} = 0 [3226]. ẋ = Ax + g(x, , t) + εp(t) [887]. DQC [870]. D ⊂ R^{3} [3407]. E [2706, 4217]. E^{3} [454, 3211]. E^{n} [457]. e^{s(e^{x}−1)} [471]. E_{0} [4608, 3535]. E_{2} [3104]. E_{7} [4069]. E_{8} [4779, 3518, 3288]. e ≥ 4 [3946]. ℓ^{1}(Z_{+}) [4554]. ℓ^{2} [4436]. ℓ^{∞}(Γ) [3792]. ℓ_{1} [4225, 4114, 2964]. ℓ_{1}^{(n)} [3242]. ℓ_{∞} [4206]. ℓ_{p} [4534, 3453]. EP [1325]. ϵ [4310]. η [3911, 2987]. F [1074, 2350, 4034, 4511, 954, 270, 472, 4399]. F (4) [3265]. f(A) [641]. f(∑_{i=0}^{m} a_{i}X^{p^{ri}}) [1715]. f(X, X^{∗}) = g(X, X^{∗}) [902]. f(x, y) = z^{2} [472]. f(X, Y)^{q} = h(X)g(X, Y) [3696]. f(x_{1}) = f(x_{2}) = ⋯ = f(x_{r}) [856]. F/[R, , R] [650]. F [V_{2} ⊕ V_{2} ⊕ V_{3}]^{C_{p}} [3691]. F^{2} = 0 [1074]. F^{n} ∪ {Indecomposable} [3917]. f_{1}(x, y)z_{1} + ⋯ + f_{k}(x, y)z_{k} + g(x, y) [3496]. F_{2^{k}}D_{8} [3845]. F_{2^{n}} [3509]. F_{4} [326, 1050]. F_{p−ε}/p [1894]. F_{p^{s}} [1715]. F_{q} [907]. F_{q}((X)) [3612]. F_{q}((t)) [3315]. fC^{N} [1391, 2556]. fR^{n} [1680]. fRP^{n} [2167]. fSP [1616]. G [1105, 3991, 2420, 3445, 4375, 1348, 4299, 4809]. G(3) [3265]. G/P [3766, 3348]. G_{2}(k) [3135]. G_{2}(q) [2501]. G_{a} [2755]. G_{δ} [3841]. Γ [4806, 3737]. Γ(N) [3275]. Γ(σ_{1}, σ_{2}, σ_{3}) [2962]. gc_{N} [3279]. C [1053]. N [1184]. p [2855]. GL(d) [756]. GL(m) [3899]. GL(n) [3292]. GL(n, q) [2180]. GL_{2} [2248]. GL_{2}(R) [940]. GL_{4} [4485]. gl_{∞} [3375, 3875]. GL_{n}(B) [1305]. gl_{N}(C_{q}) [3285]. GU(2, 2) [4485]. H [2276, 1397, 2709, 2301, 2284, 3782, 2096, 1510, 3456]. H(φ) [2190]. H^{1} [3074, 3775, 3276]. H^{1}(R^{n}) [4394]. H^{∞} [4621, 3574]. H^{p} [1855, 1370]. H^{p}, 0 < p < 1 [1764]. H_{2π} [3472]. H_{G}^{2} [4582]. i [951]. I(B ∩ C) = IB ∩ IC [1971]. I_{n}(s)sI′_{n}(s) [3295]. II_{1} [2511, 1587]. itl^{∞} [1206]. j [3139]. j_{n} [4800]. J_{ν}(z) [3096]. JB^{∗} [2670]. JH^{∗} [2594]. J ≤ H [1396]. JV [2594]. K [2083, 3829, 4293, 2184, 2232, 660, 2325, 4126, 2167, 3363, 2543, 2467, 3056, 4720, 4802, 1324, 2774, 2163, 4165, 1942, 752, 1384, 3400, 3462]. K(S(T^{∗}RP^{2n+1})) [3152]. K(X, Y) [3755]. K3 [3733, 3367, 3902, 3381, 3579, 3552, 3550]. k[X, Y, Z] [3736]. K^{1} [4120]. K^{2} = 5 [3822]. K_{0} [4021, 1889, 2955, 3364]. K_{1} [4539]. K_{2}(O) [2652]. K_{n} [1725]. k_{ω} [4771]. KK [3831]. K⟨X, Y |XY = 1⟩ [4674]. L [3902, 3558, 3487, 4374, 2450, 4272, 4170, 4468, 3371, 2712, 2762, 4236, 2130, 2114, 4239, 3192, 1773, 2031, 877]. L(l_{1}, E) [2146]. L(p) [1531]. L(s, Sym^{m}(r), π) [2130]. l − l [1905]. L^{1} [1302, 301, 2161, 3743, 3406, 3514, 1901, 395]. L^{1}(μ, E) [2377]. L^{2} [3470, 2853, 1779, 3839]. L^{2}(R) [4525]. L^{2}(R^{n}) [2778]. L^{∞} [2080, 3013, 2884]. L^{∞}(μ) [4490]. L^{∞}[0, 1] [2948]. L^{λ} [978]. L^{p} [4717, 3076, 1267, 1436, 3267, 1939, 2639, 1817, 3656, 2745, 2147, 3957, 832, 1895, 3346, 2518, 2441, 2584, 2502]. L^{P} (μ) [2392]. L^{p}(μ), 1 < p < ∞ [1496]. L_{1} [4206, 1747, 1431]. L_{1}(Ω, μ) [2739]. L_{2}(R) [4191]. L_{∞} [2170, 3786, 1747]. L_{κα} [1799]. L_{n^{α}}(z) [137]. L_{p} [1497, 2419, 2517, 3607, 2577, 3129, 3986, 2686, 2814, 3024, 3195, 3938, 3616]. L_{p}, 0 < p < 1 [3643]. L_{p}, 1 ≤ p < ∞ [3030]. L_{p}, 1 ≤ p ≠ 2 < ∞ [2888]. L_{p},; 1 < p < ∞ [2254]. l_{q} [2814]. L_{w,1} [2938]. Λ [1183]. λ(n) [586]. Λ(p) [4649]. λ = 1 [1054]. λ > 1 [2206]. ≤ 1 [574]. Lip, α [2836]. M [2429, 1746, 2282, 2146, 1116, 2702, 3011, 4112, 509, 1067, 2882, 937, 1096, 4174, 3330]. m^{2} + n^{2} [3179]. M^{3} [1410]. M_{4}^{8}(C) [1194]. M_{11} [41]. M_{12} [571, 41]. M_{2n} [664]. M_{3,1} [3701]. MA_{countable} [2961]. C^{n} [4037, 4654, 4218, 3914]. CH^{2} [4646]. CP^{1} [4382]. CP^{2} [4646]. D [4018]. F_{q}[t] [4728]. Pn [4570]. R [3906]. R^{2} [4020]. R^{N} [4241]. Z/n[x] [4447]. Z_{2} [4786]. K(X, Y) [3993]. C^{√} [4498]. D [4607, 4646]. F⊕ [4153]. K(X , Y) [4040]. M_{X} (∈, ′) [4130]. O_{N} [4636]. Q_{√}

[4402]. ∥ [3614]. C^{∗} [4806, 4171]. mB [1750]. m mod n [1592]. mLC^{n} [2822]. modℓ [3271]. mP BIB [487]. n! [3301]. n! × n! [3684]. N [4347, 1353, 1345, 1995, 1900, 365, 2414, 3987, 1579, 1351, 2526, 1268, 105, 1702, 1399, 3930, 1086, 1923, 4826, 1643, 2940, 4465, 2188, 2290, 3342, 4375, 123, 163, 1753, 46, 2731, 3542, 455, 938, 4666, 969, 4076]. n(n + d) ⋯ (n + (k − 1)d) = by^{l} [3497, 3414]. n − 1 [4654]. n − th [656]. n = 2 [2032]. N = 2A3^{n} − 1 [992]. nd^{∗}(n) [2382]. n ≥ 8 [4570]. NK_{0} [3272]. O [4498]. O(p, q) [3561]. O^{∗} [864]. O_{n} [2915]. ω [1067, 865]. \overline{∂} [2540]. P [1885, 4561, 2960, 3393, 2023, 2174, 4802, 2060, 2527, 2180, 3182, 3026, 2529, 4643, 4597, 4479, 3534, 2705, 3641, 4496, 3287, 2194, 2694, 437, 2983, 4463, 2184, 2232, 859, 4780, 4170, 4625, 4049, 4468, 633, 4099, 866, 2884, 2649, 4765, 3599, 2112, 3236, 797, 3941, 4301, 2961, 3040, 1932, 2544, 4174, 3243, 85, 3078]. p − 1 [4198]. p.o. [751]. p < ∞ [1855, 2577]. p < t [3715]. p = 2 [193]. p^{2} [2876, 2127, 539]. p^{2}q^{2} [4159]. P^{n} [1354, 1898]. P_{4} [4086]. p_{g} = 0 [3822]. PCI [3060]. PGL_{3} × G_{2} [3007]. ϕ [1364]. π [4043, 3035, 2064, 1789, 658, 3774]. π(t) − li(t) [4820]. Pic(D[α]) [2384]. Pin_{4} [3119]. pq^{4} [4159]. pq^{n} [4110]. PSL(2, 2^{n}) [3509]. PSL(2, 2p) [2368]. P SU_{11}(q) [3429]. Q [869, 3091, 4775, 2363, 1952, 2047, 3843, 3089, 3036, 3116, 2421, 3451]. Q(−2p) [1868]. Q(−p) [1868]. q^{∗} [1708]. Q_{p} [3383]. Q_{p}^{#} [4263]. QFX [1304]. R [1906, 1820, 853, 940, 955, 1880, 2081]. R(S) [769]. r, λ [1409]. r = 0 [1114]. r^{∗} [1719]. R^{d} [2832]. R^{∞} [1772]. R^{n} [3462, 3430, 3387, 2694, 2641, 3226, 2246]. R_{0} [1233]. R_{1} [362, 432, 911]. R_{n,k} [2377]. RA [4216, 3210]. rank(A + B) = rank(A) + rank(B) [1001]. ρ [4804]. S [416, 4536, 1569, 4253, 3643, 3493]. S_{n}^{+} [4444]. S^{1} [2622, 3591]. S^{1} × S^{1} [1010]. S^{3} [2622]. S^{4} [3308]. S^{′∗} [3218]. S_{2} [3302]. S_{2}(Γ_{0}(dp^{2})) [4374]. S_{n} [3988, 2915]. S_{p} [2241]. σ [4796, 2352, 925, 2756, 2376, 2809]. {σ, τ} [4522]. σ− [1344]. Σf(n)μ(n)n^{−s} [1466]. sl(2) [1112]. SL(2, 2p) [2368]. SL(n) [2112]. SL^{+}(2, Z) [1186]. SL_{2} [4465, 1950]. SL_{2}(R) [3510]. SL × E^{n} [3851]. SO(3) [2851]. SO_{0}(p, q)/SO(p) × SO(q) [3145]. SouslinSouslinH ⊂ SouslinH [390]. Spec(R[[X]]) [2859]. SU(2) [3446, 3435]. SU^{∗}(2n)/Sp(n) [2808]. ∑(S) [769].

∑_{n=0}^{∞}(−1)^{n}(1 − c^{−n−1})^{k_{z}n} [1437]. ∑∞_{{n=1 \atop n≡l(k)}} (f(n)/n^{s}) [1658].

∑_{i=1}^{n} p_{i}f_{i}(p_{i})/f_{i}(q_{i}) ≤ 1 [1682]. ∑_{1}^{∞} f(n)/n [270]. ∑_{i=1}^{n}(f(p_{i})/f(q_{i}))p_{i} ≤ 1

[1139]. ∑_{k=−n}^{n} ε_{k}k [509]. ∑_{m<n/16} σ(m)σ(n − 16m) [3619]. ∑_{p≤x} p^{a} \underline{x} ; p^{k}

[2268]. ∑ g(n)/f(n) [1812]. ∑ X_{k}/d^{k} [1654]. T [4187, 561, 2673, 2859]. T (k) [3724]. T^{2} [1768]. T^{3} [1768]. T^{N} [3387]. T_{0} [1988]. T_{1} [1802, 2648]. t_{k} [2342]. t_{n} [3740]. tan nθ = n tan θ [2126]. τ [3704]. R^{N} [4301]. θ [2155]. θ^{r} [892]. × [3370]. Toppair [1814]. Topsbast [1814]. Triod × I × I [3006]. U [2389]. U(2, p^{2n}) [193]. U(G) [4033]. U_{n} × U_{m} [2968]. U_{q}(sl(2)) [4201]. U_{q}(sl_{n+1})^{+} [3492]. u_{tt} − ∂σ_{i}(u_{x_{i}})/∂x_{i} − D_{N}u_{t} = f [1300]. V [2825, 4349]. V (G) [4591]. V_{p} [1490]. φ [3569]. VMO [3276]. W [4396, 1310, 1323]. W (X, Y) [3755]. W^{m, \frac{n}{m}} (R^{n}) [4793]. WC(K, X)^{∗} [3102]. X [3673, 4130, 686, 1665, 97]. X(5, 3) [3577]. x(x + d)(x + 2d) + y(y + d)(y + 2d) = z(z + d)(z + 2d) [1107]. x(xy) = yx [793]. x′′′ + ax′′ + bx′ + g(x) = 0 [398]. X, , Y [578]. x^{(m)} = (−1)^{m}Q(t)x [1622]. X^{∗} [2960]. x^{2} + y^{6} = z^{e} [3946]. x^{2} − Dy^{2} = −c [3304]. X^{2} − DY^{2} = c [3304]. x^{3} + 117y^{3} = 5 [803]. X^{5} + aX + b [3314]. x^{6} + ax + b [4785]. X^{k} + Y^{k} = Z^{k} [1984]. x′Ax · y′Ay − (x’Ay)^{2} [17]. Ξ [3672, 3703]. ξt_{n}1 [1375]. xy + yz + xz = n [2928]. xyz = x + y + z [2545]. y′′ + λpy = 0 [1327]. y^{2} = x^{n} + k [3921]. Z [2966, 1751, 3197]. Z^{2} [4430, 3596]. Z_{2} [2576, 3670]. Z_{4} [3733]. Z_{n}[x] [1425]. Z_{p} [1843, 3929, 2123, 2399]. ζ [81]. ZS_{3} [984]. |f(x + iy)| = |f(x) + f(iy)| [355]. |f(x + iy)| = |f(x) − f(iy)| [355]. |Γ(s)| [3878]. |q| = 1 [3036].

-Actions [3591, 2885]. -adapted [3445]. -Adic [3599, 2527, 2194, 4669, 4463, 859, 4780, 4170, 4468, 3078, 2884, 2112]. -Algebra [3357, 3461, 4701, 2593, 2670, 4064, 4089, 2439, 4077]. -algebras [2376, 3863, 3268, 3186, 3832, 3197, 3671, 3836, 3436, 3601, 3364, 2594, 1557, 4350, 956, 4806, 3771, 4496, 2955, 908, 1884, 4187, 4782, 2985, 2516, 4717, 1943, 2679, 3374, 2628, 2359, 4774, 1226, 4487, 2650, 4220, 968, 4302, 2869]. -algèbres [2594, 2679]. -analogs [1952]. -approximation [2853, 2161]. -associative [2706]. -balls [1900]. -barrelled [1569]. -Bases [3774]. -Bessel [2421]. -beta [2363]. -Betti [4436]. -Bloch [3639, 3746]. -body [4475]. -Borel [4775]. -bounded [1096]. -boundedness [3076]. -Bundles [3119, 3991, 2851]. -cell [1339]. -choosability [4303]. -Chromatic [4034]. -circulants [197]. -Class [3287]. -Clean [3699]. -closed [2948, 1397, 1233]. -cobordisms [1510]. -cocycles [2576]. -cohomology [1432, 3346]. -color [2414]. -compact [1636]. -compactness [2155, 2961]. -completeness [2685]. -Complex [3119]. -connectedness [2404]. -contact [4126]. -Continuity [3737]. -contractions [2419]. -Convergence [3743]. -convex [1702]. -Convexity [3174]. -Core [4034]. -critical [3784]. -crossed [2929]. -Crossed-Products [3747]. -cube [365, 1643, 1184]. -cutset [2526]. -cycled [1942]. -cycles [1253, 4396]. -cyclic [437]. -Dedekind [4809]. -Derivations [4085, 2902]. -Dialgebras [3569]. -Differential [3493, 2930]. -Dimensional [3987, 3342, 3804, 4086, 2389, 4535, 4390, 764, 4659, 2433, 2882, 4098, 1223, 2901, 4452]. -discrete [2809]. -distance [4346]. -distributive [1885, 1611]. -divergence [4399]. -divisible [869]. -divisors [4143]. -Einstein [4804]. -embedded [4174]. -equation [2540]. -exponential [3036]. -Extensions [3509, 3583, 1988, 2123, 4786]. -factor [2511]. -factorizations [2258]. -Families [3724]. -fermion [105]. -Fields [3500]. -finite [4796]. -fold [4465]. -Forms [3693, 3243]. -fractions [2131]. -frames [2673, 2966]. -free [1880, 2081, 4347, 4086]. -full [1789]. -Function [3703, 3704, 3011, 466, 4236, 1348, 81]. -Functions [3487, 4374, 4272, 4170, 4468, 2712, 2762, 2130, 4239, 3192, 1773]. -generated [2731]. -gons [1994, 163]. -graded [3929]. -graph [172]. -graphs [955]. -group [633, 3857, 1932]. -groupe [3857]. -Groups [3393, 3272, 3941, 4597, 2008, 982, 2522, 797, 2544, 3364, 2352, 4120, 2945, 1310, 864, 3056]. -Harmonic [3243, 2983, 2940, 3330]. -Hessian [4802]. -Hölder [4414]. -homogeneity [4076]. -homology [4554]. -homotopy [3670]. -Hypercentre [4153]. -Ideals [4511, 1751, 2429, 1746, 2146, 2702]. -idéaux [1746]. -injectivity [769]. -integers [1512]. -integral [937, 1354, 1898]. -Invariant [3139, 1719, 4043, 2622]. -invariants [2987]. -Jacobi [2567]. -Kronecker [4310]. -Laplace [4625]. -Laplacian [4479, 3534, 3641, 4301]. -large [2060]. -lattice [2649]. -like [1753]. -limits [2083]. -lines [2560]. -linkedness [2756]. -Local [3892, 2184, 2232]. -Log-Convex [3843]. -manifold [2771, 4412, 3818]. -Manifolds [3409, 3864, 3591, 4424, 2819, 4139, 2987, 3006, 2959, 1389, 1740, 3851, 1843]. -mappings [1067, 865]. -matrices [873, 1307, 1328]. -matrix [508]. -matroids [678]. -Mean-Square [3089, 3116]. -metrics [3698]. -minimal [561, 4498]. -mixing [1364]. -module [2276]. -Modules [3014, 4607]. -monoids [2031, 4021]. -Monotone [3643]. -motives [4491]. -moves [2585]. -multiplication [4349, 2004]. -Multiplier [3889]. -Multipliers [3889]. -Nets [3131]. -nilpotence [3026]. -nilpotency [4174]. -noetherian [730]. -Norm [3102, 4496, 4049, 4152]. -normed [4426]. -norms [2441, 2584, 2665, 3514, 3406]. -operator [4099]. -orbits [3829]. -Orderings [3566]. -orders [853]. -orthogonal [1554]. -paracompact [1067]. -Parallel [3911, 3169, 4018]. -parameter [586, 1353]. -parametrization [4498]. -partite [123]. -permutable [4112]. -plane [2163]. -Point [3930]. -points [2709, 4561]. -Polynomials [3758]. -polytopes [2404]. -powers [4114]. -prefrattini [954]. -Primary [3389, 2403]. -product [822]. -projective [1820]. -quasi [4375]. -quasinormally [4253]. -quotient [1906]. -radical [2986]. -Rank [3419, 2652, 4643]. -reflexive [925]. -regular [1379]. -regularity [2437]. -Riemann [4816]. -ring [3060]. -Rings [3687, 1708, 686]. -rises [1323]. -rotund [1935]. -Sections [3400]. -semidirect [2284]. -semigroup [4608]. -Semigroups [3535]. -semilattices [1335]. -seminorm [4114]. -semipermutable [4253]. -sequences [2431]. -sequential [1302]. -Series [3451, 2174, 3902, 3558, 3371, 2114]. -Sets [3091, 1105, 416, 2389, 2301, 3782]. -Sextic [4140]. -simple [2276]. -simplex [455]. -societies [752]. -space [4165, 2732, 2518, 2096, 822, 2960, 1900]. -spaces [866, 3607, 2005, 3009, 2392, 2420, 978, 1802, 4771, 2350, 2648, 1895]. -special [969]. -sphere [2469, 2569, 2997]. -Squashed [3741]. -Stability [3197]. -Stable [3481]. -structures [658]. -subgroups [1700]. -Subsets [3841]. -subspaces [2732]. -summands [2282]. -summing [1213]. -Surface [3462]. -Surfaces [3552]. -Sylow [2180, 2529]. -symmetric [4720, 2321, 1116]. -symmetries [4430]. -systems [1529, 1409, 3596]. -term [877, 3104]. -ternary [1955]. -th [3930]. -theoretical [4293]. -Theory [3363, 4217, 3831, 2184, 2232, 2325, 2167, 2467]. -thin [2694]. -topological [362]. -topologies [2825]. -torsion [2060]. -Torsors [3518]. -torus [2843, 2494]. -towers [2399]. -Traces [4517]. -transitive [728]. -tuples [1324]. -Type [3614, 2658, 2453]. -ultraspherical [2047]. -uniform [1885, 2543]. -universality [3670]. -valued [1442, 2502, 3435, 1817]. -Vector [3456].

1-smooth [4765]. 14 [1098]. 15 [1161]. 16 [1102]. 17 [1413]. 1960/61 [1317].

1984 [3850].

2-class [4786]. 2-Dimensional [3629]. 2-Forms [3488]. 21 [1619].

3 [1317]. 3-cells [4677]. 3-Forms [3488]. 3-Manifold [3441]. 3-manifolds [4709].

4 [1317]. 4-cycles [117]. 4-manifolds [4680].

90 [295].

A. [684, 2567, 154]. abacus [114, 145]. Abbena [2946]. Abbott [697]. Abel [1901, 1486, 3173, 4430, 204, 169]. Abel-type [1486]. abelian [2448, 2316, 1210, 4796, 2008, 1779, 2531, 2166, 4649, 4669, 627, 2999, 2160, 2275, 1677, 2985, 805, 1703, 2107, 1114, 797, 1227, 626, 4243, 4037, 2994, 1798, 285, 3788, 3588, 3504, 4137, 3289, 3612, 4314, 4629, 4095, 1406, 1453, 3982, 4652, 4007, 4656, 3922]. Abelianization [4027]. abéliens [1798, 2444, 3612]. Abian [1760]. abnormality [1251]. Abramson [340]. absence [106]. Absolute [2000, 2411, 1277, 2562, 1095, 1481, 559, 1982, 1100, 2580, 2375, 1670, 2018, 368, 2044]. Absolutely [1424, 1287, 3027, 3794, 1429, 1861, 1928]. absorption [1062]. Abstract [3150, 4063, 4353, 3972, 3729, 41, 1573, 861, 1042, 1333, 401]. abundant [1234]. abundant-like [1234]. ACC [2596]. accelerating [215]. accessible [2402, 1121]. accessibles [1121]. according [2011]. accretive [1953]. Accumulation [1462]. accuracy [489]. Acknowledgement [2850]. Act [4348]. Acting [4621, 3628, 4169, 346]. Action [3733, 4463, 2290]. Actions [4155, 4201, 3520, 3906, 3591, 3739, 3384, 3347, 3552, 4529, 2008, 2187, 2275, 4754, 2885]. acts [2595]. Acute [3801]. acyclic [1550]. Ad [4445]. Ad-nilpotent [4445]. Adams [3104, 4793]. Adams-type [4793]. adapted [3445]. Addenda [1917]. Addendum [4107, 1413, 550, 231, 2850, 491, 268, 2081]. addition [1989]. Additive [3505, 2001, 4428, 3840, 3315, 2430, 2385, 1800, 1066, 1922, 1628, 2209, 2449]. additives [1628]. Adequate [3872]. adherences [2301]. Adic [3599, 2527, 2194, 4669, 4463, 859, 4780, 4170, 4468, 3078, 2884, 2112]. Adjacencies [3460]. Adjacency [3463, 1200]. Adjoint [1517, 3873, 3530, 2871, 1324, 3043, 3061, 2028, 688, 1540, 369, 243, 1864, 2958, 2974, 4638, 2490]. Adjoints [1192, 3869]. Adjonction [1707]. Adjunction [3250]. Admissibility [4154]. Admissible [3519, 3890, 892, 1283]. admits [2519]. admitting [2177, 1978]. AF [3580]. AF-Skeletons [3580]. Affability [3596]. Affine [4037, 4459, 4201, 3533, 3770, 4065, 3285, 4463, 3648, 1556, 344, 2753, 2796, 2519, 4552, 4710, 2780, 433, 112, 152, 225, 538, 126, 1570, 2786, 2609]. affinely [1977]. after [3345]. again [99]. AH [3443]. Ahlfors [3834]. aléatoires [97]. aleatory [97]. Alexander [3120, 4370]. Alexandroff [4119]. Alexandrov [3971, 4545]. Algebra [3557, 4420, 4476, 3369, 3845, 3889, 3285, 3705, 4409, 380, 4088, 3357, 3510, 3461, 4639, 3363, 3350, 3602, 3317, 4674, 4701, 2977, 2593, 2670, 257, 2088, 841, 2072, 2756, 1246, 4064, 1607, 1194, 4799, 1515, 2232, 2984, 229, 4089, 1361, 766, 1332, 1276, 2439, 387, 542, 853, 1960, 1613, 1938, 4077, 1370, 2962]. Algebraic [3230, 2174, 4058, 3367, 3141, 4614, 3928, 3576, 4388, 2394, 4618, 4538, 2625, 4610, 4431, 3543, 3333, 4583, 4652, 4032, 2235, 2225, 132, 2517, 2213, 1203, 1975, 3016, 4768, 2883, 610, 1830, 736, 2733, 3219, 2202, 466, 3053, 3251]. Algebraically [3925, 700]. Algebraicity [3432]. Algebras [4401, 3148, 4175, 3257, 4611, 4106, 3263, 3264, 3321, 3638, 3368, 3266, 3863, 4437, 3142, 3268, 3903, 4160, 4521, 3790, 3232, 3748, 4110, 4083, 3865, 3520, 3906, 3929, 4111, 3410, 3186, 3832, 4085, 4115, 4270, 3197, 3223, 3562, 4426, 3608, 4167, 3671, 4537, 3125, 4294, 3836, 4559, 3580, 4378, 4501, 3937, 4659, 4214, 3858, 4592, 3895, 3494, 3421, 4028, 3436, 3377, 4216, 3601, 3647, 3238, 2181, 4652, 4219, 3648, 3364, 3678, 3255, 2320, 4244, 4764, 2249, 2065, 2594, 567, 922, 2978, 2753, 2979, 2740, 1461, 121, 3829, 2272, 1557, 4350, 2907, 2276, 2520, 2892, 582, 2080, 2773, 2893]. algebras [2954, 956, 1604, 4806, 3771, 2555, 3028, 311, 2780, 4496, 2637, 3288, 2955, 2692, 80, 124, 1125, 2558, 2967, 908, 960, 1673, 2940, 2281, 1884, 2360, 4187, 2823, 4782, 3159, 2600, 2376, 2338, 2985, 1925, 2516, 4717, 2914, 1943, 804, 1196, 2033, 1542, 2507, 2679, 3374, 1999, 2883, 2454, 556, 3201, 2040, 2918, 1919, 3085, 2628, 2359, 2564, 2847, 4774, 729, 1695, 2395, 2666, 1226, 4740, 1588, 2992, 4487, 1424, 2702, 2412, 1554, 1524, 1177, 854, 966, 273, 2339, 2650, 4220, 968, 1778, 1651, 4302, 2869, 1929]. algèbre [2977, 1361]. Algèbres [3288, 2320, 4244, 2594, 2679]. algébrique [3251, 2733, 3219]. Algebroids [4467]. algorithm [2351, 1867, 1464, 2910, 75, 820]. Algorithms [4497, 3945, 2064]. Allen [4516]. Allison [4779]. Almost [4457, 1902, 3719, 3685, 4226, 4678, 3751, 4604, 3213, 526, 125, 900, 4195, 1313, 3981, 4562, 4003, 3783, 4546, 4129, 3424, 4566, 970, 2154, 3765, 857, 1242, 488, 906, 1927, 4681, 1303, 1582, 4695, 294, 1168, 372, 664, 423, 1563, 2108, 1279, 4798, 652, 2067, 731, 879, 238]. almost-periodic [1279, 731, 879]. almost-quaternion [2108]. Along [4611, 4435, 3999, 3894]. Alster [4156]. alter [468]. alternant [199]. Alternating [1398]. alternative [1728]. AM [3883, 3620]. AM-Compact [3883, 3620]. Amalgamated [3001]. amalgamation [1694, 2411, 2226, 3078]. Amalgamations [3712, 990]. amalgams [3772, 2270]. Amenability [4401, 4175, 4083, 3377, 4626, 4128, 4221, 1411, 3255, 4131, 2773, 1777]. Amenable [4590, 3626, 3397, 4187, 4754, 3228, 3824, 2508, 1729]. Amitsur [1066]. Among [3403]. Ampère [4403]. Ample [3273, 3250, 3114, 1396]. amplitude [2693]. Amsterdam [3673]. analogous [1478]. analogs [1952]. Analogue [3581, 3234, 3365, 2751, 675, 2504, 59, 2723, 4763, 646, 483, 122, 558, 403]. analogues [4803, 1928]. Analogy [1358]. Analyses [4582]. Analysis [4556, 4427, 4297, 4008, 739, 3027, 1153, 215, 1261, 542, 1337, 29, 55, 91, 613].

Analytic [4458, 3261, 3105, 4776, 4461, 4441, 3113, 3973, 1039, 1102, 3100, 3089, 3116, 1991, 4343, 3847, 1771, 1099, 1440, 883, 4442, 1817, 2760, 1542, 546, 2131, 569, 24, 2883, 2353, 2132, 1521, 81, 464, 1666, 2919, 720, 1280, 1949, 743, 2374, 109, 2312, 2787, 1873, 2749]. analytique [1771]. analytiques [2459]. ancillary [722]. and/or [1963]. Anderson [2149]. André [4718]. André-Oort [4718]. Angle [3486]. Angular [3392, 2810, 2099, 329]. Anick [3177, 3820]. anisotropes [2876]. Anisotropic [4451, 3835, 2876]. Ann [1317]. Anneau [4087, 1889, 1950, 452, 754]. Anneaux [2292, 2042, 1143, 128, 120]. Annihilating [2426, 3871, 2626, 4322, 4344, 2730]. Annihilating-ideal [4322]. Annihilation [4239, 1002]. annihilator [255, 290, 860, 968]. Annihilators [4286, 2596, 1108, 677, 2463]. annulation [3166]. annulus [2185]. Answer [3336, 3034, 4768]. antecedent [2565]. Anti [4265, 4638, 513]. anti-automorphism [513]. Anti-distributive [4265]. Anti-dual [4638]. Antiautomorphism [3363]. Antichains [3750, 4151]. Anticommutative [4537]. Anticommutativity [4022]. antiderivatives [4698]. Antidirected [1858]. antipodal [373]. Antipode [3928]. antisymmetric [1022]. Any [3449]. Anzai [2843]. Aperiodic [4506]. Appell [326, 1050]. Application [3393, 576, 4229, 4026, 3943, 1735, 2735, 278, 1437, 713, 1463, 1625, 2662, 162, 267, 2342, 466, 1337, 1264, 587, 4783, 4690]. Applications [3261, 4572, 4160, 4264, 4227, 3889, 3627, 4120, 3129, 4295, 3225, 1510, 3434, 3433, 3957, 3524, 3980, 3731, 1591, 2547, 515, 2549, 2364, 1582, 1853, 898, 2118, 573, 933, 1363, 1670, 1854, 572, 1726, 542, 3803, 1899]. applied [1987]. appliquée [1987]. Applying [460]. Appolonius [474]. Approach [4403, 3439, 3431, 3997, 3877, 4652, 3806, 4739, 2574, 4793, 2618, 2575, 1880, 2081, 2998, 1860, 113, 191, 370, 1797]. Approximants [4477]. Approximate [4175, 3906, 2692, 4209, 4199, 2954]. approximated [2576]. approximately [857, 1717, 1604, 831, 910, 2628]. Approximating [4532, 3360]. Approximation [4455, 2419, 2100, 4058, 3745, 3686, 840, 4689, 4021, 4332, 2366, 2556, 3247, 3286, 3522, 4269, 4371, 3973, 3016, 1359, 3974, 3568, 988, 2079, 3219, 3861, 4489, 1899, 1477, 1668, 3555, 3882, 3315, 1951, 2669, 1665, 2853, 2836, 1926, 3041, 1428, 578, 2517, 1122, 1398, 1848, 1746, 2557, 1218, 2577, 1260, 1916, 1832, 119, 2824, 2825, 2173, 4754, 1565, 2727, 2161, 587]. Approximations [3380, 1968, 40, 565, 1518, 2757, 1916, 3, 2204, 401, 1027]. approximative [2018]. Arakelian [3568]. Arbitrarily [4640, 4791]. Arbitrary [3614, 4563, 2857, 3823]. Arc [816, 449]. arc-preserving [449]. Archimedean [3687, 2673, 4465, 4698, 2432, 3227, 3980, 3388, 3571, 3677]. Arcs [3286, 3212, 258, 198, 598, 1198]. Area [4461, 3118, 2470]. areas [718]. Arens [4354, 4639]. Arguesian [1963]. Argument [3156, 4523, 4613, 3877, 4710, 1215, 544]. arguments [2532, 1923]. Arhangel’ski [4585]. arise [3182, 4745]. Arising [4497, 3657, 4453, 2668, 2669, 4678, 1357, 1474, 4746]. Arithmetic [3092, 3624, 3454, 3111, 4142, 4317, 4188, 1761, 4794, 4194, 4563, 314, 3921, 95, 520, 4679, 1438, 334, 13, 348, 1499, 1856, 1171, 877, 1618, 492, 1770, 367, 1776, 2464, 158]. Arithmetical [770, 2430, 1061, 2551, 2434, 1644, 113, 191, 370, 1647]. Arnold [3760]. Arrangements [3744, 4157, 4288, 4154, 3830, 2718]. Arrays [3131]. Arrow [3697]. Arthur [3225]. article [1466]. Artin [2855, 4332, 2450, 705, 2236, 2630]. Artinian [4038, 1889, 1301, 3886, 654, 3613, 3132, 1695]. Artinianness [4287, 3915]. artiniennes [654]. artiniens [1889]. AS-regular [4659]. ascending [860]. Askey [2363, 2935]. Aspect [3310]. aspects [2034, 169]. asphericity [4672]. Associahedra [3270]. Associated [3440, 4186, 4518, 4229, 3671, 3128, 3566, 3872, 3300, 3540, 4483, 4339, 4482, 3461, 3236, 3881, 3391, 4223, 3466, 3550, 4244, 1698, 2925, 1080, 284, 410, 2551, 1465, 1625, 4087, 2760, 2028, 3200, 2957, 271, 263, 3817, 2112, 1862, 2442, 3056, 3204]. Associative [3148, 3638, 4160, 4522, 4652, 841, 2706, 685, 476, 1999, 729, 714, 1524]. associativity [913]. associé [4087]. associées [4244]. Assouad [4606, 4108]. assumptions [358]. Asymptotic [3900, 4158, 3810, 2048, 3606, 2373, 13, 1166, 4770, 2860, 3029, 4753, 1992, 3994, 4142, 3955, 2029, 2173, 3133, 3860, 4805, 2990, 150, 3616, 4381, 1603, 1606, 2722, 2376, 1472, 1599, 1888]. Asymptotic- [3616]. Asymptotically [4380, 723, 1827, 2295]. Asymptotics [4330, 3799, 2603, 4435, 4057]. asymptotique [2268]. asynchronous [1164]. Asynchronously [3351]. ation [468]. Atomic [3932]. Atomless [2781]. Atoms [2039]. Attached [3271, 4442]. attachement [3251]. attaching [1595]. attachment [3251]. attaining [3062, 3854]. attraction [863]. attractivity [2703, 2939]. attractor [3775]. attractors [2797]. Aubin [3144]. Auerbach [3896]. Auslander [4190, 1695]. auteurs [3022, 3088, 3147, 3205, 3256, 3258, 3378, 3438, 3498, 3556, 3618, 3680, 3682]. Author [3022, 3088, 3147, 3205, 3256, 3258, 3378, 3438, 3498, 3556, 3618, 3680, 3682, 2733]. Autocorrelation [4716, 3386]. Automatic [2466, 3351]. automizers [2994]. automorphes [3582]. Automorphic [4497, 3582, 2117, 720, 1280, 970, 3260]. Automorphism [3747, 4295, 4125, 3645, 513, 1918, 2533, 2250, 3612, 2021, 2181, 303, 3070, 2089]. automorphismes [248]. Automorphisms [4061, 3329, 4048, 389, 3361, 3037, 1587, 3941, 2629, 3238, 4198, 3081, 700, 1535, 2013, 2333, 1703, 1392, 2678, 1910, 3058]. Automorphy [4457]. Autonomous [3972, 1242, 2202]. Autoparallel [79]. Autour [2733]. Average [3487, 4210, 4074, 4691, 2771, 4820, 1770]. averaged [1274]. Averages [4512, 4271, 720, 4549, 4400, 4747, 2712, 2762]. Averaging [3127, 1423, 2620]. avoid [520]. avoidance [1408, 3823]. Avoiding [3689, 4092, 4462]. awfully [4685]. axe [2972]. Axially [3144]. axiom [1028, 1212]. Axiomatic [230, 2337]. Axiomatizability [3607]. Axioms [279, 1799, 674, 609, 2561, 343]. axis [2972]. Axler [2313]. Azarin [1326]. Azumaya [4764, 2272, 2507, 772].

B [734, 4493]. b.i.b [1543]. Bachmuth [3349]. Back [4509, 4569, 4628, 4687, 4706]. Bäcklund [3198]. backs [4789]. Backward [3150, 2927]. Bad [3529]. Baer [3862, 3073, 3739]. Báez [4469]. Baire [1356, 1563]. baked [31]. Bakelman [4545]. Baker [3451]. Baker-Type [3451]. balanced [2054, 2741, 224]. Ball [3574, 3665, 3476, 3224, 4391, 3100, 3914, 3372, 3102, 4549, 802, 2641, 2642, 2696, 3079, 2449, 2032, 3952]. ballot [267]. Balls [4267, 4084, 4271, 3567, 4400, 2266, 3356, 2188, 4778, 2976, 1973, 2291, 1900]. Banach [3883, 4701, 1774, 3620, 2546, 3261, 3483, 2977, 4611, 1449, 1480, 628, 1498, 3105, 2964, 4478, 1417, 1289, 3653, 3686, 3748, 503, 1765, 3750, 1678, 2619, 2800, 1607, 4084, 442, 3045, 3773, 2637, 1123, 2692, 3014, 2407, 2360, 2863, 2417, 2432, 2844, 1249, 2346, 1431, 3020, 546, 556, 4192, 4359, 1174, 3980, 1935, 2973, 963, 4798, 3943, 3677, 1575, 2142, 2830, 2449, 3616, 683, 777, 1370, 3255]. Band [3075]. Band-limited [3075]. Bands [3401, 2581]. Banff [3850]. bang [298]. bang-bang [298]. Barnes [2723, 889, 1591]. barrelled [1344, 1569]. barrier [843]. Barthe [4146]. Barycentres [1735]. base [1258, 1530, 3805]. Based [3104, 943, 2823, 3021, 2920]. Bases [3140, 4493, 4494, 3774, 3490, 3594, 3896, 3388, 2320, 1289, 1258, 761, 68, 3047, 2223, 2910, 4736, 678, 959, 2051, 1548]. Basic [3223, 4274, 4320, 2602, 2320, 2002, 501, 1507, 4676]. Basin [3420]. Basis [3691, 3714, 3349, 1593, 288, 1678, 1919]. Bass [2710]. Batyrev [3416]. Baxter [4522, 4746]. BCM [4508, 4509, 4568, 4569, 4589, 4627, 4628, 4648, 4686, 4687, 4705, 4706, 4766]. Be [4006, 2002, 3024, 1084, 1070, 1207, 2656, 2463, 2986, 3784]. Beatty [38, 50, 868, 1446, 3899, 4632, 4688]. Beckman [2601]. before [1062]. behaving [1677]. Behavior [3010, 3738, 3295, 3133, 3860, 2521, 1187, 2099, 2745]. Behaviors [4518]. behaviour [1968, 2573, 3011]. Beilinson [4653]. being [4685]. Bell [1179]. belonging [1332]. Below [3639, 3746, 2788, 2850]. Beltrami [4247, 2387]. Bending [3211]. Benson [3118]. Berger [3422]. Bergman [2110, 4663, 2452, 3536, 4738, 3224, 3100, 3042, 3914, 3019, 4189, 2468, 3742]. Berndt [1644]. Bernoulli [2299, 4062, 2024, 2038, 1472, 3342]. Bernstein [2923, 4266, 635, 616, 3773, 1018, 1667, 3524, 988, 2035]. Bernstein-type [2923, 988]. Bers [4295]. Bertini [4510]. Besicovitch [1908]. Besov [2139, 4250, 4339, 4587, 4400]. Besov-Type [4400]. Bessel [4781, 3096, 3113, 2803, 2931, 3660, 2421, 564]. Best [1273, 3634, 1665, 1926, 1428, 1122, 1518, 1218, 1916]. beta [814, 933, 2363]. Better [3598]. Betti [4436, 1779, 4826, 4212, 4242]. Between [3483, 3746, 3685, 3748, 4164, 4310, 4192, 3880, 3530, 4343, 2546, 415, 3652, 1790, 2097, 2019, 1533, 579, 4714, 2874, 2520, 2179, 2908, 3794, 1326, 2597, 1356, 2622, 737, 813, 249, 2184, 2119, 458, 3754, 2403, 2600, 4717, 2914, 1736, 243, 2488, 1047, 1315, 2650, 2490, 2361, 2869]. Beurling [4469, 2059, 3834, 3344, 4073, 3585]. Beyond [4429, 4688]. Bezdek [3726]. Bezout [2091, 1079, 2292]. Bhattacharya [2956]. Bi [4306, 4745, 4522, 4412]. Bi-orderability [4306, 4412]. Bi-orders [4745]. bialgebras [4522]. Bianchi [2589]. biased [1848]. Bibasic [4183]. BIBD [1054]. bicohomology [2394]. Bicyclic [2710, 2842, 4362]. Bieberbach [4121]. Biflatness [3421]. Bifurcation [4382, 2326, 1469, 2725, 2711, 2645, 2747]. bifurcations [2947]. Big [3659, 3700]. Biharmonic [3662]. Biholomorphic [4548]. Bihomogeneous [4755]. Bijective [2414]. bilateral [2217]. Bilinear [3842, 3748, 4274, 2246, 2113, 1897, 453, 2106, 23, 110, 111]. bimeasures [2298]. Bimodule [3447]. Bimodules [3241, 4782]. Binary [2526, 3208, 4047, 4540, 4712, 472, 1286]. binomial [716, 2912, 204, 608, 1893, 2140, 1239]. binomiaux [2182, 2416]. biomorphisms [1435]. Biorthogonal [3444, 3239, 1952, 1033]. Bipartite [4101, 1961, 561, 4792, 33, 236, 1625, 117]. biplanar [172]. bipolynomial [2065]. Bipositive [2217]. Biprojectivity [3421]. Biquadratic [3269, 775, 1479]. Birch [2403]. Bireflectionality [2375]. biregular [1552, 1526]. Birkhoff [2351, 72, 1867, 75, 4411, 1630]. Birkoff [646]. bisector [1501, 1584]. Bisectors [3486]. bisymmetry [913]. Bisztriczky [3717]. Bitangent [4637]. bitangents [212]. bivariate [1183, 1018, 1667, 1055]. Bivectors [1838]. Black [3417]. Blackburn [591]. Blaschke [3536, 2423, 2370, 4630, 4472]. BLO [4622]. Bloch [3639, 3746, 4164, 3383, 4653, 3947]. Bloch-to- [3383]. Bloch-type [3947]. Block [2054, 4581, 224, 3922, 2206, 4726, 147, 851]. block-graphs [147]. blocks [3159]. Blow [2749]. Blowflies [4604]. Blowing [2437]. Blowing-up [2437]. Blowings [3273]. Blowings-Up [3273]. BMO [2143]. BMOA [2617]. Bochner [3339, 4547]. Bockstein [3122]. Bodies [3719, 3720, 4113, 4182, 3726, 3400, 3721, 4790, 1621, 2957, 36, 14]. body [4475, 2957, 2746]. Bogdanov [2457]. Bogomolov [2751]. Bohr [4803, 4251, 1336]. Bombieri [2923]. Bonnet [1978]. Book [4634]. Boolean [44, 121, 599, 4575, 196, 1305, 2281, 1919, 940, 1929]. Boolean-valued [2281]. Borel [1413, 1135, 1481, 1790, 2613, 4775, 2293, 144, 374, 3695, 640, 735, 3481, 3840]. Borel-type [1413, 1135, 1481, 1790, 2613]. Borelian [1356]. boréliennes [1356]. borne [3857, 827]. bornés [1891]. Borsuk [391, 2026]. Borwein [3825]. Bosonic [3285]. both [4733]. Bottle [4139, 4685]. Bottles [3345]. boules [2266, 2976]. boules-unités [2976]. Bound [3107, 4421, 2286, 3994, 4092, 3131, 3542, 4171, 4103, 4102, 1787, 1626, 1219, 4763, 2700, 2763, 783, 2808]. boundaries [2541]. Boundary [3439, 2099, 4177, 4574, 3930, 3738, 2661, 3997, 3657, 1172, 4319, 1488, 4527, 4505, 4543, 4218, 4454, 1716, 2094, 3183, 1517, 2228, 2307, 1182, 4770, 311, 4753, 402, 2660, 3795, 934, 2335, 3815, 3017, 1173, 773, 938, 1863, 1315, 2787, 1668]. Boundary-value [1488]. Bounded [1912, 3948, 4437, 3413, 4649, 1716, 3024, 706, 1320, 442, 2055, 931, 1627, 2200, 815, 767, 835, 3756, 363, 2353, 1761, 1908, 1523, 2650, 1096, 1891]. Boundedness [3639, 3746, 3932, 1188, 3089, 3957, 750, 3116, 3076, 4684, 2640, 458, 1732].

Bounds [903, 1664, 656, 1643, 4635, 4462, 4142, 4069, 4526, 3389, 943, 227, 2252, 1214, 1327, 858, 2475, 727, 2234]. Bouquet [1873]. Bourgain [2940]. Bourget [3679]. Bourguignon [4695]. Box [3417]. Boxall [4614]. Boyd [4066]. bracket [171]. Braid [3327]. Braided [4136]. braids [1119]. branch [2098]. Branched [3356, 2098]. Branching [1676, 4270, 4465, 2621, 2777, 4252, 1221, 3612]. Branges [4759]. Brans [1294]. Brascamp [4146]. Brascamp-Lieb [4146]. Brassil [4503]. Brauer [2907, 4248, 3381, 4538, 4539, 2297, 3070, 2590]. breadth [1502]. Break [3269]. Brelot [249]. Brianchon [3515]. Brief [4275]. Brill [4712]. Briot [1873]. Brouwer [2072, 350]. Browder [3901]. Brown [563, 3431, 3051]. Brownian [1652, 2965, 2970, 3404, 2583, 3635]. Brualdi [2351]. Brunel [2538]. Bruns [2471]. Büchi [4563]. Budapest [1317]. Buildings [3531]. Buium [3178]. Bulk [4313]. Bull [1413, 1098, 1161, 1160, 1619, 1102]. Bump [3848, 4236]. Bundle [4570, 3702, 3213, 3766, 2798, 2167]. Bundles [4605, 3991, 3502, 3505, 3095, 3411, 3910, 3250, 3114, 4341, 4342, 4544, 4434, 3119, 2851, 3167, 1211, 2208, 3172, 2332, 372]. Burde [1479]. Burgess [4542]. Burnside [3160, 4358, 3920]. Burr [814].

C [759, 1914, 4493]. C. [1351, 1957]. Cable [4662]. Cahn [4516]. Calabi [3454, 3605, 4750, 3529, 4507]. calcul [2484, 1089]. calculation [1089]. Calculus [3705, 4372, 4398, 3761, 4547, 2851, 979, 1161, 2484, 2985]. Calderon [1026, 2143]. calibers [2676]. Calkin [2484]. Camillo [1886]. Camina [4094, 4299]. Can [2463, 2002, 2656, 2986, 1567, 1576]. Canad [1413, 1098, 1161, 1160, 1619, 1102]. canary [2714]. Cancellation [3899]. cannot [1070, 1207]. Canonical [3367, 4293, 4235, 4612, 4320, 3403, 3874, 4507, 842, 1307, 2084, 919]. canonically [2337]. Cantelli [735]. Cantor [3207, 3773, 3187, 4023, 3596, 1445, 4562, 183]. Cantor-Vitali [183]. capacities [785]. Capacity [4011, 3926, 4675]. Capelli [216]. caractères [2320, 2977]. Caractérisation [1346, 2952, 961, 2972]. Caractérisations [3020, 452]. Carathéodory [945, 1138, 4334, 353, 1150, 1735]. cardinal [2138, 208, 1439, 2153]. cardinalities [1343, 1875, 1056]. Cardinality [3940, 2318, 625, 2164, 3035, 539]. Cardinals [3697, 2005, 884]. Carleman [747, 915, 2824]. Carleson [4343]. Carmichael [3683, 4060, 4091, 1478]. Carpenter [4132]. carrés [129]. carrying [1303]. Cartan [3875, 4385, 249, 3756, 3375]. Cartesian [4328, 2531, 674, 2275, 960, 2188, 1782, 1284, 1057]. Cartier [4143]. Cascade [3945]. Case [3475, 3227, 659, 2107, 1183, 1018, 2992, 822, 1055, 3820]. Casimir [3959]. Casselman [3235]. CAT [4801, 4348]. Catching [1605]. Categorical [392, 2212]. catégorie [3251, 961]. Categories [4159, 3408, 4136, 4517, 3712, 654, 2402, 296, 196, 2337, 947, 2378, 654, 2561, 130, 1044, 1284, 2566]. Category [4326, 3503, 4139, 4256, 4027, 1165, 1471, 791, 960, 3251, 961, 1749]. Catenoids [3587]. Cauchy [2963, 976, 199, 761, 408, 2378, 2969, 680, 460, 3586, 1340, 2143]. Cayley [3393, 1162, 3631, 4323]. Cech [2331, 39, 607]. Cell [3333, 241, 1339]. cells [4677]. cellular [3251]. Cellulated [3722]. cellules [3251]. Center [4080, 3132, 3035, 650, 1877]. centers [2636, 4810]. Central [4383, 3207, 778, 4374, 3998, 2930, 1670, 3449, 3976, 1008, 1889, 3288, 1635, 1015, 1311, 1531, 624, 854, 966]. centrales [3288]. Centralizing [2238, 1392, 2015, 2678]. Centrally [3726]. centraux [1889]. centred [2839]. Centrification [4740]. centroids [2087]. Cercles [3328]. Certain [1579, 4035, 4058, 4366, 3142, 3232, 3623, 3454, 3186, 4021, 4084, 2144, 4114, 3933, 4463, 4538, 3339, 3371, 3915, 3596, 3597, 4214, 3331, 4095, 4190, 3964, 4449, 1977, 912, 4219, 1507, 3335, 3550, 4701, 2095, 328, 2369, 675, 2065, 1480, 1436, 2267, 972, 2088, 1733, 3043, 637, 906, 2892, 1603, 1800, 2163, 1214, 240, 1746, 4806, 829, 2451, 873, 1942, 1308, 672, 251, 4732, 1039, 1102, 1331, 2411, 570, 2711, 555, 2941, 2123, 1792, 2844, 2600, 459, 1359, 1421, 1360, 815, 767, 835, 1723, 358, 24, 1836, 1077, 3071, 3857, 1474, 816, 3005, 2114, 1706, 464, 1280, 90, 2074]. certain [2395, 1063, 3166, 1446, 1741, 3086, 2849, 2768, 2442, 3056, 2789, 154, 2735]. certaines [248, 1836]. certains [1746, 2600, 1359, 3857, 3166]. Cesàro [4361, 4010, 4621, 400, 2061]. cevian [1238]. Chain [4203, 4599, 83, 2050, 1396, 1246, 2187, 860, 2789, 2611]. chainable [816]. chaines [364, 364]. Chains [4561, 4151, 1501, 1584, 266, 2807, 4769]. champs [3274]. Chandra [3614]. Change [4332, 3798, 3805]. Changes [4012, 4633]. changing [4301]. Chaotic [4481]. Character [4401, 4131, 4383, 4597, 4083, 4251, 4186, 3511, 4299, 3983, 4007, 4722, 2133, 2261, 195, 3200, 1224, 1737, 172, 2590]. Characterisation [4470, 2847, 1729]. characterise [1567]. Characteristic [4058, 3182, 4575, 3830, 4388, 4311, 3478, 4130, 2851, 903, 2687, 1039, 1102, 3040, 4661, 2044]. characteristics [4699, 4668, 1977]. Characterization [4457, 3425, 4426, 3143, 2640, 3870, 3429, 4311, 3914, 3401, 3855, 4051, 4431, 3526, 4101, 3585, 3319, 2047, 786, 2524, 2993, 2671, 2069, 2528, 1552, 2952, 779, 582, 2212, 2216, 2423, 147, 2407, 1975, 4812, 423, 2822, 2784, 1925, 2516, 961, 2942, 3066, 3020, 516, 358, 2433, 2581, 2972, 1585, 2701, 1911, 4724, 597, 725, 1589, 671, 1316, 2159, 273, 1388, 1505, 1808, 2887, 1346]. Characterizations [4369, 4164, 4333, 4271, 2337, 2004, 3565, 3911, 3912, 4411, 4192, 866, 3964, 4751, 4769, 4415, 3742, 4549, 4731, 3146, 4400, 2685, 3062, 1029, 527, 486, 310, 4804, 395, 452]. Characterize [4122]. Characterizing [4200, 3353, 3069, 410, 2655, 639]. Characters [4106, 4176, 3010, 4374, 2866, 4239, 2320, 2977, 3192]. charged [175, 174]. Chasing [3676]. Chebotarev [4060]. Chebyshev [1353, 1122, 1270, 1398, 1518, 1848, 3852, 4050, 4222, 4816]. Chen [4193, 3962]. Chern [3939, 3960]. Chevalley [4521]. Chinese [1593]. Chirally [4662]. Choices [3460, 282, 1979]. choosability [4303]. Choquet [402]. Chow [3459, 3282, 4275, 4750]. Chowla [2478, 2115, 638]. Christ [4425]. Chromatic [4034, 292, 4357, 4492, 281]. Circle [3906, 3537, 3694, 3133, 2975, 4776, 2901, 5, 492]. circles [198, 259]. circuits [365, 1487, 1184, 85]. Circulant [4410, 4363, 4758, 1464]. circulants [197]. Circular [3635, 2890]. Circumscribed [3809]. Circumscribing [523]. CIV [2660, 2335]. Clamped [3908]. Clark [985]. Clarke [3030]. Class [3499, 3620, 3883, 3245, 4247, 4601, 4016, 4597, 4552, 3930, 2659, 3287, 3722, 3395, 4186, 3740, 4378, 4214, 3067, 3306, 3939, 3960, 3569, 3542, 3309, 3135, 4658, 3423, 3391, 3255, 205, 3008, 2794, 2180, 327, 1920, 2088, 779, 2704, 2982, 3830, 719, 590, 2619, 1582, 1962, 3170, 1625, 1167, 1992, 1978, 348, 2691, 287, 399, 1402, 1101, 2113, 873, 1219, 2660, 2129, 2711, 480, 2880, 388, 3199, 2335, 3815, 1851, 1421, 2092, 1868, 2945, 2131, 2647, 247, 36, 1628, 2845, 4786, 3756, 2763, 2730, 3005, 1335, 2807, 2847, 963, 1053, 853, 2666, 3192, 1151, 1822, 1254, 1490, 1599]. class [1055, 1989, 367, 1776, 1092, 854, 966, 273, 49, 1932, 2468, 1764]. class- [1053]. classe [3237, 49]. Classes [3193, 4347, 3546, 3108, 4640, 3474, 4047, 4538, 3446, 3432, 3643, 3912, 4095, 4154, 3659, 3806, 4263, 178, 2633, 2740, 1997, 2155, 410, 3043, 955, 1616, 1931, 2635, 2228, 2779, 661, 2215, 1039, 1102, 2664, 1810, 1811, 2644, 1249, 2804, 4811, 1384, 767, 835, 1041, 4727, 615, 3857, 1233, 876, 1521, 2112, 1949, 1433, 2106, 3840, 1256, 2490, 1614]. Classical [4419, 3506, 4410, 3694, 3235, 4432, 4330, 589, 2425, 2684, 2787, 4261]. classical-type [2425]. Classification [4420, 4159, 3337, 3906, 3186, 3564, 4337, 4659, 2700, 3645, 2589, 3119, 3318, 4454, 1747]. classifications [27]. Classifying [3408, 3002, 2361]. Claws [4492]. Clean [3699, 4730]. Clifford [4699]. Clifford-cyclotomic [4699]. clique [2818]. Cliques [4492]. Cliquishness [3992]. Close [4531, 3281, 1402, 2451, 4685, 1888]. close-to-convex [1402]. close-to-normal [1888]. Closed [4246, 3093, 4371, 4466, 2915, 4214, 4033, 4103, 4102, 3947, 2091, 2427, 1902, 2836, 3370, 700, 1397, 1765, 2799, 1070, 1207, 1700, 2455, 1247, 960, 2879, 1683, 4773, 4685, 2677, 1113, 2507, 1233, 363, 2328, 1284, 1119, 2809, 4783, 2948, 1023]. Closure [4114, 3873, 2609, 3023, 1440, 3768, 3752, 884, 1420, 1292, 2496]. Closures [1931, 3869, 3340, 2391, 2768]. Cluster [4441, 249, 1709]. CM [3440]. Co [4344, 4345, 4554, 2150, 4488, 4566, 2420, 960, 254, 4546]. Co-annihilating-ideal [4344]. co-complete [960]. co-Hopf [2420]. Co-Kähler [4566, 4546]. Co-maximal [4345]. co-null [254]. Co-rank [2150]. coadjoint [4752]. Coarse [4577, 2825]. coarseness [1643]. Coaxer [4377]. cobordism [2886]. cobordisms [1510]. Coburn [3422]. Cockcroft [3181]. Cocommutative [3264]. Cocycles [4437, 2576]. Cocyclic [2233]. Codazzi [3589]. Codegrees [4597]. codes [4694, 1275]. Codimension [3426]. coding [2807]. Coefficient [3734, 1817, 3445, 2724, 3019, 2475, 1949, 4217, 2722, 1721, 2967, 546, 1091]. Coefficients [3090, 3123, 4134, 4233, 3856, 4528, 3899, 356, 2552, 2323, 2450, 740, 1807, 1619, 1537, 2300, 1403, 2514, 1088, 2182, 2416, 720, 1893, 2140, 1020, 1648, 2682, 1477, 2269]. Coessential [4027]. coevaluation [2233]. Coexistence [1543]. cofibres [2357]. cofinite [2820]. Cofiniteness [3907]. coglueing [1814]. Cohen [4365, 3107, 4484, 2934, 3456]. Cohen-Macaulay [3107]. coherence [2522, 1284]. Coherent [1660, 3646, 1598]. cohesive [4681]. Cohomogeneity [4315]. Cohomological [4043, 3611, 3026, 1825]. Cohomologically [4133]. cohomologie [3166]. cohomologies [2840]. Cohomology [4038, 4365, 3104, 3904, 4307, 4287, 3886, 3907, 3591, 3562, 4184, 4517, 3933, 3477, 3611, 3613, 3915, 3346, 3646, 4239, 4242, 4300, 3806, 1997, 1066, 2219, 3814, 2123, 2744, 2971, 1432, 3166]. Coincidence [3909, 1731, 3348, 429, 1944, 2072, 350]. Coinvariants [3808, 2761]. Cokernels [4358]. colimits [541]. Colin [985]. collections [1393]. collectionwise [786]. collinear [1183]. collinearity [1153]. collineation [2737]. Colombeau [3678]. color [2414]. colorability [1782]. colorable [1087]. Coloring [801, 4211, 3164, 901, 2058]. Colorings [3617, 1550]. colors [801]. Colour [955, 2212]. coloured [2771]. colourings [101]. Comaximal [4127]. Combinations [4496, 379, 918]. combinatiorial [1306]. Combinatorial [3744, 1483, 190, 316, 673, 551, 895, 359, 811, 780, 1372, 617, 2492, 274, 874, 2034, 795, 519]. Combinatorially [4288, 4150]. Combinatorics [3410]. combined [175]. Comment [4590, 2242, 782, 3715, 1730]. Comments [2351, 1347, 1049]. Common [4585, 3092, 4142, 3400, 4527, 1551, 2655, 666, 1149, 53]. Commutant [3150, 3317, 4731]. Commutants [4185]. Commutatif [4087]. commutatifs [615]. Commutation [4228, 4125]. Commutative [4344, 2057, 3185, 4331, 4064, 3607, 1493, 4373, 4127, 4488, 3464, 3466, 841, 2164, 4087, 1125, 1282]. Commutativity [3321, 4787, 4048, 515, 1457, 1601, 2367, 2813, 2815, 3032, 456, 1753, 1674, 1720, 1524, 2068, 2497]. commutativity-preserving [2813]. Commutator [3349, 1416, 591, 2145, 2974, 1388]. Commutators [4178, 3113, 4355, 3143, 3539, 1781, 3965, 2474, 2186]. Commuting [4456, 3244, 1324, 4519, 2924, 3183, 3042, 1984]. Commutivity [3317]. comonotone [1951]. Compact [4383, 3620, 3883, 4403, 3485, 4082, 3559, 3842, 4178, 4015, 4350, 4353, 4354, 3971, 2623, 2676, 4141, 623, 4145, 4089, 3227, 4629, 4639, 4128, 4007, 1778, 3985, 4198, 839, 4155, 2448, 2546, 2240, 1719, 2429, 1679, 3074, 2520, 2773, 3012, 1746, 1852, 662, 2692, 2389, 2432, 540, 3815, 1251, 1756, 4754, 2728, 1090, 3082, 2648, 1636, 2564, 911, 1993, 1822, 1227, 1956, 2830, 2766, 4815, 3054, 2361, 1287, 1729, 2245, 2263, 3055, 3058, 3057, 2503]. Compacta [3214, 2822]. Compactification [4534, 3333, 2331, 2078, 607, 891, 920]. compactifications [1793, 1462, 1975, 1169, 1954]. Compactness [3105, 1927, 3430, 2062, 3437, 1105, 2593, 2218, 2155, 1404, 1335, 2961, 2936, 2810]. compacts [1746]. companion [919]. Comparability [3110]. Comparaison [2326, 2896]. Comparing [2548]. Comparison [3990, 4064, 2896, 2708, 2641, 3406, 3514, 1853, 1326, 1837, 2264, 2505, 2326]. compatibilité [2484]. Compatible [4332, 658, 2152, 849, 991]. Competition [4453]. Complement [3482]. Complementarity [1106]. Complementary [4511, 3161, 4688, 1141, 1978]. Complemented [3948, 3395, 2732, 127, 2782, 3956, 1775, 4783]. Complements [1739, 2270]. Complete [1681, 4684, 3867, 4133, 4443, 1763, 4659, 4093, 3664, 3243, 561, 2069, 292, 635, 1165, 236, 1625, 2861, 4812, 2409, 960, 1617, 2281, 1155, 2643, 2134, 2169, 247, 305, 2227, 1495, 1504, 2605, 911, 2681, 1253, 172]. Completely [4649, 2513, 1990, 162, 1861, 2917, 607, 2650]. complètement [635]. Completeness [4040, 3533, 4191, 2685, 2669, 2045, 2144, 1150]. completing [694]. Completion [4760, 1767, 2782, 2879, 2378, 1541]. Completions [975, 1485, 1760, 2412]. Complex [3352, 4384, 4403, 4650, 3426, 3654, 4427, 4168, 4464, 4337, 4207, 3935, 3297, 3565, 3594, 3911, 3213, 3595, 4124, 4524, 3589, 3866, 4018, 4257, 4329, 4544, 3512, 4586, 3119, 3700, 3847, 1579, 951, 2993, 1006, 2853, 1997, 3829, 2997, 1931, 1122, 4753, 2223, 2910, 1969, 2281, 2184, 2232, 2863, 2784, 2341, 2279, 2727, 2917, 1585, 2108, 1833, 1917, 383, 414, 2159, 1946, 4761, 2400, 2467, 4733, 2611]. Complexes [4105, 4385, 4348, 3670, 3916, 881, 4808, 627, 2826, 1998, 626, 261, 2611]. Complexifying [4155]. Complexity [3707, 4516, 3298]. Complicated [2947]. Component [4453, 2858]. Components [4616, 1804, 816]. Comportement [2268]. Composed [4287]. composite [2903]. Composition [3366, 3105, 3639, 4405, 4164, 3994, 4407, 3383, 180, 3947, 2546, 345, 2150, 4822, 4723, 2810]. Compositions [4070, 1351, 267, 88, 16]. composits [3047]. Compound [3870]. Compound-Exponential [3870]. Compressed [4657]. computability [113, 191, 370]. Computation [189, 4258, 113, 191, 370, 2032]. computed [2656]. Computing [3740]. comultiplication [630]. concave [2938]. concavity [4747]. Concentration [3523]. concept [1515, 875]. Concepts [1355]. concernant [129]. Concerning [4275, 3855, 469, 2873, 2613, 1834, 1581, 217, 487, 235, 165, 264, 1203, 129, 931, 75, 985, 1262, 2077, 2354, 2314, 807, 3053, 2586, 1295]. Concordance [3604, 4633, 4480]. Condensed [3320, 3658, 2091, 2391]. Condenser [4011, 4675]. Condition [3592, 3711, 3302, 4319, 4282, 4364, 3220, 3743, 2094, 1396, 1679, 1108, 884, 860, 119, 317, 2806, 1000, 3784]. Conditional [2542, 1029, 1350, 412, 2483]. conditioned [2621]. conditioning [1311]. Conditions [1268, 3930, 2674, 3372, 3449, 3866, 1818, 4380, 4221, 1947, 4802, 2050, 1813, 3024, 1517, 644, 1182, 1990, 2362, 2187, 2675, 2640, 1420, 433, 1958, 817, 989, 967, 2789]. Conductor [3544]. Cone [3669, 4114, 4667, 3344, 2408]. Cones [3109, 3977, 3254, 1006, 1106, 1415, 604]. Conference [3850]. Configurations [3718, 2841, 1862]. confined [87]. conformal [4663, 653]. Conformally [2408, 4733]. Congruence [3271, 2654, 3474, 4135, 3063, 3395, 4372, 143, 2133, 582, 2411, 948, 1628, 314, 367, 605]. congruences [2704, 1078, 4668, 2116]. Congruent [3813, 4780, 1973]. conic [73]. Conical [4520, 3966]. conics [1080]. conilpotency [574]. Coniveau [3573, 4539]. Conjecture [3080, 4573, 4642, 3850, 4667, 3581, 3726, 4396, 4653, 3302, 3416, 4190, 4503, 4001, 4360, 3164, 4303, 551, 1351, 1232, 2905, 2195, 2854, 1581, 2855, 2478, 759, 734, 1372, 4772, 3753, 738, 2819, 4718, 2742, 360, 2403, 2695, 2107, 595, 705, 391, 771, 2236, 1223, 2294, 2344, 226, 867, 2586, 2049, 4571, 4135, 3235, 3982, 3572].

Conjectures [4019, 4729, 3793, 1957, 2563, 638]. Conjugacy [3546, 4047, 4186, 3446, 2941, 1473, 4095, 4564, 2849, 2633, 2954, 2455, 2664, 1041, 2945, 2490, 1614]. Conjugate [3990, 3248, 3330, 3115, 1060, 2388, 2448, 687, 503, 2307, 1123, 287, 375, 2317, 1714, 2221, 2863, 1366]. conjugation [2792]. Connected [1196, 4124, 4096, 2101, 4473, 1947, 4680, 3259, 275, 1013, 1160, 2707, 280, 531, 373, 1977]. Connectedness [2892, 2193, 2331, 2404, 815, 1706, 993]. connectifiable [2831]. Connecting [3443, 4501, 501]. Connection [2105, 3962, 4193, 4329, 2682, 4244, 2019, 2519, 1168, 4789]. Connections [3702, 3991, 4192, 3437, 2851, 2925]. connective [2184, 2325]. Connectivity [4439, 341, 1048]. Conner [2477]. Connes [4258]. connexe [3259]. connexion [4244, 929]. Conorm [2679]. Conorme [2679]. Consecutive [3637, 3764, 4531, 2121, 2179, 733, 159, 484]. consequences [2340, 2538]. Consistency [949, 241]. consistent [258]. constancy [250, 2689]. Constant [3719, 3545, 3353, 4520, 4189, 3634, 4052, 4414, 382, 2427, 4790, 490, 2343, 2345, 1249, 546, 4768, 182, 2135, 1887, 1059]. constantes [2326]. Constantinescu [249]. Constants [4131, 4557, 2326, 1678, 300, 164, 1928, 512]. Constrained [4588, 2853, 1628]. Constraint [3701]. Constraints [3637, 3247, 3764, 1138]. constructed [1631]. Constructing [513, 2653, 4228, 3751, 3214, 4000, 2198, 2264]. Construction [3650, 3588, 281, 1810, 793, 3614, 3284, 3278, 520, 1, 3087, 1717, 1069, 3184, 2950, 475, 1541, 4779, 2117, 1054, 85]. Constructions [3969, 4606, 476, 369]. Constructive [4132, 2681, 790, 910]. Contact [4620, 3895, 4634, 2671, 2835, 2503, 3199, 2715, 4126]. contain [333]. contained [240, 1761, 2449]. Containing [3275, 2964, 404, 3025, 2619, 2017, 1924, 1431, 373, 598, 877, 1424]. contenant [2182]. Content [1588, 4581, 3949, 1692]. contingent [1303, 1582]. Continua [3455, 3846, 4119, 2398, 2707, 1783, 2578, 816, 2683, 3069]. continuation [2131, 1854, 81, 1991]. Continued [4245, 4636, 3176, 3449, 2927, 1714, 2662, 2393, 1446, 519]. continues [3428]. Continuité [2320]. Continuity [4081, 2797, 3737, 4413, 4623, 4675, 3102, 1394, 2546, 2906, 1849, 2861, 2407, 2466, 4698, 526, 1612, 1095, 2320]. continuity-like [2906]. Continuous [2240, 4650, 4438, 4551, 1793, 2596, 663, 3996, 3175, 3894, 4608, 1586, 3963, 3463, 2975, 3819, 4217, 2049, 328, 2047, 2546, 3060, 1986, 2085, 2218, 1902, 944, 706, 2329, 2162, 2306, 2381, 857, 1462, 3027, 3794, 1398, 34, 68, 3428, 1291, 2707, 80, 124, 423, 831, 910, 1563, 70, 2943, 1882, 1934, 2209, 823, 652, 93, 2073, 1287, 2567]. continuously [1684]. continus [1143]. contour [1061, 1644]. contractification [1546]. Contraction [4262, 1330, 548, 1534, 980, 455, 807, 634]. Contractions [1350, 4296, 1497, 2419, 1005, 1170, 1895, 692, 1704, 1796, 1467]. contractive [1034, 1142, 1248, 1206, 703, 717, 3803]. contragredient [2858]. Control [3439, 4053, 298]. controllable [1564]. Controlled [3188, 4470]. Convergence [3901, 3926, 3276, 1274, 2131, 3981, 3228, 667, 543, 3945, 3716, 3743, 2244, 1580, 1986, 2085, 2793, 4801, 40, 308, 440, 576, 637, 1679, 2396, 1302, 1496, 1271, 402, 2662, 2822, 2078, 847, 215, 3, 1364, 2486, 2079, 1159, 714, 2241, 2920, 1490, 312, 1094, 238, 1355]. Convergent [4225, 4031, 3424]. Convergents [4562]. Converges [3624]. Converging [3502]. converse [557, 1623, 430, 1064, 2587]. conversion [477]. Convex [4283, 3719, 2964, 3969, 3843, 4015, 3548, 4496, 4165, 4442, 4389, 3296, 3854, 3726, 3567, 3400, 3977, 3630, 1445, 3480, 4173, 3819, 3721, 3555, 952, 1698, 2310, 1106, 4790, 3787, 1033, 1621, 457, 1244, 1702, 662, 1883, 304, 1402, 2030, 2343, 2407, 1519, 1994, 3776, 1553, 2821, 1520, 985, 560, 3018, 997, 1292, 806, 2957, 1731, 4631, 45, 1252, 14, 2262, 2746, 2314, 1157, 312, 2192, 2086, 1821, 3054]. Convex-normal [4389]. convexes [1620, 1899]. Convexity [4086, 42, 3174, 2251, 1415, 1122, 4747, 2341, 1277, 280, 531, 3038, 1448, 1412]. Convolution [3468, 3619, 2410, 3622, 3452, 3453, 2296, 832, 3938, 3084, 4413, 3218, 3421, 3981, 3149, 2738, 3792, 1747, 719, 2637, 1499, 2434, 1684, 2719, 1797, 3021, 1024]. convolution-type [1747]. Convolutions [2607, 4173, 1738, 334, 573, 572, 352]. Conway [595, 867]. cooking [31]. Coordinated [3309]. Coordinates [3713, 4489, 2944]. Coordinatization [3257]. copies [2980]. coproducts [1877]. copure [4280]. Copy [3221]. Coquecigrues [3569]. Core [4034, 4369, 3308, 3648]. coreflective [547]. Cornea [249]. corners [1716]. Corona [4799, 3580]. corps [1715, 4756, 1509, 3612, 3857, 2972]. Correction [4182, 3497, 268, 1189, 1102, 145, 2762, 1914, 776, 3527, 613, 657, 1939]. Corrections [922, 921, 835, 836, 837, 2584, 60]. correlated [845, 1062, 2210, 1914]. Correlations [3951, 2084]. Correspondence [3551, 2513, 4792, 413, 1294]. correspondence-covers [4792]. Correspondences [4463, 3253]. corresponding [2363]. Corrigendum [3764, 2205, 1161, 2081, 4193, 4361, 4609]. Coset [1539, 207]. Cosine [4327, 1215, 532, 2374]. Cosmetic [4662]. cosymplectiques [248]. Cotangent [4613, 2167]. Cotype [3425]. Countable [4555, 3397, 4213, 4076, 1745, 4776, 4694, 2392, 1070, 1207, 1530, 1072, 2831, 1548]. countably [1344, 1993, 2449, 3057]. counter [1284]. counter-example [1284]. Counterexample [771, 734, 1633, 1419, 1884, 360, 2961]. Counterexamples [2870, 1430, 2695, 2314, 2192, 2811]. Counting [4497, 3342, 4091, 3402, 3460, 4447, 2114]. Coupled [4620, 4572, 3664]. Courant [4467]. courbe [2963, 2297]. Courbes [3184]. covariance [2057]. covariances [1204]. covector [4714]. Cover [4508, 4509, 4568, 4569, 4589, 4627, 4628, 4648, 4686, 4687, 4705, 4706, 4766, 4212, 1408, 441, 607]. Cover-avoidance [1408]. Covered [3978]. Covering [3193, 4084, 4465, 3728, 1545, 3721, 1458, 802, 1940, 930, 2171, 553, 2461, 1253, 1590]. Coverings [3095, 3730, 3552, 4790, 2304, 457, 2030, 2999, 2119, 2098, 14, 2488, 2764]. Covers [3785, 3356, 4792, 2175, 2654, 871, 1906, 1059]. Cox [83, 4515]. Coxeter [4825, 3575, 2601, 2899, 3055, 4452]. CR [2751, 3621, 2890, 3418]. CR-structures [2751]. Craig [3888]. Creating [3554]. critère [2079]. Criteria [3323, 3032, 3304, 3976, 4593, 4690, 1003, 1395, 272, 1534, 2149, 1012, 1851, 600, 368]. Criterion [4046, 4469, 4260, 2218, 3026, 1923, 1125, 1675, 1662, 2014, 1183, 2917, 2012, 3052, 640, 3038, 544]. Critical [1550, 3703, 3560, 3955, 3404, 4448, 3428, 2946, 3802, 2933, 2294, 2747, 4762, 1409, 3784]. critical-point [3428]. critiques [3428]. croissance [3259]. Cross [3713, 4700]. Cross-Ratio [3713]. cross-ratios [4700]. Crossed [3747, 3329, 2760, 2929, 4754]. Crossing [4633]. Cryptic [4150]. cryptography [2559]. Crystallinity [3298]. CS [2937]. CS-modules [2937]. CSL [2666]. Cube [4348, 365, 1643, 1403, 510, 331, 330, 332, 1184]. cubes [1527]. Cubic [2064, 4039, 4351, 4226, 4068, 4318, 4502, 3632, 4382, 3544, 2252, 536, 1193, 2693, 394, 64, 2545]. cubics [2120]. cuboid [1528, 1650, 1228]. cuboïde [1650]. cuboids [1823]. Cuntz [3268, 4220]. Cup [2219]. cure [973]. Curvature [4246, 3559, 3950, 4536, 4520, 4231, 3448, 1983, 4257, 4126, 4645, 4052, 4664, 3406, 3514, 4720, 2835, 3789, 2694, 4767, 2916, 32, 1977]. Curve [4441, 4427, 3114, 4124, 3894, 3696, 4434, 4763, 55, 653, 2467]. Curved [4475, 4246, 3448, 3188]. Curves [4304, 382, 4510, 4570, 4605, 4246, 3271, 3440, 3749, 3867, 4020, 4483, 3595, 4524, 3975, 3478, 4341, 4342, 4504, 3236, 3334, 3312, 4665, 4473, 4055, 3335, 4130, 3456, 2851, 2872, 3149, 3167, 327, 2551, 2855, 3184, 2557, 3752, 1203, 2317, 3013, 4712, 3813, 2119, 2697, 2265, 4811, 2473, 2805, 4768, 2582, 3084, 1830, 1983, 2488, 2719, 2035, 4805, 1567, 4807, 2053, 1791, 1059, 4661, 2125, 2049]. Cusp [3121, 3688, 4233, 3899]. cusped [2987]. cuspidal [2473]. Cut [2901, 4673, 2718]. cut-off [4673]. Cutler [593]. cutpoint [497]. cutset [2526]. Cutting [1843]. CW [881]. CW-complexes [881]. cycle [766, 2457, 900]. cycled [1942]. Cycles [418, 3459, 4255, 4615, 4123, 4653, 2082, 3631, 4279, 3543, 4055, 2570, 1656, 1400, 2451, 2693, 4396, 117, 222, 1048, 598, 1253, 2720, 1185]. Cyclic [4348, 3272, 3287, 4517, 3954, 3656, 3628, 3382, 2726, 3128, 3384, 3460, 3217, 3542, 3552, 3544, 3103, 3880, 4104, 1694, 320, 1376, 2415, 437, 755, 2710, 2744, 3159, 198, 2826, 4666, 1198, 990]. cyclically [1696]. Cyclicity [4513, 2855]. cyclicly [2092]. cyclics [3060]. Cyclotomic [4573, 3159, 1872, 2852, 1992, 4699, 1444]. cyclotomiques [3338]. Czipszer [587].

D [2149, 3826, 3760, 4057]. D-symmetric [3826]. D. [577, 734, 2149, 803, 411, 360, 889, 974, 1547]. d.g [1494]. Dahlberg [3344]. d’Alembert [3485]. d’algèbres [2600]. damped [3775]. damping [2720]. d’André [3166]. d’anneaux [178, 615, 2465, 49]. dans [2320, 2042, 1121, 1746, 3582, 1359, 1361, 2679, 120]. Darboux [2256, 30, 299]. DARMA [3031]. Data [3267, 4293, 4356, 2443, 530, 2969, 2147, 949, 31]. d’automorphismes [3612]. Davison [2434]. Decay [3427, 3919, 3075, 3154, 2147, 3757, 4721]. Decaying [4330]. Deciding [4617]. Decomposability [4553]. decomposable [1798]. décomposables [1798]. Decomposition [3515, 4764, 4521, 3283, 3932, 1708, 1725, 933, 3333, 3633, 1314, 4581, 2065, 2922, 2633, 2415, 2352, 2937, 2489, 2967, 908, 2417, 1510, 4090, 2279, 1882, 1612, 670, 210]. Decompositions [4293, 2357, 4238, 3602, 4033, 4506, 2002, 1297, 1482, 1625, 2861, 804, 1180]. Dedekind [2685, 320, 3501, 4756, 4809]. Dedication [3717]. defect [3159]. Defectless [3925]. deficiency [2568, 1730]. definability [2145]. Definable [4213]. Define [4555]. Defined [3637, 3708, 3306, 3314, 3764, 2513, 308, 2010, 1074, 4785, 2200, 2561, 63, 3817]. définis [2010]. Definite [3129, 3879, 594, 3824, 903, 946, 4673, 181, 472, 92]. Definiteness [3341, 4629, 1646]. Definition [112, 3118, 2951, 1212, 1898, 1801, 881, 910]. definitions [3771, 41]. Definitizable [2904]. Deformation [3816, 2962]. Deformations [4491, 2635, 3297, 3211, 4474, 2751, 2979, 2394, 2135, 2135]. Deformed [3394]. Deforming [3871]. Degeneracies [3847]. Degeneracy [3488]. Degenerate [4403, 4438, 4478, 4240, 3149, 3867, 1488, 1340]. Degenerations [4537, 4341]. dégénérées [1410]. degré [3288]. Degree [4034, 3206, 3442, 3287, 3098, 4718, 3511, 4615, 4623, 4504, 3553, 4362, 1690, 2124, 3288, 2747, 4719, 2067, 592]. Degrees [4131, 4186, 1250, 3659, 2159, 1240, 2590]. Dehn [3106, 4116, 3345, 3554, 4075, 3351]. Delay [4438, 4478, 4340, 1187, 1012, 2722, 2939, 2362, 2495, 1890, 2214, 3052]. Delayed [4381, 4057]. Delays [4297, 2303, 3783]. Deligne [4288, 4676]. Delone [3298]. delta [1261]. Demi [371, 364]. Demi-groupes [371, 364]. demi-groups [371, 364]. Demonstration [699, 699]. Demuskin [2399]. Dendrites [3286, 485]. Denjoy [2094, 910, 2359]. dénombrables [2304]. Dense [4603, 4206, 4213, 4076, 250, 4776, 4737, 2413, 1313]. denseness [1758]. densest [259, 331, 330, 332]. densité [4756]. Densities [1014, 3419, 2267, 632, 892, 1283]. Density [4383, 3767, 105, 3372, 4502, 3217, 1521, 4100, 4002, 227, 2690, 4756, 2554, 1654, 1806, 2210, 1020]. dentability [1774]. denumerable [2304]. Dependence [1518, 3113, 3963, 1958, 1666, 450]. dependencies [749]. Dependent [3081, 2446, 1734, 1980, 1979]. Depths [4219]. Derivable [283]. derivates [614]. Derivation [4148, 1194]. Derivations [3989, 1959, 2311, 2978, 3736, 4286, 4111, 4405, 4085, 2336, 3136, 2767, 2949, 1974, 2205, 2213, 2895, 1596, 1685, 2273, 517, 1661, 2897, 2439, 2454, 2500, 2902, 919, 2629, 1778, 2948]. Derivative [3703, 3536, 4337, 3894, 3512, 95, 636, 1987, 1246, 1260, 1650, 1764, 2810]. derivative-free [1987]. Derivatives [3483, 4227, 4269, 4464, 4630, 4472, 4600, 4609, 4549, 2280, 2906, 2197, 1829, 2196, 2366, 2691, 129, 329, 188, 550, 524, 1348, 2680, 4794, 833, 2117, 2143]. dérivé [1650]. Derived [1383, 1757, 3647, 1528, 2223, 287, 375, 399, 2945, 1280, 1228, 2566]. dérivée [636]. dérivées [1987, 129]. Derksen [3751]. Desarguesian [1570]. Descent [4180, 3608, 4683, 1152]. describe [1081]. Description [3927]. Descriptions [2687]. descriptive [2951]. design [4087, 494, 209]. Designs [3606, 4289, 487, 2054, 2206, 1038, 1543, 851, 224]. Detect [4370]. Detecting [3543]. Detection [1913]. Determinant [4291, 4520, 477, 76, 2165]. Determinantal [4316, 4397, 92]. determinants [2036]. Determinateness [1028]. Determination [3588, 1608, 2541, 1004, 2132]. Determined [4288, 4115, 4600, 4609, 690, 944, 706, 309, 409]. Determining [2480]. deterministic [2376]. detours [495]. deux [1620, 178, 2896, 2556, 129, 615]. deviation [358, 79]. deviations [1670]. DeWitt [4182, 4113]. DG [4488]. DG-rings [4488]. DGLA [4677]. d’hamiltoniens [3237]. d’hypersurfaces [2135]. Diaconis [3848]. Diagonal [3808, 3474, 3436, 676, 4811, 537, 220, 1387, 454]. diagonally [4693, 1434]. Diagonals [936, 1681, 2954, 1530, 596]. Diagram [4382]. Diagrams [4530]. Dialgebras [3569]. dialogue [218, 1349]. Diameter [4553, 4252, 508, 4323, 533, 2234]. Dichotomy [3877]. Dicke [1294]. Dickson [1162, 2290, 2227]. Diffeomorphism [4235]. diffeomorphisms [2801, 1839]. Difference [3576, 4356, 2347, 2620, 2705, 2221, 4732, 2942, 3066, 745, 836, 1115]. Differences [3652, 3545, 3092, 647, 4525, 1608, 1841, 1523]. Different [3214, 4821, 2834]. Differentiability [3640, 3480]. Differentiable [3366, 102, 250, 1941, 1099, 294, 652, 1201]. Differential [3900, 3990, 4550, 4438, 4478, 4551, 3930, 3972, 4067, 4255, 4430, 4470, 3493, 4026, 4340, 4593, 3964, 2425, 3942, 2412, 4381, 3763, 1578, 1579, 1003, 1930, 557, 1771, 356, 1517, 1187, 272, 2996, 407, 1390, 1514, 2307, 644, 2776, 788, 1166, 2512, 1607, 1923, 2451, 2495, 4732, 3795, 2014, 2725, 2711, 458, 1195, 2358, 934, 1828, 1705, 1890, 2214, 2514, 459, 2930, 1881, 1985, 3, 861, 1042, 1333, 271, 1115, 1474, 1488, 938, 1366, 3052, 2563, 1336, 2748, 2202, 1672, 3053, 1064, 1315, 3783, 667, 855, 878, 879, 1412]. differential-difference [1115]. Differential-free [4470]. Differentially [3961, 2412]. differentiation [1371]. différentielle [1771]. différentielles [2996]. differing [2121]. Diffraction [3262, 3879]. Diffusion [3963, 4381, 1639, 3203]. diffusions [262]. Digits [4062, 3097, 3301, 4748]. Digraph [3393, 2017]. Digraphs [4422, 318, 2017, 437, 1599]. Dihedral [3808, 3533, 3314, 3058, 4198, 2535]. Dilation [4063]. Dilations [1497, 4181, 4339, 1895, 3803]. Dilworth [2003, 3162]. Dimension [4155, 4365, 4108, 4159, 4643, 4110, 4331, 4134, 4065, 3707, 4045, 1501, 1584, 4409, 4230, 3611, 3299, 4069, 3132, 3523, 4059, 4397, 3464, 4363, 2544, 1376, 3604, 257, 4714, 3810, 3222, 1559, 1216, 2156, 2223, 1825, 2187, 1886, 2624, 2840, 3189, 2371, 2033, 4757, 756, 2465, 1495, 1492, 1911, 3039, 585, 1821, 1780]. dimension-free [3189]. dimension-linking [257]. Dimension-preserving [2544]. Dimensional [3987, 3574, 3733, 3557, 3266, 3903, 3353, 3605, 3929, 3907, 4085, 4113, 4182, 4167, 3643, 3342, 3629, 3919, 4054, 4435, 3390, 3804, 2388, 2612, 1997, 3789, 1010, 1717, 4552, 4681, 1536, 1931, 4266, 1604, 2656, 2674, 4086, 2861, 2389, 4534, 843, 134, 2759, 4535, 4390, 4337, 4208, 2744, 4559, 764, 427, 580, 766, 4481, 2287, 986, 4659, 4213, 1077, 2433, 2882, 1948, 2918, 2731, 4098, 1223, 2884, 558, 3802, 2901, 2992, 2199, 4398, 3313, 4452, 1846, 1929]. Dimensions [4038, 4167, 3214, 4488, 4242, 4300, 2271, 4825, 1083, 1245, 3834, 2539, 1513]. diminish [939]. Dinculeanu [937]. d’indéterminées [2042]. d’interpolation [2874, 850]. Diophantine [3497, 4058, 924, 1107, 1664, 3414, 4068, 4318, 705, 3176, 2400, 3946]. Dipolar [4025]. Dirac [307, 1261]. Direct [1297, 1482, 4186, 3299, 3343, 3977, 661, 609, 794, 1954]. Directed [4572, 1858, 744]. direction [1153]. Directional [4433]. directionally [3151]. Dirichlet [2762, 3468, 1716, 4402, 2094, 3902, 2436, 2510, 4513, 2450, 4799, 4356, 1236, 4579, 4317, 2712, 1332, 1836, 3005, 180, 4703, 1280, 4002, 352, 3021, 1824, 1773]. disc [4670]. Disconjugacy [644, 2347, 1903, 855]. disconnection [3069]. discontinu [1089]. Discontinuity [967]. Discontinuous [30, 1437, 1089]. Discrete [4283, 4404, 4512, 3275, 4308, 4273, 3935, 3877, 3313, 3879, 4595, 4031, 1876, 3466, 2481, 1455, 4284, 1745, 1909, 1536, 4808, 2576, 2303, 4673, 2391, 1675, 3015, 2843, 483, 2494, 964, 2809, 403]. Discrete-Argument [3877]. Discreteness [3352, 3673]. Discriminant [3314, 48, 1962, 1153, 17]. Discriminants [4610, 3595, 2699, 23]. Discs [4442, 3728]. Disease [4297, 973]. disfocality [2347]. Disjoint [4404, 1070, 1207, 1625, 2759, 222, 373]. Disk [953, 4461, 4088, 1332, 1276, 1447, 428]. Dismantlability [2818]. Dismantlable [4196]. disparity [1193]. Dispersive [3919]. displacements [87]. dissections [163]. dissipative [271]. Distance [4651, 4654, 3358, 4092, 4346, 2218, 2520, 2179, 2601]. distant [1311]. distinct [833]. Distinguished [4200, 3292, 4056]. Distortion [2801]. Distribution [3651, 4640, 3457, 3293, 2473, 608, 3305, 964, 3898, 4632, 2382, 716, 2430, 4758, 28, 779, 467, 2553, 2120, 286, 4748, 1153, 504, 639, 741, 933, 435, 2173, 1183, 117, 4743, 472, 680, 302, 1239, 183]. Distributional [4179, 2574]. Distributions [4014, 3870, 3129, 739, 1319, 880, 253, 827, 286, 4826, 1654, 504, 814, 310, 412, 1609, 597, 481, 1427]. Distributive [4265, 2259, 3850, 2031, 1990, 711, 1885, 2411, 3066, 1611]. distributivity [770, 2681]. Ditor [3825]. Divergence [1970, 1022, 4399]. divergence-free [1022]. divergent [1982]. Diviseurs [2042, 1995]. Divisibility [4224, 3659, 221, 2659, 1992, 1868, 277]. Divisible [3542, 4007, 869, 2401, 2918, 1893, 2140]. Division [2907, 3983, 2910, 1574, 1720, 1177, 1178]. Divisor [2909, 3912, 4317, 4029, 4824, 4335, 4205, 4144, 2627, 2077, 4101]. Divisorial [2046, 2154]. Divisors [4203, 2677, 2464, 106, 2019, 2179, 4118, 4143, 448, 1770, 2667, 3179, 2042]. d’Iwasawa [3338]. Dixmier [4571]. do [333, 4745, 90]. d’objectifs [1707]. dodecahedron [801]. does [1001]. Dold [1595]. dolls [1571]. Domain [4224, 3536, 3644, 3634, 4218, 2154, 1920, 1765, 1518, 2384, 932, 714]. domaine [827]. Domains [3023, 3320, 4283, 3841, 4349, 4351, 3770, 3953, 4519, 3190, 3491, 4295, 3658, 3137, 4507, 3392, 2379, 1947, 1395, 2091, 2812, 320, 2050, 1079, 2792, 4741, 1559, 3154, 2401, 2046, 2860, 3046, 2391, 2030, 2343, 2004, 3776, 2200, 2460, 1224, 671, 1021, 4809]. dominant [1434]. dominated [2587, 1467]. dominating [1487]. Domination [3108, 4138, 4205, 4144, 4096, 4029, 2066]. dont [2459, 1889]. Doob [2922]. d’opérateurs [2244, 1746, 3101]. Double [4161, 3454, 4228, 3510, 199, 1691, 4795, 2602, 3177, 2096, 3820, 421, 543, 4697]. doubles [1740]. Doubling [3396, 4030, 4594]. Doubly [4298, 2351, 1867, 116, 75, 936, 1630, 3780, 1117]. Douglas [2080, 4500]. Dow [3091]. Down [3500]. dragons [1725]. drift [3798]. Drinfel’d [2632, 2833, 3182, 3753]. droite [754]. Dual [4347, 4657, 4369, 4015, 4232, 1965, 4192, 2592, 1580, 669, 301, 2282, 2637, 3007, 268, 244, 2407, 1955, 1919, 1531, 1226, 4778, 4638, 1229, 4730]. Duality [3620, 3883, 4459, 3506, 4117, 1106, 1678, 1814, 2761, 2881, 479, 2034, 2936]. dualizing [1216]. Duals [4035, 4511, 1498, 2978, 2306, 978, 1886]. Duarte [4469]. Dubins [2509]. Dubois [3707, 2371]. due [1414, 3848, 920]. Duke [3487]. Dulmage [1630]. d’un [2211, 1889, 2438, 1089, 2135, 452, 754]. d’une [3259, 2963, 2977, 3020, 2079, 49, 2297]. Dunford [2555, 3706, 1869]. Duo [515, 67]. Dynamic [3884, 4210, 3923, 4195]. Dynamical [3508, 4297, 3386, 3630, 3466, 4453, 4718, 1331, 3015]. Dynamics [4053, 4679, 3203]. Dynkin [612]. Dyson [4702]. Dziubański [3757].

E-Courant [4467]. E. [146, 2309]. Each [3221, 418]. EALA [3309]. Earle [1592]. easy [1081, 2684]. EC [4659]. Eckmann [378]. Edelstein [1034, 1248, 962]. Edge [4123, 536, 728, 2771, 928, 281, 1782, 222, 4527]. edge-chromatic [281]. edge-colorability [1782]. edge-disjoint [222]. Edge-magic [4123]. Edge-of-the-wedge [4527]. editor [1051]. effacements [1265]. effect [3029]. Effective [3669, 4732]. Efficient [2368]. Efroymson [3707, 2371]. égaux [97]. Egg [4218]. Egoroff [202]. Egyptian [1295]. Eiconal [3927]. Eigencurve [3534]. Eigenfunction [3154]. Eigenfunctions [3987, 3439, 3695, 2387, 2640, 2745]. Eigenvalue [3997, 926]. Eigenvalues [4081, 3353, 4574, 3908, 3999, 4319, 3550, 56, 2924, 4692, 2838, 682, 1166, 1327, 2641, 150, 1093, 2505]. eigenvectors [3101]. eight [1931]. eight-dimensional [1931]. Eilenberg [3382]. Einstein [3991, 3789, 4535, 4536, 4559, 4804, 3985, 4646]. Einstein-like [3789]. Eisenstein [4072, 3355, 4071, 3942, 2115, 368]. Ekedahl [4466]. EKOR [4676]. elation [1482]. elations [2128]. electromagnetic [357, 175]. electromagnetic-gravitational [357]. electrophoretic [2443]. Element [4305, 4160, 4013, 4558, 4324, 4729, 2255, 4774, 2255]. Elementary [56, 295, 3458, 796, 557, 2197, 2008, 157, 2, 2531, 2863, 1195, 916, 1406, 1630, 797, 941, 2032, 1547, 110, 2547]. Elements [4385, 3828, 4043, 4558, 4445, 4412, 553, 2174, 1318, 2510, 1659, 1108, 3771, 2926, 2274, 2273, 1866, 846, 1249, 1473, 1117, 4774, 1158, 1387, 2886, 826]. Elliott [4501]. ellipsoid [523]. ellipsoids [523]. Elliptic [3008, 3104, 3271, 3440, 3752, 3579, 4667, 3371, 3595, 4091, 4124, 3478, 2582, 3236, 4450, 4434, 4241, 4055, 3335, 1454, 1716, 2551, 2855, 682, 701, 787, 4770, 3172, 2708, 4763, 3813, 3815, 2029, 2644, 2141, 2991, 2603, 4805, 600, 1021, 1640, 2467, 4661, 2125, 4690, 2996]. elliptiques [2996]. embed [2989, 3006]. Embeddability [3390]. embeddable [1247]. Embedded [4637, 4253, 4473, 4693, 4819, 2536, 4174]. Embedding [2398, 2402, 4014, 2206, 3095, 4250, 1943, 4579, 4703, 4634, 3070, 465, 2840, 1321, 806, 611]. Embeddings [4490, 4422, 3280, 3521, 3506, 2372, 2683]. emboîté [4819]. Emden [1199]. empirical [467]. Empty [2288, 241, 1994]. Enclosure [682]. Encountered [3609]. End [4092]. End-to-End [4092]. Endomorphism [1147, 2893, 1343, 3814, 2716, 689, 837, 776]. Endomorphisms [3903, 3329, 3403, 3525, 715, 1405]. Endpoint [4395, 4482, 4545, 2163]. endpoints [3027]. Ends [2461]. Energy [3427, 4313]. Eneström [2037]. Engel [2570, 982]. enlacement [3428]. Enlarged [3465]. Enlargements [3896]. Ensemble [3530, 1356]. Ensembles [781, 2304]. entières [2952, 1235]. Entire [3452, 3745, 4614, 3927, 4269, 4555, 4533, 3656, 4356, 4067, 3164, 2995, 2070, 2201, 2366, 1391, 2969, 2744, 2824, 2727, 2991, 228, 354, 681]. entrance [2621]. entre [2874, 1356, 2600]. entries [223]. Entropy [4496, 3431, 2162]. Enumerating [4422]. enumeration [2334, 1539, 2700, 89, 207]. Envelope [4403, 4059, 1068, 416]. Enveloping [4476, 3266, 2962]. Eötvös [1317]. epi [1031, 534, 2793, 296, 2680]. epi-convergence [2793]. epi-reflector [1031]. Epidemic [4053]. Epimorphisms [1084, 1209, 602]. epireflections [1169]. Equal [2103, 3478, 3449, 4129, 256, 882, 97, 733, 1875]. equalities [1416]. equality [884, 1682, 585]. Equalizing [1088]. Equation [3990, 3323, 3427, 4226, 3283, 3414, 3497, 3642, 4516, 3738, 3995, 4312, 4391, 4356, 3657, 3226, 4481, 4522, 4025, 4359, 3218, 4195, 3919, 3240, 4381, 3946, 3984, 3763, 1649, 1930, 1971, 58, 675, 902, 2504, 56, 1771, 1107, 1187, 1242, 3011, 1300, 758, 2126, 644, 1215, 788, 803, 408, 2124, 2722, 2939, 3775, 813, 1073, 1357, 2928, 1856, 1403, 532, 934, 1358, 1985, 2147, 263, 1115, 1859, 1488, 2425, 4746, 2540, 1984, 175, 1315, 575, 4762, 707, 890, 544, 2545, 1771, 1859]. Equational [1105, 1343, 661, 876, 1335, 1919]. Equations [3900, 4620, 3884, 4403, 3485, 4438, 4478, 4551, 4247, 3927, 4351, 4016, 3275, 3930, 3655, 4268, 3707, 3407, 3972, 3973, 3955, 4559, 4625, 4068, 4318, 4430, 3976, 4026, 4340, 4593, 3963, 4448, 4543, 3942, 4380, 3664, 4008, 328, 1578, 1395, 1454, 557, 1698, 1716, 675, 924, 1664, 356, 1137, 272, 2996, 3027, 721, 915, 407, 1390, 701, 787, 1514, 2655, 2373, 1214, 2347, 2307, 1012, 1853, 2620, 2776, 957, 1166, 2512, 1607, 1923, 2362, 2451, 2495, 2758, 355, 601, 2708, 200, 251, 4732, 3795, 2014, 2725, 2711, 458, 1195, 3065, 1828, 1980, 1705, 2695, 1851, 2029, 2214, 2514, 459, 2698]. equations [1881, 745, 836, 3, 2253, 1042, 1333, 546, 1860, 169, 3800, 1854, 3176, 1474, 938, 1366, 3052, 852, 2563, 1280, 2748, 2202, 466, 1693, 1672, 3053, 1064, 1315, 3783, 600, 1021, 1945, 824, 667, 913, 3760, 1199, 2720, 174, 2400, 3203, 1873, 855, 878, 879, 1027, 2996]. equi [208]. equi-cardinal [208]. equiaffinities [1025]. Equicardinality [678]. equicontinuity [1097]. equidistant [404]. Equilateral [4152, 3786]. Equilibrated [1842]. équilibrés [1842]. Equilibria [4475]. Equivalence [3532, 3010, 4694, 958, 3112, 4336, 4188, 4000, 3986, 1460, 1813, 1679, 1931, 2484, 4736, 4738, 4782, 2826, 2359, 583, 1094, 2484]. Equivalences [3578, 4417]. Equivalent [3771, 3046, 4121, 4429, 785, 2612, 2272, 2805, 1887]. Equivariant [4307, 4042, 3591, 4184, 2476, 4450, 2645]. Equivariantly [4440]. ERD [2763]. Erdos [190, 278, 316, 673, 84, 1450, 3812, 4812, 3434, 950, 3697, 545]. Erdos-Renyi [1450]. Ergodic [3519, 4590, 3248, 4000, 2834, 2690, 2385, 2170, 2624, 1921, 1704, 1908, 1996, 2104, 2254, 2199, 1545, 2587, 2786, 1467]. Ergodicity [3439, 4240, 2063]. Errata [1207, 1160, 1619, 1917]. Erratum [3883, 3294, 3875, 4072, 4103, 3319]. Error [3487, 2048, 336, 1512]. Escaping [4611]. Escott [107]. espace [1258, 1899]. espaces [2874, 1746, 2556, 2841, 1359, 1346]. Essential [3105, 4643, 4185, 1802, 3710, 3299, 3383, 4069, 4097, 3317, 2606, 4731]. Essentially [4495, 4097, 3716, 2101]. Estimate [3968, 4039, 3487, 3202, 4545, 2163, 3013, 3189, 1472, 2845, 3757, 434, 4723, 2591, 2903]. Estimates [3427, 3473, 4355, 3539, 3797, 3339, 3226, 4482, 4070, 3919, 3450, 3451, 4197, 2448, 3149, 2618, 3796, 1949, 489, 2832]. Estimateurs [827]. Estimation [3598, 2134, 987, 231, 203]. estimator [757]. estimators [892, 1283, 1338]. Eta [3501, 3157, 4259, 2657, 2246]. étale [2467]. États [1121]. étroite [2079]. Euclidean [3516, 2398, 3531, 2288, 4651, 1080, 3520, 4495, 3952, 4443, 3343, 2032, 969, 4567]. Euler [4388, 1743, 4699, 3018, 1472, 3939, 3960, 2806, 4668, 319, 1151, 1741, 2586, 3760]. Eulerian [1653, 1528, 291, 1650]. eulérien [1650]. Evaluation [3501, 487, 3028, 1039, 1102, 2009, 2735]. evaluations [2866, 2585]. Even [4624, 3997, 2171, 3874, 4504, 4549, 625, 69, 2275, 1828]. Eventual [2347]. Every [4650, 4752, 3576, 2865, 2831, 4324, 1755, 2427, 404, 4774, 729]. Everywhere [3981, 3424, 250, 409, 652]. Evolution [3999, 3877, 3860, 1698, 546, 3800]. Exact [4243, 3483, 253, 4256, 435, 4199, 880, 2196, 541, 2033, 2712, 2762, 1183]. exactes [2712]. exactly [2700]. Exactness [3769, 378, 232]. Example [1129, 264, 1897, 2863, 1284, 722]. Examples [2235, 4515, 3190, 3980, 127, 1722]. exceeding [744, 46]. Exceptional [3785, 3224, 3491, 4075, 4435, 3602, 1898, 506]. Excluding [3397]. exclusive [2398]. Exemples [2465]. Existence [2750, 1844, 3929, 3606, 4289, 3930, 3590, 3476, 3642, 3407, 3972, 4767, 4617, 934, 804, 4625, 1042, 24, 3403, 4301, 3616, 4221, 1578, 2752, 1456, 1265, 687, 2083, 1534, 1300, 1390, 361, 3029, 3199, 1173, 629, 489, 707, 497, 4733, 879]. Exotic [3610]. expanders [4767]. Expansion [4245, 3672, 3449, 976, 2115, 137]. Expansions [3558, 3594, 4612, 2048, 2860, 2696]. expansive [1953]. Expansiveness [3891]. expectation [1029, 1350, 97, 2483]. Expected [3667, 1062]. experimental [31]. explained [825]. Explicit [205, 4660, 924, 4677, 4124, 3632, 3333, 4104, 2211, 3038]. explicite [2211]. explicites [1995]. Exponent [3272, 3929, 4448, 3389, 138, 4762].

Exponential [2448, 3168, 4179, 4494, 3870, 3215, 3050, 4171, 3986, 2909, 2070, 2201, 1807, 2127, 504, 639, 1260, 2134, 3036, 3778, 2210, 354, 1523, 941, 1316, 2790]. Exponentiality [3071]. Exponentials [4759, 4787]. Exponents [3245, 3955, 953, 1316]. export [4708]. export/import [4708]. exposed [264]. Expressed [3902]. Expression [4469, 176, 76]. expressions [205, 487]. Extendable [3094, 4290]. Extended [3285, 2087, 3648, 1877, 1732]. Extending [3227, 2558, 1393]. extensible [2092]. Extension [3484, 4585, 3703, 3231, 4614, 4109, 4264, 3905, 4352, 3888, 442, 1321, 1363, 3597, 1252, 525, 2458, 1669, 1367, 4002, 3819, 4587, 4548, 2873, 2836, 1378, 133, 349, 144, 426, 1291, 4669, 1026, 1948, 583, 683, 777]. Extensional [3214]. Extensions [3824, 381, 234, 3621, 3807, 3925, 1034, 3269, 3588, 3770, 3549, 3291, 573, 4335, 3128, 3711, 3998, 3509, 4314, 3359, 3583, 3175, 3694, 572, 4095, 1511, 4190, 3494, 3982, 4652, 4638, 120, 4077, 4302, 39, 1745, 1909, 1266, 2852, 96, 1341, 1642, 1826, 1557, 1397, 1988, 1692, 265, 2839, 2757, 748, 708, 1802, 2123, 2358, 1921, 385, 1077, 4786, 1432, 668, 1785, 2363, 912, 4719, 514, 657, 1240, 1591, 2157, 2768, 1096, 2544]. exterior [1938]. Exteriors [3710]. Extra [1993]. Extrapolation [3267, 1796]. Extremal [4614, 3068, 3548, 4298, 3978, 3165, 1945, 405, 1620, 56, 2642, 71, 188, 2305, 447]. extrémale [1620]. Extreme [4455, 953, 614, 2080, 2709, 3079, 1609, 892, 1283]. extremum [245].

F [360]. F. [4019]. faces [560]. Facial [2914]. facteur [2010]. Factor [4219, 618, 2511, 2010, 848, 3057]. factorable [642]. Factorial [3190, 2090, 2812, 2235]. factoriality [2889]. Factoring [4296]. Factorisation [4170, 181]. Factorizable [4150]. Factorization [2792, 4352, 3770, 4497, 2664, 3580, 4028, 3436, 4432, 132, 1086, 930, 1195, 1540, 1680, 324]. Factorizations [3266, 4394, 1351, 347, 3828, 2258, 1840]. Factorized [2023, 2424]. Factors [2456, 3535, 4407, 3859, 2000, 345, 1243, 1717, 2124, 185, 1153, 562, 3754, 3079, 1587, 2628, 2807, 624, 592]. Facts [4591]. failure [2954]. Faith [1373, 1957]. Faithful [4270, 2415]. Falbel [3713]. Falkner [3657, 707]. Faltings [4365, 4669]. familiar [296]. Families [3749, 3867, 3724, 4140, 4093, 3877, 3860, 3552, 2445, 3335, 1288, 1750, 2998, 886, 586, 4814, 1279, 1263, 3840, 297, 238, 2049]. famille [1798]. Family [3733, 3888, 4208, 895, 1798, 286, 1362, 310, 118]. Fan [4014, 1686]. Fano [4644, 3876]. Farthest [3358]. Farthest-Point [3358]. Fast [3963]. Fatigue [357]. Fatou [2271, 2539]. Faulkner [4779]. features [4765]. Fefferman [4433]. Fefferman-Stein [4433]. Fejér [1691, 119, 1159, 1423, 187, 2920]. Felix [883]. Fell [920]. Fermat [3322, 3749, 426, 3098, 3398, 2482, 2537, 4321, 1489]. fermion [105]. Fermionic [3285]. feuilletages [1410]. fiber [4678, 2819, 3820]. Fibered [3504]. Fibers [4157, 4288, 4480]. Fibonacci [1298, 1967, 1894]. fibrations [1814, 1595, 1503, 392]. fibre [2096]. fibred [2287]. fibres [2357]. Field [4013, 3287, 3708, 3112, 4465, 4612, 4528, 3334, 3314, 2174, 994, 1733, 700, 4756, 1962, 3752, 42, 1074, 2677, 1838, 2845, 4786, 2763, 1585, 583, 175, 1475, 511, 368, 856, 2044]. Fields [4243, 4510, 4245, 3500, 3925, 3245, 3588, 3326, 2755, 4575, 3287, 3097, 3098, 4311, 3477, 4169, 3595, 4653, 4502, 3217, 4001, 3542, 4382, 3982, 3544, 4489, 3944, 3847, 3008, 2772, 2855, 132, 3274, 2456, 2553, 2674, 2659, 3047, 1992, 4669, 4785, 2742, 859, 516, 2699, 4786, 1872, 2700, 2917, 2227, 4700, 610, 2488, 2747, 2157, 775, 2545, 1576]. fifth [1285]. Figueroa [2737]. Figures [3483]. Fillable [4424]. Filling [3483]. Fillings [3106, 3345]. Film [4419]. filter [2301]. Filters [4311, 415, 3031]. Filtration [3573, 3954, 4539, 4669]. filtrations [2734]. Fine [4390, 4780, 4647, 4277, 249]. fini [1715]. finitary [960, 2868]. Finite [4243, 4131, 4080, 2090, 201, 4510, 4683, 2994, 4245, 3452, 4437, 3106, 3142, 4438, 4385, 3903, 3245, 4013, 2175, 3929, 3689, 4575, 882, 4331, 3246, 3850, 1418, 665, 4085, 3098, 711, 3395, 2075, 3564, 4186, 4167, 4292, 4169, 4253, 2985, 3595, 3132, 2826, 1077, 3461, 225, 2764, 2884, 4528, 2628, 1998, 3334, 4216, 4153, 4196, 4656, 3495, 3944, 2157, 4174, 4302, 2089, 1715, 1134, 2736, 2388, 451, 344, 4744, 2195, 1376, 4796, 690, 404, 579, 994, 3073, 1793, 2573, 1733, 606, 1659, 235, 1717, 59, 1536, 1718, 1604, 409, 1605, 1633, 1936, 2861, 3063, 2187, 2558, 2723]. finite [2413, 2198, 2999, 2160, 4810, 3776, 4773, 2744, 765, 2463, 2145, 986, 1127, 1838, 1041, 516, 10, 2715, 2227, 695, 1502, 2918, 100, 2488, 1224, 1785, 266, 2807, 1492, 2074, 2827, 1281, 1263, 1730, 853, 818, 2992, 1393, 1822, 1545, 965, 1408, 2442, 878, 368, 856, 2260]. Finite-Dimensional [4085, 1077, 2884, 2388, 1717, 1536, 1604, 2861, 2744, 2918]. Finite-to-One [3495]. Finitely [4602, 4252, 3331, 3346, 1103, 1459, 2615, 2178, 2654, 2177, 4515, 1803, 1539, 2151, 1775, 2989, 853, 2209]. Finiteness [4365, 2675, 2141, 2369, 2811, 4732, 2663]. Finsler [4103, 4368, 4535, 3913, 4500, 3448, 3756, 4004, 4544, 1255, 4664, 4102, 4162]. First [1745, 3096, 3971, 2281, 3634, 1548, 4222, 923, 1187, 2798, 309, 4770, 4753, 1072, 2723, 1366, 3052]. Fitting [3511, 1041, 55]. Five [4208, 427, 1862, 134, 89, 331]. Five-dimensional [4208]. Fixed [4245, 3183, 1713, 2380, 3995, 4209, 2172, 2188, 3525, 4626, 1567, 4221, 1230, 4130, 1350, 1005, 628, 1377, 1534, 1301, 946, 2013, 1718, 1633, 980, 1025, 2818, 1037, 2286, 672, 439, 1832, 847, 997, 2295, 3077, 1754, 1704, 1796, 717, 1818, 1149, 489, 2830, 1547, 2610, 2903, 1467]. fixed-clique [2818]. fixed-points [2903]. Fixing [3917, 2759, 2885]. FK [2445]. FK-spaces [2445]. Flag [4105, 3950, 1997]. Flags [4372, 2088]. Flat [3987, 3409, 4121, 3913, 3698, 4242, 4300, 2408, 1498, 2925, 2595, 2149, 929, 2802, 4337, 195, 4733]. flats [1225]. Flip [1009]. Flip-flops [1009]. flops [1009]. Floral [3718]. Flow [4325, 3654, 3157, 3999, 4523, 4767, 2983, 2168, 2958]. Flows [4061, 4181, 3337, 3855, 3376, 1768, 1827, 589, 2848, 3056]. Floyd [1783]. Fluid [3609]. Fluid-Structure [3609]. Flynn [4772]. FM [2673]. FM-spaces [2673]. focal [2321]. Foias [2484]. fold [1458, 4465]. foliations [2088, 840, 1010, 1978, 2916, 1410, 2609]. Folkman [1631]. Følner [3626]. fonction [3259, 4756]. fonctionnel [2484]. fonctionnelle [1771]. fonctionnelles [3428]. Fonctions [2996, 1030, 2459, 2387, 635, 1987, 1356, 1235, 2712, 1628, 2079, 1899]. Fong [1573]. Footnote [470]. forced [2307, 3775, 575]. Forces [4620]. forests [2252, 4696]. Form [3412, 3290, 4655, 4293, 3565, 3911, 4429, 4524, 3496, 4793, 2519, 1558, 1624, 268, 244, 1307, 2784, 3158, 1964, 2785, 63, 472, 2680, 2132, 2117, 919, 1960, 1197, 2106, 18, 23, 82, 111, 3179, 6, 12, 20]. Formal [4058, 4013, 4440, 3381, 455, 2042, 819]. formally [1509]. formally-real [1509]. forme [1812]. formed [2356]. formellement [1509]. formelles [2042, 3399]. formes [3582]. Forms [3426, 3654, 3547, 3121, 3688, 3689, 4041, 4042, 3488, 4134, 3578, 4497, 3910, 4047, 4337, 3233, 3300, 3540, 3693, 3385, 4233, 3962, 4193, 4645, 3662, 3229, 3312, 1554, 3899, 3451, 4259, 3243, 2993, 173, 1801, 1849, 2779, 3033, 909, 2338, 2491, 3582, 2971, 2626, 698, 2082, 2582, 2117, 720, 1280, 3805, 110, 1286, 1979]. Formula [4243, 4660, 4309, 3411, 3971, 3415, 1345, 4313, 3611, 3225, 4059, 4592, 3435, 4104, 1603, 1606, 171, 2877, 2221, 2644, 591, 2898, 2346, 2845, 293, 153, 470, 912, 2106, 1599, 2355, 1776, 699]. Formulae [3949, 4124, 2520, 502, 1166, 150, 1507, 2269]. Formulas [4353, 3994, 4274, 3628, 13, 1886, 2712, 2762]. formule [2963, 699]. formules [2712]. Foster [154]. foundations [674, 2990]. Four [4443, 4372, 4559, 4211, 4129, 801, 1083, 2212, 2041, 1198]. Four-dimensional [4559]. Four-genus [4129]. Four-space [4443]. Four-uniform [4211]. Fourier [461, 3968, 1558, 503, 4020, 3889, 4134, 1100, 287, 375, 399, 3113, 1662, 4607, 71, 164, 4233, 462, 4525, 1970, 4592, 720, 66, 640, 3839, 1020, 1490, 1476, 3879, 2115, 543, 4008, 2591, 3743]. Fourth [3990, 3913, 4222]. Fourth-order [3990]. Fowler [1199]. Fox [1414, 2857, 1788]. FPF [1957, 2791]. FPF-rings [2791]. Fractal [3212, 2422, 2223]. Fraction [4245, 3449, 519]. Fractional [3842, 4226, 4355, 4395, 4449, 1371, 1887]. Fractions [4636, 760, 2927, 1619, 1537, 1714, 2662, 2131, 2393, 3176, 1749, 1446, 1295, 2479]. framed [981, 2255]. framées [2255]. Frames [3093, 4109, 4063, 3502, 3564, 4582, 2673, 4725, 2966]. framings [4709]. Frankel [3332]. Frattini [1694, 633, 990]. Fréchet [2320, 2873, 2571, 4650, 2183, 3388]. Fredholm [4181, 2958]. Free [3967, 3622, 4158, 4348, 3790, 2008, 4598, 4165, 2052, 2043, 2194, 3537, 2275, 960, 384, 4619, 3078, 3361, 4540, 4074, 1892, 4007, 3761, 3349, 4078, 1694, 2736, 2037, 2316, 1459, 4347, 972, 1416, 2178, 1181, 2891, 2474, 133, 1987, 4708, 2499, 2817, 2839, 4086, 763, 1880, 2081, 627, 562, 3189, 2726, 1144, 1781, 1703, 4470, 2918, 1225, 1631, 2989, 90, 797, 1424, 1850, 990, 2849, 626, 1022, 1769, 654]. free-by-finite [2736]. free-by-nilpotent [1459]. Freely [4348]. Freeness [4158, 1904]. French [1715, 2320, 4244, 3259, 1030, 2977, 1620, 2255, 2042, 1842, 1771, 1798, 2594, 1007, 2211, 2326, 1889, 2304, 2952, 2387, 2996, 1035, 1463, 1258, 3274, 3222, 2874, 3184, 371, 2266, 4756, 636, 635, 1121, 1987, 1509, 2484, 1746, 2896, 1356, 3428, 364, 3288, 2010, 4087, 2638, 2876, 2438, 129, 3251, 2600, 1171, 97, 1650, 3582, 3612, 1359, 996, 1346, 654, 1361, 961, 1691, 2712, 3020, 1089, 3004, 2679, 1658, 1628, 1836, 4819, 3857, 2972, 1859, 850, 1891, 3101, 2135, 2733, 3219, 3166, 49, 128, 452, 754, 2976, 1410, 699]. Frequency [3427, 4538]. frequently [4670]. Freud [587]. Frey [3982]. Freyd [3904]. Fricke [2247]. Friedberg [4236]. Friedrichs [1469]. friendship [1439, 1250]. Frieze [4284, 1110]. Frobeniu [3600]. Frobenius [4159, 393, 3610, 1195, 4665, 1151, 602, 3467, 686]. Front [4508, 4568, 4589, 4627, 4648, 4686, 4705, 4766]. Fuchsian [684, 4234, 2589]. Fuglede [2497]. Fuks [734]. Full [1811, 3551, 1425, 2951, 1789, 3753, 2198, 1205, 1587]. Fullness [1896]. fully [1307, 473, 1935]. Function [3469, 3651, 4036, 3968, 4012, 3703, 4551, 3245, 3501, 3096, 3111, 3704, 3328, 4269, 3708, 3097, 1684, 3127, 1705, 3672, 4067, 3358, 4091, 4124, 3958, 341, 3306, 3542, 3630, 3229, 4489, 4547, 4416, 1930, 800, 3259, 2388, 2448, 2430, 3008, 2427, 2669, 1941, 2797, 2551, 2855, 3011, 780, 467, 2552, 4756, 1398, 2532, 311, 659, 2659, 2366, 2230, 4044, 4532, 3752, 171, 176, 2317, 2158, 1743, 2912, 2313, 504, 623, 1404, 1260, 1443, 2027, 2335, 1925, 1542, 1723, 2210, 4236, 1348, 1252, 2454, 1864, 610, 1366, 1367, 81, 681, 466, 325, 326, 2702, 1386, 3053, 1490, 469, 564]. function [1050, 1512, 183, 1815, 1764, 2735, 2044]. function-derivative [1260]. Function-theoretic [1705]. Functional [3485, 4351, 4179, 4016, 4359, 1666, 1280, 4008, 328, 1649, 1912, 58, 675, 2504, 1771, 1032, 915, 2655, 2484, 1215, 2512, 408, 1607, 2124, 3030, 1923, 355, 601, 813, 1073, 1357, 1403, 200, 251, 2725, 532, 846, 2985, 169, 1646, 3783, 824, 544]. functional-differential [2512, 1607, 2725]. Functionals [4426, 3627, 3854, 3062, 528, 1399, 978, 3428, 2001, 1188, 1145, 4795, 464]. Functions [4456, 3574, 4283, 3138, 3366, 3484, 3621, 3557, 3665, 4384, 4477, 3452, 3105, 4650, 3969, 3745, 1855, 4614, 4327, 3547, 4461, 4642, 3277, 3487, 4065, 3247, 3476, 4269, 4290, 4555, 4533, 3295, 3656, 3224, 3892, 3797, 4143, 4356, 4169, 3643, 3100, 3300, 4122, 4254, 4395, 1332, 3373, 3175, 3584, 4214, 3894, 3660, 4594, 3961, 3163, 3714, 4194, 4630, 3144, 3480, 3145, 3311, 4450, 1523, 4301, 965, 3678, 3679, 3351, 3316, 4454, 3392, 3965, 839, 1394, 1952, 1060, 2546, 1030, 2430, 3059, 1241, 2872, 3824, 943, 1986, 2085, 2218, 2669, 2995, 1061, 1902, 4755, 2853, 2964, 3151, 250, 410, 944]. functions [994, 706, 2329, 2306, 2381, 2103, 4716, 857, 449, 2953, 779, 1462, 2996, 2059, 1702, 1099, 1603, 1800, 2938, 719, 4808, 3794, 635, 1987, 157, 1122, 441, 68, 265, 409, 2228, 13, 2489, 4374, 1356, 1922, 2070, 2201, 4673, 1013, 1160, 2366, 2450, 1291, 1441, 4753, 348, 2691, 812, 1402, 2113, 1470, 1391, 663, 30, 1817, 601, 2434, 463, 1442, 2409, 2221, 4272, 2616, 3796, 423, 1617, 2940, 3776, 1532, 3189, 2642, 2969, 2803, 3065, 1655, 1644, 3049, 2696, 4170, 1261, 874, 70, 4468, 119, 2931, 2712, 2762, 3036, 1542, 2943, 1861, 188, 1363, 2824, 546, 1731]. functions [569, 1043, 770, 167, 645, 916, 1628, 4814, 2727, 1046, 832, 2130, 4670, 1334, 796, 3083, 2729, 341, 2353, 2114, 1983, 2488, 2262, 4711, 773, 938, 988, 1521, 852, 228, 354, 464, 2421, 1666, 2919, 4721, 2605, 5, 1949, 4239, 2305, 55, 91, 4798, 3053, 3192, 743, 888, 2374, 2209, 1647, 2192, 1315, 3783, 314, 367, 397, 823, 1316, 652, 501, 2787, 1773, 2063, 2143, 1024, 2749, 731, 970, 238, 337, 401, 1231, 4702, 2903, 4690]. Functoriality [3573]. Functors [2160, 4357, 1066, 1383, 369, 1645, 1864]. Fundamental [4116, 3835, 4252, 3762, 2523, 2072, 1964, 369, 2680, 1960, 2018, 126, 85]. Furstenberg [2843, 2494]. Further [2931, 2327, 1826, 3757, 91]. Fusion [4136]. Fuzzy [2224, 3737].

G [2567]. G. [1805, 75, 1914, 587, 2567]. Gâteaux [3640, 3480]. gage [2951]. Galois [4624, 3182, 4679, 3440, 4463, 2119, 859, 4780, 3611, 4502, 2488, 4765, 3236, 2044]. Galois-theoretic [4765]. Galton [1121]. Galvin [4019]. game [37, 2462]. Games [4156]. gamma [1532, 435, 1386]. Ganea [3503]. Gap [502, 4602, 1259]. Gaps [3637, 1736, 3764, 4718, 1761]. Garsia [3479]. Gases [4025]. Gasket [3695]. Gassner [3211]. gathered [2765]. Gatti [4090]. gauge [3173]. Gauss [3531, 3209, 1978, 3355, 4443, 2182, 2416, 2390, 153]. Gaussian [1619, 1537, 2694, 1204, 4002, 3530]. Gel’fand [2033, 4656]. genera [2119]. General [3273, 3908, 3342, 3939, 1395, 4691, 272, 2385, 1905, 294, 166, 1217, 176, 2998, 2221, 1111, 1635, 1958, 756, 263, 2917, 3801, 2806, 542, 43, 3192, 1197, 1021, 446]. généralisation [1691, 2438]. généralisé [2679]. Generalised [3533, 3958, 2536, 2891, 2909, 462, 1477, 501]. généralisée [1359]. généralisés [1007]. generality [4712]. Generalization [3811, 171, 2438, 4066, 3434, 3433, 2309, 4815, 83, 2057, 1377, 199, 37, 298, 2655, 603, 115, 1783, 843, 1037, 552, 2956, 1711, 219, 1697, 1691, 655, 875, 2165, 2482, 2354, 1444, 153, 277, 942, 1264, 154]. Generalizations [3266, 4576, 3600, 1735, 2037, 1724, 295, 408, 1525]. Generalized [1657, 3321, 567, 922, 3826, 3485, 2753, 3827, 3270, 4014, 4421, 3232, 2855, 4307, 4204, 4286, 4309, 3907, 4405, 4268, 4355, 3143, 3608, 4406, 3954, 1205, 3382, 4467, 2971, 4482, 3727, 3131, 1753, 4412, 4791, 4190, 4278, 3837, 4525, 3418, 2537, 2680, 3695, 3614, 4299, 2061, 4361, 4238, 2202, 3678, 4415, 4646, 855, 4104, 1578, 1694, 2736, 1007, 1325, 2856, 719, 2554, 1326, 3833, 9, 15, 685, 1141, 2912, 62, 1837, 4337, 2726, 660, 1236, 385, 3200, 2009, 794, 435, 1333, 998, 1731, 2679, 645, 2007, 2191, 1046, 3759, 692, 1148, 939, 680, 2441, 2584, 2848, 1991, 564, 3039, 990]. generalized [2849, 913, 1199, 1359]. Generalizing [4505]. Generate [4015, 90]. Generated [3326, 3246, 3751, 3111, 4515, 4252, 3331, 4194, 3346, 4078, 1103, 1436, 2415, 1459, 2615, 2178, 1181, 471, 2654, 1962, 871, 1101, 1803, 2145, 2151, 2731, 1775, 2989, 1701, 2630]. Generating [3206, 3904, 193, 348, 4524, 3117, 4261, 800, 1470, 4713, 325]. Generator [3363]. Generator-Interchanging [3363]. Generators [4109, 2750, 841, 842, 884, 2900, 2374, 139, 3055]. Generic [4203, 3094, 2278, 3716, 1850]. Genericity [3043]. genuine [511]. Genus [3275, 3290, 4392, 3114, 4130, 4198, 2499, 927, 2265, 2899, 4791, 4129, 85]. Genus-2 [4130]. Genus-Zero [3275]. Geodesic [4252, 1294, 589]. Geodesics [3470, 3409, 3864, 4103, 4102, 2237, 78]. Geography [4680]. Geometric [3718, 3408, 4161, 1621, 4264, 4180, 4333, 4336, 3431, 3158, 3579, 3401, 3284, 2787, 3278, 2397, 2889, 2925, 1976, 3000, 4725, 62, 245, 1946, 1506]. Geometrical [3713, 104]. geometries [4715, 1963, 1306, 151]. Geometry [3516, 4368, 3454, 4042, 3283, 3296, 4577, 4467, 4004, 3514, 1080, 2235, 559, 1192, 80, 124, 1357, 112, 152, 2375, 271, 4794, 266, 126, 2684, 79]. germ [2135]. germe [2135]. Germs [2112]. getting [880]. Ghosts [4326]. Gibbs [4379, 1407, 512]. Gioia [470]. Girth [1865, 64]. Giuga [3572]. Given [3326, 4138, 4210, 3544, 1208, 2017, 2664, 1924, 805, 2957, 2117, 666, 774, 3053, 1527]. giving [1343]. Gleason [1801, 1925]. Glimm [4089]. Global [4456, 2794, 309, 4226, 2557, 3592, 4167, 4430, 3713, 4489, 1578, 3008, 4802, 2703, 1390, 1245, 1559, 361, 2939, 3775]. globale [2556]. Globalization [3292]. Globally [4381]. Goebel [1377]. Going [3500, 3286]. Going-Down [3500]. Goldbach [59]. Goldberg [551, 4309, 148]. Goldie [1461, 2856, 1216, 2187, 1886, 2348, 643]. Goldie-like [1461]. gons [1994, 163]. good [2036]. Gordan [2247]. Gordon [2147]. Gorenstein [3107, 4678, 3028, 4256, 3162, 4300, 4417]. Gorenstein-projective [4256]. Goss [2024]. Gottlieb [2477]. Gould [1738]. Goursat [229]. governed [1930]. gra [1625]. Grade [3302]. Graded [3257, 3265, 4287, 3929, 3566, 3915, 3663, 2548, 2907, 3028, 2558]. Gradient [4325, 4391, 2832, 2618, 2848]. Gradings [3263, 3638, 3922]. Graham [1581, 3412, 647]. Gram [3515, 142]. Grammar [2406]. Graph [4034, 4262, 4224, 4572, 3670, 4331, 3606, 4270, 3710, 4373, 4144, 3714, 4127, 4487, 2427, 208, 533, 536, 333, 1269, 2206, 33, 236, 898, 1625, 1247, 510, 1250, 4713, 64, 2899, 222, 305, 1253, 774, 2234, 47, 172, 3784]. Graphs [4344, 4345, 4346, 4651, 3268, 4014, 4328, 4575, 3850, 4086, 928, 4335, 752, 4186, 3671, 4205, 4210, 3912, 3511, 4357, 4212, 3727, 4277, 4029, 4101, 4238, 4322, 4323, 4362, 4363, 4303, 1232, 1208, 4637, 581, 1550, 1656, 1902, 2252, 3370, 561, 728, 971, 313, 333, 1690, 1858, 1163, 1009, 1439, 955, 1610, 136, 165, 240, 1625, 1193, 1400, 2818, 744, 1087, 927, 751, 147, 291, 430, 1942, 496, 281, 495, 1562, 765, 1487, 1782, 1723, 243, 900, 2003, 1048, 319, 820, 101, 2143, 1185]. Gras [4001]. Grassmann [3705, 2541, 2016]. Grassmannians [4168, 4464, 4257, 4329, 3512]. gravitational [357, 175]. Green [4367, 157, 2228, 2335, 2390]. Greenleaf [771]. Greiner [3418, 3759]. Griffiths [3289]. Grigorchuk [4788]. Gromov [3280, 2822, 3693]. Gross [4025]. Grothendieck [4143, 3362, 4215, 1063]. Ground [4448, 4241, 4762]. Group [4155, 4080, 3263, 3321, 3638, 3181, 3808, 3153, 4386, 3324, 4350, 4603, 3845, 4642, 4423, 3520, 4266, 3155, 3549, 3931, 3578, 4137, 4022, 3381, 4116, 3564, 3415, 3445, 930, 4118, 4408, 3739, 4187, 4143, 479, 1275, 4639, 4666, 4299, 3308, 3346, 3333, 3552, 3647, 4007, 3349, 3762, 3240, 2523, 2379, 1134, 2257, 1004, 2871, 2174, 1208, 106, 451, 2316, 2633, 2527, 3485, 1920, 649, 1081, 1552, 778, 972, 2178, 1414, 2838, 2220, 1659, 2907, 2474, 2773, 438, 1085, 2817, 2895, 2779, 1140, 1141, 366, 691, 2052, 2043, 2194, 2480, 1744, 1752, 1721, 2187, 2558, 2967, 2160, 2229]. group [984, 2426, 4699, 2710, 2783, 2842, 62, 684, 2878, 1040, 2984, 859, 4685, 2250, 554, 2028, 2145, 2625, 1943, 571, 805, 112, 768, 1127, 4713, 650, 1703, 1041, 633, 63, 1860, 2826, 2383, 756, 2433, 2535, 3082, 3857, 917, 1432, 538, 1091, 1365, 3085, 1019, 2847, 90, 2974, 1871, 273, 1388, 1095, 139, 3070, 1932, 2490, 1614, 2044, 3260]. Group-Frames [3564]. group-graded [2558]. group-valued [1095]. Groupe [2297, 3857, 2733]. groupes [1798, 2444, 371, 364, 2010, 1950, 3612, 961]. Groupoid [4054, 1616]. groupoids [1057]. Groups [4345, 4131, 4383, 3393, 4624, 4037, 4611, 4061, 4284, 3532, 4436, 4158, 4108, 3788, 3559, 4348, 4285, 3245, 2356, 4460, 3904, 3842, 4327, 4404, 4306, 3272, 3441, 4308, 3928, 4598, 3354, 3533, 4228, 4248, 4112, 3282, 3246, 4353, 2177, 4590, 4354, 4137, 3289, 2113, 3890, 4045, 4047, 4186, 4118, 3626, 3397, 4444, 4141, 3935, 4465, 3628, 4252, 4253, 4295, 3998, 4647, 3329, 4275, 4121, 4619, 3511, 4232, 3583, 4629, 3331, 4125, 4412, 4095, 3235, 4320, 4278, 3645, 3216, 4027, 4413, 3346, 3599, 3528, 3941, 4432, 3228, 4239, 4128, 4153, 4196, 3389, 4031, 3364, 4656, 303, 4056, 3351, 3985]. Groups [4078, 4174, 4198, 4452, 3617, 4529, 2568, 2257, 2736, 2023, 2424, 2448, 4777, 3824, 451, 2316, 2737, 2994, 1210, 1798, 2060, 2002, 285, 345, 1191, 2195, 4796, 1459, 4825, 4679, 1482, 3026, 2352, 4693, 972, 1416, 3073, 1298, 2506, 2573, 1181, 2688, 235, 4597, 2857, 2008, 3074, 1465, 371, 2773, 2858, 1440, 1536, 2775, 1779, 2013, 2499, 2531, 2508, 2166, 1430, 4788, 4673, 364, 192, 921, 1825, 2010, 2781, 2052, 2955, 2153, 2533, 2389, 1484, 763, 982, 3032, 627, 2503, 2275, 2472, 2783, 3835, 1972, 1677, 2663, 2880, 4120, 1866, 4780, 4782, 385, 554, 3612, 1221, 2226]. groups [961, 2009, 2237, 3078, 986, 1075, 1781, 1262, 2325, 2501, 2945, 2826, 1310, 2522, 4754, 2433, 832, 2728, 1574, 864, 2476, 41, 3071, 3161, 2493, 1136, 1511, 2285, 2883, 1674, 1696, 3037, 1785, 2021, 1904, 4765, 2847, 1176, 624, 797, 2074, 2827, 425, 3818, 1730, 2368, 4746, 1118, 2399, 1475, 1063, 1227, 52, 3086, 1425, 2181, 514, 657, 2849, 626, 967, 1408, 2589, 2590, 2442, 2868, 2260, 2361, 3055, 2324, 2544, 3056, 3204, 3058, 2089]. Growth [4614, 4252, 4647, 3226, 4470, 4670, 4452, 3259, 4825, 2059, 1190, 2933, 4805, 3055]. Grünbaum [897, 2315, 2049]. GTW [4257, 4329]. Guessing [3674]. guide [1426]. Gunning [3290]. Gupta [4788].

H [1788, 3835, 3781]. H-sets [3781]. H-type [3835]. H. [697, 19]. Haar [3859, 4221]. Hadamard [4325, 3531, 2739, 4797, 2036, 26]. Hahn [683, 777]. Hajnal [190, 278, 4585]. Half [4010, 3190, 4049, 2462, 3860, 2812, 1359, 4795, 4724, 3839, 31]. half-baked [31]. Half-Factorial [3190, 2812]. Half-Line [3860, 4724]. half-normed [1359]. Half-Plane [4010]. Half-silvered [2462]. half-space [3839]. Halin [3727]. Hall [1222, 4]. Halmos [3051]. Hamburger [792]. Hamel [3714]. Hamilton [263, 175, 174]. Hamilton-Jacobi [175, 174]. Hamiltonian [365, 1656, 904, 1163, 1009, 1400, 291, 4615, 3727, 588, 3347, 1184, 4241, 47, 1185]. Hamiltonians [1698]. Hammerstein [758, 3655]. Hankel [2738, 3452, 3666, 2508, 2760, 660, 2562, 1732, 4731]. Harary [1232]. Hard [4275]. Hardy [4010, 3425, 1602, 1642, 1826, 3735, 3887, 3932, 1500, 3430, 3339, 2743, 2470, 4394, 3759, 1453, 293, 3838, 3422, 821, 2051, 4595, 4451, 1764, 2447, 2810]. Hardy-Type [3430]. Harish [3614]. Harish-Chandra [3614]. Harmonic [2721, 4290, 4356, 3330, 3567, 3961, 3376, 4257, 3204, 3243, 4454, 2059, 2997, 262, 2128, 2656, 2409, 2983, 2940, 2642, 2969, 269, 2082, 1312, 1983, 2305, 699]. Harmonicity [3685]. harmonics [1744]. harmonique [699]. Hartman [1047]. Hartog [3692]. Hartogs [3491]. Hasimoto [3198]. Hasse [2119, 3350]. Hauptmoduls [3588]. HAUS [1367]. Hausdorff [2584, 2218, 2669, 1577, 96, 825, 4689, 3549, 1072, 2624, 1169, 2879, 2822, 499, 342, 1322, 607, 2441, 3918, 4581, 2542, 2830, 421, 543, 1900, 2270]. Having [3751, 3711, 3633, 3550, 4824, 250, 1328, 2578, 2820, 2296, 3754, 859, 2700, 2847, 666]. Heat [3109, 3585, 2808]. Hecke [3104, 4047, 3233, 2114, 4342]. Hedgehogs [4234]. Height [3708, 3381, 3511, 2154, 1997]. Heights [3367, 3403, 4821]. Heilbronn [3812]. Heinz [186]. Heisenberg [3369, 4460, 4642, 4266, 1744, 3415, 3835, 3240, 4078]. Helices [3198, 3876]. Helicoidal [4162]. helix [2126]. Helly [1553, 1460]. Helly-type [1553]. Helmholtz [3738]. Hemivariational [3590]. Henselian [493, 505]. Henstock [3450, 799]. Hereditarily [4145, 2707]. Hereditary [4145, 2249, 1610, 2893, 1167, 1627]. Heredity [3995]. Hermite [1060, 4797, 2292, 4503, 1159, 1423, 187, 2920]. Hermitian [3991, 4777, 4202, 3685, 488, 2129, 3385, 3567, 1128, 3161, 1174, 4003, 4507]. Hermitian-Einstein [3991]. Hernández [3757]. Herstein [1140, 1661]. Herzog [2195]. Hesse [73]. Hessian [4802]. Heuristics [4243]. hexagonal [2481]. High [3427, 4623, 3882, 1210, 2287, 4791]. High-Degree [4623]. Higher [3557, 4436, 3605, 3951, 4552, 2656, 3459, 4233, 919, 4435, 3862, 1454, 2064, 4825, 4793, 1890, 525, 3802, 2115, 2949]. Higher-Dimensional [3605, 4552, 2656]. higher-order [4793]. Highest [4051]. highly [3192]. Hilbert [1455, 2668, 3180, 4821, 4683, 4744, 2574, 3519, 295, 4423, 2452, 4308, 1870, 3280, 3505, 4333, 3287, 789, 4826, 931, 4667, 2985, 2188, 2250, 4483, 4446, 65, 2460, 2314, 2865, 4028, 2605, 2254, 1568, 3241, 4816, 1349]. Hilbertianity [4580]. Hill [3011]. Hilton [552, 378]. Hincin [3622]. Hirota [3775]. Hirschfeld [306]. Hitting [1654]. Hjelmslev [1556, 1965, 2062, 1570]. HNN [3128]. HNN-Extensions [3128]. Hochschild [2840, 2472]. Hodge [4304, 4669, 3432, 2734, 4507]. Hölder [1947, 3087, 4438, 4551, 2982, 1508, 1591, 2868, 4414]. holomorphe [2556]. Holomorphic [4456, 3484, 4650, 3685, 3502, 4290, 4371, 3973, 3871, 3693, 4214, 3348, 4544, 3861, 3700, 3183, 2557, 4753, 3776, 2200, 2188, 3083, 4711, 4675]. Holomorphicity [3692]. Holomorphy [3953, 3130]. Holonomic [4607]. hom [1645, 4522]. Hom-bialgebras [4522]. homeomorphic [3810]. homeomorphism [1118]. Homéomorphismes [2874]. Homeomorphisms [3532, 3924, 4107, 3885, 3791, 1768, 3188, 24]. Homogeneity [3156, 3357, 4213, 3789, 4076]. Homogeneous [4155, 4459, 3846, 4425, 3442, 4353, 2208, 4418, 4635, 4535, 4536, 4119, 4255, 3937, 4629, 4600, 4609, 1296, 356, 1289, 701, 787, 528, 2302, 4338, 2200, 2696, 4622, 2944, 485, 2747, 414, 3965]. Homogenization [4312]. Homological [4038, 4249, 4354, 4277, 4215, 4488, 386, 229, 230, 1780]. homologies [2128, 269]. Homology [3230, 4633, 4704, 4276, 4321, 579, 3044, 2573, 4554, 2999, 4480, 1433, 4488, 969]. Homomorphic [3940, 61]. Homomorphisms [1533, 4501, 4358, 2883, 2874, 2637, 2984, 478]. Homothetic [3725, 2980, 1257]. homotopical [2477]. Homotopy [3120, 3532, 3825, 3522, 3578, 3628, 3649, 579, 3670, 4808, 2622, 4749, 456, 4120, 2066, 2009, 2826, 2717, 378, 232]. homotopy-commutativity [456]. Hook [3949]. Hook-content [3949]. Hopf [3264, 2065, 2834, 2326, 4110, 2420, 1469, 4444, 2725, 2711, 781, 3177, 3418, 2395, 4505, 4543, 4740, 273, 3070]. Hopf-type [4543]. Hopficity [1459]. Horizontal [3878]. horn [3154]. horn-like [3154]. Hospital [2038]. Howe [3551]. Howson [3001]. HRT [4642]. Huber [3690]. Hughes [2944, 2766]. hull [3787, 475, 677, 1292, 417, 715]. Hulls [3807, 320, 704, 2913, 1309, 213]. Huppert [2493]. Hurewicz [2587]. Hurwitz [3469, 4161, 4388, 2501, 4237, 2735]. Hyper [4483]. Hyper-Hilbert [4483]. hyperadjugates [219]. Hyperbolic [4263, 4663, 4825, 4285, 3690, 1514, 4576, 3609, 3935, 4393, 4481, 3234, 4452, 2981, 4813, 2721, 1110, 1980, 2987, 3055, 4733]. hypercentral [2897]. Hypercentre [4153]. Hypercovering [3934]. hypercubes [134]. Hypercyclic [4037, 4670]. Hypercyclicity [4404]. hypercylinder [4767]. Hyperelliptic [3230, 3184]. hyperelliptiques [3184]. Hyperfinite [4000, 3859, 1243, 2511]. Hypergeometric [4274, 3211, 3958, 501, 1507]. Hypergraph [3244, 4347]. Hypergraphs [4439, 4211, 1610, 1589]. Hypergroup [3797, 3796]. Hypergroups [4378, 4221]. Hyperplane [3744, 4157, 2805]. Hyperplanes [3621, 4225, 3689, 3496, 1821]. Hypersurface [3484, 3482, 3766, 2734]. Hypersurfaces [2255, 4039, 3654, 4574, 1025, 4168, 4464, 4337, 3565, 3911, 4524, 4431, 3589, 3866, 4018, 4257, 4329, 4397, 3512, 4586, 4565, 4474, 4646, 2408, 2993, 4720, 2658, 981, 2694, 2784, 3158, 1966, 2680, 2135, 2159, 4733]. hypo [1009]. hypo-Hamiltonian [1009]. hypoellipticity [2794]. Hypothesis [3904, 4469, 4613].

i.i.d [4747]. Icosahedral [3310]. Ideal [4285, 4409, 4033, 4344, 1755, 1971, 1671, 2593, 2311, 3831, 1418, 2144, 2384, 3064, 2580, 2880, 255, 1276, 2336, 3817, 729, 4322, 2767, 2348]. ideal-related [3831]. ideal-triangularizability [3064]. Ideals [4511, 4286, 3907, 3185, 4133, 3627, 3132, 4214, 4639, 3117, 3663, 3481, 4032, 3317, 1755, 1933, 4821, 4775, 2429, 2149, 1746, 2146, 2046, 708, 290, 322, 830, 860, 2516, 1847, 2625, 1751, 2678, 1999, 2730, 2731, 2764, 1728, 2630, 2702, 1910, 725, 1739, 1425, 754, 4783, 1023]. Ideaux [754, 1746]. Idempotent [3889, 1847, 4540, 3479, 2926, 1940, 862]. Idempotents [1616, 3853, 778, 2360, 1237, 1256]. Identité [1356, 1658]. Identities [1466, 4405, 947, 397, 2414, 895, 2692, 2930, 2492, 998, 795, 3040]. Identity [4458, 3848, 3397, 1795, 1655, 793, 204, 1726, 1525, 2068, 1356, 1658]. if [195]. II [613, 4302, 4456, 4680, 4175, 2568, 4605, 3702, 1664, 3826, 4349, 50, 1845, 193, 2452, 1604, 3327, 15, 4135, 3704, 573, 1685, 3490, 2929, 3157, 4378, 781, 3914, 3066, 4318, 986, 3019, 191, 1854, 51, 3307, 4700, 1587, 3714, 228, 2117, 2140, 21, 2992, 1998, 55, 1544, 622, 1512, 823, 854, 966, 3239, 2446, 2348, 2263, 3317, 1231, 6, 12, 4697]. III [3023, 1243, 617, 898, 3850, 3361, 370, 159, 22, 2681, 91, 6, 20]. ill [1844]. ill-posed [1844]. Image [2470, 4499, 3597]. Images [4444, 3940, 3236, 1053, 1522, 3315, 2240, 2707, 1924, 2040, 1545, 2809, 1688, 1846]. Imaginary [2780, 4001, 1962, 2553, 2803, 3805]. Imbedding [4587, 461, 839]. immanant [2277]. immersion [2148]. Immersions [3426, 1839, 1964, 2785]. imperfect [4669]. implicational [1101]. Implications [4458]. implicative [2039]. implicit [659, 2725, 2711]. Implies [4599, 3930, 3405, 1114, 2806]. Imply [3673]. import [4708]. impossible [521]. imprimitivity [2057]. Improved [3970, 2014]. Improvements [4142]. Impulsive [3923, 3783]. Inadmissibility [757]. INC-extensions [1692]. incidence [3085]. Inclusion [4036, 1982, 1111, 3465, 2436, 1905, 2118]. inclusions [4782]. Incomplete [4289, 596, 2261, 224]. Incompressibility [4338]. Indecomposability [4046]. Indecomposable [3459, 2491, 3730, 4216, 2339, 4055, 696, 1307, 2203, 321, 651]. Indefinite [157, 173]. Independence [4573, 3623, 1306, 2604, 1865, 1337, 1055]. Independent [1799, 4792, 716, 2075, 435, 2446]. indeterminates [2042]. Index [4034, 3022, 3088, 3147, 3205, 3256, 3258, 3378, 3438, 3498, 3556, 3618, 3680, 3682, 4138, 4292, 4523, 3461, 3544, 4362, 2144, 3015, 2074, 1523, 3821, 18, 82]. indexed [1864]. indicable [2688]. Indicators [4151]. Indicatrices [3296, 1255]. individual [1704, 1908]. Induced [3885, 4056, 3947, 2252, 2967, 4782, 2765]. inducing [2021]. induction [2527]. Inductive [1449, 3520, 3906, 3443, 4501, 3374, 3238, 3363]. inequalites [1500]. Inequalities [3636, 1455, 284, 4082, 3073, 4479, 4227, 3590, 3887, 2070, 4726, 3549, 3908, 4655, 3212, 3127, 3936, 4277, 645, 3938, 3961, 3962, 4193, 3838, 4433, 3731, 4223, 2319, 1579, 1003, 4802, 2923, 4744, 1006, 4797, 1602, 1724, 1621, 4793, 4710, 2875, 1214, 3772, 1883, 1817, 1442, 555, 3099, 1890, 483, 1697, 3778, 317, 4795, 2991, 3759, 2933, 228, 821, 2374, 2086, 1591, 4735, 1412]. inequalitiy [156]. Inequality [3517, 4585, 3267, 4179, 3537, 4316, 4622, 3524, 4594, 3144, 4281, 3530, 3966, 1912, 4803, 4722, 4797, 2837, 186, 48, 1642, 1826, 2982, 2036, 4671, 1139, 662, 2345, 2956, 1617, 1532, 2743, 560, 377, 916, 1076, 1128, 2277, 108, 1017, 1018, 1667, 2349, 2262, 1386, 1197, 92, 1791, 1508, 1591, 403, 4079, 3396, 3191, 3603]. inertia [2129]. Infection [4620]. Infeld [838]. inférieure [3857]. inférieurs [97]. Infimum [3897]. Infinite [3732, 3574, 3193, 3268, 3623, 3272, 3535, 4113, 4182, 3287, 4534, 4140, 116, 3995, 4647, 4484, 4026, 3601, 1506, 3771, 2929, 646, 2264, 114, 145, 4757, 2107, 2826, 2535, 2827, 425, 376, 1424, 3783, 878, 2089, 2042]. Infinite-dimensional [4534]. Infinitely [4247, 3751, 3945, 217, 1099].

Infinitesimal [1756, 4522, 4580]. Infinity [3641, 3286, 2042, 2451]. inflated [1239]. inflection [2931]. inflexions [1203]. influence [1134]. informal [113, 191, 370]. information [1676, 757, 2765]. Infranil [3816]. Infranil-Orbifold [3816]. Infrasequential [1673]. Ingham [1638]. Inherently [2823]. Inhomogeneous [4250, 4053]. Initial [4441, 2443, 934, 2147]. initial-boundary [934]. injections [2076]. Injective [4674, 704, 3792, 2365, 3953, 4117, 4206, 3464, 2204, 1068, 320, 1456, 1265, 416, 995, 1082, 475, 2911, 1329, 2913, 384, 214, 246, 677, 860, 195, 1309, 1129, 213, 417, 1494, 834, 1131, 261, 321, 651, 35, 715]. Injectively [4249]. injectives [995, 1907, 1769]. Injectivity [3443, 3592, 4406, 4501, 1335, 769, 277]. Inner [3665, 4477, 4011, 3863, 3535, 3476, 3644, 1334, 2729, 2954, 2655, 4754, 4721]. input [1564]. input-output [1564]. Inseparable [3592, 583]. Insolvability [161]. Instabilité [1509, 3101]. instability [272, 1509, 3101]. Integer [3625, 3576, 3281, 4041, 3994, 3384, 4610, 3117, 823, 2891, 2179, 2928, 562, 16, 472, 1871, 867, 2567, 2269]. Integer-Valued [3117, 823]. Integers [4531, 3490, 4818, 4002, 775, 2174, 2852, 2225, 1438, 1581, 882, 957, 4748, 1880, 2081, 2598, 3754, 27, 733, 336, 1841, 4728, 901, 1512, 2464, 1286, 2479]. integrability [2487, 2136]. integrable [2094]. Integral [3499, 4080, 3023, 451, 4224, 4159, 3842, 4178, 4461, 4425, 978, 1853, 3655, 3411, 3210, 4355, 4635, 2229, 4655, 3995, 3342, 4049, 1224, 4472, 4449, 1863, 3362, 3554, 3316, 3550, 3076, 2257, 2050, 976, 2951, 1664, 2102, 747, 721, 400, 1354, 1898, 636, 761, 34, 2532, 2401, 2778, 3833, 2656, 3047, 3030, 2480, 2391, 1486, 2912, 748, 708, 984, 1500, 2783, 2842, 2878, 741, 4742, 1705, 831, 910, 2496, 2173, 143, 937, 852, 2363, 2828, 1998, 821, 4723, 460, 799, 1505, 1477, 1614, 2143, 1027]. intégrale [636]. Integrality [3811, 2525]. Integrally [2091, 2328, 2507]. Integrals [3379, 407, 3143, 3211, 3253, 4070, 4485, 3450, 4416, 1266, 2103, 4684, 1982, 2875, 3065, 1831, 1887, 2735, 970]. integrand [1978]. Integration [3900, 3212, 4547, 337, 1231, 1061, 2102, 2196, 1330, 1644, 731]. Integrations [4235]. Integro [4195, 4381]. Integro-Differential [4381]. Integro-dynamic [4195]. Interaction [3609, 4312]. interactions [2097]. Interchange [3988]. Interchanging [3363]. interesting [780]. interfaces [2631]. interior [1517]. interior-point [1517]. interlacing [4800]. Intermediate [3538, 2986]. Internal [1645, 2069, 1448]. interpolating [1916, 1360]. Interpolation [3745, 3748, 3325, 1391, 4147, 3372, 3568, 1312, 4623, 2518, 1939, 1463, 2874, 2048, 1235, 726, 1126, 2353, 850, 3817, 376, 1159, 1423, 187, 2920, 443, 55]. interpolatory [1565, 444]. Interpretation [4336, 3308, 519]. interpretations [2492]. Intersection [4401, 4133, 2243, 1491, 2178, 523, 3054, 2291, 1257]. Intersections [3741, 2090, 625, 2169, 4682, 2605, 224, 1791]. intertwining [3428, 1540]. Interval [3641, 3823, 4347, 2267, 2162, 2624, 376, 1937]. intervals [2430, 250, 2045, 1973, 878]. intricate [2258]. introduced [249]. Introduction [80, 3225, 7, 21, 22, 124]. Intuitionistic [3737]. Invariance [3701, 923, 3648, 2921, 2643, 650]. Invariant [3139, 3788, 3559, 2418, 3291, 3627, 3157, 3127, 4378, 3307, 3387, 3677, 3553, 4586, 4129, 3985, 1060, 1593, 1318, 1719, 1787, 2267, 2797, 2657, 4043, 2860, 2622, 2246, 981, 185, 2927, 2624, 2745, 4822, 2243, 2932, 1794, 648, 1422, 818, 4003, 2405, 714, 1910, 1227, 2959, 1338]. Invariants [3107, 3751, 4370, 3740, 3693, 3384, 4320, 3103, 581, 2925, 3073, 2111, 2506, 2529, 4784, 4714, 2755, 1962, 2153, 3338, 2119, 2968, 2290, 1145, 2987, 3201, 3177, 2934, 2959, 111, 3072]. Inverse [3295, 3609, 4358, 4150, 4054, 1002, 2851, 2752, 2183, 4767, 1709, 1363, 2679, 695, 2567, 926]. Inverses [4369, 3093, 3248, 3583, 4415, 1325, 768, 1508, 1591]. Inversion [719, 3415, 3418, 1318, 351, 43, 66, 153]. inversions [800]. inversive [2601]. Invertibility [3341]. Invertible [1480, 2510, 3836, 2311, 2149, 2865, 2767]. invertibles [2892]. Involution [3638, 4445, 3136, 3241, 3318, 1783, 2273]. Involutions [3246, 3709, 4022, 3890, 2759, 3917, 755, 1177]. involve [749]. Involving [4604, 3955, 4316, 4656, 4417, 467, 1969, 2173, 720, 2305]. Iosifescu [177]. Irrational [1474, 3494, 3238, 3363, 2687]. irrationalities [1714]. Irrationals [3380, 2528]. Irreducibility [3734, 1151, 368, 1715]. Irreducible [4200, 3863, 4177, 4528, 3898, 2513, 870, 1011, 2559, 518, 804, 1112, 1362, 756, 125, 1629, 1117, 1840, 853, 1613, 592]. irreducibles [2820]. Irréductibilité [1715]. Irregular [4499, 439, 1040]. Irregularities [4632]. isochrones [3237]. Isolated [3146, 2734, 2831]. Isometric [3521, 2844, 4097, 2040, 3242, 3803]. Isometries [4202, 3909, 3892, 2867, 2419, 4715, 3787, 1484, 663, 1756, 1334]. Isometry [4375, 3891]. Isomorphic [2379, 4650, 1721, 1365, 345, 3196, 2337, 2863, 853]. Isomorphism [3393, 2740, 4258, 243, 2769]. Isomorphisms [4555, 2333, 885, 1542, 2348, 1033, 4776, 1084, 2466, 4717, 1146, 1651]. isoparametric [3158]. Isoperimetric [4277, 560, 764, 377, 1791]. Isospectral [4121, 3816, 2653]. isotopy [2606]. Isotropic [3426, 4442, 4372, 1964, 2785, 4664]. isotype [2741]. isotypic [2858]. Issue [4508, 4509, 4568, 4569, 4589, 4627, 4628, 4648, 4686, 4687, 4705, 4706, 4766]. Iterated [3941, 4679, 2797, 2656, 1891, 252]. iterates [2767]. iteration [308, 215, 1089, 29, 1089]. Iterations [3655, 1434, 2614, 847]. Iterative [4410, 1034, 962]. itérés [1891]. ITPFI [1717]. IV [215, 160, 22]. Ivory [355, 1357]. Iwasawa [2344].

J [577, 803]. J. [2578, 230, 782, 850, 1019, 587, 1547]. Jackson [2421]. Jacobi [2567, 2410, 4458, 3654, 3279, 3173, 4168, 2559, 219, 3233, 3540, 263, 3176, 3895, 4321, 3589, 3866, 4018, 325, 4586, 175, 174]. Jacobi-Like [3540, 3233]. Jacobian [3080, 3749, 3982]. Jacobians [3903, 3353, 3098]. Jacobson [4674, 563, 1382, 1290, 4088, 4558, 4148]. Jacquet [3967]. Jain [1914]. James [4577, 3225, 4411, 3177]. Jarden [3982]. Jensen [1738, 1132]. Jeśmanowicz’ [4135]. John [4622]. Johnson [1630]. Joins [661]. Joint [3317, 2924]. Jones [4614, 2585]. Jordan [1396, 1461, 4521, 4204, 4085, 4427, 4293, 4090, 3020, 2283, 3303, 919, 653, 3602, 2868]. Jordan-Chevalley [4521]. Joubert [4503]. Journé [4425]. Juhász [4585]. Julia [1441, 329].

K. [148]. K3 [4538, 4392]. Kac [3434, 2753]. Kadec [2938, 3615]. Kahane [3191]. Kähler [4403, 622, 4546, 3664, 4566, 3804]. Kahn [411]. Kakeya [2037]. Kakutani [1818]. Kampen [2332]. Kantor [3868]. Kantorovitch [939]. Kaplansky [2292]. Karamata [2547]. Kasparov [4077]. Katetov [2428, 39]. Kawamoto [3490]. Kegel [3034]. Kelley [2578]. Kepler [2247]. Kernel [4425, 4028, 3585, 2652, 2808, 2032]. Kernels [3379, 3499, 3735, 4531, 3109, 758, 3833, 2296, 4759, 852]. Kestelman [3825]. Khinchin [3191]. Khinchine [4281]. Khovanov [4276]. Kiguradze [3884]. Kiguradze-type [3884]. Kind [4430, 1918, 133, 890]. Kinds [4222]. Kirchberg [3368, 3831]. Kirchhoff [4625, 4448, 4762, 4362]. Kirchhoff-type [4762]. Kirillov [2033]. Kirk [1377]. Kirszbraun [4660]. Kiss [1038]. Kiss-precise [1038]. Klain [4815]. Klee [2938, 3615, 808]. Klein [1182, 883, 4139, 4685, 2147, 3345, 2247]. Kloosterman [3123, 3563, 2300, 1058]. Knapp [3156]. Kneser [1454, 4796, 4238]. knock [702]. knock-out [702]. Knopp [2262]. Knörrer [4565]. Knot [4306, 4116, 4278, 4129, 4784]. Knots [3120, 4633, 3504, 4662, 4075, 4435, 2287, 4791, 4697]. knotting [2606]. Kobayashi [2302]. Kohn [4739, 3835]. Koizumi [726]. Kolakoski [3394]. Kolakoski- [3394]. Kolmogorov [4707, 3643]. Komlos [2539]. König [4347]. Kontsevich [3669]. Korselt [4260]. Kostrikin [3648]. Köthe [3571]. Kottman [4414]. Krall [1985]. Krasnosel’ski [2614, 1460]. Krasnosel’skii [1956]. Krein [3629, 2904]. Krieger [3268]. Krivine [4547]. Krohn [4506]. Kronecker [4310, 1917, 1555, 1833]. Kronecker-type [1917, 1833]. Krull [2154, 1920, 772, 2465, 3464, 3862]. Kuiper [2550]. Kummer [2438, 3416, 3958, 852, 2344, 2768]. Kummer-Type [3958]. Künneth [3599, 699]. Kuo [2550]. Kuratowski [268, 244]. Kurzweil [3450]. Kwong [2837]. Ky [1686].

L [697, 3412, 2578, 146, 1547, 154]. L. [115, 171, 54, 525, 974]. Labelings [4123]. Lack [3583]. lacks [2142]. lacunary [2872, 443]. ladders [436]. Lagrange [263, 3976]. Lagrangian [3587, 3246, 3489, 4704, 4336, 4588]. Laguerre [3796, 3797, 444, 137]. Lambek [230]. Landau [2837, 293]. landmark [3582]. Landsberg [4664]. Lang [4039, 3753]. Langlands [3260]. Laplace [3666, 4247, 2387, 1371, 2495, 351, 420, 3295, 3609, 4625, 77]. Laplacian [4739, 4479, 3154, 3534, 3641, 2641, 4520, 2745, 4301]. Laplacians [3244, 3835]. Large [3637, 3734, 4493, 3473, 4640, 4312, 1924, 4492, 3764, 1930, 4824, 2994, 2060, 1800, 1450, 3032, 322, 2116, 2810]. larger [3170, 3162]. Largest [4210, 4005, 739]. L’arithmétique [1171]. Lascenov [2920]. Lasnev [2340]. Last [1489]. latent [951]. Latin [1681, 1940, 694, 596, 862, 2354, 155]. Lattice [3262, 2189, 2965, 676, 1895, 454, 4567, 4453, 2752, 1458, 28, 2220, 1795, 2654, 2781, 2782, 276, 2250, 427, 580, 537, 4743, 876, 1502, 2649, 1176, 877, 220, 2765, 1928, 4783, 2769]. lattice-ordered [2781]. lattice-valued [276]. Lattices [3620, 3850, 3888, 3395, 4377, 1117, 4196, 4057, 1533, 2770, 582, 732, 1616, 1963, 196, 1990, 127, 711, 3063, 2187, 2877, 2039, 387, 343, 1116, 1775, 279, 2628, 1186, 1313, 1281, 853, 648, 1314, 798, 2193, 4688, 2324, 3883]. Laurent [976, 2241, 3239]. Lavrentiev [2239]. Law [4665, 943, 4739, 1450, 685, 863, 2974, 1479]. Laws [3168, 3216, 1417]. Lawton [4387, 4066]. Lax [1711]. Layer [3657, 2099]. Layers [4616]. LCM [2792]. LCQ [4141]. Least [4556, 4142, 1737, 1551, 1152, 3813]. Leaves [3592, 2609]. leaving [1910]. Lebesgue [2299, 1266, 4351, 300, 908, 1691, 164, 2136, 1505, 512]. Lebow [3422]. Leffler [3247]. Lefschetz [3832, 4275, 3646]. Left [3686, 4603, 3441, 4598, 761, 4116, 4075, 3985, 4033, 3319, 4697, 409, 1777, 729, 2767, 103, 2348, 2887]. Left-orderability [4075]. Left-orderable [4116]. Left-ordered [4598]. Left-orderings [4603]. Legendre [4005]. Legendrian [2579, 3662]. Lehmer [3294, 3249]. Leibniz [4599, 4537]. Lelong [1452]. Lemma [4011, 4460, 4264, 3418, 4505, 4543, 4218, 2481, 2271, 340, 1963, 828, 2723, 2438, 1711, 329, 1205, 70, 3004, 2539, 2976, 1979]. lemmas [231, 203]. lemmata [1369, 1544]. lemme [2438, 3004, 2976, 2109]. lemniscate [1]. Length [4643, 4171, 3647, 3513, 4103, 4102, 418, 690, 3828, 744, 2290, 1597, 1041, 1502, 939, 1281, 1059, 2280]. length-preserving [2280]. lengths [2945]. lens [1538, 3072]. Leo [893]. Leonetti [2956]. Leopold [838]. Leopoldt [2852, 2344]. Leopoldt-type [2852]. Leray [3407]. L’espérance [97]. less [1205, 97, 2700]. Letter [1051]. Level [3369, 4502, 3862, 2796, 1983, 710, 2542]. Levi [2364, 1969, 4418]. Levi-like [2364]. Leviatan [974]. Levine [1612]. Levitzki [1104, 1572]. Lewis [803]. Lewittes [1547]. Li [3856]. Lichtenbaum [2742]. Lie [3148, 4155, 4244, 4611, 3265, 2311, 3266, 3153, 2753, 2979, 2796, 3559, 4385, 4160, 4521, 3790, 3842, 3829, 4286, 1440, 3354, 2780, 3285, 3868, 4184, 2503, 1515, 4464, 3853, 4337, 4207, 2273, 384, 4559, 2028, 766, 804, 1196, 3078, 2325, 1146, 2336, 2678, 3071, 1999, 2883, 3201, 3895, 3479, 4018, 1554, 1613, 3647, 1739, 3512, 4586, 3136, 4652, 3389, 3648, 3985, 2490, 2361, 2324]. Lie-algebra [1613]. Lieb [4146]. Liénard [575, 2720]. life [530]. life-test [530]. Lift [3551, 4240]. lifted [2002]. Lifting [4203, 708, 1237]. Liftings [4200]. light [2408]. Lightness [3885]. Like [3688, 3540, 1241, 1461, 2906, 2364, 3789, 4460, 2519, 3154, 1234, 1677, 3233, 1753, 668, 256]. likelihood [757, 1338]. Limanov [3751]. Limit [3924, 4107, 1319, 3624, 3520, 3443, 2451, 1561, 4255, 4615, 3855, 1609, 4658, 3238, 3363, 2720, 1032, 467, 503, 2954, 1848, 2693, 1743, 1635, 1916, 1015, 815, 2457, 1363, 1311, 1670, 3857, 2583, 4774, 261]. limitations [2651, 32]. Limited [4183, 3075]. Limites [364]. Limiting [4640, 4826, 4518, 2521, 253]. Limits [3987, 3906, 4501, 3374, 4358, 4072, 4071, 1449, 2083, 2392, 409, 364, 609, 130, 695, 2539, 252]. Lindelöf [2921, 1745, 3444, 1786, 4006, 2067]. Lindelöfness [2207]. Line [4288, 3505, 3860, 3838, 4154, 3146, 3221, 2189, 3011, 3830, 1087, 535, 291, 1328, 4724, 1808]. Lineability [4698]. linéaire [2211]. linéaires [1707, 3101]. Linear [3990, 3948, 1006, 4573, 3232, 3324, 3623, 3273, 4041, 528, 349, 1607, 3413, 2129, 380, 3836, 4252, 3643, 4067, 2804, 3341, 4255, 4375, 1145, 3079, 4095, 4340, 3403, 4195, 3528, 3919, 4564, 3451, 1930, 2921, 2612, 749, 2211, 1664, 2795, 1428, 1517, 1187, 1381, 1138, 721, 298, 1765, 2908, 4752, 761, 34, 265, 2347, 644, 3061, 1218, 42, 2708, 1188, 1827, 2466, 2014, 639, 555, 1273, 1014, 480, 2914, 518, 767, 835, 1736, 986, 140, 756, 1564, 108, 3101, 2864, 1312, 3052, 2011, 736, 918, 1279, 55, 91, 1693, 1092, 824, 1316, 2650, 2053, 3821, 2442, 855, 878, 2936, 2869]. Linéarisation [3222, 3274]. linearization [3274, 3222, 1448, 2935]. linearized [3821]. Linearly [3867]. Lines [4637, 3978, 1600, 2560, 1723, 4682]. Link [3710, 2599, 2606]. Linkage [3126, 2338]. linkedness [2756]. Linking [3593, 257]. Links [4370, 3593, 3249, 3294, 4709, 3169, 2999, 2661, 2098, 2342, 2585]. Lions [850]. Liouville [3586, 56, 4012, 3027, 3475, 3874, 3219, 2505]. Lipschitz [922, 4401, 4325, 567, 2329, 4083, 3640, 958, 2230, 490, 3892, 3974, 4122, 119, 3175, 2867, 3742, 3965]. Lipschitzian [3151, 1870, 2380, 1520, 2680, 2610, 3986]. lipschitziennes [2079]. Liquid [4419]. lissité [4819]. Little [1516]. Littlewood [3735, 2998, 4271, 3541]. Livingston [4573, 4360]. Lizorkin [4250, 3539, 4339, 4400]. Lizorkin-Type [4400]. Ljusternik [2026]. Local [4243, 3900, 2331, 4038, 4365, 3090, 3276, 4287, 3886, 3907, 3283, 3933, 4311, 4409, 3892, 4465, 1404, 3298, 3613, 3915, 539, 917, 1406, 1531, 4149, 989, 1904, 1680, 2747, 4488, 4154, 3762, 4242, 3317, 4300, 2270, 3806, 2593, 1920, 1641, 309, 262, 2170, 2124, 1190, 4669, 2840, 2184, 2232, 2711, 3814, 1950, 1042, 2248, 2104, 2542, 993, 3317]. localement [654]. localisés [1889]. Localization [4386, 4191, 3350, 1671, 2573]. localizations [2753, 1889, 844]. Localized [4120, 2357, 3203]. Localizing [4523]. Locally [3736, 3410, 4165, 4354, 2076, 4145, 4629, 3977, 4240, 4128, 2766, 4007, 1712, 2448, 2752, 1719, 1033, 2329, 2688, 1244, 3074, 2773, 2707, 2386, 1451, 1519, 1832, 654, 997, 2971, 806, 4754, 2082, 522, 1252, 1224, 1935, 2827, 1393, 312, 1227, 1953, 1821, 1287, 1729, 2245, 2263]. Locating [3242]. location [396]. Loci [3489]. Locus [4397, 4130, 2321]. Loewy [2395]. Loewy-series [2395]. Log [3843, 4538, 4747, 4644, 4735]. Log-concavity [4747]. log-convexity [4747]. log-majorization [4735]. Logarithmic [3703, 3537, 4030, 620, 1790, 1807, 2691, 1486, 1648]. Logarithms [4573, 3482, 1603]. logic [7, 21, 22]. Logical [2034]. Logistic [4268]. Lomonosov [3160]. long [520, 2185, 1272, 1351]. longueur [2963]. Loop [3210, 3998, 4216, 4749]. Loops [4203, 3709, 4216, 2455, 3177, 1687, 3820]. Lorch [115, 525, 974]. Lorentz [3789, 3409, 3108, 2938, 3109, 2658, 2001, 3506, 3000, 4273, 2475, 4783]. Lorentz-Schatten [3108]. Lorentzian [4234]. Lototsky [2007, 2191]. Low [3426, 4516, 4311, 3389, 3840]. Low-pass [4311]. Lower [2252, 4142, 4092, 3542, 727, 4103, 4102, 227, 1626, 1219, 476, 4763, 2763, 3857]. LS [3251]. LS-catégorie [3251]. Lucas [4824, 1583, 1624, 2309]. LUR [1575, 2142]. Lusin [202]. Lusternik [3503]. Lusztig [4676]. Luzin [4371]. Luzin-type [4371]. Lyapunov [2326, 1214, 3860, 3783, 712]. Lyme [4297]. Lyra [79].

M [738, 684]. M. [340, 1222, 177, 94]. Macaulay [4365, 3107, 1401, 2934, 3456]. Macbeath [684]. machine [207]. Mackey [2057]. MacNeille [2782]. Maeda [1035]. Magic [765, 134, 4123]. Magma [4228]. Magnetic [4223]. Magnus [3003]. Mahler [3902, 1834, 2225, 3041, 2552, 3625, 3576, 3281, 3897]. main [3029, 2403]. Maitland [564]. Majorations [1995]. Majorization [4631, 4735]. Majorization-minimization [4631]. majorizations [4797]. Makar [3751]. Malcev [989]. Mal′ [1238]. Mandelbrojt [1669]. Manifold [3441, 3213, 3861, 4103, 4102, 3765, 2771, 4752, 3061, 2148, 664, 522, 4412, 2946, 3818, 78, 1977, 585]. Manifold-Valued [3861]. Manifolds [4155, 4366, 4403, 4246, 3409, 3791, 3685, 3864, 3454, 3605, 4418, 3591, 3851, 4536, 4119, 4121, 4580, 3448, 3978, 3895, 4126, 3964, 4634, 4003, 4544, 3600, 4546, 3664, 4566, 3804, 3243, 4680, 2408, 4709, 2671, 2835, 2890, 1997, 449, 2653, 2008, 4424, 1582, 294, 1978, 1168, 981, 2819, 4139, 372, 3188, 3199, 3815, 1756, 2987, 522, 1223, 4804, 1843, 3006, 2541, 2959, 383, 414, 622, 2016, 1389, 1740]. Manin [3416]. manipulations [1470]. Manis [2100]. Mann [2614, 2354]. Mannheim [327]. Mannheim-curves [327]. Many [4247, 4602, 2398, 3025, 2511, 2177, 3754, 1405, 853, 2683]. Map [3531, 3122, 3871, 4104, 2546, 1601, 3028, 2624, 2461, 1447, 1938]. mapping [102, 548, 2183, 1247, 1248, 2295, 453]. Mappings [4262, 3483, 3640, 3522, 1170, 4470, 4548, 2521, 2238, 2890, 2795, 1034, 1381, 1465, 980, 2118, 1067, 2691, 1784, 1689, 1519, 672, 723, 1142, 1713, 1520, 1832, 2200, 2188, 997, 526, 1563, 2015, 2227, 703, 865, 3802, 2701, 807, 963, 1263, 634, 1230, 2610, 2749, 252, 2869]. Maps [4037, 3141, 3992, 3885, 3231, 3353, 3471, 3324, 3685, 3767, 3443, 3413, 4496, 3592, 4443, 4406, 3871, 3954, 3996, 3836, 4428, 3341, 3567, 4578, 3597, 3435, 3463, 3348, 2247, 4398, 3495, 3700, 4414, 2299, 2513, 2813, 2334, 4663, 1896, 2267, 3183, 2815, 2162, 2997, 2908, 3794, 603, 4737, 2129, 1897, 1906, 3016, 2914, 2233, 3079, 2932, 2101, 2011, 4746, 2975, 429, 1944, 118, 1254, 2339, 2650, 297, 2588, 1742, 1876, 1200, 2361, 2280, 1929]. Marcinkiewicz [2549, 3833, 3965]. Marcus [149, 177, 3530]. Marczewski [3523]. Marginally [4493]. Marica [782]. Markoff [2905]. Markov [4691, 1657, 1164, 2621, 2418, 646, 2376, 266, 2807]. Markovian [1204]. Marshall [1591]. Martingale [3425, 3127]. Martingales [3089, 3961, 3116, 1076]. Marvin [149]. Masks [3945]. Mason [1805]. Massey [4248]. Matching [4421, 3253]. Math [1413, 1098, 1161, 1160, 1619, 1102, 1317]. Mathematical [2168, 1735, 97, 7, 21, 22]. mathematics [2097, 2406, 218]. mathématique [97]. mathematization [1349]. Mathieu [921, 192, 571, 41]. Matrices [3232, 3108, 3625, 3474, 4578, 4410, 140, 4361, 3918, 3463, 3922, 826, 3550, 2924, 903, 2351, 3024, 4758, 1867, 1181, 105, 1464, 507, 2508, 4726, 829, 185, 873, 1219, 1307, 437, 1328, 116, 75, 2129, 646, 254, 181, 86, 936, 1128, 631, 753, 1630, 434, 90, 2061, 2441, 2584, 2665, 1984, 1434, 1387, 1475, 1871, 92, 641, 670, 3040]. Matrix [3263, 486, 2814, 4797, 3684, 3354, 3593, 4205, 3936, 3694, 3569, 3021, 3920, 1002, 951, 1003, 1657, 902, 1549, 1267, 1202, 1137, 1461, 1324, 1939, 3196, 2856, 2396, 2452, 138, 2418, 789, 737, 142, 1125, 424, 1819, 3049, 2914, 768, 2986, 1574, 1293, 508, 2846, 387, 223, 919, 1158, 1693, 1064, 469, 2339, 1093, 231, 203, 4735, 2278]. matrixes [739]. matroid [577]. matroids [1976, 678, 450]. matter [4508, 4509, 4568, 4569, 4589, 4627, 4628, 4648, 4686, 4687, 4705, 4706, 4766, 1427]. Mauldin [3671]. Maurer [4385]. max [2023]. Maxima [3968, 2933]. Maximal [3167, 3925, 4597, 3850, 4788, 4355, 3187, 4518, 4407, 3125, 3797, 3339, 4254, 4395, 3400, 3306, 4070, 4099, 797, 425, 3982, 4433, 718, 3661, 666, 3944, 4223, 4345, 4821, 2872, 690, 3196, 2754, 131, 2705, 3772, 2274, 1442, 3796, 255, 290, 1276, 1041, 948, 862, 1233, 1948, 817, 3817, 1282, 725, 2589]. maximal-resolvability [817]. maximality [1125, 2698]. maximaux [2556]. Maximizing [4138]. Maximum [4034, 4246, 4514, 1369, 1544, 1004, 943, 275, 1108, 438, 366, 2616, 757, 3004, 681, 1338]. Maxwell [4559]. May [4006]. Mazur [3054, 2291]. McCarthy [60, 8]. McCoy [563, 3711]. McQuillan [1547]. Mean [3809, 2059, 2139, 3089, 2916, 937, 3116, 2920, 3317, 3406, 2523, 1813, 3195, 2256, 813, 4767, 1670, 2973, 681, 2289]. mean-value [813]. meaning [296, 3158]. Means [4461, 3111, 4378, 4540, 3964, 4472, 2397, 375, 1111, 4742, 119, 1991, 1227, 1407, 421, 543]. Measurable [441, 4608, 2005, 871, 3840, 252, 337, 1231]. Measure [3532, 3902, 4351, 3625, 3548, 4518, 4622, 4147, 4026, 4594, 3897, 4343, 4221, 2870, 1615, 1719, 1834, 3041, 1639, 2552, 2756, 166, 2621, 144, 442, 2006, 2624, 569, 2025, 818, 941, 2542, 2449]. Measure-Preserving [3532]. measure-valued [2621]. measurement [1352]. Measurements [3841]. Measures [4601, 3576, 3281, 3486, 4020, 4163, 3124, 3291, 3956, 4298, 4321, 3981, 3839, 4379, 3879, 4596, 2788, 2850, 4399, 1267, 1436, 2873, 1787, 2225, 1324, 1939, 944, 706, 338, 2797, 1676, 409, 2896, 1419, 619, 2927, 1560, 276, 1014, 1921, 2264, 1363, 3084, 1312, 2719, 1863, 1132, 1816, 1095, 1287, 2245, 2263, 237]. mechanical [140]. Mechanics [3684, 184, 589]. meet [2220, 3066]. meets [4060]. Mehler [912]. Meixner [3672, 2393]. Melikov [512]. Mellin [3121, 3277, 2647, 1878]. Melzak [226]. Membership [3536, 4663]. Mendelsohn [4265, 1630]. Mennicke [2568]. Mercerian [2007]. Merge [4506]. Meromorphic [3247, 3392, 4814, 1174, 2903]. Merten [4073]. mesures [2896, 1899]. metabelian [2817, 4786, 3037]. metacyclic [2257, 2783, 2699]. metanilpotent [1041]. Metaplectic [3253, 3086]. Metastable [2148, 2715]. Method [4419, 3380, 4560, 3856, 4623, 2547, 1061, 2326, 880, 2307, 144, 1190, 2198, 2660, 1644, 3798, 2007, 1631, 1638, 31, 2115, 1946, 512, 4816]. méthode [2326]. Methods [4410, 620, 1135, 1413, 1481, 1790, 1813, 2613, 1486, 1705, 2264, 1322]. Metric [4262, 3470, 4350, 4331, 4534, 4209, 4122, 4622, 4147, 3493, 3916, 4594, 4192, 3962, 3897, 4363, 3765, 4715, 628, 4663, 2671, 1758, 2293, 2216, 2389, 1520, 3199, 1757, 692, 4193, 1934, 1149, 1150, 2809, 1374, 1688, 289, 1846]. Metrics [3559, 3950, 4535, 3913, 4315, 4500, 4645, 4052, 3819, 3985, 4664, 3789, 2860, 1168, 2409, 3756, 3698]. Metrizability [1990, 1909, 1530]. metrizable [1495]. Metrization [2056, 1389]. Meyer [2922]. middle [1610]. midpoint [1445]. Migration [3635]. Mikusiński [511]. Milgram [1711]. Mills [3437]. Milnor [4157, 4288, 3954]. min [2023]. min- [2023]. Minima [3723]. Minimal [4061, 4016, 3624, 4017, 3561, 4443, 4208, 3049, 4232, 4524, 3596, 3896, 4099, 3614, 3134, 3647, 3316, 4162, 2593, 561, 4825, 490, 2153, 884, 1683, 4498, 1921, 64, 2900, 2746, 774, 1739, 2588, 3056]. Minimally [3344, 1701, 932]. minimaux [754]. minimax [11]. minimization [4631]. Minimizers [4398]. Minimizing [3799]. minimum [1208, 469, 754]. Minkowski [2524, 4553, 3486, 3296, 3728, 32, 2700, 3118, 4815, 1349]. Minkowskian [3517]. Minority [4617]. Mirimanoff [3586]. mirror [4704]. mirrors [2462]. missing [1605]. Mitchell [724]. Mittag [3247]. Mittag-Leffler [3247]. Mixed [3121, 3655, 3887, 2147, 4399, 2260, 1602, 2724, 2696, 1275, 1938]. Mixing [4691, 3170, 1364]. MKZ [3598]. Möbius [3283, 436, 874, 4124, 4665]. Mock [3584, 4702]. mod [2855, 3236]. Model [4419, 4457, 3394, 1209, 3951, 4604, 3538, 4297, 3305, 4822, 3964, 4630, 3363, 3635, 2443, 4781, 62, 1205, 168]. Models [4385, 4530, 3750, 4453, 2703, 4677, 2337, 1810, 1338, 4661]. Moderately [3724]. modern [1515]. modes [1094]. modest [2482]. modification [498]. modified [4811, 78]. Modular [4262, 3139, 3267, 4159, 3547, 3275, 3605, 4134, 4497, 3581, 3300, 3540, 3729, 3307, 4563, 4626, 3312, 3942, 3103, 968, 2125, 4683, 2615, 4714, 1963, 2779, 2194, 127, 2187, 1803, 2968, 2971, 4811, 2383, 2582, 2847, 1281, 4807, 3805]. Modularity [4653, 3529, 582]. Module [2637, 4431, 2276, 2879, 677, 195, 1129, 387, 453, 2566, 715]. Module-theoretic [4431]. Modules [3967, 4038, 4365, 4420, 1376, 3687, 4287, 3886, 3907, 3768, 3909, 3014, 3933, 4117, 4607, 4407, 1143, 1627, 1677, 3507, 3613, 3915, 4051, 4059, 3940, 4098, 4359, 3614, 1492, 3661, 3464, 4242, 4300, 4417, 4674, 869, 3806, 2249, 2632, 2833, 320, 1456, 2415, 3182, 2796, 2891, 2937, 1065, 1084, 1429, 2259, 2401, 1290, 3046, 2780, 2637, 3753, 1401, 2802, 2841, 1329, 2820, 1820, 3814, 384, 2985, 848, 2646, 1310, 2151, 1882, 1278, 2203, 2918, 1494, 2989, 1701, 277, 1613, 1131, 730, 1632, 3166, 1892, 1850, 854, 966, 321, 651, 2769]. Moduli [3669, 3171, 4390, 3297, 3126, 4446, 4342, 4645, 3881, 783, 743]. Modulo [3624, 4372, 4666, 1268, 2705, 1086, 2876]. modulus [2251, 2261, 2396, 2127, 681, 1369, 1544]. Molchanov [1675]. Moment [616, 499, 3435, 1577, 1809, 358, 2217]. Moments [3097, 4233, 440, 576, 4747, 792, 2114]. momentum [413]. monads [1769]. Monge [4403]. Monic [4041]. mono [296]. Monochromatic [2980]. Monodromy [4157]. monoid [386, 2731]. Monoidal [3408, 4517]. monoïdes [1007]. monoids [2031, 1007, 72, 4021, 1343, 4773]. Monomial [4519, 4219, 3456]. Monomial-type [4519]. Monopolies [4210]. Monotone [604, 3643, 4057, 265, 2577, 1861, 938, 1568, 1672]. monotones [635]. monotonic [635, 2624, 162, 796, 2071]. Monotonicity [3469, 3878, 328, 811, 2307, 2038, 525, 464]. Monotony [198]. Montassier [4303]. Montel [2392]. Moody [2753]. Moon [551]. Moore [2420, 4749, 3177, 2608]. Moran [4601, 4017]. Mordell [3753, 3579, 4805]. Morita [4771, 2359, 2936]. Morphism [4256]. Morphismes [1361]. Morphisms [3141, 4027, 3016, 1361]. Morrey [4355, 4799, 4655, 4147, 4795, 4594, 4751, 4343]. Morse [4808]. Mortality [4604]. Mosco [4801]. Moser [1083, 896, 4710, 115, 171, 54, 135, 3144, 893]. Moser-Aubin [3144]. Most [1099, 1271, 1328, 1821]. Motion [3520, 3404, 1652, 2965, 2970, 2583]. motions [4675]. motives [4491]. Motivic [4230, 4507, 4239]. mots [1171]. Moulton [1556]. Mountain [3080]. moves [2342, 2585]. moyenne [827]. moyennes [2712]. MR2414205 [3764]. Multi [3327, 4398, 980, 773, 1254]. Multi-dimensional [4398]. multi-point [773]. Multi-Sided [3327]. multi-valued [980, 1254]. Multidimensional [3778, 4803, 529, 2715]. multidirectional [2973]. multiformes [2459]. Multifunctions [1709, 1217, 717]. Multilinear [3149, 4512, 4394, 2118, 2349]. multinomial [2323, 122]. multiparameter [2690, 2104]. Multiple [3650, 1470, 4462, 3955, 3815, 4024, 3460, 1045, 3780, 4448, 4301, 4399, 4567, 227, 4684, 2127, 741, 1996]. Multiples [4142, 1551, 1707, 53]. Multiplication [3746, 630, 3595, 3661, 888, 1632, 2356, 4349, 3184, 2004, 4738, 2400, 2467]. multiplications [3184, 1361]. Multiplicative [3637, 3622, 3152, 4666, 1132, 3365, 3764, 4744, 1603, 1188, 1015, 1655, 397, 1024, 4730, 926]. Multiplicities [4106, 3208, 1536]. Multiplicity [3967, 3417, 3423, 3761, 160]. Multiplicity-Free [3761]. Multiplier [3889, 4028, 4451, 4799, 2313, 1226, 1651]. Multipliers [3666, 3472, 2597, 3889, 4639, 1878, 3220, 2110, 1927, 3074, 2639, 4718, 2724, 3019, 2647, 2932, 2468]. multiply [796]. multipoint [3795]. multipolar [2623]. multiquadratic [3857]. multiquadratiques [3857]. Multiresolution [4582]. Multisublinear [4254, 4395]. Multivalued [4262, 3641, 1713, 963]. Multivariable [4035, 1960]. Multivariate [3623, 4485, 3882, 739, 639, 933]. multizeta [4728]. Müntz [4490, 1030, 2572, 4005]. mushy [2631]. Musielak [4594]. must [1084]. mutations [2661]. Mutual [2913]. Mutually [4112, 3674, 2398, 862, 1862]. MVBV [3743]. my [60, 124, 145].

n [468]. Nachbin [3168]. Nagata [4108]. Nagumo [1000]. Nagy [2484]. Naiïm [249]. Napoleon [3234]. narrow [217]. Natural [3291, 1906, 194, 1171]. nature [938]. naturels [1171]. Navier [3407]. nc [4823]. Near [3703, 3791, 478, 3405, 1456, 1457, 599, 2634, 2675, 1310, 539, 1114, 1494, 1131, 1526]. Near-Homeomorphisms [3791]. near-repetitions [2634]. Near-ring [478, 1456, 1310, 1494, 1131, 1526]. near-rings [1457, 599, 2675, 539]. Nearly [2011, 3804, 3786, 1859]. nearly-periodic [1859]. Nearness [2486, 2105, 1767]. Nearnesses [1988]. nearring [2986].

Nearrings [2943]. Necessary [4802, 3122, 3449, 3220, 1679, 612]. Necessity [468]. need [1233]. Nef [4667]. Negative [3323, 1549, 1626, 1619, 1537, 2514, 2141, 223, 4728, 965, 3821]. Negatively [4246, 3448]. négligeables [2304]. Negligibility [3261]. negligible [2304]. Nehari [4762, 4241, 1591]. Neighborhood [4611, 2018]. neighbours [1527]. Nerves [3408]. Nest [2555, 2892]. nests [2888]. net [2137]. Nets [1080, 3626, 3131, 1266, 1273, 283, 339, 1057]. Network [4620]. Neuigebauer [1336]. Neumann [4437, 4479, 3154, 1246, 1604, 4496, 80, 124, 4292, 1154, 3815, 4378, 3017, 4319, 2020, 1778, 4454]. Neutral [3972, 2512, 2722, 2362, 2758, 2725, 2695, 2214, 3066, 1262, 1115, 2563, 2748]. Nevanlinna [2468]. Newforms [4374]. Newton [4511, 3380]. nice [1129, 1695, 1688]. Nicholson [4604]. Nielsen [2107]. Nikishin [4352]. Nikodým [1246]. Nikodýn [4192]. nil [1759, 323, 2974, 4730]. nil-clean [4730]. nil-ring [2974]. Nilmanifolds [3248, 2279]. nilpotence [3026, 2052, 2194]. Nilpotency [1974, 2205, 2180, 4749, 2847, 4174]. Nilpotent [2424, 4305, 3231, 4350, 4654, 3736, 3410, 3223, 4373, 3959, 1728, 2195, 1459, 1829, 1108, 2775, 2177, 2926, 364, 1125, 2663, 846, 4445, 2283, 1730]. Nilpotent-by-Noetherian [2424]. nilpotents [1011, 364]. Nilradical [3481]. nilstufe [2260]. Nine [4211]. Nirenberg [4622]. Niven [491, 1120, 1637, 445]. no [1108, 744, 1619, 4442, 535, 2386, 1994, 598, 624, 877, 666, 1910]. no-three-in-line [535]. Nöbeling [3791]. Nodal [3644, 4474, 2265]. node [1625]. node-disjoint [1625]. nodes [1731]. Noether [3107, 4712, 317, 3646]. noetherian [730, 2424, 1755, 2391, 1627, 260, 1847, 2033, 935, 1385, 3464]. noeuds [3604]. nombre [97, 1995]. nombres [2952, 97, 3219]. Non [4456, 4155, 3468, 4284, 2995, 3107, 3196, 3440, 3441, 2673, 3867, 3886, 3504, 1122, 3831, 4290, 3722, 4515, 4649, 3607, 3972, 2983, 3609, 4337, 3935, 4465, 4252, 1236, 4067, 3956, 4232, 4231, 4481, 4622, 3301, 589, 3227, 3361, 4631, 3976, 4000, 3461, 3403, 4594, 3962, 4193, 3980, 4666, 3818, 3165, 4004, 489, 3388, 3571, 3677, 4007, 3554, 4596, 1427, 3391, 4733, 3466, 4057, 3882, 3965, 3456, 3259, 1456, 1549, 1458, 953, 356, 1187, 687, 1243, 1717, 1398, 2484, 1123, 1619, 1537, 1493, 4698, 2432, 427, 580, 1114, 615, 1183, 832, 3005, 417, 1521, 223, 720, 1280, 853, 109]. non [1316, 47, 172, 3821, 1410, 497, 1229, 337, 1231]. Non-abelian [4649, 1114]. non-analytic [720, 1280, 109]. Non-Archimedean [2673, 4465, 3980, 3388, 3571, 3677, 4698, 2432]. Non-Artinian [3886]. Non-Autonomous [3972]. non-biplanar [172]. Non-branching [4252]. Non-CM [3440]. Non-Cohen [3456]. non-commutatifs [615]. Non-Commutative [3607, 3466]. Non-Commuting [4456]. Non-Compact [3227, 4155]. Non-complemented [3956]. non-continuous [1398]. Non-convex [4631]. Non-convexity [1122]. non-cutpoint [497]. Non-cyclic [4666]. Non-degenerate [3867]. Non-Discrete [3935, 4284]. Non-doubling [4594]. Non-existence [489, 4733, 1456, 687]. Non-extendable [4290].

Non-Extremal [3165]. non-extreme [953]. Non-Finite [3461]. Non-finitely [4515]. Non-flat [4337]. non-Hamiltonian [47]. Non-homogeneous [4622, 3965, 356]. non-hyperfinite [1243]. Non-Integral [3554]. Non-isomorphic [3196, 853]. non-ITPFI [1717]. non-lattice [1458, 427, 580]. Non-Left-Orderable [3441]. non-linear [1187, 1316, 3821]. non-measurable [337, 1231]. Non-Metric [3962, 4193]. Non-Monotone [4057]. non-negative [1549, 1619, 1537, 223]. non-null [1183]. Non-Orientable [3554]. Non-Polytopal [3722]. Non-Principal [3976]. Non-Product [4000]. non-rational [417]. Non-real [2995]. non-reflexive [1123, 1229]. Non-Riemannian [4231, 4004]. Non-right-orderable [3818]. Non-Sidon [4649]. non-singular [1493]. Non-smooth [589, 1123]. Non-spectral [4596]. Non-splitting [3831]. Non-static [1427]. Non-Stationary [3609]. Non-strictly [4481]. Non-Strongly [3361]. Non-Trace [3391]. Non-Trivial [3403, 417]. Non-uniform [3882]. non-unimodular [832]. Non-uniqueness [2983, 1236]. Non-Vanishing [3468, 3005]. Non-Zero [3301, 4067]. Nonaveraging [2156]. nonbinary [2511]. noncentral [739]. Noncircular [4768]. Nonclassical [1881]. Noncoercive [3590]. Noncommutative [3325, 3660, 748, 708, 2744, 4805]. Noncompact [3585, 2251, 2867]. Noncompactness [4163, 4026]. noncompatibility [2484]. nonconnective [2184]. Nonconstant [3894]. Nonconvexity [3200]. Nondegeneracy [3868]. nondegenerate [1410]. nondeterminism [2542]. nonemptiness [3061]. Nonexistence [4540, 2991, 2108, 2933, 2750, 1852, 709, 1741]. Nonexpansive [3901, 1519, 1827, 1784, 723, 2172, 1832, 1254]. nonextinction [2621]. nonfinitely [2823]. Nonflat [4524]. nonhomogeneous [3065]. noninert [2700]. nonisometric [2653]. Nonisomorphic [2272, 1721]. nonisomorphism [2485]. Nonlinear [3900, 3590, 3641, 4268, 3795, 4209, 4024, 3240, 3986, 1012, 1923, 2451, 1828, 1173, 1854, 1366, 1568, 1315, 2720, 1027]. Nonlinearities [4255, 3963, 2747]. Nonlocal [4381, 3203]. nonmetrizable [929]. nonmonogenic [4785]. Nonnegative [3559, 3401, 2935, 3798]. Nonnegativity [4486]. nonnormal [2383, 1764]. Nonoccurrence [2303]. nonorientable [2819]. Nonoscillation [2776, 600]. Nonoscillatory [1622]. nonoverlapping [1176]. nonparallel [2798]. Nonparallelizability [2016]. nonparametric [1055]. nonpolynomial [824]. Nonpolynomially [4442]. nonpositively [3188]. nonprimitive [4743]. nonreciprocal [2225]. nonrecurrent [2418]. nonresidues [2435]. nonseparable [2732]. nonseparating [1956]. nonsimply [3259]. nonsimply-connected [3259]. Nonsingular [1768, 2937, 2246, 2769]. nonspin [4680]. Nonstandard [2383, 2870, 1810, 1261]. nontangential [2370]. nontorsion [4788]. nontrivial [3780, 1910]. Nonunital [4220]. Nonvanishing [4374]. nonzero [1011, 2657, 860]. Nörlund [2686, 905, 1905, 1100]. Norm [3105, 4650, 3887, 3889, 2639, 3854, 3383, 3449, 3102, 2374, 3062, 2923, 3370, 3772, 4496, 3834, 2724, 3099, 2696, 4049, 2147, 108, 2114, 3817, 4152].

Norm-attaining [3854]. norm-counting [2114]. Normal [4570, 2795, 981, 3210, 3490, 2417, 4145, 4314, 4194, 3072, 2953, 758, 721, 2798, 2654, 286, 74, 4389, 684, 2171, 1473, 86, 2383, 2173, 863, 1687, 597, 568, 2142, 1198, 1888, 2768, 6, 12, 20]. Normal-preserving [2795]. normality [786, 358, 3057]. Normalization [4466, 2894]. normaux [1950]. Normed [4015, 3627, 3343, 4375, 3975, 2614, 1428, 802, 2655, 349, 1218, 4426, 1273, 1359, 1736]. normés [1359]. Norms [3667, 4115, 3191, 3976, 3615, 3514, 2524, 1602, 2534, 2441, 2584, 2665, 4778, 1201, 3406]. Note [3260, 3059, 4038, 534, 1665, 4437, 561, 3106, 4285, 3950, 1098, 2617, 1035, 3325, 3736, 3928, 136, 507, 3625, 4387, 3769, 4019, 1299, 762, 3489, 144, 4514, 3381, 402, 4443, 1595, 3933, 372, 4662, 4143, 4294, 385, 4275, 518, 4468, 1703, 4277, 1156, 999, 4237, 3419, 4172, 4149, 293, 3543, 466, 4658, 2209, 4128, 1599, 4217, 1915, 3572, 3721, 3881, 3862, 4077, 484, 642, 856, 1286, 1479, 3946, 3243, 4303, 3847, 894, 379, 461, 1699, 2751, 563, 784, 1068, 2430, 2509, 2793, 4801, 2889, 1459, 533, 971, 40, 1428, 559, 1690, 944, 972, 2162, 1298, 1341, 48, 457]. note [449, 1181, 2429, 2772, 1738, 1397, 3074, 1450, 881, 697, 828, 4756, 33, 493, 1071, 1429, 1182, 2282, 2895, 1634, 1468, 1216, 1700, 1529, 1327, 2515, 1101, 1203, 751, 1471, 1073, 437, 476, 1596, 1685, 1451, 959, 456, 1220, 2024, 541, 2370, 2137, 1972, 2968, 1520, 2743, 2744, 255, 246, 473, 2371, 433, 1806, 2898, 4480, 688, 1686, 782, 1431, 1075, 769, 1597, 1861, 26, 63, 86, 1754, 998, 1146, 2283, 705, 1206, 65, 107, 391, 485, 863, 2416, 795, 1667, 1293, 695, 962, 1749, 16, 135, 1494, 2488, 1322, 1636, 4711, 1785, 2021, 611, 689, 776, 837, 2035, 2421]. note [680, 2377, 549, 1226, 583, 834, 1159, 19, 352, 1638, 4798, 1238, 1393, 568, 1490, 634, 1824, 798, 602, 52, 707, 612, 1888, 2053, 1879, 1953, 651, 2588, 544, 2590, 1742, 1740, 1894, 1651, 2324, 237, 2447, 85, 1035]. Notes [1694, 3808, 2351, 1867, 1814, 215, 46, 51, 159, 160, 161, 2353, 29, 55, 91, 613, 514, 657, 23]. Notice [2569]. Notion [4009]. nouveau [3582]. nowhere [1941, 1099, 1072, 2413]. Nuclear [1046, 3421]. Nuclearity [2330]. nucleus [482]. Nuij [4393]. Null [2658, 3254, 254, 1183]. nullification [3166]. Nullstellensatz [2242]. Number [3367, 3719, 3588, 3624, 3326, 4640, 3689, 3097, 4138, 3593, 3994, 4144, 3342, 3301, 4502, 3542, 4666, 4279, 3982, 4074, 4492, 405, 1323, 275, 2310, 749, 690, 1787, 2019, 1733, 132, 235, 2704, 292, 2552, 2, 240, 2659, 3047, 1626, 1992, 2274, 927, 1062, 4785, 3813, 1345, 97, 274, 874, 1041, 2522, 2845, 2763, 3067, 305, 46, 51, 159, 160, 161, 2285, 189, 3162, 1875, 223, 1893, 2140, 853, 2116, 774, 1770, 965, 1527, 1285, 856, 2479, 3179]. Numbers [3683, 2121, 4062, 4436, 3193, 3010, 4161, 4573, 3832, 4086, 4091, 2945, 3758, 4194, 4562, 4473, 4632, 4242, 4452, 1653, 194, 2921, 2905, 2952, 471, 1269, 1450, 1779, 1234, 1583, 1624, 4812, 2281, 858, 97, 1144, 1472, 1868, 1965, 99, 160, 1474, 2003, 4794, 727, 3219, 315, 1179, 168, 1478, 53]. Numerators [4062]. Numerical [3232, 380, 3341, 3174, 3741, 3629, 1157, 3633, 1476, 3828, 215, 3200, 1172, 29, 55, 91, 613, 3038]. Numerical-valued [1476]. Nyman [4469]. obeying [685]. object [1715, 3612, 1424]. objectifs [1707]. Objects [4550, 1509, 2674, 3857, 2972, 1531]. Obreanu [170]. observation [1016, 94]. observations [943]. Obstacles [3630]. Obstruction [3274, 3998, 4683, 3274]. Obstructions [3197, 3871, 1900]. occupation [2777, 1274]. Octic [3454]. Octonion [4115]. Odd [3287, 69, 4523, 3449, 4054, 749, 2759, 2296, 2287, 3177]. odd-dimensional [2759, 2287]. ODE [2947]. off [4673, 4591]. Oka [4823]. Olkin [1591]. omitting [1441]. One [3901, 3733, 3484, 3734, 3667, 3269, 3720, 4308, 3907, 3889, 3114, 4499, 766, 4315, 4502, 3551, 2731, 2161, 4376, 3313, 3423, 3495, 2789, 2154, 2753, 3789, 2754, 2838, 1010, 2982, 1430, 1969, 744, 843, 2639, 4520, 322, 830, 768, 2186, 310, 1364, 2700, 3067, 679, 4489]. One-Dimensional [3733, 3907, 766, 2731, 3313, 1010]. One-Parameter [3901, 4308, 4376, 310]. one-relator [1430]. one-sample [1364]. One-sided [2161, 2982, 830, 768]. ones [1137]. only [749, 2177, 1619, 1537, 430, 195, 2089]. Onofri [3966]. Onofri-type [3966]. onto [1852, 429, 1447]. Oort [4718, 4466]. Open [1013, 1160, 4371, 3781, 4634, 428, 2925, 1247, 2301, 2011, 3782, 2976]. opérateur [2387, 1089]. Opérateurs [1891, 1746]. Operations [3104, 4228, 2184, 2232, 2290, 1262]. Operator [4803, 4621, 3735, 4327, 3654, 3865, 4330, 3769, 3560, 3869, 4168, 3339, 3911, 3330, 3510, 4296, 3383, 3089, 4049, 3936, 4199, 4411, 4097, 3418, 1794, 3116, 3589, 3866, 4018, 3377, 3819, 4586, 4364, 3076, 1436, 2387, 2220, 4684, 1765, 2575, 1746, 3772, 2150, 1675, 3834, 1330, 4707, 2376, 1705, 2881, 1172, 2253, 1089, 3759, 4099, 988, 2865, 3817, 66, 2374, 1133]. Operators [3499, 3620, 3883, 4010, 3948, 4384, 3452, 3105, 3453, 4202, 3826, 4082, 3142, 3639, 3746, 4178, 3232, 3093, 3748, 4654, 4308, 4181, 4164, 3706, 4166, 4355, 3932, 3187, 4635, 4518, 3430, 4293, 4519, 4206, 4391, 4141, 3127, 3836, 4169, 3233, 3956, 3341, 4482, 4522, 3957, 3873, 3629, 3959, 4070, 3598, 3421, 3436, 4449, 4433, 4379, 4595, 4638, 4548, 3391, 4416, 4223, 3947, 2546, 1480, 2794, 2738, 2672, 2686, 2814, 3024, 3195, 604, 1324, 257, 1517, 1325, 2429, 3792, 3043, 682, 1845, 747, 1747, 4741, 946, 4689, 2800, 1213, 1746, 4770, 3833, 1852, 2555, 2657, 1634, 2055, 2692, 3013, 884, 2413]. operators [2308, 4738, 2760, 931, 3064, 3189, 1684, 480, 3099, 2432, 2358, 2985, 3042, 2804, 767, 688, 835, 1113, 1540, 861, 413, 1077, 2141, 2729, 4822, 2562, 3051, 1891, 3101, 1174, 2865, 2958, 2649, 4703, 2564, 1336, 1279, 2289, 2666, 1226, 1869, 818, 2254, 2518, 4724, 1568, 2020, 2093, 888, 1446, 511, 481, 1092, 2312, 4723, 1953, 4731, 2483, 2904, 2810]. Opial [483, 1697, 1197, 403]. Oppenheim [2277]. Optimal [4477, 4708, 1958, 3133, 4053, 3315, 3984, 2211]. optimales [2211]. Optimization [3373, 4631, 2158]. Orbifold [3816]. Orbisurfaces [3690]. Orbit [3340, 4752]. Orbital [3253, 1186]. Orbits [3153, 4180, 3528, 1758, 3829, 1141, 4737, 4682, 2848, 2490]. Order [3884, 344, 3452, 4347, 4478, 4551, 3720, 3747, 3473, 3849, 3930, 4555, 3355, 4271, 3997, 152, 3132, 3584, 4096, 4593, 3874, 1585, 3351, 4549, 3716, 3882, 1296, 3763, 1578, 1579, 1454, 1004, 3990, 1930, 1698, 1208, 2794, 1663, 2064, 4776, 1187, 3789, 2771, 4793, 438, 2045, 1399, 1214, 644, 1012, 2620, 2776, 788, 1166, 2657, 366, 1923, 134, 3795, 2281, 2014, 504, 741, 458, 1828, 1890, 554, 459, 198, 1196, 1881, 1985, 1127, 1609, 2581, 709, 2285, 494, 938, 2012, 3052, 434, 736, 1760, 1049, 1336, 2748, 2734, 1198, 1512, 1640, 667, 2886, 155, 209, 878, 4198, 4223, 2949, 3850]. Order-interval [4347]. order-isomorphisms [4776]. Order-Preserving [3716]. order-theoretical [2581]. Order-Two [3747]. orderability [4306, 4412, 4075]. Orderable [3441, 2688, 4116, 1511, 3818, 4697]. Ordered [4367, 3131, 1967, 1570, 1351, 975, 2978, 4598, 1718, 2508, 1594, 2274, 2818, 2781, 2707, 814, 2078, 473, 1136, 1696, 1176, 151, 128]. Ordering [3988, 773, 820]. Orderings [4603, 3566]. Orders [3194, 4614, 4598, 3209, 3774, 3125, 4729, 3196, 4745, 1789, 2286, 2864, 621, 853, 1339]. ordinaires [2996]. Ordinary [4593, 1830, 4665, 1578, 557, 272, 2996, 1390, 2307, 788, 1166, 2358, 1828, 3]. ordonnées [128]. Ore [4335]. Orientable [3554]. orientations [430]. Oriented [1976]. Origin [4615]. Originating [3684, 3354]. Orlicz [4594, 2709, 3000, 4655, 2534, 3570, 4587, 2543, 2447]. Ornstein [3170, 2502]. Orthogonal [4158, 3949, 4421, 4556, 2027, 792, 4653, 886, 3131, 3694, 1760, 4056, 3944, 3530, 4261, 288, 3029, 1978, 1484, 2998, 2662, 805, 1881, 862, 709, 2354, 2363, 2425, 1554, 2682, 155, 210, 6, 12, 20]. Orthogonality [4411, 2410, 2816, 2421, 103]. Orthogonally [1504]. ortholattice [2615, 1803]. orthomodular [2770, 1314]. Orthonormal [4063]. Oscillating [3123, 4330, 2300]. Oscillation [1003, 1395, 3884, 2703, 1012, 2939, 2495, 1828, 1980, 1851, 2514, 4593, 878, 3317, 3763, 2139, 2620, 1182, 1923, 2362, 2758, 2014, 458, 2012, 3052, 1021, 1640, 544]. Oscillations [2214, 2748, 1930]. Oscillators [4572]. Oscillatory [4635, 3065, 4483, 1187, 2722]. osculating [198]. Osculatory [693]. Osgood [4334]. Ostrowski [2038]. Outer [3906, 1816, 4415, 2788, 2850]. Outline [7, 21, 22]. output [1564]. ouvertes [2976]. oval [2041]. Overconvergent [4044]. overgroups [2915]. Overlapping [3724]. overpartitions [4702]. Overrings [2812, 1079, 2090, 1598, 1660].

p [4669]. P. [129, 525, 950, 443, 1317]. Pach [3726]. Packable [4813]. Packing [1253, 2481, 227, 662, 100, 553, 259, 331, 330, 332, 419, 1527, 1900]. Packings [3133, 693, 953, 427, 580, 428]. Padé [1926]. Pair [4299, 3229, 404, 789, 3007, 591, 1541, 3020, 139]. paire [3020]. paired [1684]. Pairing [4523]. Pairings [2717]. Pairs [4200, 3868, 4389, 4543, 4656, 4399, 1181, 3015, 373, 90, 4638, 484]. Pairwise [3944, 2054, 911, 1337]. Palais [3592]. Paley [96, 3735, 2998, 4271]. Pankov [4241]. Pants [4172]. Paouri [4079]. Paper [4103, 2351, 2205, 3812, 8, 60, 1805, 124, 411, 566, 145, 1156, 239, 177, 670, 2567]. Pappian [1795]. Pappus [433, 538, 558]. para [2207]. para-Lindelöfness [2207]. Parabolic [3702, 3991, 3735, 3864, 3143, 4229, 3172, 934]. paracompact [1067]. Paracompactness [1098, 1594, 4145, 809, 2207]. paracontact [3765]. Parallel [1600, 4168, 3911, 4329, 3169, 1964, 4018]. Parallelism [4199]. parallelogram [1417]. parallelohedra [89]. Parameter [3901, 1930, 1353, 4308, 788, 1166, 74, 310, 586, 987, 4376, 2071]. parametric [2390]. parametrization [4498]. parametrizations [2125]. parametrized [1969]. Paraproduct [4595]. Parareductive [2800]. parfaits [1346]. parity [2553]. Parrott [4638]. Parseval [4109, 4063, 4582]. part [714]. Partial [3244, 3094, 4227, 4137, 3329, 4316, 1339, 4600, 4609, 2206, 4756, 4726, 2286, 934, 1705, 745, 836, 795, 2354, 2591]. Partially [1082, 1119, 1718, 2274, 843, 128, 652]. partially-ordered [128]. particle [2777, 1062, 175, 174]. particles [2834, 4716]. particules [2841]. partiellement [128]. partielles [4756]. partite [123]. Partition [167, 2414, 606, 979, 1161, 1124, 930, 2698, 2021]. partition-inducing [2021]. Partitioning [1973, 496]. Partitions [194, 4357, 3799, 520, 2019, 2189, 3045, 2598, 1562, 1275, 10, 1875, 1967, 1631, 1934]. parts [2102, 2403, 1925]. Parzen [3882]. Pascalian [2041]. Pasch [1028]. Pass [3080, 4311]. passage [923]. past [1311]. paste [2901]. pasting [1843]. Pastro [3239]. Patch [4604]. Path [4238, 744, 2502]. paths [365, 676, 537, 1680, 220, 454]. pattern [3823]. patterns [1110, 3033, 949]. PBW [4493]. Peano [2196, 1783, 755, 1017]. Pearl [670]. pedestal [3020]. Peetre [850]. Peirce [3137]. Pell [3323, 3976]. pencil [3049]. Pencils [4393]. Penrose [3048]. Percolation [3048]. Perfect [3652, 3165, 3319, 1846, 649, 3752, 2911, 1346, 1275, 315, 1285, 2887]. Perfectly [1289]. Perforated [3364]. Perimeter [3799, 690]. Perimeter-Minimizing [3799]. Period [4245]. Periodic [4419, 4478, 565, 3904, 1759, 3111, 1142, 4602, 4604, 3995, 3997, 4024, 4340, 887, 4195, 2848, 1766, 4380, 4588, 2095, 2995, 1242, 1182, 2939, 2638, 2862, 1768, 672, 398, 270, 119, 1859, 3780, 3085, 1279, 4798, 2947, 731, 879]. Periodicity [4565, 1927, 3783, 4690]. périodique [2638]. périodiques [1859]. Periods [3805, 2747]. Permanent [4291, 477, 1464, 873, 1219]. permanental [3040]. Permanents [1328, 197, 784]. permutable [1134, 4112]. Permutation [3393, 994, 4444, 3307, 3103, 205, 285, 1191, 2478, 1465, 1141, 554, 179]. Permutations [1653, 87, 800, 2162, 914, 69]. Permutative [4636]. Permuting [1659]. Perron [2094, 4159, 2102, 636, 831, 2136, 4452]. Perspective [4161]. Perturbation [3653, 3093, 4556, 4615, 4760, 788, 1884, 1115, 109]. Perturbations [4292, 3996, 3391, 1244, 2520, 3800, 2933, 4816]. perturbed [2099, 3011, 1851]. Peterson [4184]. Pettis [2555, 3706, 1869]. PFA [3849]. Phase [4430, 3065, 4795, 4711]. Phases [4635]. Phenomena [3534]. Phenomenon [4527, 2239, 1407, 512]. Philip [4]. Photo [3681]. Photon [3827]. Phragmén [1786]. physics [2097, 1515, 1860]. PI [2033]. Picard [3700, 3367, 565, 3655, 4782, 4429]. PID [2090]. pieces [1973]. Piecewise [4161, 4623, 1930, 2624, 2231]. Pierce [1641]. Pigeonhole [3194]. Pillai [4763]. Pisier [4671, 3530]. Pisot [2122, 4013, 3508, 3758, 3313]. Pitaevskii [4025]. Pitt [1972]. PL [2606]. Places [3566]. Planar [3979, 4303, 3847, 313, 1193, 1883, 496, 1048, 2086]. Planarity [4237]. Plancherel [4353]. Plane [4304, 4010, 4510, 3577, 3592, 4168, 4483, 4049, 4234, 3234, 4257, 3631, 4504, 4433, 4473, 4545, 3512, 4434, 1579, 3149, 2334, 1080, 404, 1621, 28, 3810, 2997, 2163, 662, 1110, 1203, 2317, 3013, 4464, 2821, 4682, 167, 182, 521, 2727, 4743, 4329, 2684, 1946, 242]. Planes [3486, 3246, 3728, 3975, 4215, 1556, 2737, 344, 133, 4810, 1965, 2062, 2944, 225, 2901, 2766, 1570]. plans [3029, 2342]. Plaque [3891]. plasma [1860]. Plate [3908]. plumbing [2287]. Pluriharmonic [3042]. plurisubharmonic [2616]. Plus [3474]. plusieurs [1030]. Pohozaev [4762]. Pohozaev-type [4762]. Poincaré [4460, 4479, 4288, 3396, 4699, 2761, 1223, 1938]. Poinsot [2247]. Poinsot-Type [2247]. Point [3951, 3094, 4288, 3930, 3995, 4520, 4209, 3358, 3856, 3386, 3874, 4505, 4543, 4626, 3879, 4221, 4130, 742, 1005, 628, 1377, 404, 1517, 946, 1718, 1633, 2041, 980, 3428, 1037, 2407, 2286, 1713, 2380, 2172, 3582, 997, 3077, 1754, 4629, 4819, 679, 2946, 2885, 773, 1366, 2866, 877, 717, 1818, 1149, 2830, 1547, 1230, 2610]. Point-Correlations [3951]. point-of-continuity [2407]. point-symmetric [742]. point-to-set [1230]. Pointlikes [4506]. Points [4455, 3990, 3261, 4641, 4651, 3273, 3708, 4602, 4667, 4499, 3342, 3402, 3696, 3102, 4489, 3921, 3335, 1350, 2833, 2995, 3322, 3183, 404, 687, 28, 1462, 1534, 1301, 1913, 2080, 264, 3428, 1441, 1402, 4826, 983, 672, 1142, 2709, 439, 2098, 1832, 847, 2188, 2265, 815, 1172, 2295, 2473, 2931, 3779, 3079, 4743, 1704, 4561, 2294, 1796, 1567, 1198, 2831, 1410, 2058, 2903, 1467]. Pointwise [3968, 1534, 3108, 3640, 1090, 1263, 1023, 4549, 1, 1496, 402, 2304]. Poisson [3642, 4391, 1358, 4235, 3450, 2032, 3316, 4416]. Pol [3558]. polar [1313, 549]. polar-dense [1313]. polarities [1645]. Polarized [4392, 606, 3061]. pole [2734]. Poles [3138]. Polish [3523]. Pollaczek [3672, 2393]. Pólya [4722, 2345]. Polyak [3553]. Polycharacters [3264]. polycyclic [1136]. Polydisk [4164, 2940]. Polyèdres [1842]. Polygon [3334, 25]. Polygonal [3126, 2369, 2811, 2663, 2941]. Polygons [3531, 4641, 4462, 3297, 3211, 952, 3171, 4738, 718]. polyhedra [1842, 2310, 560]. polyhedral [570]. polyhedron [1447]. Polynômes [2876, 1715, 2952, 2896, 996, 2465]. polynômiale [1359]. Polynomial [3120, 3090, 4201, 3767, 3326, 4288, 1401, 4635, 3249, 3294, 4578, 4212, 4430, 4615, 4446, 3373, 4148, 3525, 3675, 4382, 2753, 1483, 3274, 2552, 2896, 1847, 1359, 2475, 998, 394, 4682, 2727, 833, 2294, 1444, 2087, 2629, 2585, 2611, 29]. Polynomiality [4161]. polynomially [2564]. Polynomials [4200, 4458, 4624, 3734, 3667, 3843, 3473, 3279, 141, 4041, 4532, 3563, 4140, 3852, 3212, 3740, 4393, 3672, 4067, 3342, 4050, 3360, 3541, 3694, 2988, 3758, 3837, 4623, 4447, 4528, 4486, 406, 3117, 4563, 3983, 4074, 4584, 4005, 3586, 3239, 4600, 4609, 592, 4222, 4261, 3315, 1715, 500, 527, 2047, 2410, 2037, 1733, 2952, 779, 2478, 2517, 1692, 528, 2554, 616, 2909, 1625, 1759, 1086, 1619, 1537, 2876, 2426, 2024, 2662, 396, 2559, 2027, 76, 792, 996, 1421, 1360, 2626, 1881, 1985, 1088, 156, 188, 2393, 550, 524, 2627, 2435, 3067, 2227, 2025, 444, 939, 228, 783]. polynomials [918, 1840, 2606, 2425, 710, 325, 1159, 1151, 2920, 1833, 1917, 2071, 2682, 2935, 1477, 447, 446, 506, 484, 642, 137, 3204, 4816]. polynomiaux [3274, 3237]. Polystable [3991]. Polytopal [3270, 3722]. polytope [211]. Polytopes [3809, 4553, 4389, 3729, 2404, 304]. polytopes* [3018]. polyverbal [1262]. ponctuel [4819]. ponctuelle [2244]. ponctuellements [2304]. Pontrjagin [2088]. ponts [373]. Poonen [4772]. Poorly [2576]. population [2703]. populations [964]. porosity [2156]. Ported [3168]. Portraits [4430]. Pöschl [4716]. posed [577, 1844, 974]. posedness [4551, 4226]. poset [2633]. Posets [4347, 3575]. positifs [996]. Positive [4058, 946, 3590, 4673, 4268, 3910, 4426, 2409, 4311, 4520, 3129, 3657, 3226, 3341, 4296, 3360, 4629, 4645, 3879, 4129, 3240, 3923, 2869, 3456, 594, 716, 1497, 4680, 2513, 3824, 903, 1664, 1615, 1202, 173, 2489, 3771, 4726, 2928, 3064, 480, 2491, 2914, 2514, 996, 2626, 181, 1128, 1895, 2991, 1646, 2865, 223, 680, 4668, 4774, 1737, 1945, 1092, 2339, 92, 2567]. Positivity [3703, 3527]. Positivstellensatz [3227]. possible [1273, 1251, 521]. Possibly [4220]. Potential [4330, 3560, 3358, 2481, 4762]. Potentials [4594, 4716, 2208, 4795]. pour [3022, 3088, 3147, 3205, 3256, 3258, 3378, 3438, 3498, 3556, 3618, 3682, 1030, 2266, 827, 3428, 2712, 1658, 1995, 2976]. Poussin [855]. Power [4058, 3323, 4013, 4349, 3844, 4286, 3047, 4044, 4518, 729, 1601, 1437, 1813, 2261, 656, 2772, 882, 2894, 1271, 1626, 1807, 1190, 2275, 504, 819, 2241, 1991, 1648, 52, 303, 2936]. Power-associative [729]. Powers [3652, 4775, 2499, 3853, 4610, 4213, 4484, 2774, 1606, 4114, 4818, 223, 1618, 158, 256]. PP [2716]. PP-endomorphism [2716]. practical [2199]. Pre [3790, 3998, 2222]. Pre-Lie [3790]. Pre-Quantization [3998]. Pre-Tits [2222]. precise [1038]. Precompact [3931, 1244]. Preduals [4225, 3865, 4166, 3859, 2381]. Preenvelopes [4117]. prefrattini [954]. preimages [1793]. premiers [2952, 97]. prenant [2952]. Preparation [4149]. Prescribed [3138, 4611, 4720, 3531, 4574, 3689, 406, 1387, 489, 826]. presence [757]. presentation [1414, 571, 1871]. Presentations [1085, 2817, 3046, 2368]. presented [1539]. Preservation [3010]. preserve [3079]. Preserved [4021]. Preserving [3532, 3471, 3232, 4496, 3836, 3341, 3916, 3463, 3716, 2813, 2795, 2815, 449, 2113, 1561, 554, 2804, 1200, 2544, 2280]. Presheaves [3538]. presque [248, 1859]. presque-périodiques [1859]. Pretzel [3249, 3294]. primaires [2444]. primal [2593]. primality [992]. Primary [2002, 2352, 3132, 3389, 1755, 2403, 3177]. Prime [4476, 3734, 3624, 4204, 3185, 3287, 3098, 3290, 4068, 4318, 3132, 1751, 3918, 3136, 4632, 2548, 2367, 1959, 3081, 1461, 2952, 2261, 618, 2046, 1345, 1627, 3754, 2273, 255, 388, 540, 830, 860, 1661, 97, 2463, 2930, 516, 1392, 2015, 2678, 2435, 2700, 3067, 1322, 1727, 52, 2348, 2949]. prime-power [2261, 52]. prime-producing [2435, 3067]. Primes [3271, 4640, 3398, 4073, 2154, 3025, 4748, 2515, 899, 2039, 2988, 46, 159, 2467]. primitif [754]. Primitive [4641, 4160, 3663, 971, 28, 2855, 636, 138, 4763, 2900, 754]. Principal [4605, 3702, 3991, 4775, 3534, 3976, 1613, 1418, 2384, 2580, 3200, 2972, 2348, 826]. Principe [3004]. Principle [4262, 3091, 4109, 4334, 4514, 4602, 548, 298, 184, 2616, 1786, 479, 413, 3004, 542, 1979]. Principles [4267, 4692, 1320, 2058]. prior [757]. priority [2850]. Privaloff [422]. pro [4765, 2544]. pro- [4765, 2544]. Probabilistic [1516, 4146, 2618, 1797, 2032]. probabilités [2079]. Probability [2210, 253, 2264, 310, 104, 2245, 2263]. Problem [3620, 3883, 4305, 3827, 4368, 4266, 3505, 3412, 3704, 3908, 4418, 4533, 4356, 4317, 4123, 3874, 4503, 4563, 3422, 3677, 3920, 4343, 4596, 190, 278, 405, 673, 2851, 84, 1716, 577, 2094, 1106, 1577, 72, 1844, 1083, 2574, 896, 431, 235, 1035, 474, 732, 59, 1718, 306, 872, 2, 170, 206, 617, 897, 4770, 2166, 1140, 1109, 4753, 1969, 115, 790, 2638, 2158, 2694, 1788, 54, 535, 1940, 930, 129, 1734, 2642, 2969, 646, 146, 1809, 2970, 1236, 764, 2022, 934, 3017, 274, 874, 745, 836, 71, 1748, 647, 263, 65, 107, 104, 98, 2627, 267, 1115, 4700, 2537, 2077]. problem [100, 2294, 877, 950, 2315, 245, 2305, 1130, 3053, 1064, 735, 492, 653, 419, 808, 109, 1448, 1264, 2961, 443, 587, 1317, 2217, 3821, 1340, 299, 2631, 926]. problème [1035, 2638, 129]. Problems [4475, 4479, 4494, 3930, 3609, 3997, 4024, 280, 4173, 205, 557, 2228, 2307, 1430, 1308, 2660, 4725, 3795, 504, 2335, 3815, 2645, 499, 1262, 1173, 531, 2168, 1488, 3780, 2505]. procedures [215]. Process [3988, 1652, 1121, 2621, 2715]. Processes [3519, 3129, 4691, 2385, 1639, 1204, 1635, 2104, 2502, 2786]. processus [1121]. producing [2435, 2988, 3067, 429, 1944]. Product [3140, 4328, 3623, 2474, 4112, 3536, 3869, 4291, 4296, 4500, 4212, 4000, 3941, 3731, 4656, 4055, 4077, 4664, 1755, 1933, 581, 1378, 1297, 1482, 440, 1397, 1536, 2655, 2950, 311, 2557, 619, 2423, 899, 2598, 2877, 4812, 372, 2472, 2188, 2264, 688, 794, 1088, 2210, 26, 2439, 522, 864, 46, 1948, 2322, 2500, 2885, 2865, 833, 3803, 822, 2231, 1180, 969, 1576, 2044]. Productively [4006]. Products [3732, 4108, 3471, 3747, 4598, 4248, 2128, 4440, 1419, 2153, 4185, 4186, 3626, 3446, 4338, 4088, 3329, 3299, 3252, 4629, 4612, 3977, 3526, 4644, 4630, 3675, 4472, 4701, 1694, 2369, 2736, 2811, 1435, 2633, 1656, 1297, 1098, 4741, 4678, 1606, 2246, 661, 2760, 2219, 2929, 2370, 1777, 3001, 2663, 2726, 2941, 2643, 2625, 4811, 3078, 733, 935, 1782, 4754, 1954, 2602, 556, 2114, 2284, 1176, 4774, 568, 2765, 1741, 990, 2849, 1093, 891, 1506, 501, 2073, 809, 3057, 2483, 1185]. Profinite [1065, 4713, 1904]. program [2211, 114, 145, 207]. programmation [1707]. programme [2211]. Programming [1735]. progression [1631, 877, 492, 1770, 2464]. progression-free [1631]. Progressions [3092, 3624, 4142, 4317, 3921, 520, 1761, 1618]. Projection [3720, 4654, 4189, 211, 1852]. Projections [3787, 2298, 3447, 3974, 3979, 1657, 789, 2628, 1314, 312]. Projective [3230, 2770, 3273, 3692, 4618, 3526, 3589, 3866, 4018, 2440, 4434, 4565, 3700, 4417, 3456, 2592, 2612, 3167, 2334, 2997, 1963, 2392, 133, 2216, 711, 2759, 4810, 1820, 152, 4256, 2203, 2108, 225, 607, 2901, 2565, 854, 966, 2159, 261, 2966, 364]. Projectively [3913, 3698, 2612, 2805]. projectives [364, 2558]. projectivity [2166]. Proof [4079, 4457, 4132, 3518, 2478, 4272, 3458, 3508, 4146, 3003, 4050, 3596, 3235, 4623, 551, 1460, 548, 2069, 2404, 307, 186, 2102, 2197, 1244, 1382, 1372, 883, 1126, 3173, 2313, 1874, 1195, 1762, 1655, 3018, 230, 2457, 1026, 1630, 11, 390, 653, 2355, 2684, 1547, 643]. proofs [2414, 1516, 3149, 56, 216, 2787]. Propagation [3218, 1514]. Proper [3940, 3060, 3810, 2229, 1910, 1946]. Properties [4455, 785, 3469, 3138, 4605, 3181, 4572, 3686, 1610, 4552, 4640, 4021, 4249, 3706, 4576, 3250, 4607, 3955, 4208, 3872, 3330, 4339, 3673, 3727, 2439, 3201, 4626, 717, 1716, 2105, 2251, 2547, 56, 2738, 486, 2256, 2364, 1438, 38, 50, 309, 2373, 2228, 221, 957, 185, 2722, 1560, 3000, 2579, 424, 985, 1980, 2200, 2171, 1309, 1360, 2092, 4480, 4811, 188, 412, 4754, 323, 267, 213, 444, 5, 2502, 1132, 1239, 1255, 1590, 1408, 3069, 2193, 3054, 110]. Property [3194, 4347, 3153, 3653, 4177, 4063, 4043, 4165, 4354, 4409, 4187, 3580, 3130, 4192, 4376, 4504, 4151, 4217, 1097, 3315, 1288, 1750, 1593, 2526, 2834, 2906, 1379, 1483, 2387, 780, 2266, 2775, 2619, 3170, 349, 4806, 2555, 2041, 286, 2818, 2861, 30, 2407, 813, 1743, 2286, 2578, 1330, 3001, 2803, 2844, 2643, 2698, 3077, 45, 182, 354, 1336, 2377, 1445, 624, 2074, 1981, 314, 1338, 683, 777, 2291, 3583]. Proportionally [4557]. propres [2387, 3101]. propriété [2387, 2266]. Prototypical [4526]. Providing [3598]. proving [106]. Proximal [3087, 2689, 2346, 2701]. proximately [2172, 2610]. Prüfer [4349, 2046, 3046, 3132, 671]. Pseudo [2525, 3093, 3853, 4338, 3643, 2944, 1385, 3467, 1559, 1583, 1624, 1329, 1548]. pseudo-bases [1548]. Pseudo-Dimensional [3643]. Pseudo-Frobenius [3467]. Pseudo-homogeneous [4338, 2944]. Pseudo-Idempotents [3853]. pseudo-injective [1329]. Pseudo-integrality [2525]. Pseudo-Inverses [3093]. pseudo-Lucas [1583, 1624]. Pseudo-Noetherian [1385]. pseudo-valuation [1559]. Pseudoconvex [4418, 4519, 3390, 2860]. Pseudodifferential [3957]. pseudometric [2302]. pseudopolyhedra [45]. pseudoreflection [2858]. Pseudoumbilical [2453]. pseudovarieties [4773]. Pucci [4545]. pull [4789]. pull-backs [4789]. Pullback [2595]. Pullback-flat [2595]. Pullbacks [1044, 2668]. punctured [3171]. Purdy [1130]. Pure [696, 4117, 3357, 4771, 3313, 3879, 1210, 1429]. pure-high [1210]. Purely [2929, 3601, 2803, 583, 2209]. purifiables [2444]. Pursuit [4421]. pyramidal [924].

QF [2646]. QF- [2646]. Quadratic [3323, 3380, 3245, 3588, 3687, 3689, 1399, 4374, 4047, 2338, 3385, 4296, 4430, 4659, 2435, 4001, 4563, 4239, 3229, 4259, 1461, 2528, 173, 1801, 4756, 1962, 2553, 2779, 2659, 691, 1714, 2491, 63, 2845, 4786, 2700, 2763, 2917, 2988, 3067, 3176, 2602, 472, 1737, 2935, 4380, 3805, 18, 1286, 1937, 3179]. Quadratics [4012]. quadratique [4756]. quadrature [498]. Quadric [4586, 4565, 4733]. quadrilateral [73]. Quandle [3837]. quantification [1320]. quantiles [1049]. Quantitative [3090, 2903]. Quantities [4231, 4004]. Quantization [3211, 3998]. Quantizing [3223]. Quantum [4383, 3650, 4436, 4158, 2979, 3439, 4350, 3928, 4496, 4354, 3112, 4810, 4444, 4025, 3645, 4639, 4072, 4071, 3309, 4487, 4240, 4128, 4054, 3318, 4716, 184]. Quarles [2601]. Quartic [3577, 3217, 212, 851]. Quartics [4637]. Quasi [2893, 4736, 3124, 3891, 1329, 3739, 3191, 3235, 4280, 1118, 1201, 2936, 3716, 686, 4777, 995, 1380, 2655, 393, 665, 2215, 2075, 2337, 1451, 1763, 1143, 2843, 2152, 499, 849, 1541, 4375, 750, 1521, 2494, 187, 1254, 602, 1177, 120, 991, 715]. quasi- [2337]. quasi-analytic [1521]. Quasi-Baer [3739]. quasi-continus [1143]. Quasi-Convergence [3716]. Quasi-copure [4280]. Quasi-differentiable [1201]. quasi-discrete [2843, 2494]. quasi-division [1177]. Quasi-duality [2936]. Quasi-equivalence [4736]. Quasi-Frobenius [686, 393, 602]. quasi-Hausdorff [499]. Quasi-hereditary [2893]. quasi-Hermite [187]. quasi-Hermitian [4777]. quasi-independent [2075]. Quasi-injective [1329, 715]. quasi-injectives [995]. quasi-inner-product [2655]. Quasi-isometry [3891]. Quasi-Measures [3124]. quasi-nonexpansive [1254]. Quasi-Norms [3191]. quasi-radicales [120]. Quasi-split [3235]. Quasi-uniform [1118, 1380, 665, 1451, 1763, 1541, 750]. quasi-uniformities [991]. quasi-uniformity [2215, 2152, 849]. Quasianalytic [4149]. quasiconformal [2521]. Quasicontinuity [3992]. quasicontinuous [4731]. Quasiconvexity [4100]. Quasidiagonal [3863, 3359]. Quasigroups [805, 685, 793, 794, 913]. Quasilinear [4543, 2991]. Quasinilpotency [4376]. quasinilpotent [3064]. quasinormable [2571]. quasinormally [4253]. Quasiradial [3339]. quasisimilarity [1845, 2093, 2312]. quasistrong [1903]. Quasisymmetrically [4017]. Quaternion [4460, 25, 2338, 2108]. Quaternion-like [4460]. Quaternionic [4285, 3161, 3038]. quaternions [2891]. Quelques [2712, 452, 2244, 1987]. Question [3280, 3336, 2571, 4679, 3034, 985, 1661, 2477, 974, 3178]. Questions [3354]. Quillen [2742, 3166]. Quintic [3314, 4704]. Quot [4819]. Quotient [4366, 4533, 1103, 2772, 1304, 1906, 2022, 650, 435, 2210, 948, 668, 1148, 1687, 1894]. quotient-like [668]. Quotients [4495, 3489, 4619, 3571, 4259, 1007, 44, 1081, 131, 2080, 9, 15, 4748, 812, 517, 540, 1155, 2600, 260, 4771, 875, 2458, 1586, 2606, 1282, 1063, 321, 651, 1374].

R [1575]. R. [2351, 4019, 306, 3412, 1109, 738, 1788]. RA [3709]. Rabinowitsch [2917]. Rabinowitz [4768]. Radial [4169, 3012]. Radially [3475, 1427]. Radical [4088, 4558, 1233, 4059, 3648, 2871, 1104, 2548, 1341, 2636, 1290, 2862, 476, 2028, 654, 3020, 2986, 3085, 1256, 4719, 452, 1240, 654]. Radical-free [654]. radicales [120]. Radicals [4367, 3398, 4148, 3959, 563, 2151, 2405]. Radii [2854, 3232, 3723]. Radius [4011, 4064, 3644, 4592, 3115, 1549, 1271]. Rado [316, 4019, 3697, 4761]. Radon [1246, 1419, 3415, 3979, 4526, 4192, 619]. raising [303, 1929]. Ramanujan [3732, 3139, 3704, 3740, 2393, 2327, 519]. Ramification [3399, 3612, 3334]. ramified [3192]. Ramsay [278, 233]. Ramsey [278, 894, 713, 1269, 858, 4817, 1144, 4279, 727, 3943, 4151, 3659]. Randers [3950, 4315, 4052, 3985, 4162]. Random [3809, 4606, 923, 1319, 2267, 872, 843, 845, 1062, 1734, 4747, 4696, 2210, 167, 117, 1914, 2807, 2583, 1337, 302, 2446, 242]. Randomized [3882]. Randomness [4665]. rang [3857]. Range [4621, 3093, 3970, 2118, 4206, 3341, 3174, 3629, 3947, 1834, 2622, 1259, 2984, 3200, 1172, 1157, 1732, 2828, 1227, 3038, 2449, 1133]. Ranges [3232, 3447, 1794]. Rank [3142, 4654, 829, 3579, 3580, 1384, 3551, 3419, 4342, 3982, 4489, 3389, 1002, 3167, 2652, 1379, 2329, 4643, 2754, 2838, 2551, 33, 2817, 185, 2955, 1306, 1485, 2150, 3813, 1205, 2985, 2186, 1364, 631, 3857, 753, 4668, 2540, 2115, 585, 1932, 4702]. Rank-One [3551, 4489]. ranked [2056]. Rankin [4468, 4485]. Ranks [3749, 4291, 2611, 4805]. Rao [3131]. Rasmussen [4129]. rate [1159]. Rates [3945, 4452, 4825]. Ratio [3713, 286, 1865]. Rational [3138, 4641, 4651, 3380, 3867, 3522, 4143, 4667, 3384, 3937, 3402, 3645, 3696, 3921, 4416, 1926, 1968, 1122, 1518, 2576, 1619, 1537, 4812, 3049, 2265, 1823, 2473, 188, 1565, 4818, 417, 434, 1984, 1479, 2044]. Rationality [3340, 4090, 4583, 4700, 610]. Rationally [3555]. Rationals [3140, 2122]. rationnels [3322]. ratios [2122, 4700]. Rav [1295]. Ray [3202, 3719]. raying [4790]. Rayleigh [1148]. Raymond [2477]. Rays [3978]. RCD [4514, 4252]. RD [4250]. RD-spaces [4250]. reaction [3203]. reaction-diffusion [3203]. Real [3230, 3261, 3352, 3141, 3426, 3546, 3654, 3576, 4555, 4442, 3592, 4045, 4372, 3973, 3398, 4168, 4464, 4337, 4207, 3565, 3911, 3580, 4524, 4653, 3130, 4612, 3962, 4193, 4527, 3589, 3866, 4018, 4257, 4329, 3632, 4504, 3463, 3838, 3512, 4586, 4646, 3221, 500, 2427, 2993, 2469, 2569, 2995, 212, 1462, 488, 3184, 141, 1440, 1509, 980, 2659, 2955, 1203, 423, 2281, 2784, 3158, 859, 2677, 3016, 1925, 2785, 2824, 2845, 4786, 2700, 2763, 3071, 168, 4733, 2545]. Real-Analytic [3261, 4442]. real-valued [2427, 1462]. realizability [570]. realizable [1163, 751]. realization [1699]. Realizations [3270, 2247, 4507, 3049]. realize [2826]. Reals [4484, 4694]. Rearrangement [3627, 3127, 1060]. Rearrangement-Invariant [3627]. Rearrangements [4023]. Reciprocal [2225, 158, 1258]. Reciprocals [1800, 1551, 1603, 2517, 302]. reciprocities [201]. Reciprocity [3744, 3010, 3458, 1479]. Réciproque [1258]. Reconstruction [3653, 4618, 949]. Recouvrements [2304]. rectangles [694]. rectangular [631, 753]. Rectangularity [2231]. Recurrence [3843, 4327, 1719, 792]. Recurrences [3208, 143]. Recurrent [3409]. recursive [1811]. redefining [2356]. Reduced [2319, 3648, 2950, 1584, 1762, 3050, 1760, 1647, 969]. Reducibilities [3551]. reducibility [326, 1050]. Reducible [1619, 1537, 2120, 2092]. Reducing [3345]. Reduction [4409, 3416, 3478, 3729, 3529, 2302, 179, 1410, 1410]. reductive [1634]. redundancy [4760]. Ree [1075]. Reeb [3654, 4168, 4329]. réelles [3184]. réels [1509, 2556]. Refined [4530, 4230, 4823]. refinement [1630]. refinements [1942]. reflecting [843]. Reflection [4334, 4320, 2565, 3204]. Reflections [3530, 193]. Reflective [3014]. reflector [1031]. Reflexive [61, 1835, 2646, 925, 2520, 442, 1123, 2877, 1575, 2142, 1229]. reflexivity [3062]. Region [3703, 2370, 2631]. Regions [3430, 2718, 2037, 1716, 4708]. règle [1987]. regression [1205, 1337]. Regular [4305, 4163, 2017, 4607, 1272, 4141, 1562, 2644, 3777, 3384, 3779, 2007, 1857, 1282, 2247, 1374, 2750, 418, 2427, 2402, 1191, 1379, 1082, 211, 25, 2175, 2858, 2954, 1290, 300, 1814, 603, 2242, 4737, 1683, 1154, 2698, 2295, 2131, 4659, 1044, 3082, 607, 1052, 1175, 2112, 1722, 5, 2518, 3803, 1991, 118, 1946, 2831, 93, 2467, 1024, 237]. regular-closed [1683]. Regularity [4354, 4707, 4254, 4395, 3493, 4639, 2631, 3151, 3183, 2069, 1165, 2437, 1097]. Regularized [4622]. Regularizers [1113]. regulating [1193]. Regulator [4055]. Reich [1037]. Reichstein [4503]. Reidemeister [3504, 4435]. Reinhardt [4741]. Reiten [4190, 1695]. rejective [4769]. Related [3903, 4391, 4378, 4613, 3400, 3583, 3311, 4173, 4221, 4303, 1653, 675, 1137, 2133, 3831, 1883, 2158, 3013, 1500, 4725, 1170, 2803, 3065, 2804, 1751, 317, 3759, 2461, 228, 2294, 877, 710, 1130, 821, 1647, 1179, 4815, 1773]. Relation [3701, 4207, 3530, 1790, 2622, 737, 813, 249, 2184, 458, 2403, 3018, 1047, 583]. Relational [3649]. Relations [4036, 4367, 579, 3843, 4714, 4135, 2119, 3374, 4125, 4000, 3403, 1315, 606, 61, 1111, 389, 2488]. Relationship [3748, 4310, 2520, 2400, 2490]. relationships [415, 143]. Relative [4475, 4249, 4672, 3649, 1769, 3768, 2166, 1804, 2644, 2898, 14, 1592]. Relatively [4664]. relativistic [174]. Relativity [4408, 1582]. relator [1430]. relaxation [1930]. Rellich [217, 465]. Remainder [1593]. remainders [1975]. Remark [3621, 4571, 124, 3776, 3581, 254, 3144, 4472, 4449, 1840, 2837, 149, 1804, 304, 1784, 748, 554, 1823, 2341, 162, 1127, 148, 1222, 703, 772, 2719, 1598, 432, 4, 621, 2395, 1368, 2051, 901, 512, 2361, 712, 2567, 731, 202, 1007, 2135]. Remarkable [1821, 327]. Remarks [95, 4477, 2178, 896, 60, 170, 1849, 685, 2201, 812, 931, 4444, 3855, 239, 648, 187, 3086, 4032, 299, 316, 1352, 614, 810, 8, 233, 914, 192, 921, 2030, 411, 566, 329, 2599, 146, 861, 130, 796, 2236, 1175, 807, 91, 289, 3428, 1836, 1859]. Remarque [1466, 1007, 1235, 2135]. Remarques [3322, 3428, 1836, 1859]. Removable [3221]. removal [4760]. remplissage [3328]. Renewal [4620]. renorming [2739, 2076, 2386]. Renyi [1450]. repeated [4719]. repère [3582]. repetitions [800, 2634]. replacement [2137]. replete [2006]. representable [1286, 2479]. Representation [4160, 4386, 4015, 3687, 4423, 3561, 509, 463, 420, 3211, 4121, 1421, 2649, 996, 2031, 2211, 121, 1381, 1558, 2858, 2777, 2926, 789, 2912, 351, 848, 1112, 2475, 756, 1836, 3082, 1863, 466, 3803, 1526, 2211, 1836, 996]. representational [1080]. Representations [3988, 3369, 3863, 3271, 4369, 3440, 3285, 4270, 4497, 3014, 3292, 2503, 4314, 4636, 4256, 2728, 3307, 3675, 3309, 3236, 3135, 3103, 4056, 3182, 1536, 4780, 846, 518, 766, 804, 1196, 1362, 2186, 160, 179, 1998]. Representatives [3474]. Representing [1433, 3803]. Reproducing [4028]. repulsive [847]. Rescaled [3999]. Residual [2084, 2369, 2811, 2052, 2194, 2663]. Residue [3323, 4640, 3724, 3435, 3284, 3278, 2705, 691, 4669, 3050, 4727, 1647, 4661]. residues [656, 394, 1737, 484, 1937]. Resolution [3254, 3009]. Resolutions [3148, 1189]. resolvability [817]. Resolvable [3606]. Resolvent [3427]. Resolving [4249]. resonance [3017]. Resonances [3417]. resonant [3780]. Respect [4615, 3634, 2396, 1760, 1150]. restreintes [1628]. Restricted [282, 4475, 4351, 4532, 10, 3202, 3979, 4528, 4281, 2083, 1215]. Restricting [4020]. Restriction [3766, 4169, 3462, 2697]. restrictions [208, 1381]. Result [3090, 3668, 4432, 2852, 1342, 2213, 149, 1572, 2758, 1373, 684, 1573, 2471, 2309, 148, 2715, 525, 1293, 1914, 2077, 1669, 941, 879, 1412]. Resultant [3852]. Resultants [4050, 4222]. Results [3500, 3156, 4144, 4322, 1241, 1380, 2704, 3793, 165, 1496, 4695, 4772, 1992, 845, 2804, 2325, 2191, 1854, 2991, 1503, 1489, 1422, 1387, 3760, 2787]. retarded [1672]. Retract [4583]. retractable [2769]. retraction [2569]. Retractions [4325]. Retracts [1718, 1907, 2411, 2018]. retrieval [4711]. Return [3970]. Reuleaux [718]. Reverse [3396, 2982]. reversibility [102]. Reversible [4335, 1164]. revisited [2922, 2818, 4707, 3067, 1058]. Revisiting [4809]. Revolution [3521, 4208]. Rham [3362]. Rhodes [4506]. Ribbon [4136]. RIC [4421]. Riccati [1853]. Ricci [2993, 4695, 4464, 3999, 4231, 4329, 4546, 4566, 3406, 3514]. Riemann [3651, 4036, 4263, 1266, 3703, 2487, 637, 3505, 2532, 4388, 3171, 3172, 3328, 4044, 4334, 4371, 4532, 2533, 3672, 4469, 4613, 3165, 3058, 4198, 4816]. Riemannian [2408, 2653, 906, 658, 3283, 1978, 1168, 664, 2321, 3815, 4231, 4580, 2916, 522, 4126, 3585, 78, 4004]. Riesz [1516, 2685, 1437, 2613, 2364, 1789, 2045, 2955, 1111, 3339, 4394, 4482, 4795, 4594, 94, 1879, 4032, 4223]. Riesz-summability [1437]. Right [3686, 1136, 1696, 103, 1318, 2688, 131, 409, 2347, 2508, 1167, 255, 290, 860, 2463, 2728, 3082, 3818, 4769, 754]. Right-minimum [754]. Right-ordered [1136]. Rigid [3751, 2683, 2608]. Rigidity [4693, 4576, 3347]. Rinehart [3895]. Ring [3807, 3107, 3185, 4331, 3210, 3477, 424, 4373, 3130, 3132, 2454, 4190, 4127, 4215, 1103, 2871, 1456, 3060, 1601, 1008, 3152, 44, 2753, 1081, 1889, 907, 2013, 2636, 2177, 393, 9, 15, 2164, 1305, 4087, 1752, 2862, 984, 2761, 1820, 1866, 255, 246, 322, 473, 448, 846, 1155, 260, 1955, 1309, 2463, 3035, 4750, 2028, 1473, 768, 1127, 875, 478, 1310, 2248, 1114, 453, 3162, 213, 417, 1494, 1091, 2974, 940, 834, 1131, 602, 1425, 1526, 452, 754, 2068, 2886, 1614, 2611]. Rings [1103, 4344, 3989, 4365, 4550, 1918, 4201, 1641, 2529, 4349, 3687, 1108, 618, 4204, 3155, 3110, 3751, 4115, 4137, 4515, 4022, 4184, 4335, 3445, 3490, 4118, 4205, 3739, 3628, 517, 3566, 3711, 4428, 4558, 3777, 4445, 4048, 4358, 4148, 4149, 3117, 819, 1910, 3464, 3663, 4323, 3136, 3699, 4324, 3467, 3319, 2095, 2379, 2257, 1755, 2668, 1068, 4821, 2042, 106, 451, 2316, 515, 1457, 2238, 2367, 2570, 1959, 2527, 2415, 2852, 2235, 2815, 3081, 1162, 1396, 649, 870, 1081, 1552, 1379, 778, 972, 2596, 1082, 925, 1341, 2891, 3196, 131, 2772, 1535, 1829, 2856, 599, 1301, 235, 760]. rings [2894, 196, 1304, 493, 4678, 1071, 2706, 1759, 2895, 696, 1418, 1140, 2365, 2242, 2052, 2043, 2194, 812, 2480, 2391, 1721, 1485, 2675, 476, 2911, 1708, 2802, 2160, 2229, 1220, 2840, 2426, 2580, 2710, 2783, 2842, 2878, 424, 3034, 1627, 3814, 2273, 1762, 290, 388, 540, 860, 1154, 2022, 1661, 554, 1847, 2625, 2897, 2930, 2333, 2496, 2825, 2283, 2646, 1751, 539, 1147, 1392, 2015, 2678, 1385, 323, 2535, 1574, 1753, 2846, 1405, 917, 1406, 1674, 2716, 2731, 1586, 1710, 1857, 1957, 1091, 1365, 2764, 3085, 2989, 1760, 1052, 1175, 1504, 1728, 1720, 1722, 2087, 2629, 2328]. rings [2630, 505, 4488, 2934, 1739, 602, 1850, 1178, 273, 67, 49, 128, 93, 2348, 35, 2789, 2936, 154, 2949, 1024, 2791, 686, 2887, 4730]. rises [1323]. Robillard [858]. robin [189]. Robinson [738]. Roelcke [3931]. Rogers [3725, 2327]. Rokhlin [4217]. Rolewicz [3336]. Rolle [1260]. Romberg [498]. Room [494, 629, 209]. Root [3023, 3265, 3010, 3326, 3770, 3913, 3525, 3311, 951, 2799, 2908, 3823]. rooted [2310]. Roots [3090, 3507, 3541, 739, 2855, 141, 2198, 4763, 2900]. Rota [4522]. rotatable [1038]. Rotation [3494, 3238, 3363]. Rotationally [4247]. Rotations [3248]. Rotors [4276, 2599]. rotund [2076, 2386, 1935]. rotundity [2543]. Rough [3379, 3499, 4178, 4355, 3833, 4723]. Roughness [4376]. round [189]. round-robin [189]. Rovnyak [4759]. row [1893, 2140]. Rubel [3053]. Ruelle [4379]. ruin [1734]. Rule [4599, 1987, 1246, 2144, 2038]. Ruled [3230, 4524, 3402, 2755]. Rules [4465]. running [1766]. Russian [1571]. Ryser [1293].

S [1098]. S. [1156, 239, 177, 1669, 3336]. Saddle [3856]. Saddle-Point [3856]. SAGBI [3691]. Salisbury [2970]. same [3828, 286, 302]. Sample [2502, 241, 1670, 1364, 949, 489]. Sampling [3882]. Samuel [868]. Sanderson [360]. Sandwich [2672]. sans [1987, 654]. Sasakian [3864, 4804, 3662]. satisfaction [1810, 1811]. satisfaisant [2972]. satisfied [1386, 3053]. satisfy [1232]. Satisfying [4221, 1074, 793, 119, 2972, 3053, 314]. Sato [3581]. saturatedness [2144]. saturation [1811]. sausage [2854]. Scalar [3950, 3448, 4340, 2114, 4645, 4052, 2835, 1853, 659, 1045]. scale [74]. Scales [3636, 3884, 620, 4195, 3923]. Scaling [4065]. Scattering [3417, 4072, 4071, 3513]. scenery [2583]. Schaefer [4772]. Schatten [3108, 4496]. Schauder [68, 2417, 3388]. schema [4819, 4819]. Scheme [4667, 3762, 907]. schemes [4826, 584]. Scherk [129]. Schild [4408]. Schlömilch [2882]. Schmidt [772, 3350]. Schmüdgen [3130]. Schnirelmann [2026, 3503]. Schönheim [2195, 782]. Schreier [385, 3611]. Schroder [169]. Schrödinger [4420, 4476, 4226, 4330, 3560, 3065, 1851, 4482, 4724, 1945, 4380, 3821, 4223, 3984]. Schubert [3705, 4372, 3761]. Schur [1036, 3159, 3365]. Schütte [4152]. Schützenberger [3793]. Schwarz [4011, 4264, 3905, 3004, 1369, 1544, 4218, 2976, 2109]. Schwerdtfeger [19]. Sci [1317]. score [123, 189, 1961]. scrolls [2560]. SDEs [3798]. search [1795]. secant [1653]. Second [1698, 3884, 2835, 3545, 4478, 2857, 2657, 4271, 4430, 3584, 4593, 3351, 4222, 4223, 3763, 739, 1930, 2794, 1582, 1214, 1012, 2620, 2776, 788, 1166, 3795, 2014, 458, 1955, 1881, 1985, 1964, 2277, 2680, 2012, 1336, 2748, 1638, 1640, 890, 878]. Second-Order [4271, 1698, 4223, 1336, 1640]. Secondary [3939, 3960, 2635].

Sect [1317]. Sectionally [3395, 55, 91]. Sections [4229, 3725, 3343, 3400, 2612, 4744, 2805]. Sectorial [1059]. Sedenions [3546]. Segal [4175]. segments [2459]. Segrè [3910]. Seifert [3411, 3593, 2819, 4480]. Selberg [4468, 4485, 3423, 2115]. selecting [1016]. selection [542]. selectors [3840]. Self [4346, 4552, 1623, 1141, 4065, 3578, 3407, 4462, 4120, 4232, 4092, 2958, 2975, 4638, 2786, 3530, 4596, 35, 1324, 1082, 3043, 4737, 430, 2911, 246, 1540, 2826, 1919, 2372, 4746, 834, 1791]. Self-Adjoint [3530, 2958, 4638, 3043, 1540]. Self-Affine [4065, 4552, 2786]. Self-Avoiding [4092, 4462]. Self-complementary [1141]. Self-converse [1623, 430]. self-dual [1919]. self-embeddings [2372]. Self-Equivalences [3578]. Self-homotopy [4120, 2826]. Self-injective [35, 1082, 2911]. self-intersections [1791]. Self-maps [2975, 4737, 4746]. Self-Similar [3407, 4596]. Selfadjoint [1168, 3341]. selfmaps [1206]. Selmer [4780, 4647]. Semi [3765, 1301, 4330, 3641, 1880, 2081, 4519, 3299, 2350, 342, 3962, 4193, 4126, 4003, 4261, 106, 649, 618, 2463, 1359, 766, 3066, 341, 343, 279, 1524, 2348, 1097]. Semi- [2081, 2350]. Semi-classical [4330, 4261]. Semi-commuting [4519]. semi-connectivity [341]. Semi-Direct [3299]. semi-distributive [3066]. semi-equicontinuity [1097]. semi-group [106]. Semi-Hausdorff [342]. Semi-Infinity [3641]. Semi-invariant [4003]. semi-lattices [343, 279]. semi-normés [1359]. semi-perfect [649]. semi-prime [618, 2463]. semi-principal [2348]. Semi-Riemannian [4126]. Semi-simple [1301, 766, 1524]. Semi-slant [3765]. Semi-Symmetric [3962, 4193]. semiabelian [4737]. Semibounded [2358]. Semiclassical [3987]. semicommuting [1535]. semicontinuity [1689]. semicontinuous [2672]. semiconvex [4734]. Semicritical [2022]. Semicrossed [4185, 4088]. semidefinite [4726, 1128]. Semiderivations [2367, 4204]. Semidirect [3626, 1777, 1954, 2284, 2044]. semidirectly [4773]. Semiflows [3716]. Semigroup [4015, 3410, 3562, 3125, 4209, 1954, 3216, 2978, 1719, 3828, 810, 1922, 1539, 1321, 389, 948, 4608, 1680, 1368, 3803, 2245, 2263]. semigroup-valued [1922]. Semigroups [3901, 4325, 4367, 4572, 2170, 3535, 3329, 4314, 3741, 4150, 4376, 3633, 4626, 3920, 4054, 4220, 3986, 2752, 604, 925, 1302, 1011, 590, 1870, 1167, 1247, 2198, 723, 2380, 3064, 1493, 4773, 2172, 1144, 1147, 4682, 876, 689, 837, 1282, 818, 2374, 346, 1287, 1411, 2610, 776]. semihereditary [1071]. semilattices [2240, 704, 1835, 1885, 1335, 1611, 1727]. Semilinear [4247, 3955, 1640, 1395, 2099, 2373, 2029, 3800, 2603, 1945]. semimodular [3063]. seminorm [4114]. semiperfect [1820, 1278, 2791]. semipermutable [4253]. Semiprime [1829, 2856, 4085, 2897, 4445, 4048, 1068, 2238, 2815, 1396]. Semiprimitive [3661]. Semiprojectivity [3601]. semiregular [2989]. Semiregularity [3871]. semirings [1011, 1321]. Semisimple [4110, 3354, 3333, 4216, 3829]. Semisimplicity [2839]. Semistable [4661]. Semitopological [4626, 4529]. Sendov [2294]. Separability [3128, 4093, 2736, 2674, 2726, 2941, 2849].

Separable [4650, 3357, 4619, 4447, 2293, 3012, 806, 2846, 621, 854, 966, 2849, 1846]. Separate [3692, 2662]. separated [4781, 997]. séparateurs [371]. Separating [2249, 3794, 3083, 3781, 3782, 2058, 2466, 371]. Separation [1320, 609, 3492, 1043, 2443, 977, 1212]. Separators [2707, 3134]. Sequence [4611, 3610, 57, 3570, 4632, 2382, 2000, 2687, 1438, 1483, 1581, 2704, 946, 1608, 2472, 1792, 1924, 123, 1996, 378, 1841, 232, 1369, 1544, 634, 1133, 4783, 2543, 158]. Sequences [1323, 3637, 3545, 3025, 4225, 1163, 3111, 4084, 4183, 3890, 4209, 692, 319, 3570, 3943, 3899, 4031, 2521, 3764, 4824, 1615, 2122, 1690, 38, 616, 1308, 570, 815, 162, 1277, 796, 189, 2431, 1761, 1908, 1695, 2445, 1446, 1375, 421, 4688, 2217]. Sequential [4040, 3849, 547, 1302, 1038]. sequentially [1344, 1023]. Series [3468, 4058, 3902, 3558, 4013, 4349, 3844, 4044, 3113, 4274, 3371, 4072, 4071, 3942, 3451, 3679, 3424, 4197, 3743, 2174, 2042, 1008, 1437, 1813, 2436, 308, 2487, 2772, 400, 1558, 503, 1466, 2894, 1271, 2450, 1626, 1807, 1100, 287, 375, 399, 1470, 1190, 1662, 1259, 1831, 270, 71, 164, 57, 1836, 2882, 3161, 3005, 2602, 1970, 2114, 180, 1280, 2395, 640, 819, 2241, 1991, 1490, 1648, 2115, 1946, 543, 1507, 2886, 3055, 2936, 2790, 2042, 3399, 1836]. Serre [3806, 2472, 2106]. Set [3394, 3352, 3092, 4133, 4649, 3836, 3854, 4562, 3570, 3062, 2224, 690, 1353, 250, 980, 1356, 1922, 2144, 1441, 2274, 2413, 249, 555, 1404, 448, 1360, 1292, 389, 1363, 1252, 2701, 666, 717, 3053, 2209, 482, 346, 1230, 85]. set-valued [2701]. Sets [4080, 3207, 4457, 3924, 4107, 308, 3091, 4015, 4017, 3951, 3094, 841, 3548, 4065, 4290, 4441, 4555, 4183, 4251, 3187, 4023, 4310, 4557, 1994, 3224, 3491, 3298, 3855, 3674, 3344, 3305, 3386, 3568, 3979, 3163, 4002, 4030, 4173, 4584, 4152, 3944, 1257, 3466, 3555, 1288, 1750, 2498, 2174, 975, 1105, 1986, 2085, 1909, 749, 416, 2836, 1615, 1533, 404, 410, 2304, 1799, 457, 4694, 1898, 2293, 2774, 882, 1718, 1270, 1605, 258, 1326, 2705, 1193, 2156, 1883, 2223, 2818, 2623, 2075, 2389, 751, 2286, 672, 932, 3776, 1709, 2676, 1553, 2821, 1040, 1014, 985, 1924, 2301, 815, 1249]. sets [647, 886, 1754, 569, 247, 182, 862, 10, 363, 2717, 2604, 2900, 1706, 2314, 3802, 1157, 3781, 3782, 1875, 598, 1445, 797, 877, 710, 4734, 1944, 1822, 2542, 1491, 867, 1956, 1545, 2086, 4815, 3054, 1668, 1876, 1961]. Setting [3466, 4698, 3801, 4668, 2587]. seven [580, 64, 2865, 1513]. Several [4456, 4550, 4384, 3745, 3759, 3630, 1030, 2271, 3041, 2366, 4753, 573, 3834, 537, 572, 2025, 2390, 2539, 1840, 1833, 1917]. Severi [4539]. Sextic [4624, 4140]. Sextuples [3773]. SF-rings [2789]. Sha [3960]. shadows [4784]. Shahidi [3235]. Shaken [3725]. Shape [3911, 2340]. Shaped [3430, 36]. shapedness [3829]. Shared [4067, 3392]. Sharp [4510, 4421, 4710, 4635, 4025, 4189, 3961, 2234]. sharpened [4793, 1197]. Sharpness [3156]. sheaf [262, 2219, 1178, 1526]. Sheaves [4390, 4446, 214, 848]. shell [4742, 2746]. Sherk [2247]. Sherman [1047]. Sheu [4054]. Shift [3150, 4657, 3112, 2308, 1446]. Shifted [4560, 4410]. Shifts [4035, 4613, 3252, 2511, 4670]. Shih [3184]. Shimi [1377]. Shimura [4653]. Shokurov [4143]. Short [4079, 3470, 4457, 2069, 4146, 4050, 3596, 4172, 4217, 3862, 2430, 4722, 1244, 1372, 1126, 3018]. shorter [643]. shrinking [1981]. Shuffle [4104]. Shuffling [2864]. Shum [1697]. Side [3484]. Sided [3327, 2982, 322, 830, 768, 2161, 2789, 4506]. sides [73]. Sidki [4788]. Sidon [2498, 4310, 4557, 4649]. Siegel [2971]. Sierpiński [3695, 2586]. Sieve [4560, 2919]. Sign [4012, 2228, 4301, 2278]. Sign-changing [4301]. sign-stability [2278]. signature [4680]. Silver [3676]. silvered [2462]. Similar [3407, 1033, 4596]. Similarité [2600]. Similarity [2888, 3909, 2804, 2600]. Similitude [4432]. Simple [4420, 3024, 3670, 2401, 3197, 4272, 3507, 3132, 4051, 3449, 3876, 3364, 3602, 3146, 3060, 2979, 2102, 1889, 4793, 1301, 2276, 732, 3771, 696, 2010, 2802, 2929, 1762, 290, 766, 804, 1196, 2457, 247, 323, 2846, 1019, 1775, 2412, 1524, 3288]. Simplectic [3222]. simplement [3259]. simpler [880]. simples [1889, 3288]. simplex [455]. Simplexes [4267]. Simplices [3952, 3786, 2288]. Simplicial [4385, 4554, 3562, 1189, 3538, 3916, 2126]. Simplicité [2010]. Simplicity [4137, 2561, 4125, 2010]. Simplified [4059]. simply [1947, 4680]. Simpson [4390]. simultanée [3219]. Simultaneous [2577, 3304, 3568, 2564, 1428, 1218, 251, 1723, 587, 3219]. SIN [4413]. Single [4651]. Single-Distance [4651]. Singquandle [4784]. Singular [4304, 3379, 3499, 4605, 4178, 4425, 658, 3560, 1484, 3593, 3143, 3295, 4390, 4625, 4070, 4526, 4398, 3697, 1002, 3076, 2443, 4784, 3011, 3027, 758, 721, 1913, 507, 2875, 2778, 762, 2758, 983, 1493, 2358, 934, 2029, 794, 948, 1115, 4721, 4724, 1387, 1198, 4762, 4723, 1093, 1410, 826]. singularités [3259, 1410]. Singularities [3218, 3254, 3146, 3966, 3259, 2579, 3083, 1830, 2734, 1410]. Singularity [3874, 2296]. singularly [2099, 2073]. singuliers [1410]. Sitter [4720]. Six [3209, 580, 134, 332]. sixth [4818]. Size [4469, 4096, 489]. Sizes [3395, 4023, 4186, 2054, 2144, 2945]. Sjögren [3344]. Skan [3657, 707]. Skein [3044]. Skeletons [3580]. sketches [2651]. Skew [4286, 4137, 4410, 4148, 4565, 2273, 2087, 2629, 1475, 3040, 2442]. skew-symmetric [3040]. Slab [4356]. slant [3765]. Slater [2492]. Slice [2932]. Slices [4084, 3224]. slimming [927]. Slope [3610]. Slow [4636]. Smale [3592]. Small [4105, 173, 4574, 3689, 2774, 3708, 2693, 4557, 4252, 4696, 4068, 4318, 2116, 605, 1098, 2293, 2820, 2250, 4727, 569, 809]. Smaller [3940]. Smallest [4013, 4005]. Smirnov [253, 3306]. Smith [2165, 2077]. Smooth [3483, 4637, 3719, 3141, 4112, 3045, 3974, 2534, 4470, 3894, 3802, 4197, 4824, 2739, 1123, 1654, 3016, 2560, 589, 4765, 1722, 1229, 4816, 2049]. Smoothability [1774]. Smoothing [1010, 3013, 2192]. Smoothness [1716, 1560, 4819, 1850]. snake [1874]. SNC [4644]. Sobolev [1394, 2319, 3059, 4793, 2597, 4271, 3537, 4594, 1488, 3634, 4549]. Sobolev-type [1488]. societies [752]. socle [3020]. socles [2440]. Socratic [218]. soft [1589]. solenoidal [2740]. solid [1712, 2713]. Solitar [2107]. Solitons [4546, 4566, 4695, 4804]. Soluble [2688, 1825, 2533, 982, 2226, 1041, 1408, 2442]. soluble-by-finite [2442]. Solution [577, 3827, 1403, 4391, 3874, 3920, 3240, 356, 1083, 758, 732, 790, 1856, 1236, 1723, 2315, 1448, 707]. Solutions [4419, 2443, 58, 4438, 4478, 3427, 4247, 3927, 4016, 701, 3590, 4268, 3642, 4516, 3407, 3972, 3973, 3738, 3995, 3955, 4024, 3657, 3226, 4625, 4068, 4318, 3304, 3963, 4448, 4195, 4543, 4301, 4241, 2400, 3923, 4454, 4588, 328, 1912, 1716, 902, 924, 2211, 1664, 565, 1844, 1242, 1622, 1300, 787, 157, 1214, 2722, 2928, 200, 4732, 3835, 398, 3815, 459, 1881, 1985, 1042, 1333, 1859, 2991, 472, 3780, 887, 2202, 4746, 466, 2116, 1672, 2947, 4762, 1945, 824, 667, 1199, 3203, 1873, 605, 856, 299, 2545, 1027]. Solvability [3398, 3847, 4802, 1152, 3017, 1854]. Solvable [3223, 1866, 3511, 3528, 1659, 2823, 3071, 3201, 2074, 1475, 1739, 2590]. solving [2660, 1358]. Some [316, 2685, 527, 3732, 952, 3180, 557, 3138, 415, 1241, 515, 628, 3266, 2738, 1032, 227, 2155, 614, 96, 2549, 272, 1438, 1724, 1380, 905, 1642, 1826, 810, 811, 38, 50, 3905, 1465, 3793, 1765, 165, 3354, 1496, 4695, 4772, 8, 60, 233, 957, 914, 192, 921, 1883, 185, 2455, 3562, 2722, 127, 2215, 2030, 4355, 1560, 411, 566, 3250, 4607, 3656, 329, 845, 4519, 2599, 504, 814, 4144, 146, 3099, 3872, 2200, 3299, 1360, 4339, 2762, 4539, 2325, 861, 130, 3855, 2393, 310, 1966, 323, 317, 521, 2191, 3694, 267, 4125, 2340, 2236, 3216, 2322, 4804, 817]. Some [1503, 4194, 852, 444, 1631, 228, 2363, 1049, 1175, 807, 395, 1722, 1422, 1149, 4322, 4004, 1387, 1283, 2934, 1239, 1255, 1525, 890, 203, 1735, 4606, 3390, 4596, 110, 4032, 289, 4735, 4261, 3392, 2414, 1952, 2257, 2037, 2925, 2097, 296, 2364, 347, 880, 1243, 487, 915, 253, 467, 2013, 1987, 1165, 13, 221, 898, 3772, 408, 4736, 2480, 674, 200, 1500, 129, 2783, 2842, 3432, 2643, 27, 4750, 1697, 2492, 1543, 3800, 1726, 4670, 1957, 1322, 1091, 2933, 2563, 406, 2441, 2584, 2665, 5, 1949, 892, 2947, 622, 2449, 1179, 4154, 2684, 2787, 2589, 111, 1979, 2712, 452]. Some [231]. somewhat [2171]. somme [1620, 129]. Sommes [1628, 1812, 2182, 4756]. sont [2459, 1889]. soon [853]. sorting [1976]. sous [2444, 2556, 1950, 2733]. sous-espaces [2556]. sous-groupe [2733]. sous-groupes [2444, 1950]. Souslin [1426]. Space [4262, 3574, 3425, 4650, 3426, 4225, 3735, 3654, 4603, 3669, 3280, 787, 3578, 4229, 3224, 4273, 4337, 3297, 3126, 3565, 3357, 3911, 4089, 4622, 4524, 4446, 3629, 3962, 4193, 4099, 3589, 3866, 4018, 4630, 3102, 4645, 3662, 3118, 3700, 4162, 2685, 2299, 2427, 2993, 2398, 4720, 2288, 1580, 2593, 4663, 2925, 2510, 2364, 1767, 2617, 1258, 2997, 1765, 1931, 1870, 872, 2621, 1852, 2658, 442, 4165, 2721, 789, 1123, 4443, 1072, 2386, 2819, 4812, 791, 931, 2321, 2879, 2863, 1273, 2784, 3158, 2744, 2696, 997, 1251, 1347, 804, 1249, 4394, 1541, 2346, 2713, 1276, 3779, 2785, 1042, 546, 1252]. space [750, 341, 692, 1948, 2108, 607, 2732, 2884, 2314, 2865, 1935, 1495, 1934, 2518, 1568, 2096, 3839, 822, 1575, 2142, 1255, 2449, 2960, 1448, 2159, 3820, 2608, 2831, 969, 1513, 1229, 1900, 2904]. Space-time [4273, 1251, 1347]. Spaces [4490, 3180, 4010, 3516, 4036, 4156, 3261, 3483, 4325, 3531, 4459, 4402, 3557, 4366, 4651, 3452, 3105, 4202, 4478, 4551, 3408, 2306, 3653, 3231, 3639, 3746, 3686, 4350, 3273, 4599, 3750, 3769, 1070, 1207, 4513, 4163, 3444, 4164, 4353, 4495, 4084, 4440, 4165, 3773, 4333, 3692, 3706, 3971, 4514, 3536, 4166, 3506, 3124, 4355, 4534, 3932, 4250, 4271, 3607, 574, 4655, 4390, 3954, 3656, 3539, 3892, 4209, 3127, 3339, 4252, 4169, 3854, 4577, 4145, 3956, 3914, 4122, 4579, 4339, 4622, 3343, 4147, 3893, 4375, 4213, 4629, 3523, 3938, 3526, 4594, 4237, 4192, 3980, 4028, 4645, 3480, 4626, 3570, 3585, 3943]. Spaces [4240, 3388, 3571, 3677, 3600, 4006, 4595, 3616, 3659, 4565, 3881, 4343, 4364, 4451, 3742, 4549, 4567, 3965, 4400, 3947, 1394, 2005, 1497, 1774, 1060, 2546, 3786, 2110, 2105, 234, 547, 2669, 1745, 2318, 1449, 1480, 628, 4801, 1498, 1267, 1436, 2571, 2614, 2964, 3009, 1758, 1033, 1428, 1939, 1417, 1793, 2381, 1855, 1380, 1289, 4793, 1381, 3810, 1244, 802, 1302, 2874, 2293, 4781, 2673, 1988, 2190, 2392, 1765, 4681, 2938, 196, 2045, 2420, 1678, 2139, 3012, 2619, 2800, 3794, 529, 761, 1165, 2655, 2575, 349, 68, 166, 978, 2597, 1746, 2282, 665, 3171, 2216, 3045, 2230, 1067, 2183, 2001]. spaces [2302, 1804, 1530, 1218, 4749, 4736, 2055, 2298, 2006, 2215, 1391, 674, 2407, 1560, 1451, 1519, 2160, 1538, 1802, 630, 2759, 4799, 4738, 2709, 3000, 1763, 755, 1344, 2056, 2940, 2724, 1683, 623, 1404, 2078, 1684, 2417, 1520, 2432, 1792, 2844, 2171, 2152, 2470, 2264, 1359, 4771, 2881, 1346, 3002, 3042, 518, 2350, 3755, 1563, 849, 806, 1431, 1736, 3019, 1277, 2943, 2534, 1757, 1966, 357, 1046, 2082, 4795, 1090, 342, 362, 4822, 2648, 695, 1895, 3177, 1864, 2101, 1174, 866, 865, 1636, 2461, 2428, 3817, 4703, 2973, 432, 963, 911, 1053, 1993, 1149, 2445, 4798, 1118, 1822, 1369, 1544]. spaces [941, 312, 2051, 1569, 2830, 4751, 1821, 2231, 2809, 4761, 4783, 2543, 1374, 2867, 2063, 1096, 1548, 1780, 1808, 1688, 2468, 2361, 289, 3072, 1592, 2447, 2810, 1355, 1846, 1929]. Spacings [3754]. span [2148]. spanning [1487, 900, 1492, 774]. Sparse [4526, 4584, 2423]. Spatial [4312, 3297, 3404]. Special [3487, 2580, 4653, 3541, 4807, 155, 1918, 347, 1438, 1483, 2455, 1219, 1856, 281, 4750, 822, 273, 969, 1548]. Specialisation [4236]. Specializations [3303]. specific [2400]. specified [2003, 592]. Spectra [4575, 4379, 4787, 3027, 1675, 2141, 2518]. Spectral [4458, 4758, 4181, 3610, 3157, 4208, 4523, 4592, 3239, 1549, 2754, 2472, 679, 1174, 2958, 4596, 2063, 3020]. spectrales [3020]. Spectrality [4601]. Spectrally [3413]. Spectrum [3557, 4476, 3684, 3521, 4409, 4097, 3661, 3313, 3513, 4606, 2452, 2576, 1343, 2843, 1502, 2289, 2494, 2073]. Speculations [1834]. Sperner [828, 3162, 1130]. Sphere [4084, 4208, 3115, 3144, 4634, 3966, 2469, 2569, 2997, 663, 2275, 2969, 427, 580, 2167]. Spheres [4247, 4633, 4267, 3722, 4535, 3345, 4054, 1352, 4739, 693, 2874, 1303, 3171, 2453, 4516, 4480, 1736, 2536, 2604, 2718, 2885, 1862, 331, 330, 332, 1973, 2874]. Spherical [2523, 4790, 4512, 3578, 1744, 3126, 4645, 3145, 3311, 4742, 2433]. Spin [4530, 588]. spines [1223]. spinor [1347]. spinors [2721]. spiny [594, 465]. spiral [1241]. spiral-like [1241]. Spiralling [242, 167]. Splicing [1702]. Spline [4623, 2048]. splines [1702]. Split [3331, 2003, 4737, 3235, 1785, 853]. Splitting [3033, 3610, 4410, 4093, 2249, 3831, 2472, 1872, 1432, 514, 657]. Spreading [3750, 4053]. Spreads [669, 2137]. Springer [4386]. Sprott [722]. spurious [1016]. Square [3683, 3326, 3561, 3096, 3089, 3133, 3116, 1647, 386, 2474, 1583, 1468, 813, 1073, 2198, 562, 373, 259, 4074]. square-free [562, 4074]. Square-reduced [1647]. Squared [4492]. squarefree [336, 2285]. Squares [3545, 4556, 3994, 3360, 2121, 1681, 1163, 1342, 1400, 1152, 1940, 129, 596, 862, 629, 2354, 965, 155, 2269]. Squashed [3741]. squences [50]. Stability [288, 1242, 3534, 3197, 2343, 3031, 4604, 1331, 4340, 4341, 4359, 3662, 376, 4381, 3986, 4453, 2504, 309, 2619, 2303, 2844, 459, 2645, 1564, 1915, 3760, 1199, 2278]. stabilizer [2129]. Stabilizers [3528, 1416]. Stable [4326, 3520, 3172, 4616, 3015, 3129, 2287, 4357, 3374, 1710, 4342, 3420, 4486, 2540, 3481, 943, 1032, 2622, 1827, 2184, 2232, 2644, 3002, 863]. Stably [2891, 4302, 1892, 1850]. Stackings [4396]. Stage [4297, 2703]. Stage-Structured [4297, 2703]. Stalking [1426]. stalks [1641]. Standard [4634, 2954, 1196, 2383]. Stanley [4219]. Star [3829, 3430, 36]. Star-Shaped [3430, 36]. Star-shapedness [3829]. starlike [2691]. stars [1269, 1625]. Stasheff [3954]. State [3357, 4448, 4241, 4419, 4762]. States [4496, 4072, 4071, 1121]. static [1427]. Stationary [3609, 3674, 1921]. statistic [722]. Statistical [3684, 2396, 1338, 231, 203]. statistics [253, 504, 741, 1364, 1609, 1049, 2990, 612]. Steady [4419]. Steady-state [4419]. Steenrod [2332, 2290]. Steepest [1152]. Stefan [2631]. Stein [4424, 4433]. Steiner [431, 98, 709, 666]. Steinness [4471]. Steklov [4574]. Step [3316]. Stepanoff [879]. Stephenson [4542]. Stepping [3635]. Stepping-Stone [3635]. steps [676, 537, 220, 454]. Stiefel [4366, 2008, 2106]. Stieltjes [779, 4800, 1809, 4592]. Stirling [99]. Stochastic [1639, 4312, 4298, 3404, 2351, 1867, 2385, 33, 116, 75, 1635, 936, 3800, 1630, 469]. Stokes [3407]. Stolz [3536, 932]. Stone [3635, 2331, 39, 1302, 956, 607, 798]. Stonean [2281]. stopping [1958]. straight [2189, 1723]. Strata [4466, 4676]. Strategies [4053]. Stratification [4237]. Stratifications [4307, 2566]. Stratified [3893]. stream [1766]. Strengthening [4272]. strict [328, 1005]. Strictly [3854, 3630, 3763, 2860, 4481]. string [1844]. strings [1676, 2988]. Strip [3703]. stripes [1269]. strips [2030]. Strong [4158, 2815, 4326, 1557, 3412, 4514, 3112, 1903, 2359, 3228, 4433, 3423, 3647, 1256, 1021, 4638, 1774, 2813, 2834, 1656, 1481, 905, 2396, 4808, 632, 375, 374, 799, 1879]. Strongly [4455, 3843, 3296, 4121, 3361, 4631, 3658, 4098, 1672, 3364, 3716, 275, 2595, 2953, 1700, 2709, 4685, 1277, 1175, 2608]. structural [412]. Structure [2095, 3393, 3654, 3845, 4042, 4604, 3609, 4408, 4168, 3297, 3854, 3551, 3645, 453, 3589, 3866, 4018, 4030, 3570, 2765, 4586, 4220, 4507, 4701, 1712, 2138, 1620, 178, 3152, 136, 165, 1303, 1304, 294, 1936, 664, 1074, 2417, 1347, 851, 2142, 1888, 49, 1370]. Structured [4297, 2703]. Structures [3352, 4061, 3538, 4498, 4644, 4099, 3616, 3390, 3649, 2751, 248, 488, 906, 658, 2177, 665, 958, 1101, 2914, 1355]. Study [4295]. Studying [3207]. Sturm [56, 977, 3027, 3874, 2505]. Sub [522, 3835]. sub-Laplacians [3835]. Sub-manifolds [522]. Subadditivity [4082]. Subaffine [584]. Subalgebra [3481, 766, 1332, 999, 729]. Subalgebras [4383, 3546, 3279, 3410, 4292, 4294, 3375, 3875, 1276, 2846]. Subbarao [1124, 2077, 470]. subbundles [3167]. Subcartesian [4599, 3893, 3600, 3779]. Subcategories [4249, 1367]. subcategory [547]. subclass [1233]. Subcritically [4424]. Subderivatives [3087]. Subdiagonal [2992].

Subdifferential [3151, 3030, 1689]. Subdifferentials [3370, 3465]. Subdirectly [870, 1011]. Subdivisions [3916]. Subelliptic [3240]. Subequalizers [746, 1069]. subexponential [3075]. Subfactors [3650, 3327, 3361, 2992]. subfields [2176, 859]. subfoliations [2635]. subgradient [2689, 2346, 2701]. subgraphs [292, 236, 305]. Subgroup [2736, 3128, 4591, 4196, 3349, 2002, 1191, 2726, 2145, 633, 2733, 1388]. Subgroups [4345, 3275, 3442, 3128, 4253, 3331, 4666, 3312, 4489, 4174, 1134, 1694, 2994, 1210, 2060, 2180, 2178, 2529, 954, 2857, 1440, 842, 1700, 4788, 2052, 2043, 2194, 2967, 2741, 684, 591, 2009, 2233, 1041, 2383, 425, 568, 3039, 990, 2589]. Subhankulov [1525]. Subharmonic [2757, 2616]. Sublaplacian [4081]. Sublattices [3850, 1281, 265]. sublinear [3763]. Subloops [3210, 1687]. Submanifold [3939]. Submanifolds [2798, 3962, 4193, 3662, 3603, 3765, 2469, 2569, 4704]. Submersions [4003]. submodule [762]. Submodules [4280, 3422, 1376, 2967, 2151, 2458]. Subnormal [4035, 1845, 2052]. Suborbit [3393]. Subordinacy [3027]. Subordinate [3961, 488, 906]. subplanes [225]. Subproducts [4727]. subretracts [2580]. subring [2013, 1299]. subrings [2164]. subsemigroups [1729]. subsequence [2690]. subsequences [1561, 1819]. subset [2011]. Subsets [3262, 3841, 4371, 4089, 625, 4776, 2293, 2799, 1070, 1207, 662, 2879, 623, 2713, 2727, 373, 666, 2765, 2446]. Subsocles [2741]. Subspace [3093, 3993, 3677, 3772, 2877, 648, 4783]. Subspaces [3261, 3948, 4143, 3129, 3387, 4032, 2392, 2006, 2639, 2470, 885, 2824, 1757, 2243, 2932, 2732, 2884, 1422]. substantial [1411]. Substitution [3313]. substructures [2108]. subtournaments [275, 335, 1629]. subtrees [1858, 744]. Subvarieties [4602]. Successive [3723, 40, 3, 833]. Sucheston [1573]. Sufficient [2362, 3449, 3896, 4802, 1679, 817, 612, 3784]. suite [2079]. suites [2244]. Sullivan [1566]. Sum [3619, 4062, 3097, 3869, 4558, 3215, 3873, 3525, 4032, 4324, 716, 4729, 1620, 1551, 1268, 1606, 2926, 829, 2127, 608, 631, 159, 753, 797, 964, 2959, 1058, 2667]. sum-free [797]. Sum-Of-Digits [3097]. Summability [983, 1819, 1657, 2000, 620, 1135, 1413, 1437, 1481, 1790, 1813, 2613, 905, 2396, 1905, 1982, 300, 1100, 287, 375, 399, 1486, 1662, 1561, 162, 374, 1145, 1748, 2191, 1322, 1706, 2322, 2806, 640, 1879, 1946, 512]. summable [2487, 400, 1558, 1831, 1277]. summands [2282, 1892]. Summation [3679, 164, 293, 1638, 1507]. summations [573, 572]. summing [1213, 2649, 1928]. Sums [3207, 3123, 4553, 3410, 3209, 2323, 4114, 3355, 3563, 4251, 627, 3994, 4274, 2360, 4294, 3343, 4613, 3252, 3360, 3977, 4321, 3983, 1618, 626, 158, 2269, 1319, 4722, 313, 2133, 2261, 637, 1342, 882, 4756, 1270, 2909, 1594, 2515, 2300, 1734, 3034, 2559, 3050, 2134, 1147, 795, 160, 4818, 1312, 1875, 1776, 1093, 1946, 256, 2479, 2591, 1628]. Sumsets [4734]. sun [1353]. Super [4015, 3564, 3404, 4380, 1798, 2965]. Super-Brownian [3404, 2965]. super-decomposable [1798]. super-décomposables [1798]. Super-Frames [3564]. Super-quadratic [4380]. superadditive [2104].

Superalgebras [3265, 3303]. Superconductivity [4313]. supercritical [2777, 4696]. Supercuspidal [3292, 3135]. Supergroups [4113, 4182, 4207]. superharmonic [2489]. Superlinear [3884, 4241, 2776, 215, 2012]. superposition [184]. superprocess [2777]. superrecurrence [2530]. superreflexive [2449]. Superreflexivity [4192]. Supersingular [3282, 3289]. supersolvable [2493]. supersymmetric [4716]. Support [1402, 839, 3059, 1921, 2314]. Supported [3945, 4704, 2719]. supporting [2741]. Supports [4298]. suppressible [928]. suppression [1707]. supramenable [4777]. Sure [4226]. Surface [3902, 3381, 3855, 3462, 3308, 3919, 3550, 85]. Surfaces [3230, 3785, 4243, 3733, 4036, 3367, 3521, 4334, 4371, 4443, 3710, 4538, 4392, 3579, 4667, 3371, 4235, 3416, 3632, 3552, 3165, 3312, 3554, 3316, 4198, 4162, 382, 4263, 1663, 2175, 4813, 2422, 3172, 2453, 2533, 3198, 2560, 2805, 1433, 4807, 3822, 3058]. Surgeries [3554, 3880, 4435, 4697]. Surgery [3668, 4116, 3851, 4662, 3880, 4075]. Surjective [4715, 2908, 3787, 3792]. Surjectivity [3271, 3440]. survey [1190]. Suslin [3458]. Suslinian [3455, 3846]. suspension [2096, 3820]. suspensions [1211, 456]. Sverdrup [1205]. sweeping [2834]. switching [2303]. Sylow [2180, 2529, 4666]. Sylvester [381, 1036]. Symbol [4643]. symbols [3042, 3176]. symétrique [4087]. symétriques [827]. Symmetric [4247, 3654, 4530, 3625, 3474, 4440, 3475, 2335, 3726, 4615, 698, 3660, 3962, 3144, 3570, 3585, 4240, 4656, 4261, 742, 4720, 2633, 4755, 3826, 2506, 1733, 4087, 4677, 1220, 1617, 2321, 62, 859, 2971, 2915, 1966, 645, 916, 1077, 2082, 4193, 1116, 425, 2106, 3040, 82, 1427]. Symmetrically [3627, 2650]. Symmetries [3368, 3722, 4487, 4813, 2664, 3049, 4430]. symmetrization [4815]. symmetrized [4758]. Symmetry [3310, 2816, 1384, 2561]. Symplectic [3949, 3474, 3895, 4644, 4680, 4692, 4752, 1221, 2886]. symplectique [3222]. Synchronous [1164]. synthesis [679, 2063]. System [3641, 3973, 3997, 3859, 3311, 4301, 4241, 3923, 4453, 4588, 1166, 3061, 2457, 2253, 666, 3803, 2947]. Systems [4620, 3265, 3273, 4265, 3443, 3444, 3868, 4270, 3385, 4255, 4615, 3596, 3386, 4540, 1078, 4026, 4320, 3404, 3964, 3313, 4154, 1873, 3466, 4057, 2057, 1033, 356, 2797, 2099, 2206, 105, 1164, 4781, 2799, 2908, 2777, 2303, 1529, 2693, 1331, 3015, 1015, 2698, 3050, 140, 588, 271, 1564, 2222, 709, 2603, 2363, 1130, 1647, 1491, 2053, 1409, 3823]. Szego [525, 3858, 4282]. Szekeres [84, 545]. Szpilrajn [3523].

T [3583]. T. [2970]. Table [4106]. Tableaux [3949, 4493]. tails [2790]. Takagi [2422]. Take [3734]. Takens [2457]. taking [2952]. Talenti [810, 1368]. Tamari [3575]. Tanaka [4337]. Tangent [1415, 3213, 1653, 3766, 488, 906, 294, 1168, 372, 664, 1862]. Tangential [3894]. Tangles [3356]. Tannakian [4459]. targets [4671]. Tarry [107]. Tate [2632, 4669, 2403, 3581, 2297]. Tate-conjecture [2403]. Tauberian [1413, 1135, 1481, 1813, 2613, 4739, 4206, 2806, 3365]. Tautological [3701]. Taylor [300, 1662, 4670]. Tchebychev [76]. Technique [3207]. Techniques [3160, 2620, 1502]. Ted [3717]. Teichmüller [1804, 1592]. telephone [973]. telescope [1471]. Telescoping [4197]. Teller [4716]. tempered [2296]. Tension [3919]. Tensor [1435, 3561, 4464, 4231, 556, 4329, 3675, 2073, 2993, 1536, 311, 2877, 1074, 2625, 935, 756, 3817, 1977, 2765, 1022]. Tensors [4235, 1384]. tent [4703]. tenth [65]. Term [3104, 3624, 3487, 2942, 3066, 336, 877]. Terminal [4441]. Terms [4824, 4604, 4337, 4617, 3062, 2993, 410, 1403, 2822, 4394, 158]. Ternary [4259, 749, 1955]. Terquem [1264]. tessellations [2981]. test [1034, 241, 530, 1837, 1046, 962]. Testing [4306]. Tests [4497, 74, 983, 1055, 855]. Tetrahedra [3342, 377]. th [1579, 951, 2774, 1399, 3930, 1923, 938, 1345]. th-order [938]. Their [3670, 4182, 3248, 4337, 3645, 3731, 4035, 2994, 2978, 1939, 2364, 3951, 4227, 1746, 4113, 4115, 2927, 573, 372, 4810, 540, 2645, 243, 267, 572, 2458, 4668, 2605, 1893, 2140, 302, 2400, 4600, 4609, 1257]. theme [904, 2572, 1555]. Theorem [3990, 4510, 3744, 4660, 4570, 4585, 3825, 4132, 3092, 4160, 3653, 4614, 3687, 3970, 4352, 4388, 4590, 4334, 3707, 3506, 4443, 4250, 4427, 4272, 3490, 4066, 3934, 4295, 3777, 4429, 3003, 3523, 4542, 3462, 3115, 3420, 4149, 4527, 4028, 3527, 3646, 3348, 3943, 4152, 4217, 4638, 1735, 4451, 4221, 3437, 278, 316, 894, 839, 2751, 1578, 2057, 1454, 2031, 2138, 1005, 1030, 2037, 2509, 2065, 381, 359, 976, 977, 1457, 2834, 2527, 4796, 724, 2873, 1377, 2436, 2613, 1758, 73, 307, 422, 2549, 2329, 2197, 576, 1244, 2550, 637, 217, 465, 713, 2690, 2385, 1371, 945, 1138, 1302, 1258, 946]. theorem [825, 1905, 1390, 2266, 507, 1440, 2072, 1382, 697, 1120, 635, 361, 956, 2170, 233, 979, 1161, 2489, 2185, 1182, 2212, 1468, 1217, 659, 1469, 426, 2183, 1936, 1247, 883, 1783, 1124, 2332, 1036, 268, 244, 1037, 355, 1357, 2819, 726, 908, 552, 1126, 755, 3173, 2313, 1248, 1635, 2711, 1972, 1321, 1553, 1260, 1874, 1195, 1762, 4823, 1015, 934, 4817, 2371, 591, 229, 219, 1806, 4771, 1221, 2038, 230, 2226, 1691, 2601, 655, 1686, 2279, 806, 269, 1026, 2825, 1173, 1311, 1670, 593, 32, 1222, 462, 937, 2340, 51, 2972, 2493, 629, 122, 135, 1674, 850, 1630, 1704]. theorem [1908, 1996, 2538, 2477, 2390, 772, 1174, 538, 2354, 11, 1566, 1047, 558, 1452, 2342, 2973, 1637, 2583, 1720, 2733, 1960, 505, 1796, 2104, 2199, 126, 353, 350, 1818, 324, 2565, 339, 942, 1824, 1833, 1917, 2830, 889, 545, 2587, 1915, 612, 670, 4815, 587, 2786, 4761, 1389, 3821, 1340, 497, 1547, 1295, 2868, 2497, 445, 491, 299, 712, 2324, 2610, 643, 2769, 2270, 1467, 1516, 1593, 3622, 3518, 295, 4387, 3690, 3475, 3692, 3725, 3382, 3508, 3160, 3434, 3433, 3234, 3858, 3695, 3332, 4073, 3599, 1489, 3585, 3600, 3862, 3700, 3365]. Théorème [1030, 1258, 2266, 635, 1691, 2972, 850, 2733]. Theorems [3257, 3884, 3905, 3283, 3995, 3693, 4146, 4579, 3344, 4544, 3678, 461, 83, 952, 2100, 923, 515, 628, 1032, 1135, 1413, 1460, 1481, 2672, 96, 905, 811, 682, 467, 1982, 1853, 2343, 2495, 2998, 2708, 2801, 1708, 1713, 4799, 2641, 2380, 2172, 1828, 1510, 997, 1943, 2482, 3051, 453, 2322, 1453, 180, 2314, 216, 1149, 1369, 1544, 1491, 1989, 1956, 2026, 2327, 2684, 890, 1230, 202, 3860]. theoretic [3173, 2313, 424, 1705, 1860, 2433, 4431, 4765]. theoretical [4293, 2581]. Theoretically [4133]. théorie [1463, 3428, 3582, 699]. Theories [4550, 1031, 420, 1695]. Theory [4304, 3900, 3788, 3439, 4160, 3668, 3354, 4289, 3522, 3112, 4408, 3657, 3358, 3694, 3461, 3307, 3376, 3363, 4217, 405, 2481, 2870, 1351, 1209, 2224, 2364, 235, 1463, 298, 3768, 3831, 294, 1633, 2620, 2, 898, 3428, 351, 1308, 1501, 1584, 2281, 2599, 1884, 2184, 2232, 741, 1897, 4707, 2029, 3582, 479, 274, 874, 3, 2325, 357, 2340, 46, 51, 159, 160, 161, 1706, 1503, 464, 2990, 2167, 231, 203, 1294, 2467, 699]. There [1439]. Theta [3300, 3584, 4702]. Thickness [3517, 4328]. Thin [1538, 3344, 1326, 2694, 932]. Thinnest [1527]. Third [4551, 982, 4222, 644, 1123, 459, 2747, 667, 3039]. Third-Engel [982]. Thom [4258]. Thompson [3181]. Thomsen [333, 4501]. Those [99]. though [4685]. Three [3266, 1352, 4109, 3624, 4494, 4266, 4516, 3356, 3995, 4337, 3912, 3529, 4129, 3390, 739, 1663, 1458, 3789, 535, 1306, 3813, 985, 198, 558, 2565, 2959]. Three-balls [3356]. Three-Dimensional [3390, 4266, 4337, 3789]. three-fold [1458]. Three-genus [4129]. Three-spheres [4516]. Three-Term [3624]. Threefolds [3876, 3529, 4704]. Threshold [4210, 4025, 4340]. Thresholds [4606]. Thurston [2946, 4778]. Tight [4725, 1789, 3029, 1252]. Tiles [4552]. Tilings [4567, 3048]. tilting [2249]. Time [3636, 4283, 3884, 4195, 3923, 923, 2777, 4273, 1654, 1980, 3798, 1251, 1347]. Times [3970, 4691, 1274]. Titchmarsh [3011]. Tits [2222, 4541, 1019]. Toda [2255]. Toeplitz [4741, 825, 4111, 4519, 3042, 4410, 2729, 2562, 3051, 3494, 2093, 2312, 3317, 4731]. Toms [3633]. topics [4815]. topoi [1522]. Topological [4156, 2669, 4552, 2045, 3112, 4607, 2579, 1777, 2152, 4619, 849, 3855, 3308, 4645, 2829, 4764, 234, 1986, 2085, 581, 486, 1378, 2162, 1988, 292, 761, 665, 1247, 2153, 2215, 1189, 1560, 1835, 1673, 448, 4817, 1925, 1965, 2062, 2713, 2943, 2728, 3082, 362, 2604, 2718, 1495, 1911, 1118, 1569, 2830, 1426, 1096]. Topologically [1864, 2843, 2135]. topologically-constant [2135]. Topologies [4114, 1712, 1679, 2220, 1772, 2825, 341, 1794, 549]. topologique [2135]. Topology [4637, 4366, 3168, 3916, 3386, 4100, 4473, 383, 622, 3482, 3555, 1947, 2670, 871, 1936, 2713, 324, 511]. topos [2224, 1424]. Toposian [1171]. toposienne [1171]. Toral [3238, 2730]. Tori [3987, 4520, 4208, 3645, 3309, 4583, 3318]. Toric [3477, 3881]. torision [626]. torision-free [626]. Toroidal [3880]. Torsion [4105, 2833, 4550, 2316, 3504, 4118, 3610, 2878, 4811, 4412, 4278, 2918, 4007, 4435, 3335, 2060, 2671, 972, 1181, 3768, 627, 1383, 1677, 3756, 4791, 1225, 2035, 1695]. Torsion-free [2316, 2918, 972]. torsionless [1068, 2328]. Torsors [4009, 3518]. Torus [4226, 4442, 4139, 2388, 2334, 2819, 2843, 2494, 2829]. Total [318, 4331, 4299, 4486, 3527, 1103, 2835, 4745, 1816, 1815]. totalement [2556]. Totally [859, 2507, 4653, 1024, 2469, 2569, 2978, 2824, 363, 2545]. Totaro [3518]. Totatives [3457, 3293]. Totient [3683, 2586]. touch [73]. Touching [2821]. tournament [1004, 438, 258, 366, 691, 790, 123, 335, 598]. Tournaments [3194, 2880, 4560, 1629, 418, 742, 702, 1623, 206, 125, 117, 189]. tours [3338]. tous [2952]. Tower [4321]. Towers [3287, 4647, 4786, 4805, 2399]. Trace [2816, 4353, 2779, 3225, 3461, 2666, 3634, 3362, 3391, 3033, 2644, 2898, 2132]. Trace-class [2666]. Traceless [4398]. Traces [4517, 3627, 4316, 4665, 2553, 4726, 2846]. Tracial [4021]. Tracially [3359]. Tracy [3848]. trajectories [288]. Transcendental [4016]. Transfer [3307, 4379, 2266, 2974]. transfert [2266]. transfinite [53]. Transform [3968, 3666, 3519, 3415, 4483, 3202, 4526, 3879, 1371, 351, 420, 462, 1797, 2254, 2241, 2567]. Transformation [4423, 530, 4274, 4187, 3958, 2754, 2838, 42, 2198, 1040, 66, 77, 153, 1732, 1878, 2828, 2494, 1776, 2935, 967, 501, 6, 12, 20]. Transformations [3575, 4273, 3239, 3617, 1381, 502, 3170, 34, 2576, 1901, 2503, 3198, 2843, 660, 499, 112, 2107, 57, 108, 2602, 1453, 736, 43, 3021, 346, 210, 289]. Transforming [3324]. Transforms [3666, 3121, 4308, 3277, 4020, 3295, 3609, 4607, 4394, 4482, 4525, 3979, 4223, 461, 1455, 719, 2163, 2495, 463, 1819, 815, 1723, 3839, 1476]. Transitive [2055, 536, 728, 1629, 2181]. transitively [346]. transitivity [2829, 913]. translate [1257]. Translates [1615, 4642, 3724, 3025]. Translation [4384]. Translations [4404, 4078]. translative [2030]. Translativity [374]. transonic [2168]. Transport [3827, 3984, 1045]. transportable [909]. transpose [737]. transposed [2137]. Transvection [3788]. transvections [2664]. Transversals [3724, 3872, 3400, 4792]. transversely [2609]. trapèze [1987]. trapezoid [1987]. Trapped [4025]. Traveling [4453]. Travelling [4057]. treatment [1205]. Tree [4167, 4125, 2372, 2714, 1426]. Trees [4598, 4138, 4279, 2526, 2965, 4696, 429, 774]. trefoil [1085]. triacontagon [211]. Triangle [4285, 3935, 3510, 284, 4693, 1724, 3801]. Triangles [4541, 690, 240]. Triangular [3232, 163, 3922, 2111, 375, 560, 2186, 2665]. Triangularizability [3405, 3064]. Triangularization [4578, 2564]. Triangularizing [1475]. Triangulated [4417]. triangulation [3801]. triangulations [2771]. Tridiagonal [3474]. Triebel [4250, 3539, 4339, 4400]. Trigonometric [3472, 2487, 2517, 601, 1831]. trigonometry [2981]. trilinear [1403]. Trinomial [3314]. Trinomials [3898]. tripartitions [521]. Triple [4288, 4265, 3868, 1528, 2206, 4811, 666]. Triples [844, 4094]. Tripotent [4558]. Tripotents [4324]. Trivectors [1513, 1931, 1597]. Trivial [3155, 4370, 3445, 4125, 3403, 417]. Triviality [4605]. trivializing [1510]. Trott [239]. Trudinger [4710, 4655, 4795]. Trudinger-type [4655]. truncated [74, 608, 964]. tubes [217]. Tubular [4348]. Tuples [4177, 1324]. Turaev [2959]. Turán [1317, 176, 4560, 3433, 443]. Turning [3874]. Turnpike [4757]. Twelve [3209, 3355]. Twin [4125]. Twist [4075, 4435, 4697]. Twisted [4570, 3369, 3749, 3504, 4374, 4370, 3890, 4095, 3631, 2316, 1297]. Twists [3577, 3192]. Two [4155, 4131, 4637, 4347, 3992, 3903, 3353, 4328, 3747, 4109, 3094, 4575, 3576, 2875, 4139, 2578, 4168, 4464, 4409, 4294, 3540, 4170, 4297, 4296, 4481, 4629, 3873, 2538, 3866, 4257, 4329, 3309, 3731, 3512, 2086, 4078, 4324, 4453, 3966, 4506, 2811, 4729, 1620, 1997, 4714, 3027, 2474, 241, 3793, 2499, 2228, 1849, 788, 1166, 1468, 1430, 2896, 2674, 1441, 2124, 1978, 2148, 601, 1328, 2840, 129, 523, 27, 986, 768, 1262, 2210, 2824, 2439, 631, 753, 2243, 4818, 2500, 1875, 1737, 216, 2630, 2199, 638, 2505, 3367]. Two-Component [4453]. Two-Dimensional [3353, 4481, 1997, 2674, 986, 2199]. two-generated [2630]. Two-graphs [4637]. two-parameter [788]. Two-Plane [4168, 4257, 3512, 4464, 4329]. Two-Point [3094, 4629]. two-sample [241]. Two-sided [4506, 768]. two-square [1468]. two-type [2148]. Two-Variable [3992, 4170]. Two-weighted [2875]. Tychonoff [1217]. Tychonov [1381]. Type [3732, 4402, 3666, 3927, 4368, 3520, 4425, 3327, 3655, 4164, 3887, 3247, 4167, 3870, 3995, 3430, 3797, 3157, 3935, 3339, 4393, 4579, 4625, 4500, 4235, 3361, 4659, 3958, 3344, 4000, 3523, 4148, 3524, 3418, 3598, 4527, 3614, 3589, 4448, 4449, 4433, 3585, 2247, 3313, 4241, 3349, 3451, 3220, 3742, 4507, 4400, 4162, 2751, 3884, 2923, 2852, 1135, 1413, 1481, 1790, 2613, 1298, 4793, 1243, 2754, 2059, 1747, 253, 1466, 2072, 4808, 34, 2307, 393, 2658, 2070, 2201, 2453, 603, 4371, 2148, 1619, 1537, 573, 1073, 1486, 4518, 1074, 3796, 4655, 3776, 4519, 3835, 1553, 3189, 1684, 2784, 847, 3042, 2601]. type [1075, 1854, 572, 2763, 2988, 3051, 3759, 1174, 1488, 538, 988, 2135, 4703, 354, 1280, 1336, 2425, 4543, 1423, 350, 1523, 118, 1833, 1917, 4762, 2327, 183, 4751, 4735, 3966, 1027, 3947]. Types [4493, 3670, 3578, 4184, 328, 1296, 1699, 3002]. Typical [3924, 4107, 1941, 1111]. typically [3754].

Uchimura [2019]. Uchiyama [1156]. UFD [3751]. UHF [3520]. Uhlenbeck [2502]. Ulam [4359, 2026]. Ultracontractive [4572]. ultrafilter [1767]. ultrafilters [1272]. Ultrametric [3659]. ultrapower [2138]. ultrapowers [1272]. Ultraspherical [3473, 4378, 527, 2047, 2071]. umbilics [2536]. UMD [4671, 3189]. UMD-valued [3189]. unavoidable [4814]. unbiased [987]. Unbounded [338, 3873, 2727, 2500, 2607]. Uncertainty [4267]. Uncomplemented [3993, 3755]. unconditional [959]. Uncountable [4061, 1909, 1745, 2643]. Uncountably [2511, 2398]. Underlying [4122, 2144]. undetermined [356]. undirected [751]. unfoldings [2645]. Unicity [4443, 2905, 2457]. unicoherent [1783]. Uniconnected [4746]. Unicyclic [1232, 1400]. unified [3806, 1880, 2081]. Uniform [2207, 3557, 3041, 3473, 2938, 3280, 4806, 3305, 3858, 3377, 1546, 2291, 1688, 1394, 1986, 1380, 1679, 665, 1922, 2144, 3013, 1451, 1885, 1763, 2600, 4211, 1541, 750, 2241, 1118, 2543, 238, 401, 3882, 2874]. uniformément [1891]. uniformes [2874, 2600]. uniformities [991]. uniformity [2215, 2152, 849]. Uniformization [4463, 4471]. Uniformly [4777, 3969, 4437, 1870, 1934, 1005, 2938, 1291, 2076, 2386, 931, 1827, 2295, 1935, 2610, 1891]. Uniformly-bounded [1891]. unifying [1515]. unilateral [2197]. unimodal [2803]. Unimodular [3541, 832, 139]. Union [4530, 1070, 1207, 2556, 2759, 2463, 2824]. unions [2075]. Unipotent [1574, 3762, 1866, 2112, 1475]. Unique [3865, 4240, 132, 3768, 2215, 2152, 849, 1233, 1680]. Uniquely [1087, 2782]. Uniqueness [2267, 2381, 2176, 3930, 4166, 4272, 4293, 4625, 1333, 3163, 1091, 991, 3424, 3391, 4057, 3392, 1578, 2174, 40, 1300, 1390, 2983, 1236, 1818, 1340, 299]. Unisolvence [529]. Unit [3845, 4461, 4084, 3476, 4391, 3777, 3914, 3694, 3525, 3102, 1352, 4776, 802, 2874, 1692, 2641, 2940, 2188, 3079, 2248, 2845, 4670, 2847, 5, 4778, 2976, 2479, 2545]. Unit-Regular [3777]. unitaires [2010]. Unital [3197]. Unitaries [3470]. Unitary [3324, 4109, 4136, 4738, 4320, 4323, 315, 3043, 1302, 2010, 737, 1285]. unité [2874]. unités [2976]. Units [2257, 3323, 3155, 1752, 3445, 2783, 4428, 2535, 4608, 3008, 2527, 2052, 2043, 2194, 2480, 2160, 2229, 984, 2710, 2842, 2878, 2123, 1127, 1406]. unity [3045, 2607, 1934]. Univ [1317]. Univalent [3163]. Univalently [3947]. Universal [3574, 4061, 4476, 3665, 4384, 3266, 4436, 3844, 3476, 4555, 2648, 4321, 3388, 4217, 2809, 1031, 2208, 2967, 324, 4789, 4702, 2962]. Universality [4282, 3670]. Universally [4044]. universe [1749]. universels [2292]. Unlabelled [4422]. Unruly [2460, 4821]. Unstable [2512, 4120, 3002]. upcrossings [1076]. Upper [2111, 1689, 4069, 3922, 632, 1327, 1743, 2808]. urchin [594, 465]. Urysohn [2318, 1027]. usco [2588]. useful [741, 231, 203]. Using [4200, 3988, 4228, 1260, 3360, 3806, 1702, 1631]. usual [1968].

V [161]. Valdivia [2571]. valencies [313]. valeurs [2459, 2952, 3004]. Valiron [1190]. Vallée [855]. valuated [1780]. Valuation [3989, 2496, 1559, 493, 2391]. Valuations [3566, 2100, 2316, 765]. Value [3651, 3734, 3487, 3930, 3997, 4067, 4584, 4454, 2256, 2120, 2228, 2307, 4770, 4753, 813, 2660, 3795, 4820, 934, 2335, 3017, 1173, 2845, 937, 1488, 773, 938, 2973, 892, 1283, 1315, 1946]. Valued [3925, 3540, 3089, 3694, 3116, 3435, 3117, 3861, 4803, 2546, 2427, 1436, 1462, 3794, 2621, 980, 1922, 4134, 2230, 1817, 1442, 276, 2281, 1404, 3189, 3004, 2701, 2502, 1254, 1476, 823, 1095, 2157, 2063]. Values [3439, 4614, 3547, 3326, 4286, 4374, 3403, 4074, 3451, 3392, 2311, 903, 2952, 173, 3068, 2213, 2532, 442, 2144, 1441, 814, 2627, 1252, 2562, 4798, 4728, 1387, 1093, 2787, 826]. Vandermonde [573, 572]. Vanishes [3894]. Vanishing [3468, 4570, 4248, 3933, 2840, 4536, 3693, 4591, 4525, 3646, 4658, 4544, 4728, 3512, 3678, 2671, 3005, 822]. Variable [3992, 4595, 2753, 1969, 4170]. Variables [4456, 4384, 3745, 3736, 4332, 3492, 3309, 3496, 716, 1030, 923, 1319, 3041, 2366, 4753, 1734, 4747, 435, 2210, 2025, 2390, 1840, 1337, 2116, 302, 1833, 1917, 2446]. Variants [3503, 3760, 1508, 4260, 1591]. variates [608, 964]. Variation [3971, 2835, 706, 309, 2071, 1816, 1815]. Variational [4692, 169, 2671, 1315, 150, 799]. Variations [904, 2572, 1555, 3812, 3634, 4398, 4507, 2851, 981]. variété [3259]. variétés [3322]. Varieties [3749, 3282, 3708, 3289, 4184, 4338, 3477, 4539, 4653, 3402, 3982, 4215, 3881, 2752, 4683, 2592, 2612, 2654, 3830, 4737, 4669, 2337, 2411, 2677, 4750, 2971, 2942, 3066, 2581, 1238, 4676]. Variety [3352, 3713, 3259, 763, 2145]. Various [4613, 1862]. varisolvent [1848]. Vary [3090]. Vassiliev [3553]. Vaught [1372]. VB [4467]. VB-Courant [4467]. vecteurs [3274, 3101]. Vector [3574, 3367, 2506, 3107, 1211, 3502, 3641, 4134, 3910, 3250, 3477, 4169, 3114, 3540, 3956, 3089, 4341, 3116, 4382, 4626, 4544, 3103, 3847, 3456, 3167, 1580, 338, 2306, 2381, 4714, 1258, 3274, 1931, 3794, 442, 2230, 2055, 930, 2160, 276, 3004, 987, 1495, 2747, 1863, 1569, 2063, 1780, 1355]. Vector-Valued [3540, 3089, 3116, 4134, 3794, 2230, 3004, 2063]. vectoriel [1258]. vectorielles [3004]. Vectors [3656, 3944, 749, 2418, 1384]. velocities [2834]. velocity [2674]. verbally [4685]. Verdier [3934]. Verma [2780]. Verschiebung [4130]. Version [4179, 3862, 3530, 1573, 2344]. Versions [4011, 2548, 3169]. Versus [4279, 4760, 2136, 2231]. Vertex [728, 536, 1400, 4672]. vertex-unicyclic [1400]. Verticality [2560]. Vertices [4211, 1250]. Very [3273, 322, 1857]. Via [4325, 3646, 4419, 3701, 4660, 4787, 4386, 4109, 4689, 3522, 4044, 4270, 2495, 4426, 4271, 4427, 1501, 1584, 1261, 4631, 4582, 2765, 1735, 4549, 4400]. vibrating [1844]. Vibration [4481]. View [3984]. VII [1093]. Vilenkin [3074, 3424, 2591]. Viniberghi [4090]. Vinogradov [4722, 2345]. Virasoro [3369]. Viro [3553, 2959]. Virtually [3272, 4713]. Virus [4053]. Viscosity [4658]. Visibility [4009]. Vision [4618]. visits [1062]. Vitali [1590, 183]. Vladimirskii [1244]. VMO [4229]. Voiculescu [3431]. Volterra [3995, 4703, 4195, 890]. Volume [3517, 4742, 3896, 3731, 4508, 4509, 4568, 4569, 4589, 4627, 4628, 4648, 4686, 4687, 4705, 4706, 4766, 2519, 2656]. volume-like [2519]. Vore [1109]. Vortex [3664]. vs [1557].

W [411, 889]. W. [3905]. Waerden [3092, 51]. Wald [317]. Walk [4092, 872, 843, 845, 1062, 167, 1914, 242]. Wall [2907]. Wallace [2961]. Walsh [3524]. Walter [2567]. Waring [3704]. Warped [4500, 4664]. Wasserstein [3971]. Water [3919]. Watons [1121]. Watson [77, 2355]. Wave [3427, 4312, 3919, 3391, 675, 2504, 1514]. Wavefronts [4453]. Wavelet [4109, 4063, 4582]. wavelets [3075, 3757]. Waves [4057, 1766]. way [27, 2680]. WCG [3480]. Weak [3180, 2546, 3105, 4040, 1417, 3686, 3276, 2773, 2894, 3506, 4518, 3797, 3339, 4394, 4188, 1364, 3102, 3228, 3615, 3601, 4489, 297, 2948, 3220, 3255, 2739, 3370, 1557, 1559, 956, 1849, 791, 3796, 3189, 1590, 1412]. Weak-type [4518, 3189]. Weakly [1296, 4015, 3502, 4781, 3444, 4519, 4373, 3374, 4639, 1052, 3390, 1005, 2365, 1220, 1980, 1279, 3054]. Weakness [2670]. Webster [4337]. Wedge [4527]. Wedge-of-the-edge [4527]. wedges [1754]. Weibull [597]. Weierstrass [576, 502, 956, 463, 2265, 2473, 4149]. Weierstrass-transforms [463]. Weight [4420, 4407, 3540, 4499, 4051, 4301, 1398, 2027, 518, 804]. Weighted [3515, 4035, 3262, 2872, 3105, 3195, 3453, 4404, 3951, 2620, 2777, 4164, 4267, 3772, 4114, 4355, 3932, 1442, 4518, 3656, 2697, 4580, 3252, 3253, 2727, 3957, 88, 3938, 4581, 2397, 1319, 1813, 2837, 2875, 3031, 1675, 1916, 4717, 2470, 3042, 2237, 4670, 3051, 2114, 1908, 2289, 1991, 2810]. Weights [3396, 1455, 914, 3778, 3759]. Weil [4039, 3010, 3579, 3432, 4823, 4805]. Well [4551, 4226]. Well-posedness [4551, 4226]. Welsh [577]. Weyl [4739, 4386, 3011, 3890, 4407]. Wheels [4279]. where [2180, 1305, 4774, 940]. Which [2122, 3410, 430, 3331, 3006, 4324, 1755, 2427, 3060, 520, 953, 1353, 669, 1081, 1889, 2519, 1084, 2420, 1270, 1304, 1070, 1207, 1723, 3079, 2248, 1114, 1129, 1224, 1935, 2011, 90, 1893, 2140, 729, 2541, 302, 1822, 2449, 1946, 652, 303, 2089]. Whitehead [3503]. Whitney [4736, 146, 2437, 135, 2106, 1742]. Whittaker [3277]. whole [2952, 1946]. Whose [4034, 3353, 3940, 3589, 4018, 3714, 4007, 2002, 749, 1376, 3370, 2225, 73, 1641, 618, 1530, 1168, 2955, 2391, 2802, 4558, 1249, 2826, 1629, 3894, 2989, 2767, 1764]. Widom [3848]. Width [3719, 3809, 382, 4790, 2343, 1249, 4768, 182]. Widths [3643, 14]. Wiener [1652, 1972, 781, 942, 1490]. Wigert [779]. Wilks [1183]. Willard [1098]. Williams [3671]. Willmore [3603]. Wilson [3485, 2482, 2363, 2935]. Wiman [1190]. Windows [3882]. Wintner [2014]. Wirtinger [3636]. Within [4527]. Without [4177, 3460, 4492, 839, 976, 3183, 1011, 2634, 2229, 2709, 846, 2295, 3130, 2494, 2831]. Witt [2426, 2119, 2333, 2476]. Witten [3693]. Wolff [4799]. Wolfowitz [317]. Wonderful [3333]. Wong [759, 566]. word [732]. Words [2634, 3410, 1171]. Work [3225]. Wreath [4108, 864, 1176, 3941, 4656, 1297]. written [1070, 1207]. Wronski [4406]. Wythoff [37, 2462].

X [4790, 3719, 3202]. X-Ray [3202, 3719]. X-raying [4790].

Y. [759]. Yamabe [4368, 4546]. Yang [4522, 483, 1697, 4746, 3437]. Yau [3454, 3605, 4750, 3529, 4507]. Young [96, 3549, 2349, 2270]. Yudovich [4658].

Zariski [4737]. Zassenhaus [2478, 655]. Zayed [2567]. Zelmanowitz [1886]. Zero [2110, 2037, 3369, 3667, 4351, 4741, 3275, 3096, 4290, 4335, 4118, 4205, 4144, 3580, 3912, 3301, 4213, 3631, 4029, 4101, 2568, 2612, 106, 1268, 1533, 2690, 4681, 2908, 2955, 2694, 2313, 3776, 448, 4067, 876, 1948, 1730, 2767, 1846, 1929, 2044]. Zero-dimensional [4213, 2612, 4681, 1948, 1846, 1929]. Zero-Divisor [3912, 4029, 4335, 4205, 4144, 4101]. Zero-divisors [4118, 106]. Zero-free [2037]. Zero-One [3667]. zero-sets [1533]. zeroes [2370, 1773]. Zeros [3689, 4532, 3536, 918, 1568, 4005, 500, 1437, 1733, 4756, 4800, 1626, 1807, 1328, 396, 2803, 1089, 550, 524, 406, 783, 833, 2920, 2071, 1648, 4756, 1089]. Zeta [3469, 3651, 3328, 4044, 4532, 3630, 3229, 4450, 4756, 2532, 176, 2158, 2559, 2488, 2735]. Zeta-function [4044, 4532, 2158]. zeta-functions [2488]. zeta-polynomials [2559]. Zettl [2837]. Ziegler [4046]. Zilber [3382]. Zindler [3975]. ZL [4131]. ZL-amenability [4131]. Zorn [1979]. Zygmund [4202, 726, 4339].

References

Blum:1958:PCL

[1] Richard Blum. On a pointwise construction of the lemniscate. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 1(1):1– 4, January 1958. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Erdos:1958:EPN

[2] Paul Erdős. On an elementary problem in number theory. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 1(1):5– 8, January 1958. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Luxemburg:1958:CSA

[3] W. A. J. Luxemburg. On the convergence of successive approximations in the theory of ordinary differential equations. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 1(1):9–20, January 1958. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Robinson:1958:RPH

[4] G. de B. Robinson. A remark by Philip Hall. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 1(1):21–23, January 1958. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Rooney:1958:SPF

[5] P. G. Rooney. On some properties of functions regular in the unit circle. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 1(1):25–29, January 1958. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Zassenhaus:1958:NFOa

[6] Hans Zassenhaus. On a normal form of the orthogonal transformation. I, II, III. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 1(1):31–40, January 1958. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Robinson:1958:OIMa

[7] Abraham Robinson. Outline of an introduction to mathematical logic. I. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 1(1):41–54, January 1958. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Erdos:1958:SRP

[8] Paul Erdős. Some remarks on a paper of McCarthy. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 1(1):71–75, ???? 1958. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic). See remarks and corrections [60].

Findlay:1958:GRQa

[9] G. D. Findlay and J. Lambek. A generalized ring of quotients. I. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 1(1):77– 85, ???? 1958. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Miksa:1958:RPF

[10] F. L. Miksa, L. Moser, and M. Wyman. Restricted partitions of finite sets. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 1(1):87–96, ???? 1958. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Peck:1958:YAP

[11] J. E. L. Peck. Yet another proof of the minimax theorem. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 1(1):97– 100, ???? 1958. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Zassenhaus:1958:NFOb

[12] H. J. Zassenhaus. On a normal form of the orthogonal transformation, II. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 1(1):101–112, ???? 1958. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Erdos:1958:AFS

[13] Paul Erdős. Asymptotic formulas for some arithmetic functions. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 1(1): 149–153, ???? 1958. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Moser:1958:RWC

[14] W. O. J. Moser. On the relative widths of coverings by convex bodies. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 1(1):154, 168, 174, ???? 1958. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Findlay:1958:GRQb

[15] G. D. Findlay and J. Lambek. A generalized ring of quotients II. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 1(1): 155–167, ???? 1958. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Narayana:1958:NCI

[16] T. V. Narayana and G. E. Fulton. A note on the compositions of an integer. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 1(1):169–173, ???? 1958. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Schwerdtfeger:1958:D

[17] H. Schwerdtfeger. On the discriminant x′Ax · y′Ay − (x’Ay)^{2}. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 1(1): 175–179, ???? 1958. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Wild:1958:IQF

[18] Jonathan Wild. On the index of a quadratic form. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 1(1):180, ???? 1958. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Scherk:1958:NHS

[19] Peter Scherk. On a note by H. Schwerdtfeger. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 1(1):181–182, ???? 1958. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Zassenhaus:1958:NFOc

[20] Hans Zassenhaus. On a normal form of the orthogonal transformation III. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 1(1):183–191, ???? 1958. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Robinson:1958:OIMb

[21] Abraham Robinson. Outline of an introduction to mathematical logic. II. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 1(2):113–127, May 1958. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Robinson:1959:OIM

[22] Abraham Robinson. Outline of an introduction to mathematical logic. III, IV. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 2(??):33–42, ???? 1959. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic). See Canad. Math. Bull. 1 (1958), 193– 208.

Wild:1959:NDB

[23] Jonathan Wild. Notes: On the discriminants of a bilinear form. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 2(??): 46–48, ???? 1959. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Melzak:1959:ECA

[24] Z. A. Melzak. Existence of certain analytic homeomorphisms. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 2(??): 71–75, ???? 1959. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Crowe:1959:RQP

[25] Donald W. Crowe. A regular quaternion polygon. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 2(??):77–79, ???? 1959. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Marcus:1959:NHP

[26] M. Marcus and N. A. Khan. A note on the Hadamard product. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 2(??): 81–83, ???? 1959. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Lambek:1959:STW

[27] J. Lambek and L. Moser. On some two way classifications of integers. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 2(??):85–89, ???? 1959. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Chalk:1959:DPL

[28] J. H. H. Chalk and Paul Erdős. On the distribution of primitive lattice points in the plane. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 2(??):91–96, ???? 1959. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Schwerdtfeger:1959:NNA

[29] Hans Schwerdtfeger. Notes on numerical analysis. I. Polynomial iteration. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 2(??):97–110, ???? 1959. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Halperin:1959:DFD

[30] Israel Halperin. Discontinuous functions with the Darboux property. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 2(??):111–118, ???? 1959. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Stanton:1959:HBM

[31] R. G. Stanton. The half-baked method for cooking of experimental data. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 2(??):125–150, ???? 1959. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Melzak:1959:MTC

[32] Z. A. Melzak. Minkowski’s theorem with curvature limitations. I. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 2(??): 151–158, ???? 1959. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Dulmage:1959:NSR

[33] A. L. Dulmage and N. S. Mendelsohn. A note on the stochastic rank of a bipartite graph. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 2(??):159–162, ???? 1959. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Ellis:1959:CLT

[34] H. W. Ellis and T. E. Mott. On continuous linear transformations of integral type. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 2(??):163–166, ???? 1959. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Wong:1959:SIR

[35] E. T. Wong and R. E. Johnson. Self-injective rings. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 2(??):167– 173, ???? 1959. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Melzak:1959:CSS

[36] Z. A. Melzak. A class of star-shaped bodies. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 2(??):175–180, ???? 1959. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Connell:1959:GWG

[37] Ian G. Connell. A generalization of Wythoff’s game. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 2(??):181– 189, ???? 1959. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Connell:1959:SPB

[38] Ian G. Connell. Some properties of Beatty sequences. I. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 2(??):190– 197, ???? 1959. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Banaschewski:1959:KSC

[39] Bernhard Banaschewski. On the Katětov and Stone–Čech extensions. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 2(1):1–4, January 1959. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Brauer:1959:NUC

[40] Fred Brauer. A note on uniqueness and convergence of successive approximations. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 2(1):5–8, January 1959. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Moser:1959:ADM

[41] W. O. J. Moser. Abstract definitions for the Mathieu groups M_{11} and M_{12}. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 2(1):9–13, January 1959. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Goldman:1959:CFL

[42] A. J. Goldman and M. Marcus. Convexity of the field of a linear transformation. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 2(1):15–18, January 1959. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Rooney:1959:IGT

[43] P. G. Rooney. On the inversion of general transformations. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 2(1): 19–24, January 1959. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Brainerd:1959:RQB

[44] B. Brainerd and J. Lambek. On the ring of quotients of a Boolean ring. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 2(1):25–29, January 1959. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Melzak:1959:PCP

[45] Z. A. Melzak. A property of convex pseudopolyhedra. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 2(1):31–32, January 1959. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Moser:1959:NNT

[46] Leo Moser. Notes on number theory. I. On the product of the primes not exceeding n. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 2(2):119–121, May 1959. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Tutte:1960:NHG

[47] W. T. Tutte. A non-Hamiltonian graph. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 3(??):1–5, ???? 1960. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Chalk:1960:NDI

[48] J. H. H. Chalk. A note on a discriminant inequality. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 3(??):7–9, ???? 1960. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Thierrin:1960:SCA

[49] Gabriel Thierrin. Sur la structure d’une classe d’anneaux (French) [On the structure of a class of rings]. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 3(??):11–16, ???? 1960. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Connell:1960:SPB

[50] Ian G. Connell. Some properties of Beatty squences. II. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 3(??):17–22, ???? 1960. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Moser:1960:NNT

[51] Leo Moser. Notes on number theory. II. On a theorem of van der Waerden. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 3(??):23–25, ???? 1960. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Struik:1960:NPP

[52] Ruth Rebekka Struik. A note on prime-power groups. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 3(??):27–30, ???? 1960. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Zakon:1960:CMT

[53] Elias Zakon. On common multiples of transfinite numbers. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 3(??): 31–33, ???? 1960. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Guttman:1960:PM

[54] Irwin Guttman. On a problem of L. Moser. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 3(??):35–39, ???? 1960. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Schwerdtfeger:1960:NNA

[55] Hans Schwerdtfeger. Notes on numerical analysis. II. Interpolation and curve fitting by sectionally linear functions. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 3(??):41–57, ???? 1960. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic). See correction [613].

Beesack:1960:EPE

[56] Paul R. Beesack. Elementary proofs of the extremal properties of the eigenvalues of the Sturm–Liouville equation. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 3(??):59–77, ???? 1960. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Macphail:1960:SST

[57] M. S. Macphail. Sequence and series transformations. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 3(??):80–83, ???? 1960. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Baker:1960:SFE

[58] I. N. Baker. Solutions of the functional equation (f(x))^{2} − f(x^{2}) = h(x). Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 3(??):113–120, ???? 1960. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Dixon:1960:FAG

[59] J. D. Dixon. A finite analogue of the Goldbach problem. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 3(??):121– 126, ???? 1960. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Erdos:1960:RCM

[60] Paul Erdős. Remarks and corrections to my paper “Some remarks on a paper of McCarthy”. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 3(??):127–129, ???? 1960. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic). See [8].

Findlay:1960:RHR

[61] G. D. Findlay. Reflexive homomorphic relations. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 3(??):131–132, ???? 1960. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Johnson:1960:GMG

[62] Norman W. Johnson. A geometric model for the generalized symmetric group. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 3(??):133–142, ???? 1960. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Marcus:1960:NGD

[63] Marvin Marcus and Nisar A. Khan. A note on a group defined by a quadratic form. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 3(??):143–148, ???? 1960. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

McGee:1960:MCG

[64] W. F. McGee. A minimal cubic graph of girth seven. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 3(??):149– 152, ???? 1960. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Melzak:1960:NHT

[65] Z. A. Melzak. A note on Hilbert’s tenth problem. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 3(??):153–156, ???? 1960. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Rooney:1960:IOF

[66] P. G. Rooney. On an inversion operator for the Fourier transformation. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 3(??):157–165, ???? 1960. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Thierrin:1960:DR

[67] G. Thierrin. On duo rings. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 3(??):167–172, ???? 1960. CODEN CM-BUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Ellis:1960:SBS

[68] H. W. Ellis and D. G. Kuehner. On Schauder bases for spaces of continuous functions. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 3(??):173–184, ???? 1960. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Halperin:1960:OEP

[69] Israel Halperin. Odd and even permutations. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 3(??):185, ???? 1960. CO-DEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Lipman:1960:LCF

[70] J. Lipman. A lemma on continuous functions. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 3(??):186–187, ???? 1960. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Lorch:1960:EPF

[71] Lee Lorch. On an extremal problem in Fourier series. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 3(??):188, ???? 1960. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Brainerd:1960:BPM

[72] Barron Brainerd. On Birkhoff’s problem 73 for monoids. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 3(??):217– 220, ???? 1960. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Brown:1960:HTQ

[73] W. G. Brown. Hesse’s theorem for a quadrilateral whose sides touch a conic. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 3(??):221–224, ???? 1960. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Guttman:1960:TSP

[74] I. Guttman. Tests for the scale parameter of the truncated normal. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 3(??):225–236, ???? 1960. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Johnson:1960:AGB

[75] Diane M. Johnson, A. L. Dulmage, and N. S. Mendelsohn. On an algorithm of G. Birkhoff concerning doubly stochastic matrices. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 3(??): 237–242, ???? 1960. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Korsak:1960:DET

[76] A. Korsak and C. Schubert. A determinant expression of Tchebychev polynomials. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 3(??):243–246, ???? 1960. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Rooney:1960:WLT

[77] P. G. Rooney. On the Watson and Laplace transformation. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 3(??): 247–253, ???? 1960. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Sen:1960:GMR

[78] D. K. Sen. On geodesics of a modified Riemannian manifold. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 3(??):

255–261, ???? 1960. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Vanstone:1960:ADG

[79] J. R. Vanstone. Autoparallel deviation in the geometry of Lyra. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 3(??): 263–271, ???? 1960. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Halperin:1960:INA

[80] Israel Halperin. Introduction to von Neumann algebras and continuous geometry. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 3(??):273–288, ???? 1960. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic). See remark [124].

Prowse:1960:ACF

[81] Ebenezer Prowse. On the analytic continuation of the ζ-function. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 3 (??):289–292, ???? 1960. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Wild:1960:ISF

[82] Jonathan Wild. On the index of a symmetric form. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 3(??):293– 295, ???? 1960. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Al-Dhahir:1961:GCC

[83] M. W. Al-Dhahir. A generalization of Cox’s chain of theorems. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 4(??): 1–6, ???? 1961. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Atkinson:1961:PES

[84] F. V. Atkinson. On a problem of Erdős and Szekeres. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 4(??):7–12, ???? 1961. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

deBuda:1961:NCS

[85] R. G. de Buda. A note on the construction of a set of fundamental circuits on a surface of genus p. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 4(??):13–21, ???? 1961. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Marcus:1961:NNM

[86] Marvin Marcus and Nisar Khan. A note on normal matrices. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 4(??): 23–27, ???? 1961. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Mendelsohn:1961:PCD

[87] N. S. Mendelsohn. Permutations with confined displacements. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 4(??): 29–38, ???? 1961. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Moser:1961:WC

[88] L. Moser and E. L. Whitney. Weighted compositions. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 4(??):39–43, ???? 1961. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Moser:1961:EFP

[89] William Moser. An enumeration of the five parallelohedra. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 4(??): 45–47, ???? 1961. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Ree:1961:CPM

[90] Rimhak Ree. On certain pairs of matrices which do not generate a free group. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 4(??):49–52, ???? 1961. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Schwerdtfeger:1961:NNA

[91] Hans Schwerdtfeger. Notes on numerical analysis. III. Further remarks on sectionally linear functions. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 4(??):53–55, ???? 1961. CODEN CM-BUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Thompson:1961:DIP

[92] R. C. Thompson. A determinantal inequality for positive definite matrices. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 4(??):57–62, ???? 1961. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Utumi:1961:CRR

[93] Y. Utumi. On continuous regular rings. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 4(??):63–69, ???? 1961. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Scherk:1961:OMR

[94] P. Scherk. On an observation of M. Riesz. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 4(??):70–71, ???? 1961. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Barbeau:1961:RAD

[95] E. J. Barbeau. Remarks on an arithmetic derivative. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 4(??):117– 122, ???? 1961. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Bullen:1961:SEH

[96] P. S. Bullen. Some extensions of the Hausdorff–Young and Paley theorems. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 4(??):123–138, ???? 1961. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Lachapelle:1961:LMD

[97] Benoît Lachapelle. L’espérance mathématique du nombre de nombres premiers aléatoires inférieurs ou égaux à x. (French) [The mathematical expectation of the number of aleatory prime numbers less than or equal to x]. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 4(??):139–142, ???? 1961. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Melzak:1961:PS

[98] Z. A. Melzak. On the problem of Steiner. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 4(??):143–148, ???? 1961. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Mendelsohn:1961:SNA

[99] N. S. Mendelsohn. Those Stirling numbers again. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 4(??):149–151, ???? 1961. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Oler:1961:FPP

[100] Norman Oler. A finite packing problem. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 4(??):153–155, ???? 1961. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Tutte:1961:CG

[101] W. T. Tutte. On the colourings of graphs. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 4(??):157–160, ???? 1961. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Atkinson:1961:RDM

[102] F. V. Atkinson. The reversibility of a differentiable mapping. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 4(??): 161–181, ???? 1961. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Wild:1961:RLO

[103] Jonathan Wild. Right and left orthogonality. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 4(??):182–184, ???? 1961. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Melzak:1961:PGP

[104] Z. A. Melzak. On a problem in geometrical probability. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 4(??):185– 186, ???? 1961. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Coleman:1961:DMF

[105] A. J. Coleman. Density matrices of n-fermion systems. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 4(??):209– 212, ???? 1961. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Banaschewski:1961:PAZ

[106] Bernhard Banaschewski. On proving the absence of zero-divisors for semi-group rings. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 4(??):225–231, ???? 1961. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Melzak:1961:NTE

[107] Z. A. Melzak. A note on the Tarry–Escott problem. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 4(??):233– 237, ???? 1961. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Moyls:1961:NIL

[108] B. N. Moyls and N. A. Khan. A norm inequality for linear transformations. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 4(??):239–242, ???? 1961. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Swanson:1961:NAP

[109] C. A. Swanson. On a non-analytic perturbation problem. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 4(??):243– 248, ???? 1961. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Wild:1961:SEP

[110] Jonathan Wild. Some elementary properties of bilinear forms. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 4(??): 249–259, ???? 1961. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Wild:1961:SIB

[111] Jonathan Wild. On some invariants of a bilinear form. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 4(??):261– 264, ???? 1961. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Lipman:1961:DAG

[112] Joe Lipman. Definition of affine geometry by a group of transformations. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 4(??):265–278, ???? 1961. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Melzak:1961:IAA

[113] Z. A. Melzak. An informal arithmetical approach to computability and computation. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 4(??):279–293, ???? 1961. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Lambek:1961:HPI

[114] Joachim Lambek. How to program an infinite abacus. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 4(??):295– 302, ???? 1961. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic). See correction [145].

Gould:1961:GPL

[115] H. W. Gould. A generalization of a problem of L. Lorch and L. Moser. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 4(??):303–305, ???? 1961. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Isbell:1962:IDS

[116] J. R. Isbell. Infinite doubly stochastic matrices. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 5(??):1–4, ???? 1962. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Moon:1962:DCR

[117] J. W. Moon and L. Moser. On the distribution of 4-cycles in random bipartite tournaments. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 5(??):5–12, ???? 1962. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Sherk:1962:FRM

[118] F. A. Sherk. A family of regular maps of type {6, 6}. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 5(??):13–20, ???? 1962. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Lorch:1962:AFM

[119] Lee Lorch. On approximation by Fejér means to periodic functions satisfying a Lipschitz condition. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 5(??):21–27, ???? 1962. CODEN CM-BUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Thierrin:1962:ERQ

[120] G. Thierrin. Extensions radicales et quasi-radicales dans les anneaux. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 5(??):29–35, ???? 1962. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Bruns:1962:RBA

[121] Günter Bruns. On the representation of Boolean algebras. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 5(??):37–41, ???? 1962. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Narayana:1962:AMT

[122] T. V. Narayana. An analogue of the multinomial theorem. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 5(??): 43–50, ???? 1962. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Moon:1962:SSP

[123] J. W. Moon. On the score sequence of an n-partite tournament. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 5(??): 51–58, ???? 1962. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Halperin:1962:RMP

[124] Israel Halperin. Remark to my paper: Introduction to von Neumann algebras and continuous geometry. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 5(??):59, ???? 1962. CODEN CM-BUA3. ISSN 0008-4395 (print), 1496-4287 (electronic). See [80].

Moon:1962:AAT

[125] J. W. Moon and L. Moser. Almost all tournaments are irreducible. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 5 (??):61–65, ???? 1962. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Scherk:1962:FTA

[126] P. Scherk. On the fundamental theorem of affine geometry. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 5(??): 67–69, ???? 1962. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Gratzer:1962:SEC

[127] G. Grätzer and María J. Wonenburger. Some examples of complemented modular lattices. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 5(??):111–121, ???? 1962. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Thierrin:1962:APO

[128] G. Thierrin. Sur les anneaux partiellement ordonnées. (French) [On partially-ordered rings]. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 5(??):123–128, ???? 1962. CODEN CM-BUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Iosifescu:1962:PPS

[129] Marius Iosifescu and Solomon Marcus. Sur un problème de P. Scherk, concernant la somme des carrés de deux dérivées. (French) [On a problem of P. Scherk, concerning the some of squares of two derivatives]. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 5(??): 129–132, ???? 1962. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Maranda:1962:SRL

[130] J.-M. Maranda. Some remarks on limits in categories. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 5(??):133– 146, ???? 1962. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Christensen:1962:MRR

[131] Joanne Christensen. On maximal rings of right quotients. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 5(??):147– 149, ???? 1962. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Connell:1962:ANF

[132] Ian G. Connell. On algebraic number fields with unique factorization. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 5(??):151–166, ???? 1962. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Ditor:1962:NKF

[133] Seymour Ditor. A new kind of free extension for projective planes. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 5 (??):167–170, ???? 1962. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Hendricks:1962:FSD

[134] J. R. Hendricks. The five and six dimensional magic hypercubes of order 3. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 5(??):171–190, ???? 1962. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Narayana:1962:NTM

[135] T. V. Narayana and H. M. Pettigrew. A note on a theorem of Moser and Whitney. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 5(??):191–220, ???? 1962. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Dirac:1962:NSG

[136] G. A. Dirac. Note on the structure of graphs. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 5(??):221–227, ???? 1962. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Wyman:1962:ELP

[137] Max Wyman. An expansion of the Laguerre polynomials, L_{n^{α}}(z). Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 5 (??):229–240, ???? 1962. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Dulmage:1962:EPM

[138] A. L. Dulmage and N. S. Mendelsohn. The exponent of a primitive matrix. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 5(??):241–244, ???? 1962. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Trott:1962:PGU

[139] Stanton M. Trott. A pair of generators for the unimodular group. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 5(??): 245–252, ???? 1962. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic). See remarks [239].

Marcus:1962:MLM

[140] Marvin Marcus. Matrices in linear mechanical systems. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 5(??):253– 257, ???? 1962. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Dixon:1962:PRR

[141] J. D. Dixon. Polynomials with real roots. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 5(??):259–263, ???? 1962. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Halperin:1962:GM

[142] Israel Halperin. On the Gram matrix. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 5(??):265–280, ???? 1962. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Mendelsohn:1962:CRI

[143] N. S. Mendelsohn. Congruence relationships for integral recurrences. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 5(??):281–284, ???? 1962. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Fox:1962:NBM

[144] G. Fox. Note on the Borel method of measure extension. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 5(??):285– 296, ???? 1962. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Lambek:1962:CMP

[145] J. Lambek. Correction to my paper “How to program an infinite abacus”. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 5(??):297–304, ???? 1962. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic). See [114].

Kimura:1962:SRP

[146] N. Kimura. Some remarks on a problem of E. L. Whitney. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 5(??): 305–309, ???? 1962. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Harary:1963:CBG

[147] Frank Harary. A characterization of block-graphs. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 6(??):1–6, ???? 1963. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Marcus:1963:ARR

[148] Marvin Marcus. Another remark on a result of K. Goldberg. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 6(??): 7–9, ???? 1963. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Dulmage:1963:RRM

[149] A. L. Dulmage and N. S. Mendelsohn. A remark on a result of Marvin Marcus. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 6(??):11–13, ???? 1963. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Swanson:1963:AVF

[150] C. A. Swanson. Asymptotic variational formulae for eigenvalues. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 6(??): 15–25, ???? 1963. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Scherk:1963:OG

[151] P. Scherk. On ordered geometries. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 6(??):27–36, ???? 1963. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Lipman:1963:OAP

[152] Joe Lipman. Order in affine and projective geometry. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 6(??):37–43, ???? 1963. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Rooney:1963:GIF

[153] P. G. Rooney. A generalization of an inversion formula for the Gauss transformation. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 6(??):45–53, ???? 1963. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Yaqub:1963:GCR

[154] Adil Yaqub. A generalization of certain rings of A. L. Foster. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 6(??): 55–60, ???? 1963. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Weisner:1963:SOL

[155] Louis Weisner. Special orthogonal Latin squares of order 10. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 6(??): 61–63, ???? 1963. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Malik:1963:IP

[156] M. A. Malik. An inequalitiy for polynomials. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 6(??):65–69, ???? 1963. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Duff:1963:IGF

[157] G. F. D. Duff and R. A. Ross. Indefinite Green’s functions and elementary solutions. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 6(??):71–103, ???? 1963. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Whitney:1963:SRP

[158] E. L. Whitney. Sums of reciprocal powers of terms in arithmetic sequence. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 6(??):109–112, ???? 1963. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Moser:1963:NNTa

[159] L. Moser. Notes on number theory. III. On the sum of consecutive primes. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 6(??):159–161, ???? 1963. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Moser:1963:NNTb

[160] L. Moser. Notes on number theory. IV. On the multiplicity of representations of numbers by sums. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 6(??):163–166, ???? 1963. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Moser:1963:NNTc

[161] L. Moser. Notes on number theory. V. Insolvability of \binom{2n}{n} = \binom{2a}{a}\binom{2b}{b}. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 6(??):167–169, ???? 1963. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Lorch:1963:RCM

[162] Lee Lorch and Leo Moser. A remark on completely monotonic sequences, with an application to summability. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 6(??):171–173, ???? 1963. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Moon:1963:TDG

[163] J. W. Moon and L. Moser. Triangular dissections of n-gons. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 6(??): 175–178, ???? 1963. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Lorch:1963:LCS

[164] Lee Lorch and Donald J. Newman. The Lebesgue constants for (γ, r) summation of Fourier series. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 6(??):179–182, ???? 1963. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Dirac:1963:SRC

[165] G. A. Dirac. Some results concerning the structure of graphs. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 6(??): 183–210, ???? 1963. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Ellis:1963:GMS

[166] H. W. Ellis and D. O. Snow. On (L^{1})^{∗} for general measure spaces. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 6(??):211–229, ???? 1963. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Melzak:1963:PFS

[167] Z. A. Melzak. Partition functions and spiralling in plane random walk. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 6(??):231–237, ???? 1963. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Trott:1963:MRN

[168] Stanton M. Trott. A model of the real numbers. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 6(??):239–255, ???? 1963. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

McKiernan:1963:VAA

[169] M. A. McKiernan. Variational aspects of the Abel and Schroder functional equations. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 6(??):257–265, ???? 1963. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Erdos:1963:RPO

[170] Paul Erdős and A. Rényi. Remarks on a problem of Obreanu. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 6(??): 267–273, ???? 1963. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Gould:1963:GBF

[171] H. W. Gould. Generalization of a bracket function formula of L. Moser. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 6(??):275–318, ???? 1963. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Tutte:1963:NBC

[172] W. T. Tutte. The non-biplanar character of the complete 9-graph. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 6 (??):319–330, ???? 1963. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Chalk:1963:SPV

[173] J. H. H. Chalk. Small positive values of indefinite quadratic forms. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 6 (??):331–339, ???? 1963. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Vanstone:1963:HJE

[174] J. R. Vanstone. The Hamilton-Jacobi equations for a relativistic charged particle. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 6(??):341–350, ???? 1963. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Sen:1963:HJE

[175] D. K. Sen. The Hamilton-Jacobi equation for a charged particle in a combined gravitational and electromagnetic field. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 6(??):351–358, ???? 1963. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Gould:1963:GTE

[176] H. W. Gould. A general Turán expression for the zeta function. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 6(??): 359–366, ???? 1963. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Neugebauer:1963:PMI

[177] C. J. Neugebauer. On a paper by M. Iosifescu and S. Marcus. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 6(??): 367–371, ???? 1963. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Barbeau:1963:SDC

[178] A. Barbeau. Sur la structure de deux classes d’anneaux. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 6(??):373– 383, ???? 1963. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Puttaswamaiah:1963:RPR

[179] B. M. Puttaswamaiah. On the reduction of permutation representations. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 6(??):385–395, ???? 1963. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Penry:1963:CTD

[180] M. D. Penry. Composition theorems on Dirichlet series. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 6(??):397– 403, ???? 1963. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Majindar:1963:FPD

[181] Kulendra N. Majindar. On a factorisation of positive definite matrices. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 6(??):405–407, ???? 1963. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Melzak:1963:PPS

[182] Z. A. Melzak. A property of plane sets of constant width. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 6(??): 409–415, ???? 1963. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Sumner:1964:DFC

[183] D. B. Sumner. A distribution function of Cantor-Vitali type. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 7(??): 65–75, ???? 1964. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Duff:1964:PSQ

[184] G. F. D. Duff. On the principle of superposition in quantum mechanics. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 7(??):77–83, ???? 1964. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Gomory:1964:SPR

[185] R. E. Gomory, A. J. Hoffman, and N. C. Hsu. Some properties of the rank and invariant factors of matrices. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 7(??):85–96, ???? 1964. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Bullen:1964:NPI

[186] P. S. Bullen. A new proof of an inequality of Heinz. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 7(??):97–100, ???? 1964. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Sharma:1964:RQH

[187] A. Sharma. Remarks on quasi-Hermite–Fejér interpolation. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 7(??): 101–119, ???? 1964. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Malik:1964:EPD

[188] M. A. Malik. On extremal properties of the derivatives of polynomials and rational functions. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 7(??):121–131, ???? 1964. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Narayana:1964:CNS

[189] T. V. Narayana and D. H. Bent. Computation of the number of score sequences in round-robin tournaments. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 7(??):133–176, ???? 1964. CO-DEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Abbott:1964:CPE

[190] H. L. Abbott and L. Moser. On a combinatorial problem of Erdős and Hajnal. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 7(??):177–181, ???? 1964. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Melzak:1964:IAA

[191] Z. A. Melzak. An informal arithmetical approach to computability and computation. II. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 7(??):183–200, ???? 1964. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Garbe:1964:SRM

[192] Dietmar Garbe and Jens L. Mennicke. Some remarks on the Mathieu groups. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 7(??):201–212, ???? 1964. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic). See corrections [921].

Crowe:1964:GRI

[193] D. W. Crowe. Generating reflections for U(2, p^{2n}). II. p = 2. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 7(??): 213–217, ???? 1964. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Angel:1964:PNN

[194] Myer Angel. Partitions of the natural numbers. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 7(??):219–236, ???? 1964. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Lambek:1964:MFI

[195] J. Lambek. A module is flat if and only if its character module is injective. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 7(??):237–243, ???? 1964. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Doctor:1964:CBL

[196] Hoshang P. Doctor. The categories of Boolean lattices, Boolean rings and Boolean spaces. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 7(??):245–252, ???? 1964. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Minc:1964:PC

[197] Henryk Minc. Permanents of (0, 1)-circulants. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 7(??):253–263, ???? 1964. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Lane:1964:MOC

[198] N. D. Lane, K. K. Singh, and P. Scherk. Monotony of the osculating circles of arcs of cyclic order three. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 7(??):265–271, ???? 1964. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Carlson:1964:GCD

[199] David Carlson and Chandler Davis. A generalization of Cauchy’s double alternant. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 7(??):273–278, ???? 1964. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Howroyd:1964:SSF

[200] T. D. Howroyd. The solutions of some functional equations. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 7(??): 279–282, ???? 1964. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Artiaga:1964:FR

[201] Lucio Artiaga. Finite reciprocities. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 7(??):283–290, ???? 1964. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Zakon:1964:RTL

[202] Elias Zakon. A remark on the theorems of Lusin and Egoroff. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 7(??): 291–295, ???? 1964. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Tiao:1964:SUM

[203] George C. Tiao and Irwin Guttman. Some useful matrix lemmas in statistical estimation theory. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 7(??):297–300, ???? 1964. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic). See addendum [231].

Majindar:1964:ABI

[204] Kulendra N. Majindar. On Abel’s binomial identity. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 7(??):301– 344, ???? 1964. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Abramson:1964:EEC

[205] Morton Abramson. Explicit expressions for a class of permutation problems. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 7(??):345–350, ???? 1964. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Erdos:1964:PT

[206] Paul Erdős and L. Moser. A problem on tournaments. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 7(??):351– 356, ???? 1964. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Trotter:1964:MPC

[207] H. F. Trotter. A machine program for coset enumeration. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 7(??):357– 368, ???? 1964. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Beatty:1964:ECR

[208] J. C. Beatty and R. E. Miller. On equi-cardinal restrictions of a graph. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 7(??):369–375, ???? 1964. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Weisner:1964:RDO

[209] Louis Weisner. A Room design of order 10. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 7(??):377–378, ???? 1964. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Wonenburger:1964:DOT

[210] María J. Wonenburger. A decomposition of orthogonal transformations. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 7(??):379–383, ???? 1964. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Chilton:1964:PRP

[211] B. L. Chilton. On the projection of the regular polytope {5, 3, 3} into a regular triacontagon. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 7(??):385–398, ???? 1964. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Blum:1964:QRB

[212] Richard Blum and A. P. Guinand. A quartic with 28 real bitangents. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 7(??):399–404, ???? 1964. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Osofsky:1964:RPI

[213] B. L. Osofsky. On ring properties of injective hulls. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 7(??):405– 413, ???? 1964. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Kleisli:1964:IS

[214] H. Kleisli and Y. c. Wu. On injective sheaves. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 7(??):415–423, ???? 1964. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Kotze:1964:NNA

[215] Wesley Kotzé. Notes on numerical analysis. IV. On accelerating iteration procedures with superlinear convergence. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 7(??):425–430, ???? 1964. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Rowlinson:1964:NPT

[216] Elizabeth Rowlinson. New proofs for two theorems of Capelli. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 7(??): 431–433, ???? 1964. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Clark:1964:RTC

[217] Colin Clark. On Rellich’s theorem concerning infinitely narrow tubes. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 7(??):435–440, ???? 1964. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Renyi:1964:SDM

[218] A. Rényi. A Socratic dialogue on mathematics. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 7(??):441–462, ???? 1964. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Lawrence:1964:GJT

[219] Peter A. Lawrence. A generalization of Jacobi’s theorem to hyperadjugates. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 7(??):467–469, ???? 1964. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Rohatgi:1964:LPD

[220] V. K. Rohatgi. On lattice paths with diagonal steps. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 7(??):470– 512, ???? 1964. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Erdos:1964:SDP

[221] Paul Erdős. On some divisibility properties of \binom{2n}{n}. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 7(??):513–518, ???? 1964. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Moon:1964:EDC

[222] J. W. Moon. On edge-disjoint cycles in a graph. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 7(??):519–523, ???? 1964. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Pullman:1964:NPE

[223] N. Pullman. On the number of positive entries in the powers of a nonnegative matrix. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 7(??):525–537, ???? 1964. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Stanton:1964:BIB

[224] R. G. Stanton and D. A. Sprott. Block intersections in balanced incomplete block designs. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 7(??):539–548, ???? 1964. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Ostrom:1964:FPP

[225] T. G. Ostrom and F. A. Sherk. Finite projective planes with affine subplanes. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 7(??):549–559, ???? 1964. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Shephard:1964:CM

[226] G. C. Shephard. On a conjecture of Melzak. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 7(??):561–563, ???? 1964. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Blundon:1964:SLB

[227] W. J. Blundon. Some lower bounds for density of multiple packing. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 7(??):565–572, ???? 1964. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Rahman:1964:SIP

[228] Q. I. Rahman. Some inequalities for polynomials and related entire functions. II. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 7(??):573–595, ???? 1964. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Lambek:1964:GTH

[229] Joachim Lambek. Goursat’s theorem and homological algebra. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 7(??): 597–608, ???? 1964. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Leicht:1964:APJ

[230] J. B. Leicht. Axiomatic proof of J. Lambek’s homological theorem. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 7 (??):609–613, ???? 1964. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Tiao:1964:ASU

[231] George C. Tiao and Irwin Guttman. Addendum to “Some useful matrix lemmas in statistical estimation theory”. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 7(??):615, ???? 1964. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic). See [203].

Scherk:1964:EHS

[232] P. Scherk. On the exactness of the homotopy sequence. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 7(??):617– 618, ???? 1964. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Erdos:1964:SRR

[233] Paul Erdős. Some remarks on Ramsay’s theorem. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 7(??):619–622, ???? 1964. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Banaschewski:1964:ETS

[234] Bernhard Banaschewski. Extensions of topological spaces. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 7(1):1–22, January 1964. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Connell:1964:NTP

[235] Ian G. Connell. A number theory problem concerning finite groups and rings. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 7(1):23–34, January 1964. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Erdos:1964:SCB

[236] Paul Erdős and J. W. Moon. On subgraphs of the complete bipartite graph. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 7(1):35–39, January 1964. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Zakon:1964:NRM

[237] Elias Zakon. A note on regular measures. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 7(1):41–44, January 1964. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Zakon:1964:AUC

[238] Elias Zakon. On almost uniform convergence of families of functions. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 7(1):45–48, January 1964. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Moser:1964:RPT

[239] W. O. J. Moser. Remarks on a paper by S. Trott. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 7(1):49–51, January 1964. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic). See [139].

Erdos:1964:NTC

[240] Paul Erdős. On the number of triangles contained in certain graphs. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 7(1):53–56, January 1964. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Csorgo:1964:CTS

[241] M. Csörgö and Irwin Guttman. On the consistency of the two-sample empty cell test. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 7(1):57–63, January 1964. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Wright:1965:SPR

[242] E. M. Wright. Spiralling in plane random walk. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 8(??):1–6, ???? 1965. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Menon:1965:IBG

[243] V. V. Menon. The isomorphism between graphs and their adjoint graphs. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 8(??):7–16, ???? 1965. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Harary:1965:DFK

[244] Frank Harary and William T. Tutte. A dual form of Kuratowski’s theorem. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 8(??):17–20, ???? 1965. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic). See correction and addendum [268].

Schaer:1965:GEP

[245] J. Schaer and A. Meir. On a geometric extremum problem. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 8(??): 21–28, ???? 1965. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Koh:1965:NSI

[246] Kwangil Koh. A note on a self injective ring. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 8(??):29–32, ???? 1965. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

McLaughlin:1965:CCS

[247] T. G. McLaughlin. On a class of complete simple sets. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 8(??):33–38, ???? 1965. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Akbar-Zadeh:1965:ACS

[248] H. Akbar-Zadeh. Sur les automorphismes de certaines structures presque cosymplectiques. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 8(??):39–58, ???? 1965. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Jackson:1965:RBC

[249] H. L. Jackson. On the relation between a cluster set introduced by Constantinescu and Cornea, and the fine cluster set of Cartan, Brelot, and Naiïm. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 8(??):59–72, ???? 1965. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Bruckner:1965:DFH

[250] A. M. Bruckner and John L. Leonard. On differentiable functions having an everywhere dense set of intervals of constancy. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 8(??):73–76, ???? 1965. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Howroyd:1965:CSF

[251] T. D. Howroyd. On certain simultaneous functional equations. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 8(??): 77–82, ???? 1965. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Zakon:1965:ILM

[252] Elias Zakon. On iterated limits of measurable mappings. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 8(??):83–92, ???? 1965. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Csorgo:1965:ELP

[253] Miklós Csörgö. Exact and limiting probability distributions of some Smirnov type statistics. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 8(??):93–104, ???? 1965. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Macphail:1965:RCN

[254] M. S. Macphail. Remark on co-null matrices. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 8(??):105–108, ???? 1965. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Koh:1965:NPR

[255] Kwangil Koh. A note on a prime ring with a maximal annihilator right ideal. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 8(??):109–192, ???? 1965. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Wright:1965:ESL

[256] E. M. Wright. Equal sums of like powers. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 8(??):193–202, ???? 1965. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Bruck:1965:ADL

[257] R. H. Bruck. The algebra of dimension-linking operators. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 8(??):203– 268, ???? 1965. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Erdos:1965:SCA

[258] Paul Erdős and J. W. Moon. On sets of consistent arcs in a tournament. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 8(??):269–272, ???? 1965. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Schaer:1965:DPC

[259] J. Schaer. The densest packing of 9 circles in a square. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 8(??):273– 278, ???? 1965. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Lambek:1965:RQN

[260] J. Lambek. On the ring of quotients of a Noetherian ring. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 8(??): 279–290, ???? 1965. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Tewari:1965:IPL

[261] Krishna Tewari. On the injective and projective limit of complexes. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 8(??):291–306, ???? 1965. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Dawson:1965:LDH

[262] Donald A. Dawson. The local diffusions of a harmonic sheaf. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 8(??): 307–316, ???? 1965. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

McKiernan:1965:GHJ

[263] M. A. McKiernan. A general Hamilton–Jacobi equation and associated problem of Lagrange. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 8(??):317–322, ???? 1965. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Edelstein:1965:ECE

[264] M. Edelstein. An example concerning exposed points. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 8(??):323– 328, ???? 1965. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Ellis:1965:EMF

[265] H. W. Ellis. On extensions of monotone functions from linear sublattices. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 8(??):329–344, ???? 1965. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Pullman:1965:GFM

[266] N. Pullman. The geometry of finite Markov chains. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 8(??):345–358, ???? 1965. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Mohanty:1965:SPC

[267] S. G. Mohanty. Some properties of compositions and their application to the ballot problem. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 8(??):359–372, ???? 1965. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Harary:1965:CAD

[268] F. Harary and W. T. Tutte. Correction and addendum to: “A dual form of Kuratowski’s theorem”. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 8(??):373–374, ???? 1965. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic). See [244].

Marsden:1965:THH

[269] J. E. Marsden. A theorem on harmonic homologies. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 8(??):375– 412, ???? 1965. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Livingston:1965:SP

[270] Arthur E. Livingston. The series ∑_{1}^{∞} f(n)/n for periodic f. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 8(??): 413–432, ???? 1965. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

McKiernan:1965:DGA

[271] M. A. McKiernan. A differential geometry associated with dissipative systems. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 8(??):433–452, ???? 1965. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Burton:1965:SGI

[272] T. A. Burton. Some general instability criteria for ordinary differential equations. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 8(??):453–458, ???? 1965. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Takahashi:1965:CGR

[273] Shuichi Takahashi. A characterization of group rings as a special class of Hopf algebras. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 8(??):465–476, ???? 1965. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Lindstrom:1965:CPN

[274] Bernt Lindström. On a combinatorial problem in number theory. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 8(??): 477–490, ???? 1965. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Beineke:1965:MNS

[275] Lowell W. Beineke and Frank Harary. The maximum number of strongly connected subtournaments. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 8(??):491–498, ???? 1965. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Hrycay:1965:VLV

[276] R. Hrycay. On vector lattice-valued measures. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 8(??):499–504, ???? 1965. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Sanderson:1965:GDI

[277] D. F. Sanderson. A generalization of divisibility and injectivity in modules. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 8(??):505–514, ???? 1965. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Abbott:1965:ARR

[278] H. L. Abbott. An application of Ramsay’s [Ramsey’s] theorem to a problem of Erdős and Hajnal. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 8(??):515–517, ???? 1965. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Potts:1965:ASL

[279] D. H. Potts. Axioms for semi-lattices. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 8(??):519–564, ???? 1965. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Melzak:1965:PCC

[280] Z. A. Melzak. Problems connected with convexity. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 8(??):565–574, ???? 1965. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Kalbfleisch:1965:CSE

[281] J. G. Kalbfleisch. Construction of special edge-chromatic graphs. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 8(??): 575–584, ???? 1965. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Abramson:1965:RC

[282] M. Abramson. Restricted choices. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 8(??):585–600, ???? 1965. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Ostrom:1965:DN

[283] T. G. Ostrom. Derivable nets. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 8(??):601–614, ???? 1965. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Blundon:1965:IAT

[284] W. J. Blundon. Inequalities associated with the triangle. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 8(??):615– 626, ???? 1965. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Bercov:1965:APG

[285] R. Bercov and L. Moser. On Abelian permutation groups. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 8(??): 627–630, ???? 1965. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Fox:1965:FDS

[286] Charles Fox. A family of distributions with the same ratio property as normal distribution. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 8(??):631–636, ???? 1965. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Govil:1965:SCD

[287] Narendra K. Govil. On the summability of a class of the derived conjugate series of a Fourier series. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 8(??):637–646, ???? 1965. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Burton:1965:SBO

[288] T. A. Burton. Stability on the basis of orthogonal trajectories. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 8(??): 647–658, ???? 1965. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Wong:1965:SRT

[289] James S. W. Wong. Some remarks on transformations in metric spaces. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 8(??):659–666, ???? 1965. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Koh:1965:SRM

[290] Kwangil Koh. On simple rings with maximal annihilator right ideals. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 8(??):667–700, ???? 1965. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Harary:1965:EHG

[291] Frank Harary and C. St. J. A. Nash-Williams. On Eulerian and Hamiltonian graphs and line graphs. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 8(??):701–710, ???? 1965. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Dirac:1965:CNT

[292] G. A. Dirac. Chromatic number and topological complete subgraphs. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 8(??):711–716, ???? 1965. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Pearson:1965:NHL

[293] T. L. Pearson. Note on the Hardy–Landau summation formula. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 8(??): 717–720, ???? 1965. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Eliopoulos:1965:GTD

[294] Hermes A. Eliopoulos. On the general theory of differentiable manifolds with almost tangent structure. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 8(??):721–748, ???? 1965. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Connell:1965:EGH

[295] Ian G. Connell. Elementary generalizations of Hilbert’s Theorem 90. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 8(??):749–758, ???? 1965. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Burgess:1965:MME

[296] W. Burgess. The meaning of mono and epi in some familiar categories. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 8(??):759–770, ???? 1965. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Taylor:1965:WFM

[297] J. C. Taylor. Weak families of maps. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 8(??):771–782, ???? 1965. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Datko:1965:GBB

[298] Richard Datko. A generalization of the bang-bang principle of linear control theory. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 8(??):783–790, ???? 1965. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Wong:1965:RUT

[299] James S. W. Wong. Remarks on the uniqueness theorem of solutions of the Darboux problem. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 8(??):791–796, ???? 1965. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Forbes:1965:LCR

[300] R. L. Forbes. Lebesgue constants for regular Taylor summability. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 8(??): 797–808, ???? 1965. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Ellis:1965:D

[301] H. W. Ellis. On the dual of L^{1}. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 8(??):809–818, ???? 1965. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Seshadri:1965:RVW

[302] V. Seshadri. On random variables which have the same distribution as their reciprocals. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 8(??):819–824, ???? 1965. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Trotter:1965:GWR

[303] H. F. Trotter. Groups in which raising to a power is an automorphism. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 8(??):825–828, ???? 1965. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Glass:1965:RCP

[304] A. S. Glass. A remark on convex polytopes. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 8(??):829–830, ???? 1965. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Moon:1965:NCS

[305] J. W. Moon. On the number of complete subgraphs of a graph. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 8(??): 831–834, ???? 1965. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Edelstein:1965:PRH

[306] M. Edelstein. On a problem by R. A. Hirschfeld. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 8(??):835, ???? 1965. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Brown:1965:NPT

[307] William G. Brown. A new proof of a theorem of Dirac. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 8(4):459–464, June 1965. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Brauer:1966:SCS

[308] George Brauer. Sets of convergence for series defined by iteration. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 9 (??):83–88, ???? 1966. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Datko:1966:GSD

[309] R. Datko. Global stability determined by local properties and the first variation. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 9(??):89–94, ???? 1966. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Mathai:1966:SCO

[310] A. M. Mathai. Some characterizations of the one-parameter family of probability distributions. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 9(??):95–102, ???? 1966. CODEN CM-BUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Fox:1966:BTP

[311] A. S. Fox. On the boundary and tensor product of function algebras. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 9(??):103–106, ???? 1966. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Simpson:1966:CPL

[312] J. E. Simpson. On convergence of projections in locally convex spaces. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 9(??):107–110, ???? 1966. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Bowen:1966:SVP

[313] Rufus Bowen. On sums of valencies in planar graphs. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 9(??):111– 142, ???? 1966. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Subbarao:1966:AFS

[314] M. V. Subbarao. Arithmetic functions satisfying a congruence property. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 9(??):143–146, ???? 1966. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Subbarao:1966:UPN

[315] M. V. Subbarao and L. J. Warren. Unitary perfect numbers. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 9(??): 147–154, ???? 1966. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Abbott:1966:SRC

[316] H. L. Abbott. Some remarks on a combinatorial theorem of Erdős and Rado. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 9(??):155–160, ???? 1966. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Mehra:1966:SIR

[317] K. L. Mehra and J. S. W. Wong. Some inequalities related to the Wald– Wolfowitz–Noether condition. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 9(??):161–170, ???? 1966. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Chartrand:1966:TD

[318] Gary Chartrand and M. James Stewart. Total digraphs. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 9(??):171– 176, ???? 1966. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Sabidussi:1966:SEG

[319] Gert Sabidussi. Sequences of Euler graphs. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 9(??):177–182, ???? 1966. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Banaschewski:1966:IHC

[320] B. Banaschewski. On the injective hulls of cyclic modules over Dedekind domains. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 9(??):183–186, ???? 1966. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Tiwary:1966:QII

[321] A. K. Tiwary. On the quotients of indecomposable injective modules. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 9(??):187–190, ???? 1966. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Koh:1966:VLO

[322] Kwangil Koh. On very large one sided ideals of a ring. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 9(??):191– 196, ???? 1966. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

McWorter:1966:SSP

[323] W. A. McWorter. Some simple properties of simple nil rings. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 9(??): 197–200, ???? 1966. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Sharpe:1966:UFT

[324] R. Sharpe, M. Beattie, and J. Marsden. A universal factorization theorem in topology. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 9(??):201–208, ???? 1966. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Saxena:1966:GFJ

[325] R. K. Saxena. A generating function for Jacobi polynomials. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 9(??): 209–214, ???? 1966. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Saxena:1966:RAF

[326] R. K. Saxena. On the reducibility of Appell’s function F_{4}. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 9(??): 215–222, ???? 1966. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Blum:1966:RCM

[327] Richard Blum. A remarkable class of Mannheim-curves. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 9(??):223– 228, ???? 1966. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Aczel:1966:SMC

[328] J. Aczél. On strict monotonicity of continuous solutions of certain types of functional equations. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 9(??):229–232, ???? 1966. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Jackson:1966:SRA

[329] H. L. Jackson. Some remarks on angular derivatives and Julia’s lemma. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 9(??):233–264, ???? 1966. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Schaer:1966:DPSa

[330] J. Schaer. On the densest packing of spheres in a cube. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 9(??):265– 270, ???? 1966. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Schaer:1966:DPF

[331] J. Schaer. The densest packing of five spheres in a cube. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 9(??):271– 274, ???? 1966. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Schaer:1966:DPSb

[332] J. Schaer. The densest packing of six spheres in a cube. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 9(??):275– 280, ???? 1966. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Brown:1966:GDC

[333] W. G. Brown. On graphs that do not contain a Thomsen graph. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 9(??): 281–286, ???? 1966. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Davison:1966:AC

[334] T. M. K. Davison. On arithmetic convolutions. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 9(??):287–296, ???? 1966. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Moon:1966:ST

[335] J. W. Moon. On subtournaments of a tournament. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 9(??):297–302, ???? 1966. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Moser:1966:ETS

[336] L. Moser and R. A. MacLeod. The error term for the squarefree integers. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 9(??):303–306, ???? 1966. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Zakon:1966:INM

[337] Elias Zakon. Integration of non-measurable functions. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 9(??):307– 330, ???? 1966. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Byers:1966:UVM

[338] William Byers. Unbounded vector measures. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 9(??):331–342, ???? 1966. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Shimrat:1966:TN

[339] M. Shimrat. A theorem on nets. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 9(??):343–346, ???? 1966. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Church:1966:LMA

[340] C. A. Church, Jr. On a lemma of M. Abramson. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 9(??):347–348, ???? 1966. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Naimpally:1966:FST

[341] Somashekhar Amrith Naimpally. Function space topologies for connectivity and semi-connectivity functions. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 9(??):349–352, ???? 1966. CO-DEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Murdeshwar:1966:SHS

[342] M. G. Murdeshwar and S. A. Naimpally. Semi-Hausdorff spaces. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 9(??): 353–356, ???? 1966. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Padmanabhan:1966:ASL

[343] R. Padmanabhan. On axioms for semi-lattices. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 9(??):357–406, ???? 1966. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Beattie:1966:OFA

[344] Mary Beattie, J. E. Marsden, and R. W. Sharpe. Order in finite affine planes. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 9(??):407–412, ???? 1966. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Bercov:1966:GAC

[345] R. Bercov. On groups with all composition factors isomorphic. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 9(??): 413–416, ???? 1966. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Tully:1966:STA

[346] E. J. Tully, Jr. On semigroups of transformations acting transitively on a set. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 9(??):417–420, ???? 1966. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Carlitz:1966:SSF

[347] L. Carlitz and L. Moser. On some special factorizations of (1−x^{n})/(1−x). Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 9(??):421–426, ???? 1966. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Gioia:1966:GFC

[348] A. A. Gioia and M. V. Subbarao. Generating functions for a class of arithmetic functions. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 9(??):427–432, ???? 1966. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Elliott:1966:LNS

[349] George Elliott and Israel Halperin. Linear normed spaces with extension property. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 9(??):433–442, ???? 1966. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Schirmer:1966:BTC

[350] Helga Schirmer. A Brouwer type coincidence theorem. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 9(??):443– 446, ???? 1966. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Heinig:1966:NIR

[351] H. P. Heinig. A new inversion and representation theory for the Laplace transform. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 9(??):447–456, ???? 1966. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Segal:1966:NDC

[352] S. L. Segal. A note on Dirichlet convolutions. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 9(??):457–462, ???? 1966. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Scherk:1966:CT

[353] P. Scherk. On Carathéodory’s theorem. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 9(??):463–466, ???? 1966. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Rahman:1966:PEF

[354] Q. I. Rahman. A property of entire functions of exponential type. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 9(??): 467–472, ???? 1966. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Haruki:1966:FEI

[355] Hiroshi Haruki. On the functional equations |f(x + iy)| = |f(x) + f(iy)| and |f(x + iy)| = |f(x) − f(iy)| and on Ivory’s theorem. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 9(??):473– 480, ???? 1966. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Brauer:1966:SNH

[356] Fred Brauer. The solution of non-homogeneous systems of differential equations by undetermined coefficients. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 9(??):481–488, ???? 1966. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

McKiernan:1966:FSE

[357] M. A. McKiernan. Fatigue spaces in electromagnetic-gravitational theory. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 9(??):489–508, ???? 1966. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

McGillivray:1966:CDN

[358] W. R. McGillivray and C. L. Kaller. A characterization of deviation from normality under certain moment assumptions. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 9(??):509–514, ???? 1966. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Bassett:1966:CT

[359] Paul G. Bassett. A combinatorial theorem. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 9(??):515–516, ???? 1966. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Kleiner:1966:CCD

[360] Israel Kleiner. A counterexample to a conjecture of D. F. Sanderson. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 9(??):517–518, ???? 1966. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Eisen:1966:GET

[361] M. Eisen. A global existence theorem. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 9(??):519–520, ???? 1966. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Murdeshwar:1966:TS

[362] M. G. Murdeshwar and S. A. Naimpally. R_{1}-topological spaces. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 9(??): 521–524, ???? 1966. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Murdeshwar:1966:CTB

[363] M. G. Murdeshwar. On closed, totally bounded sets. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 9(??):525– 526, ???? 1966. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Gagnon:1966:LPC

[364] F. P. Gagnon. Limites projectives de chaïnes de demi-groupes nilpotents. (French) [Projective limits of chaines of nilpotent demi-groups]. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 9(??): 527–556, ???? 1966. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Abbott:1966:HCP

[365] H. L. Abbott. Hamiltonian circuits and paths on the n-cube. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 9(??): 557–562, ???? 1966. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Goldberg:1966:MOG

[366] Myron Goldberg and J. W. Moon. On the maximum order of the group of a tournament. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 9(??):563–570, ???? 1966. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Subbarao:1966:CCA

[367] M. V. Subbarao. A congruence for a class of arithmetic functions. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 9 (??):571–574, ???? 1966. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Williams:1966:ECA

[368] Kenneth S. Williams. Eisenstein’s criteria for absolute irreducibility over a finite field. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 9(??):575–580, ???? 1966. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Maranda:1966:FCA

[369] J.-M. Maranda. On fundamental constructions and adjoint functors. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 9(??):581–592, ???? 1966. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Melzak:1966:IAA

[370] Z. A. Melzak. An informal arithmetical approach to computability and computation. III. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 9(??):593–610, ???? 1966. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Derderian:1966:DGS

[371] J.-C. Derdérian and G. Thierrin. Demi-groupes séparateurs. (French) [Separating demi-groups]. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 9(??):611–620, ???? 1966. CODEN CM-BUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Houh:1966:NAP

[372] Chorng shi Houh. Note on almost product manifolds and their tangent bundles. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 9(??):621–630, ???? 1966. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Morrison:1966:DCS

[373] D. Morrison, M. Kiang, and J. Wright. On disjoint connected subsets of a square containing pairs of antipodal ponts. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 9(??):631–638, ???? 1966. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Lorch:1966:TSB

[374] Lee Lorch. Translativity for strong Borel summability. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 9(??):639– 646, ???? 1966. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Govil:1966:SST

[375] Narendra K. Govil. On the strong summability by triangular means of the derived Fourier series and its conjugate series. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 9(??):647–654, ???? 1966. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Saxena:1966:SII

[376] R. B. Saxena. Stability of interpolation on an infinite interval. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 9(??): 655–666, ???? 1966. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Melzak:1966:IIT

[377] Z. A. Melzak. An isoperimetric inequality for tetrahedra. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 9(??):667– 670, ???? 1966. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Pears:1966:EEH

[378] A. R. Pears. On the exactness of the Eckmann–Hilton homotopy sequence. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 9(??):671–674, ???? 1966. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Abramson:1966:NC

[379] M. Abramson and W. Moser. A note on combinations. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 9(??):675– 756, ???? 1966. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Kahan:1966:NLA

[380] W. M. Kahan. Numerical linear algebra. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 9(??):757–801, ???? 1966. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Balomenos:1966:EST

[381] Richard H. Balomenos, William E. Bonnice, and Robert J. Silverman. Extensions of Sylvester’s theorem. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 9(1):1–14, ???? 1966. CODEN CM-BUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Armstrong:1966:CSC

[382] William W. Armstrong. Curves on surfaces of constant width. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 9(1):15– 22, ???? 1966. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Srinivasacharyulu:1966:TCM

[383] K. Srinivasacharyulu. Topology of complex manifolds. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 9(1):23–28, ???? 1966. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Kleiner:1966:FIL

[384] Israel Kleiner. Free and injective Lie modules. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 9(1):29–42, ???? 1966. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Kuyk:1966:NGS

[385] Willem Kuyk. Note on generalized Schreier extensions of groups. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 9(1):43– 48, ???? 1966. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Bonyun:1966:SHM

[386] Rosemary Bonyun. On the square of a homological monoid. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 9(1):49– 56, ???? 1966. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Nobusawa:1966:LMM

[387] Nobuo Nobusawa. On lattices in a module over a matrix algebra. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 9(1):57– 62, ???? 1966. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Koh:1966:CPR

[388] Kwangil Koh and A. C. Mewborn. A class of prime rings. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 9(1):63– 72, ???? 1966. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Magill:1966:ASA

[389] Kenneth D. Magill, Jr. Automorphisms of the semigroup of all relations on a set. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 9(1):73–78, ???? 1966. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Simons:1966:P

[390] S. Simons. A proof that SouslinSouslinH ⊂ SouslinH. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 9(1):79–82, ???? 1966. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Melzak:1967:NBC

[391] Z. A. Melzak. A note on the Borsuk conjecture. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 10(??):1–4, ???? 1967. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Seip:1967:CF

[392] Ulrich Seip. Categorical fibrations. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 10(??):5–18, ???? 1967. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Feller:1967:TQF

[393] Edmund H. Feller. A type of quasi-Frobenius ring. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 10(??):19–28, ???? 1967. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

McCann:1967:RCP

[394] K. McCann and K. S. Williams. On the residues of a cubic polynomial (modp). Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 10(??):29–38, ???? 1967. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Retherford:1967:SC

[395] J. R. Retherford. Some characterizations of c_{0} and l^{1}. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 10(??):39– 52, ???? 1967. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Joyal:1967:LZP

[396] A. Joyal, G. Labelle, and Q. I. Rahman. On the location of zeros of polynomials. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 10(??):53–64, ???? 1967. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Subbarao:1967:IMF

[397] M. V. Subbarao and A. A. Gioia. Identities for multiplicative functions. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 10(??):65–74, ???? 1967. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Kemp:1967:PS

[398] R. R. D. Kemp. On periodic solutions of x′′′ + ax′′ + bx′ + g(x) = 0. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 10(??):75–78, ???? 1967. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Govil:1967:SCD

[399] Narendra K. Govil. On the summability of a class of derived Fourier series. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 10(??):79–84, ???? 1967. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Cross:1967:ICS

[400] George Cross. An integral for Cesàro summable series. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 10(??):85–98, ???? 1967. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Zakon:1967:UAA

[401] Elias Zakon. On uniform approximations of abstract functions. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 10(??): 99–108, ???? 1967. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Grossman:1967:NPC

[402] Marvin W. Grossman. Note on pointwise convergence on the Choquet boundary. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 10(??):109–114, ???? 1967. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Wong:1967:DAO

[403] James S. W. Wong. A discrete analogue of Opial’s inequality. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 10(??): 115–118, ???? 1967. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Brown:1967:FPS

[404] W. G. Brown. On finite plane sets containing for every pair of points an equidistant point. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 10(??):119–172, ???? 1967. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Abbott:1967:EPN

[405] H. L. Abbott and B. Gardner. An extremal problem in number theory. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 10(??):173–178, ???? 1967. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Rahman:1967:PSP

[406] Q. I. Rahman and Mohd. Ali Khan. Polynomials with some prescribed zeros. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 10(??):179–190, ???? 1967. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Datko:1967:IDE

[407] R. Datko. Integrals of the differential equations ẍ + f(s)x = 0. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 10(??): 191–196, ???? 1967. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Fischer:1967:SNG

[408] P. Fischer and Gy. Muszély. On some new generalizations of the functional equation of Cauchy. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 10(??):197–206, ???? 1967. CODEN CM-BUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Ellis:1967:MDF

[409] H. W. Ellis and R. L. Jeffery. On measures determined by functions with finite right and left limits everywhere. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 10(??):207–226, ???? 1967. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Bruckner:1967:CCF

[410] A. M. Bruckner. On characterizing classes of functions in terms of associated sets. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 10(??):227–232, ???? 1967. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Hoo:1967:SRP

[411] C. S. Hoo. Some remarks on a paper of D. W. Kahn. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 10(??):233–238, ???? 1967. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Mathai:1967:SPC

[412] A. M. Mathai. On the structural properties of the conditional distributions. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 10(??):239–246, ???? 1967. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Marsden:1967:CPM

[413] J. E. Marsden. A correspondence principle for momentum operators. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 10(??):247–250, ???? 1967. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Srinivasacharyulu:1967:CHM

[414] K. Srinivasacharyulu. On complex homogeneous manifolds. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 10(??): 251–256, ???? 1967. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Baggs:1967:SRB

[415] Ivan Baggs. Some relationships between filters. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 10(??):257–260, ???? 1967. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Berthiaume:1967:IES

[416] P. Berthiaume. The injective envelope of S-sets. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 10(??):261–274, ???? 1967. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Osofsky:1967:NTR

[417] B. L. Osofsky. A non-trivial ring with non-rational injective hull. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 10(??): 275–282, ???? 1967. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Alspach:1967:CEL

[418] Brian Alspach. Cycles of each length in regular tournaments. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 10(??): 283–286, ???? 1967. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Stanton:1967:PP

[419] R. G. Stanton, H. C. Williams, and C. R. Zarnke. A packing problem. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 10(??):287–332, ???? 1967. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Heinig:1967:RTL

[420] H. P. Heinig. Representation theories for the Laplace transform. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 10(??): 333–346, ???? 1967. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Ustina:1967:HMD

[421] F. Ustina. The Hausdorff means for double sequences. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 10(??):347–352, ???? 1967. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Bullen:1967:TP

[422] P. S. Bullen. On a theorem of Privaloff. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 10(??):353–360, ???? 1967. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Husain:1967:CRA

[423] T. Husain and T. D. Dwivedi. A characterization of real almost continuous functions. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 10(??):361–364, ???? 1967. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Kaye:1967:RTP

[424] S. M. Kaye. Ring theoretic properties of matrix rings. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 10(??):365–374, ???? 1967. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Richman:1967:MSI

[425] Fred Richman. Maximal subgroups of infinite symmetric groups. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 10(??): 375–382, ???? 1967. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Gandhi:1967:EFT

[426] Chinthayamma Gandhi and J. M. Gandhi. An extension of Fermat’s theorem. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 10(??):383–386, ???? 1967. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Leech:1967:FDN

[427] John Leech. Five dimensional non-lattice sphere packings. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 10(??):387– 394, ???? 1967. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Wilker:1967:ODP

[428] John B. Wilker. Open disk packings of a disk. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 10(??):395–416, ???? 1967. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Schirmer:1967:CPM

[429] Helga Schirmer. Coincidence producing maps onto trees. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 10(??):417– 424, ???? 1967. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Harary:1967:WGO

[430] Frank Harary, Edgar Palmer, and Cedric Smith. Which graphs have only self-converse orientations? Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 10(??):425–430, ???? 1967. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Cockayne:1967:SP

[431] E. J. Cockayne. On the Steiner problem. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 10(??):431–450, ???? 1967. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Rao:1967:RS

[432] B. V. Rao. A remark on R_{1} spaces. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 10(??):451–452, ???? 1967. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Lane:1967:NAP

[433] N. D. Lane. A note on affine Pappus conditions. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 10(??):453–458, ???? 1967. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Putz:1967:EOR

[434] Randee Putz. An estimate for the order of rational matrices. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 10(??): 459–462, ???? 1967. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Malik:1967:EDQ

[435] Henrick John Malik. Exact distribution of the quotient of independent generalized gamma variables. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 10(??):463–492, ???? 1967. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Guy:1967:ML

[436] Richard K. Guy and Frank Harary. On the Möbius ladders. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 10(??): 493–496, ???? 1967. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Heap:1967:NCM

[437] B. R. Heap and M. S. Lynn. A note on p-cyclic matrices and digraphs. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 10(??):497–502, ???? 1967. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Dixon:1967:MOG

[438] John D. Dixon. The maximum order of the group of a tournament. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 10(??):503–506, ???? 1967. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Kaul:1967:IFP

[439] S. K. Kaul. On irregular fixed points. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 10(??):507–512, ???? 1967. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Chan:1967:CPM

[440] L. K. Chan. On the convergence of product moments. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 10(??):513–518, ???? 1967. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Eames:1967:MCF

[441] W. Eames and L. E. May. Measurable cover functions. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 10(??):519–524, ???? 1967. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Fox:1967:EBV

[442] Geoffrey Fox. Extension of a bounded vector measure with values in a reflexive Banach space. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 10(??):525–530, ???? 1967. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Varma:1967:PPT

[443] A. K. Varma. On a problem of P. Turán on lacunary interpolation. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 10(??):531–558, ???? 1967. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Prasad:1967:SIP

[444] J. Prasad and R. B. Saxena. Some interpolatory properties of Laguerre polynomials. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 10(??):559–572, ???? 1967. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Williams:1967:TN

[445] Kenneth S. Williams. On a theorem of Niven. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 10(??):573–578, ???? 1967. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic). See addendum [491].

Williams:1967:GP

[446] Kenneth S. Williams. On general polynomials. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 10(??):579–584, ???? 1967. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Williams:1967:EP

[447] Kenneth S. Williams. On extremal polynomials. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 10(??):585–594, ???? 1967. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Koh:1967:SZD

[448] Kwangil Koh. On the set of zero divisors of a topological ring. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 10(??): 595–596, ???? 1967. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Charlton:1967:NAP

[449] H. J. Charlton. A note on arc-preserving functions for manifolds. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 10(??): 597–598, ???? 1967. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Welsh:1967:DM

[450] D. J. A. Welsh. On dependence in matroids. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 10(??):599–634, ???? 1967. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Banaschewski:1967:IGR

[451] B. Banaschewski. Integral group rings of finite groups. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 10(??):635–642, ???? 1967. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Thierrin:1967:QCD

[452] G. Thierrin. Quelques caractérisations du radical d’un anneau. (French) [Some characterizations of the radical of a ring]. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 10(??):643–648, ???? 1967. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Nobusawa:1967:STM

[453] Nobuo Nobusawa. Structure theorems of a module over a ring with a bilinear mapping. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 10(??):649–652, ???? 1967. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Stocks:1967:LPD

[454] D. R. Stocks, Jr. Lattice paths in E^{3} with diagonal steps. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 10(??): 653–658, ???? 1967. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Peterson:1967:FCS

[455] Bruce B. Peterson. Formal contraction of the n-simplex. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 10(??):659– 664, ???? 1967. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Hoo:1967:NHC

[456] C. S. Hoo. A note on the homotopy-commutativity of suspensions. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 10 (??):665–668, ???? 1967. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Chalk:1967:NCC

[457] J. H. H. Chalk. A note on the coverings of E^{n} by convex sets. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 10(??): 669–674, ???? 1967. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Kartsatos:1967:RBB

[458] A. G. Kartsatos. On the relation between boundedness and oscillation of differential equations of second order. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 10(??):675–680, ???? 1967. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Lalli:1967:SSC

[459] B. S. Lalli. On stability of solutions of certain differential equations of the third order. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 10(??):681–688, ???? 1967. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Taylor:1967:ACI

[460] J. C. Taylor. Applying the Cauchy integral. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 10(??):689–694, ???? 1967. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Adams:1967:NFT

[461] Robert A. Adams. A note on Fourier transforms and imbedding theorems. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 10(??):695–698, ???? 1967. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Mitra:1967:TGF

[462] S. C. Mitra. On a theorem in the generalised Fourier transform. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 10(??): 699–710, ???? 1967. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Heinig:1967:RFW

[463] H. P. Heinig. Representation of functions as Weierstrass-transforms. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 10 (??):711–722, ???? 1967. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Rahman:1967:MCF

[464] Q. I. Rahman. On the monotonicity of certain functionals in the theory of analytic functions. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 10(??):723–730, ???? 1967. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Clark:1967:RET

[465] Colin Clark. Rellich’s embedding theorem for a “spiny urchin”. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 10(??): 731–734, ???? 1967. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Sabat:1967:NAF

[466] J. B. Sabat. Note on an application of the δ-function in the representation of solutions of algebraic equations. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 10(??):735–738, ???? 1967. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Csorgo:1967:SLT

[467] Miklós Csörgö. On some limit theorems involving the empirical distribution function. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 10(??):739–742, ???? 1967. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Dunham:1967:NAA

[468] C. Dunham. Necessity of alter(n)ation. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 10(??):743–744, ???? 1967. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Sinkhorn:1967:CMF

[469] Richard Sinkhorn. Concerning the minimum function of a stochastic matrix. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 10(??):745–748, ???? 1967. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Segal:1967:FFG

[470] S. L. Segal. Footnote to a formula of Gioia and Subbarao. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 10(??): 749–750, ???? 1967. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Chinthayamma:1967:NG

[471] Chinthayamma and J. M. Gandhi. On numbers generated by e^{s(e^{x}−1)}. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 10(??):751–754, ???? 1967. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Nobusawa:1967:DIS

[472] Nobuo Nobusawa. On the distribution of integer solutions of f(x, y) = z^{2} for a definite binary quadratic form f. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 10(??):755–756, ???? 1967. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Koh:1967:NFO

[473] Kwangil Koh. A note on a fully ordered ring. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 10(??):757–758, ???? 1967. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Coxeter:1968:PA

[474] H. S. M. Coxeter. The problem of Appolonius. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 11(??):1–18, ???? 1968. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Fleischer:1968:NCI

[475] Isidore Fleischer. A new construction of the injective hull. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 11(??): 19–22, ???? 1968. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Heinicke:1968:NLR

[476] A. G. Heinicke. A note on lower radical constructions for associative rings. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 11(??):23–30, ???? 1968. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Botta:1968:CDP

[477] Peter Botta. On the conversion of the determinant into the permanent. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 11(??):31–34, ???? 1968. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Malone:1968:NRH

[478] Joseph J. Malone, Jr. Near-ring homomorphisms. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 11(??):35–42, ???? 1968. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Lightstone:1968:GTP

[479] A. H. Lightstone. Group theory and the principle of duality. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 11(??): 43–50, ???? 1968. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

King:1968:CPL

[480] J. P. King. A class of positive linear operators. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 11(??):51–60, ???? 1968. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Struble:1968:OD

[481] Raimond A. Struble. On operators and distributions. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 11(??):61– 64, ???? 1968. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Strang:1968:NS

[482] Gilbert Strang. The nucleus of a set. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 11(??):65–72, ???? 1968. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Lee:1968:DAI

[483] Cheng ming Lee. On a discrete analogue of inequalities of Opial and Yang. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 11(??):73–78, ???? 1968. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Williams:1968:NPC

[484] Kenneth S. Williams. Note on pairs of consecutive residues of polynomials. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 11(??):79–84, ???? 1968. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Melzak:1968:NHD

[485] Z. A. Melzak. A note on homogeneous dendrites. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 11(??):85–94, ???? 1968. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Bonnett:1968:MCT

[486] D. A. Bonnett and J. R. Porter. Matrix characterizations of topological properties. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 11(??):95–106, ???? 1968. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Cleroux:1968:ESE

[487] Robert Cléroux. On the evaluation of some expressions concerning the mP BIB designs. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 11(??):107–114, ???? 1968. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Closs:1968:RAH

[488] Mike P. Closs. On real almost Hermitian structures subordinate to almost tangent structures. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 11(??):115–134, ???? 1968. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Singh:1968:NEE

[489] Rajinder Singh. Non-existence of estimates of prescribed accuracy in fixed sample size. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 11(??):135–140, ???? 1968. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Goebel:1968:MLC

[490] K. Goebel. On the minimal Lipschitz constant. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 11(??):141–144, ???? 1968. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Williams:1968:ATN

[491] Kenneth S. Williams. Addendum: “On a theorem of Niven”. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 11(??): 145–174, ???? 1968. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic). See [445].

Smith:1968:CPA

[492] R. A. Smith. The circle problem in an arithmetic progression. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 11(??): 175–184, ???? 1968. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Endler:1968:NHV

[493] Otto Endler. A note on Henselian valuation rings. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 11(??):185–190, ???? 1968. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

OShaughnessy:1968:RDO

[494] C. D. O’Shaughnessy. A room design of order 14. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 11(??):191–194, ???? 1968. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Kapoor:1968:DG

[495] S. F. Kapoor, H. V. Kronk, and D. R. Lick. On detours in graphs. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 11(??):195–202, ???? 1968. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Hedetniemi:1968:PPG

[496] Stephen Hedetniemi. On partitioning planar graphs. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 11(??):203–212, ???? 1968. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Ward:1968:NCE

[497] L. E. Ward. On the non-cutpoint existence theorem. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 11(??):213–216, ???? 1968. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Manohar:1968:MRQ

[498] R. Manohar and C. Turnbull. On a modification of Romberg quadrature. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 11(??):217–224, ???? 1968. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Leviatan:1968:MPQ

[499] Dany Leviatan. Moment problems and quasi-Hausdorff transformations. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 11(??):225–236, ???? 1968. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Ahmad:1968:PRZ

[500] Mansoor Ahmad. On polynomials with real zeros. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 11(??):237–240, ???? 1968. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Verma:1968:TCP

[501] Arun Verma. A transformation connecting products of generalised basic hypergeometric functions. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 11(??):241–248, ???? 1968. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Ditzian:1968:GFW

[502] Z. Ditzian. Gap formulae for the Weierstrass transformations. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 11(??): 249–254, ???? 1968. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Dayal:1968:BLF

[503] S. Dayal. On Banach limit of Fourier series and conjugate series. I. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 11(??):255–262, ???? 1968. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Kabe:1968:SDP

[504] D. G. Kabe. Some distribution problems of order statistics from exponential and power function distributions. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 11(??):263–274, ???? 1968. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Sankaran:1968:THR

[505] N. Sankaran. A theorem on Henselian rings. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 11(??):275–278, ???? 1968. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Williams:1968:EP

[506] Kenneth S. Williams. On exceptional polynomials. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 11(??):279–282, ???? 1968. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Djokovic:1968:NTS

[507] D. Z. Djoković. Note on a theorem on singular matrices. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 11(??):283– 284, ???? 1968. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Murty:1968:DM

[508] U. S. R. Murty. Diameter of a (0, , 1)-matrix. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 11(??):285–288, ???? 1968. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Entringer:1968:R

[509] R. C. Entringer. Representation of m as ∑_{k=−n}^{n} ε_{k}k. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 11(??):289–294, ???? 1968. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Karaganis:1968:CG

[510] Jerome J. Karaganis. On the cube of a graph. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 11(??):295–296, ???? 1968. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Struble:1968:GTF

[511] Raimond A. Struble. A genuine topology for the field of Mikusiński operators. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 11(??):297–300, ???? 1968. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Ustina:1968:RGP

[512] Fred Ustina. A remark on the Gibbs phenomenon and Lebesgue constants for a summability method of Melikov. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 11(??):301–366, ???? 1968. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Ayoub:1968:CAA

[513] Christine Ayoub. Constructing an automorphism from an antiautomorphism. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 11(??):367–370, ???? 1968. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Tang:1968:NSE

[514] C. Y. Tang. Notes on splitting extensions of groups. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 11(??):371–374, ???? 1968. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic). See correction [657].

Bell:1968:DRS

[515] Howard E. Bell. Duo rings: Some applications to commutativity theorems. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 11(??):375–380, ???? 1968. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Maxson:1968:NCF

[516] Carlton J. Maxson. A new characterization of finite prime fields. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 11(??): 381–382, ???? 1968. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Kleiner:1968:RQR

[517] Israel Kleiner. Rings of quotients of rings of derivations. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 11(??):383– 398, ???? 1968. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Lemire:1968:NWS

[518] F. W. Lemire. Note on weight spaces of irreducible linear representations. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 11(??):399–404, ???? 1968. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Szekeres:1968:CIR

[519] G. Szekeres. A combinatorial interpretation of Ramanujan’s continued fraction. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 11(??):405–408, ???? 1968. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Berlekamp:1968:CPW

[520] E. R. Berlekamp. A construction for partitions which avoid long arithmetic progressions. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 11(??):409–414, ???? 1968. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Meyers:1968:SPS

[521] Leroy F. Meyers. Some possible and some impossible tripartitions of the plane. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 11(??):415–422, ???? 1968. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Mishra:1968:SML

[522] R. S. Mishra. Sub-manifolds of a locally product Riemannian manifold. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 11(??):423–436, ???? 1968. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Kahan:1968:CEA

[523] W. M. Kahan. Circumscribing an ellipsoid about the intersection of two ellipsoids. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 11(??):437–441, ???? 1968. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Meir:1968:ZDP

[524] A. Meir and A. Sharma. On zeros of derivatives of polynomials. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 11(??): 443–446, ???? 1968. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic). See addendum [550].

Muldoon:1968:ERL

[525] M. E. Muldoon. Extension of a result of L. Lorch and P. Szego on higher monotonicity. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 11(??):447–452, ???? 1968. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Lin:1968:ACM

[526] Shwu yeng T. Lin. Almost continuity of mappings. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 11(??):453–456, ???? 1968. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Al-Salam:1968:SCU

[527] N. A. Al-Salam and W. A. Al-Salam. Some characterizations of the ultraspherical polynomials. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 11(??):457–464, ???? 1968. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Dunkl:1968:LFH

[528] Charles F. Dunkl. Linear functionals on homogeneous polynomials. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 11 (??):465–468, ???? 1968. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Dunham:1968:UMS

[529] Charles B. Dunham. Unisolvence on multidimensional spaces. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 11(??): 469–474, ???? 1968. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Draper:1968:TLT

[530] Norman R. Draper and Irwin Guttman. Transformation of life-test data. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 11(??):475–488, ???? 1968. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Melzak:1968:MPC

[531] Z. A. Melzak. More problems connected with convexity. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 11(??):489– 494, ???? 1968. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Kannappan:1968:FEC

[532] P. L. Kannappan. A functional equation for the cosine. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 11(??):495–498, ???? 1968. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Bondy:1968:NDG

[533] J. A. Bondy. A note on the diameter of a graph. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 11(??):499–502, ???? 1968. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Baron:1968:NE

[534] S. Baron. Note on epi in T_{0}. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 11(??):503–526, ???? 1968. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Guy:1968:NTL

[535] Richard K. Guy and Patrick A. Kelly. The no-three-in-line problem. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 11(??):527–532, ???? 1968. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Bouwer:1968:EVT

[536] I. Z. Bouwer. An edge but not vertex transitive cubic graph. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 11(??): 533–536, ???? 1968. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Mohanty:1968:LPS

[537] S. G. Mohanty and B. R. Handa. On lattice paths with several diagonal steps. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 11(??):537–546, ???? 1968. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Pare:1968:PTT

[538] R. Paré. A Pappus type theorem in the affine group. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 11(??):547–554, ???? 1968. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Maxson:1968:LNR

[539] Carlton J. Maxson. Local near-rings of cardinality p^{2}. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 11(??):555–562, ???? 1968. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Koh:1968:CPR

[540] Kwangil Koh. On compact prime rings and their rings of quotients. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 11(??):563–568, ???? 1968. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Isbell:1968:NEC

[541] John R. Isbell. A note on exact colimits. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 11(??):569–572, ???? 1968. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Rice:1968:GSP

[542] Norman M. Rice. A general selection principle, with applications in analysis and algebra. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 11(??):573–584, ???? 1968. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Ustina:1968:CHM

[543] Fred Ustina. Convergence of the Hausdorff means of double Fourier series. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 11(??):585–592, ???? 1968. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Waltman:1968:NOC

[544] Paul Waltman. A note on an oscillation criterion for an equation with a functional argument. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 11(??):593–596, ???? 1968. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Subbarao:1968:TES

[545] M. V. Subbarao. On a theorem of Erdős and Szekeres. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 11(??):597–598, ???? 1968. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Marsden:1968:BSA

[546] J. Marsden. A Banach space of analytic functions for constant coefficient equations of evolution. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 11(??):599–602, ???? 1968. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Baron:1968:CSS

[547] S. Baron. The coreflective subcategory of sequential spaces. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 11(??): 603–604, ???? 1968. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Boyd:1968:APC

[548] D. W. Boyd and J. S. W. Wong. Another proof of the contraction mapping principle. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 11(??):605–606, ???? 1968. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Robertson:1968:NPT

[549] J. M. Robertson and H. C. Wiser. A note on polar topologies. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 11(??): 607–610, ???? 1968. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Meir:1968:AZD

[550] A. Meir and A. Sharma. Addendum to “On zeros of derivatives of polynomials”. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 11(??):611–654, ???? 1968. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic). See [524].

Alspach:1968:CPC

[551] Brian Alspach. A combinatorial proof of a conjecture of Goldberg and Moon. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 11(??):655–662, ???? 1968. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Hoo:1968:GTH

[552] C. S. Hoo. A generalization of a theorem of Hilton. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 11(??):663–670, ???? 1968. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Oler:1968:EPC

[553] N. Oler. Elements of packing and covering. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 11(??):671–678, ???? 1968. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Lagrange:1968:RGR

[554] R. H. Lagrange and A. H. Rhemtulla. A remark on the group rings of order preserving permutation groups. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 11(??):679–680, ???? 1968. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Kalbfleisch:1968:CSL

[555] J. G. Kalbfleisch and R. G. Stanton. On a certain set of linear inequalities. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 11(??):681–690, ???? 1968. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Natzitz:1968:TPB

[556] Boaz Natzitz. Tensor products of Banach algebras. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 11(??):691–702, ???? 1968. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Anonymous:1968:SEC

[557] Anonymous. Some elementary converse problems in ordinary differential equations. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 11(??):703–716, ???? 1968. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Phadke:1968:TDA

[558] B. B. Phadke. A three dimensional analogue of Pappus’s theorem. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 11 (??):717–718, ???? 1968. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Brossard:1968:NAG

[559] Roland Brossard. A note on absolute geometry. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 11(??):719–722, ???? 1968. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Komhoff:1968:IIC

[560] Magelone Kömhoff. An isoperimetric inequality for convex polyhedra with triangular faces. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 11(??):723–728, ???? 1968. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Bouwer:1968:NMC

[561] I. Z. Bouwer and I. Broere. Note on t-minimal complete bipartite graphs. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 11(??):729–732, ???? 1968. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

James:1968:FSF

[562] R. D. James. The factors of a square-free integer. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 11(??):733–736, ???? 1968. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Anderson:1968:NJB

[563] T. Anderson and A. Heinicke. A note on the Jacobson and Brown–McCoy radicals. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 11(??):737–738, ???? 1968. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Srivastava:1968:MGB

[564] T. N. Srivastava and Y. P. Singh. On Maitland’s generalized Bessel function. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 11(??):739–742, ???? 1968. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Burton:1968:PSP

[565] T. A. Burton. Periodic solutions by Picard’s approximations. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 11(??): 743–746, ???? 1968. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Houston:1968:SRP

[566] R. Houston and S. K. Kaul. Some remarks on a paper of Wong. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 11(??): 747–749, ???? 1968. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Bishop:1969:GLA

[567] E. R. Bishop. Generalized Lipschitz algebras. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 12(??):1–20, ???? 1969. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic). See corrections [922].

Shores:1969:NPN

[568] T. S. Shores. A note on products of normal subgroups. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 12(??):21–24, ???? 1969. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

May:1969:FAS

[569] L. E. May. On C^{∞} functions analytic on sets of small measure. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 12(??): 25–30, ???? 1969. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Jucovic:1969:PRC

[570] E. Jucovic. On polyhedral realizability of certain sequences. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 12(??): 31–40, ???? 1969. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Leech:1969:PMG

[571] John Leech. A presentation of the Mathieu group M_{12}. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 12(??):41–44, ???? 1969. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Mohanty:1969:EVT

[572] S. G. Mohanty and B. R. Handa. Extensions of Vandermonde type convolutions with several summations and their applications. I. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 12(??): 45–62, ???? 1969. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Handa:1969:EVT

[573] B. R. Handa. Extensions of Vandermonde type convolutions with several summations and their applications. II. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 12(??):63–74, ???? 1969. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Hoo:1969:SC

[574] C. S. Hoo. Spaces A×B of conilpotency ≤ 1. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 12(??):75–78, ???? 1969. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Stevens:1969:FLE

[575] R. R. Stevens. On the forced Liénard equation. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 12(??):79–85, ???? 1969. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Chan:1969:AWT

[576] L. K. Chan and E. R. Mead. Application of a Weierstrass theorem to the convergence of moments. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 12(??):86–128, ???? 1969. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Bean:1969:SMP

[577] D. W. T. Bean. Solution to a matroid problem posed by D. J. A. Welsh. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 12(??):129–132, ???? 1969. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Brown:1969:A

[578] B. Brown. An approximation to X, , Y. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 12(??):133–138, ???? 1969. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Brown:1969:RBF

[579] B. Brown. Relations between finite homology and homotopy. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 12(??): 139–150, ???? 1969. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Leech:1969:SSD

[580] John Leech. Six and seven dimensional non-lattice sphere packings. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 12 (??):151–156, ???? 1969. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Behzad:1969:TIP

[581] M. Behzad and S. E. Mahmoodian. On topological invariants of the product of graphs. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 12(??):157–166, ???? 1969. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Day:1969:CMC

[582] Alan Day. A characterization of modularity for congruence lattices of algebras. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 12(??):167–174, ???? 1969. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Rygg:1969:NER

[583] P. Rygg and B. Lehman. A note on an equivalence relation on a purely inseparable field extension. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 12(??):175–178, ???? 1969. CODEN CM-BUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Hoechsmann:1969:SS

[584] Klaus Hoechsmann. Subaffine schemes. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 12(??):179–182, ???? 1969. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Thomas:1969:EMR

[585] R. S. D. Thomas. The equality of a manifold’s rank and dimension. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 12(??):183–184, ???? 1969. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Mathsen:1969:PF

[586] Ronald M. Mathsen. λ(n)-parameter families. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 12(??):185–192, ???? 1969. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Varma:1969:ATJ

[587] A. K. Varma. An application of a theorem of J. Czipszer and G. Freud to a problem of simultaneous approximation. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 12(??):193–202, ???? 1969. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Marsden:1969:HSS

[588] J. E. Marsden. Hamiltonian systems with spin. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 12(??):203–208, ???? 1969. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Marsden:1969:NSG

[589] J. E. Marsden. Non-smooth geodesic flows and classical mechanics. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 12 (??):209–212, ???? 1969. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Djokovic:1969:CS

[590] D. Z. Djoković. On a class of semigroups. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 12(??):213–216, ???? 1969. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Lam:1969:CFP

[591] Tsit yuen Lam. A commutator formula of a pair of subgroups and a theorem of Blackburn. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 12(??):217–220, ???? 1969. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Williams:1969:PIF

[592] Kenneth S. Williams. Polynomials with irreducible factors of specified degree. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 12(??):221–224, ???? 1969. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Megibben:1969:TC

[593] Charles K. Megibben. On a theorem of Cutler. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 12(??):225–228, ???? 1969. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Adams:1969:PDS

[594] Robert Adams. −D is positive definite on a “spiny urchin”. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 12(??): 229–232, ???? 1969. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Marica:1969:CC

[595] John Marica. On a conjecture of Conway. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 12(??):233–234, ???? 1969. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Marica:1969:IDL

[596] J. Marica and J. Schönheim. Incomplete diagonals of Latin squares. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 12(??):235–256, ???? 1969. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Seshadri:1969:CNW

[597] V. Seshadri. A characterization of the normal and Weibull distributions. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 12(??):257–260, ???? 1969. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Reid:1969:SAC

[598] K. B. Reid. On sets of arcs containing no cycles in a tournament. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 12(??): 261–264, ???? 1969. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Clay:1969:BNR

[599] James R. Clay and Donald A. Lawver. Boolean near-rings. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 12(??): 265–274, ???? 1969. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Swanson:1969:NCE

[600] C. A. Swanson. Nonoscillation criteria for elliptic equations. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 12(??): 275–280, ???? 1969. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Haruki:1969:TFE

[601] Hiroshi Haruki. On two functional equations for the trigonometric functions. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 12(??):281–286, ???? 1969. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Storrer:1969:NQF

[602] H. H. Storrer. A note on quasi-Frobenius rings and ring epimorphisms. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 12(??):287–292, ???? 1969. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Garbe:1969:GRM

[603] Dietmar Garbe. A generalization of the regular maps of type 4, , 4_{b,,c} and 3, , 6_{b,,c}. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 12(??):293–298, ???? 1969. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Boyd:1969:MSO

[604] David W. Boyd. Monotone semigroups of operators on cones. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 12(??): 299–310, ???? 1969. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Williams:1969:SSC

[605] Kenneth S. Williams. Small solutions of the congruence ax^{2} + by^{2} ≡ c(modk). Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 12(??):311–320, ???? 1969. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Chvatal:1969:FPP

[606] V. Chvátal. On finite polarized partition relations. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 12(??):321–326, ???? 1969. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Park:1969:PCS

[607] Young Lim Park. On the projective cover of the Stone–Čech compactification of a completely regular Hausdorff space. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 12(??):327–333, ???? 1969. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Malik:1969:DST

[608] Henrick John Malik. Distribution of the sum of truncated binomial variates. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 12(??):334–336, ???? 1969. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Herrlich:1969:SAD

[609] Horst Herrlich. Separation axioms and direct limits. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 12(??):337–338, ???? 1969. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Nobusawa:1969:RAF

[610] Nobuo Nobusawa. On rationality of algebraic function fields. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 12(??): 339–342, ???? 1969. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Pittas:1969:NE

[611] P. A. Pittas. A note on embedding. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 12(??):343–346, ???? 1969. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Tan:1969:NTD

[612] Peter Tan. A note on a theorem of Dynkin on necessary and sufficient statistics. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 12(??):347–352, ???? 1969. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Schwerdtfeger:1969:CNN

[613] Hans Schwerdtfeger. Correction to: “Notes on numerical analysis. II”. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 12(??):353–384, ???? 1969. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic). See [55].

Bruckner:1969:SRE

[614] A. M. Bruckner. Some remarks on extreme derivates. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 12(??):385–388, ???? 1969. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Menard:1969:DCA

[615] Jean Menard. Sur deux classes d’anneaux non-commutatifs. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 12(??): 389–400, ???? 1969. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Eisenberg:1969:MSB

[616] Sheldon M. Eisenberg. Moment sequences and the Bernstein polynomials. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 12(??):401–412, ???? 1969. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Erdos:1969:CPI

[617] Paul Erdős. On a combinatorial problem. III. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 12(??):413–416, ???? 1969. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Courter:1969:RAW

[618] R. C. Courter. Rings all of whose factor rings are semi-prime. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 12(??): 417–426, ???? 1969. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Godfrey:1969:PRM

[619] M. C. Godfrey and M. Sion. On product of Radón measures. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 12(??): 427–444, ???? 1969. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Borwein:1969:SLM

[620] D. Borwein and R. Phillips. Scales of logarithmic methods of summability. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 12(??):445–452, ???? 1969. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Roggenkamp:1969:RSO

[621] Klaus W. Roggenkamp. A remark on separable orders. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 12(??):453–456, ???? 1969. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Srinivasacharyulu:1969:TSK

[622] K. Srinivasacharyulu. Topology of some Kähler manifolds. II. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 12(??): 457–460, ???? 1969. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Kaul:1969:CSF

[623] S. K. Kaul. Compact subsets in function spaces. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 12(??):461–466, ???? 1969. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Rhemtulla:1969:PGN

[624] A. H. Rhemtulla. A property of groups with no central factors. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 12(??): 467–470, ???? 1969. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Berlekamp:1969:SIE

[625] E. R. Berlekamp. On subsets with intersections of even cardinality. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 12 (??):471–474, ???? 1969. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Tarwater:1969:SCT

[626] J. D. Tarwater and R. C. Entringer. Sums of complexes in torisionfree abelian groups. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 12(??):475–478, ???? 1969. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Heilbronn:1969:SCT

[627] H. Heilbronn and P. Scherk. Sums of complexes in torsion free abelian groups. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 12(??):479–480, ???? 1969. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Belluce:1969:SFP

[628] L. P. Belluce and W. A. Kirk. Some fixed point theorems in metric and Banach spaces. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 12(??):481–492, ???? 1969. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Mullin:1969:ETR

[629] R. C. Mullin and E. Nemeth. An existence theorem for room squares. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 12(??):493–498, ???? 1969. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Hoo:1969:MSC

[630] C. S. Hoo. Multiplication on spaces with comultiplication. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 12(??): 499–507, ???? 1969. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Meyer:1969:RST

[631] C. D. Meyer. On the rank of the sum of two rectangular matrices. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 12(??): 508, ???? 1969. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Eames:1969:SUD

[632] W. Eames. On strong upper densities. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 12(??):509–510, ???? 1969. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Manely:1969:FSG

[633] P. L. Manely. The Frattini subgroup of a p-group. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 12(??):511–512, ???? 1969. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Singh:1969:NSC

[634] S. P. Singh and W. Russell. A note on a sequence of contraction mappings. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 12(??):513–516, ???? 1969. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Dubuc:1969:TBF

[635] Serge Dubuc. Le théorème de Bernstein sur les fonctions complètement monotones. (French) [The Bernstein theorem on complete monotonic functions]. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 12(??):517–520, ???? 1969. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Dubuc:1969:DPI

[636] Serge Dubuc. La dérivée de la primitive de l’intégrale de Perron. (French) [The derivative of the primitive of Perron’s integral]. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 12(??):521–522, ???? 1969. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Chui:1969:CTC

[637] Charles K. Chui. A convergence theorem for certain Riemann sums. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 12(??):523–544, ???? 1969. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Williams:1969:TCC

[638] Kenneth S. Williams. On two conjectures of Chowla. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 12(??):545–566, ???? 1969. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Kabe:1969:CML

[639] D. G. Kabe. On characterizing the multivariate linear exponential distribution. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 12(??):567–572, ???? 1969. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Sahney:1969:NCB

[640] B. N. Sahney and P. D. Kathal. A new criterion for Borel summability of Fourier series. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 12(??):573–580, ???? 1969. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Thompson:1969:M

[641] R. C. Thompson. On the matrices A and f(A). Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 12(??):581–588, ???? 1969. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Williams:1969:NFP

[642] Kenneth S. Williams. Note on factorable polynomials. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 12(??):589–596, ???? 1969. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Zelmanowitz:1969:SPG

[643] Julius Zelmanowitz. A shorter proof of Goldie’s theorem. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 12(??):597– 602, ???? 1969. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Erbe:1969:DCT

[644] Lynn Erbe. Disconjugacy conditions for the third order linear differential equation. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 12(??):603–614, ???? 1969. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Menon:1969:IGS

[645] K. V. Menon. Inequalities for generalized symmetric functions. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 12(??): 615–624, ???? 1969. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Kim:1969:ABP

[646] Choo whan Kim. An analogue of Birkoff’s problem 111 for infinite Markov matrices. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 12(??):625–634, ???? 1969. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Marica:1969:DSP

[647] J. Marica and J. Schönheim. Differences of sets and a problem of Graham. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 12(??):635–638, ???? 1969. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Rosenthal:1969:RIS

[648] Peter Rosenthal. Remarks on invariant subspace lattices. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 12(??):639– 644, ???? 1969. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Burgess:1969:SPG

[649] W. D. Burgess. On semi-perfect group rings. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 12(??):645–652, ???? 1969. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

MacHenry:1969:IQG

[650] Trueman MacHenry. On the invariance of a quotient group of the center of F/[R, , R]. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 12(??):653–660, ???? 1969. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Vamos:1969:NQI

[651] P. Vámos. A note on the quotients of indecomposable injective modules. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 12(??):661–667, ???? 1969. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Toralballa:1969:CFW

[652] L. V. Toralballa. On continuous functions which are partially differentiable almost everywhere. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 12(??):668–672, ???? 1969. CODEN CM-BUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Spoar:1969:CPJ

[653] G. Spoar and N. D. Lane. A conformal proof of a Jordan curve problem. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 12(??):673–674, ???? 1969. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Leduc:1969:CLA

[654] P.-Y. Leduc. Catégories localement artiniennes sans radical. (French) [Radical-free locally Artinian categories]. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 12(??):675–676, ???? 1969. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Ligh:1969:GTZ

[655] Steve Ligh. A generalization of a theorem of Zassenhaus. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 12(??):677– 678, ???? 1969. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Chowla:1969:BPR

[656] S. Chowla and H. London. Bounds on the n−th power residues (mod, p). Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 12(??):679–680, ???? 1969. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Tang:1969:CNS

[657] C. Y. Tang. Correction to: “Notes on splitting extensions of groups”. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 12(??):681–704, ???? 1969. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic). See [514].

Duggal:1969:SRS

[658] K. L. Duggal. Singular Riemannian structures compatible with πstructures. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 12(??):705–720, ???? 1969. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Freedman:1969:IFT

[659] H. I. Freedman. The implicit function theorem in the scalar case. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 12(??): 721–732, ???? 1969. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Koh:1969:GHT

[660] E. L. Koh. On the generalized Hankel and K transformations. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 12(??): 733–740, ???? 1969. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Gratzer:1969:JDP

[661] G. Grätzer, H. Lakser, and J. Plonka. Joins and direct products of equational classes. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 12(??):741–744, ???? 1969. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Folkman:1969:PIC

[662] J. H. Folkman and R. L. Graham. A packing inequality for compact convex subsets of the plane. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 12(??):745–752, ???? 1969. CODEN CM-BUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Hadwiger:1969:CFS

[663] H. Hadwiger and P. Mani. Continuous functions on the sphere and isometries. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 12(??):753–758, ???? 1969. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Houh:1969:RMA

[664] C. S. Houh. On a Riemannian manifold M_{2n} with an almost tangent structure. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 12(??):759–770, ???? 1969. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Fletcher:1969:FTS

[665] P. Fletcher. Finite topological spaces and quasi-uniform structures. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 12 (??):771–776, ???? 1969. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Sauer:1969:MSG

[666] N. Sauer and J. Schönheim. Maximal subsets of a given set having no triple in common with a Steiner triple system on the set. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 12(??): 777–778, ???? 1969. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Swick:1969:CST

[667] K. E. Swick. Convergence of solutions of third order differential equations. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 12(??):779–792, ???? 1969. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Park:1969:QLE

[668] Young L. Park. On quotient-like extensions of C(X). Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 12(??):793–800, ???? 1969. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Bruen:1969:SWD

[669] A. Bruen and J. C. Fisher. Spreads which are not dual spreads. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 12(??): 801–804, ???? 1969. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Thompson:1969:PPD

[670] R. C. Thompson. On Pearl’s paper “A decomposition theorem for matrices”. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 12(??):805–808, ???? 1969. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Storrer:1969:CPD

[671] H. H. Storrer. A characterization of Prüfer domains. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 12(??):809–812, ???? 1969. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Holmes:1969:FPP

[672] R. D. Holmes. On fixed and periodic points under certain sets of mappings. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 12(??):813–822, ???? 1969. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Abbott:1969:CPE

[673] H. L. Abbott and D. Hanson. On a combinatorial problem of Erdős. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 12(??):823–830, ???? 1969. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Gupta:1969:SAF

[674] Haragauri N. Gupta. On some axioms in foundations of Cartesian spaces. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 12(??):831–836, ???? 1969. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Baker:1969:AWE

[675] John A. Baker. An analogue of the wave equation and certain related functional equations. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 12(??):837–846, ???? 1969. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Goodman:1969:LPD

[676] E. Goodman and T. V. Narayana. Lattice paths with diagonal steps. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 12(??):847–857, ???? 1969. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Koh:1969:AIH

[677] Kwangil Koh. On the annihilators of the injective hull of a module. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 12(??):858–860, ???? 1969. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Higgs:1969:EBM

[678] Denis Higgs. Equicardinality of bases in B-matroids. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 12(??):861–862, ???? 1969. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Natzitz:1969:SSO

[679] B. Natzitz. On spectral synthesis for one point. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 12(??):863–864, ???? 1969. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Rao:1969:NGP

[680] B. Raja Rao and M. L. Garg. A note on the generalized (positive) Cauchy distribution. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 12(??):865–868, ???? 1969. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Reddy:1969:MMM

[681] A. R. Reddy. On the maximum modulus and the mean modulus of an entire function. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 12(??):869–872, ???? 1969. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Clements:1970:ETE

[682] John C. Clements. Enclosure theorems for eigenvalues of elliptic operators. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 13(??):1–8, ???? 1970. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

To:1970:HBEa

[683] Ting-On To. On the Hahn–Banach extension property. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 13(??):9–14, ???? 1970. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Jonsson:1970:RMM

[684] W. Jónsson. On a result of A. M. Macbeath on normal subgroups of a Fuchsian group. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 13(??):15–16, ???? 1970. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Ford:1970:RQO

[685] Gary G. Ford. Remarks on quasigroups obeying a generalized associative law. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 13(??):17–22, ???? 1970. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Zaks:1970:QFR

[686] Abraham Zaks. Quasi-Frobenius X-rings. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 13(??):23–30, ???? 1970. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Butler:1970:NEC

[687] G. J. Butler, L. H. Erbe, and R. M. Mathsen. On the non-existence of conjugate points. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 13(??):31–38, ???? 1970. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Lin:1970:NAP

[688] Chia shiang Lin. A note on the adjoint of the product of operators. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 13(??):39–46, ???? 1970. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Platt:1970:NESa

[689] Craig Platt. A note on endomorphism semigroups. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 13(??):47–48, ???? 1970. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic). See corrections [776, 837].

Bollobas:1970:MNT

[690] Béla Bollobás. The maximal number of triangles of maximal perimeter length determined by a finite set. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 13(??):49–50, ???? 1970. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Goldberg:1970:GQR

[691] Myron Goldberg. The group of the quadratic residue tournament. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 13 (??):51–54, ???? 1970. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Norris:1970:SCG

[692] C. W. Norris. Sequences of contractions in a generalized metric space. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 13(??):55–58, ???? 1970. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Boyd:1970:OPS

[693] David W. Boyd. Osculatory packings by spheres. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 13(??):59–64, ???? 1970. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Lindner:1970:CLR

[694] Charles C. Lindner. On completing Latin rectangles. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 13(??):65– 68, ???? 1970. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Nadler:1970:NIL

[695] S. B. Nadler, Jr. A note on inverse limits of finite spaces. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 13(??): 69–70, ???? 1970. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Fieldhouse:1970:PSI

[696] David J. Fieldhouse. Pure simple and indecomposable rings. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 13(??): 71–78, ???? 1970. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Douglas:1970:NTH

[697] Robert J. Douglas. A note on a theorem of H. L. Abbott. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 13(??): 79–82, ???? 1970. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Menon:1970:SF

[698] K. V. Menon. Symmetric forms. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 13(??):83–88, ???? 1970. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

vanQue:1970:DFK

[699] Ngô van Quê. Démonstration de la formule Künneth par la théorie harmonique. (French) [Demonstration by harmonic theory of the Künneth formula]. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 13(??):89–94, ???? 1970. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Charnow:1970:AAC

[700] A. Charnow. The automorphisms of an algebraically closed field. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 13(??): 95–98, ???? 1970. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Dubinsky:1970:SHE

[701] E. Dubinsky and T. Husain. Solutions of homogeneous elliptic equations. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 13(??):99–104, ???? 1970. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Capell:1970:KT

[702] P. Capell and T. V. Narayana. On knock-out tournaments. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 13(??): 105–110, ???? 1970. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Ng:1970:RCM

[703] Kai wang Ng. A remark on contractive mappings. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 13(??):111–114, ???? 1970. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Bruns:1970:IHS

[704] G. Bruns and H. Lakser. Injective hulls of semilattices. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 13(??):115–118, ???? 1970. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Maxwell:1970:NAD

[705] George Maxwell. A note on Artin’s Diophantine conjecture. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 13(??): 119–120, ???? 1970. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Burry:1970:MDC

[706] J. H. W. Burry and H. W. Ellis. On measures determined by continuous functions that are not of bounded variation. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 13(??):121–124, ???? 1970. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Tam:1970:NES

[707] K. Kuen Tam. A note on the existence of a solution of the Falkner–Skan equation. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 13(??):125–128, ???? 1970. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Hoechsmann:1970:LIN

[708] K. Hoechsmann. Lifting ideals in noncommutative integral extensions. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 13(??):129–130, ???? 1970. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Mullin:1970:NOS

[709] R. C. Mullin and E. Nemeth. On the nonexistence of orthogonal Steiner systems of order 9. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 13(??):131–136, ???? 1970. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Rosenfeld:1970:PRL

[710] M. Rosenfeld. On polynomials with related level sets. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 13(??):137–138, ???? 1970. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Gratzer:1970:FPD

[711] G. Grätzer and B. Wolk. Finite projective distributive lattices. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 13(??): 139–140, ???? 1970. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Wong:1970:RTL

[712] James S. W. Wong. A remark on a theorem of Lyapunov. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 13(??): 141–144, ???? 1970. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Cockayne:1970:ART

[713] E. J. Cockayne. An application of Ramsey’s theorem. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 13(??):145–146, ???? 1970. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Sember:1970:APC

[714] John J. Sember. The associative part of a convergence domain is invariant. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 13(??):147–148, ???? 1970. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Wong:1970:EQI

[715] Edward T. Wong. Endomorphisms of the quasi-injective hull of a module. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 13(??):149–150, ???? 1970. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Ahuja:1970:DSI

[716] J. C. Ahuja. On the distribution of sum of independent positive binomial variables. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 13(??):151–168, ???? 1970. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Schirmer:1970:PFP

[717] Helga Schirmer. Properties of the fixed point set of contractive multifunctions. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 13(??):169–174, ???? 1970. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Sallee:1970:MAR

[718] G. T. Sallee. Maximal areas of Reuleaux polygons. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 13(??):175–180, ???? 1970. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Ditzian:1970:ICC

[719] Z. Ditzian. Inversion of a class of convolution transforms of generalized functions. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 13(??):181–186, ???? 1970. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Rao:1970:AIF

[720] V. Venugopal Rao. Averages involving Fourier coefficients of non-analytic automorphic forms. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 13(??):187–198, ???? 1970. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Costley:1970:SNL

[721] Charles G. Costley. On singular normal linear integral equations. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 13(??): 199–204, ???? 1970. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Tan:1970:SEA

[722] Peter Tan. On Sprott’s example of an ancillary statistic. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 13(??):205– 208, ???? 1970. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Holmes:1970:ANS

[723] R. D. Holmes and P. P. Narayanaswamy. On asymptotically nonexpansive semigroups of mappings. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 13(??):209–214, ???? 1970. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Bishop:1970:TM

[724] E. R. Bishop. On a theorem of Mitchell. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 13(??):215–218, ???? 1970. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Siddiqi:1970:CMI

[725] Jamil A. Siddiqi. On a characterization of maximal ideals. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 13(??): 219–220, ???? 1970. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Heinig:1970:ITZ

[726] H. P. Heinig. On an interpolation theorem of Zygmund and Koizumi. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 13(??):221–226, ???? 1970. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Robillard:1970:LBR

[727] Pierre Robillard. Lower bounds for the Ramsey numbers. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 13(??):227– 230, ???? 1970. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Bouwer:1970:VET

[728] I. Z. Bouwer. Vertex and edge transitive, but not 1-transitive, graphs. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 13(??):231–238, ???? 1970. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Rodabaugh:1970:PAA

[729] D. J. Rodabaugh. Power-associative algebras in which every subalgebra is a left ideal. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 13(??):239–244, ???? 1970. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Simis:1970:NM

[730] Aron Simis. On ℵ_{α}-noetherian modules. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 13(??):245–248, ???? 1970. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Zaidman:1970:RIA

[731] S. Zaidman. A remark on integration of almost-periodic functions. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 13 (??):249–252, ???? 1970. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Day:1970:SSW

[732] Alan Day. A simple solution to the word problem for lattices. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 13(??): 253–254, ???? 1970. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

MacLeod:1970:EPC

[733] R. A. MacLeod and I. Barrodale. On equal products of consecutive integers. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 13(??):255–260, ???? 1970. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Demers:1970:CCD

[734] Luc Demers. A counterexample to a conjecture of D. B. Fuks. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 13(??): 261–272, ???? 1970. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Shuster:1970:BCP

[735] Jonathan Shuster. On the Borel–Cantelli problem. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 13(??):273–276, ???? 1970. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Putz:1970:OAL

[736] Randee Putz. The order of algebraic linear transformations. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 13(??): 277–278, ???? 1970. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Gordon:1970:URB

[737] William R. Gordon. A unitary relation between a matrix and its transpose. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 13(??):279–280, ???? 1970. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Giroux:1970:CRM

[738] André Giroux. Sur une conjecture de R. M. Robinson. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 13(??):281–298, ???? 1970. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Al-Ani:1970:NDS

[739] Sabri Al-Ani. On the noncentral distributions of the second largest roots of three matrixes in multivariate analysis. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 13(??):299–304, ???? 1970. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Gandhi:1970:C

[740] J. M. Gandhi. The coefficients of (sinhx)/cosx. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 13(??):305–310, ???? 1970. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Kabe:1970:MIU

[741] D. G. Kabe. On a multiple integral useful in order statistics distribution theory. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 13(??):311–316, ???? 1970. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Alspach:1970:PST

[742] Brian Alspach. On point-symmetric tournaments. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 13(??):317–324, ???? 1970. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Sherman:1970:MAF

[743] Malcolm J. Sherman. On the moduli of analytic functions. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 13(??): 325–328, ???? 1970. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Graham:1970:SDG

[744] R. L. Graham. On subtrees of directed graphs with no path of length exceeding one. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 13(??):329–332, ???? 1970. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Long:1970:PPD

[745] Calvin T. Long. On a problem in partial difference equations. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 13(??): 333–336, ???? 1970. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic). See corrections [836].

Lambek:1970:S

[746] Joachim Lambek. Subequalizers. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 13(??):337–350, ???? 1970. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Costley:1970:CIO

[747] Charles G. Costley. On Carleman integral operators. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 13(??):351–358, ???? 1970. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Heinicke:1970:RAN

[748] A. G. Heinicke. A remark about noncommutative integral extensions. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 13(??):359–362, ???? 1970. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Berlekamp:1970:STV

[749] E. R. Berlekamp. On sets of ternary vectors whose only linear dependencies involve an odd number of vectors. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 13(??):363–366, ???? 1970. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Murdeshwar:1970:BQU

[750] M. G. Murdeshwar and K. K. Theckedath. Boundedness in a quasiuniform space. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 13(??):367–370, ???? 1970. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Haff:1970:NUG

[751] C. E. Haff, U. S. R. Murty, and R. C. Wilton. A note on undirected graphs realizable as p.o. sets. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 13(??):371–374, ???? 1970. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Hell:1970:GS

[752] Pavol Hell and Jaroslav Nesetril. Graphs and k-societies. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 13(??):375– 383, ???? 1970. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Murphy:1970:RST

[753] Ian S. Murphy. The rank of the sum of two rectangular matrices. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 13(??): 384, ???? 1970. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Thierrin:1970:IDM

[754] G. Thierrin. Ideaux à droite minimaux d’un anneau primitif. (French) [Right-minimum ideals of a primitive ring]. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 13(??):385–386, ???? 1970. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Hunt:1970:TIC

[755] J. H. V. Hunt. A theorem on involutions on cyclic Peano spaces. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 13(??): 387–388, ???? 1970. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Matthews:1970:DIT

[756] J. A. J. Matthews and G. de B. Robinson. On the dimension of an irreducible tensor representation of the general linear group GL(d). Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 13 (??):389–390, ???? 1970. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Kale:1970:IML

[757] B. K. Kale. Inadmissibility of the maximum likelihood estimator in the presence of prior information. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 13(??):391–414, ???? 1970. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Costley:1970:SHE

[758] Charles G. Costley. On a solution of the Hammerstein equation with singular normal kernels. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 13(??):415–422, ???? 1970. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Datta:1970:CWC

[759] D. K. Datta. On Y. C. Wong’s conjecture. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 13(??):423–424, ???? 1970. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Davison:1970:RF

[760] T. M. K. Davison. On rings of fractions. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 13(??):425–430, ???? 1970. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Dyer:1970:LCI

[761] James A. Dyer. Left Cauchy integral bases in linear topological spaces. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 13(??):431–440, ???? 1970. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Fieldhouse:1970:NSS

[762] D. Fieldhouse. Note on the singular submodule. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 13(??):441–442, ???? 1970. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Gupta:1970:FGV

[763] Chander Kanta Gupta. On free groups of the variety AN_{2}∧N_{2}A. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 13(??): 443–446, ???? 1970. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Komhoff:1970:DIP

[764] Magelone Kömhoff. On a 3-dimensional isoperimetric problem. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 13(??): 447–450, ???? 1970. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Kotzig:1970:MVF

[765] Anton Kotzig and Alexander Rosa. Magic valuations of finite graphs. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 13(??):451–462, ???? 1970. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Lemire:1970:ODR

[766] F. W. Lemire. One-dimensional representations of the cycle subalgebra of a semi-simple Lie algebra. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 13(??):463–468, ???? 1970. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Lin:1970:CCB

[767] C.-S. Lin. On certain classes of bounded linear operators. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 13(??):469– 474, ???? 1970. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic). See corrections [835].

Losey:1970:OSI

[768] Gerald Losey. Are one-sided inverses two-sided inverses in a matrix ring over a group ring? Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 13(??):475–480, ???? 1970. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Luedeman:1970:NI

[769] John K. Luedeman. A note on the ∑(S)-injectivity of R(S). Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 13(??): 481–490, ???? 1970. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

McCarthy:1970:AFD

[770] P. J. McCarthy. Arithmetical functions and distributivity. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 13(??):491– 496, ???? 1970. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Milnes:1970:CCG

[771] Paul Milnes. Counterexample to a conjecture of Greenleaf. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 13(??): 497–500, ???? 1970. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Osofsky:1970:RKS

[772] B. L. Osofsky. A remark on the Krull–Schmidt–Azumaya theorem. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 13 (??):501–506, ???? 1970. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Peterson:1970:OMP

[773] A. C. Peterson. On the ordering of multi-point boundary value functions. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 13(??):507–514, ???? 1970. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Sedlacek:1970:MGG

[774] J. Sedlácek. On the minimal graph with a given number of spanning trees. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 13(??):515–518, ???? 1970. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Williams:1970:IBF

[775] Kenneth S. Williams. Integers of biquadratic fields. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 13(??):519–525, ???? 1970. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Platt:1970:NESb

[776] Craig Platt. A note of Endomorphism Semigroups (correction). Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 13(??): 526, ???? 1970. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic). See [689].

To:1970:HBEb

[777] Ting On To. On the Hahn–Banach extension property. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 13(??):526, ???? 1970. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Burns:1970:CIG

[778] R. G. Burns. Central idempotents in group rings. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 13(??):527–528, ???? 1970. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Chihara:1970:CCD

[779] T. S. Chihara. A characterization and a class of distribution functions for the Stieltjes–Wigert polynomials. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 13(??):529–532, ???? 1970. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Consul:1970:IPC

[780] P. C. Consul and M. J. Kemp. On an interesting property of a combinatorial function. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 13(??):533–536, ???? 1970. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Letac:1970:EWH

[781] Gérard Letac. Ensembles de Wiener–Hopf. II. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 13(??):537–538, ???? 1970. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Lindner:1970:CNJ

[782] Charles C. Lindner. Comment on a note by J. Marica and J. Schönheim. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 13(??):539–540, ???? 1970. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Rahman:1970:BMZ

[783] Q. I. Rahman. A bound for the moduli of the zeros of polynomials. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 13(??):541–542, ???? 1970. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Anonymous:1971:NP

[784] Anonymous. A note on permanents. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 14(??):1–4, ???? 1971. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Adams:1971:PEC

[785] R. A. Adams. Properties of equivalent capacities. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 14(??):5–12, ???? 1971. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Alas:1971:CCN

[786] O. T. Alas. On a characterization of collectionwise normality. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 14(??): 13–16, ???? 1971. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Dubinsky:1971:SSH

[787] E. Dubinsky and T. Husain. Space of solutions of homogeneous elliptic equations. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 14(??):17–24, ???? 1971. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Faierman:1971:PTP

[788] M. Faierman. On a perturbation in a two-parameter ordinary differential equation of the second order. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 14(??):25–34, ???? 1971. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Giles:1971:MRP

[789] R. Giles and H. Kummer. A matrix representation of a pair of projections in a Hilbert space. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 14(??):35–44, ???? 1971. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Graham:1971:CST

[790] R. L. Graham and J. H. Spencer. A constructive solution to a tournament problem. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 14(??):45–48, ???? 1971. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Hoo:1971:WWC

[791] C. S. Hoo. The weak weak category of a space. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 14(??):49–52, ???? 1971. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Law:1971:ORP

[792] A. G. Law. Orthogonal recurrence polynomials and Hamburger moments. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 14(??):53–56, ???? 1971. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Lindner:1971:CQS

[793] Charles C. Lindner. Construction of quasigroups satisfying the identity x(xy) = yx. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 14(??):57–60, ???? 1971. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Lindner:1971:GSD

[794] Charles C. Lindner. The generalized singular direct product for quasigroups. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 14(??):61–64, ???? 1971. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Mohanty:1971:NCI

[795] S. G. Mohanty. A note on combinatorial identities for partial sums. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 14(??):65–68, ???? 1971. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Muldoon:1971:ERM

[796] M. E. Muldoon. Elementary remarks on multiply monotonic functions and sequences. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 14(??):69–72, ???? 1971. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Rhemtulla:1971:MSF

[797] A. H. Rhemtulla and Anne Penfold Street. Maximal sum-free sets in elementary abelian p-groups. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 14(??):73–80, ???? 1971. CODEN CM-BUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Speed:1971:NSL

[798] T. P. Speed. A note on Stone lattices. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 14(??):81–86, ???? 1971. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Thomson:1971:HSV

[799] B. S. Thomson. On the Henstock strong variational integral. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 14(??): 87–100, ???? 1971. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Anonymous:1971:GFP

[800] Anonymous. On the generating function for permutations with repetitions and inversions. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 14(??):101–102, ???? 1971. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Ballard:1971:CDF

[801] David Ballard. Coloring a dodecahedron with four colors. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 14(??):103– 106, ???? 1971. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Connett:1971:CUB

[802] J. Connett. On covering the unit ball in normed spaces. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 14(??):107–110, ???? 1971. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Finkelstein:1971:DJL

[803] R. Finkelstein and H. London. On D. J. Lewis’s equation x^{3} +117y^{3} = 5. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 14(??):111–112, ???? 1971. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Lemire:1971:EWS

[804] F. W. Lemire. Existence of weight space decompositions for irreducible representations of simple Lie algebras. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 14(??):113–116, ???? 1971. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Lindner:1971:QOG

[805] Charles C. Lindner. Quasigroups orthogonal to a given abelian group. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 14(??):117–118, ???? 1971. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Lohman:1971:ETS

[806] Robert H. Lohman. An embedding theorem for separable locally convex spaces. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 14(??):119–120, ???? 1971. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Reich:1971:SRC

[807] Simeon Reich. Some remarks concerning contraction mappings. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 14(??): 121–124, ???? 1971. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Stavrakas:1971:PK

[808] N. M. Stavrakas and R. E. Jamison. On a problem of Klee. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 14(??): 125–126, ???? 1971. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Willard:1971:PSP

[809] S. Willard. Paracompactness in small products. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 14(??):127–146, ???? 1971. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic). See note [1098].

Chrysovergis:1971:SRT

[810] A. Chrysovergis. Some remarks on Talenti’s semigroup. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 14(??):147–150, ???? 1971. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Chvatal:1971:SCT

[811] V. Chvátal and J. Komlós. Some combinatorial theorems on monotonicity. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 14(??):151–158, ???? 1971. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Gonshor:1971:RRQ

[812] Harry Gonshor. Remarks on rings of quotients of rings of functions. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 14(??):159–160, ???? 1971. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Haruki:1971:RBS

[813] Hiroshi Haruki. On a relation between the “square” functional equation and the “square” mean-value property. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 14(??):161–166, ???? 1971. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Kabe:1971:SDO

[814] D. G. Kabe. Some distributions of ordered values from Burr and beta distributions. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 14(??):167–174, ???? 1971. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Leviatan:1971:CSL

[815] Dany Leviatan and Lee Lorch. On the connectedness of the sets of limit points of certain transforms of bounded sequences. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 14(??):175–182, ???? 1971. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Nadler:1971:ACC

[816] Sam B. Nadler, Jr. Arc components of certain chainable continua. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 14 (??):183–190, ???? 1971. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Pearson:1971:SSC

[817] T. L. Pearson. Some sufficient conditions for maximal-resolvability. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 14 (??):191–196, ???? 1971. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Sachdeva:1971:FIM

[818] Usha Sachdeva. On finite invariant measure for semigroups of operators. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 14(??):197–206, ???? 1971. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Sankaran:1971:RFP

[819] N. Sankaran. Rings of formal power series. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 14(??):207–220, ???? 1971. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Sekanina:1971:AOG

[820] Milan Sekanina. On an algorithm for ordering of graphs. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 14(??):221– 224, ???? 1971. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Shum:1971:IIR

[821] D. T. Shum. On integral inequalities related to Hardy’s. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 14(??):225– 230, ???? 1971. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Singh:1971:SCV

[822] K. Singh. A special case of the vanishing of a (G, , σ)-product in a (G, , σ)- space. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 14(??):231–234, ???? 1971. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Subramanian:1971:IVC

[823] H. Subramanian. Integer-valued continuous functions. II. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 14(??):235– 238, ???? 1971. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Swiatak:1971:LFE

[824] Halina Światak. On linear functional equations with nonpolynomial C^{∞} solutions. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 14(??):239–244, ???? 1971. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Davis:1971:THT

[825] Chandler Davis. The Toeplitz–Hausdorff theorem explained. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 14(??): 245–246, ???? 1971. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

deOliveira:1971:MPP

[826] Graciano N. de Oliveira. Matrices with prescribed principal elements and singular values. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 14(??):247–250, ???? 1971. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Dion:1971:EMP

[827] J. P. Dion. Estimateurs de la moyenne pour les distributions symétriques à domaine borné. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 14(??):251–254, ???? 1971. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Drake:1971:ANS

[828] David A. Drake. Another note on Sperner’s lemma. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 14(??):255–256, ???? 1971. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Gibson:1971:RSC

[829] P. M. Gibson. Rank of the sum of certain matrices. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 14(??):257–258, ???? 1971. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Koh:1971:POS

[830] Kwangil Koh. On prime one-sided ideals. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 14(??):259–260, ???? 1971. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Kubota:1971:ACP

[831] Y. Kubota. An approximately continuous Perron integral. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 14(??):261– 264, ???? 1971. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Milnes:1971:CFN

[832] Paul Milnes. Convolution of L^{p} functions on non-unimodular groups. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 14(??):265–266, ???? 1971. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Rahman:1971:DZP

[833] Q. I. Rahman. The distinct zeros of the product of a polynomial and its successive derivatives. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 14(??):267–270, ???? 1971. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Satyanarayana:1971:NSI

[834] M. Satyanarayana. A note on a self injective ring. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 14(??):271–272, ???? 1971. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Lin:1971:CCC

[835] Chia Shiang Lin. Corrections: “On certain classes of bounded linear operators”. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 14(??):273, ???? 1971. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic). See [767].

Long:1971:CPP

[836] Calvin T. Long. Corrections: “On a problem in partial difference equations”. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 14(??):273, ???? 1971. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic). See [745].

Platt:1971:CNE

[837] Craig Platt. Corrections: “A note on endomorphism semigroups”. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 14 (??):273–300, ???? 1971. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic). See [689, 776].

Robinson:1971:LI

[838] G. de B. Robinson. Leopold Infeld. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 14(??):301–304, ???? 1971. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Adams:1971:CIT

[839] R. A. Adams and John Fournier. A compact imbedding theorem for functions without compact support. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 14(??):305–310, ???? 1971. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Cohen:1971:AF

[840] Maurice Cohen. Approximation of foliations. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 14(??):311–314, ???? 1971. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Djokovic:1971:SGC

[841] D. Z. Djoković. Sets of generators of a commutative and associative algebra. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 14(??):315–320, ???? 1971. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Fantham:1971:CGS

[842] Peter Fantham. On canonical generators of subgroups. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 14(??):321–324, ???? 1971. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Handa:1971:GOD

[843] B. R. Handa. On a generalization of one dimensional random walk with a partially reflecting barrier. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 14(??):325–332, ???? 1971. CODEN CM-BUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Heinicke:1971:TL

[844] A. G. Heinicke. Triples and localizations. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 14(??):333–340, ???? 1971. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Jain:1971:SRC

[845] G. C. Jain. Some results in a correlated random walk. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 14(??):341–348, ???? 1971. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Koh:1971:FRR

[846] Kwangil Koh. On functional representations of a ring without nilpotent elements. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 14(??):349–352, ???? 1971. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Koneru:1971:CIF

[847] Rao S. Koneru and B. S. Lalli. On convergence of iterations for fixed points of repulsive type. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 14(??):353–358, ???? 1971. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Lambek:1971:RMS

[848] J. Lambek. On the representation of modules by sheaves of factor modules. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 14(??):359–368, ???? 1971. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Lindgren:1971:TSU

[849] W. F. Lindgren. Topological spaces with a unique compatible quasiuniformity. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 14(??):369–372, ???? 1971. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Nieto:1971:TIJ

[850] J. I. Nieto. Sur un théorème d’interpolation de J. Lions et J. Peetre. (French) [On an interpolation theorem of J. Lions and J. Peetre]. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 14(??): 373–376, ???? 1971. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

OShaughnessy:1971:BSQ

[851] C. D. O’Shaughnessy. On the block structure of quartic designs. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 14(??): 377–390, ???? 1971. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Prabhakar:1971:SIE

[852] Tilak Raj Prabhakar. Some integral equations with Kummer’s functions in the kernels. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 14(??):391–404, ???? 1971. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Roggenkamp:1971:OSA

[853] Klaus W. Roggenkamp. R-orders in a split algebra have finitely many non-isomorphic irreducible lattices as soon as R has finite class number. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 14(??):405–410, ???? 1971. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Szeto:1971:CPM

[854] George Szeto. On a class of projective modules over central separable algebras. I, II. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 14(??):415–418, ???? 1971. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Willett:1971:GVP

[855] D. Willett. Generalized de la Vallée Poussin disconjugacy tests for linear differential equations. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 14(??):419–428, ???? 1971. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Williams:1971:NNS

[856] Kenneth S. Williams. Note on the number of solutions of f(x_{1}) = f(x_{2}) = ⋯ = f(x_{r}) over a finite field. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 14(??):429–432, ???? 1971. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Chaika:1971:ACA

[857] M. Chaika and S. Perlman. On approximately continuous and almost continuous functions. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 14(??):433–436, ???? 1971. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Kalbfleisch:1971:RBR

[858] J. G. Kalbfleisch. On Robillard’s bounds for Ramsey numbers. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 14(??): 437–440, ???? 1971. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Kleiman:1971:TRS

[859] Howard Kleiman. Totally real subfields of p-adic fields having the symmetric group as Galois group. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 14(??):441–442, ???? 1971. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Koh:1971:PRA

[860] Kwangil Koh and A. C. Mewborn. On prime rings with ascending chain condition on annihilator right ideals and nonzero injective right ideals. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 14(??):443–444, ???? 1971. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Malik:1971:SRA

[861] M. A. Malik. Some remarks on abstract differential operators. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 14(??): 445–448, ???? 1971. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Mendelsohn:1971:MSM

[862] N. S. Mendelsohn. On maximal sets of mutually orthogonal idempotent Latin squares. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 14(??):449–450, ???? 1971. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Menon:1971:NNA

[863] M. V. Menon and V. Seshadri. A note on normal attraction to a stable law. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 14(??):451–452, ???? 1971. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Modak:1971:WPG

[864] S. V. Modak. Wreath product of O^{∗}-groups that is not in O^{∗}. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 14(??): 453–454, ???? 1971. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Pareek:1971:SM

[865] C. M. Pareek. On spaces and ω-mappings. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 14(??):455–458, ???? 1971. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Pareek:1971:CS

[866] C. M. Pareek. Characterizations of p-spaces. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 14(??):459–460, ???? 1971. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Spohn:1971:CCI

[867] William G. Spohn, Jr. On Conway’s conjecture for integer sets. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 14(??): 461–488, ???? 1971. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Robinson:1971:SB

[868] G. de B. Robinson. Samuel Beatty. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 14(??):489–490, ???? 1971. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Armendariz:1971:DM

[869] Efraim P. Armendariz. Q-divisible modules. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 14(??):491–494, ???? 1971. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Burgess:1971:SIR

[870] W. Burgess and M. Chacron. Subdirectly irreducible DQC rings. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 14(??): 495–498, ???? 1971. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Eames:1971:TGM

[871] W. Eames. On a topology generated by measurable covers. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 14(??): 499–502, ???? 1971. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Elazar:1971:PRW

[872] R. Elazar and M. Gutterman. On a problem of random walk in space. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 14(??):503–506, ???? 1971. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Hartfiel:1971:PCC

[873] D. J. Hartfiel and J. W. Crosby. On the permanent of a certain class of (0, , 1)-matrices. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 14(??):507–512, ???? 1971. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Lindstrom:1971:MFP

[874] Bernt Lindström. On Möbius functions and a problem in combinatorial number theory. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 14(??):513–516, ???? 1971. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Luedeman:1971:GCR

[875] John K. Luedeman. A generalization of the concept of a ring of quotients. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 14(??):517–530, ???? 1971. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Nelson:1971:LEC

[876] Evelyn Nelson. The lattice of equational classes of semigroups with zero. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 14(??):531–534, ???? 1971. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Riddell:1971:LPP

[877] J. Riddell. A lattice point problem related to sets containing no l-term arithmetic progression. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 14(??):535–538, ???? 1971. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Willett:1971:OFI

[878] D. Willett. Oscillation on finite or infinite intervals of second order linear differential equations. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 14(??):539–550, ???? 1971. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Zaidman:1971:ERS

[879] S. Zaidman. An existence result for Stepanoff almost-periodic differential equations. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 14(??):551–554, ???? 1971. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Chaput:1971:SMG

[880] Luc Chaput. A simpler method for getting some exact distributions. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 14(??):555–558, ???? 1971. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Deruaz:1971:NAD

[881] Marcel Déruaz and Mary O’Keefe. A note about the definition of CWcomplexes. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 14(??):559–560, ???? 1971. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Djokovic:1971:FSI

[882] D. Z. Djoković. Finite sets of integers with equal power sums. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 14(??): 561–564, ???? 1971. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Griffith:1971:APT

[883] Gareth J. Griffith. An analytic proof of a theorem of Felix Klein. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 14(??): 565–568, ???? 1971. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Hall:1971:CEC

[884] Japheth Hall, Jr. A condition for equality of cardinals of minimal generators under closure operators. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 14(??):569–570, ???? 1971. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Lohman:1971:IS

[885] Robert H. Lohman. Isomorphisms in subspaces of c_{0}. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 14(??):571–572, ???? 1971. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Marica:1971:OFS

[886] John Marica. Orthogonal families of sets. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 14(??):573–574, ???? 1971. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Ponzo:1971:PS

[887] Peter J. Ponzo. Periodic solutions for ẋ = Ax + g(x, , t) + εp(t). Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 14(??): 575–578, ???? 1971. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Shultz:1971:MOF

[888] Harris S. Shultz. Multiplication of operators by C^{∞} functions. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 14(??): 579–582, ???? 1971. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Stitzinger:1971:TDW

[889] Ernest L. Stitzinger. On a theorem of D. W. Barnes. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 14(??):583–584, ???? 1971. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Thompson:1971:STV

[890] D. E. Thompson. Some theorems on a Volterra equation of the second kind. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 14(??):585–590, ???? 1971. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Todd:1971:CP

[891] Christopher Todd. On the compactification of products. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 14(??):591– 592, ???? 1971. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Singh:1971:AES

[892] R. Singh. Admissible estimators of θ^{r} in some extreme value densities. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 14(3):411–414, ???? 1971. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Wyman:1972:LM

[893] M. Wyman. Leo Moser (1921–1970). Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 15(??):1–8, ???? 1972. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Abbott:1972:NRT

[894] H. L. Abbott. A note on Ramsey’s theorem. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 15(??):9–10, ???? 1972. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Andrews:1972:FCI

[895] G. E. Andrews, M. V. Subbarao, and M. Vidyasagar. A family of combinatorial identities. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 15(??):11–18, ???? 1972. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Chvatal:1972:RPM

[896] V. Chvátal. Remarks on a problem of Moser. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 15(??):19–22, ???? 1972. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Erdos:1972:PG

[897] Paul Erdős. On a problem of Grünbaum. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 15(??):23–26, ???? 1972. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Erdos:1972:SAG

[898] Paul Erdős, A. Meir, V. T. Sós, and P. Turán. On some applications of graph theory. III. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 15(??):27–32, ???? 1972. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Hanson:1972:PP

[899] Denis Hanson. On the product of the primes. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 15(??):33–38, ???? 1972. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Moon:1972:AAG

[900] J. W. Moon. Almost all graphs have a spanning cycle. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 15(??):39–42, ???? 1972. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Spencer:1972:RCI

[901] Joel Spencer. A remark on coloring integers. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 15(??):43–44, ???? 1972. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Basavappa:1972:SME

[902] P. Basavappa. On the solutions of the matrix equation f(X, X^{∗}) = g(X, X^{∗}). Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 15(??):45–50, ???? 1972. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Binding:1972:BCV

[903] P. A. Binding, W. D. Hoskins, and P. J. Ponzo. Bounds for characteristic values of positive definite matrices. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 15(??):51–56, ???? 1972. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Bondy:1972:VHT

[904] J. A. Bondy. Variations on the Hamiltonian theme. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 15(??):57–62, ???? 1972. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Cass:1972:STS

[905] Frank P. Cass. Some theorems on strong Nörlund summability. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 15(??): 63–70, ???? 1972. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Closs:1972:RSS

[906] M. P. Closs. Riemannian structures subordinate to certain almost tangent structures. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 15(??):71–78, ???? 1972. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Connell:1972:RS

[907] Ian G. Connell. The ring scheme F_{q}. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 15(??):79–86, ???? 1972. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Henle:1972:LDT

[908] Michael Henle. A Lebesgue decomposition theorem for C^{∗}-algebras. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 15 (??):87–92, ???? 1972. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Klemola:1972:TF

[909] Tapio Klemola. On transportable forms. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 15(??):93–102, ???? 1972. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Kubota:1972:CDA

[910] Y. Kubota. A constructive definition of the approximately continuous Denjoy integral. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 15(??):103–108, ???? 1972. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Richardson:1972:PCP

[911] G. D. Richardson. R_{1} pairwise compact, and pairwise complete spaces. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 15(??):109–114, ???? 1972. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Singhal:1972:CEM

[912] J. P. Singhal. Certain extensions of the Mehler formula. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 15(??):115–118, ???? 1972. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Taylor:1972:GEB

[913] M. A. Taylor. The generalized equations of bisymmetry associativity and transitivity on quasigroups. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 15(??):119–124, ???? 1972. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Fantham:1972:SRW

[914] Peter Fantham. Some remarks on weights of permutations. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 15(??): 125–128, ???? 1972. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Costley:1972:SFE

[915] Charles G. Costley. On some functional equations of Carleman. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 15(??): 129–132, ???? 1972. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Menon:1972:IES

[916] K. V. Menon. An inequality for elementary symmetric functions. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 15(??): 133–136, ???? 1972. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Nicholson:1972:LGR

[917] W. K. Nicholson. Local group rings. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 15(??):137–138, ???? 1972. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Rahman:1972:ZLC

[918] Q. I. Rahman. Zeros of linear combinations of polynomials. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 15(??): 139–142, ???? 1972. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Roberts:1972:HDJ

[919] Leslie G. Roberts. Higher derivations and the Jordan canonical form of the companion matrix. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 15(??):143–144, ???? 1972. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Wulfsohn:1972:CDF

[920] Aubrey Wulfsohn. A compactification due to Fell. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 15(??):145–146, ???? 1972. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Garbe:1972:CSR

[921] D. Garbe and J. Mennicke. Corrections: “Some remarks on the Mathieu groups”. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 15(??):147, ???? 1972. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic). See [192].

Bishop:1972:CGL

[922] E. R. Bishop. Corrections: “Generalized Lipschitz algebras”. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 15(??): 148–170, ???? 1972. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic). See [567].

Basu:1972:ITF

[923] A. K. Basu. Invariance theorems for first passage time random variables. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 15(??):171–176, ???? 1972. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Bernstein:1972:ESP

[924] Leon Bernstein. Explicit solutions of pyramidal Diophantine equations. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 15(??):177–184, ???? 1972. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Chacron:1972:RSR

[925] M. Chacron and G. Thierrin. σ-reflexive semigroups and rings. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 15(??): 185–188, ???? 1972. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

deOliveira:1972:MIE

[926] G. N. de Oliveira. On the multiplicative inverse eigenvalue problem. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 15(??):189–194, ???? 1972. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Guy:1972:SNG

[927] Richard K. Guy. The slimming number and genus of graphs. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 15(??): 195–200, ???? 1972. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Harary:1972:GSE

[928] F. Harary, J. Krarup, and A. Schwenk. Graphs suppressible to an edge. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 15(??):201–204, ???? 1972. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Helfenstein:1972:FNC

[929] H. G. Helfenstein. A flat nonmetrizable connexion. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 15(??):205–206, ???? 1972. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Herzog:1972:GPF

[930] M. Herzog and J. Schönheim. Group partition, factorization and the vector covering problem. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 15(??):207–214, ???? 1972. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Istratescu:1972:RCU

[931] I. Istratescu. Remarks concerning uniformly bounded operators on Hilbert space. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 15(??):215–218, ???? 1972. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Jackson:1972:MTS

[932] H. L. Jackson. On minimally thin sets in a Stolz domain. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 15(??): 219–224, ???? 1972. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Kabe:1972:DMB

[933] D. G. Kabe and R. P. Gupta. Decomposition of the multivariate beta distribution with applications. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 15(??):225–228, ???? 1972. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Krantzberg:1972:ETI

[934] Julius A. Krantzberg. Existence theorem for the initial-boundary value problem for a singular parabolic partial differential equation. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 15(??): 229–234, ???? 1972. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Magarian:1972:NTP

[935] E. A. Magarian and J. L. Mott. Noetherian tensor products. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 15(??): 235–238, ???? 1972. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Maxwell:1972:DDS

[936] George Maxwell. Diagonals of doubly stochastic matrices. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 15(??):239– 242, ???? 1972. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Morales:1972:MVT

[937] Pedro Morales. Mean value theorem for the m-integral of dinculeanu. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 15(??):243–252, ???? 1972. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Peterson:1972:MNB

[938] A. C. Peterson. On the monotone nature of boundary value functions for nth-order differential equations. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 15(??):253–258, ???? 1972. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Price:1972:KPD

[939] Martin E. Price. Kantorovitch polynomials diminish generalized length. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 15(??):259–262, ???? 1972. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Rosenstein:1972:WBR

[940] Joseph G. Rosenstein. On GL_{2}(R) where R is a Boolean ring. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 15(??): 263–276, ???? 1972. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Shen:1972:ERE

[941] C. Y. Shen. An elementary result on exponential measure spaces. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 15(??): 277–278, ???? 1972. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Siddiqi:1972:GTW

[942] Jamil A. Siddiqi. On a generalization of a theorem of Wiener. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 15(??): 279–284, ???? 1972. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Basu:1972:BBF

[943] A. K. Basu and M. T. Wasan. The bounds based on the functions of observations for maximum of stable law. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 15(??):285–288, ???? 1972. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Bruckner:1972:NMD

[944] A. M. Bruckner. A note on measures determined by continuous functions. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 15(??):289–292, ???? 1972. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Cook:1972:CT

[945] W. D. Cook and R. J. Webster. Carathéodory’s theorem. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 15(??):293– 294, ???? 1972. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Dash:1972:PDS

[946] A. T. Dash. Positive definite sequence of operators and a fixed point theorem. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 15(??):295–296, ???? 1972. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Herrlich:1972:IC

[947] H. Herrlich and C. M. Ringel. Identities in categories. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 15(??):297–300, ???? 1972. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

McMorris:1972:SCM

[948] F. R. McMorris. The singular congruence and the maximal quotient semigroup. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 15(??):301–304, ???? 1972. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Moore:1972:CRP

[949] Marc Moore. Consistency in the reconstruction of patterns from sample data. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 15(??):305–308, ???? 1972. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Ruzsa:1972:PPE

[950] I. Ruzsa, Jr. On a problem of P. Erdős. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 15(??):309–322, ???? 1972. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Al-Ani:1972:LRC

[951] Sabri Al-Ani. On the ith latent root of a complex matrix. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 15(??): 323–328, ???? 1972. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Altman:1972:STC

[952] E. Altman. Some theorems on convex polygons. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 15(??):329–340, ???? 1972. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Boyd:1972:DPW

[953] David W. Boyd. Disk packings which have non-extreme exponents. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 15 (??):341–344, ???? 1972. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Chambers:1972:PS

[954] Graham A. Chambers. On f-prefrattini subgroups. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 15(??):345–348, ???? 1972. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Cockayne:1972:CCG

[955] E. J. Cockayne. Colour classes for r-graphs. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 15(??):349–354, ???? 1972. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Elliott:1972:AWS

[956] George A. Elliott. Another weak Stone–Weierstrass theorem for C^{∗}- algebras. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 15(??):355–358, ???? 1972. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Faierman:1972:SPE

[957] M. Faierman. Some properties of equations in integers. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 15(??):359–362, ???? 1972. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Fraser:1972:ELS

[958] Robert B. Fraser, Jr. Equivalence of Lipschitz structures. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 15(??):363– 368, ???? 1972. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Holub:1972:NUB

[959] J. R. Holub and J. R. Retherford. A note on unconditional bases. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 15(??): 369–372, ???? 1972. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Howlett:1972:FFA

[960] C. Howlett and D. Schumacher. Free finitary algebras in a co-complete Cartesian closed category. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 15(??):373–374, ???? 1972. CODEN CM-BUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Leroux:1972:CCG

[961] Pierre Leroux. Une caractérisation de la catégorie des groupes. (French) [A characterization of the category of groups]. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 15(??):375–380, ???? 1972. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Nadler:1972:NIT

[962] Sam B. Nadler, Jr. A note on an iterative test of Edelstein. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 15(??): 381–386, ???? 1972. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Rhee:1972:CMM

[963] C. J. Rhee. On a class of multivalued mappings in Banach spaces. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 15 (??):387–394, ???? 1972. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Saleh:1972:DSV

[964] A. K. Md. Ehsanes Saleh and M. A. Rahim. Distribution of the sum of variates from truncated discrete populations. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 15(??):395–398, ???? 1972. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Stewart:1972:FFN

[965] James Stewart. Functions with a finite number of negative squares. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 15 (??):399–410, ???? 1972. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Szeto:1972:CPM

[966] George Szeto. On a class of projective modules over central separable algebras II. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 15(??):411–416, ???? 1972. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Thompson:1972:DCT

[967] D. V. Thompson. Discontinuity conditions on transformation groups. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 15(??):417–420, ???? 1972. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Tomiuk:1972:MAA

[968] B. J. Tomiuk. Modular annihilator A^{∗}-algebras. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 15(??):421–426, ???? 1972. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Wakae:1972:HSR

[969] M. Wakae and O. Hamara. On the homology of the n-special reduced product space of a Euclidean space. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 15(??):427–432, ???? 1972. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Zaki:1972:AAI

[970] M. Zaki. Almost automorphic integrals of almost automorphic functions. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 15(??):433–436, ???? 1972. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Bouwer:1972:NPG

[971] I. Z. Bouwer and G. F. LeBlanc. A note on primitive graphs. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 15(??): 437–440, ???? 1972. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Burns:1972:NGR

[972] R. G. Burns and V. W. D. Hale. A note on group rings of certain torsion-free groups. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 15(??):441–446, ???? 1972. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Hajnal:1972:CTD

[973] A. Hajnal, E. C. Milner, and E. Szemerédi. A cure for the telephone disease. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 15(??):447–450, ???? 1972. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Tzimbalario:1972:QPD

[974] Jean Tzimbalario. On a question posed by D. Leviatan and L. Lorch. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 15(??):453–468, ???? 1972. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Ballinger:1972:COS

[975] B. T. Ballinger. Completions of ordered sets. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 15(??):469–472, ???? 1972. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Beesack:1972:LEC

[976] Paul R. Beesack. The Laurent expansion without Cauchy’s integral theorem. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 15(??):473–480, ???? 1972. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Beesack:1972:SST

[977] Paul R. Beesack. On Sturm’s separation theorem. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 15(??):481–488, ???? 1972. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Ellis:1972:IFD

[978] H. W. Ellis and J. D. Hiscocks. Integral functionals in the duals of L^{λ}- spaces. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 15(??):489–500, ???? 1972. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Erdos:1972:TPC

[979] P. Erdős and E. C. Milner. A theorem in the partition calculus. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 15(??): 501–506, ???? 1972. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic). See corrigendum [1161].

Finbow:1972:FPS

[980] Arthur S. Finbow. The fixed point set of real multi-valued contraction mappings. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 15(??):507–512, ???? 1972. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Goldberg:1972:NVI

[981] Samuel I. Goldberg. Normal variations of invariant hypersurfaces of framed manifolds. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 15(??):513–522, ???? 1972. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Gupta:1972:TEG

[982] Narain Gupta. Third-Engel 2-groups are soluble. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 15(??):523–524, ???? 1972. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Hartmann:1972:STS

[983] F. W. Hartmann. Summability tests for singular points. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 15(??):525– 528, ???? 1972. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Hughes:1972:GUI

[984] I. Hughes and K. R. Pearson. The group of units of the integral group ring ZS_{3}. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 15(??):529–534, ???? 1972. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Klee:1972:QCC

[985] Victor Klee. On a question of Colin Clark concerning three properties of convex sets. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 15(??):535–538, ???? 1972. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Lorimer:1972:FTD

[986] Peter Lorimer. On the finite two-dimensional linear groups. II. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 15(??): 539–546, ???? 1972. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Mehra:1972:UEV

[987] K. L. Mehra and P. V. Ramachandramurty. On unbiased estimation of a vector parameter. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 15(??):547–550, ???? 1972. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Pethe:1972:AFB

[988] S. P. Pethe and G. C. Jain. Approximation of functions by a Bernsteintype operator. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 15(??):551–558, ???? 1972. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Pixley:1972:LMC

[989] Alden F. Pixley. Local Malcev conditions. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 15(??):559–568, ???? 1972. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Tang:1972:FSG

[990] C. Y. Tang. On the Frattini subgroups of generalized free products with cyclic amalgamations. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 15(??):569–574, ???? 1972. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Votaw:1972:UCQ

[991] Charles I. Votaw. Uniqueness of compatible quasi-uniformities. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 15(??): 575–584, ???? 1972. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Williams:1972:P

[992] H. C. Williams. The primality of N = 2A3^{n} − 1. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 15(??):585–590, ???? 1972. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Woods:1972:LC

[993] R. Grant Woods. On the local connectedness of βX − X. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 15(??): 591–594, ???? 1972. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Bruen:1972:PFF

[994] Aiden Bruen. Permutation functions on a finite field. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 15(??):595–598, ???? 1972. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Byrd:1972:WQI

[995] K. A. Byrd. When are quasi-injectives injective? Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 15(??):599–600, ???? 1972. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Leblanc:1972:RPP

[996] Raymond Leblanc. Représentation des polynômes positifs sur [−1, 1]. (French) [Representation of positive polynomials on [−1, 1]]. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 15(??): 601–602, ???? 1972. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Lee:1972:FPT

[997] Cheng Ming Lee. On fixed point theorems for mappings in a separated locally convex space. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 15(??):603–604, ???? 1972. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Martindale:1972:NGP

[998] Wallace S. Martindale III. A note on generalized polynomial identities. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 15(??):605–606, ???? 1972. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Nel:1972:NS

[999] L. D. Nel and D. Riordan. Note on a subalgebra of C(X). Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 15(??): 607–608, ???? 1972. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Rogers:1972:NC

[1000] Thomas Rogers. On Nagumo’s condition. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 15(??):609–611, ???? 1972. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Marsaglia:1972:WD

[1001] G. Marsaglia and G. P. H. Styan. When does rank(A + B) = rank(A) + rank(B)? Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 15(3):451–452, ???? 1972. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Ahsanullah:1973:SIM

[1002] M. Ahsanullah and M. Rahman. A singular inverse of a matrix by rank annihilation. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 16(??):1–4, ???? 1973. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Allegretto:1973:OCM

[1003] W. Allegretto and L. Erbe. Oscillation criteria for matrix differential inequalities. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 16(??):5–10, ???? 1973. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Alspach:1973:DMO

[1004] B. Alspach and J. L. Berggren. On the determination of the maximum order of the group of a tournament. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 16(??):11–14, ???? 1973. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Assad:1973:FPT

[1005] Nadim A. Assad. A fixed point theorem for weakly uniformly strict contractions. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 16(??):15–18, ???? 1973. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Berman:1973:LIC

[1006] Abraham Berman. Linear inequalities over complex cones. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 16(??):19–22, ???? 1973. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Berthiaume:1973:RMQ

[1007] Pierre Berthiaume. Une remarque sur les monoïdes de quotients généralisés. (French) [Remark on the monoids of generalized quotients]. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 16(??):23–26, ???? 1973. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Biggs:1973:CSR

[1008] R. G. Biggs. On the central series of a ring. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 16(??):27–32, ???? 1973. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Chvatal:1973:FFH

[1009] V. Chvátal. Flip-flops in hypo-Hamiltonian graphs. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 16(??):33–42, ???? 1973. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Cohen:1973:SOD

[1010] Maurice Cohen. Smoothing one-dimensional foliations on S^{1} × S^{1}. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 16 (??):43–44, ???? 1973. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Cornish:1973:SIS

[1011] William H. Cornish. Subdirectly irreducible semirings and semigroups without nonzero nilpotents. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 16(??):45–48, ???? 1973. CODEN CM-BUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Erbe:1973:OCS

[1012] Lynn Erbe. Oscillation criteria for second order nonlinear delay equations. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 16(??):49–56, ???? 1973. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Friedler:1973:OCF

[1013] Louis Friedler. Open, connected functions. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 16(??):57–60, ???? 1973. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic). See errata [1160].

Kelly:1973:DML

[1014] B. G. A. Kelly. Densities and measures of linear sets. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 16(??):61– 66, ???? 1973. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Komlos:1973:CLT

[1015] J. Komlós. A central limit theorem for multiplicative systems. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 16(??): 67–74, ???? 1973. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Mount:1973:SSO

[1016] K. S. Mount and B. K. Kale. On selecting a spurious observation. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 16(??): 75–78, ???? 1973. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Muldowney:1973:IP

[1017] James S. Muldowney. On an inequality of Peano. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 16(??):79–82, ???? 1973. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Mullen:1973:BIB

[1018] Kenneth Mullen. Bernstein’s inequality in the bivariate case. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 16(??): 83–86, ???? 1973. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Parrott:1973:SGJ

[1019] David Parrott. On the simple group of J. Tits. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 16(??):87–92, ???? 1973. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Siddiqi:1973:DFC

[1020] Rafat N. Siddiqi. On density of Fourier coefficients. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 16(??):93–104, ???? 1973. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Swanson:1973:SOE

[1021] C. A. Swanson. Strong oscillation of elliptic equations in general domains. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 16(??):105–110, ???? 1973. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Tupper:1973:DFA

[1022] B. O. J. Tupper. A divergence-free antisymmetric tensor. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 16(??):111– 114, ???? 1973. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Wilson:1973:PSC

[1023] Richard G. Wilson. Pointwise sequentially closed ideals in C^{∗}(X). Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 16 (??):115–118, ???? 1973. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Yocom:1973:TMF

[1024] K. L. Yocom. Totally multiplicative functions in regular convolution rings. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 16(??):119–128, ???? 1973. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Fisher:1973:HFE

[1025] J. C. Fisher. Hypersurfaces fixed by equiaffinities. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 16(??):129–132, ???? 1973. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Marsden:1973:PCE

[1026] J. Marsden. A proof of the Calderon extension theorem. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 16(??):133– 136, ???? 1973. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Zischka:1973:ASN

[1027] K. A. Zischka. On approximations to solutions of nonlinear integral equations of the Urysohn type. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 16(??):137–158, ???? 1973. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Adler:1973:DPA

[1028] Andrew Adler. Determinateness and the Pasch axiom. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 16(??):159–160, ???? 1973. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Al-Hussaini:1973:CCE

[1029] A. N. Al-Hussaini. On characterizations of conditional expectation. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 16 (??):161–164, ???? 1973. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Aupetit:1973:TMP

[1030] Bernard H. Aupetit. Théorème de Müntz pour les fonctions de plusieurs variables. (French) [Müntz’s theorem for functions of several variables]. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 16(??):165–166, ???? 1973. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Bacsich:1973:ERU

[1031] Paul D. Bacsich. An epi-reflector for universal theories. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 16(??):167–172, ???? 1973. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Beuerman:1973:SFS

[1032] D. R. Beuerman. Some functional stable limit theorems. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 16(??):173– 178, ???? 1973. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Bozel:1973:ILC

[1033] F. Bozel and T. Husain. On isomorphisms of locally convex spaces with similar biorthogonal systems. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 16(??):179–184, ???? 1973. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Bryant:1973:ECM

[1034] Jack Bryant and L. F. Guseman, Jr. Extensions of contractive mappings and Edelstein’s iterative test. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 16(??):185–192, ???? 1973. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Constantin:1973:NPM

[1035] Julien Constantin. Note sur un problème de Maeda. (French) [Note on a problem of Maeda]. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 16(??):193–194, ???? 1973. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Hanson:1973:TSS

[1036] D. Hanson. On a theorem of Sylvester and Schur. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 16(??):195–200, ???? 1973. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Hardy:1973:GFP

[1037] G. E. Hardy and T. D. Rogers. A generalization of a fixed point theorem of Reich. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 16(??):201–206, ???? 1973. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Herzberg:1973:KPS

[1038] Agnes M. Herzberg, C. W. L. Garner, and B. G. F. Springer. Kissprecise sequential rotatable designs. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 16(??):207–218, ???? 1973. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Jeffries:1973:AEC

[1039] Clark D. Jeffries. Analytic evaluation of certain characteristic classes. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 16(??):219–224, ???? 1973. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic). See correction [1102].

Kaul:1973:IST

[1040] S. K. Kaul. On the irregular sets of a transformation group. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 16(??): 225–232, ???? 1973. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Makan:1973:FLF

[1041] A. R. Makan. The fitting length of a finite soluble group and the number of conjugacy classes of its maximal metanilpotent subgroups. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 16(??): 233–238, ???? 1973. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Malik:1973:ESA

[1042] M. A. Malik. Existence of solutions of abstract differential equations in a local space. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 16(??):239–244, ???? 1973. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

May:1973:SF

[1043] L. E. May. Separation of functions. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 16(??):245–250, ???? 1973. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Meisen:1973:PRC

[1044] Jeanne Meisen. Pullbacks in regular categories. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 16(??):251–256, ???? 1973. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Melzak:1973:MST

[1045] Z. A. Melzak. Multiple scalar transport. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 16(??):257–268, ???? 1973. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Millington:1973:NSG

[1046] H. Millington. Nuclear spaces of generalized test functions. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 16(??): 269–274, ???? 1973. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Peterson:1973:RBT

[1047] A. C. Peterson. On a relation between a theorem of Hartman and a theorem of Sherman. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 16(??):275–282, ???? 1973. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Plummer:1973:CCP

[1048] M. D. Plummer and E. L. Wilson. On cycles and connectivity in planar graphs. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 16(??):283–288, ???? 1973. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Ramachandramurty:1973:SCQ

[1049] P. V. Ramachandramurty and M. Sudhakara Rao. Some comments on quantiles and order statistics. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 16(??):289–294, ???? 1973. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Srivastava:1973:RAF

[1050] H. M. Srivastava. On the reducibility of Appell’s function F_{4}. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 16(??): 295–298, ???? 1973. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Halter-Koch:1973:LE

[1051] Franz Halter-Koch. Letter to the editor. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 16(??):299–316, ???? 1973. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Ramamurthi:1973:WRR

[1052] V. S. Ramamurthi. Weakly regular rings. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 16(??):317–322, ???? 1973. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Rishel:1973:ICS

[1053] Thomas W. Rishel. Images of class-C spaces. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 16(??):323–328, ???? 1973. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Scott:1973:CB

[1054] Kenneth F. N. Scott. On the construction of BIBD with λ = 1. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 16(??):

329–336, ???? 1973. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Srivastava:1973:CNT

[1055] M. S. Srivastava. On a class of nonparametric tests for independence — bivariate case. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 16(??):337–342, ???? 1973. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Stein:1973:C

[1056] Sherman K. Stein. The cardinalities of A + A and A − A. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 16(??): 343–346, ???? 1973. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Taylor:1973:CNG

[1057] M. A. Taylor. Cartesian nets and groupoids. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 16(??):347–362, ???? 1973. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Williams:1973:KSR

[1058] Kenneth S. Williams. The Kloosterman sum revisited. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 16(??):363–366, ???? 1973. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Wetzel:1973:SCC

[1059] John E. Wetzel. Sectorial covers for curves of constant length. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 16(??): 367–376, ???? 1973. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Andersen:1973:HCF

[1060] Kenneth F. Andersen. Hermite conjugate functions and rearrangement invariant spaces. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 16(??):377–380, ???? 1973. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Berndt:1973:NMA

[1061] Bruce C. Berndt. A new method in arithmetical functions and contour integration. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 16(??):381–388, ???? 1973. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Jain:1973:ENV

[1062] G. C. Jain. On the expected number of visits of a particle before absorption in a correlated random walk. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 16(??):389–396, ???? 1973. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic). See [1914].

Small:1973:CQG

[1063] Charles Small. On certain quotients of Grothendieck groups. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 16(??): 397–400, ???? 1973. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Shreve:1973:CPM

[1064] Warren E. Shreve. A converse problem in matrix differential equations. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 16(??):401–404, ???? 1973. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Cohen:1973:PM

[1065] Gerard Elie Cohen. Profinite modules. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 16(??):405–416, ???? 1973. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Dobbs:1973:ACA

[1066] David E. Dobbs. Amitsur cohomology in additive functors. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 16(??): 417–426, ???? 1973. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Gantner:1973:PSM

[1067] T. E. Gantner. On m-paracompact spaces and ω-mappings. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 16(??): 427–428, ???? 1973. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Armendariz:1973:NSR

[1068] Efraim P. Armendariz. A note on semiprime rings with torsionless injective envelope. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 16(??):429–432, ???? 1973. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Coppotelli:1973:CS

[1069] Fred Coppotelli. A construction of subequalizers. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 16(??):433–434, ???? 1973. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Eberhart:1973:SWC

[1070] C. Eberhart, J. B. Fugate, and L. Mohler. Spaces which cannot be written as a countable disjoint union of closed subsets. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 16(??):435–438, ???? 1973. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic). See errata [1207].

Enochs:1973:NSR

[1071] E. Enochs. A note on semihereditary rings. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 16(??):439–440, ???? 1973. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Harley:1973:CNF

[1072] P. W. Harley III. A countable nowhere first countable Hausdorff space. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 16(??):441–442, ???? 1973. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Haruki:1973:NST

[1073] Shigeru Haruki. A note on a square type functional equation. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 16(??): 443–446, ???? 1973. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Houh:1973:SDT

[1074] C. S. Houh. On a structure defined by a tensor field F of type (1, , 1) satisfying F^{2} = 0. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 16(??):447–450, ???? 1973. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Lorimer:1973:NGR

[1075] Peter Lorimer. A note on groups of Ree type. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 16(??):451–452, ???? 1973. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Menon:1973:MUI

[1076] M. V. Menon and V. Seshadri. On the martingales upcrossings inequality. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 16(??):453–454, ???? 1973. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Michael:1973:FDE

[1077] I. M. Michael. Finite-dimensional extensions of certain symmetric operators. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 16(??):455–456, ???? 1973. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Mordell:1973:SC

[1078] L. J. Mordell. Systems of congruences. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 16(??):457–474, ???? 1973. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Beauregard:1973:OBD

[1079] Raymond A. Beauregard. Overrings of Bezout domains. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 16(??):475– 478, ???? 1973. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Blum:1973:NCE

[1080] R. Blum and A. P. Guinand. Nets of conics in the Euclidean plane and an associated representational geometry. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 16(??):479–496, ???? 1973. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Burgess:1973:RQG

[1081] W. D. Burgess. A ring of quotients for group rings which is easy to describe. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 16(??):497–500, ???? 1973. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Carson:1973:PSI

[1082] Andrew B. Carson. Partially self-injective regular rings. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 16(??):501– 506, ???? 1973. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Chandra:1973:SMP

[1083] Ashok K. Chandra. On the solution of Moser’s problem in four dimensions. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 16(??):507–512, ???? 1973. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Djokovic:1973:EMW

[1084] D. Z. Djoković. Epimorphisms of modules which must be isomorphisms. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 16(??):513–516, ???? 1973. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Dunwoody:1973:PTG

[1085] M. J. Dunwoody and A. Pietrowski. Presentations of the trefoil group. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 16(??):517–520, ???? 1973. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Gilmer:1973:FPM

[1086] Robert Gilmer. On factorization of polynomials modulo n. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 16(??): 521–524, ???? 1973. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Greenwell:1973:ULC

[1087] D. L. Greenwell and H. V. Kronk. Uniquely line colorable graphs. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 16 (??):525–530, ???? 1973. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

MacLeod:1973:ECP

[1088] R. A. MacLeod and F. D. K. Roberts. Equalizing the coefficients in a product of polynomials. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 16(??):531–540, ???? 1973. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Maruster:1973:CZO

[1089] S. Maruster. Sur le calcul des zéros d’un opérateur discontinu par itération. (French) [On the calculation of zeros of a discontinuous operator by iteration]. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 16(??):541–544, ???? 1973. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Morales:1973:PCS

[1090] Pedro Morales. Pointwise compact spaces. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 16(??):545–550, ???? 1973. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Parmenter:1973:UCR

[1091] M. Parmenter and S. Sehgal. Uniqueness of the coefficient ring in some group rings. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 16(??):551–556, ???? 1973. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Swetits:1973:CPL

[1092] J. Swetits and B. Wood. On a class of positive linear operators. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 16(??): 557–560, ???? 1973. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Thompson:1973:ESV

[1093] R. C. Thompson and S. Therianos. The eigenvalues and singular values of matrix sums and products. VII. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 16(??):561–570, ???? 1973. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Tomkins:1973:EMC

[1094] R. J. Tomkins. On the equivalence of modes of convergence. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 16(??): 571–576, ???? 1973. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Traynor:1973:ACG

[1095] Tim Traynor. Absolute continuity for group-valued measures. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 16(??): 577–580, ???? 1973. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Weston:1973:BET

[1096] J. H. Weston. m-bounded extensions of topological spaces. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 16(??): 581–586, ???? 1973. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Yang:1973:PRS

[1097] J. S. Yang. Property (G), regularity, and semi-equicontinuity. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 16(??): 587–594, ???? 1973. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Chew:1973:NPS

[1098] James Chew. Note on: “Paracompactness in small products” (Canad. Math. Bull. 14 (1971), 127) by S. Willard. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 16(??):595–596, ???? 1973. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic). See [809].

Darst:1973:MID

[1099] R. B. Darst. Most infinitely differentiable functions are nowhere analytic. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 16(??):597–598, ???? 1973. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Goel:1973:ANS

[1100] D. S. Goel and B. N. Sahney. On the absolute Nörlund summability of a Fourier series. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 16(??):599–602, ???? 1973. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Gratzer:1973:NIC

[1101] G. Grätzer and H. Lakser. A note on the implicational class generated by a class of structures. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 16(??):603–622, ???? 1973. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Jeffries:1973:CAE

[1102] Clark D. Jeffries. Correction: “Analytic evaluation of certain characteristic classes” (Canad. Math. Bull. 16 (1973), 219–223). Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 16(??):623, ???? 1973. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic). See [1039].

Adams:1974:RFG

[1103] John Conway Adams. Rings with a finitely generated total quotient ring. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 17(??):1–4, ???? 1974. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Anderson:1974:LR

[1104] Tim Anderson. On the Levitzki radical. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 17(??):5–10, ???? 1974. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Banaschewski:1974:ECS

[1105] B. Banaschewski. Equational compactness of G-sets. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 17(??):11– 18, ???? 1974. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Berman:1974:CPD

[1106] Abraham Berman. Complementarity problem and duality over convex cones. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 17(??):19–26, ???? 1974. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Bernstein:1974:DE

[1107] Leon Bernstein. On the Diophantine equation x(x + d)(x + 2d) + y(y + d)(y + 2d) = z(z + d)(z + 2d). Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 17(??):27–34, ???? 1974. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Cornish:1974:RNN

[1108] W. H. Cornish and P. N. Stewart. Rings with no nilpotent elements and with the maximum condition on annihilators. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 17(??):35–38, ???? 1974. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Freud:1974:PRV

[1109] Geza Freud. On the problem of R. de Vore. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 17(??):39–44, ???? 1974. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Garner:1974:FPH

[1110] C. W. L. Garner. Frieze patterns in the hyperbolic plane. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 17(??): 45–50, ???? 1974. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Jakimovski:1974:IRG

[1111] A. Jakimovski and J. Tzimbalario. Inclusion relations for general Riesz typical means. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 17(??):51–62, ???? 1974. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Lemire:1974:IR

[1112] F. W. Lemire. An irreducible representation of sl(2). Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 17(??):63– 66, ???? 1974. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Lin:1974:RCO

[1113] C. S. Lin. Regularizers of closed operators. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 17(??):67–72, ???? 1974. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

McQuarrie:1974:NAN

[1114] Bruce McQuarrie. A non-abelian near ring in which (−1)r = r implies r = 0. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 17(??):73–76, ???? 1974. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Moore:1974:SPP

[1115] Edward Moore. A singular perturbation problem and a neutral differential-difference equation. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 17(??):77–84, ???? 1974. CODEN CM-BUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Padmanabhan:1974:SL

[1116] R. Padmanabhan. On M-symmetric lattices. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 17(??):85–86, ???? 1974. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Rival:1974:LDI

[1117] Ivan Rival. Lattices with doubly irreducible elements. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 17(??):91– 96, ???? 1974. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Seyedin:1974:QUS

[1118] Massood Seyedin. Quasi-uniform spaces and topological homeomorphism groups. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 17(??):97–98, ???? 1974. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Thomas:1974:PCB

[1119] R. S. D. Thomas. Partially closed braids. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 17(??):99–108, ???? 1974. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Dressler:1974:TN

[1120] Robert E. Dressler. On a theorem of Niven. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 17(??):109–110, ???? 1974. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Dubuc:1974:EAP

[1121] Serge Dubuc. États accessibles dans un processus de Galton–Watons. (French) [Accessible states in a process of Galton–Watons]. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 17(??): 111–114, ???? 1974. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Dunham:1974:NCB

[1122] Charles B. Dunham. Non-convexity in best complex Chebyshev approximation by rational functions. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 17(??):115–116, ???? 1974. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Giles:1974:NRB

[1123] J. R. Giles. A non-reflexive Banach space has non-smooth third conjugate space. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 17(??):117–120, ???? 1974. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Gupta:1974:PTS

[1124] Hansraj Gupta. A partition theorem of Subbarao. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 17(??):121–124, ???? 1974. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Handelman:1974:MCN

[1125] D. Handelman and P. Selick. A maximality criterion for nilpotent commutative matrix algebras. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 17(??):125–126, ???? 1974. CODEN CM-BUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Hughes:1974:SPI

[1126] Edward Hughes. A short proof of an interpolation theorem. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 17(??): 127–128, ???? 1974. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Losey:1974:RUF

[1127] Gerald Losey. A remark on the units of finite order in the group ring of a finite group. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 17(??):129–130, ???? 1974. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Merris:1974:IPS

[1128] Russell Merris. An inequality for positive semidefinite Hermitian matrices. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 17(??):131–132, ???? 1974. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Michler:1974:EIM

[1129] Gerhard O. Michler. Example of an injective module which is not nice. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 17(??):133–134, ???? 1974. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Schonheim:1974:PPR

[1130] J. Schönheim. On a problem of Purdy related to Sperner systems. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 17 (??):135–136, ???? 1974. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Seth:1974:INR

[1131] V. Seth and K. Tewari. On injective near-ring modules. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 17(??):137–142, ???? 1974. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Siddiqi:1974:MPJ

[1132] Jamil A. Siddiqi. Multiplicative properties of Jensen measures. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 17(??): 143–144, ???? 1974. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Vasilescu:1974:RSO

[1133] F.-H. Vasilescu. The range sequence of an operator. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 17(??):145–158, ???? 1974. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Agrawal:1974:IFG

[1134] Ram K. Agrawal. The influence on a finite group of its permutable subgroups. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 17(??):159–166, ???? 1974. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Borwein:1974:TTB

[1135] D. Borwein and E. Smet. Tauberian theorems for Borel-type methods of summability. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 17(??):167–174, ???? 1974. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic). See addendum [1413].

Mura:1974:ROP

[1136] Roberta Botto Mura. Right-ordered polycyclic groups. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 17(??):175–178, ???? 1974. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Brenner:1974:MER

[1137] J. L. Brenner and M. J. S. Lim. The matrix equations A = XY Z and B = ZY X and related ones. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 17(??):179–184, ???? 1974. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Cook:1974:CTL

[1138] W. D. Cook. Carathéodory’s theorem with linear constraints. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 17(??): 189–192, ???? 1974. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Fischer:1974:I

[1139] Pál Fischer. On the inequality ∑_{i=1}^{n}(f(p_{i})/f(q_{i}))p_{i} ≤ 1. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 17(??):193– 200, ???? 1974. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Formanek:1974:PHG

[1140] Edward Formanek. A problem of Herstein on group rings. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 17(??): 201–202, ???? 1974. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Frucht:1974:SCG

[1141] Roberto Frucht and Frank Harary. Self-complementary generalized orbits of a permutation group. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 17(??):203–208, ???? 1974. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Hsieh:1974:PPC

[1142] Tsu Teh Hsieh and Kok Keong Tan. Periodic points and contractive mappings. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 17(??):209–212, ???? 1974. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Jeremy:1974:MAQ

[1143] Louis Jeremy. Modules et anneaux quasi-continus. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 17(??):217–228, ???? 1974. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Lallement:1974:FSR

[1144] Gérard Lallement. On free semigroups and Ramsey numbers. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 17(??): 229–232, ???? 1974. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Macphail:1974:LFS

[1145] M. S. Macphail and A. Wilansky. Linear functionals and summability invariants. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 17(??):233–242, ???? 1974. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Martindale:1974:NLI

[1146] Wallace S. Martindale III. A note on Lie isomorphisms. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 17(??):243–246, ???? 1974. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Maxson:1974:ESS

[1147] Carlton J. Maxson. Endomorphism semigroups of sums of rings. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 17(??): 247–250, ???? 1974. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Pattabhiraman:1974:GRQ

[1148] M. V. Pattabhiraman. The generalized Rayleigh quotient. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 17(??):251– 256, ???? 1974. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Sehgal:1974:SFC

[1149] V. M. Sehgal. Some fixed and common fixed point theorems in metric spaces. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 17(??):257–260, ???? 1974. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Selby:1974:CRC

[1150] M. A. Selby. On completeness with respect to the Carathéodory metric. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 17(??):261–264, ???? 1974. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Sharma:1974:ICE

[1151] A. Sharma and E. G. Straus. On the irreducibility of a class of Euler Frobenius polynomials. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 17(??):265–274, ???? 1974. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Groetsch:1974:SDL

[1152] C. W. Groetsch. Steepest descent and least squares solvability. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 17(??): 275–276, ???? 1974. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Gupta:1974:DDC

[1153] R. P. Gupta and R. D. Gupta. On the distribution of the direction and collinearity factors in discriminant analysis. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 17(??):277–282, ???? 1974. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Koh:1974:NRR

[1154] Kwangil Koh. On von Neumann regular rings. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 17(??):283–284, ???? 1974. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Koh:1974:CRQ

[1155] Kwangil Koh. On the complete ring of quotients. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 17(??):285–288, ???? 1974. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Mortimer:1974:NPU

[1156] B. C. Mortimer and K. S. Williams. Note on a paper of S. Uchiyama. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 17(??):289–294, ???? 1974. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Pollack:1974:NRC

[1157] Fredric M. Pollack. Numerical range and convex sets. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 17(??):295–296, ???? 1974. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Sawyer:1974:ME

[1158] W. W. Sawyer. On the matrix with elements 1/(r + s − 1). Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 17(??): 297–298, ???? 1974. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Saxena:1974:NRC

[1159] R. B. Saxena. A note on the rate of convergence of Hermite–Fejér interpolation polynomials. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 17(??):299–302, ???? 1974. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Friedler:1974:EOC

[1160] Louis Friedler. Errata to: “Open, connected functions” (Canad. Math. Bull. 16 (1973), 57–60). Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 17(??):303–304, ???? 1974. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic). See [1013].

Erdos:1974:CTP

[1161] P. Erdős and E. C. Milner. Corrigendum: “A theorem in the partition calculus” (Canad. Math. Bull. 15 (1972), 501–505). Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 17(??):305–306, ???? 1974. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic). See [979].

Britten:1974:CDR

[1162] Daniel J. Britten. On Cayley–Dickson rings. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 17(??):625–628, ???? 1974. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Chungphaisan:1974:SRG

[1163] V. Chungphaisan. Sequences realizable by graphs with Hamiltonian squares. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 17(??):629–632, ???? 1974. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Dawson:1974:SAR

[1164] D. A. Dawson. Synchronous and asynchronous reversible Markov systems. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 17(??):633–650, ???? 1974. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Eames:1974:CRS

[1165] W. Eames. On the complete regularity of some category spaces. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 17(??): 651–656, ???? 1974. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Faierman:1974:AFE

[1166] M. Faierman. Asymptotic formulae for the eigenvalues of a two parameter system of ordinary differential equations of the second order. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 17(??): 657–666, ???? 1974. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Feller:1974:CRH

[1167] E. H. Feller. On a class of right hereditary semigroups. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 17(??):667–670, ???? 1974. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Goel:1974:SMA

[1168] D. S. Goel. Selfadjoint metrics on almost tangent manifolds whose Riemannian connection is almost tangent. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 17(??):671–674, ???? 1974. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Hunsaker:1974:HCE

[1169] W. N. Hunsaker and S. A. Naimpally. Hausdorff compactifications as epireflections. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 17(??):675–678, ???? 1974. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Kiang:1974:MRC

[1170] Mo Tak Kiang. Mappings related to contractions. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 17(??):679–684, ???? 1974. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Labelle:1974:LTM

[1171] Gilbert Labelle. L’arithmétique toposienne des mots naturels. (French) [The Toposian arithmetic of natural words]. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 17(??):685–688, ???? 1974. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Lin:1974:BPN

[1172] C. S. Lin. Boundary points of the numerical range of an operator. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 17(??): 689–692, ???? 1974. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

McCandless:1974:ETN

[1173] W. L. McCandless. An existence theorem for nonlinear boundary value problems. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 17(??):693–702, ???? 1974. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Owusu-Ansah:1974:STH

[1174] T. Owusu-Ansah. A spectral theorem for Hermitian operators of meromorphic type on Banach spaces. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 17(??):703–708, ???? 1974. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Raphael:1974:SRR

[1175] R. Raphael. Some remarks on regular and strongly regular rings. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 17(??): 709–712, ???? 1974. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Read:1974:WPN

[1176] John A. Read. Wreath products of nonoverlapping lattice ordered groups. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 17(??):713–722, ???? 1974. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Sweet:1974:IQD

[1177] Lowell Sweet. On involutions of quasi-division algebras. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 17(??):723– 726, ???? 1974. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Szeto:1974:SDR

[1178] George Szeto. On a sheaf of division rings. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 17(??):727–732, ???? 1974. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Tanny:1974:SNR

[1179] Stephen M. Tanny. On some numbers related to the Bell numbers. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 17 (??):733–738, ???? 1974. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Varadarajan:1974:PD

[1180] K. Varadarajan. On product decompositions. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 17(??):739–746, ???? 1974. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Charnow:1974:NTF

[1181] A. Charnow. A note on torsion free groups generated by pairs of matrices. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 17(??):747–748, ???? 1974. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Faierman:1974:NKO

[1182] M. Faierman. A note on Klein’s oscillation theorem for periodic boundary conditions. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 17(??):749–756, ???? 1974. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Mikhail:1974:ENN

[1183] N. N. Mikhail and D. S. Tracy. The exact non-null distribution of Wilks’ Λ criterion in the bivariate collinear case. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 17(??):757–758, ???? 1974. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Smith:1974:HCC

[1184] D. H. Smith. Hamiltonian circuits on the N-cube. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 17(??):759–762, ???? 1974. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Zaks:1974:HCP

[1185] Joseph Zaks. Hamiltonian cycles in products of graphs. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 17(??):763–765, ???? 1974. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Rankin:1974:OL

[1186] S. A. Rankin. Orbital lattices of SL^{+}(2, Z). Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 17(1):87–90, March 1974. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Burkowski:1974:OBF

[1187] Forbes J. Burkowski and Peter J. Ponzo. The oscillatory behavior of a first order non-linear differential equation with delay. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 17(2):185– 188, June 1974. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Husain:1974:BML

[1188] T. Husain and S. B. Ng. Boundedness of multiplicative linear functionals. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 17(2):213–216, June 1974. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Hastings:1974:CST

[1189] Harold M. Hastings. Correction to “Simplicial topological resolutions”. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 17(3):307–316, ???? 1974. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Hayman:1974:LGP

[1190] W. K. Hayman. The local growth of power series: a survey of the Wiman– Valiron method. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 17(3):317–358, September 1974. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Bercov:1974:PGR

[1191] R. D. Bercov. On permutation groups with regular subgroup. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 17(3): 359–362, September 1974. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Cheung:1974:AG

[1192] Alan L. C. Cheung. Adjoints of a geometry. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 17(3):363–366, September 1974. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Fleischner:1974:RSD

[1193] Herbert Fleischner. On regulating sets and the disparity of planar cubic graphs. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 17(3):367–374, September 1974. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Goodaire:1974:DA

[1194] Edgar G. Goodaire and Roy C. Snell. The derivation algebra of M_{4}^{8}(C). Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 17(3):375–378, September 1974. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Kennedy:1974:EPF

[1195] E. W. Kennedy. An elementary proof of the Frobenius factorization theorem for differential equations. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 17(3):379–380, September 1974. CO-DEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Lemire:1974:COS

[1196] F. W. Lemire. Connected order 3 standard representations of simple Lie algebras. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 17(3):381–384, September 1974. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Shum:1974:GSF

[1197] D. T. Shum. A general and sharpened form of Opial’s inequality. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 17(3): 385–390, September 1974. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Spoar:1974:SPN

[1198] G. Spoar and N. D. Lane. On singular points of normal arcs of cyclic order four. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 17(3):391–396, September 1974. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Teufel:1974:SSG

[1199] Hugo Teufel, Jr. The stability of solutions of generalized Emden–Fowler equations. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 17(3):397–402, September 1974. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Westwick:1974:APM

[1200] R. Westwick. On adjacency preserving maps. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 17(3):403–406, September 1974. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Whitfield:1974:QDN

[1201] J. H. M. Whitfield. Quasi-differentiable norms. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 17(3):407–408, September 1974. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Brand:1974:WMP

[1202] Daniel Brand. When is a matrix positive? Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 17(3):409–410, September 1974. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Griffith:1974:NCR

[1203] Gareth J. Griffith. A note concerning the real inflexions of real, plane, algebraic curves. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 17(3):411–412, September 1974. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Johnson:1974:GPM

[1204] Dudley Paul Johnson. Gaussian processes with Markovian covariances. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 17(3):413–416, September 1974. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Kabe:1974:GSL

[1205] D. G. Kabe. Generalized Sverdrup’s lemma and the treatment of less than full rank regression model. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 17(3):417–420, September 1974. CO-DEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

McCabe:1974:NCS

[1206] T. F. McCabe. A note on contractive selfmaps of itl^{∞}. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 17(3):421–428, September 1974. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Eberhart:1974:ESW

[1207] C. Eberhart, J. B. Fugate, and L. Mohler. Errata to “Spaces which cannot be written as a countable disjoint union of closed subsets”. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 17 (3):429, September 1974. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic). See [1070].

Babai:1974:MOG

[1208] László Babai. On the minimum order of graphs with given group. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 17(4): 467–470, December 1974. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Bacsich:1974:MTE

[1209] Paul D. Bacsich. Model theory of epimorphisms. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 17(4):471–478, December 1974. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Benabdallah:1974:PHS

[1210] K. Benabdallah. On pure-high subgroups of abelian groups. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 17(4): 479–482, December 1974. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Chan:1974:VBS

[1211] W. M. Chan and P. Hoffman. Vector bundles over suspensions. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 17(4): 483–484, December 1974. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Cramer:1974:DSA

[1212] T. Cramer. A definition of separation axiom. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 17(4):485–492, December 1974. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Duncan:1974:SO

[1213] Richard Duncan. On 2-summing operators. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 17(4):493–498, December 1974. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Eliason:1974:LIB

[1214] Stanley B. Eliason. Lyapunov inequalities and bounds on solutions of certain second order equations. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 17(4):499–504, December 1974. CO-DEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Etigson:1974:CFE

[1215] L. B. Etigson. A cosine functional equation with restricted argument. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 17(4):505–510, December 1974. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Fleury:1974:NDG

[1216] Patrick Fleury. A note on dualizing Goldie dimension. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 17(4):511– 518, December 1974. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Fox:1974:GTT

[1217] Geoffrey Fox and Pedro Morales. A general Tychonoff theorem for multifunctions. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 17(4):519–522, December 1974. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Goel:1974:BSA

[1218] D. S. Goel, A. S. B. Holland, C. Nasim, and B. N. Sahney. On best simultaneous approximation in normed linear spaces. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 17(4):523–528, December 1974. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Hartfiel:1974:LBP

[1219] D. J. Hartfiel. A lower bound for the permanent on a special class of matrices. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 17(4):529–530, December 1974. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Horn:1974:NWS

[1220] P. J. Horn. A note on weakly symmetric rings. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 17(4):531–534, December 1974. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Lee:1974:BTS

[1221] C. Y. Lee. On the branching theorem of the symplectic groups. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 17(4): 535–546, December 1974. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Minc:1974:RTM

[1222] Henryk Minc. A remark on a theorem of M. Hall. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 17(4):547–548, December 1974. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Osborne:1974:DPC

[1223] R. P. Osborne. On the 4-dimensional Poincaré conjecture for manifolds with 2-dimensional spines. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 17(4):549–552, December 1974. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Pacholke:1974:IDW

[1224] Kenneth Pacholke. Integral domains which have finite character locally. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 17(4):553–558, December 1974. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Page:1974:WFT

[1225] S. Page. When flats are torsion free. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 17(4):559–562, December 1974. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Rowlands:1974:NMO

[1226] K. Rowlands. A note on multiplier operators and dual B^{∗}-algebras. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 17(4):563–566, December 1974. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Snell:1974:RIM

[1227] Roy C. Snell. The range of invariant means on locally compact abelian groups. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 17(4):567–574, December 1974. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Spohn:1974:DC

[1228] W. G. Spohn. On the derived cuboid. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 17(4):575–578, December 1974. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Whitfield:1974:NRS

[1229] J. H. M. Whitfield. A non-reflexive smooth space with a smooth dual. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 17(4):579–580, December 1974. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Wong:1974:FPT

[1230] Chi Song Wong. Fixed point theorems for point-to-set mappings. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 17(4): 581–586, December 1974. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Zakon:1974:INM

[1231] Elias Zakon. Integration of non-measurable functions. II. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 17(4):587– 592, December 1974. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Arjomandi:1974:UGS

[1232] E. Arjomandi and D. G. Corneil. Unicyclic graphs satisfy Harary’s conjecture. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 17(4):593–596, December 1974. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Mura:1974:RCN

[1233] Roberta Botto Mura. Radical classes need not have a unique maximal R_{0}-closed subclass. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 17(4):597–598, December 1974. CODEN CM-BUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Erdos:1974:ALN

[1234] Paul Erdős. On abundant-like numbers. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 17(4):599–602, December 1974. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Giroux:1974:RIF

[1235] André Giroux. Une remarque sur l’interpolation des fonctions entières. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 17(4):603–604, December 1974. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Koh:1974:NUS

[1236] E. L. Koh. Non-uniqueness of the solution to a generalized Dirichlet problem. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 17(4):605–606, December 1974. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Koh:1974:LI

[1237] Kwangil Koh. On lifting idempotents. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 17(4):607–608, December 1974. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Shafaat:1974:NMC

[1238] Ahmad Shafaat. A note on Mal′cevian varieties. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 17(4):609–610, December 1974. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Sobich:1974:SPI

[1239] Lucja Sobich and Dominik Szynal. Some properties of inflated binomial distribution. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 17(4):611–614, December 1974. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Ursell:1974:DRE

[1240] H. D. Ursell. The degrees of radical extensions. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 17(4):615–624, December 1974. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Bagsoze:1975:SRS

[1241] Türkân Bagsöze. Some results on spiral-like functions. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 18(??):633–638, ???? 1975. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Chang:1975:SAP

[1242] K. W. Chang. Stability of almost periodic solutions of an autonomous equation. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 18(??):639–642, ???? 1975. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Ching:1975:SNH

[1243] Wai Mee Ching. On some non-hyperfinite factors of type III. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 18(??): 643–648, ???? 1975. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Chu:1975:SPV

[1244] Le Quang Chu. A short proof of Vladimirskii’s theorem on precompact perturbations in locally convex spaces. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 18(??):649–656, ???? 1975. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Dobbs:1975:GD

[1245] David E. Dobbs. On the global dimensions of D + M. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 18(??):657–660, ???? 1975. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Elliott:1975:RND

[1246] George A. Elliott. On the Radon–Nikodým derivative with a chain rule in a von Neumann algebra. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 18(??):661–670, ???? 1975. CODEN CM-BUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Grant:1975:OMC

[1247] Douglass L. Grant. The open mapping and closed graph theorem for embeddable topological semigroups. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 18(??):671–674, ???? 1975. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Janos:1975:ECM

[1248] Ludvik Janos. On the Edelstein contractive mapping theorem. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 18(??): 675–678, ???? 1975. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Lewis:1975:BSW

[1249] J. E. Lewis. A Banach space whose elements are classes of sets of constant width. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 18(??):679–690, ???? 1975. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Kotzig:1975:DVF

[1250] Anton Kotzig. Degrees of vertices in a friendship graph. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 18(??):691–694, ???? 1975. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Lee:1975:APA

[1251] K. K. Lee. Another possible abnormality of compact space-time. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 18(??): 695–698, ???? 1975. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Morales:1975:ETS

[1252] Pedro Morales. Extension of a tight set function with values in a locally convex space. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 18(??):699–702, ???? 1975. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Schonheim:1975:PCC

[1253] J. Schönheim and A. Bialostocki. Packing and covering of the complete graph with 4-cycles. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 18(??):703–708, ???? 1975. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Shiau:1975:CQN

[1254] Chyi Shiau, Kok Keong Tan, and Chi Song Wong. A class of quasinonexpansive multi-valued maps. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 18(??):709–714, ???? 1975. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Srivastava:1975:SPI

[1255] T. N. Srivastava and Shôji Watanabe. Some properties of indicatrices in a Finsler space. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 18(??):715–722, ???? 1975. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Stewart:1975:SRC

[1256] Patrick N. Stewart. Strong radical classes and idempotents. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 18(??): 723–726, ???? 1975. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Wilker:1975:SHT

[1257] J. B. Wilker. Sets homothetic to their intersection with a translate. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 18(??):739–748, ???? 1975. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Couture:1975:RDT

[1258] Raymond Couture and Mario Lavoie. Réciproque du théorème des bases sur un espace vectoriel. (French) [Reciprocal of the base theorem in a vector space]. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 18(??):749–752, ???? 1975. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Hwang:1975:RGS

[1259] J. S. Hwang. The range of a gap series. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 18(??):753–754, ???? 1975. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Keener:1975:URT

[1260] L. Keener. Using Rolle’s theorem in exponential function-derivative approximation. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 18(??):755–758, ???? 1975. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Lightstone:1975:DDF

[1261] A. H. Lightstone and Kam Wong. Dirac delta functions via nonstandard analysis. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 18(??):759–762, ???? 1975. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Macedonska-Nosalska:1975:TPC

[1262] O. N. Macedońska-Nosalska. On two problems concerning neutral polyverbal operations on groups. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 18(??):763–766, ???? 1975. CO-DEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Roberts:1975:PFF

[1263] James W. Roberts. Pointwise finite families of mappings. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 18(??):767– 768, ???? 1975. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Tanny:1975:AGT

[1264] Stephen M. Tanny. An application of a generalization of Terquem’s problem. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 18(??):769–770, ???? 1975. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Barr:1975:EIE

[1265] Michael Barr. The existence of injective effacements. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 18(1):1–6, March 1975. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Bellamy:1975:ERL

[1266] O. S. Bellamy and H. W. Ellis. On extensions of the Riemann and Lebesgue integrals by nets. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 18(1):7–18, March 1975. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Binding:1975:SMM

[1267] P. Binding and P. J. Browne. L^{p} spaces from matrix measures. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 18(1): 19–26, March 1975. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Bovey:1975:CZS

[1268] J. D. Bovey, Paul Erdős, and Ivan Niven. Conditions for a zero sum modulo n. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 18(1):27–30, March 1975. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Cockayne:1975:RGN

[1269] E. J. Cockayne and P. J. Lorimer. On Ramsey graph numbers for stars and stripes. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 18(1):31–34, March 1975. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Dunham:1975:CSW

[1270] Charles B. Dunham. Chebyshev sets in C[0, , 1] which are not sums. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 18(1):35–38, March 1975. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Fournier:1975:MPS

[1271] J. J. F. Fournier and P. M. Gauthier. Most power series have radius of convergence 0 or 1. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 18(1):39–40, March 1975. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Jorgensen:1975:RUL

[1272] Murray Jorgensen. Regular ultrafilters and long ultrapowers. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 18(1): 41–44, March 1975. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Keener:1975:BPN

[1273] L. L. Keener. Best possible nets in a normed linear space. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 18(1): 45–48, March 1975. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Lamb:1975:CAO

[1274] Charles W. Lamb. Convergence of averaged occupation times. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 18(1): 49–56, March 1975. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Lindstrom:1975:GPM

[1275] Bernt Lindström. Group partitions and mixed perfect codes. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 18(1): 57–60, March 1975. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Lund:1975:MIS

[1276] Bruce Lund. The maximal ideal space of subalgebras of the disk algebra. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 18(1):61–66, March 1975. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Maddox:1975:ACS

[1277] I. J. Maddox and J. W. Roles. Absolute convexity in spaces of strongly summable sequences. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 18(1):67–76, March 1975. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Nicholson:1975:SM

[1278] W. K. Nicholson. On semiperfect modules. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 18(1):77–80, March 1975. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Rao:1975:WAP

[1279] Aribindi Satyanarayan Rao. On weakly almost-periodic families of linear operators. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 18(1):81–86, March 1975. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Rao:1975:FED

[1280] V. Venugopal Rao. Functional equations of Dirichlet series derived from non-analytic automorphic forms of a certain type. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 18(1):87–94, March 1975. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Rival:1975:SML

[1281] Ivan Rival. Sublattices of modular lattices of finite length. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 18(1): 95–98, March 1975. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Rompke:1975:RCM

[1282] Jurgen Rompke. Regular, commutative, maximal semigroups of quotients. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 18(1):99–104, March 1975. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Singh:1975:SAE

[1283] R. Singh. Some admissible estimators in extreme value densities. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 18(1): 105–110, March 1975. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Szabo:1975:CEC

[1284] M. E. Szabo. A counter-example to coherence in Cartesian closed categories. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 18(1):111–114, March 1975. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Wall:1975:FUP

[1285] Charles R. Wall. The fifth unitary perfect number. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 18(1):115–122, March 1975. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Williams:1975:NIR

[1286] Kenneth S. Williams. Note on integers representable by binary quadratic forms. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 18(1):123–126, March 1975. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Wong:1975:ACM

[1287] James C. S. Wong. Absolutely continuous measures on locally compact semigroups. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 18(1):127–132, March 1975. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Abbott:1975:PFS

[1288] H. L. Abbott. On property B of families of sets. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 18(1):133–136, March 1975. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Casazza:1975:PHB

[1289] P. G. Casazza and Bor Luh Lin. Perfectly homogeneous bases in Banach spaces. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 18(1):137–140, March 1975. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Fieldhouse:1975:JRR

[1290] David J. Fieldhouse. The Jacobson radical and regular modules. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 18(1): 141–142, March 1975. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Gardner:1975:EUC

[1291] L. T. Gardner and P. Milnes. On the extension of uniformly continuous functions. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 18(1):143–146, March 1975. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Lin:1975:CCH

[1292] C. S. Lin. On the closure of the convex hull of a set. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 18(1):147– 148, March 1975. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Murphy:1975:NMR

[1293] I. S. Murphy. A note on a matrix result of Ryser. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 18(1):149–150, March 1975. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Tupper:1975:GCB

[1294] B. O. J. Tupper. Geodesic correspondence in the Brans–Dicke theory. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 18(1):151–154, March 1975. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Webb:1975:TRC

[1295] William A. Webb. On a theorem of Rav concerning Egyptian fractions. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 18(1):155–158, March 1975. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Adams:1975:WHO

[1296] M. E. Adams. Weakly homogeneous order types. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 18(2):159–162, June 1975. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Brown:1975:DPD

[1297] Jeffrey M. Brown. Direct product decompositions of twisted wreath products. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 18(2):163–172, June 1975. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Campbell:1975:NFT

[1298] C. M. Campbell and E. F. Robertson. A note on Fibonacci type groups. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 18(2):173–176, June 1975. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Eng-Ung:1975:NS

[1299] Choo Eng-Ung. Note on a subring of C^{∗}(X). Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 18(2):177–180, June 1975. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Clements:1975:EUS

[1300] John C. Clements. On the existence and uniqueness of solutions of the equation u_{tt} − ∂σ_{i}(u_{x_{i}})/∂x_{i} − D_{N}u_{t} = f. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 18(2):181–188, June 1975. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Cohen:1975:SSA

[1301] Miriam Cohen and Susan Montgomery. Semi-simple Artinian rings of fixed points. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 18(2):189–190, June 1975. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Cooper:1975:SCS

[1302] J. B. Cooper. L^{1}-sequential convergence on Stone spaces and Stone’s theorem on unitary semigroups. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 18(2):191–194, June 1975. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Duggal:1975:SCA

[1303] K. L. Duggal. On the spheres carrying an almost contingent structure. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 18(2):195–202, June 1975. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Dunn:1975:SQR

[1304] Samuel L. Dunn. On the structure of quotient rings which are QFX rings. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 18(2):203–208, June 1975. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Gonshor:1975:WBR

[1305] H. Gonshor. On GL_{n}(B) where B is a Boolean ring. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 18(2):209– 216, June 1975. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Hall:1975:ICG

[1306] Japheth Hall, Jr. Independence in combinatiorial geometries of rank three. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 18(2):217–222, June 1975. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Hartfiel:1975:CFF

[1307] D. J. Hartfiel. A canonical form for fully indecomposable (0, , 1)-matrices. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 18(2):223–228, June 1975. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Higgins:1975:CPT

[1308] Rada Higgins. On certain problems in the theory of sequences. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 18(2): 229–232, June 1975. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Lang:1975:RPI

[1309] N. C. Lang. On ring properties of injective hulls. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 18(2):233–240, June 1975. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Mason:1975:GNR

[1310] Gordon Mason. W -groups and near-ring modules. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 18(2):241–244, June 1975. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

McLeish:1975:CLT

[1311] D. L. McLeish. A central limit theorem with conditioning on the distant past. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 18(2):245–248, June 1975. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Ortel:1975:ILS

[1312] Marvin Ortel. Interpolation by linear sums of harmonic measures. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 18 (2):249–254, June 1975. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Redfield:1975:APD

[1313] R. H. Redfield. Almost polar-dense lattices. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 18(2):255–262, June 1975. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Ruttimann:1975:DPO

[1314] G. T. Rüttimann. Decomposition of projections on orthomodular lattices. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 18(2):263–268, June 1975. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Spencer:1975:RBB

[1315] James D. Spencer. Relations between boundary value functions for a nonlinear differential equation and its variational equations. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 18(2): 269–276, June 1975. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Swiatak:1975:CEF

[1316] Halina Światak. A characterization of exponential functions with nonlinear exponents. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 18(2):277–282, June 1975. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Vertesi:1975:PPT

[1317] P. O. H. Vértesi. On a problem of P. Turán (Ann. Univ. Sci. Budapest. Eötvös Sect. Math. 3–4 (1960/61), 369–377). Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 18(2):283–288, June 1975. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Beauregard:1975:IRI

[1318] Raymond A. Beauregard. On the inversion of right invariant elements. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 18(2):289–290, June 1975. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Beuerman:1975:LDS

[1319] D. R. Beuerman. Limit distributions for sums of weighted random variables. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 18(2):291–294, June 1975. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Dawes:1975:SPB

[1320] A. M. Dawes. Separation principles and bounded quantification. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 18(2): 295–296, June 1975. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Karvellas:1975:ESE

[1321] Paul H. Karvellas. Extension of a semigroup embedding theorem to semirings. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 18(2):297–298, June 1975. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Parameswaran:1975:NSP

[1322] M. R. Parameswaran. A note on some prime Hausdorff methods of summability. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 18(2):299–316, June 1975. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Abramson:1975:SNR

[1323] Morton Abramson. Sequences by number of w-rises. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 18(3):317–320, September 1975. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Browne:1975:CTS

[1324] Patrick J. Browne. Commuting k-tuples of self adjoint operators and matrix measures. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 18(3):321–326, September 1975. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Campbell:1975:OGI

[1325] Stephen L. Campbell and Carl D. Meyer. EP operators and generalized inverses. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 18(3):327–334, September 1975. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Essen:1975:CBT

[1326] M. Essén and H. L. Jackson. A comparison between thin sets and generalized Azarin sets. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 18(3):335–346, September 1975. CODEN CM-BUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Gentry:1975:NUB

[1327] Rodney D. Gentry. A note on upper bounds for the eigenvalues of y′′ + λpy = 0. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 18(3):347–352, September 1975. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Henderson:1975:PMH

[1328] J. R. Henderson. Permanents of (0, 1)-matrices having at most two zeros per line. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 18(3):353–358, September 1975. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Jain:1975:QIP

[1329] S. K. Jain and Surjeet Singh. Quasi-injective and pseudo-injective modules. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 18(3):359–366, September 1975. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Janos:1975:CPO

[1330] Ludvik Janos. Contraction property of the operator of integration. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 18(3):367–370, September 1975. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Jeffries:1975:SCD

[1331] Clark Jeffries. Stability of certain dynamical systems. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 18(3):371–374, September 1975. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Lund:1975:FBD

[1332] Bruce Lund. Functions belonging to a Dirichlet subalgebra of the disk algebra. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 18(3):375–378, September 1975. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Malik:1975:UGS

[1333] M. A. Malik. Uniqueness of generalized solutions of abstract differential equations. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 18(3):379–382, September 1975. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Muhly:1975:IFI

[1334] Paul S. Muhly. Inner functions and isometries. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 18(3):383–386, September 1975. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Nelson:1975:IEC

[1335] Evelyn Nelson. Injectivity and equational compactness in the class of ℵ_{0}-semilattices. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 18(3):387–392, September 1975. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Rao:1975:SOD

[1336] Aribindi Satyanarayan Rao. On second-order differential operators with Bohr–Neuigebauer type property. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 18(3):393–396, September 1975. CO-DEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Schaufele:1975:API

[1337] Ronald A. Schaufele. An application of pairwise independence of random variables to regression analysis. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 18(3):397–404, September 1975. CO-DEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Tan:1975:PML

[1338] Peter Tan and Constantin Drossos. A property of maximum likelihood estimators for invariant statistical models. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 18(3):405–410, September 1975. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Tymchatyn:1975:POC

[1339] E. D. Tymchatyn. Partial orders on the 2-cell. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 18(3):411–416, September 1975. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Wang:1975:UTD

[1340] Chung Lie Wang. A uniqueness theorem on the degenerate Cauchy problem. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 18(3):417–422, September 1975. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Chacron:1975:NRE

[1341] M. Chacron, J. Lawrence, and D. Madison. A note on radical extensions of rings. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 18(3):423–424, September 1975. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Davison:1975:RSS

[1342] J. L. Davison. A result on sums of squares. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 18(3):425–426, September 1975. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Gould:1975:ESG

[1343] Matthew Gould. An equational spectrum giving cardinalities of endomorphism monoids. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 18(3):427–430, September 1975. CODEN CM-BUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Husain:1975:CSB

[1344] T. Husain and S. M. Khaleelulla. On countably, σ−, and sequentially barrelled spaces. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 18(3):431–432, September 1975. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Jones:1975:FTP

[1345] James P. Jones. Formula for the Nth prime number. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 18(3):433–434, September 1975. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Leblanc:1975:CEP

[1346] Raymond Leblanc. Caractérisation des espaces parfaits. (French) [Characterization of perfect spaces]. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 18(3):435–436, September 1975. CO-DEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Lee:1975:CSS

[1347] K. K. Lee. Comments on the spinor structure of space-time. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 18(3): 437–438, September 1975. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Melzak:1975:DF

[1348] Z. A. Melzak. On the derivatives of the G-function. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 18(3):439–442, September 1975. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Zassenhaus:1975:MHD

[1349] Hans J. Zassenhaus. On the Minkowski–Hilbert dialogue on mathematization. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 18(3):443–474, September 1975. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Al-Hussaini:1975:CFP

[1350] A. N. Al-Hussaini. Contractions with fixed points and conditional expectation. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 18(4):475–478, December 1975. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Andrews:1975:TCO

[1351] George E. Andrews. The theory of compositions. I. The ordered factorizations of n and a conjecture of C. Long. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 18(4):479–484, December 1975. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Betts:1975:TRM

[1352] J. R. Betts and A. C. Thompson. Three remarks on the measurement of unit spheres. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 18(4):485–488, December 1975. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Braess:1975:PCS

[1353] Dietrich Braess. An N-parameter Chebyshev set which is not a sun. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 18(4):489–492, December 1975. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Cross:1975:I

[1354] George Cross. The P^{n}-integral. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 18(4):493–498, December 1975. CO-DEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

deBruyn:1975:CVS

[1355] G. F. C. de Bruyn. Concepts in vector spaces with convergence structures. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 18(4):499–502, December 1975. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Fournier:1975:IEE

[1356] Gilles Fournier. Identité entre l’ensemble des fonctions de Baire et l’ensemble des fonctions boréliennes. (French) [Identity between the set of Baire functions and the set of Borelian functions]. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 18(4):503–506, December 1975. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Haruki:1975:FEA

[1357] Hiroshi Haruki. A functional equation arising from Ivory’s theorem in geometry. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 18(4):507–516, December 1975. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Kreith:1975:ASP

[1358] Kurt Kreith. Analogy in solving Poisson’s equation. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 18(4):517–522, December 1975. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Lavoie:1975:APG

[1359] Mario Lavoie. Approximation polynômiale généralisée dans certains espaces semi-normés. (French) [Generalized polynomial approximation in certain half-normed spaces]. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 18(4):523–528, December 1975. CO-DEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Leeming:1975:SPC

[1360] David J. Leeming. Some properties of a certain set of interpolating polynomials. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 18(4):529–538, December 1975. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Lemaire:1975:MMA

[1361] Claude Lemaire. Morphismes et multiplications dans une algèbre. (French) [Morphisms and multiplications in an algebra]. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 18(4):539– 542, December 1975. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Lemire:1975:NFI

[1362] F. W. Lemire. A new family of irreducible representations of A_{n}. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 18(4): 543–546, December 1975. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Mallory:1975:ESF

[1363] D. Mallory. Extension of set functions to measures and applications to inverse limit measures. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 18(4):547–554, December 1975. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Mehra:1975:WCO

[1364] K. L. Mehra. Weak convergence and one-sample rank statistics under ϕmixing. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 18(4):555–566, December 1975. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Parmenter:1975:IGR

[1365] M. M. Parmenter. Isomorphic group rings. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 18(4):567–576, December 1975. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Peterson:1975:FCP

[1366] Allan C. Peterson and Dwight V. Sukup. On the first conjugate point function for nonlinear differential equations. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 18(4):577–586, December 1975. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Porter:1975:EFS

[1367] Jack R. Porter. Extension function and subcategories of HAUS. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 18(4): 587–590, December 1975. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Seidman:1975:RTS

[1368] Thomas I. Seidman. A remark on Talenti’s semigroup. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 18(4):591– 592, December 1975. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Shawyer:1975:MMT

[1369] B. L. R. Shawyer. Maximum modulus theorems and Schwarz lemmata for sequence spaces. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 18(4):593–596, December 1975. CODEN CM-BUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Wigley:1975:BAS

[1370] Neil M. Wigley. A Banach algebra structure for H^{p}. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 18(4):597–604, December 1975. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Conlan:1975:FDT

[1371] James Conlan and E. L. Koh. A fractional differentiation theorem for the Laplace transform. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 18(4):605–606, December 1975. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Edmunds:1975:SCP

[1372] Charles C. Edmunds. A short combinatorial proof of the Vaught conjecture. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 18(4):607–608, December 1975. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Herstein:1975:RF

[1373] I. N. Herstein. On a result of Faith. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 18(4):609–610, December 1975. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Warrack:1975:RQM

[1374] Brian Warrack. Regular quotients of metric spaces. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 18(4):611–632, December 1975. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Strzelecki:1975:S

[1375] E. Strzelecki. On sequences ξt_{n} (mod 1). Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 18(5):727–738, ???? 1975. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Berry:1976:MWC

[1376] David Berry. Modules whose cyclic submodules have finite dimension. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 19(1):1–6, March 1976. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Bogin:1976:GFP

[1377] Joseph Bogin. A generalization of a fixed point theorem of Goebel, Kirk and Shimi. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 19(1):7–12, March 1976. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Broverman:1976:TEP

[1378] S. Broverman. The topological extension of a product. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 19(1):13–20, March 1976. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Burke:1976:PRR

[1379] John L. Burke. On the property (PU) for ∗-regular rank rings. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 19(1): 21–38, March 1976. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Carter:1976:SRQ

[1380] Karen S. Carter and T. L. Hicks. Some results on quasi-uniform spaces. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 19(1):39–52, March 1976. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Chew:1976:RMT

[1381] Kim Peu Chew and Kok Keong Tan. On the representation of mappings of Tychonov spaces as restrictions of linear transformations. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 19(1): 53–58, March 1976. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Dolan:1976:PJT

[1382] S. W. Dolan. A proof of Jacobson’s theorem. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 19(1):59–62, March 1976. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Hiller:1976:DFT

[1383] Howard Lyn Hiller. Derived functors of torsion. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 19(1):63–66, March 1976. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Lim:1976:RVS

[1384] Ming Huat Lim. Rank k vectors in symmetry classes of tensors. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 19(1): 67–76, March 1976. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

McDowell:1976:PNR

[1385] Kenneth P. McDowell. Pseudo-Noetherian rings. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 19(1):77–84, March 1976. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Selliah:1976:ISG

[1386] J. B. Selliah. An inequality satisfied by the gamma function. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 19(1): 85–88, March 1976. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Sing:1976:SRM

[1387] Fuk Yum Sing. Some results on matrices with prescribed diagonal elements and singular values. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 19(1):89–92, March 1976. CODEN CM-BUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Thierrin:1976:CCS

[1388] G. Thierrin. A characterization of the commutator subgroup of a group. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 19(1):93–94, March 1976. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Vincent:1976:MTM

[1389] Paul A. Vincent. A metrization theorem for 2-manifolds. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 19(1):95–104, March 1976. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Derrick:1976:GEU

[1390] W. Derrick and L. Janos. A global existence and uniqueness theorem for ordinary differential equations. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 19(1):105–108, March 1976. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Gruman:1976:ISE

[1391] Lawrence Gruman. Interpolation in spaces of entire functions in fC^{N}. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 19(1):109–112, March 1976. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Mayne:1976:CAP

[1392] Joseph H. Mayne. Centralizing automorphisms of prime rings. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 19(1): 113–116, March 1976. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Shapiro:1976:NEL

[1393] H. L. Shapiro and F. A. Smith. A note on extending locally finite collections. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 19(1):117–128, March 1976. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Adams:1976:UCF

[1394] R. A. Adams. The uniform continuity of functions in Sobolev spaces. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 19(2):129–136, June 1976. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Allegretto:1976:OCS

[1395] W. Allegretto. Oscillation criteria for semilinear equations in general domains. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 19(2):137–144, June 1976. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Britten:1976:SAJ

[1396] Daniel J. Britten. On semiprime ample Jordan rings. J ≤ H with chain condition. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 19(2):145–148, June 1976. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

DAristotle:1976:NCE

[1397] Anthony J. D’Aristotle. A note on H-closed extensions of a product. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 19(2):149–154, June 1976. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Dunham:1976:ACA

[1398] Charles B. Dunham. Alternating Chebyshev approximation with a noncontinuous weight function. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 19(2):155–158, June 1976. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Easwaran:1976:QFO

[1399] S. Easwaran. Quadratic functionals of nth order. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 19(2):159–168, June 1976. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Fleischner:1976:HCS

[1400] Herbert Fleischner and Arthur M. Hobbs. Hamiltonian cycles in squares of vertex-unicyclic graphs. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 19(2):169–172, June 1976. CODEN CM-BUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Gordon:1976:PMM

[1401] Robert Gordon. Polynomial modules over Macaulay modules. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 19(2): 173–176, June 1976. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Grassmann:1976:SPC

[1402] E. Grassmann, W. Hengartner, and G. Schober. Support points of the class of close-to-convex functions. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 19(2):177–180, June 1976. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Hemdan:1976:SCF

[1403] H. T. Hemdan. Solution of the ‘cube’ functional equation in terms of ‘trilinear’ coefficients. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 19(2):181–192, June 1976. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Kaul:1976:LCS

[1404] Saroop K. Kaul. Local compactness in set valued function spaces. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 19 (2):193–198, June 1976. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Neggers:1976:RME

[1405] Joseph Neggers. On rings with many endomorphisms. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 19(2):199– 204, June 1976. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Nicholson:1976:LRE

[1406] W. K. Nicholson and H. J. Springer. Local rings with elementary Abelian units. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 19(2):205–208, June 1976. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Swaminathan:1976:GPM

[1407] V. Swaminathan. The Gibbs phenomenon for [S, α_{n}] means and [T, α_{n}] means. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 19(2):209–212, June 1976. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Tomkinson:1976:CAP

[1408] M. J. Tomkinson. Cover-avoidance properties in finite soluble groups. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 19(2):213–216, June 1976. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Vranch:1976:CS

[1409] J. Vranch. On critical r, λ-systems. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 19(2):217–220, June 1976. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Wagneur:1976:RPS

[1410] E. Wagneur. Réduction des points singuliers des feuilletages à singularités non dégénérées de M^{3}. (French) [Reduction of singular points of foliations with nondegenerate singularities in M^{3}]. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 19(2):221–230, ???? 1976. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Wong:1976:ASS

[1411] James C. S. Wong. Amenability and substantial semigroups. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 19(2): 231–234, June 1976. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Zaidman:1976:CRW

[1412] S. Zaidman. A convexity result for weak differential inequalities. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 19(2): 235–244, June 1976. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Borwein:1976:ATT

[1413] D. Borwein and E. Smet. Addendum to: “Tauberian theorems for Boreltype methods of summability” (Canad. Math. Bull. 17 (1974), 167–173). Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 19(2):245–246, June 1976. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic). See [1135].

Campbell:1976:GPD

[1414] C. M. Campbell and E. F. Robertson. On a group presentation due to Fox. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 19(2):247–256, June 1976. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Borwein:1976:TCC

[1415] J. Borwein and R. O’Brien. Tangent cones and convexity. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 19(3): 257–262, September 1976. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Burns:1976:CES

[1416] R. G. Burns, C. C. Edmunds, and I. H. Farouqi. On commutator equalities and stabilizers in free groups. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 19(3):263–268, September 1976. CO-DEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Bynum:1976:WPL

[1417] W. L. Bynum. Weak parallelogram laws for Banach spaces. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 19(3): 269–276, September 1976. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Fisher:1976:FPI

[1418] James L. Fisher. Finite principal ideal rings. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 19(3):277–284, September 1976. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Fremlin:1976:PRM

[1419] D. H. Fremlin. Products of Radon measures: a counterexample. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 19(3): 285–290, September 1976. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Kannappan:1976:CC

[1420] Pl. Kannappan and M. A. Taylor. On closure conditions. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 19(3): 291–296, September 1976. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Leblanc:1976:RCC

[1421] Raymond Leblanc. Representation of a certain class of polynomials. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 19(3):297–302, September 1976. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Rosenthal:1976:SRR

[1422] Peter Rosenthal. Some recent results on invariant subspaces. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 19(3): 303–314, September 1976. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Saxena:1976:AIH

[1423] R. B. Saxena. Averaging interpolation of Hermite–Fejér type. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 19(3): 315–322, September 1976. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Schumacher:1976:AFA

[1424] D. Schumacher. Absolutely free algebras in a topos containing an infinite object. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 19(3):323–328, September 1976. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Suvak:1976:FIR

[1425] John A. Suvak. Full ideals and ring groups in Z_{n}[x]. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 19(3):329–336, September 1976. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Tall:1976:SST

[1426] Franklin D. Tall. Stalking the Souslin tree — a topological guide. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 19(3): 337–342, September 1976. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Wyman:1976:NSR

[1427] Max Wyman. Non-static radially symmetric distributions of matter. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 19(3):343–358, September 1976. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Brondsted:1976:NBS

[1428] Arne Brøndsted. A note on best simultaneous approximation in normed linear spaces. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 19(3):359–360, September 1976. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Enochs:1976:NAP

[1429] Edgar Enochs. A note on absolutely pure modules. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 19(3):361–362, September 1976. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Fischer:1976:CTP

[1430] J. Fischer. Counterexamples to two problems on one-relator groups. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 19(3):363–364, September 1976. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Lohman:1976:NBS

[1431] Robert H. Lohman. A note on Banach spaces containing l_{1}. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 19(3): 365–368, September 1976. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Pandya:1976:CSG

[1432] G. N. Pandya and R. D. Bercov. 1-cohomology and splitting of group extensions. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 19(3):369–372, September 1976. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Schafer:1976:RHC

[1433] James A. Schafer. Representing homology classes on surfaces. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 19(3): 373–374, September 1976. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Shivakumar:1976:IDD

[1434] P. N. Shivakumar and Kim Ho Chew. Iterations for diagonally dominant matrices. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 19(3):375–384, September 1976. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Banaschewski:1976:TPB

[1435] Bernhard Banaschewski and Evelyn Nelson. Tensor products and biomorphisms. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 19(4):385–402, December 1976. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Binding:1976:SGC

[1436] Paul Binding and Patrick J. Browne. L^{p} spaces generated by certain operator valued measures. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 19(4):403–416, December 1976. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Borwein:1976:ZPS

[1437] D. Borwein and W. Kratz. On the zeros of the power series ∑_{n=0}^{∞}(−1)^{n}(1 − c^{−n−1})^{k_{z}n} with an application to discontinuous Rieszsummability. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 19(4):417–424, December 1976. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Carlitz:1976:SAP

[1438] L. Carlitz. Some arithmetic properties of a special sequence of integers. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 19(4):425–430, December 1976. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Chvatal:1976:TFG

[1439] Václav Chvátal, Anton Kotzig, Ivo G. Rosenberg, and Roy O. Davies. There are 2^{ℵ_{α}} friendship graphs of cardinal ℵ_{α}. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 19(4):431–434, December 1976. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Djokovic:1976:CTA

[1440] D. Z. Djoković. A closure theorem for analytic subgroups of real Lie groups. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 19(4):435–440, December 1976. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Gauthier:1976:SJP

[1441] P. M. Gauthier and J. S. Hwang. The set of Julia points for functions omitting two values. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 19(4):441–444, December 1976. CODEN CM-BUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Heinig:1976:WMI

[1442] H. P. Heinig. Weighted maximal inequalities for ↕^{r}-valued functions. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 19(4):445–454, December 1976. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Klincsek:1976:AF

[1443] G. T. Klincsek. About the # function. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 19(4):455–460, December 1976. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Rao:1976:GCP

[1444] K. Nageswara Rao. A generalization of the cyclotomic polynomial. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 19(4):461–466, December 1976. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Reiter:1976:CSC

[1445] Harold Reiter. Convex sets, Cantor sets and a midpoint property. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 19(4):467–472, December 1976. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Stolarsky:1976:BSC

[1446] Kenneth B. Stolarsky. Beatty sequences, continued fractions, and certain shift operators. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 19(4):473–482, December 1976. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Strube:1976:MPD

[1447] Richard F. E. Strube. A map of a polyhedron onto a disk. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 19(4): 483–486, December 1976. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Szafron:1976:ISP

[1448] D. A. Szafron and J. H. Weston. An internal solution to the problem of linearization of a convexity space. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 19(4):487–494, December 1976. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Bellenot:1976:ILB

[1449] Steven F. Bellenot. Inductive limits of Banach spaces. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 19(4):495– 496, December 1976. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Deo:1976:NER

[1450] Chandrakant M. Deo. A note on Erdős-renyi law of large numbers. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 19(4):497–500, December 1976. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Hicks:1976:NLQ

[1451] Troy L. Hicks and Shirley M. Huffman. A note on locally quasi-uniform spaces. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 19(4):501–504, December 1976. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Pham-Gia:1976:TL

[1452] Thu Pham-Gia. On a theorem of Lelong. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 19(4):505–506, December 1976. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Pathak:1977:ATH

[1453] R. S. Pathak and J. N. Pandey. Abelian theorems for Hardy transformations. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 20(??):331–336, ???? 1977. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Allegretto:1977:KTH

[1454] W. Allegretto. A Kneser theorem for higher order elliptic equations. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 20(1):1–8, March 1977. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Andersen:1977:IWD

[1455] Kenneth F. Andersen. Inequalities with weights for discrete Hilbert transforms. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 20(1):9–16, March 1977. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Banaschewski:1977:NEI

[1456] Bernhard Banaschewski and Evelyn Nelson. On the non-existence of injective near-ring modules. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 20(1):17–24, March 1977. CODEN CM-BUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Bell:1977:CTN

[1457] Howard E. Bell. A commutativity theorem for near-rings. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 20(1):25–28, March 1977. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Blundon:1977:TFN

[1458] W. J. Blundon. A three-fold non-lattice covering. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 20(1):29–32, March 1977. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Boler:1977:NHF

[1459] James Boler. A note on the Hopficity of finitely generated free-by-nilpotent groups. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 20(1):33–34, March 1977. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Borwein:1977:PEH

[1460] J. Borwein. A proof of the equivalence of Helly’s and Krasnoselski’s theorems. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 20(1):35–37, March 1977. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Britten:1977:PGL

[1461] Daniel J. Britten. On prime Goldie-like quadratic Jordan matrix algebras. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 20(1):39–46, March 1977. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Choo:1977:APC

[1462] Eng Ung Choo. Accumulation points of continuous real-valued functions and compactifications. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 20(1):47–52, March 1977. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Cote:1977:ATI

[1463] C. Côté and R. Gervais. Une application de la théorie de l’interpolation — (0, 2). (French) [An application of interpolation theory — (0, 2)]. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 20 (1):53–66, March 1977. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Cummings:1977:APC

[1464] Larry J. Cummings and Jennifer Seberry Wallis. An algorithm for the permanent of circulant matrices. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 20(1):67–70, March 1977. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Davison:1977:SMA

[1465] J. L. Davison. Some mappings associated with the permutation groups. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 20(1):71–76, March 1977. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

DeKoninck:1977:RAI

[1466] Jean-Marie De Koninck and Armel Mercier. Remarque sur un article “Identities for series of the type Σf(n)μ(n)n^{−s}”. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 20(1):77–88, March 1977. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

delaTorre:1977:DET

[1467] A. de la Torre. A dominated ergodic theorem for contractions with fixed points. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 20(1):89–92, March 1977. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Fine:1977:NTS

[1468] Benjamin Fine. A note on the two-square theorem. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 20(1):93–94, March 1977. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Freedman:1977:BTH

[1469] H. I. Freedman. On a bifurcation theorem of Hopf and Friedrichs. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 20 (1):95–102, March 1977. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Grimson:1977:MSM

[1470] R. C. Grimson. Multiple series manipulations and generating functions. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 20(1):103–106, March 1977. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Hardie:1977:NCT

[1471] K. A. Hardie. A note on the category of the telescope. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 20(1):107– 108, March 1977. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Leeming:1977:AEB

[1472] David J. Leeming. An asymptotic estimate for the Bernoulli and Euler numbers. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 20(1):109–112, March 1977. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Lim:1977:CEN

[1473] Taw Pin Lim. Conjugacy of elements in a normal ring. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 20(1):113– 116, March 1977. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Murty:1977:INA

[1474] M. Ram Murty and V. Kumar Murty. Irrational numbers arising from certain differential equations. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 20(1):117–120, March 1977. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Sizer:1977:TSG

[1475] Walter S. Sizer. Triangularizing solvable groups of unipotent matrices over a skew field. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 20(1):121–124, March 1977. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Struble:1977:NVF

[1476] Raimond A. Struble. Numerical-valued Fourier transforms. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 20(1): 125–128, March 1977. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Tzimbalario:1977:AGP

[1477] J. Tzimbalario. Approximation by generalised polynomials with integral coefficients. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 20(1):129–132, March 1977. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Williams:1977:NAC

[1478] H. C. Williams. On numbers analogous to the Carmichael numbers. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 20(1):133–144, March 1977. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Williams:1977:NBR

[1479] Kenneth S. Williams. Note on Burde’s rational biquadratic reciprocity law. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 20(1):145–152, March 1977. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Belley:1977:IOC

[1480] J.-M. Belley. Invertible operators on certain Banach spaces. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 20(2):

153–160, June 1977. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Borwein:1977:TTS

[1481] D. Borwein and E. Smet. Tauberian theorems for strong and absolute Borel-type methods of summability. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 20(2):161–172, June 1977. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Brown:1977:DPD

[1482] Julia M. Nowlin Brown. Direct product decompositions of elation groups. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 20(2):173–182, June 1977. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Carlitz:1977:CPS

[1483] L. Carlitz. Combinatorial property of a special polynomial sequence. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 20(2):183–188, June 1977. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Gunther:1977:SIO

[1484] Georg Gunther. Singular isometries in orthogonal groups. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 20(2):189– 198, June 1977. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Handelman:1977:CRR

[1485] David Handelman. Completions of rank rings. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 20(2):199–206, June 1977. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Heagy:1977:IAT

[1486] E. C. Heagy and B. L. R. Shawyer. On integral Abel-type and logarithmic methods of summability. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 20(2):207–214, June 1977. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Lesniak-Foster:1977:SDC

[1487] Linda M. Lesniak-Foster and James E. Williamson. On spanning and dominating circuits in graphs. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 20(2):215–220, June 1977. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Pao:1977:BVP

[1488] C. V. Pao. Boundary-value problems of a degenerate Sobolev-type differential equation. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 20(2):221–228, June 1977. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Ribenboim:1977:RRF

[1489] P. Ribenboim. Recent results on Fermat’s Last Theorem. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 20(2):229– 242, June 1977. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Siddiqi:1977:NCF

[1490] Rafat N. Siddiqi. A note on convergence of Fourier series of a function on Wiener’s class V_{p}. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 20(2):243–248, June 1977. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Spencer:1977:ITS

[1491] Joel Spencer. Intersection theorems for systems of sets. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 20(2):249– 254, June 1977. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Rangaswamy:1977:MFS

[1492] K. M. Rangaswamy. Modules with finite spanning dimension. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 20(2): 255–262, June 1977. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Johnson:1977:CNS

[1493] C. S. Johnson, Jr. and F. R. McMorris. Commutative non-singular semigroups. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 20(2):263–266, June 1977. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Oswald:1977:NIM

[1494] A. Oswald. A note on injective modules over a d.g. near-ring. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 20(2):

267–270, June 1977. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Popoola:1977:DCM

[1495] J. O. Popoola and I. Tweddle. On the dimension of a complete metrizable topological vector space. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 20(2):271–276, June 1977. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Duncan:1977:SPC

[1496] Richard Duncan. Some pointwise convergence results in L^{p}(μ), 1 < p < ∞. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 20(3):277–284, September 1977. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Akcoglu:1977:DPC

[1497] M. A. Akcoglu and L. Sucheston. Dilations of positive contractions on L_{p} spaces. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 20(3):285–292, September 1977. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Bick:1977:DFB

[1498] Abraham Bick. On the duals of flat Banach spaces. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 20(3):293–300, September 1977. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Hanumanthachari:1977:AC

[1499] J. Hanumanthachari. On an arithmetic convolution. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 20(3):301–306, September 1977. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Imoru:1977:SII

[1500] Christopher Olutunde Imoru. On some integral inequalites related to Hardy’s. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 20(3):307–312, September 1977. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Janos:1977:DTB

[1501] Ludvik Janos. Dimension theory via bisector chains. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 20(3):313–318, September 1977. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Nowakowski:1977:SFL

[1502] Richard Nowakowski and Ivan Rival. The spectrum of a finite lattice: breadth and length techniques. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 20(3):319–330, September 1977. CO-DEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Piccinini:1977:SRT

[1503] Renzo A. Piccinini. Some results in the theory of fibrations. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 20(3): 337–346, September 1977. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Raphael:1977:OCR

[1504] R. Raphael and W. Stephenson. Orthogonally complete rings. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 20(3): 347–352, September 1977. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Thomson:1977:CLI

[1505] B. S. Thomson. A characterization of the Lebesgue integral. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 20(3): 353–358, September 1977. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Vagholkar:1977:IGP

[1506] M. K. Vagholkar and P. J. Staff. Infinite geometric products. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 20(3): 359–368, September 1977. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Verma:1977:CSF

[1507] Arun Verma. Certain summation formulae for basic hypergeometric series. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 20(3):369–376, September 1977. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Wang:1977:VHI

[1508] Chung Lie Wang. Variants of the Hölder inequality and its inverses. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 20(3):377–384, September 1977. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Duret:1977:ICF

[1509] Jean-Louis Duret. Instabilité des corps formellement réels. (French) [Instability of formally-real objects]. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 20(3):385–388, September 1977. CO-DEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Lawson:1977:ADT

[1510] Terry Lawson. Applications of decomposition theorems to trivializing h-cobordisms. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 20(3):389–392, September 1977. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Mura:1977:EOG

[1511] R. Botto Mura and A. H. Rhemtulla. Extensions of orderable groups. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 20(3):393–396, September 1977. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Subbarao:1977:OEF

[1512] M. V. Subbarao and D. Suryanarayana. On the order of the error function of the (k, r)-integers. II. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 20(3):397–400, September 1977. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Westwick:1977:TSS

[1513] R. Westwick. Trivectors in a space of seven dimensions. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 20(3):401–404, September 1977. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Duff:1977:HDE

[1514] G. F. D. Duff. Hyperbolic differential equations and wave propagation. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 20(4):405–428, December 1977. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Ihrig:1977:LAU

[1515] Edwin Ihrig. Lie algebra as a unifying concept in modern physics. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 20(4):429–442, December 1977. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Al-Hussaini:1977:PPL

[1516] A. N. Al-Hussaini. Probabilistic proofs of the Little Riesz Theorem. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 20(4):443–446, December 1977. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Bryan:1977:AIP

[1517] Robert Neff Bryan. Adjoint interior-point boundary conditions for linear differential operators. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 20(4):447–450, December 1977. CODEN CM-BUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Dunham:1977:DBR

[1518] Charles B. Dunham. Dependence of best rational Chebyshev approximations on the domain. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 20(4):451–454, December 1977. CODEN CM-BUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Hicks:1977:NML

[1519] Troy L. Hicks and John D. Kubicek. Nonexpansive mappings in locally convex spaces. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 20(4):455–462, December 1977. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Kirk:1977:NLM

[1520] W. A. Kirk and William O. Ray. A note on Lipschitzian mappings in convex metric spaces. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 20(4):463–466, December 1977. CODEN CM-BUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Pham-Gia:1977:DNQ

[1521] Thu Pham-Gia. Density of non quasi-analytic classes of functions. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 20(4):467–470, December 1977. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Rowe:1977:IT

[1522] K. A. Rowe. Images in topoi. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 20(4):471–478, December 1977. CO-DEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Shah:1977:FET

[1523] Shantilal N. Shah. Functions of exponential type are differences of functions of bounded index. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 20(4):479–484, December 1977. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Stitzinger:1977:CSS

[1524] Ernest L. Stitzinger. On the commutativity of semi-simple associative algebras. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 20(4):485–488, December 1977. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Suryanarayana:1977:SGI

[1525] D. Suryanarayana and David T. Walker. Some generalizations of an identity of Subhankulov. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 20(4):489–494, December 1977. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Szeto:1977:SRB

[1526] George Szeto. On sheaf representation of a biregular near-ring. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 20(4): 495–500, December 1977. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Toth:1977:TPC

[1527] L. Fejes Tóth and N. Sauer. Thinnest packing of cubes with a given number of neighbours. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 20(4):501–508, December 1977. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Chein:1977:DCE

[1528] E. Z. Chein. On the derived cuboid of an Eulerian triple. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 20(4): 509–510, December 1977. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Gardner:1977:NS

[1529] B. Gardner. A note on (r, λ)-systems. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 20(4):511–512, December 1977. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Ginsburg:1977:MSW

[1530] John Ginsburg. The metrizability of spaces whose diagonals have a countable base. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 20(4):513–514, December 1977. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Parker:1977:LCD

[1531] Willard A. Parker. Local central L(p) dual objects. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 20(4):515, December 1977. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Kabe:1978:IGF

[1532] D. G. Kabe. An inequality for gamma functions. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 21(??):243–246, ???? 1978. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Broverman:1978:HBL

[1533] S. Broverman. Homomorphisms between lattices of zero-sets. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 21(1):1– 6, March 1978. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Clarke:1978:PCC

[1534] Frank H. Clarke. Pointwise contraction criteria for the existence of fixed points. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 21(1):7–12, March 1978. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Chung:1978:SAR

[1535] L. O. Chung and Jiang Luh. On semicommuting automorphisms of rings. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 21(1):13–16, March 1978. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Djokovic:1978:MTP

[1536] Dragomir Z. Djoković. Multiplicities in the tensor product of finitedimensional representations of discrete groups. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 21(1):17–20, March 1978. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Grosswald:1978:RRF

[1537] Emil Grosswald. Reducible rational fractions of the type of Gaussian polynomials with only non-negative coefficients. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 21(1):21–30, March 1978. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic). See errata [1619].

Hoffman:1978:TLS

[1538] P. Hoffman and A. Zabrodsky. Thin lens spaces. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 21(1):31–36, March 1978. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Jura:1978:CEF

[1539] Andrzej Jura. Coset enumeration in a finitely presented semigroup. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 21 (1):37–46, March 1978. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Lin:1978:IFS

[1540] C. S. Lin and M. Radjabalipour. On intertwining and factorization by self-adjoint operators. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 21(1):47–52, March 1978. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Lindgren:1978:CPC

[1541] William F. Lindgren and Peter Fletcher. A construction of the pair completion of a quasi-uniform space. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 21(1):53–60, March 1978. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Lund:1978:IFA

[1542] Bruce Lund. Isomorphisms of function algebras and algebras of analytic functions. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 21(1):61–72, March 1978. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Majindar:1978:CSB

[1543] K. N. Majindar. Coexistence of some b.i.b. designs. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 21(1):73–78, March 1978. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Shawyer:1978:MMT

[1544] B. L. R. Shawyer. Maximum modulus theorems and Schwarz lemmata for sequence spaces. II. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 21(1):79–84, March 1978. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Steele:1978:CFS

[1545] J. Michael Steele. Covering finite sets by ergodic images. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 21(1):85–92, March 1978. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Tan:1978:UC

[1546] Kok Keong Tan. Uniform contractification. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 21(1):93–98, March 1978. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Wayman:1978:EPF

[1547] Arthur K. Wayman. An elementary proof of a fixed point theorem of J. Lewittes and D. L. McQuillan. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 21(1):99–102, March 1978. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

White:1978:FCS

[1548] H. E. White, Jr. First countable spaces that have special pseudo-bases. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 21(1):103–112, March 1978. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Berman:1978:SRN

[1549] A. Berman. The spectral radius of a non-negative matrix. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 21(1): 113–114, March 1978. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Berman:1978:CGA

[1550] David M. Berman. Critical graphs for acyclic colorings. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 21(1):115– 116, March 1978. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Borwein:1978:SRL

[1551] D. Borwein. A sum of reciprocals of least common multiples. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 21(1): 117–118, March 1978. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Burgess:1978:CBG

[1552] W. D. Burgess. A characterization of biregular group rings. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 21(1): 119–120, March 1978. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Katchalski:1978:HTT

[1553] Meir Katchalski. A helly-type theorem for convex sets. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 21(1):121– 124, March 1978. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Servedio:1978:FOL

[1554] Frank Servedio. Forms with 0-orthogonal Lie algebras. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 21(1):125– 128, March 1978. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Boyd:1978:VTK

[1555] David W. Boyd. Variations on a theme of Kronecker. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 21(2):129– 134, June 1978. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Baker:1978:MAH

[1556] Catherine Baker. Moulton affine Hjelmslev planes. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 21(2):135–142, June 1978. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Cho:1978:SEV

[1557] S. J. Cho. Strong extensions vs. weak extensions of C^{∗}-algebras. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 21(2): 143–148, June 1978. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Cross:1978:RSS

[1558] G. E. Cross. The representation of (C, k) summable series in Fourier form. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 21(2):149–158, June 1978. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Dobbs:1978:WGD

[1559] David E. Dobbs. On the weak global dimension of pseudo-valuation domains. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 21(2):159–164, June 1978. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Hegde:1978:SSP

[1560] Shankar Hegde. Some smoothness properties of measures on topological spaces. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 21(2):165–172, June 1978. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Keagy:1978:LPS

[1561] Thomas A. Keagy. Limit preserving summability of subsequences. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 21 (2):173–176, June 1978. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Kleinschmidt:1978:RPR

[1562] Peter Kleinschmidt. Regular partitions of regular graphs. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 21(2):177– 182, June 1978. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Lin:1978:ACM

[1563] Shwu Yeng T. Lin and You Feng Lin. On almost continuous mappings and Baire spaces. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 21(2):183–186, June 1978. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Megan:1978:IOS

[1564] Mihail Megan. On the input-output stability of linear controllable systems. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 21(2):187–196, June 1978. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Meir:1978:IRA

[1565] A. Meir. An interpolatory rational approximation. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 21(2):197–200, June 1978. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Penna:1978:TS

[1566] Michael A. Penna. On a theorem of Sullivan. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 21(2):201–206, June 1978. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Schirmer:1978:FPC

[1567] Helga Schirmer. Fixed points can characterise curves. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 21(2):207– 212, June 1978. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Schoneberg:1978:ZNM

[1568] R. Schöneberg. Zeros of nonlinear monotone operators in Hilbert space. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 21(2):213–220, June 1978. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Snipes:1978:BTV

[1569] Ray F. Snipes. S-barrelled topological vector spaces. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 21(2):221– 228, June 1978. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Thomas:1978:ODA

[1570] L. A. Thomas. Ordered Desarguesian affine Hjelmslev planes. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 21(2): 229–236, June 1978. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Wilker:1978:RD

[1571] J. B. Wilker. Russian dolls. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 21(2):237–240, June 1978. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Felzenszwalb:1978:RL

[1572] B. Felzenszwalb. On a result of Levitzki. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 21(2):241–242, June 1978. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Kopp:1978:AVR

[1573] P. E. Kopp. An abstract version of a result of Fong and Sucheston. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 21(2):247–248, June 1978. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Mochizuki:1978:UMG

[1574] H. Y. Mochizuki. Unipotent matrix groups over division rings. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 21(2): 249–250, June 1978. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Smith:1978:RBS

[1575] Mark A. Smith. A reflexive Banach space that is LUR and not 2R. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 21(2):251–252, June 1978. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Zorzitto:1978:HFC

[1576] Frank Zorzitto. How fields can have a product. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 21(2):253–256, June 1978. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Borwein:1978:HMP

[1577] David Borwein. The Hausdorff moment problem. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 21(3):257–266, September 1978. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Al-Abedeen:1978:GEU

[1578] Ahmed Z. Al-Abedeen and H. L. Arora. A global existence and uniqueness theorem for ordinary differential equations of generalized order. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 21(3):267–272, September 1978. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Al-Amiri:1978:COD

[1579] H. S. Al-Amiri. Certain nth order differential inequalities in the complex plane. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 21(3):273–278, September 1978. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Beattie:1978:CVS

[1580] Ronald Beattie. On the convergence vector space L_{c}(E, , F) and its dual space. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 21(3):279–284, September 1978. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Chein:1978:CGC

[1581] E. Z. Chein. On a conjecture of Graham concerning a sequence of integers. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 21(3):285–288, September 1978. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Duggal:1978:ACM

[1582] K. L. Duggal. On almost contingent manifolds of second class with applications in relativity. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 21(3):289–296, September 1978. CODEN CM-BUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Eswarathasan:1978:SPL

[1583] A. Eswarathasan. On square pseudo-Lucas numbers. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 21(3):297–304, September 1978. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Janos:1978:DTR

[1584] Ludvik Janos. Dimension theory via reduced bisector chains. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 21(3): 305–312, September 1978. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Novoa:1978:OCC

[1585] Lino Gutierrez Novoa. Order characterization of the complex field. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 21(3):313–318, September 1978. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Page:1978:CRR

[1586] S. S. Page. Continuous rings and rings of quotients. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 21(3):319–324, September 1978. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Phillips:1978:AFF

[1587] John Phillips. Automorphisms of full II_{1} factors. II. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 21(3):325–328, September 1978. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Rush:1978:CA

[1588] David E. Rush. Content algebras. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 21(3):329–334, September 1978. CO-DEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Slater:1978:CSH

[1589] Peter J. Slater. A characterization of soft hypergraphs. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 21(3):335–338, September 1978. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Thomson:1978:WVC

[1590] B. S. Thomson. On weak Vitali covering properties. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 21(3):339–346, September 1978. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Wang:1978:AVH

[1591] Chung Lie Wang. Applications of variants of the Hölder inequality and its inverses: extensions of Barnes, Marshall–Olkin, and Nehari inequalities. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 21(3):347–354, September 1978. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Zarrow:1978:ERT

[1592] Robert Zarrow. On Earle’s m mod n relative Teichmüller spaces. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 21(3): 355–360, September 1978. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Anderson:1978:CRT

[1593] David F. Anderson. The Chinese Remainder Theorem and the invariant basis property. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 21(3):361–362, September 1978. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Fleischer:1978:POS

[1594] Isidore Fleischer. Paracompactness for ordered sums. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 21(3):363–364, September 1978. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Heath:1978:NAD

[1595] Philip R. Heath and Klaus Heiner Kamps. Note on attaching Dold fibrations. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 21(3):365–368, September 1978. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Herstein:1978:ND

[1596] I. N. Herstein. A note on derivations. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 21(3):369–370, September 1978. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

MacDougall:1978:NLT

[1597] J. A. MacDougall. A note on the length of trivectors. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 21(3):371–372, September 1978. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Papick:1978:RCO

[1598] Ira J. Papick. A remark on coherent overrings. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 21(3):373–376, September 1978. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Sridharan:1978:NAF

[1599] M. R. Sridharan. Note on an asymptotic formula for a class of digraphs. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 21(3):377–384, September 1978. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Coxeter:1978:PL

[1600] H. S. M. Coxeter. Parallel lines. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 21(4):385–398, December 1978. CO-DEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Bell:1978:PMR

[1601] Howard E. Bell. On the power map and ring commutativity. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 21(4): 399–404, December 1978. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Bradley:1978:HIM

[1602] J. Scott Bradley. Hardy inequalities with mixed norms. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 21(4):405– 408, December 1978. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

DeKoninck:1978:AFR

[1603] Jean-Marie De Koninck and Aleksandar Ivić. An asymptotic formula for reciprocals of logarithms of certain multiplicative functions. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 21(4): 409–414, December 1978. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Elliott:1978:AFD

[1604] George A. Elliott. On approximately finite-dimensional von Neumann algebras. II. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 21(4):415–418, December 1978. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Ellis:1978:CMF

[1605] Martin H. Ellis. Catching and missing finite sets. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 21(4):419–426, December 1978. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Evans:1978:AFS

[1606] Ronald J. Evans. An asymptotic formula for a sum of products of powers. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 21(4):427–434, December 1978. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Fitzpatrick:1978:LFD

[1607] W. J. Fitzpatrick and L. J. Grimm. Linear functional-differential equations in a Banach algebra. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 21(4):435–440, December 1978. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Fraenkel:1978:DSD

[1608] A. S. Fraenkel, M. Mushkin, and U. Tassa. Determination of [nθ] by its sequence of differences. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 21(4):441–446, December 1978. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Mejzler:1978:LDE

[1609] D. Mejzler. Limit distributions for the extreme order statistics. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 21(4): 447–460, December 1978. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Cockayne:1978:PHH

[1610] E. J. Cockayne, S. T. Hedetniemi, and D. J. Miller. Properties of hereditary hypergraphs and middle graphs. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 21(4):461–468, December 1978. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Pawar:1978:DS

[1611] Y. S. Pawar and N. K. Thakare. 0-distributive semilattices. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 21(4): 469–476, December 1978. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Rose:1978:LDC

[1612] David Alon Rose. On Levine’s decomposition of continuity. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 21(4): 477–482, December 1978. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Servedio:1978:PIL

[1613] Frank J. Servedio. Principal irreducible Lie-algebra modules. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 21(4): 483–490, December 1978. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Williamson:1978:CCI

[1614] Alan Williamson. On the conjugacy classes in an integral group ring. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 21(4):491–496, December 1978. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Borwein:1978:TSS

[1615] D. Borwein and S. Z. Ditor. Translates of sequences in sets of positive measure. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 21(4):497–498, December 1978. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Day:1978:IGA

[1616] Alan Day. Idempotents in the groupoid of all fSP classes of lattices. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 21(4):499–502, December 1978. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Ilori:1978:ICS

[1617] Samuel A. Ilori. An inequality for complete symmetric functions. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 21(4): 503–504, December 1978. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Small:1978:SPA

[1618] Charles Small. Sums of powers in arithmetic progressions. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 21(4): 505–506, December 1978. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Grosswald:1978:ERR

[1619] Emil Grosswald. Errata: “Reducible rational fractions of the type of Gaussian polynomials with only non-negative coefficients” (Canad. Math. Bull. 21 (1978), no. 1, 21–30). Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 21(4):508, December 1978. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic). See [1537].

Bair:1979:SES

[1620] J. Bair. Sur la structure extrémale de la somme de deux convexes. (French) [On the extremal structure of the sum of two convexes]. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 22(1):1– 8, March 1979. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Chakerian:1979:GIP

[1621] G. D. Chakerian. Geometric inequalities for plane convex bodies. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 22(1):9– 16, March 1979. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Cheng:1979:NS

[1622] Sui Sun Cheng. Nonoscillatory solutions of x^{(m)} = (−1)^{m}Q(t)x. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 22(1): 17–22, March 1979. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Eplett:1979:SCT

[1623] W. J. R. Eplett. Self-converse tournaments. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 22(1):23–28, March 1979. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Eswarathasan:1979:PLN

[1624] A. Eswarathasan. On pseudo-Lucas numbers of the form 2S^{2}. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 22(1): 29–34, March 1979. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Farrell:1979:CPA

[1625] E. J. Farrell. On a class of polynomials associated with the stars of a graph and its application to node-disjoint decompositions of complete graphs and complete bipartite gra. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 22(1):35–46, March 1979. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Gawronski:1979:LBN

[1626] W. Gawronski. A lower bound for the number of negative zeros of power series. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 22(1):47–52, March 1979. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Khan:1979:MBH

[1627] M. Zubair Khan. Modules over bounded hereditary Noetherian prime rings. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 22(1):53–58, March 1979. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Mercier:1979:SFA

[1628] Armel Mercier. Sommes de fonctions additives restreintes à une class de congruence. (French) [Sums of additive functions constrained by a congruence class]. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 22(1):59–74, March 1979. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Moon:1979:TWS

[1629] J. W. Moon. Tournaments whose subtournaments are irreducible or transitive. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 22(1):75–80, March 1979. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Nishi:1979:EPJ

[1630] Akihiro Nishi. An elementary proof of Johnson–Dulmage–Mendelsohn’s refinement of Birkhoff’s theorem on doubly stochastic matrices. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 22(1): 81–86, March 1979. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Rabung:1979:SPF

[1631] John R. Rabung. Some progression-free partitions constructed using Folkman’s method. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 22(1):87–92, March 1979. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Singh:1979:MM

[1632] Surjeet Singh and Fazal Mehdi. Multiplication modules. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 22(1):93–98, March 1979. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Enos:1979:CFF

[1633] H. C. Enos. A counterexample in finite fixed point theory. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 22(1): 99–100, March 1979. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Fillmore:1979:NRO

[1634] P. A. Fillmore. A note on reductive operators. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 22(1):101–102, March 1979. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Johnson:1979:CLT

[1635] Dudley Paul Johnson. A central limit theorem for general stochastic processes. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 22(1):103–104, March 1979. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Pareek:1979:NCS

[1636] C. M. Pareek and H. Z. Hdeib. A note on [a, , b]-compact spaces. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 22(1): 105–112, March 1979. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Rao:1979:TN

[1637] R. Sitaramachandra Rao and G. Sri Rama Chandra Murty. On a theorem of Niven. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 22(1):113–116, March 1979. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Segal:1979:SNI

[1638] S. L. Segal. A second note on Ingham’s summation method. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 22(1): 117–128, March 1979. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Dawson:1979:SMD

[1639] Donald A. Dawson. Stochastic measure diffusion processes. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 22(2): 129–138, June 1979. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Swanson:1979:SSO

[1640] C. A. Swanson. Semilinear second-order elliptic oscillation. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 22(2): 139–158, June 1979. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Burgess:1979:RAW

[1641] W. D. Burgess and W. Stephenson. Rings all of whose Pierce stalks are local. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 22(2):159–164, June 1979. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Chan:1979:SEH

[1642] Ling Yau Chan. Some extensions of Hardy’s inequality. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 22(2):165– 170, June 1979. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Hartman:1979:BCC

[1643] Jehuda Hartman. Bounds on the coarseness of the n-cube. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 22(2): 171–176, June 1979. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Krishnaiah:1979:BMA

[1644] P. V. Krishnaiah and R. Sitaramachandra Rao. On Berndt’s method in arithmetical functions and contour integration. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 22(2):177–186, June 1979. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Nelson:1979:IHF

[1645] Evelyn Nelson. Internal hom functors for polarities. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 22(2):187–202, June 1979. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Pandey:1979:PDF

[1646] J. N. Pandey. On the positive definiteness of a functional. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 22(2): 203–206, June 1979. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Sivaramakrishnan:1979:SRR

[1647] R. Sivaramakrishnan. Square-reduced residue systems (modr) and related arithmetical functions. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 22(2):207–220, June 1979. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Stadtmuller:1979:ZPS

[1648] U. Stadtmüller. On the zeros of power series with logarithmic coefficients. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 22(2):221–234, June 1979. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Al-Salam:1979:FE

[1649] Nadhla A. Al-Salam and Waleed A. Al-Salam. On a functional equation. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 22(2):235–238, June 1979. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Lagrange:1979:DDC

[1650] Jean Lagrange. Sur le dérivé du cuboïde eulérien. (French) [On the derivative of the Eulerian cuboid]. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 22(2):239–242, June 1979. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Wong:1979:NIM

[1651] Pak Ken Wong. A note on isomorphisms of multiplier algebras. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 22(2): 243–256, June 1979. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Csorgo:1979:BMW

[1652] Miklós Csörgö. Brownian motion — Wiener process. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 22(3):257–280, September 1979. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Abramson:1979:PRS

[1653] Morton Abramson. Permutations related to secant, tangent and Eulerian numbers. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 22(3):281–292, September 1979. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Johnson:1979:HTD

[1654] Dudley Paul Johnson. Hitting time distributions when ∑ X_{k}/d^{k} has a smooth density. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 22(3):293–298, September 1979. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Krishna:1979:PIM

[1655] K. Krishna. A proof of an identity for multiplicative functions. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 22(3): 299–304, September 1979. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Bermond:1979:HCS

[1656] J.-C. Bermond, A. Germa, and M.-C. Heydemann. Hamiltonian cycles in strong products of graphs. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 22(3):305–310, September 1979. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Atalla:1979:GMP

[1657] Robert E. Atalla. Generalized Markov projections and matrix summability. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 22(3):311–316, September 1979. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Mercier:1979:IPF

[1658] Armel Mercier. Identité pour ∑∞_{{n=1 \atop n≡l(k)}} (f(n)/n^{s}). (French) [Identity for

∑∞_{{n=1 \atop n≡l(k)}} (f(n)/n^{s})]. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 22(3):317–326, September 1979. CODEN CM-BUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Cliff:1979:PEF

[1659] Gerald H. Cliff and Akbar H. Rhemtulla. Permuting the elements of a finite solvable group. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 22(3):327–330, September 1979. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Papick:1979:CO

[1660] Ira J. Papick. Coherent overrings. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 22(3):331–338, September 1979. CO-DEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Kovacs:1979:DPR

[1661] Amos Kovacs. On derivations in prime rings and a question of Herstein. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 22(3):339–344, September 1979. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Holland:1979:CTS

[1662] A. S. B. Holland, B. N. Sahney, and J. Tzimbalario. A criterion for Taylor summability of Fourier series. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 22(3):345–350, September 1979. CO-DEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Bisztriczky:1979:SOT

[1663] Tibor Bisztriczky. On surfaces of order three. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 22(3):351–356, September 1979. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Borosh:1979:BPI

[1664] I. Borosh and L. B. Treybig. Bounds on positive integral solutions of linear Diophantine equations. II. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 22(3):357–362, September 1979. CO-DEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Bennett:1979:NBA

[1665] Colin Bennett, Karl Rudnick, and Jeffrey D. Vaaler. Note on best approximation of x. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 22(3):363–366, September 1979. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Ramankutty:1979:FDA

[1666] P. Ramankutty. Functional dependence and analytic functions. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 22(3): 367–370, September 1979. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Mullen:1979:NBB

[1667] K. Mullen. A note on Bernstein’s bivariate inequality. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 22(3):371–376, September 1979. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Wang:1979:ABS

[1668] James Li Ming Wang. Approximation on boundary sets. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 22(3):377– 380, September 1979. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Pham-Gia:1979:ERM

[1669] T. Pham-Gia. Extension of a result of S. Mandelbrojt. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 22(3):381–390, September 1979. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

McLeish:1979:CLT

[1670] D. L. McLeish. Central limit theorem for absolute deviations from the sample mean and applications. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 22(4):391–396, December 1979. CO-DEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Beachy:1979:LI

[1671] John A. Beachy. On localization at an ideal. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 22(4):397–402, December 1979. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Sficas:1979:SMS

[1672] Y. G. Sficas. Strongly monotone solutions of retarded differential equations. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 22(4):403–412, December 1979. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Husain:1979:ITA

[1673] T. Husain. Infrasequential topological algebras. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 22(4):413–418, December 1979. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Nicholson:1979:CTR

[1674] W. K. Nicholson and Adil Yaqub. A commutativity theorem for rings and groups. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 22(4):419–424, December 1979. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Hinton:1979:MDS

[1675] Don B. Hinton. Molchanov’s discrete spectra criterion for a weighted operator. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 22(4):425–432, December 1979. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Ebanks:1979:BMI

[1676] Bruce R. Ebanks. Branching measures of information on strings. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 22(4): 433–448, December 1979. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Khan:1979:MBL

[1677] M. Zubair Khan. Modules behaving like torsion abelian groups. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 22(4): 449–458, December 1979. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Dor:1979:BCD

[1678] Leonard E. Dor. On basis constants and duality in Banach spaces. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 22(4):459–466, December 1979. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Clapp:1979:NSC

[1679] Michael H. Clapp and Ray C. Shiflett. A necessary and sufficient condition for the equivalence of the topologies of uniform and compact convergence. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 22(4):467–470, December 1979. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Putcha:1979:LUF

[1680] Mohan S. Putcha. Local unique factorization in the semigroup of paths in fR^{n}. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 22(4):471–476, December 1979. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Chang:1979:CDL

[1681] Gerard J. Chang. Complete diagonals of Latin squares. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 22(4):477– 482, December 1979. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Kardos:1979:E

[1682] Peter Kardos. On the equality ∑_{i=1}^{n} p_{i}f_{i}(p_{i})/f_{i}(q_{i}) ≤ 1. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 22(4):483– 490, December 1979. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Joseph:1979:RCM

[1683] James E. Joseph. On regular-closed and minimal regular spaces. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 22(4): 491–498, December 1979. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Kerman:1979:FSC

[1684] R. A. Kerman. Function spaces continuously paired by operators of convolution type. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 22(4):499–508, December 1979. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Herstein:1979:NDI

[1685] I. N. Herstein. A note on derivations. II. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 22(4):509–512, December 1979. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Lin:1979:NTK

[1686] Tzu Chu Lin. A note on a theorem of Ky Fan. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 22(4):513–516, December 1979. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Santhakumari:1979:QLN

[1687] C. Santhakumari. On quotient loops of normal subloops. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 22(4):517– 518, December 1979. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Willmott:1980:USN

[1688] Richard Willmott. Uniform spaces as nice images of nice uniform and metric spaces. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 23(1):1–10, March 1980. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Gregory:1980:USS

[1689] David A. Gregory. Upper semicontinuity of subdifferential mappings. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 23(1):11–20, March 1980. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Brualdi:1980:NDS

[1690] Richard A. Brualdi. A note on degree sequences of graphs. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 23(1): 21–28, March 1980. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Lessard:1980:DGD

[1691] Sabin Lessard. Une double généralisation du théorème de Fejér– Lebesgue. (French) [A double generalization of the Fejér–Lebesgue theorem]. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 23(1):29–36, March 1980. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Dobbs:1980:IEP

[1692] David E. Dobbs. On INC-extensions and polynomials with unit content. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 23(1):37–42, March 1980. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Scobey:1980:LME

[1693] P. Scobey and D. G. Kabe. On linear matrix equations. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 23(1):43–50, March 1980. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Allenby:1980:NFS

[1694] R. B. J. T. Allenby, C. Y. Tang, and S. Y. Tang. Notes on Frattini subgroups of generalized free products with cyclic amalgamation. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 23 (1):51–60, March 1980. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Roggenkamp:1980:ARS

[1695] K. W. Roggenkamp. Auslander–Reiten sequences for “nice” torsion theories of Artinian algebras. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 23(1):61–66, March 1980. CODEN CM-BUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Oltikar:1980:RCO

[1696] B. C. Oltikar. Right cyclically ordered groups. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 23(1):67–70, March 1980. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Lee:1980:SGI

[1697] Cheng Shyong Lee. On some generalization of inequalities of Opial, Yang and Shum. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 23(1):71–80, March 1980. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Aubin:1980:SOE

[1698] J.-P. Aubin and I. Ekeland. Second-order evolution equations associated with convex Hamiltonians. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 23(1):81–94, March 1980. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Adamson:1980:NRT

[1699] Alan Adamson. A note on the realization of types. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 23(1):95–98, March 1980. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Fong:1980:NSC

[1700] Paul Fong and Morton E. Harris. A note on strongly closed 2-subgroups. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 23(1):99–102, March 1980. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Rant:1980:MGM

[1701] W. H. Rant. Minimally generated modules. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 23(1):103–106, March 1980. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Cross:1980:SCF

[1702] G. E. Cross. Splicing n-convex functions using splines. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 23(1):107– 110, March 1980. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Macedonska-Nosalska:1980:NAF

[1703] Olga Macedońska-Nosalska. Note on automorphisms of a free abelian group. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 23(1):111–114, March 1980. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Olsen:1980:IET

[1704] James H. Olsen. The individual ergodic theorem for contractions with fixed points. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 23(1):115–126, March 1980. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Kreyszig:1980:FTI

[1705] Erwin Kreyszig. Function-theoretic integral operator methods for partial differential equations. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 23(2):127–136, June 1980. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Parameswaran:1980:CCS

[1706] Mangalam R. Parameswaran. On the connectedness of certain sets in summability theory. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 23(2):137–142, June 1980. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Bragard:1980:AOS

[1707] L. Bragard and J. Vangeldère. Adjonction ou suppression d’objectifs en programmation à objectifs linéaires multiples. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 23(2):143–154, June 1980. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Hill:1980:DTR

[1708] David A. Hill. Decomposition theorems for q^{∗}-rings. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 23(2):155– 160, June 1980. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Joseph:1980:MIC

[1709] James E. Joseph. Multifunctions and inverse cluster sets. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 23(2):161– 172, June 1980. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Page:1980:SR

[1710] S. S. Page. Stable rings. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 23(2):173–178, June 1980. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Inayatnoor:1980:GLM

[1711] K. Inayatnoor and M. Aslam Noor. A generalization of the Lax–Milgram lemma. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 23(2):179–184, June 1980. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Aliprantis:1980:SLS

[1712] C. D. Aliprantis and O. Burkinshaw. On the structure of locally solid topologies. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 23(2):185–192, June 1980. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Hu:1980:FPT

[1713] Thakyin Hu. Fixed point theorems for multivalued mappings. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 23(2): 193–198, June 1980. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Herzog:1980:CFC

[1714] Fritz Herzog. On the continued fractions of conjugate quadratic irrationalities. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 23(2):199–206, June 1980. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Agou:1980:IPC

[1715] S. Agou. Irréductibilité des polynômes f(∑_{i=0}^{m} a_{i}X^{p^{ri}}) sur un corps fini F_{p^{s}}. (French) [Irreducibility of polynomials f(∑_{i=0}^{m} a_{i}X^{p^{ri}}) by a finite object F_{p^{s}}]. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 23(2):207–212, June 1980. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Azzam:1980:SPB

[1716] A. Azzam. Smoothness properties of bounded solutions of Dirichlet’s problem for elliptic equations in regions with corners on the boundary. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 23(2):213–226, June 1980. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Connes:1980:CAF

[1717] Alain Connes and E. J. Woods. A construction of approximately finitedimensional non-ITPFI factors. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 23(2):227–230, June 1980. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Duffus:1980:RFP

[1718] Dwight Duffus, Werner Poguntke, and Ivan Rival. Retracts and the fixed point problem for finite partially ordered sets. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 23(2):231–236, June 1980. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Bourne:1980:IML

[1719] Samuel Bourne. On r^{∗}-invariant measure on a locally compact semigroup with recurrence. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 23(2):237–240, June 1980. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Richoux:1980:CTD

[1720] Anthony Richoux. A commutativity theorem for division rings. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 23(2): 241–244, June 1980. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Grunenfelder:1980:IGR

[1721] L. Grünenfelder and M. M. Parmenter. Isomorphic group rings with nonisomorphic coefficient rings. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 23(2):245–254, June 1980. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Roberts:1980:SES

[1722] Leslie G. Roberts. Some examples of smooth and regular rings. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 23(3): 255–260, September 1980. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Maher:1980:FWT

[1723] Laurence P. Maher, Jr. A function which transforms certain graphs into straight lines for simultaneous solution. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 23(3):261–266, September 1980. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Carroll:1980:STI

[1724] C. E. Carroll, C. C. Yang, and S. Ahn. Some triangle inequalities and generalizations. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 23(3):267–274, September 1980. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Huang:1980:DD

[1725] C. Huang and J. Schönheim. Decomposition of K_{n} into dragons. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 23(3): 275–280, September 1980. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Moser:1980:NIS

[1726] W. O. J. Moser and Richard Pollack. A new identity and some applications. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 23(3):281–290, September 1980. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Pawar:1980:PS

[1727] Y. S. Pawar and N. K. Thakare. On prime semilattices. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 23(3):291–298, September 1980. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Rich:1980:NIA

[1728] Michael Rich. Nilpotent ideals in alternative rings. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 23(3):299–304, September 1980. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Wong:1980:CLC

[1729] James C. S. Wong. A characterisation of locally compact amenable subsemigroups. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 23(3):305–312, September 1980. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Robertson:1980:CFN

[1730] Edmund F. Robertson. A comment on finite nilpotent groups of deficiency zero. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 23(3):313–316, September 1980. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Mathsen:1980:CNG

[1731] R. M. Mathsen. Coincidence of nodes for generalized convex functions. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 23(3):317–320, September 1980. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Rooney:1980:BRE

[1732] P. G. Rooney. On the boundedness and range of the extended Hankel transformation. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 23(3):321–326, September 1980. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Ceccherini:1980:NZF

[1733] P. V. Ceccherini and J. W. P. Hirschfeld. On the number of zeros over a finite field of certain symmetric polynomials. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 23(3):327–332, September 1980. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Johnson:1980:RPS

[1734] Dudley Paul Johnson. The ruin problem for sums of dependent random variables. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 23(3):333–338, September 1980. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Tijs:1980:SGC

[1735] S. H. Tijs and J. M. Borwein. Some generalizations of Carathéodory’s theorem via barycentres, with application to mathematical programming. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 23(3):339–346, September 1980. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Lohman:1980:GBS

[1736] Robert H. Lohman. Gaps between spheres in normed linear spaces. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 23(3):347–354, September 1980. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Rosen:1980:LPR

[1737] Kenneth H. Rosen. Least positive residues and the quadratic character of two. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 23(3):355–358, September 1980. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Cohen:1980:NJG

[1738] M. E. Cohen and H. S. Sun. A note on the Jensen–Gould convolutions. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 23(3):359–362, September 1980. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Stitzinger:1980:CMI

[1739] Ernest L. Stitzinger. Complements of minimal ideals in solvable Lie rings. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 23(3):363–366, September 1980. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Weintraub:1980:NDM

[1740] Steven H. Weintraub. A note on doubles of 4-manifolds. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 23(3):367– 370, September 1980. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Subbarao:1980:NCE

[1741] M. V. Subbarao. On the nonexistence of certain Euler products. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 23(3): 371–372, September 1980. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Ward:1980:NWM

[1742] L. E. Ward, Jr. A note on Whitney maps. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 23(3):373–374, September 1980. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Hausman:1980:ULP

[1743] Miriam Hausman. An upper limit property of the Euler function. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 23(3): 375–382, September 1980. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Greiner:1980:SHH

[1744] Peter C. Greiner. Spherical harmonics on the Heisenberg group. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 23(4): 383–396, December 1980. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Bell:1980:FCL

[1745] Murray Bell and John Ginsburg. First countable Lindelöf extensions of uncountable discrete spaces. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 23(4):397–400, December 1980. CO-DEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Fakhoury:1980:ICE

[1746] H. Fakhoury. Sur les M-idéaux dans certains espaces d’opérateurs et l’approximation par des opérateurs compacts. (French) [On the M-ideals in certain operator spaces and their approximation by compact operators]. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 23(4):401–412, December 1980. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Crombez:1980:TCC

[1747] G. Crombez and W. Govaerts. Towards a classification of convolutiontype operators from l_{1} to l_{∞}. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 23(4):413–420, December 1980. CO-DEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Macphail:1980:SP

[1748] M. S. Macphail. A summability problem. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 23(4):421–424, December 1980. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Nanda:1980:NUC

[1749] Sribatsa Nanda. A note on the universe of a category of fractions. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 23(4): 425–428, December 1980. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Abbott:1980:PFS

[1750] H. L. Abbott and A. C. Liu. On property mB of families of sets. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 23(4): 429–436, December 1980. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Mason:1980:PIR

[1751] Gordon Mason. Prime z-ideals of C(X) and related rings. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 23(4): 437–444, December 1980. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Groenewald:1980:UGR

[1752] N. J. Groenewald. Units of the group ring. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 23(4):445–448, December 1980. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Moore:1980:GLR

[1753] H. G. Moore. Generalized n-like rings and commutativity. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 23(4): 449–452, December 1980. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Martin:1980:NFP

[1754] John R. Martin and Sam B. Nadler, Jr. A note on fixed point sets and wedges. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 23(4):453–456, December 1980. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Anderson:1980:NRW

[1755] D. D. Anderson. Noetherian rings in which every ideal is a product of primary ideals. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 23(4):457–460, December 1980. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Liang:1980:IIC

[1756] Chao Chu Liang. Infinitesimal isometries on compact manifolds. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 23(4): 461–464, December 1980. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Martin:1980:DSM

[1757] Harold W. Martin. Derived subspaces of metric spaces. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 23(4):465– 468, December 1980. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Boyarsky:1980:TDO

[1758] Abraham Boyarsky and Manny Scarowsky. A theorem on the denseness of orbits in metric spaces. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 23(4):469–472, December 1980. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Felzenszwalb:1980:PNP

[1759] Bernardo Felzenszwalb and Antonino Giambruno. Periodic and nil polynomials in rings. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 23(4):473–476, December 1980. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Rai:1980:OCR

[1760] R. K. Rai. Orthogonal completions of reduced rings with respect to abian order. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 23(4):477–490, December 1980. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Nathanson:1980:APC

[1761] Melvyn B. Nathanson. Arithmetic progressions contained in sequences with bounded gaps. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 23(4):491–494, December 1980. CODEN CM-BUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Klein:1980:SPT

[1762] Abraham A. Klein. A simple proof of a theorem on reduced rings. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 23(4):495–496, December 1980. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Huffman:1980:CQU

[1763] Shirley M. Huffman, Troy L. Hicks, and John W. Carlson. Complete quasi-uniform spaces. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 23(4):497–498, December 1980. CODEN CM-BUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Yamashita:1980:NFW

[1764] Shinji Yamashita. A nonnormal function whose derivative is of Hardy class H^{p}, 0 < p < 1. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 23(4):499–500, December 1980. CODEN CM-BUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Dierolf:1980:SSD

[1765] Peter Dierolf and Susanne Dierolf. Some spaces are not the domain of a closed linear operator in a Banach space. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 23(4):501–503, December 1980. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Shinbrot:1981:PWR

[1766] Marvin Shinbrot. Periodic waves in a running stream. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 24(1):3–12, March 1981. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Carlson:1981:UCN

[1767] John W. Carlson. An ultrafilter completion of a nearness space. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 24(1): 13–22, March 1981. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Hartzman:1981:NPF

[1768] C. S. Hartzman. Nonsingular periodic flows on T^{3} and periodic homeomorphisms of T^{2}. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 24(1):23–28, March 1981. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Wolff:1981:RIF

[1769] Harvey Wolff. Relative injectives and free monads. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 24(1):29–36, March 1981. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Smith:1981:AND

[1770] R. A. Smith and M. V. Subbarao. The average number of divisors in an arithmetic progression. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 24(1):37–42, March 1981. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Belair:1981:EDF

[1771] Jacques Bélair. Sur une équation différentielle fonctionnelle analytique. (French) [On a functional analytic differential equation]. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 24(1):43–46, March 1981. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Ho:1981:T

[1772] Chung Wu Ho. On the topologies of R^{∞}. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 24(1):47–52, March 1981. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Weinstein:1981:ZFR

[1773] Lenard Weinstein. The zeroes of functions related to Dirichlet Lfunctions. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 24(1):53–58, March 1981. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Anantharaman:1981:SSS

[1774] R. Anantharaman, T. Lewis, and J. H. M. Whitfield. Smoothability, strong smoothability and dentability in Banach spaces. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 24(1):59–68, March 1981. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Poguntke:1981:FGS

[1775] Werner Poguntke. On finitely generated simple complemented lattices. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 24(1):69–72, March 1981. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Subbarao:1981:TFC

[1776] M. V. Subbarao and V. C. Harris. A transformation formula for a class of arithmetic sums. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 24(1):73–78, March 1981. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Junghenn:1981:TLA

[1777] H. D. Junghenn. Topological left amenability of semidirect products. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 24(1):79–86, March 1981. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Tsui:1981:CDN

[1778] Sze Kai Tsui. Compact derivations on von Neumann algebras. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 24(1): 87–90, March 1981. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Donnelly:1981:BNA

[1779] Harold Donnelly. On L^{2}–Betti numbers for abelian groups. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 24(1): 91–96, March 1981. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

White:1981:HDV

[1780] Errin Erb White. On the homological dimension of valuated vector spaces. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 24(1):97–100, March 1981. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Lyndon:1981:CFG

[1781] R. C. Lyndon and M. J. Wicks. Commutators in free groups. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 24(1): 101–106, March 1981. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Mahamoodian:1981:ECC

[1782] E. S. Mahamoodian. On edge-colorability of Cartesian products of graphs. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 24(1):107–108, March 1981. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Grover:1981:GFT

[1783] A. K. Grover and J. H. V. Hunt. A generalization of Floyd’s theorem on unicoherent Peano continua with involution. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 24(1):109–112, March 1981. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Goebel:1981:RNM

[1784] Kazimierz Goebel and Małgorzata Koter. A remark on nonexpansive mappings. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 24(1):113–116, March 1981. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Pettet:1981:NSE

[1785] Martin R. Pettet. A note on split extensions of finite groups. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 24(1): 117–120, March 1981. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Humphries:1981:PLP

[1786] Michael D. Humphries. The Phragmén–Lindelöf principle. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 24(1):121– 122, March 1981. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Boyarsky:1981:BNI

[1787] Abraham Boyarsky. A bound on the number of invariant measures. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 24 (1):123–128, March 1981. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Gupta:1981:PRH

[1788] Narain Gupta. A problem of R. H. Fox. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 24(2):129–136, June 1981. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Davis:1981:FTR

[1789] Gary Davis and Stephen H. McCleary. The π-full tight Riesz orders on A(Ω). Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 24(2):137–152, June 1981. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Borwein:1981:RBL

[1790] D. Borwein and B. Watson. On the relation between the logarithmic and Borel-type summability methods. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 24(2):153–160, June 1981. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Vogt:1981:IIC

[1791] Andrew Vogt. The isoperimetric inequality for curves with selfintersections. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 24(2):161–168, June 1981. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Kizmaz:1981:CSS

[1792] H. Kizmaz. On certain sequence spaces. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 24(2):169–176, June 1981. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Cain:1981:CPS

[1793] George L. Cain, Jr. Continuous preimages of spaces with finite compactifications. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 24(2):177–180, June 1981. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Ong:1981:OTI

[1794] Sing Cheong Ong. Operator topologies and invariant operator ranges. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 24(2):181–186, June 1981. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Day:1981:SPL

[1795] Alan Day. In search of a Pappian lattice identity. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 24(2):187–198, June 1981. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Sato:1981:ETC

[1796] Ryotaro Sato. An extrapolation theorem for contractions with fixed points. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 24(2):199–204, June 1981. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Rivest:1981:PAC

[1797] Louis-Paul Rivest. A probabilistic approach to the convolution transform. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 24(2):205–212, June 1981. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Benabdallah:1981:FGA

[1798] K. Benabdallah and A. Birtz. Sur une famille de groupes abéliens superdécomposables. (French) [On a family of Abelian super-decomposable groups]. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 24(2):213–218, June 1981. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Caicedo:1981:ISA

[1799] Xavier Caicedo. Independent sets of axioms in L_{κα}. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 24(2):219– 224, June 1981. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

DeKoninck:1981:RCL

[1800] J.-M. De Koninck, P. Erdős, and A. Ivić. Reciprocals of certain large additive functions. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 24(2):225–232, June 1981. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Davison:1981:GDQ

[1801] T. M. K. Davison. On Gleason’s definition of quadratic forms. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 24(2): 233–236, June 1981. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Hoffmann:1981:EES

[1802] Rudolf-E. Hoffmann. Essential extensions of T_{1}-spaces. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 24(2):237– 240, June 1981. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Herrmann:1981:FGM

[1803] Christian Herrmann. A finitely generated modular ortholattice. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 24(2): 241–244, June 1981. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Gilman:1981:RAC

[1804] Jane Gilman. A remark about components of relative Teichmüller spaces. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 24(2):245–246, June 1981. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Feigelstock:1981:PGM

[1805] Shalom Feigelstock. On a paper of G. Mason. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 24(2):247–248, June 1981. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Lawrence:1981:NDT

[1806] John Lawrence. A note on the density theorem. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 24(2):249–256, June 1981. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Gawronski:1981:ZPS

[1807] W. Gawronski and U. Stadtmüller. On the zeros of power series with exponential logarithmic coefficients. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 24(3):257–272, September 1981. CO-DEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Whitfield:1981:CLS

[1808] J. H. M. Whitfield and S. Yong. A characterization of line spaces. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 24(3): 273–278, September 1981. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Klambauer:1981:SMP

[1809] G. Klambauer. The Stieltjes moment problem. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 24(3):279–282, September 1981. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Kotlarski:1981:CSC

[1810] H. Kotlarski, S. Krajewski, and A. H. Lachlan. Construction of satisfaction classes for nonstandard models. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 24(3):283–294, September 1981. CO-DEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Lachlan:1981:FSC

[1811] A. H. Lachlan. Full satisfaction classes and recursive saturation. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 24(3): 295–298, September 1981. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Mercier:1981:SF

[1812] Armel Mercier. Sommes de la forme ∑ g(n)/f(n). Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 24(3):299–308, September 1981. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Borwein:1981:TCE

[1813] David Borwein. Tauberian conditions for the equivalence of weighted mean and power series methods of summability. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 24(3):309–316, September 1981. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Furey:1981:NRF

[1814] Rosalita Furey and Philip Heath. Notes on Toppair, Topsbast and regular fibrations, coglueing and duality. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 24(3):317–330, September 1981. CO-DEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Thomson:1981:TVF

[1815] B. S. Thomson. On the total variation of a function. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 24(3):331–340, September 1981. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Thomson:1981:OMT

[1816] B. S. Thomson. Outer measures and total variation. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 24(3):341–346, September 1981. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Harris:1981:CIV

[1817] Lawrence A. Harris. Coefficient inequalities for L^{p}-valued analytic functions. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 24(3):347–350, September 1981. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Schirmer:1981:CUF

[1818] Helga Schirmer. Conditions for the uniqueness of the fixed point in Kakutani’s theorem. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 24(3):351–358, September 1981. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Keagy:1981:SMT

[1819] Thomas A. Keagy. Summability of matrix transforms of subsequences. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 24(3):359–364, September 1981. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Ketkar:1981:PMS

[1820] R. D. Ketkar and N. Vanaja. R-projective modules over a semiperfect ring. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 24(3):365–368, September 1981. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Tsirulnikov:1981:RHL

[1821] Bella Tsirulnikov. Remarkable hyperplanes in locally convex spaces of dimension at most c. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 24(3):369–372, September 1981. CODEN CM-BUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Sharma:1981:CSW

[1822] P. L. Sharma. A class of spaces in which compact sets are finite. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 24(3): 373–376, September 1981. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Leech:1981:RRC

[1823] John Leech. A remark on rational cuboids. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 24(3):377–378, September 1981. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Smith:1981:NDT

[1824] R. A. Smith. A note on Dirichlet’s theorem. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 24(3):379–384, September 1981. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Gildenhuys:1981:CDS

[1825] D. Gildenhuys and R. Strebel. On the cohomological dimension of soluble groups. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 24(4):385–392, December 1981. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Chan:1981:SFE

[1826] Ling Yau Chan. Some further extensions of Hardy’s inequality. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 24(4): 393–400, December 1981. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Janos:1981:NUA

[1827] Ludvik Janos, Roger C. McCann, and J. L. Solomon. Nonexpansive uniformly asymptotically stable flows are linear. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 24(4):401–408, December 1981. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Kreith:1981:OTN

[1828] Kurt Kreith and Takaŝi Kusano. Oscillation theorems for nonlinear ordinary differential equations of even order. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 24(4):409–414, December 1981. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Chung:1981:SRN

[1829] L. O. Chung and Jiang Luh. Semiprime rings with nilpotent derivatives. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 24(4):415–422, December 1981. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Orecchia:1981:OSA

[1830] Ferruccio Orecchia. Ordinary singularities of algebraic curves. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 24(4): 423–432, December 1981. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Lee:1981:IST

[1831] Cheng Ming Lee. On integrals and summable trigonometric series. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 24(4):433–440, December 1981. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Kirk:1981:AFP

[1832] W. A. Kirk and Claudio Morales. On the approximation of fixed points of locally nonexpansive mappings. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 24(4):441–446, December 1981. CO-DEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Smyth:1981:KTT

[1833] C. J. Smyth. A Kronecker-type theorem for complex polynomials in several variables. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 24(4):447–452, December 1981. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic). See addenda and errata [1917].

Boyd:1981:SCR

[1834] David W. Boyd. Speculations concerning the range of Mahler’s measure. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 24(4):453–470, December 1981. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Hong:1981:RTS

[1835] S. S. Hong and L. D. Nel. Reflexive topological semilattices. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 24(4): 471–482, December 1981. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Mercier:1981:RRC

[1836] Armel Mercier. Remarques sur la représentation de certaines séries de Dirichlet. (French) [Remarks on the representation of certain Dirichlet series]. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 24(4):483–484, December 1981. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Johnston:1981:GCT

[1837] Elgin Johnston. A generalized comparison test. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 24(4):485–488, December 1981. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

MacDougall:1981:BFF

[1838] J. A. MacDougall. Bivectors over a finite field. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 24(4):489–490, December 1981. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Ho:1981:WID

[1839] Chung Wu Ho. When are immersions diffeomorphisms? Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 24(4):491– 492, December 1981. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Rodman:1981:RIF

[1840] Leiba Rodman. Remark on irreducible factorizations of polynomials in several variables. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 24(4):493–496, December 1981. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Pollington:1981:IDS

[1841] Andrew Pollington and Charles Vanden Eynden. The integers as differences of a sequence. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 24(4):497–499, December 1981. CODEN CM-BUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Bantegnie:1982:PEF

[1842] Robert Bantegnie. Polyèdres équilibrés. (French) [Equilibrated polyhedra]. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 25(1):3–12, March 1982. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Prevot:1982:CPM

[1843] Kenneth Prevot. Cutting and pasting Z_{p}-manifolds. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 25(1):13–28, March 1982. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Campbell:1982:ESI

[1844] L. L. Campbell. Existence of solutions of an ill-posed problem for the vibrating string. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 25(1):29–36, March 1982. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Conway:1982:QSO

[1845] John B. Conway. On quasisimilarity for subnormal operators. II. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 25(1): 37–40, March 1982. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

vanMill:1982:PIZ

[1846] Jan van Mill and R. Grant Woods. Perfect images of zero-dimensional separable metric spaces. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 25(1):41–47, March 1982. CODEN CM-BUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Lanski:1982:IIN

[1847] Charles Lanski. Idempotent ideals and Noetherian polynomial rings. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 25(1):48–53, March 1982. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Dunham:1982:LBV

[1848] Charles B. Dunham. The limit of biased varisolvent Chebyshev approximation. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 25(1):54–58, March 1982. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Espelie:1982:RTW

[1849] M. Solveig Espelie and James E. Joseph. Remarks on two weak forms of continuity. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 25(1):59–63, March 1982. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Swift:1982:SGR

[1850] Robert W. Swift. On the smoothness of generic rings for stably free modules. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 25(1):64–70, March 1982. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Kusano:1982:OCC

[1851] Takaŝi Kusano and Manabu Naito. Oscillation criteria for a class of perturbed Schrödinger equations. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 25(1):71–77, March 1982. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Feder:1982:NPS

[1852] Moshe Feder. On the nonexistence of a projection onto the space of compact operators. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 25(1):78–81, March 1982. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Erbe:1982:ICT

[1853] Lynn Erbe. Integral comparison theorems for scalar Riccati equations and applications. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 25(1):82–97, March 1982. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Milojevic:1982:SCT

[1854] P. S. Milojević. On the solvability and continuation type results for nonlinear equations with applications. II. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 25(1):98–109, March 1982. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Campbell:1982:FAS

[1855] Douglas M. Campbell. Functions in all H^{p} spaces, p < ∞. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 25(1): 110–113, March 1982. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Hausman:1982:SSA

[1856] Miriam Hausman. The solution of a special arithmetic equation. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 25(1): 114–117, March 1982. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Page:1982:RRV

[1857] S. S. Page. Regular rings are very regular. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 25(1):118, March 1982. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Burr:1982:ASD

[1858] Stefan A. Burr. Antidirected subtrees of directed graphs. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 25(1):119– 120, March 1982. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

NGuerekata:1982:RSP

[1859] Gaston Mandata N’Guérékata. Remarques sur les solutions presquepériodiques de l’équation [(d/dt) − A]x = 0. (French) [Remarks on the nearly-periodic solutions of the equation [(d/dt) − A]x = 0]. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 25(1): 121–128, March 1982. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Marsden:1982:GTA

[1860] Jerrold E. Marsden. A group theoretic approach to the equations of plasma physics. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 25(2):129–142, June 1982. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Mahajan:1982:NCA

[1861] Arvind Mahajan and Dieter K. Ross. A note on completely and absolutely monotone functions. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 25(2):143–148, June 1982. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Rigby:1982:FMT

[1862] J. F. Rigby. Five mutually tangent spheres and various associated configurations. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 25(2):149–163, June 1982. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Saab:1982:IRB

[1863] Paulette Saab. Integral representation by boundary vector measures. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 25(2):164–168, June 1982. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Niefield:1982:TIF

[1864] S. B. Niefield. Topologically indexed function spaces and adjoint functors. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 25(2):169–178, June 1982. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Hopkins:1982:GIR

[1865] Glenn Hopkins and William Staton. Girth and independence ratio. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 25 (2):179–186, June 1982. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Klein:1982:SGU

[1866] Abraham A. Klein. Solvable groups of unipotent elements in a ring. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 25(2):187–190, June 1982. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Brualdi:1982:NBA

[1867] Richard A. Brualdi. Notes on the Birkhoff algorithm for doubly stochastic matrices. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 25(2):191–199, June 1982. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic). See comments [2351].

Leonard:1982:DCN

[1868] Philip A. Leonard and Kenneth S. Williams. On the divisibility of the class numbers of Q(\sqrt{−p}) and Q(\sqrt{−2p}) by 16. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 25(2):200–206, June 1982. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Saab:1982:DPO

[1869] Elias Saab. On Dunford–Pettis operators. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 25(2):207–209, June 1982. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Downing:1982:ULS

[1870] David J. Downing and William O. Ray. Uniformly Lipschitzian semigroups in Hilbert space. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 25(2):210–214, June 1982. CODEN CM-BUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Smythe:1982:PGI

[1871] N. Smythe. A presentation for a group of integer matrices. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 25(2): 215–221, June 1982. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Mollin:1982:CSF

[1872] R. A. Mollin. Cyclotomic splitting fields. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 25(2):222–229, June 1982. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Westreich:1982:SBB

[1873] David Westreich and Esther Podolak. Systems of Briot–Bouquet equations with analytic solutions. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 25(2):230–233, June 1982. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Keener:1982:NPS

[1874] Lee L. Keener. A new proof of the snake theorem. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 25(2):234–237, June 1982. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Prodinger:1982:NPT

[1875] Helmut Prodinger. On the number of partitions of 1, ⋯, n into two sets of equal cardinalities and equal sums. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 25(2):238–241, June 1982. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Warren:1982:DSD

[1876] Richard H. Warren. Discrete sets and discrete maps. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 25(2):242– 244, June 1982. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Martindale:1982:ECC

[1877] Wallace S. Martindale III. The extended center of coproducts. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 25(2): 245–256, June 1982. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Rooney:1982:MMT

[1878] P. G. Rooney. Multipliers for the Mellin transformation. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 25(3):257– 262, September 1982. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Thorpe:1982:NSR

[1879] B. Thorpe. A note on strong Riesz summability. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 25(3):263–272, September 1982. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Hardy:1982:SFF

[1880] G. E. Hardy and M. V. Subbarao. Semi r-free and r-free integers — a unified approach. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 25(3):273–290, September 1982. CODEN CM-BUA3. ISSN 0008-4395 (print), 1496-4287 (electronic). See corrigendum and addendum [2081].

Littlejohn:1982:NOP

[1881] Lance L. Littlejohn and Samuel D. Shore. Nonclassical orthogonal polynomials as solutions to second order differential equations. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 25(3): 291–295, September 1982. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Muller:1982:DCM

[1882] Bruno J. Müller and S. Tariq Rizvi. On the decomposition of continuous modules. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 25(3):296–301, September 1982. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Gardner:1982:SIR

[1883] R. J. Gardner, S. Kwapień, and D. P. Laurie. Some inequalities related to planar convex sets. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 25(3):302–310, September 1982. CODEN CM-BUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Johnson:1982:CPT

[1884] B. E. Johnson. A counterexample in the perturbation theory of C^{∗}- algebras. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 25(3):311–316, September 1982. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Hoo:1982:DUS

[1885] C. S. Hoo and K. P. Shum. 0-distributive and P -uniform semilattices. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 25(3):317–324, September 1982. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Haack:1982:DCZ

[1886] Joel K. Haack. The duals of the Camillo–Zelmanowitz formulas for Goldie dimension. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 25(3):325–334, September 1982. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Roberts:1982:FIE

[1887] K. L. Roberts. On fractional integrals equivalent to a constant. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 25(3): 335–338, September 1982. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Tan:1982:NAN

[1888] Kok Keong Tan. A note on asymptotic normal structure and closeto-normal structure. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 25(3):339–343, September 1982. CODEN CM-BUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Burgess:1982:ADL

[1889] W. D. Burgess and J.-M. Goursaud. K_{0} d’un anneau dont les localisés centraux sont simples artiniens. (French) [K_{0} of a ring in which the central localizations are simple Artinian]. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 25(3):344–347, September 1982. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Ladas:1982:DDI

[1890] G. Ladas and I. P. Stavroulakis. On delay differential inequalities of higher order. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 25(3):348–354, September 1982. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Nieto:1982:OIU

[1891] José I. Nieto. Opérateurs à itérés uniformément bornés. (French) [Uniformly-bounded iterated operators]. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 25(3):355–360, September 1982. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Swift:1982:FSS

[1892] Robert W. Swift. Free summands of stably free modules. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 25(3): 361–362, September 1982. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Robbins:1982:NBC

[1893] Neville Robbins. On the number of binomial coefficients which are divisible by their row number. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 25(3):363–365, September 1982. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Williams:1982:NFQ

[1894] H. C. Williams. A note on the Fibonacci quotient F_{p−ε}/p. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 25(3): 366–370, September 1982. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Nagel:1982:LDP

[1895] Rainer Nagel and Günther Palm. Lattice dilations of positive contractions on L^{p}-spaces. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 25(3):371–374, September 1982. CODEN CM-BUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Boyarsky:1982:FM

[1896] Abraham Boyarsky and William Byers. Fullness of maps. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 25(3): 375–376, September 1982. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Kalton:1982:ETB

[1897] N. J. Kalton. An example in the theory of bilinear maps. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 25(3): 377–384, September 1982. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Cross:1982:ESD

[1898] G. E. Cross. The exceptional sets in the definition of the P^{n}-integral. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 25(4):385–391, December 1982. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Temam:1982:AFC

[1899] R. Temam. Approximation de fonctions convexes sur un espace de mesures et applications. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 25(4):392–413, December 1982. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Wilker:1982:HOP

[1900] J. B. Wilker. Hausdorff obstructions to packing (N −1)-balls in N-space. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 25(4):414–420, December 1982. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Fridy:1982:AT

[1901] J. A. Fridy. Abel transformations into l^{1}. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 25(4):421–427, December 1982. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Berner:1982:ACF

[1902] Andrew J. Berner. Almost continuous functions with closed graphs. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 25(4):428–434, December 1982. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

London:1982:SQD

[1903] David London and Binyamin Schwarz. Strong and quasistrong disconjugacy. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 25(4):435–440, December 1982. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Pletch:1982:LFP

[1904] Andrew Pletch. Local freeness of profinite groups. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 25(4):441–446, December 1982. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

DeFranza:1982:GIT

[1905] J. DeFranza. A general inclusion theorem for l − l Nörlund summability. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 25(4):447–455, December 1982. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Karnik:1982:NCQ

[1906] S. M. Karnik and S. Willard. Natural covers and R-quotient maps. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 25(4):456–461, December 1982. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Feigelstock:1982:RI

[1907] Shalom Feigelstock and Aaron Klein. Retracts and injectives. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 25(4): 462–467, December 1982. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Olsen:1982:IWE

[1908] James H. Olsen. The individual weighted ergodic theorem for bounded Besicovitch sequences. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 25(4):468–471, December 1982. CODEN CM-BUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Bell:1982:UDS

[1909] Murray Bell and John Ginsburg. Uncountable discrete sets in extensions and metrizability. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 25(4):472–477, December 1982. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Shamsuddin:1982:RAL

[1910] Ahmad Shamsuddin. Rings with automorphisms leaving no nontrivial proper ideals invariant. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 25(4):478–486, December 1982. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Rode:1982:CTD

[1911] Gerd Rodé. A characterization of the topological dimension. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 25(4): 487–490, December 1982. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Albert:1982:BSF

[1912] Michael Albert and John A. Baker. Bounded solutions of a functional inequality. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 25(4):491–495, December 1982. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Cowling:1982:DSP

[1913] V. F. Cowling and J. P. King. Detection of singular points. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 25(4): 496–497, December 1982. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Nain:1982:CRG

[1914] R. B. Nain and Kanwar Sen. Correction to a result of G. C. Jain on a correlated random walk. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 25(4):498–499, December 1982. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic). See [1062].

Szekelyhidi:1982:NST

[1915] László Székelyhidi. Note on a stability theorem. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 25(4):500–501, December 1982. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Keener:1982:BIA

[1916] Lee L. Keener. The best interpolating approximation is a limit of best weighted approximations. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 25(4):502–503, December 1982. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Smyth:1982:AEK

[1917] C. J. Smyth. Addenda and errata for: “A Kronecker-type theorem for complex polynomials in several variables”. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 25(4):504, December 1982. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic). See [1833].

Bergen:1983:RSK

[1918] Jeffrey Bergen and I. N. Herstein. Rings with a special kind of automorphism. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 26(1):3–8, March 1983. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Padmanabhan:1983:SDE

[1919] R. Padmanabhan. A self-dual equational basis for Boolean algebras. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 26(1):9–12, March 1983. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Bouvier:1983:LCG

[1920] A. Bouvier. The local class group of a Krull domain. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 26(1):13–19, March 1983. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Krebs:1983:EES

[1921] William B. Krebs and James B. Robertson. On ergodic extensions of stationary measures with minimal support. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 26(1):20–25, March 1983. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Fox:1983:USV

[1922] Geoffrey Fox and Pedro Morales. On uniform semigroup-valued additive set functions. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 26(1):26–34, March 1983. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Grace:1983:OCO

[1923] S. R. Grace and B. S. Lalli. An oscillation criterion for nth order nonlinear differential equations with functional arguments. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 26(1):35–40, March 1983. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Komjath:1983:LSC

[1924] Péter Komjáth. Large sets not containing images of a given sequence. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 26(1):41–43, March 1983. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Kulkarni:1983:TCG

[1925] S. H. Kulkarni and B. V. Limaye. A topological characterization of Gleason parts of real function algebras. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 26(1):44–49, March 1983. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Borwein:1983:PBR

[1926] Peter B. Borwein. On Padé and best rational approximation. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 26(1): 50–57, March 1983. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Crombez:1983:CAP

[1927] G. Crombez. Compactness and almost periodicity of multipliers. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 26(1): 58–62, March 1983. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Szulga:1983:LAA

[1928] J. Szulga. On lattice analogues of absolutely summing constants. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 26(1): 63–69, March 1983. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

vanMill:1983:BAR

[1929] Jan van Mill. Boolean algebras and raising maps to zero-dimensional spaces. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 26(1):70–79, March 1983. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Anand:1983:ROG

[1930] K. K. Anand. On relaxation oscillations governed by a second order differential equation for a large parameter and with a piecewise linear function. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 26(1):80–91, March 1983. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Djokovic:1983:CEC

[1931] Dragomir Z. Djoković. Closures of equivalence classes of trivectors of an eight-dimensional complex vector space. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 26(1):92–100, March 1983. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Webb:1983:RGC

[1932] U. H. M. Webb. On the rank of a p-group of class 2. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 26(1):101– 105, March 1983. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Anderson:1983:PI

[1933] David F. Anderson and David E. Dobbs. On the product of ideals. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 26(1):106–114, March 1983. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Robinson:1983:UCP

[1934] Stewart M. Robinson and Zachary Robinson. Uniformly continuous partitions of unity on a metric space. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 26(1):115–117, March 1983. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Polak:1983:BSW

[1935] T. Polak and Brailey Sims. A Banach space which is fully 2-rotund but not locally uniformly rotund. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 26(1):118–120, March 1983. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Ginsburg:1983:STF

[1936] John Ginsburg. A structure theorem in finite topology. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 26(1):121– 122, March 1983. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Williams:1983:QRI

[1937] Kenneth S. Williams. On the quadratic residues (modp) in the interval (0, , p/4). Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 26(1):123–128, March 1983. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Vanstone:1983:PMM

[1938] J. R. Vanstone. The Poincaré map in mixed exterior algebra. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 26(2): 129–136, June 1983. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Browne:1983:SMM

[1939] Patrick J. Browne and Claude Costa. L^{p} spaces from matrix measures: a correction and their interpolation. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 26(2):137–143, June 1983. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Heinrich:1983:CPI

[1940] Katherine Heinrich. A covering problem for idempotent Latin squares. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 26(2):144–148, June 1983. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Cater:1983:TND

[1941] F. S. Cater. A typical nowhere differentiable function. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 26(2):149– 151, June 1983. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Hartman:1983:CRC

[1942] Jehuda Hartman and Meir Katchalski. On k-cycled refinements of certain graphs. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 26(2):152–156, June 1983. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Lee:1983:ETG

[1943] Tan Yu Lee. Embedding theorems in group C^{∗}-algebras. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 26(2):157– 166, June 1983. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Schirmer:1983:CSC

[1944] Helga Schirmer. Coincidence sets of coincidence producing maps. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 26 (2):167–170, June 1983. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Swanson:1983:EPS

[1945] C. A. Swanson. Extremal positive solutions of semilinear Schrödinger equations. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 26(2):171–178, June 1983. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Tomm:1983:RSM

[1946] Ludwig Tomm. A regular summability method which sums the geometric series to its proper value in the whole complex plane. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 26(2): 179–188, June 1983. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Aharonov:1983:HCT

[1947] Dov Aharonov. Hölder conditions and the topology of simply connected domains. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 26(2):189–191, June 1983. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Ohta:1983:MEZ

[1948] Haruto Ohta. The maximal extension of a zero-dimensional product space. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 26(2):192–201, June 1983. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Samaris:1983:CES

[1949] Nicolas Samaris. Coefficient estimates of some classes of analytic functions. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 26(2):202–208, June 1983. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Lacroix:1983:SGN

[1950] N. H. J. Lacroix and C. Levesque. Sur les sous-groupes normaux de SL_{2} sur un anneau local. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 26(2):209–219, June 1983. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Beatson:1983:CA

[1951] R. K. Beatson and D. Leviatan. On comonotone approximation. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 26(2): 220–224, June 1983. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Al-Salam:1983:ASB

[1952] W. A. Al-Salam and A. Verma. q-analogs of some biorthogonal functions. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 26(2):225–227, June 1983. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Torrejon:1983:NLE

[1953] Ricardo Torrejón. A note on locally expansive and locally accretive operators. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 26(2):228–232, June 1983. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Milnes:1983:SCD

[1954] Paul Milnes. Semigroup compactifications of direct and semidirect products. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 26(2):233–240, June 1983. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Landesman:1983:SDT

[1955] E. M. Landesman and Bernard Russo. The second dual of a C^{∗}-ternary ring. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 26(2):241–246, June 1983. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Stavrakas:1983:KTN

[1956] N. Stavrakas. Krasnosel’skii theorems for nonseparating compact sets. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 26(2):247–256, June 1983. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Page:1983:FRS

[1957] S. S. Page. FPF rings and some conjectures of C. Faith. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 26(3): 257–259, September 1983. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Longnecker:1983:OSU

[1958] M. Longnecker. Optimal stopping under general dependence conditions. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 26(3):260–266, September 1983. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Bergen:1983:DPR

[1959] Jeffrey Bergen. Derivations in prime rings. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 26(3):267–270, September 1983. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Salzberg:1983:MFF

[1960] Pablo M. Salzberg. A multivariable form of the fundamental theorem of algebra. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 26(3):271–272, September 1983. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Wayland:1983:BSS

[1961] Keith Wayland. Bipartite score sets. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 26(3):273–279, September 1983. CO-DEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Dummit:1983:FGD

[1962] D. S. Dummit, R. Gold, and H. Kisilevsky. The field generated by the discriminant of the class invariants of an imaginary quadratic field. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 26(3): 280–282, September 1983. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Day:1983:LPG

[1963] Alan Day. A lemma on projective geometries as modular and/or Arguesian lattices. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 26(3):283–290, September 1983. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Maeda:1983:IIP

[1964] Sadahiro Maeda. Isotropic immersions with parallel second fundamental form. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 26(3):291–296, September 1983. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Lorimer:1983:DNT

[1965] J. W. Lorimer. Dual numbers and topological Hjelmslev planes. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 26(3): 297–302, September 1983. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Matsuyama:1983:SHS

[1966] Yoshio Matsuyama. Some hypersurfaces of symmetric spaces. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 26(3): 303–311, September 1983. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Prodinger:1983:OFP

[1967] Helmut Prodinger. Ordered Fibonacci partitions. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 26(3):312–316, September 1983. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Borwein:1983:UBR

[1968] Peter B. Borwein. The usual behaviour of rational approximations. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 26(3):317–323, September 1983. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Gilligan:1983:PLP

[1969] Bruce Gilligan. On a parametrized Levi problem involving one complex variable. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 26(3):324–327, September 1983. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Oberlin:1983:DFS

[1970] Daniel M. Oberlin. Divergence of Fourier series. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 26(3):328–330, September 1983. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Anderson:1983:IE

[1971] D. D. Anderson. On the ideal equation I(B ∩ C) = IB ∩ IC. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 26(3): 331–332, September 1983. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Karanikas:1983:NWP

[1972] C. Karanikas. A note on a Wiener–Pitt theorem on Wiener groups. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 26(3):333–336, September 1983. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Wagon:1983:PIS

[1973] Stanley Wagon. Partitioning intervals, spheres and balls into congruent pieces. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 26(3):337–340, September 1983. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Chung:1983:ND

[1974] L. O. Chung and Jiang Luh. Nilpotency of derivations. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 26(3):341–346, September 1983. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic). See corrigendum [2205].

Hatzenbuhler:1983:ACR

[1975] James Hatzenbuhler, Don A. Mattson, and Walter S. Sizer. An algebraic characterization of remainders of compactifications. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 26(3):347–350, September 1983. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Cordovil:1983:OMG

[1976] Raul Cordovil. Oriented matroids and geometric sorting. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 26(3):351– 354, September 1983. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Sharma:1983:CCC

[1977] Ramesh Sharma. Certain characteristics of the curvature tensor of an affinely connected manifold. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 26(3):355–357, September 1983. CO-DEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Gil-Medrano:1983:GBI

[1978] O. Gil-Medrano and A. M. Naveira. The Gauss–Bonnet integrand for a class of Riemannian manifolds admitting two orthogonal complementary foliations. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 26(3):358–364, September 1983. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Wolk:1983:PDC

[1979] Elliot S. Wolk. On the principle of dependent choices and some forms of Zorn’s lemma. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 26(3):365–367, September 1983. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Kreith:1983:OPW

[1980] Kurt Kreith. Oscillation properties of weakly time dependent hyperbolic equations. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 26(3):368–384, September 1983. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Rudin:1983:SP

[1981] Mary Ellen Rudin. The shrinking property. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 26(4):385–388, December 1983. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Diamond:1983:ITA

[1982] Harvey Diamond, Brian Kuttner, and Louise A. Raphael. Inclusion theorems for the absolute summability of divergent integrals. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 26(4): 389–398, December 1983. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Ortel:1983:CLC

[1983] Marvin Ortel and Walter Schneider. Curvature of level curves of harmonic functions. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 26(4):399–405, December 1983. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Rush:1983:ECR

[1984] David E. Rush. The equation X^{k} + Y^{k} = Z^{k} in commuting rational matrices. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 26(4):406–409, December 1983. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Littlejohn:1983:KPS

[1985] Lance L. Littlejohn. The Krall polynomials as solutions to a second order differential equation. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 26(4):410–417, December 1983. CODEN CM-BUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Beer:1983:UCC

[1986] Gerald Beer. On uniform convergence of continuous functions and topological convergence of sets. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 26(4):418–424, December 1983. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Dubuc:1983:RDT

[1987] Serge Dubuc and Fabian Todor. La règle du trapèze appliquée à quelques fonctions sans dérivées. (French) [The trapezoid rule applied to some derivative-free functions]. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 26(4):425–429, December 1983. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Dean:1983:NET

[1988] Alice M. Dean. Nearnesses and T_{0}-extensions of topological spaces. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 26(4):430–437, December 1983. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Srivastava:1983:CAT

[1989] H. M. Srivastava, J.-L. Lavoie, and Richard Tremblay. A class of addition theorems. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 26(4):438–445, December 1983. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Gierz:1983:MCC

[1990] G. Gierz, J. D. Lawson, and A. R. Stralka. Metrizability conditions for completely distributive lattices. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 26(4):446–453, December 1983. CO-DEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Shawyer:1983:ACP

[1991] Bruce L. R. Shawyer and Ludwig Tomm. Analytic continuation of power series by regular generalized weighted means. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 26(4):454–463, December 1983. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Gerth:1983:ARC

[1992] Frank Gerth III. Asymptotic results for class number divisibility in cyclotomic fields. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 26(4):464–472, December 1983. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Saks:1983:ECC

[1993] Victor Saks. Extra countably compact spaces. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 26(4):473–481, December 1983. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Horton:1983:SNE

[1994] J. D. Horton. Sets with no empty convex 7-gons. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 26(4):482–484, December 1983. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Nicolas:1983:MEP

[1995] J.-L. Nicolas and G. Robin. Majorations explicites pour le nombre de diviseurs de N. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 26(4):485–492, December 1983. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Olsen:1983:MSE

[1996] James H. Olsen. A multiple sequence ergodic theorem. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 26(4):493– 497, December 1983. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Broughton:1983:HTD

[1997] S. Allen Broughton, Michael Hoffman, and William Homer. The height of two-dimensional cohomology classes of complex flag manifolds. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 26(4): 498–502, December 1983. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Schafer:1984:FCI

[1998] James A. Schafer. Finite complexes and integral representations. II. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 27(1):3–9, March 1984. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Murphy:1984:LIA

[1999] G. J. Murphy. Lie ideals in associative algebras. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 27(1):10–15, March 1984. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Baron:1984:ASF

[2000] S. Baron. Absolute summability factors in a sequence. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 27(1):16–30, March 1984. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Ghatage:1984:AFL

[2001] Pratibha G. Ghatage. Additive functionals on Lorentz spaces. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 27(1): 31–37, March 1984. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Benabdallah:1984:PGW

[2002] K. Benabdallah and A. Laroche. Primary groups whose basic subgroup decompositions can be lifted. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 27(1):38–42, March 1984. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Nara:1984:SGS

[2003] Chiê Nara and Iwao Sato. Split graphs with specified Dilworth numbers. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 27(1):43–47, March 1984. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Houston:1984:CMD

[2004] Evan G. Houston, Saroj B. Malik, and Joe L. Mott. Characterizations of ∗-multiplication domains. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 27(1):48–52, March 1984. CODEN CM-BUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Aguirre:1984:SMC

[2005] José A. Facenda Aguirre. (mHM)-spaces and measurable cardinals. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 27 (1):53–57, March 1984. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Grassi:1984:SRM

[2006] Peter Grassi. On subspaces of replete and measure replete spaces. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 27 (1):58–64, March 1984. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Miller:1984:RMG

[2007] J. F. Miller and H. B. Skerry. Regular and Mercerian generalized Lototsky method. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 27(1):65–71, March 1984. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Cusick:1984:FAE

[2008] Larry W. Cusick. Free actions by elementary abelian 2-groups on Stiefel manifolds. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 27(1):72–77, March 1984. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Lim:1984:ESG

[2009] K. L. Lim. On evaluation subgroups of generalized homotopy groups. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 27(1):78–86, March 1984. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Giordano:1984:SGU

[2010] Thierry Giordano and Pierre de la Harpe. Simplicité des groupes unitaires définis par un facteur simple. (French) [Simplicity of unitary groups defined by a simple factor]. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 27(1):87–95, March 1984. CODEN CM-BUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Pomerol:1984:NOA

[2011] J.-Ch. Pomerol. Nearly open, according to a subset, linear maps which are open. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 27(1):96–101, March 1984. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Philos:1984:SOS

[2012] Ch. G. Philos. A second order superlinear oscillation criterion. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 27(1): 102–112, March 1984. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Dowlen:1984:FSS

[2013] Mary Dowlen. The fixed subring of some groups of ring automorphisms. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 27(1):113–116, March 1984. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Johnson:1984:IWO

[2014] George W. Johnson and Ju Rang Yan. An improved Wintner oscillation criterion for second order linear differential equations. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 27(1):117–121, March 1984. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Mayne:1984:CMP

[2015] Joseph H. Mayne. Centralizing mappings of prime rings. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 27(1):122– 126, March 1984. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Trew:1984:NGM

[2016] S. Trew and P. Zvengrowski. Nonparallelizability of Grassmann manifolds. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 27(1):127–128, March 1984. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Harary:1984:RDC

[2017] Frank Harary and Razmik Karabed. Regular digraphs containing a given digraph. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 27(2):129–133, June 1984. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Sanjurjo:1984:FAA

[2018] J. M. R. Sanjurjo. On fundamental approximative absolute neighborhood retracts. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 27(2):134–142, June 1984. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Bressoud:1984:UCB

[2019] D. M. Bressoud and M. V. Subbarao. On Uchimura’s connection between partitions and the number of divisors. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 27(2):143–145, June 1984. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Seddighi:1984:NO

[2020] Karim Seddighi. von Neumann operators in B_{1}(Ω). Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 27(2):146–156, June 1984. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Pettet:1984:NPI

[2021] Martin R. Pettet. A note on partition-inducing automorphism groups. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 27(2):157–159, June 1984. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Kosler:1984:SRQ

[2022] Karl A. Kosler. Semicritical rings and the quotient problem. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 27(2):

160–170, June 1984. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Amberg:1984:FGM

[2023] Bernhard Amberg. Factorized groups with max, min and min-p. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 27(2): 171–178, June 1984. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Ireland:1984:NBG

[2024] K. Ireland and D. Small. A note on Bernoulli–Goss polynomials. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 27(2): 179–184, June 1984. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Myerson:1984:MPS

[2025] Gerald Myerson. A measure for polynomials in several variables. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 27(2): 185–191, June 1984. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Steinlein:1984:TBU

[2026] H. Steinlein. On the theorems of Borsuk–Ulam and Ljusternik– Schnirelmann–Borsuk. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 27(2):192–204, June 1984. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Koornwinder:1984:OPW

[2027] Tom H. Koornwinder. Orthogonal polynomials with weight function (1− x)^{α}(1+x)^{η} +Mδ(x+1)+Nδ(x−1). Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 27(2):205–214, June 1984. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Laue:1984:ALR

[2028] Hartmut Laue. On the associated Lie ring and the adjoint group of a radical ring. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 27(2):215–222, June 1984. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Kusano:1984:ATS

[2029] Takaŝi Kusano and Charles A. Swanson. Asymptotic theory of singular semilinear elliptic equations. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 27(2):223–232, June 1984. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Groemer:1984:SRT

[2030] H. Groemer. Some remarks on translative coverings of convex domains by strips. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 27(2):233–237, June 1984. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Anderson:1984:RTD

[2031] Marlow Anderson and C. C. Edwards. A representation theorem for distributive l-monoids. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 27(2):238–240, June 1984. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Vauthier:1984:EPC

[2032] Jacques Vauthier. An elementary probabilistic computation of the Poisson kernel for the n = 2 and 3 Euclidean ball. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 27(2):241–246, June 1984. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Lenagan:1984:GKD

[2033] T. H. Lenagan. Gel’fand–Kirillov dimension is exact for Noetherian PI algebras. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 27(2):247–250, June 1984. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

McKee:1984:LAC

[2034] T. A. McKee. Logical aspects of combinatorial duality. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 27(2):251– 256, June 1984. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Price:1984:NTB

[2035] Martin E. Price. A note on the torsion of Bernstein curves. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 27(3): 257–259, September 1984. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Dixon:1984:HGH

[2036] John D. Dixon. How good is Hadamard’s inequality for determinants? Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 27(3):260–264, September 1984. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Aziz:1984:ZFR

[2037] Abdul Aziz and Q. G. Mohammad. Zero-free regions for polynomials and some generalizations of Eneström–Kakeya theorem. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 27(3):265–272, September 1984. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Lee:1984:MTB

[2038] Cheng Ming Lee. A monotonicity theorem and a Bernoulli–l’Hospital– Ostrowski rule. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 27(3):273–278, September 1984. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Hoo:1984:API

[2039] C. S. Hoo. Atoms, primes and implicative lattices. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 27(3):279–285, September 1984. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

ODonovan:1984:IIA

[2040] Donal P. O’Donovan and Kenneth R. Davidson. Isometric images of C^{∗} algebras. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 27(3):286–294, September 1984. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Fernandes:1984:OFP

[2041] Olga Fernandes. On an oval with the four point Pascalian property. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 27(3):295–300, September 1984. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Ballet:1984:DAS

[2042] B. Ballet. Diviseurs dans les anneaux de séries formelles en une infinité d’indéterminées. (French) [Divisors in rings of formal series in an infinity of indeterminates]. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 27(3):301–308, September 1984. CODEN CM-BUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Goncalves:1984:FSU

[2043] Jairo Zacarias Gonçalves. Free subgroups of units in group rings. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 27(3): 309–312, September 1984. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

vandenDries:1984:AGG

[2044] Lou van den Dries and Paulo Ribenboim. The absolute Galois group of a rational function field in characteristic zero is a semidirect product. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 27(3):313–315, September 1984. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Dodds:1984:TCO

[2045] P. G. Dodds. Topological completeness of order intervals in Riesz spaces. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 27(3):316–323, September 1984. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Fontana:1984:DPI

[2046] Marco Fontana, James A. Huckaba, and Ira J. Papick. Divisorial prime ideals in Prüfer domains. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 27(3):324–328, September 1984. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Al-Salam:1984:CCU

[2047] Waleed Al-Salam, Wm. R. Allaway, and Richard Askey. A characterization of the continuous q-ultraspherical polynomials. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 27(3):329–336, September 1984. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Dikshit:1984:AEE

[2048] H. P. Dikshit, A. Sharma, and J. Tzimbalario. Asymptotic error expansions for spline interpolation. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 27(3):337–344, September 1984. CO-DEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Zamfirescu:1984:CFS

[2049] Tudor Zamfirescu and Andreana Zucco. Continuous families of smooth curves and Grünbaum’s conjecture. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 27(3):345–350, September 1984. CO-DEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Barucci:1984:CCI

[2050] Valentina Barucci and David E. Dobbs. On chain conditions in integral domains. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 27(3):351–359, September 1984. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Sjogren:1984:RBH

[2051] Peter Sjögren. A remark on bases in Hardy spaces. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 27(3):360–364, September 1984. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Goncalves:1984:FGS

[2052] Jairo Z. Gonçalves. Free groups in subnormal subgroups and the residual nilpotence of the group of units of group rings. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 27(3):365–370, September 1984. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Tannenbaum:1984:NAL

[2053] Allen Tannenbaum. A note about linear systems on curves. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 27(3): 371–374, September 1984. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Colbourn:1984:BSP

[2054] Charles J. Colbourn, Kevin T. Phelps, and Vojtech Rödl. Block sizes in pairwise balanced designs. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 27(3):375–380, September 1984. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Grabiner:1984:TVS

[2055] Sandy Grabiner. Transitive vector spaces of bounded operators. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 27(3): 381–384, September 1984. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Ishikawa:1984:MRS

[2056] Fumie Ishikawa. Metrization of ranked spaces. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 27(4):385–389, December 1984. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Ali:1984:CSC

[2057] S. Twareque Ali. Commutative systems of covariance and a generalization of Mackey’s imprimitivity theorem. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 27(4):390–397, December 1984. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Watson:1984:SPC

[2058] W. Stephen Watson. Separating points and coloring principles. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 27(4): 398–404, December 1984. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Collier:1984:MGH

[2059] Manning G. Collier and John A. Kelingos. Mean growth of harmonic functions of Beurling type. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 27(4):405–409, December 1984. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Benabdallah:1984:LST

[2060] K. Benabdallah and S. Yoshioka. On p-large subgroups of p-torsion groups. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 27(4):410–416, December 1984. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Rhaly:1984:GCM

[2061] H. C. Rhaly, Jr. Generalized Cesàro matrices. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 27(4):417–422, December 1984. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic). See corrigendum [4361].

Lorimer:1984:CTH

[2062] J. W. Lorimer. Compactness in topological Hjelmslev planes. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 27(4): 423–429, December 1984. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Weit:1984:SSE

[2063] Yitzhak Weit. On spectral synthesis and ergodicity in spaces of vectorvalued functions. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 27(4):430–435, December 1984. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Borwein:1984:CHO

[2064] J. M. Borwein and P. B. Borwein. Cubic and higher order algorithms for π. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 27(4):436–443, December 1984. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Baker:1984:DTC

[2065] Andrew Baker. A decomposition theorem for certain bipolynomial Hopf algebras. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 27(4):444–447, December 1984. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Kwasik:1984:HD

[2066] Slawomir Kwasik. On homotopy domination. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 27(4):448–451, December 1984. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Willard:1984:ALD

[2067] S. Willard and U. N. B. Dissanayake. The almost Lindelöf degree. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 27(4):452–455, December 1984. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Tong:1984:CRI

[2068] Jing Cheng Tong. On the commutativity of a ring with identity. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 27(4): 456–460, December 1984. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Brandenburg:1984:SPI

[2069] Harald Brandenburg and Adam Mysior. Short proof of an internal characterization of complete regularity. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 27(4):461–462, December 1984. CO-DEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Frappier:1984:IEF

[2070] Clément Frappier. Inequalities for entire functions of exponential type. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 27(4):463–471, December 1984. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Spigler:1984:MVZ

[2071] Renato Spigler. On the monotonic variation of the zeros of ultraspherical polynomials with the parameter. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 27(4):472–477, December 1984. CO-DEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Dodson:1984:BTC

[2072] M. M. Dodson. A Brouwer type coincidence theorem and the fundamental theorem of algebra. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 27(4):478–480, December 1984. CODEN CM-BUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

White:1984:TPS

[2073] Denis A. W. White. Tensor products and singularly continuous spectrum. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 27(4):481–484, December 1984. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Rhemtulla:1984:FIP

[2074] A. H. Rhemtulla and H. Smith. A finite index property of certain solvable groups. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 27(4):485–489, December 1984. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Grow:1984:FUQ

[2075] David Grow and William C. Whicher. Finite unions of quasi-independent sets. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 27(4):490–493, December 1984. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Godefroy:1984:LUR

[2076] G. Godefroy, S. Troyanski, J. Whitefield, and V. Zizler. Locally uniformly rotund renorming and injections into c_{0}(Γ). Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 27(4):494–500, December 1984. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Nowak:1984:RSS

[2077] Werner Georg Nowak. On a result of Smith and Subbarao concerning a divisor problem. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 27(4):501–504, December 1984. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Kent:1984:CCO

[2078] D. C. Kent and G. D. Richardson. A compactification for convergence ordered spaces. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 27(4):505–513, December 1984. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Rihaoui:1984:APF

[2079] I. Rihaoui. Approximation par des fonctions lipschitziennes et critère de convergence étroite d’une suite de probabilités. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 27(4):514–516, December 1984. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Deeb:1984:EPQ

[2080] Waleed Deeb and Rahman Younis. On the extreme points of quotients of L^{∞} by Douglas algebras. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 27(4):517–522, December 1984. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Hardy:1984:CAS

[2081] G. E. Hardy and M. V. Subbarao. Corrigendum and addendum: “semir free and r-free integers — a unified approach”. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 27(4):523, December 1984. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic). See [1880].

Millson:1985:CHF

[2082] John J. Millson. Cycles and harmonic forms on locally symmetric spaces. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 28(1):3–38, March 1985. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Cegrell:1985:ERL

[2083] Urban Cegrell. On the existence of restricted K-limits. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 28(1):39–45, March 1985. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Kshirsagar:1985:RCC

[2084] Anant M. Kshirsagar and R. P. Gupta. Residual canonical correlations. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 28(1):46–51, March 1985. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Beer:1985:MCC

[2085] Gerald Beer. More on convergence of continuous functions and topological convergence of sets. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 28(1):52–59, March 1985. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Tsintsifas:1985:TIP

[2086] George Tsintsifas. Two inequalities for planar convex sets. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 28(1): 60–66, March 1985. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Rosen:1985:ECS

[2087] Jerry D. Rosen and Mary Peles Rosen. Extended centroids of skew polynomial rings. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 28(1):67–76, March 1985. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Carreras:1985:PAC

[2088] F. J. Carreras and A. M. Naveira. On the Pontrjagin algebra of a certain class of flags of foliations. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 28(1):77–83, March 1985. CODEN CM-BUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Zimmerman:1985:FGW

[2089] Jay Zimmerman. Finite groups which are automorphism groups of infinite groups only. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 28(1):84–90, March 1985. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Anderson:1985:FIP

[2090] D. D. Anderson and David F. Anderson. Finite intersections of PID or factorial overrings. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 28(1):91–97, March 1985. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Anderson:1985:ICC

[2091] David F. Anderson, Jimmy T. Arnold, and David E. Dobbs. Integrally closed condensed domains are Bezout. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 28(1):98–102, March 1985. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Lehman:1985:ACC

[2092] B. Lehman. Another class of cyclicly extensible and reducible properties. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 28(1):103–106, March 1985. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Seddighi:1985:QTO

[2093] K. Seddighi. On quasisimilarity for Toeplitz operators. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 28(1):107– 112, March 1985. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Benedicks:1985:DPD

[2094] M. Benedicks and W. F. Pfeffer. The Dirichlet problem with Denjoy– Perron integrable boundary condition. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 28(1):113–119, March 1985. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Abu-Khuzam:1985:SCP

[2095] Hazar Abu-Khuzam and Adil Yaqub. Structure of certain periodic rings. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 28(1):120–123, March 1985. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Selick:1985:FDS

[2096] Paul Selick. The fibre of the double suspension is an H-space. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 28(1): 124–128, March 1985. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Bott:1985:SRI

[2097] Raoul Bott. On some recent interactions between mathematics and physics. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 28(2):129–164, June 1985. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Kinoshita:1985:BPB

[2098] Shin’ichi Kinoshita. On the branch points in the branched coverings of links. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 28(2):165–173, June 1985. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Chang:1985:BAL

[2099] K. W. Chang and G. X. Liu. Boundary and angular layer behavior in singularly perturbed semilinear systems. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 28(2):174–183, June 1985. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Arapovic:1985:ATM

[2100] Miroslav Arapović. Approximation theorems for Manis valuations. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 28 (2):184–189, June 1985. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Nishiura:1985:CME

[2101] Eizo Nishiura. Connected maps and essentially connected spaces. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 28 (2):190–194, June 1985. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Bullen:1985:SPI

[2102] P. S. Bullen. A simple proof of integration by parts for the Perron integral. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 28(2):195–199, June 1985. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Cater:1985:EIF

[2103] F. S. Cater. Equal integrals of functions. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 28(2):200–204, June 1985. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Sato:1985:LET

[2104] Ryotaro Sato. A local ergodic theorem for multiparameter superadditive processes. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 28(2):205–211, June 1985. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Baboolal:1985:CPN

[2105] D. Baboolal, H. L. Bentley, and R. G. Ori. Connection properties in nearness spaces. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 28(2):212–217, June 1985. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Snaith:1985:SWC

[2106] Victor Snaith. Stiefel–Whitney classes of a symmetric bilinear form — a formula of Serre. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 28(2):218–222, June 1985. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Macedonska-Nosalska:1985:ACS

[2107] Olga Macedonska-Nosalska. The abelian case of Solitar’s conjecture on infinite Nielsen transformations. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 28(2):223–230, June 1985. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Onder:1985:NAQ

[2108] Turgut Önder. Nonexistence of almost-quaternion substructures on the complex projective space. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 28(2):231–232, June 1985. CODEN CM-BUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Yamashita:1985:LS

[2109] Shinji Yamashita. Sur le lemme de Schwarz. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 28(2):233–236, June 1985. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Axler:1985:ZMB

[2110] Sheldon Axler. Zero multipliers of Bergman spaces. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 28(2):237–242, June 1985. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Campbell:1985:UTI

[2111] H. E. A. Campbell. Upper triangular invariants. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 28(2):243–256, June 1985. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Repka:1985:GAR

[2112] Joe Repka. Germs associated to regular unipotent classes in p-adic sl(n). Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 28(3):257–266, September 1985. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Greub:1985:GPC

[2113] W. H. Greub, J. Malzan, and J. R. Vanstone. Groups preserving a class of bilinear functions. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 28(3):267–271, September 1985. CODEN CM-BUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Odoni:1985:SPC

[2114] R. W. K. Odoni. Scalar products of certain Hecke L-series and moments of weighted norm-counting functions. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 28(3):272–279, September 1985. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Terras:1985:CSM

[2115] Audrey Terras. The Chowla–Selberg method for Fourier expansion of higher rank Eisenstein series. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 28(3):280–294, September 1985. CO-DEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Schmidt:1985:SSC

[2116] Wolfgang M. Schmidt. Small solutions of congruences in a large number of variables. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 28(3):295–305, September 1985. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Rankin:1985:CAF

[2117] R. A. Rankin. The construction of automorphic forms from the derivatives of a given form. II. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 28(3):306–316, September 1985. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Fong:1985:RIM

[2118] C. K. Fong. Range inclusion for multilinear mappings: applications. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 28(3):317–320, September 1985. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Kani:1985:RBG

[2119] Ernst Kani. Relations between the genera and between the Hasse–Witt invariants of Galois coverings of curves. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 28(3):321–327, September 1985. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Elliott:1985:VDR

[2120] P. D. T. A. Elliott. The value distribution of reducible cubics. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 28(3): 328–336, September 1985. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Barbeau:1985:NDC

[2121] E. J. Barbeau. Numbers differing from consecutive squares by squares. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 28(3):337–342, September 1985. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Boyd:1985:WRR

[2122] David Boyd. Which rationals are ratios of Pisot sequences? Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 28(3): 343–349, September 1985. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Kisilevsky:1985:CUC

[2123] H. Kisilevsky. The cohomology of the units in certain Z_{p}-extensions. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 28(3):350–354, September 1985. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Gerardin:1985:FED

[2124] Paul Gérardin and Wen-Ch’ing Winnie Li. A functional equation for degree two local factors. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 28(3):355–371, September 1985. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Zagier:1985:MPE

[2125] D. Zagier. Modular parametrizations of elliptic curves. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 28(3):372–384, September 1985. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Coxeter:1985:SHE

[2126] H. S. M. Coxeter. The simplicial helix and the equation tan nθ = n tan θ. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 28(4):385–393, December 1985. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Heath-Brown:1985:MES

[2127] D. R. Heath-Brown. A multiple exponential sum to modulus p^{2}. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 28(4): 394–396, December 1985. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Ellers:1985:PEH

[2128] Erich W. Ellers. Products of elations and harmonic homologies. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 28(4): 397–400, December 1985. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Johnson:1985:LMH

[2129] Charles R. Johnson and Stephen Pierce. Linear maps on Hermitian matrices: the stabilizer of an inertia class. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 28(4):401–404, December 1985. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Moreno:1985:F

[2130] C. J. Moreno and F. Shahidi. The L-functions L(s, Sym^{m}(r), π). Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 28(4): 405–410, December 1985. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Masson:1985:CAC

[2131] D. Masson. Convergence and analytic continuation for a class of regular C-fractions. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 28(4):411–421, December 1985. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Perlis:1985:ADT

[2132] Robert Perlis. On the analytic determination of the trace form. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 28(4): 422–430, December 1985. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Chalk:1985:CRC

[2133] J. H. H. Chalk. On a congruence related to character sums. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 28(4): 431–439, December 1985. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Loxton:1985:ECE

[2134] J. H. Loxton and R. C. Vaughan. The estimation of complete exponential sums. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 28(4):440–454, December 1985. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Pnevmatikos:1985:RDG

[2135] Spyros N. Pnevmatikos. Une remarque sur les déformations d’un germe d’hypersurfaces à type topologique constant. (French) [Remark on the deformations of a germ of hypersurfaces of topologically-constant type]. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 28(4):455–462, December 1985. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Schurle:1985:PIV

[2136] Arlo W. Schurle. Perron integrability versus Lebesgue integrability. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 28(4):463–468, December 1985. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Johnson:1985:NNR

[2137] N. L. Johnson. A note on net replacement in transposed spreads. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 28(4): 469–471, December 1985. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Apter:1985:CST

[2138] Arthur W. Apter. A cardinal structure theorem for an ultrapower. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 28(4):472–473, December 1985. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Dorronsoro:1985:MOB

[2139] José R. Dorronsoro. Mean oscillation and Besov spaces. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 28(4):474– 480, December 1985. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Robbins:1985:NBC

[2140] Neville Robbins. On the number of binomial coefficients which are divisible by their row number. II. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 28(4):481–486, December 1985. CO-DEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Murata:1985:FNS

[2141] Minoru Murata. Finiteness of negative spectra of elliptic operators. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 28(4):487–491, December 1985. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Smith:1985:RLB

[2142] Mark A. Smith and Barry Turett. A reflexive LUR Banach space that lacks normal structure. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 28(4):492–494, December 1985. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Yabuta:1985:CIC

[2143] Kôzô Yabuta. Cauchy integral of Calderón on the graphs of functions with BMO derivatives. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 28(4):495–500, December 1985. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Franek:1985:CVC

[2144] Frantisek Franek. Certain values of completeness and saturatedness of a uniform ideal rule out certain sizes of the underlying index set. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 28(4): 501–504, December 1985. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Lawrence:1985:DCS

[2145] John Lawrence. The definability of the commutator subgroup in a variety generated by a finite group. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 28(4):505–507, December 1985. CO-DEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Fleming:1986:I

[2146] D. J. Fleming and D. M. Giarrusso. M-ideals in L(l_{1}, E). Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 29(1):

3–10, March 1986. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Marshall:1986:MND

[2147] Bernard Marshall. Mixed norm decay for the Klein–Gordon equation with initial data in L^{p}. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 29(1):11–19, March 1986. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Glover:1986:MIS

[2148] Henry Glover and William Homer. Metastable immersion, span and the two-type of a manifold. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 29(1):20–24, March 1986. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Dobbs:1986:CDD

[2149] David E. Dobbs. On the criteria of D. D. Anderson for invertible and flat ideals. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 29(1):25–32, March 1986. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Harrington:1986:CRC

[2150] David J. Harrington. Co-rank of a composition operator. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 29(1):33–36, March 1986. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

McCasland:1986:RSF

[2151] Roy L. McCasland and Marion E. Moore. On radicals of submodules of finitely generated modules. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 29(1):37–39, March 1986. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Kunzi:1986:TSU

[2152] Hans-Peter A. Künzi. Topological spaces with a unique compatible quasiuniformity. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 29(1):40–43, March 1986. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Grant:1986:PCI

[2153] Douglass L. Grant and W. W. Comfort. Products and cardinal invariants of minimal topological groups. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 29(1):44–49, March 1986. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Arnold:1986:AKD

[2154] J. T. Arnold and Ryuki Matsuda. An almost Krull domain with divisorial height one primes. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 29(1):50–53, March 1986. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Broverman:1986:SCC

[2155] S. Broverman. Some classes of θ-compactness. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 29(1):54–59, March 1986. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Foran:1986:NSD

[2156] James Foran. Nonaveraging sets, dimension and porosity. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 29(1):60–63, March 1986. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Warner:1986:FEV

[2157] Seth Warner. Finite extensions of valued fields. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 29(1):64–69, March 1986. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Guiasu:1986:OPR

[2158] Silviu Guiasu. An optimization problem related to the zeta-function. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 29(1):70–73, March 1986. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Szurek:1986:CHD

[2159] Michal Szurek. On characterization of hypersurfaces of degrees 2 and 3 in the complex projective space. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 29(1):74–78, March 1986. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Hoechsmann:1986:FFV

[2160] Klaus Hoechsmann. Functors on finite vector spaces and units in abelian group rings. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 29(1):79–83, March 1986. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Pinkus:1986:OSA

[2161] A. Pinkus and V. Totik. One-sided L^{1}-approximation. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 29(1):84–90, March 1986. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Byers:1986:NPT

[2162] Bill Byers. A note on permutations and topological entropy of continuous maps of the interval. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 29(1):91–95, March 1986. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Drury:1986:EEC

[2163] S. W. Drury. An endpoint estimate for certain k-plane transforms. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 29 (1):96–101, March 1986. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Gilmer:1986:CSC

[2164] Robert Gilmer and William Heinzer. On the cardinality of subrings of a commutative ring. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 29(1):102–108, March 1986. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

McCarthy:1986:GSD

[2165] P. J. McCarthy. A generalization of Smith’s determinant. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 29(1):109– 113, March 1986. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Feigelstock:1986:PRP

[2166] S. Feigelstock and R. Raphael. A problem on relative projectivity for abelian groups. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 29(1):114–122, March 1986. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Snaith:1986:TCS

[2167] Victor Snaith. The K-theory of the cotangent sphere bundle of fRp^{n}. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 29(1):123–128, March 1986. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Morawetz:1986:MPT

[2168] Cathleen Synge Morawetz. Mathematical problems in transonic flow. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 29(2):129–139, June 1986. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Marinari:1986:CI

[2169] Maria Grazia Marinari and Mario Raimondo. On complete intersections. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 29(2):140–145, June 1986. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Emilion:1986:SLE

[2170] R. Émilion. Semigroups in L_{∞} and local ergodic theorem. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 29(2): 146–153, June 1986. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Kunzi:1986:ECP

[2171] Hans-Peter Künzi and Peter Fletcher. Even covering properties and somewhat normal spaces. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 29(2):154–159, June 1986. CODEN CM-BUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Kiang:1986:FPT

[2172] Mo Tak Kiang and Kok-Keong Tan. Fixed point theorems for proximately nonexpansive semigroups. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 29(2):160–166, June 1986. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

McClure:1986:AAI

[2173] J. P. McClure and R. Wong. Asymptotic approximation of an integral involving the normal distribution. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 29(2):167–176, June 1986. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Aubertin:1986:AES

[2174] Bruce Aubertin. Algebraic elements and sets of uniqueness in the group of integers of a p-series field. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 29(2):177–184, June 1986. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Cusick:1986:FRC

[2175] Larry W. Cusick. Finite regular covers of surfaces. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 29(2):185–190, June 1986. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Deveney:1986:US

[2176] James K. Deveney and John N. Mordeson. Uniqueness of subfields. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 29 (2):191–196, June 1986. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Feigelstock:1986:GAO

[2177] Shalom Feigelstock. Groups admitting only finitely many nilpotent ring structures. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 29(2):197–203, June 1986. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Burns:1986:RIF

[2178] R. G. Burns, Wilfried Imrich, and Brigitte Servatius. Remarks on the intersection of finitely generated subgroups of a free group. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 29(2): 204–207, June 1986. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

DeKoninck:1986:DBC

[2179] Jean-Marie De Koninck and Aleksandar Ivić. On the distance between consecutive divisors of an integer. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 29(2):208–217, June 1986. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Bialostocki:1986:NCS

[2180] Arie Bialostocki. The nilpotency class of the p-sylow subgroups of gl(n, q) where (p, q) = 1. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 29(2):218–223, June 1986. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Sweet:1986:ATA

[2181] L. G. Sweet and J. A. MacDougall. Algebras with transitive automorphism groups. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 29(2):224–226, June 1986. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Mercier:1986:SCC

[2182] Armel Mercier. Sommes contenant des coefficients binomiaux de Gauss. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 29(2):227–237, June 1986. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Gautrin:1986:IMT

[2183] Henri-François Gautrin, Khaldoun Imam, Tapio Klemola, and Jean-Marc Terrier. An inverse mapping theorem in Fréchet spaces. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 29(2):238– 245, June 1986. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Johnson:1986:RBS

[2184] Keith Johnson. The relation between stable operations for connective and nonconnective p-local complex K-theory. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 29(2):246–256, June 1986. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Evans:1986:LAT

[2185] Benny Evans. The long annulus theorem. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 29(3):257–267, September 1986. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Ma:1986:ROC

[2186] Tsoy-Wo Ma. On rank one commutators and triangular representations. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 29(3):268–273, September 1986. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Grzeszczuk:1986:GDC

[2187] Piotr Grzeszczuk and Edmund R. Puczylowski. Goldie dimension and chain conditions for modular lattices with finite group actions. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 29(3): 274–280, September 1986. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Kuczumow:1986:FPH

[2188] T. Kuczumow and A. Stachura. Fixed points of holomorphic mappings in the Cartesian product of n unit Hilbert balls. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 29(3):281–286, September 1986. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Brownrigg:1986:LPS

[2189] D. R. K. Brownrigg. Lattice partitions with a straight line. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 29(3): 287–294, September 1986. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Deeb:1986:S

[2190] W. Deeb and M. Marzuq. H(φ) spaces. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 29(3):295–301, September 1986. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Miller:1986:SRG

[2191] J. F. Miller. Some results on generalized Lototsky summability. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 29(3): 302–307, September 1986. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Smith:1986:CSC

[2192] Patrick Adrian Neale Smith. Counterexamples to smoothing convex functions. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 29(3):308–313, September 1986. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Vainio:1986:CPL

[2193] R. Vainio. Connectedness properties of lattices. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 29(3):314–320, September 1986. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Goncalves:1986:FSR

[2194] Jairo Zacarias Gonçalves. Free subgroups and the residual nilpotence of the group of units of modular and p-adic group rings. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 29(3):321–328, September 1986. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Berger:1986:HSC

[2195] Marc A. Berger, Alexander Felzenbaum, and Aviezri Fraenkel. The Herzog–Schönheim conjecture for finite nilpotent groups. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 29(3):329– 333, September 1986. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Cross:1986:IEP

[2196] G. E. Cross. The integration of exact Peano derivatives. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 29(3):334– 340, September 1986. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Cater:1986:EPT

[2197] F. S. Cater. An elementary proof of a theorem on unilateral derivatives. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 29(3):341–343, September 1986. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Higgins:1986:MCS

[2198] Peter M. Higgins. A method for constructing square roots in finite full transformation semigroups. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 29(3):344–351, September 1986. CO-DEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Scarowsky:1986:PTD

[2199] Manny Scarowsky and Abraham Boyarsky. A practical two-dimensional ergodic theorem. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 29(3):352–357, September 1986. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Kubota:1986:SPB

[2200] Yoshihisa Kubota. Some properties of bounded holomorphic mappings defined on bounded homogeneous domains. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 29(3):358–364, September 1986. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Frappier:1986:REF

[2201] Clément Frappier. Remarks on entire functions of exponential type. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 29(3):365–371, September 1986. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Rubel:1986:GSA

[2202] Lee A. Rubel. Generalized solutions of autonomous algebraic differential equations. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 29(3):372–374, September 1986. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

ONeill:1986:IPM

[2203] John D. O’Neill. On indecomposable projective modules. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 29(3):375– 377, September 1986. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Varadarajan:1986:IA

[2204] K. Varadarajan. Injective approximations. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 29(3):378–382, September 1986. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Chung:1986:CPN

[2205] L. O. Chung and Jiang Luh. Corrigendum to the paper: “Nilpotency of derivations”. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 29(3):383–384, September 1986. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic). See [1974].

Colbourn:1986:EPG

[2206] C. J. Colbourn, R. C. Hamm, C. C. Lindner, and C. A. Rodger. Embedding partial graph designs, block designs, and triple systems with λ > 1. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 29(4):385–391, December 1986. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Alas:1986:UPU

[2207] O. T. Alas. Uniform paracompactness and uniform para-Lindelöfness. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 29(4):392–397, December 1986. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Fegan:1986:HBU

[2208] H. D. Fegan. Homogeneous bundles and universal potentials. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 29(4): 398–406, December 1986. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Siebe:1986:NCP

[2209] Wilfried Siebe. Note on continuous and purely finitely additive set functions. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 29(4):407–412, December 1986. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Malik:1986:PDF

[2210] Henrick J. Malik and Roger Trudel. Probability density function of the product and quotient of two correlated exponential random variables. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 29(4):413–418, December 1986. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Bilodeau:1986:RES

[2211] Martin Bilodeau. Sur une représentation explicite des solutions optimales d’un programme linéaire. (French) [On an explicit representation of optimal solutions of a linear program]. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 29(4):419–425, December 1986. CO-DEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Fawcett:1986:CCF

[2212] Barry Fawcett. A categorical characterization of the four colour theorem. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 29(4):426–431, December 1986. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

DiVincenzo:1986:RDA

[2213] Onofrio M. Di Vincenzo. A result on derivations with algebraic values. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 29(4):432–437, December 1986. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Ladas:1986:OND

[2214] G. Ladas and Y. G. Sficas. Oscillations of neutral delay differential equations. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 29(4):438–445, December 1986. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Gregori:1986:SCT

[2215] V. Gregori and J. Ferrer. Some classes of topological spaces with unique quasi-uniformity. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 29(4):446–449, December 1986. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Frank:1986:CPM

[2216] Rolfdieter Frank. A characterization of projective metric spaces. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 29(4): 450–455, December 1986. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Vinuesa:1986:BSB

[2217] Jaime Vinuesa and Rafael Guadalupe. Bipositive sequences — the bilateral moment problem. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 29(4):456–462, December 1986. CODEN CM-BUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Beer:1986:HDC

[2218] Gerald Beer. Hausdorff distance and a compactness criterion for continuous functions. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 29(4):463–468, December 1986. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Jardine:1986:CPS

[2219] J. F. Jardine. Cup products in sheaf cohomology. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 29(4):469–477, December 1986. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Clark:1986:MOL

[2220] Bradd Clark and Victor Schneider. The meet operator in the lattice of group topologies. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 29(4):478–481, December 1986. CODEN CM-BUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Hiriart-Urruty:1986:GFC

[2221] J.-B. Hiriart-Urruty. A general formula on the conjugate of the difference of functions. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 29(4):482–485, December 1986. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Moran:1986:PTS

[2222] Siegfried Moran. Pre-Tits systems. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 29(4):486–494, December 1986. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Gilbert:1986:FDS

[2223] William J. Gilbert. The fractal dimension of sets derived from complex bases. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 29(4):495–500, December 1986. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Barr:1986:FST

[2224] Michael Barr. Fuzzy set theory and topos theory. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 29(4):501–508, December 1986. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Boyd:1987:RAI

[2225] David W. Boyd. Reciprocal algebraic integers whose Mahler measures are nonreciprocal. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 30(1):3–8, March 1987. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Leinen:1987:ATS

[2226] Felix Leinen. An amalgamation theorem for soluble groups. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 30(1): 9–18, March 1987. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Mullen:1987:DPF

[2227] Gary L. Mullen and Harald Niederreiter. Dickson polynomials over finite fields and complete mappings. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 30(1):19–27, March 1987. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Eloe:1987:SPG

[2228] P. W. Eloe. Sign properties of Green’s functions for two classes of boundary value problems. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 30(1):28–35, March 1987. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Hoechsmann:1987:IGR

[2229] K. Hoechsmann and S. K. Sehgal. Integral group rings without proper units. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 30(1):36–42, March 1987. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Galindo:1987:VVL

[2230] Pablo Galindo. On vector-valued Lipschitz function spaces. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 30(1): 43–48, March 1987. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Tsuda:1987:RVP

[2231] Koichi Tsuda. Rectangularity versus piecewise rectangularity of product spaces. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 30(1):49–56, March 1987. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Johnson:1987:ASO

[2232] Keith Johnson. The algebra of stable operations for p-local complex Ktheory. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 30(1):57–62, March 1987. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Lim:1987:CMC

[2233] K. L. Lim. Cocyclic maps and coevaluation subgroups. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 30(1):63–71, March 1987. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Smyth:1987:SBD

[2234] W. F. Smyth. Sharp bounds on the diameter of a graph. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 30(1):72–74, March 1987. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Bochnak:1987:EFR

[2235] Jacek Bochnak and Wojciech Kucharz. Examples of factorial rings in algebraic geometry. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 30(1):75–79, March 1987. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Murty:1987:SRA

[2236] M. Ram Murty and S. Srinivasan. Some remarks on Artin’s conjecture. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 30(1):80–85, March 1987. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Lipschutz:1987:WGG

[2237] Seymour Lipschutz. On weighted geodesics in groups. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 30(1):86–91, March 1987. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Bell:1987:CMS

[2238] H. E. Bell and W. S. Martindale III. Centralizing mappings of semiprime rings. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 30(1):92–101, March 1987. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Loewen:1987:LP

[2239] Philip D. Loewen. On the Lavrentiev phenomenon. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 30(1):102–108, March 1987. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Bell:1987:CIC

[2240] Murray Bell and Jan Pelant. Continuous images of compact semilattices. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 30(1):109–113, March 1987. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Settu:1987:TUC

[2241] S. A. Settu. S_{p} transform and uniform convergence of Laurent and power series. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 30(1):114–123, March 1987. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Gentle:1987:CNR

[2242] Ron Gentle. Comment on the Nullstellensatz for regular rings. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 30(1): 124–128, March 1987. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Nakazi:1987:ITI

[2243] Takahiko Nakazi. Intersection of two invariant subspaces. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 30(2):129– 133, June 1987. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Assani:1987:CPQ

[2244] I. Assani. Sur la convergence ponctuelle de quelques suites d’opérateurs. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 30(2):134–141, June 1987. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Wong:1987:SPMa

[2245] James C. S. Wong. On the semigroup of probability measures of a locally compact semigroup. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 30(2):142–146, June 1987. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Gilkey:1987:EIN

[2246] Peter B. Gilkey. The eta invariant and nonsingular bilinear products on R^{n}. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 30(2):147–154, June 1987. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Schulte:1987:KPT

[2247] E. Schulte and J. M. Wills. Kepler–Poinsot-type realizations of regular maps of Klein, Fricke, Gordan and Sherk. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 30(2):155–164, June 1987. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Mason:1987:LRW

[2248] A. W. Mason. On GL_{2} of a local ring in which 2 is not a unit. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 30(2): 165–176, June 1987. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Assem:1987:SST

[2249] Ibrahim Assem. Separating splitting tilting modules and hereditary algebras. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 30(2):177–181, June 1987. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Kunzi:1987:HLS

[2250] Urs-Martin Künzi. A Hilbert lattice with a small automorphism group. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 30(2):182–185, June 1987. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Banas:1987:MNC

[2251] Józef Banaś. On modulus of noncompact convexity and its properties. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 30(2):186–192, June 1987. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Bondy:1987:LBI

[2252] J. A. Bondy, Glenn Hopkins, and William Staton. Lower bounds for induced forests in cubic graphs. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 30(2):193–199, June 1987. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Magajna:1987:SOE

[2253] Bojan Magajna. A system of operator equations. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 30(2):200–209, June 1987. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Sato:1987:EHT

[2254] Ryotaro Sato. On the ergodic Hilbert transform for operators in L_{p},; 1 < p < ∞. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 30(2):210–214, June 1987. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Baker:1987:HFE

[2255] A. Baker, N. Ray, and L. Schwartz. Hypersurfaces framées et l’élément β_{1} de Toda. (French) [Framed hypersurfaces and the Toda’s element β_{1}]. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 30(2):215–222, June 1987. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Bullen:1987:DMV

[2256] P. S. Bullen and D. N. Sarkhel. On Darboux and mean value properties. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 30(2):223–230, June 1987. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Allen:1987:UIG

[2257] P. J. Allen and C. Hobby. Units in integral group rings of some metacyclic groups. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 30(2):231–240, June 1987. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Hilton:1987:HIF

[2258] A. J. W. Hilton. How intricate are (2s + 1)-factorizations? Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 30(2): 241–247, June 1987. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Erdogdu:1987:DM

[2259] V. Erdogdu. Distributive modules. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 30(2):248–254, June 1987. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Wilson:1987:MGF

[2260] George V. Wilson. Mixed groups of finite nilstufe. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 30(2):255–256, June 1987. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Chalk:1987:ICS

[2261] J. H. H. Chalk. On incomplete character sums to a prime-power modulus. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 30(3):257–266, September 1987. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Pecaric:1987:KIC

[2262] J. E. Pecarić and P. R. Beesack. On Knopp’s inequality for convex functions. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 30(3):267–272, September 1987. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Wong:1987:SPMb

[2263] James C. S. Wong. On the semigroup of probability measures of a locally compact semigroup. II. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 30(3):273–281, September 1987. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Lamb:1987:CMC

[2264] Charles W. Lamb. A comparison of methods for constructing probability measures on infinite product spaces. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 30(3):282–285, September 1987. CO-DEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Lax:1987:WPR

[2265] R. F. Lax and Carl Widland. Weierstrass points on rational nodal curves of genus 3. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 30(3):286–294, September 1987. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Deville:1987:TTP

[2266] Robert Deville. Un théorème de transfert pour la propriété des boules. (French) [A transfer theorem for the property of balls]. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 30(3):295–300, September 1987. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Boyarsky:1987:UID

[2267] Abraham Boyarsky. Uniqueness of invariant densities for certain random maps of the interval. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 30(3):301–308, September 1987. CODEN CM-BUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Mercier:1987:CA

[2268] Armel Mercier. Comportement asymptotique de ∑_{p≤x} p^{a} \underline{x} ; p^{k}. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 30(3): 309–317, September 1987. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Yiu:1987:SSF

[2269] Paul Y. H. Yiu. Sums of squares formulae with integer coefficients. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 30(3):318–324, September 1987. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

deSquire:1987:LCH

[2270] Maria L. Torres de Squire. Local complements to the Hausdorff–Young theorem for amalgams. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 30(3):325–333, September 1987. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Balder:1987:MFL

[2271] Erik J. Balder. More on Fatou’s lemma in several dimensions. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 30(3): 334–339, September 1987. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Childs:1987:NEA

[2272] Lindsay N. Childs. Nonisomorphic equivalent Azumaya algebras. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 30(3):340–343, September 1987. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Killam:1987:LDS

[2273] Eleanor Killam. Lie derivations on skew elements in prime rings with involution. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 30(3):344–350, September 1987. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Ginsburg:1987:NME

[2274] John Ginsburg. On the number of maximal elements in a partially ordered set. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 30(3):351–357, September 1987. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Hoffman:1987:FAA

[2275] Michael Hoffman. Free actions of abelian groups on a Cartesian power of an even sphere. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 30(3):358–362, September 1987. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Cohen:1987:SMA

[2276] Miriam Cohen. H-simple H-module algebras. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 30(3):363–366, September 1987. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Merris:1987:OIS

[2277] Russell Merris. Oppenheim’s inequality for the second immanant. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 30(3):367–369, September 1987. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Yamada:1987:GMS

[2278] Takeo Yamada. Generic matrix sign-stability. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 30(3):370–376, September 1987. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Loeb:1987:DTC

[2279] Jean-Jacques Loeb, Karl Oeljeklaus, and Wolfgang Richthofer. A decomposition theorem for complex nilmanifolds. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 30(3):377–378, September 1987. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Yamashita:1987:DLP

[2280] Shinji Yamashita. Derivatives and length-preserving maps. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 30(3): 379–384, September 1987. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Jech:1987:FOT

[2281] Thomas Jech. First order theory of complete Stonean algebras (Booleanvalued real and complex numbers). Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 30(4):385–392, December 1987. CO-DEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Feeman:1987:NSD

[2282] Timothy Feeman. A note on M-summands in dual spaces. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 30(4): 393–398, December 1987. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Martindale:1987:NNJ

[2283] Wallace S. Martindale III. A note on nilpotent Jordan rings. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 30(4): 399–401, December 1987. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Peschke:1987:SP

[2284] Georg Peschke. H-semidirect products. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 30(4):402–411, December 1987. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Murty:1987:NGS

[2285] M. Ram Murty and S. Srinivasan. On the number of groups of squarefree order. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 30(4):412–420, December 1987. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Hoft:1987:BSP

[2286] Hartmut Höft. Bound sets in partial orders and the fixed point property. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 30(4):421–428, December 1987. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Lines:1987:SPH

[2287] Daniel Lines. Stable plumbing for high odd-dimensional fibred knots. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 30(4):429–435, December 1987. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Barany:1987:ESE

[2288] Imre Bárány and Zoltán Füredi. Empty simplices in Euclidean space. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 30(4):436–445, December 1987. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Rhoades:1987:SWM

[2289] B. E. Rhoades. The spectrum of weighted mean operators. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 30(4): 446–449, December 1987. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Kulich:1987:ADI

[2290] James Kulich. The action of the Dickson invariants on length n Steenrod operations. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 30(4):450–454, December 1987. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Whitfield:1987:UMI

[2291] J. H. M. Whitfield and V. Zizler. Uniform Mazur’s intersection property of balls. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 30(4):455–460, December 1987. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Carbonne:1987:ABH

[2292] Philippe Carbonne. Anneaux de Bezout, Hermite et Kaplansky “universels”. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 30(4):461–470, December 1987. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Daniels:1987:BSM

[2293] P. Daniels and G. Gruenhage. Borel sets in metric spaces with small separable subsets. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 30(4):471–475, December 1987. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Rahman:1987:PRC

[2294] Q. I. Rahman and Q. M. Tariq. On a problem related to the conjecture of Sendov about the critical points of a polynomial. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 30(4):476–480, December 1987. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Lin:1987:UAR

[2295] Pei-Kee Lin. A uniformly asymptotically regular mapping without fixed points. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 30(4):481–483, December 1987. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Kerman:1988:COK

[2296] R. A. Kerman. Convolution with odd kernels having a tempered singularity. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 31(1):3–12, March 1988. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Touibi:1988:GBC

[2297] C. Touibi. Groupe de Brauer d’une courbe de Tate. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 31(1):13–18, March 1988. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Graham:1988:PSB

[2298] Colin C. Graham and Bertram M. Schreiber. Projections in spaces of bimeasures. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 31(1):19–25, March 1988. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Alufohai:1988:BML

[2299] Imohimi C. Alufohai. Bernoulli maps of a Lebesgue space. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 31(1): 26–31, March 1988. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Hajela:1988:KSO

[2300] D. Hajela, A. Pollington, and B. Smith. On Kloosterman sums with oscillating coefficients. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 31(1):32–36, March 1988. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Krystock:1988:SOF

[2301] Robert L. Krystock. On H-sets and open filter adherences. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 31(1): 37–44, March 1988. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Gilligan:1988:KPR

[2302] Bruce Gilligan. On the Kobayashi pseudometric reduction of homogeneous spaces. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 31(1):45–51, March 1988. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Freedman:1988:NSS

[2303] H. I. Freedman and K. Gopalsamy. Nonoccurrence of stability switching in systems with discrete delays. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 31(1):52–58, March 1988. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Burke:1988:RPD

[2304] Maxim R. Burke. Recouvrements ponctuellements dénombrables par des ensembles négligeables. (French) [Pointwise denumerable coverings by negligible sets]. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 31(1):59–62, March 1988. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Schmidt:1988:EPI

[2305] E. J. P. Georg Schmidt. On an extremal problem involving harmonic functions. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 31(1):63–69, March 1988. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Cambern:1988:SCV

[2306] Michael Cambern and Peter Greim. Spaces of continuous vector functions as duals. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 31(1):70–78, March 1988. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Eloe:1988:MFM

[2307] P. W. Eloe and P. L. Saintignon. A method of forced monotonicity for conjugate type boundary value problems for ordinary differential equations. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 31(1):79–84, March 1988. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Holub:1988:SO

[2308] J. R. Holub. On shift operators. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 31(1):85–94, March 1988. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

MacLeod:1988:GRL

[2309] R. A. MacLeod. Generalization of a result of E. Lucas. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 31(1):95–98, March 1988. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Bender:1988:NRC

[2310] Edward A. Bender and Nicholas C. Wormald. The number of rooted convex polyhedra. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 31(1):99–102, March 1988. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Bergen:1988:DIV

[2311] Jeffrey Bergen and L. Carini. Derivations with invertible values on a Lie ideal. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 31(1):103–110, March 1988. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Takahashi:1988:QAT

[2312] Katsutoshi Takahashi. On quasisimilarity for analytic Toeplitz operators. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 31(1):111–116, March 1988. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Izuchi:1988:FTP

[2313] Keiji Izuchi. A function theoretic proof of Axler’s zero multiplier theorem. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 31(1):117–120, March 1988. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Phelps:1988:CCS

[2314] R. R. Phelps. Counterexamples concerning support theorems for convex sets in Hilbert space. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 31(1):121–128, March 1988. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Salamon:1988:SPG

[2315] Peter Salamon and Paul Erdős. The solution to a problem of Grünbaum. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 31(2):129–138, June 1988. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Bastos:1988:TFA

[2316] Gervasio G. Bastos and T. M. Viswanathan. Torsion-free abelian groups, valuations and twisted group rings. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 31(2):139–146, June 1988. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Guadalupe:1988:CFP

[2317] Jose J. Guadalupe and Ma. Luisa Rezola. The conjugate function in plane curves. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 31(2):147–152, June 1988. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Bella:1988:CUS

[2318] A. Bella and F. Cammaroto. On the cardinality of Urysohn spaces. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 31(2):153–158, June 1988. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Adams:1988:RSI

[2319] R. A. Adams. Reduced Sobolev inequalities. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 31(2):159–167, June 1988. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Akkar:1988:CCA

[2320] M. Akkar, M. El Azhari, and M. Oudadess. Continuité des caractères dans les algèbres de Fréchet à bases. (French) [Continuity of characters in basic Fréchet algebras]. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 31(2):168–174, June 1988. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Jimenez:1988:FLR

[2321] J. Alfredo Jiménez. The focal locus of a Riemannian 4-symmetric space. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 31(2):175–181, June 1988. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Parameswaran:1988:SNP

[2322] Mangalam R. Parameswaran. Some new product theorems in summability. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 31(2):182–186, June 1988. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Fixman:1988:SMC

[2323] Uri Fixman. Sums of multinomial coefficients. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 31(2):187–189, June 1988. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Wu:1988:NTL

[2324] T. S. Wu. A note on a theorem on lattices in Lie groups. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 31(2): 190–193, June 1988. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Magalhaes:1988:SRC

[2325] L. Magalhães. Some results in the connective K-theory of Lie groups. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 31(2):194–199, June 1988. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Bonin:1988:CMC

[2326] Guy Bonin and Josée Legault. Comparaison de la méthode des constantes de Lyapunov et de la bifurcation de Hopf. (French) [Comparison of Lyapunov’s method of constants and Hopf bifurcation]. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 31(2):200– 209, June 1988. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Subbarao:1988:FTR

[2327] M. V. Subbarao and A. K. Agarwal. Further theorems of the Rogers– Ramanujan type theorems. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 31(2):210–214, June 1988. CODEN CM-BUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Rush:1988:ICT

[2328] David E. Rush. Integrally closed torsionless rings. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 31(2):215–216, June 1988. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Butler:1988:RTL

[2329] G. J. Butler, J. G. Timourian, and C. Viger. The rank theorem for locally Lipschitz continuous functions. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 31(2):217–226, June 1988. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Rowe:1988:N

[2330] K. A. Rowe. Nuclearity. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 31(2):227–235, June 1988. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Baboolal:1988:LCS

[2331] D. Baboolal. Local connectedness of the Stone–Čech compactification. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 31(2):236–240, June 1988. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Hansen:1988:SBV

[2332] Vagn Lundsgaard Hansen. On Steenrod bundles and the van Kampen theorem. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 31(2):241–249, June 1988. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Leep:1988:IAW

[2333] David Leep and Murray Marshall. Isomorphisms and automorphisms of Witt rings. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 31(2):250–256, June 1988. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Bender:1988:EMT

[2334] E. A. Bender, E. R. Canfield, and R. W. Robinson. The enumeration of maps on the torus and the projective plane. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 31(3):257–271, September 1988. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Kreith:1988:SGF

[2335] Kurt Kreith. Symmetric Green’s function for a class of CIV boundary value problems. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 31(3):272–279, September 1988. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Mauceri:1988:DLI

[2336] Silvana Mauceri and Paola Misso. Derivations on a Lie ideal. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 31(3): 280–286, September 1988. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Hebert:1988:CAC

[2337] Michel Hébert. Characterizations of axiomatic categories of models canonically isomorphic to (quasi-)varieties. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 31(3):287–300, September 1988. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Krishnamani:1988:QFL

[2338] Vatsala Krishnamani and Joseph Yucas. Quadratic forms and linkage of quaternion algebras. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 31(3):301–307, September 1988. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Tanahashi:1988:IPM

[2339] Kôtarô Tanahashi and Jun Tomiyama. Indecomposable positive maps in matrix algebras. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 31(3):308–317, September 1988. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Moron:1988:SCL

[2340] M. Alonso Moron. Some consequences of Lasnev’s theorem in shape theory. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 31(3):318–321, September 1988. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Leranoz:1988:RCC

[2341] Camino Leránoz. A remark on complex convexity. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 31(3):322–324, September 1988. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Przytycki:1988:PTL

[2342] Józef H. Przytycki. Plans’ theorem for links: an application of t_{k} moves. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 31(3):325–327, September 1988. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Groemer:1988:STC

[2343] H. Groemer. Stability theorems for convex domains of constant width. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 31(3):328–337, September 1988. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Sands:1988:KIV

[2344] Jonathan W. Sands. Kummer’s and Iwasawa’s version of Leopoldt’s conjecture. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 31(3):338–346, September 1988. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Hildebrand:1988:CPV

[2345] Adolf Hildebrand. On the constant in the Pólya–Vinogradov inequality. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 31(3):347–352, September 1988. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Loewen:1988:PSF

[2346] Philip D. Loewen. The proximal subgradient formula in Banach space. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 31(3):353–361, September 1988. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Eloe:1988:EDR

[2347] P. W. Eloe. Eventual disconjugacy and right disfocality of linear difference equations. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 31(3):362–373, September 1988. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Wolfson:1988:IPG

[2348] Kenneth G. Wolfson. Isomorphisms of prime Goldie semi-principal left ideal rings. II. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 31(3):374–379, September 1988. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Oberlin:1988:MYI

[2349] Daniel M. Oberlin. A multilinear Young’s inequality. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 31(3):380–384, September 1988. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Levy:1988:SS

[2350] Ronnie Levy. Semi-F -spaces. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 31(4):385–393, December 1988. CO-DEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Bongiovanni:1988:CPR

[2351] G. Bongiovanni, D. P. Bovet, and A. Cerioli. Comments on a paper of R. A. Brualdi: “Notes on the Birkhoff algorithm for doubly stochastic matrices”. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 31(4):394–398, December 1988. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic). See [1867].

Brunker:1988:PDG

[2352] Don Brunker and Denis Higgs. Primary decomposition for Σ-groups. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 31(4):399–403, December 1988. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Nakazi:1988:NIB

[2353] Takahiko Nakazi. Notes on interpolation by bounded analytic functions. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 31(4):404–408, December 1988. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Parker:1988:PGM

[2354] E. T. Parker and Lawrence Somer. A partial generalization of Mann’s theorem concerning orthogonal Latin squares. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 31(4):409–413, December 1988. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Stempak:1988:NPW

[2355] Krzysztof Stempak. A new proof of a Watson’s formula. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 31(4):414– 418, December 1988. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Chandler:1988:GFR

[2356] K. A. Chandler. Groups formed by redefining multiplication. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 31(4): 419–423, December 1988. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Oprea:1988:DLF

[2357] John Oprea. Decompositions of localized fibres and cofibres. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 31(4): 424–431, December 1988. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Krall:1988:SES

[2358] Allan M. Krall. Semibounded extensions of singular ordinary differential operators. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 31(4):432–438, December 1988. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Putnam:1988:SME

[2359] Ian F. Putnam. Strong Morita equivalence for the Denjoy C^{∗}-algebras. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 31(4):439–447, December 1988. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Kalton:1988:SIB

[2360] N. J. Kalton. Sums of idempotents in Banach algebras. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 31(4):448– 451, December 1988. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Wojtkowiak:1988:RMB

[2361] Zdzisław Wojtkowiak. A remark on maps between classifying spaces of compact Lie groups. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 31(4):452–458, December 1988. CODEN CM-BUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Grove:1988:SCO

[2362] E. A. Grove, G. Ladas, and J. Schinas. Sufficient conditions for the oscillation of delay and neutral delay equations. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 31(4):459–466, December 1988. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Rahman:1988:SEA

[2363] Mizan Rahman. Some extensions of Askey–Wilson’s Q-beta integral and the corresponding orthogonal systems. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 31(4):467–476, December 1988. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Buskes:1988:LLP

[2364] G. Buskes and I. Labuda. On levi-like properties and some of their applications in Riesz space theory. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 31(4):477–486, December 1988. CO-DEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Gardner:1988:IWI

[2365] B. J. Gardner and P. N. Stewart. Injective and weakly injective rings. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 31(4):487–494, December 1988. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Frih:1988:AFD

[2366] El Mostapha Frih and Paul M. Gauthier. Approximation of a function and its derivatives by entire functions of several variables. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 31(4): 495–499, December 1988. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Bell:1988:SCP

[2367] H. E. Bell and W. S. Martindale III. Semiderivations and commutativity in prime rings. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 31(4):500–508, December 1988. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Robertson:1989:EPG

[2368] E. F. Robertson and P. D. Williams. Efficient presentations of the groups PSL(2, 2p) and SL(2, 2p). Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 32(1):3–10, March 1989. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Allenby:1989:RFC

[2369] R. B. J. T. Allenby and C. Y. Tang. On the residual finiteness of certain polygonal products. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 32(1):11–17, March 1989. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Jevtic:1989:NBP

[2370] Miroljub Jevtić. A note on Blaschke products with zeroes in a nontangential region. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 32(1):18–23, March 1989. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Kucharz:1989:NDE

[2371] Wojciech Kucharz. A note on the Dubois–Efroymson dimension theorem. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 32(1):24–29, March 1989. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Ross:1989:TSE

[2372] David Ross. Tree self-embeddings. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 32(1):30–33, March 1989. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Edelson:1989:APS

[2373] Allan L. Edelson. Asymptotic properties of semilinear equations. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 32(1): 34–46, March 1989. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Siddiqi:1989:NIG

[2374] Jamil A. Siddiqi and Abdelkader El Koutri. Norm inequalities for generators of analytic semigroups and cosine operator functions. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 32(1): 47–53, March 1989. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Ljubic:1989:BAG

[2375] Dragoslav Ljubić. Bireflectionality in absolute geometry. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 32(1):54–63, March 1989. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Krengel:1989:DAA

[2376] Ulrich Krengel and Michael Lin. On the deterministic and asymptotic σ-algebras of a Markov operator. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 32(1):64–73, March 1989. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Rao:1989:NP

[2377] T. S. S. R. K. Rao. A note on the R_{n,k} property for L^{1}(μ, E). Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 32(1): 74–77, March 1989. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Kent:1989:CCC

[2378] D. C. Kent and G. D. Richardson. Cauchy completion categories. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 32(1): 78–84, March 1989. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Adjaero:1989:IGR

[2379] Isabelle Adjaero and Eugene Spiegel. Isomorphic group rings over domains. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 32(1):85–89, March 1989. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Ishihara:1989:FPT

[2380] Hajime Ishihara. Fixed point theorems for Lipschitzian semigroups. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 32 (1):90–97, March 1989. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Cambern:1989:UPS

[2381] Michael Cambern and Peter Greim. Uniqueness of preduals for spaces of continuous vector functions. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 32(1):98–104, March 1989. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Abbott:1989:DS

[2382] H. L. Abbott and M. V. Subbarao. On the distribution of the sequence nd^{∗}(n). Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 32(1):105–108, March 1989. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Mason:1989:NNS

[2383] A. W. Mason. Nonstandard, normal subgroups and nonnormal, standard subgroups of the modular group. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 32(1):109–113, March 1989. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Gilmer:1989:PID

[2384] Robert Gilmer and William Heinzer. On Pic(D[α]) for a principal ideal domain D. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 32(1):114–116, March 1989. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Comez:1989:SET

[2385] Dogan Çömez. A stochastic ergodic theorem for general additive processes. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 32(1):117–121, March 1989. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Haydon:1989:NSN

[2386] Richard Haydon and Václav Zizler. A new space with no locally uniformly rotund renorming. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 32(1):122–128, March 1989. CODEN CM-BUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Chipot:1989:PFP

[2387] M. Chipot and V. Oliker. Sur une propriété des fonctions propres de l’opérateur de Laplace–Beltrami. (French) [on a property of eigenfunctions of the Laplace–Beltrami operator]. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 32(2):129–139, June 1989. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Asmar:1989:CFF

[2388] Nakhlé Asmar. The conjugate function on the finite-dimensional torus. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 32(2):140–148, June 1989. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Grubb:1989:SCD

[2389] D. J. Grubb. U-sets in compact, 0-dimensional, metric groups. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 32(2): 149–155, June 1989. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Ortel:1989:PGG

[2390] M. Ortel and W. Schneider. A parametric Gauss–Green theorem in several variables. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 32(2):156–165, June 1989. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Gottlieb:1989:CND

[2391] Christian Gottlieb. On condensed Noetherian domains whose integral closures are discrete valuation rings. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 32(2):166–168, June 1989. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Diaz:1989:MSC

[2392] J. C. Díaz. Montel subspaces in the countable projective limits of L^{P} (μ)- spaces. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 32(2):169–176, June 1989. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Masson:1989:SCF

[2393] David R. Masson. Some continued fractions of Ramanujan and Meixner– Pollaczek polynomials. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 32(2):177–181, June 1989. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Fox:1989:ADB

[2394] Thomas F. Fox. Algebraic deformations and bicohomology. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 32(2): 182–189, June 1989. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Rohl:1989:RLS

[2395] Frank Röhl. A remark on the Loewy-series of certain Hopf algebras. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 32(2):190–193, June 1989. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Connor:1989:SMS

[2396] Jeff Connor. On strong matrix summability with respect to a modulus and statistical convergence. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 32(2):194–198, June 1989. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Alzer:1989:WGM

[2397] Horst Alzer. On weighted geometric means. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 32(2):199–206, June 1989. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Baker:1989:EUM

[2398] B. J. Baker and Michael Laidacker. Embedding uncountably many mutually exclusive continua into Euclidean space. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 32(2):207–214, June 1989. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Simons:1989:TDG

[2399] Lloyd D. Simons. Z_{p}-towers in Demuskin groups. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 32(2):215–222, June 1989. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Wajngurt:1989:SSD

[2400] Clara Wajngurt. Solutions of specific Diophantine equations and their relationship to complex multiplication. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 32(2):223–229, June 1989. CO-DEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Facchini:1989:SDM

[2401] Alberto Facchini. Simple divisible modules over integral domains. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 32 (2):230–240, June 1989. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Barr:1989:EAR

[2402] Michael Barr. Embedding of accessible regular categories. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 32(2):241– 247, June 1989. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Kolster:1989:RBP

[2403] Manfred Kolster. A relation between the 2-primary parts of the main conjecture and the Birch–Tate-conjecture. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 32(2):248–251, June 1989. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Brondsted:1989:NPC

[2404] A. Brøndsted and G. Maxwell. A new proof of the d-connectedness of d-polytopes. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 32(2):252–254, June 1989. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Sands:1989:IR

[2405] A. D. Sands. On invariant radicals. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 32(2):255–256, June 1989. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Lambek:1989:GM

[2406] J. Lambek. Grammar as mathematics. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 32(3):257–273, September 1989. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Hare:1989:DCB

[2407] D. E. G. Hare. A dual characterization of Banach spaces with the convex point-of-continuity property. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 32(3):274–280, September 1989. CO-DEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Asperti:1989:CFR

[2408] A. C. Asperti and M. Dajczer. Conformally flat Riemannian manifolds as hypersurfaces of the light cone. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 32(3):281–285, September 1989. CO-DEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Herron:1989:PHF

[2409] David A. Herron and Joel L. Schiff. Positive harmonic functions and complete metrics. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 32(3):286–297, September 1989. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Allaway:1989:COJ

[2410] Wm. R. Allaway. Convolution orthogonality and the Jacobi polynomials. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 32(3):298–308, September 1989. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Jipsen:1989:ARA

[2411] Peter Jipsen and Henry Rose. Absolute retracts and amalgamation in certain congruence distributive varieties. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 32(3):309–313, September 1989. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Seibt:1989:DCD

[2412] Peter Seibt. Differential completions and differentially simple algebras. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 32(3):314–319, September 1989. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Herrero:1989:SFO

[2413] Domingo A. Herrero. The set of finite operators is nowhere dense. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 32(3):320–326, September 1989. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Agarwal:1989:BPS

[2414] A. K. Agarwal. Bijective proofs of some n-color partition identities. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 32(3):327–332, September 1989. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Birkenmeier:1989:DRG

[2415] Gary F. Birkenmeier. A decomposition of rings generated by faithful cyclic modules. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 32(3):333–339, September 1989. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Mercier:1989:NCB

[2416] Armel Mercier. Une note sur les coefficients binomiaux de Gauss. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 32(3): 340–343, September 1989. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Khamsi:1989:NSB

[2417] M. A. Khamsi. Normal structure for Banach spaces with Schauder decomposition. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 32(3):344–351, September 1989. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Foguel:1989:IVN

[2418] S. R. Foguel. Invariant vectors of a nonrecurrent Markov matrix. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 32(3): 352–359, September 1989. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Akcoglu:1989:ACI

[2419] M. A. Akcoglu and D. Boivin. Approximation of L_{p}-contractions by isometries. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 32(3):360–364, September 1989. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Doman:1989:MSW

[2420] Ryszard Doman. Moore G-spaces which are not co-Hopf G-spaces. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 32(3):365–368, September 1989. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Rahman:1989:NOJ

[2421] Mizan Rahman. A note on the orthogonality of Jackson’s q-Bessel functions. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 32(3):369–376, September 1989. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Dubuc:1989:TFS

[2422] Benoit Dubuc. On Takagi fractal surfaces. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 32(3):377–384, September 1989. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Guillory:1989:CSB

[2423] Carroll Guillory. A characterization of a sparse Blaschke product. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 32(4):385–390, December 1989. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Amberg:1989:NNF

[2424] Bernhard Amberg, Silvana Franciosi, and Francesco de Giovanni. Nilpotent-by-Noetherian factorized groups. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 32(4):391–403, December 1989. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Ronveaux:1989:DEC

[2425] A. Ronveaux and F. Marcellán. Differential equation for classical-type orthogonal polynomials. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 32(4):404–411, December 1989. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Hurrelbrink:1989:APG

[2426] Jürgen Hurrelbrink. Annihilating polynomials for group rings and Witt rings. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 32(4):412–416, December 1989. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Baggs:1989:RSW

[2427] Ivan Baggs. A regular space on which every real-valued function with a closed graph is constant. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 32(4):417–424, December 1989. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Porter:1989:KS

[2428] Jack Porter and Mohan Tikoo. On Katetov spaces. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 32(4):425–433, December 1989. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Cho:1989:NIC

[2429] Chong-Man Cho. A note on M-ideals of compact operators. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 32(4): 434–440, December 1989. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Babu:1989:NDF

[2430] Gutti Jogesh Babu and Paul Erdős. A note on the distribution function of additive arithmetical functions in short intervals. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 32(4):441–445, December 1989. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Nash:1989:S

[2431] John C. M. Nash. On B_{4}-sequences. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 32(4):446–449, December 1989. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Kiyosawa:1989:SCO

[2432] Takemitsu Kiyosawa. On spaces of compact operators in non-Archimedean Banach spaces. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 32(4):450–458, December 1989. CO-DEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Miller:1989:GTC

[2433] Andy Miller. A group theoretic characterization of the 2-dimensional spherical groups. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 32(4):459–466, December 1989. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Haukkanen:1989:DCA

[2434] Pentti Haukkanen. On the Davison convolution of arithmetical functions. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 32(4):467–473, December 1989. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Mollin:1989:QNP

[2435] R. A. Mollin and H. C. Williams. Quadratic nonresidues and primeproducing polynomials. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 32(4):474–478, December 1989. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Borwein:1989:ITD

[2436] David Borwein. An inclusion theorem for Dirichlet series. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 32(4):479– 481, December 1989. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Kuo:1989:BWR

[2437] Tzee Char Kuo, David J. A. Trotman, and Pei Xin Li. Blowing-up and Whitney (a)-regularity. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 32(4):482–485, December 1989. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Hoechsmann:1989:GLK

[2438] Klaus Hoechsmann. Généralisation d’un lemme de Kummer. (French) [generalization of a lemma of Kummer]. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 32(4):486–489, December 1989. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Mathieu:1989:PPT

[2439] Martin Mathieu. Properties of the product of two derivations of a C^{∗}- algebra. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 32(4):490–497, December 1989. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Stewart:1989:PS

[2440] Patrick N. Stewart. Projective socles. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 32(4):498–499, December 1989. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Rhoades:1989:NSG

[2441] B. E. Rhoades. l^{p}-norms of some generalized Hausdorff matrices. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 32(4): 500–504, December 1989. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic). See corrections [2584].

Wehrfritz:1989:SLG

[2442] B. A. F. Wehrfritz and M. Shirvani. On skew linear groups associated with certain soluble-by-finite groups. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 32(4):505–508, December 1989. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Avrin:1990:SSI

[2443] Joel D. Avrin. Solutions with singular initial data for a model of electrophoretic separation. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 33(1):3–10, March 1990. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Benabdallah:1990:SGP

[2444] K. Benabdallah and C. Piché. Sur les sous-groupes purifiables des groupes abéliens primaires. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 33(1):11–17, March 1990. CODEN CM-BUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Sember:1990:FSF

[2445] John Sember. Families of sequences of 0’s and 1’s in FK-spaces. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 33(1): 18–23, March 1990. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Wang:1990:DRV

[2446] Y. H. Wang. Dependent random variables with independent subsets. II. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 33(1):24–28, March 1990. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Zhang:1990:NHO

[2447] Jian Zhong Zhang. A note on Hardy–Orlicz spaces. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 33(1):29–33, March 1990. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Asmar:1990:EEC

[2448] Nakhlé Habib Asmar. Exponential estimates for the conjugate function on locally compact abelian groups. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 33(1):34–44, March 1990. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Sun:1990:BSW

[2449] Yeneng Sun. A Banach space in which a ball is contained in the range of some countably additive measure is superreflexive. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 33(1):45–49, March 1990. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Funakura:1990:CAF

[2450] Takeo Funakura. On coefficients of Artin L functions as Dirichlet series. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 33(1):50–54, March 1990. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Guinez:1990:LCC

[2451] Víctor Guíñez, Eduardo Sáez, and Iván Szántó. Limit cycles close to infinity of certain nonlinear differential equations. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 33(1):55–59, March 1990. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Davis:1990:SBH

[2452] Chandler Davis and Pratibha Ghatage. On the spectrum of the Bergman–Hilbert matrix. II. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 33(1):60–64, March 1990. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Garay:1990:PTS

[2453] Oscar J. Garay. Pseudoumbilical 2-type surfaces in spheres. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 33(1):

65–68, March 1990. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Nandakumar:1990:RDF

[2454] N. R. Nandakumar. Ring derivations on function algebras. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 33(1): 69–72, March 1990. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Goodaire:1990:SSC

[2455] Edgar G. Goodaire and D. A. Robinson. Some special conjugacy closed loops. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 33(1):73–78, March 1990. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Deveney:1990:FF

[2456] James K. Deveney and Joe Yanik. Factors of fields. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 33(1):79–83, March 1990. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Li:1990:SPU

[2457] Chengzhi Li, Christiane Rousseau, and Xian Wang. A simple proof for the unicity of the limit cycle in the Bogdanov–Takens system. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 33(1): 84–92, March 1990. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Okoh:1990:EST

[2458] F. Okoh and F. Zorzitto. Extension that are submodules of their quotients. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 33(1):93–99, March 1990. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Baribeau:1990:FAM

[2459] Line Baribeau. Sur les fonctions analytiques multiformes dont les valeurs sont des segments. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 33(1):100–105, March 1990. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Mott:1990:UHD

[2460] J. L. Mott and M. Zafrullah. Unruly Hilbert domains. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 33(1):106– 109, March 1990. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Peschke:1990:ESR

[2461] Georg Peschke. Ends of spaces related by a covering map. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 33(1): 110–118, March 1990. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Porta:1990:HSM

[2462] Horacio Porta and Kenneth B. Stolarsky. Half-silvered mirrors and Wythoff’s game. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 33(1):119–125, March 1990. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Lanski:1990:CSP

[2463] Charles Lanski. Can a semi-prime ring be a finite union of right annihilators? Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 33(1):126–128, March 1990. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Varbanec:1990:DIA

[2464] P. D. Varbanec and P. Zarzycki. Divisors of integers in arithmetic progression. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 33(2):129–134, June 1990. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Ouertani:1990:EDK

[2465] Ahmed Ouertani. Exemples de dimension de Krull d’anneaux de polynômes. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 33(2):135–138, June 1990. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Jarosz:1990:ACS

[2466] Krzysztof Jarosz. Automatic continuity of separating linear isomorphisms. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 33(2):139–144, June 1990. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Wingberg:1990:ETE

[2467] Kay Wingberg. On the étale K-theory of an elliptic curve with complex multiplication for regular primes. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 33(2):145–150, June 1990. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Wojtaszczyk:1990:MBS

[2468] P. Wojtaszczyk. On multipliers into Bergman spaces and Nevanlinna class. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 33(2):151–161, June 1990. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Bashir:1990:TRS

[2469] M. A. Bashir. On totally real submanifolds in a 6-sphere. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 33(2):162– 166, June 1990. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic). See retraction [2569].

Kwon:1990:IAW

[2470] E. G. Kwon. Image area and the weighted subspaces of Hardy spaces. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 33(2):167–174, June 1990. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Kummer:1990:RB

[2471] Martin Kummer, Richard C. Churchill, and David L. Rod. On a result of Bruns. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 33(2):175–180, June 1990. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Jannsen:1990:SHS

[2472] Uwe Jannsen. The splitting of the Hochschild–Serre spectral sequence for a product of groups. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 33(2):181–183, June 1990. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Little:1990:DWP

[2473] John B. Little. Distribution of Weierstrass points on rational cuspidal curves. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 33(2):184–189, June 1990. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Comerford:1990:PTC

[2474] Jonell A. Comerford and Y. Lee. Product of two commutators as a square in a free group. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 33(2):190–196, June 1990. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Lubinsky:1990:CBL

[2475] D. S. Lubinsky and Z. Ziegler. Coefficient bounds in the Lorentz representation of a polynomial. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 33(2):197–206, June 1990. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Morales:1990:EWG

[2476] Jorge F. Morales. Equivariant Witt groups. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 33(2):207–218, June 1990. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Oprea:1990:HCR

[2477] John Oprea. A homotopical Conner–Raymond theorem and a question of Gottlieb. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 33(2):219–229, June 1990. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Cohen:1990:PCC

[2478] Stephen D. Cohen. Proof of a conjecture of Chowla and Zassenhaus on permutation polynomials. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 33(2):230–234, June 1990. CODEN CM-BUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Yokota:1990:NIR

[2479] Hisashi Yokota. On number of integers representable as sums of unit fractions. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 33(2):235–241, June 1990. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Goodaire:1990:DUS

[2480] E. G. Goodaire, E. Jespers, and M. M. Parmenter. Determining units in some integral group rings. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 33(2):242–246, June 1990. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Aharonov:1990:HPL

[2481] Dov Aharonov. The hexagonal packing lemma and discrete potential theory. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 33(2):247–252, June 1990. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Moser:1990:MGT

[2482] W. O. J. Moser. A (modest) generalization of the theorems of Wilson and Fermat. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 33(2):253–256, June 1990. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Zaharopol:1990:PCE

[2483] Radu Zaharopol. On products of conditional expectation operators. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 33(3):257–260, September 1990. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Esterle:1990:NCE

[2484] J. Esterle and F. Zarouf. Sur la non compatibilité de l’équivalence de Calkin avec le calcul fonctionnel de Nagy–Foias. (French) [On the noncompatibility of Calkin’s equivalence with Nagy–Foias functional calculus]. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 33(3):261–263, September 1990. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Mingo:1990:N

[2485] James A. Mingo. On the nonisomorphism of C_{λ}^{∗}(G) ⊗_{max} C_{h}^{∗}o(G) and C^{∗}(C_{λ}^{∗}(G), C_{h}^{∗}o(G)). Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 33(3):264–267, September 1990. CODEN CM-BUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Naimpally:1990:NC

[2486] S. A. Naimpally and Mohan L. Tikoo. Nearness convergence. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 33(3): 268–272, September 1990. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Bullen:1990:IRS

[2487] P. S. Bullen and S. N. Mukhopadhyay. The integrability of Riemann summable trigonometric series. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 33(3):273–281, September 1990. CO-DEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Pacheco:1990:NRB

[2488] Amílcar Pacheco. A note on relations between the zeta-functions of Galois coverings of curves over finite fields. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 33(3):282–285, September 1990. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Eriksson-Bique:1990:DTP

[2489] Sirkka-Liisa Eriksson-Bique. A decomposition theorem for positive superharmonic functions. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 33(3):286–296, September 1990. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Wildberger:1990:RBA

[2490] N. J. Wildberger. On a relationship between adjoint orbits and conjugacy classes of a Lie group. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 33(3):297–304, September 1990. CO-DEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Kruskemper:1990:IPQ

[2491] Martin Krüskemper. Indecomposable positive quadratic forms. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 33(3): 305–310, September 1990. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Lewis:1990:SCI

[2492] R. P. Lewis. On some combinatorial interpretations of Slater’s identities. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 33(3):311–313, September 1990. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Mukherjee:1990:SGT

[2493] N. P. Mukherjee and Prabir Bhattacharya. On supersolvable groups and a theorem of Huppert. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 33(3):314–315, September 1990. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Rouhani:1990:FTT

[2494] H. Rouhani. A Furstenberg transformation of the 2-torus without quasidiscrete spectrum. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 33(3):316–322, September 1990. CODEN CM-BUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Gyori:1990:OTD

[2495] I. Győri, G. Ladas, and L. Pakula. Oscillation theorems for delay differential equations via Laplace transforms. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 33(3):323–326, September 1990. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Lucas:1990:VRI

[2496] Thomas G. Lucas. Valuation rings and integral closure. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 33(3):327–330, September 1990. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Whitley:1990:FCT

[2497] Robert Whitley. Fuglede’s commutativity theorem and ⋂ R(T −λ). Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 33(3):331–334, September 1990. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Abbott:1990:SS

[2498] H. L. Abbott. Sidon sets. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 33(3):335–341, September 1990. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Edmunds:1990:PGT

[2499] Charles C. Edmunds and Gerhard Rosenberger. Powers of genus two in free groups. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 33(3):342–344, September 1990. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Pedersen:1990:PTU

[2500] Steen Pedersen. The product of two (unbounded) derivations. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 33(3): 345–348, September 1990. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Malle:1990:HG

[2501] Gunter Malle. Hurwitz groups and G_{2}(q). Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 33(3):349–357, September 1990. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Schmuland:1990:SPP

[2502] B. Schmuland. Sample path properties of l^{p}-valued Ornstein–Uhlenbeck processes. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 33(3):358–368, September 1990. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Herz:1990:RLG

[2503] Carl Herz. Representations of Lie groups by contact transformations. I. Compact groups. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 33(4):369–375, December 1990. CODEN CM-BUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Bean:1990:SFA

[2504] Michael Bean and John A. Baker. The stability of a functional analogue of the wave equation. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 33(4):376–385, December 1990. CODEN CM-BUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Yang:1990:CET

[2505] Yi Song Yang. A comparison of eigenvalues of two Sturm–Liouville problems. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 33(4):386–390, December 1990. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Campbell:1990:VIS

[2506] H. E. A. Campbell, I. Hughes, and R. D. Pollack. Vector invariants of symmetric groups. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 33(4):391–397, December 1990. CODEN CM-BUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Macoosh:1990:TIC

[2507] R. Macoosh and R. Raphael. Totally integrally closed Azumaya algebras. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 33(4):398–403, December 1990. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Exel:1990:HMR

[2508] Ruy Exel. Hankel matrices over right ordered amenable groups. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 33(4): 404–415, December 1990. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Babu:1990:NDT

[2509] Gutti Jogesh Babu. A note on Dubins’ theorem. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 33(4):416–418, December 1990. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Brown:1990:IED

[2510] Leon Brown. Invertible elements in the Dirichlet space. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 33(4):419– 422, December 1990. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Enomoto:1990:UMN

[2511] Masatoshi Enomoto, Marie Choda, and Yasuo Watatani. Uncountably many nonbinary shifts on the hyperfinite II_{1}-factor. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 33(4):423–427, December 1990. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Feldstein:1990:UNF

[2512] Alan Feldstein and Zdzisław Jackiewicz. Unstable neutral functionaldifferential equations. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 33(4):428–433, December 1990. CODEN CM-BUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Anantharaman-Delaroche:1990:CPM

[2513] C. Anantharaman-Delaroche. On completely positive maps defined by an irreducible correspondence. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 33(4):434–441, December 1990. CO-DEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Ladas:1990:ODE

[2514] G. Ladas and C. Qian. Oscillation in differential equations with positive and negative coefficients. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 33(4):442–451, December 1990. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Granville:1990:NSP

[2515] Andrew Granville. A note on sums of primes. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 33(4):452–454, December 1990. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Kusuda:1990:CIA

[2516] Masaharu Kusuda. A characterization of ideals of C^{∗}-algebras. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 33(4):

455–459, December 1990. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

DeVore:1990:ART

[2517] Ronald A. DeVore, D. Leviatan, and Xiang Ming Yu. L_{p} approximation by reciprocals of trigonometric and algebraic polynomials. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 33(4): 460–469, December 1990. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Saxe:1990:ISR

[2518] Karen Saxe. Interpolation and spectra of regular L^{p}-space operators. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 33(4):470–481, December 1990. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Chi:1990:ACW

[2519] D. P. Chi and Y. D. Yoon. On an affine connection which admits a volume-like form. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 33(4):482–488, December 1990. CODEN CM-BUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Davidson:1990:RBD

[2520] Kenneth R. Davidson. The relationship between distance formulae and compact perturbations for reflexive algebras. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 33(4):489–493, December 1990. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Aebischer:1990:LBS

[2521] Beat Aebischer. The limiting behavior of sequences of quasiconformal mappings. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 33(4):494–502, December 1990. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

McCool:1990:CNG

[2522] James McCool. The coherence number of 2-groups. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 33(4):503–508, December 1990. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Adachi:1991:SMF

[2523] Toshiaki Adachi. Spherical mean and the fundamental group. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 34(1):3– 11, March 1991. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Anderson:1991:CMN

[2524] C. L. Anderson. A characterization of the Minkowski norms. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 34(1): 12–14, March 1991. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Anderson:1991:PI

[2525] David F. Anderson, Evan G. Houston, and Muhammad Zafrullah. Pseudo-integrality. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 34(1):15–22, March 1991. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Arpin:1991:BTC

[2526] Peter Arpin and John Ginsburg. Binary trees and the n-cutset property. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 34(1):23–30, March 1991. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Bhandari:1991:ITU

[2527] A. K. Bhandari and S. K. Sehgal. An induction theorem for units of p-adic group rings. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 34(1):31–35, March 1991. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Brown:1991:CQI

[2528] Tom C. Brown. A characterization of the quadratic irrationals. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 34(1): 36–41, March 1991. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Campbell:1991:RIS

[2529] H. E. A. Campbell, I. Hughes, and R. D. Pollack. Rings of invariants and p-sylow subgroups. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 34(1):42–47, March 1991. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Dajani:1991:S

[2530] Karma Dajani. On superrecurrence. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 34(1):48–57, March 1991. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Evans:1991:EAC

[2531] Anthony B. Evans. On elementary abelian Cartesian groups. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 34(1): 58–59, March 1991. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Ewell:1991:VRZ

[2532] John A. Ewell. On values of the Riemann zeta function at integral arguments. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 34(1):60–66, March 1991. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Gromadzki:1991:SGA

[2533] Grzegorz Gromadzki. On soluble groups of automorphism of Riemann surfaces. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 34(1):67–73, March 1991. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Maleev:1991:SNO

[2534] R. P. Maleev and S. L. Troyanski. Smooth norms in Orlicz spaces. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 34 (1):74–82, March 1991. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Mirowicz:1991:UGR

[2535] Maciej Mirowicz. Units in group rings of the infinite dihedral group. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 34(1):83–89, March 1991. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Noakes:1991:GUE

[2536] J. L. Noakes. Generalised umbilics on embedded spheres. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 34(1):90–95, March 1991. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Noda:1991:GFP

[2537] R. Noda, T. Sakai, and M. Morimoto. Generalized Fermat’s problem. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 34(1):96–104, March 1991. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Olsen:1991:TCB

[2538] James H. Olsen. Two consequences of Brunel’s theorem. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 34(1):105– 108, March 1991. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Page:1991:KLF

[2539] Frank H. Page, Jr. Komlos limits and Fatou’s lemma in several dimensions. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 34(1):109–112, March 1991. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Rupp:1991:SRE

[2540] Rudolf Rupp. Stable rank and the \overline{∂}-equation. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 34(1):113–118, March 1991. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Sankaran:1991:DGM

[2541] Parameswaran Sankaran. Determination of Grassmann manifolds which are boundaries. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 34(1):119–122, March 1991. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Shieh:1991:CLN

[2542] Narn-Rueih Shieh. Conditional local nondeterminism and Hausdorff measure of level sets. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 34(1):123–127, March 1991. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Wang:1991:URS

[2543] Ting Fu Wang and Shu Tao Chen. K-uniform rotundity of sequence Orlicz spaces. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 34(1):128–135, March 1991. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Wurfel:1991:DPE

[2544] Tilmann Würfel. Dimension-preserving extensions of pro-p-groups. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 34 (1):136–140, March 1991. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Zhang:1991:USE

[2545] Liang Cheng Zhang and Jonathan Gordon. On unit solutions of the equation xyz = x + y + z in not totally real cubic fields. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 34(1):141– 144, March 1991. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Asthagiri:1991:WCC

[2546] Rajappa K. Asthagiri. Weak continuity of a composition map between spaces of compact operators and Banach valued continuous functions. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 34(2):145–146, June 1991. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Baran:1991:KME

[2547] Mirosław Baran. A Karamata method. I. Elementary properties and applications. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 34(2):147–157, June 1991. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Beattie:1991:CGV

[2548] M. A. Beattie, S.-X. Liu, and P. N. Stewart. Comparing graded versions of the prime radical. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 34(2):158–164, June 1991. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Bullen:1991:SAT

[2549] P. S. Bullen and R. Výborný. Some applications of a theorem of Marcinkiewicz. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 34(2):165–174, June 1991. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Chu:1991:KKT

[2550] Ch’uan I. Chu. On the Kuiper–Kuo theorem. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 34(2):175–180, June 1991. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Clark:1991:AFA

[2551] David Clark. An arithmetical function associated with the rank of elliptic curves. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 34(2):181–185, June 1991. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Dobrowolski:1991:MMP

[2552] Edward Dobrowolski. Mahler’s measure of a polynomial in function of the number of its coefficients. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 34(2):186–195, June 1991. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Dummit:1991:PDT

[2553] D. S. Dummit. The parity distribution of traces in imaginary quadratic fields. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 34(2):196–201, June 1991. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Dyn:1991:DGP

[2554] N. Dyn, D. S. Lubinsky, and Boris Shekhtman. On density of generalized polynomials. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 34(2):202–207, June 1991. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Feeman:1991:NAO

[2555] Timothy G. Feeman. Nest algebras of operators and the Dunford–Pettis property. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 34(2):208–214, June 1991. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Frih:1991:AHG

[2556] El Mostapha Frih. Approximation holomorphe globale sur l’union de deux sous-espaces totalement réels maximaux de fC^{n}. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 34(2):215– 219, June 1991. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Frih:1991:GHA

[2557] El Mostapha Frih and Paul M. Gauthier. Global holomorphic approximation on the product of curves in C^{n}. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 34(2):220–223, June 1991. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Harris:1991:EPF

[2558] Morton E. Harris. On extending projectives of finite group-graded algebras. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 34(2):224–228, June 1991. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Koblitz:1991:JSI

[2559] Neal Koblitz. Jacobi sums, irreducible zeta-polynomials, and cryptography. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 34(2):229–235, June 1991. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Lanteri:1991:VLS

[2560] Antonio Lanteri. Verticality of (−1)-lines in scrolls over smooth surfaces. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 34(2):236–239, June 1991. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Lowen-Colebunders:1991:SCD

[2561] E. Lowen-Colebunders and Z. G. Szabó. Simplicity of categories defined by symmetry axioms. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 34(2):240–248, June 1991. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Nakazi:1991:AVT

[2562] Takahiko Nakazi. Absolute values of Toeplitz operators and Hankel operators. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 34(2):249–253, June 1991. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Qian:1991:SCN

[2563] Chuan Xi Qian. On some conjectures on neutral differential equations. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 34(2):254–259, June 1991. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Radjabalipour:1991:STA

[2564] M. Radjabalipour. Simultaneous triangularization of algebras of polynomially compact operators. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 34(2):260–264, June 1991. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Sherk:1991:PAT

[2565] F. A. Sherk. The projective antecedent of the three reflection theorem. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 34(2):265–274, June 1991. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Wiedemann:1991:SDM

[2566] Alfred Wiedemann. On stratifications of derived module categories. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 34 (2):275–280, June 1991. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Xie:1991:RPG

[2567] T. F. Xie and S. P. Zhou. A remark on a paper of G. G. Walter and A. I. Zayed: “The continuous (α, η)-Jacobi transform and its inverse when α + η + 1 is a positive integer”. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 34(2):281–288, June 1991. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Albar:1991:MGD

[2568] Muhammad A. Albar and Abdul-Aziz A. Al-Shuaibi. On mennicke groups of deficiency zero. II. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 34(3):289–293, September 1991. CO-DEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Bashir:1991:NRT

[2569] M. A. Bashir. Notice of retraction: “On totally real submanifolds in a 6sphere”. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 34(3):294, September 1991. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic). See [2469].

Bell:1991:REC

[2570] H. E. Bell and A. A. Klein. On rings with Engel cycles. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 34(3):295–300, September 1991. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Bonet:1991:QVQ

[2571] José Bonet. A question of Valdivia on quasinormable Fréchet spaces. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 34(3):301–304, September 1991. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Borwein:1991:VMT

[2572] Peter B. Borwein. Variations on Müntz’s theme. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 34(3):305–310, September 1991. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Casacuberta:1991:BHL

[2573] Carles Casacuberta. The behaviour of homology in the localization of finite groups. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 34(3):311–320, September 1991. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Chaudhry:1991:HPD

[2574] M. A. Chaudhry and J. N. Pandey. The Hilbert problem — a distributional approach. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 34(3):321–328, September 1991. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Effros:1991:NAO

[2575] Edward G. Effros and Zhong-Jin Ruan. A new approach to operator spaces. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 34(3):329–337, September 1991. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Fieldsteel:1991:PAC

[2576] Adam Fieldsteel. Poorly approximated Z_{2}-cocycles for transformations with rational discrete spectrum. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 34(3):338–342, September 1991. CO-DEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Huotari:1991:SMA

[2577] Robert Huotari and Salem Sahab. Simultaneous monotone L_{p} approximation, p < ∞. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 34(3):343–350, September 1991. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Ingram:1991:TCH

[2578] W. T. Ingram and D. D. Sherling. Two continua having a property of J. L. Kelley. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 34(3):351–356, September 1991. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Izumiya:1991:TPL

[2579] Shyuichi Izumiya. Topological properties of Legendrian singularities. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 34(3):357–363, September 1991. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Jespers:1991:SPI

[2580] Eric Jespers. Special principal ideal rings and absolute subretracts. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 34(3):364–367, September 1991. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Mistlbacher:1991:OTC

[2581] August Mistlbacher. An order-theoretical characterization of varieties of bands. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 34(3):368–374, September 1991. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Murty:1991:ECM

[2582] M. Ram Murty. Elliptic curves and modular forms. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 34(3):375–384, September 1991. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Remillard:1991:LTB

[2583] Bruno Rémillard and Donald A. Dawson. A limit theorem for Brownian motion in a random scenery. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 34(3):385–391, September 1991. CO-DEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Rhoades:1991:CNS

[2584] B. E. Rhoades. Corrections: “l^{p}-norms of some generalized Hausdorff matrices”. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 34(3):392, September 1991. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic). See [2441].

Silver:1991:MLJ

[2585] Daniel S. Silver. Delta-moves on links and Jones polynomial evaluations. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 34(3):393–400, September 1991. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Subbarao:1991:SCC

[2586] M. V. Subbarao and L.-W. Yip. On Sierpiński’s conjecture concerning the Euler totient. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 34(3):401–404, September 1991. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Szabo:1991:CDE

[2587] László I. Szabó. The converse of the dominated ergodic theorem in Hurewicz setting. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 34(3):405–411, September 1991. CODEN CM-BUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Verona:1991:NMU

[2588] Andrei Verona and Maria Elena Verona. A note on minimal usco maps. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 34(3):412–416, September 1991. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Vulakh:1991:CMF

[2589] L. Ya. Vulakh. Classification of maximal Fuchsian subgroups of some Bianchi groups. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 34(3):417–422, September 1991. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Wang:1991:NBC

[2590] You-Qiang Wang. A note on Brauer character degrees of solvable groups. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 34(3):423–425, September 1991. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Young:1991:EPS

[2591] Wo-Sang Young. On an estimate of the partial sums of Vilenkin–Fourier. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 34(3):426–432, September 1991. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Ballico:1991:DPV

[2592] E. Ballico. On the dual of projective varieties. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 34(4):433–439, December 1991. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Beckhoff:1991:MPI

[2593] Ferdinand Beckhoff. The minimal primal ideal space of a C^{∗}-algebra and local compactness. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 34(4):440–446, December 1991. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Bensebah:1991:AAF

[2594] Ali Bensebah. JV-algèbres et JH^{∗}-algèbres. (French) [JV-algebras and JH^{∗}-algebras]. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 34(4):447–455, December 1991. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Bulman-Fleming:1991:PFA

[2595] Sydney Bulman-Fleming. Pullback-flat acts are strongly flat. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 34(4): 456–461, December 1991. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Camillo:1991:CRA

[2596] Victor Camillo and Mohamed F. Yousif. Continuous rings with ACC on annihilators. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 34(4):462–464, December 1991. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Fabec:1991:MBS

[2597] R. C. Fabec. Multipliers between Sobolev spaces. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 34(4):465–473, December 1991. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Harris:1991:PPI

[2598] V. C. Harris and M. V. Subbarao. On product partitions of integers. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 34(4):474–479, December 1991. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Jin:1991:SRR

[2599] Gyo Taek Jin and Dale Rolfsen. Some remarks on rotors in link theory. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 34(4):480–484, December 1991. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Koua:1991:SEC

[2600] Konin Koua. Similarité entre certains quotients de A^{+} et des quotients d’algèbres uniformes. (French) [Similarity between certain quotients of A^{+} and quotients of uniform algebras]. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 34(4):485–491, December 1991. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Lester:1991:BQT

[2601] J. A. Lester. A Beckman–Quarles type theorem for Coxeter’s inversive distance. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 34(4):492–498, December 1991. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Nassrallah:1991:BDS

[2602] Bassam Nassrallah. Basic double series, quadratic transformations and products of basic series. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 34(4):499–513, December 1991. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Noussair:1991:ASE

[2603] Ezzat S. Noussair and Charles A. Swanson. Asymptotics for semilinear elliptic systems. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 34(4):514–519, December 1991. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Pakula:1991:IST

[2604] Lewis Pakula and Sol Schwartzman. Independence for sets of topological spheres. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 34(4):520–524, December 1991. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Reid:1991:CIT

[2605] Les Reid, Leslie G. Roberts, and Moshe Roitman. On complete intersections and their Hilbert functions. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 34(4):525–535, December 1991. CO-DEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Rolfsen:1991:PLI

[2606] Dale Rolfsen. PL link isotopy, essential knotting and quotients of polynomials. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 34(4):536–541, December 1991. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Ramaiah:1991:CUU

[2607] V. Sita Ramaiah and M. V. Subbarao. Convolutions with unbounded unity. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 34(4):542–546, December 1991. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Tzannes:1991:MSR

[2608] V. Tzannes. A Moore strongly rigid space. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 34(4):547–552, December 1991. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Wolak:1991:CLT

[2609] Robert A. Wolak. Closure of leaves in transversely affine foliations. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 34(4):553–558, December 1991. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Xu:1991:FPT

[2610] Hong Kun Xu. A fixed point theorem for semigroups of proximately uniformly Lipschitzian mappings. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 34(4):559–562, December 1991. CO-DEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Zelewski:1991:RCC

[2611] Piotr M. Zelewski. Ranks of chain complexes over the complex polynomial ring. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 34(4):563–565, December 1991. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Ballico:1992:PVP

[2612] E. Ballico. On projective varieties with projectively equivalent zerodimensional linear sections. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 35(1):3–13, March 1992. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Borwein:1992:TTC

[2613] David Borwein. A Tauberian theorem concerning Borel-type and Riesz summability methods. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 35(1):14–20, March 1992. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Borwein:1992:KMI

[2614] Jonathan Borwein, Simeon Reich, and Itai Shafrir. Krasnosel’ski–Mann iterations in normed spaces. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 35(1):21–28, March 1992. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Bruns:1992:FGM

[2615] Günter Bruns and Michael Roddy. A finitely generated modular ortholattice. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 35(1):29–33, March 1992. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Huaihui:1992:MPS

[2616] Chen Huaihui and P. M. Gauthier. A maximum principle for subharmonic and plurisubharmonic functions. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 35(1):34–39, March 1992. CO-DEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Choa:1992:NSB

[2617] Jun Soo Choa. Note on the space BMOA. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 35(1):40–45, March 1992. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Cranston:1992:PAG

[2618] M. Cranston. A probabilistic approach to gradient estimates. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 35(1): 46–55, March 1992. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Dowling:1992:SPC

[2619] Patrick N. Dowling. A stability property of a class of Banach spaces not containing c_{0}. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 35(1):56–60, March 1992. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Erbe:1992:WAT

[2620] Lynn H. Erbe and Peng Xiang Yan. Weighted averaging techniques in oscillation theory for second order difference equations. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 35(1):61–69, March 1992. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Evans:1992:ESM

[2621] Steven N. Evans. The entrance space of a measure-valued Markov branching process conditioned on nonextinction. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 35(1):70–74, March 1992. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Gilbert:1992:RBI

[2622] Shirley M. F. Gilbert and Peter Zvengrowski. A relation between S^{1} and S^{3}-invariant homotopy in the stable range. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 35(1):75–80, March 1992. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

GowriSankaran:1992:CMS

[2623] Kohur GowriSankaran and Ramasamy Jesuraj. Compact multipolar sets. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 35(1):81–83, March 1992. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Hofbauer:1992:HDE

[2624] Franz Hofbauer and Peter Raith. The Hausdorff dimension of an ergodic invariant measure for a piecewise monotonic map of the interval. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 35(1): 84–98, March 1992. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Lawrence:1992:AIG

[2625] John Lawrence. Algebraic ideals in group rings and tensor products. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 35(1):99–102, March 1992. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Lewis:1992:APP

[2626] D. W. Lewis. Annihilating polynomials and positive forms. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 35(1): 103–107, March 1992. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Mercier:1992:DPV

[2627] Armel Mercier and Werner Georg Nowak. A divisor problem for values of polynomials. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 35(1):108–115, March 1992. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Power:1992:FLP

[2628] S. C. Power. Finite lattices of projections in factors and approximately finite C^{∗}-algebras. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 35(1):116–125, March 1992. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Rosen:1992:ADS

[2629] Mary P. Rosen and Jerry D. Rosen. Automorphisms and derivations of skew polynomial rings. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 35(1):126–132, March 1992. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Rush:1992:ART

[2630] David E. Rush. Artin rings with two-generated ideals. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 35(1):133– 135, March 1992. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Yin:1992:RIS

[2631] Hong-Ming Yin. Regularity of the interfaces in the Stefan problem with a mushy region. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 35(1):136–144, March 1992. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Bae:1992:TDM

[2632] Sunghan Bae and Pyung-Lyun Kang. On Tate–Drinfel’d modules. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 35 (2):145–151, June 1992. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Bedard:1992:PCC

[2633] François Bédard and Alain Goupil. The poset of conjugacy classes and decomposition of products in the symmetric group. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 35(2):152–160, June 1992. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Currie:1992:WNR

[2634] J. Currie and A. Bendor-Samuel. Words without near-repetitions. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 35 (2):161–166, June 1992. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Dominguez:1992:DSC

[2635] Demetrio Domínguez. Deformations of secondary classes for subfoliations. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 35(2):167–173, June 1992. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Du:1992:CRR

[2636] Xian Kun Du. The centers of a radical ring. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 35(2):174–179, June 1992. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Ghahramani:1992:MHD

[2637] F. Ghahramani and J. P. McClure. Module homomorphisms of the dual modules of convolution Banach algebras. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 35(2):180–185, June 1992. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Guennoun:1992:PPF

[2638] Zine E. A. Guennoun. Sur un problème périodique. (French) [On a periodic problem]. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 35(2):186–193, June 1992. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Hare:1992:NOM

[2639] Kathryn E. Hare. Norm one multipliers on subspaces of L^{p}. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 35(2): 194–203, June 1992. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Howard:1992:CEB

[2640] Ralph Howard and Margaret Reese. Characterization of eigenfunctions by boundedness conditions. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 35(2):204–213, June 1992. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Hui:1992:CTE

[2641] Kin Ming Hui. Comparison theorems for the eigenvalues of the Laplacian in the unit ball in R^{n}. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 35(2):214–217, June 1992. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Khavinson:1992:EPH

[2642] Dmitry Khavinson. An extremal problem for harmonic functions in the ball. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 35(2):218–220, June 1992. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Koszmider:1992:CIP

[2643] Piotr Koszmider. On the complete invariance property in some uncountable products. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 35(2):221–229, June 1992. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Lai:1992:REC

[2644] K. F. Lai. Regular elliptic classes and the stable relative trace formula. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 35(2):230–236, June 1992. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Lari-Lavassani:1992:SEB

[2645] Ali Lari-Lavassani and Yung-Chen Lu. On the stability of equivariant bifurcation problems and their unfoldings. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 35(2):237–246, June 1992. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Masaike:1992:RMQ

[2646] Kanzo Masaike. Reflexive modules over QF-3 rings. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 35(2):247–251, June 1992. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

McBride:1992:CMM

[2647] A. C. McBride and W. J. Spratt. A class of Mellin multipliers. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 35(2): 252–260, June 1992. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Mysior:1992:UCS

[2648] Adam Mysior. Universal compact T_{1}-spaces. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 35(2):261–266, June 1992. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Pomares:1992:RLS

[2649] Beatriz Porras Pomares. Representation of p-lattice summing operators. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 35(2):267–277, June 1992. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Tang:1992:SCB

[2650] Wai-Shing Tang. Symmetrically completely bounded linear maps between C^{∗}-algebras. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 35(2):278–286, June 1992. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Barr:1992:LS

[2651] Michael Barr and Charles Wells. On the limitations of sketches. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 35(3): 287–294, September 1992. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Berger:1992:KR

[2652] Ruth I. Berger. The kernel of C(N)oC(N(\sqrt{−1})) and the 4-rank of K_{2}(O). Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 35(3):295–302, September 1992. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Chen:1992:CIN

[2653] Sheng Chen. Constructing isospectral but nonisometric Riemannian manifolds. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 35(3):303–310, September 1992. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Day:1992:CNC

[2654] Alan Day and J. B. Nation. Congruence normal covers of finitely generated lattice varieties. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 35(3):311–320, September 1992. CODEN CM-BUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Ebanks:1992:CGF

[2655] B. R. Ebanks, Pl. Kannappan, and P. K. Sahoo. A common generalization of functional equations characterizing normed and quasi-inner-product spaces. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 35(3):321–327, September 1992. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Faucette:1992:HDH

[2656] William M. Faucette. Higher-dimensional harmonic volume can be computed as an iterated integral. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 35(3):328–340, September 1992. CO-DEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Fegan:1992:SOO

[2657] H. D. Fegan. Second order operators with nonzero eta invariant. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 35(3): 341–353, September 1992. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Ferrandez:1992:NTH

[2658] Angel Ferrández and Pascual Lucas. Null 2-type hypersurfaces in a Lorentz space. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 35(3):354–360, September 1992. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Friesen:1992:CND

[2659] Christian Friesen. Class number divisibility in real quadratic function fields. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 35(3):361–370, September 1992. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Iraniparast:1992:MSC

[2660] Nezam Iraniparast. A method of solving a class of CIV boundary value problems. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 35(3):371–375, September 1992. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Jin:1992:BLM

[2661] Gyo Taek Jin and Ki Hyoung Ko. Boundary links and mutations. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 35(3): 376–380, September 1992. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Jones:1992:ASC

[2662] William B. Jones, W. J. Thron, and Nancy J. Wyshinski. An application of separate convergence for continued fractions to orthogonal polynomials. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 35(3):381–389, September 1992. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Kim:1992:RFP

[2663] Goansu Kim and C. Y. Tang. On the residual finiteness of polygonal products of nilpotent groups. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 35(3):390–399, September 1992. CO-DEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Knuppel:1992:FST

[2664] Frieder Knüppel. Factorization into symmetries and transvections of given conjugacy classes. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 35(3):400–409, September 1992. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Rhoades:1992:NST

[2665] B. E. Rhoades. bv_{0}-norms for some triangular matrices. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 35(3):410– 415, September 1992. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Rosenoer:1992:TCO

[2666] Shlomo Rosenoer. Trace-class operators in CSL algebras. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 35(3):416– 422, September 1992. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Yokota:1992:SD

[2667] Hisashi Yokota. On a sum of divisors. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 35(3):423–430, September 1992. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Anderson:1992:HRA

[2668] David F. Anderson, David E. Dobbs, and Marco Fontana. Hilbert rings arising as pullbacks. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 35(4):431–438, December 1992. CODEN CM-BUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Beer:1992:TCF

[2669] Gerald Beer. Topological completeness of function spaces arising in the Hausdorff approximation of functions. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 35(4):439–448, December 1992. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Bensebah:1992:WTA

[2670] Ali Bensebah. Weakness of the topology of a JB^{∗}-algebra. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 35(4): 449–454, December 1992. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Blair:1992:VCC

[2671] D. E. Blair and D. Perrone. A variational characterization of contact metric manifolds with vanishing torsion. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 35(4):455–462, December 1992. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Borwein:1992:STS

[2672] J. M. Borwein and M. Théra. Sandwich theorems for semicontinuous operators. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 35(4):463–474, December 1992. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

DeGrande-DeKimpe:1992:NAF

[2673] N. De Grande-De Kimpe, C. Perez-Garcia, and W. H. Schikhof. Non-Archimedean t-frames and FM-spaces. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 35(4):475–483, December 1992. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Foldes:1992:CSO

[2674] Stephan Foldes. Conditions for the separability of objects in twodimensional velocity fields. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 35(4):484–491, December 1992. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Heatherly:1992:FCN

[2675] H. E. Heatherly and J. D. P. Meldrum. Finiteness conditions for nearrings. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 35(4):492–496, December 1992. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Kalamidas:1992:CSC

[2676] N. D. Kalamidas and G. D. Spiliopoulos. Compact sets in C_{p}(X) and calibers. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 35(4):497–502, December 1992. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Kucharz:1992:DVR

[2677] W. Kucharz. Divisors on varieties over a real closed field. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 35(4): 503–509, December 1992. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Mayne:1992:CAL

[2678] Joseph H. Mayne. Centralizing automorphisms of Lie ideals in prime rings. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 35(4):510–514, December 1992. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Mbekhta:1992:CIG

[2679] Mostafa Mbekhta. Conorme et inverse généralisé dans les C^{∗}-algèbres. (French) [conorm and generalized inverse in C^{∗}-algebras]. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 35(4):515– 522, December 1992. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Noll:1992:GSF

[2680] Dominikus Noll. Generalized second fundamental form for Lipschitzian hypersurfaces by way of second epi derivatives. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 35(4):523–536, December 1992. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Rosebrugh:1992:CCD

[2681] Robert Rosebrugh and R. J. Wood. Constructive complete distributivity. III. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 35(4):537–547, December 1992. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Szwarc:1992:CCO

[2682] Ryszard Szwarc. Connection coefficients of orthogonal polynomials. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 35(4):548–556, December 1992. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Terasawa:1992:RCM

[2683] Jun Terasawa. Rigid continua with many embeddings. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 35(4):557– 559, December 1992. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Thas:1992:EPS

[2684] C. Thas. An easy proof for some classical theorems in plane geometry. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 35(4):560–568, December 1992. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Abramovich:1993:SCD

[2685] Y. A. Abramovich and A. W. Wickstead. Some characterizations of Dedekind α-completeness of a Riesz space. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 36(1):3–7, March 1993. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Borwein:1993:NO

[2686] David Borwein. Nörlund operators on l_{p}. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 36(1):8–14, March 1993. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Brown:1993:DCS

[2687] Tom C. Brown. Descriptions of the characteristic sequence of an irrational. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 36(1):15–21, March 1993. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Chiswell:1993:SRO

[2688] I. M. Chiswell and P. H. Kropholler. Soluble right orderable groups are locally indicable. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 36(1):22–29, March 1993. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Clarke:1993:PSC

[2689] F. H. Clarke and R. M. Redheffer. The proximal subgradient and constancy. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 36(1):30–32, March 1993. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Cogswell:1993:MZD

[2690] Kurt D. Cogswell. A multiparameter, zero density subsequence ergodic theorem. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 36(1):33–37, March 1993. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Gluchoff:1993:LDF

[2691] Alan D. Gluchoff. On logarithmic derivatives of functions in a class of starlike mappings. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 36(1):38–44, March 1993. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Gronbaek:1993:AIB

[2692] Niels Grønbæk and George A. Willis. Approximate identities in Banach algebras of compact operators. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 36(1):45–53, March 1993. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Guinez:1993:SAL

[2693] V. Guíñez, E. Sáez, and I. Szántó. Small amplitude limit cycles for cubic systems. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 36(1):54–63, March 1993. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Guo:1993:PHZ

[2694] Kang Hui Guo. On the p-thin problem for hypersurfaces of R^{n} with zero Gaussian curvature. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 36(1):64–73, March 1993. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Krisztin:1993:CCN

[2695] T. Krisztin, R. M. Mathsen, and Yuan Tong Xu. Counterexamples to a conjecture for neutral equations. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 36(1):74–77, March 1993. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Kwon:1993:HEM

[2696] E. G. Kwon. On homogeneous expansions of mixed norm space functions in the ball. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 36(1):78–86, March 1993. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Lakey:1993:WRC

[2697] Joseph D. Lakey. Weighted restriction for curves. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 36(1):87–95, March 1993. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Lefmann:1993:MPP

[2698] Hanno Lefmann and Hamza Si-Kaddour. On a maximality property of partition regular systems of equations. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 36(1):96–102, March 1993. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Mayer:1993:DMF

[2699] Daniel C. Mayer. Discriminants of metacyclic fields. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 36(1):103–107, March 1993. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Mollin:1993:CER

[2700] R. A. Mollin and H. C. Williams. Classification and enumeration of real quadratic fields having exactly one noninert prime less than a Minkowski bound. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 36(1):108–115, March 1993. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Poliquin:1993:CPS

[2701] R. A. Poliquin. A characterization of proximal subgradient set-valued mappings. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 36(1):116–122, March 1993. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Seddighi:1993:IFA

[2702] K. Seddighi and H. Zahedani. M-ideals and function algebras. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 36(1): 123–128, March 1993. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Cao:1993:OGA

[2703] Yulin Cao and H. I. Freedman. Oscillation and global attractivity in stage-structured population models. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 36(2):129–138, June 1993. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Costa:1993:SRC

[2704] Antone Costa. A sequence of results on class number congruences. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 36 (2):139–143, June 1993. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Fabrykowski:1993:MRD

[2705] J. Fabrykowski. On maximal residue difference sets modulo p. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 36(2): 144–146, June 1993. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Feigelstock:1993:AR

[2706] Shalom Feigelstock. E-associative rings. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 36(2):147–153, June 1993. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Grispolakis:1993:SCI

[2707] J. Grispolakis, J. Nikiel, J. N. Simone, and E. D. Tymchatyn. Separators in continuous images of ordered continua and hereditarily locally connected continua. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 36(2):154–163, June 1993. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Headley:1993:CTL

[2708] V. B. Headley. Comparison theorems for linear elliptic equations. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 36 (2):164–172, June 1993. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Hudzik:1993:OSS

[2709] Henryk Hudzik. Orlicz spaces without strongly extreme points and without H-points. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 36(2):173–177, June 1993. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Jespers:1993:BBC

[2710] E. Jespers, G. Leal, and M. M. Parmenter. Bicyclic and Bass cyclic units in group rings. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 36(2):178–182, June 1993. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Kaczynski:1993:LHB

[2711] Tomasz Kaczynski and Wiesław Krawcewicz. A local Hopf bifurcation theorem for a certain class of implicit differential equations. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 36(2): 183–189, June 1993. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Louboutin:1993:QFE

[2712] Stéphane Louboutin. Quelques formules exactes pour des moyennes de fonctions L de Dirichlet. (French) [Some exact formulas for averages of Dirichlet L-functions]. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 36(2):190–196, June 1993. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic). Voir correction [2762].

Lucchetti:1993:TSS

[2713] Roberto Lucchetti, Anna Torre, and Roger J.-B. Wets. A topology for the solid subsets of a topological space. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 36(2):197–208, June 1993. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Mekler:1993:CT

[2714] Alan H. Mekler and Saharon Shelah. The canary tree. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 36(2):209– 215, June 1993. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Mountford:1993:MRF

[2715] T. S. Mountford. A metastable result for the finite multidimensional contact process. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 36(2):216–226, June 1993. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Nicholson:1993:PER

[2716] W. K. Nicholson. On PP-endomorphism rings. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 36(2):227–230, June 1993. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Oda:1993:PHS

[2717] Nobuyuki Oda. Pairings of homotopy sets over and under B. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 36(2): 231–240, June 1993. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Pakula:1993:RCA

[2718] Lewis Pakula. Regions cut by arrangements of topological spheres. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 36 (2):241–244, June 1993. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Pan:1993:RCM

[2719] Yibiao Pan. A remark on convolution with measures supported on curves. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 36(2):245–250, June 1993. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Urbina:1993:LCL

[2720] A. M. Urbina, G. León de la Barra, M. León de la Barra, and M. Cañas. Limit cycles of Liénard equations with nonlinear damping. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 36(2):251– 256, June 1993. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Gaillard:1993:HSH

[2721] Pierre-Yves Gaillard. Harmonic spinors on hyperbolic space. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 36(3): 257–262, September 1993. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Graef:1993:SAP

[2722] John R. Graef and Paul W. Spikes. Some asymptotic properties of solutions of a neutral delay equation with an oscillatory coefficient. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 36(3): 263–272, September 1993. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Helversen-Pasotto:1993:BFL

[2723] Anna Helversen-Pasotto and Patrick Solé. Barnes’ first lemma and its finite analogue. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 36(3):273–282, September 1993. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Jevtic:1993:CMM

[2724] Miroljub Jevtić and Ivan Jovanović. Coefficient multipliers of mixed norm spaces. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 36(3):283–285, September 1993. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Kaczynski:1993:HBI

[2725] Tomasz Kaczynski and Hua Xing Xia. Hopf bifurcation for implicit neutral functional-differential equations. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 36(3):286–295, September 1993. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Kim:1993:CSS

[2726] Goansu Kim. Cyclic subgroup separability of generalized free products. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 36(3):296–302, September 1993. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Mhaskar:1993:WPA

[2727] H. N. Mhaskar. Weighted polynomial approximation of entire functions on unbounded subsets of the complex plane. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 36(3):303–313, September 1993. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Milnes:1993:RCR

[2728] Paul Milnes. Representations of compact right topological groups. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 36(3):314–323, September 1993. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Murphy:1993:IFT

[2729] Gerard J. Murphy. Inner functions and Toeplitz operators. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 36(3): 324–331, September 1993. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Nakos:1993:IAT

[2730] George Nakos. On ideals annihilating the toral class of BP_{∗}((BZ/p^{k})^{n}). Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 36(3):332–343, September 1993. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Okon:1993:ODM

[2731] James S. Okon and J. Paul Vicknair. One-dimensional monoid rings with n-generated ideals. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 36(3):344–350, September 1993. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Patterson:1993:CSN

[2732] Wanda M. Patterson. Complemented c_{0}-subspaces of a nonseparable C(K)-space. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 36(3):351–357, September 1993. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Roy:1993:ADT

[2733] Damien Roy and Michel Waldschmidt. Autour du théorème du sousgroupe algébrique. (French) [Author of the theorem of the algebraic subgroup]. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 36(3):358–367, September 1993. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Scherk:1993:POH

[2734] John Scherk. On the pole order and Hodge filtrations of isolated hypersurface singularities. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 36(3):368–372, September 1993. CODEN CM-BUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Yue:1993:AHZ

[2735] Zhang Nan Yue and Kenneth S. Williams. Application of the Hurwitz zeta function to the evaluation of certain integrals. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 36(3):373–384, September 1993. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Allenby:1993:SSG

[2736] R. B. J. T. Allenby and C. Y. Tang. Subgroup separability of generalized free products of free-by-finite groups. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 36(4):385–389, December 1993. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Batten:1993:CGF

[2737] Lynn M. Batten and Peter M. Johnson. The collineation groups of Figueroa planes. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 36(4):390–397, December 1993. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Betancor:1993:SPH

[2738] J. J. Betancor and I. Marrero. Some properties of Hankel convolution operators. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 36(4):398–406, December 1993. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Borwein:1993:WHS

[2739] Jonathan M. Borwein and Simon Fitzpatrick. A weak Hadamard smooth renorming of L_{1}(Ω, μ). Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 36(4):407–413, December 1993. CODEN CM-BUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Brenken:1993:ICS

[2740] Berndt Brenken. Isomorphism classes of solenoidal algebras. I. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 36(4): 414–418, December 1993. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Hill:1993:SSI

[2741] Paul Hill and Charles Megibben. Subsocles supporting isotype and balanced subgroups. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 36(4):419–425, December 1993. CODEN CM-BUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Jardine:1993:LQC

[2742] J. F. Jardine. The Lichtenbaum–Quillen conjecture for fields. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 36(4): 426–441, December 1993. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Klemes:1993:NHI

[2743] Ivo Klemes. A note on Hardy’s inequality. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 36(4):442–448, December 1993. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Klimek:1993:NEC

[2744] Slawomir Klimek and Andrzej Lesniewski. A note on the entire cyclic cohomology of a finite-dimensional noncommutative space. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 36(4):449– 457, December 1993. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Kwon:1993:BEI

[2745] E. G. Kwon. L^{p} behavior of the eigenfunctions of the invariant Laplacian. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 36(4):458–465, December 1993. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Peri:1993:MCS

[2746] Carla Peri. On the minimal convex shell of a convex body. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 36(4): 466–472, December 1993. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Rousseau:1993:LBC

[2747] C. Rousseau and B. Toni. Local bifurcation of critical periods in vector fields with homogeneous nonlinearities of the third degree. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 36(4): 473–484, December 1993. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Ruan:1993:OSO

[2748] Shi Gui Ruan. Oscillations of second order neutral differential equations. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 36(4):485–496, December 1993. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Yoshinaga:1993:BAM

[2749] Etsuo Yoshinaga. Blow analytic mappings and functions. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 36(4):497– 506, December 1993. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Adamek:1994:ENR

[2750] J. Adámek. Existence and nonexistence of regular generators. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 37(1):3– 7, March 1994. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Akahori:1994:NAB

[2751] Takao Akahori and Kimio Miyajima. A note on the analogue of the Bogomolov type theorem on deformations of CR-structures. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 37(1): 8–12, March 1994. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Auinger:1994:LEV

[2752] Karl Auinger. On the lattice of existence varieties of locally inverse semigroups. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 37(1):13–20, March 1994. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Bremner:1994:GAK

[2753] Murray Bremner. Generalized affine Kac–Moody Lie algebras over localizations of the polynomial ring in one variable. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 37(1):21–28, March 1994. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Choksi:1994:MST

[2754] J. R. Choksi and M. G. Nadkarni. The maximal spectral type of a rank one transformation. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 37(1):29–36, March 1994. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Deveney:1994:FIR

[2755] James K. Deveney and David R. Finston. Fields of G_{a} invariants are ruled. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 37(1):37–41, March 1994. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Dow:1994:LMA

[2756] Alan Dow and Juris Steprans. The σ-linkedness of the measure algebra. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 37(1):42–45, March 1994. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Gauthier:1994:SEA

[2757] P. M. Gauthier. Subharmonic extensions and approximations. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 37(1): 46–53, March 1994. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Gyori:1994:ORS

[2758] István Győri and Janos Turi. An oscillation result for singular neutral equations. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 37(1):54–65, March 1994. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Hou:1994:IFD

[2759] Duo Hou and Bruce Torrence. Involutions fixing the disjoint union of odd-dimensional projective spaces. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 37(1):66–74, March 1994. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Imina:1994:HOA

[2760] Yoshiki Imina and Kichi-Suke Saito. Hankel operators associated with analytic crossed products. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 37(1):75–81, March 1994. CODEN CM-BUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Kane:1994:PDR

[2761] Richard Kane. Poincaré duality and the ring of coinvariants. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 37(1): 82–88, March 1994. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Louboutin:1994:CSE

[2762] Stéphane Louboutin. Correction: “Some exact formulas for averages of Dirichlet L-functions”. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 37(1):89, March 1994. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic). See [2712].

Mollin:1994:LBC

[2763] R. A. Mollin, L.-C. Zhang, and Paula Kemp. A lower bound for the class number of a real quadratic field of ERD type. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 37(1):90–96, March 1994. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Parmenter:1994:FCR

[2764] M. M. Parmenter. Finite coverings of rings by ideals. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 37(1):97–99, March 1994. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Smith:1994:SIS

[2765] Brent Smith. Structure induced on subsets of a lattice via information gathered from tensor products. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 37(1):100–111, March 1994. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Stroppel:1994:LCH

[2766] Markus Stroppel. Locally compact Hughes planes. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 37(1):112–123, March 1994. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Tilly:1994:DWI

[2767] Ben Tilly. Derivations whose iterates are zero or invertible on a left ideal. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 37(1):124–132, March 1994. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Waterhouse:1994:NCC

[2768] William C. Waterhouse. The normal closures of certain Kummer extensions. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 37(1):133–139, March 1994. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Zhou:1994:LIT

[2769] Zheng Ping Zhou. A lattice isomorphism theorem for nonsingular retractable modules. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 37(1):140–144, March 1994. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Bruns:1994:POL

[2770] Günter Bruns and Michael Roddy. Projective orthomodular lattices. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 37(2):145–153, June 1994. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Casali:1994:AEO

[2771] Maria Rita Casali. The average edge order of 3-manifold coloured triangulations. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 37(2):154–161, June 1994. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Chu:1994:NQF

[2772] Huah Chu and Yi Chuan Lang. A note on quotient fields of power series rings. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 37(2):162–164, June 1994. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Despic:1994:WAG

[2773] M. Despić and F. Ghahramani. Weak amenability of group algebras of locally compact groups. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 37(2):165–167, June 1994. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Ding:1994:SSP

[2774] Ping Ding and A. R. Freedman. Small sets of kth powers. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 37(2): 168–173, June 1994. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Dokuchaev:1994:PNG

[2775] Michael Dokuchaev. On a property of nilpotent groups. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 37(2):174– 177, June 1994. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Erbe:1994:NSO

[2776] L. H. Erbe, H. X. Xia, and J. H. Wu. Nonoscillation of second order superlinear differential equations. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 37(2):178–186, June 1994. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Evans:1994:WOT

[2777] Steven N. Evans and Neil O’Connell. Weighted occupation time for branching particle systems and a representation for the supercritical superprocess. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 37(2):187–196, June 1994. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Fan:1994:SI

[2778] Da Shan Fan. A singular integral on L^{2}(R^{n}). Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 37(2):197–201, June 1994. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Fine:1994:TCQ

[2779] Benjamin Fine. Trace classes and quadratic forms in the modular group. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 37(2):202–212, June 1994. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Futorny:1994:IVM

[2780] V. M. Futorny. Imaginary Verma modules for affine Lie algebras. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 37(2):

213–218, June 1994. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Glass:1994:ALO

[2781] A. M. W. Glass. Atomless lattice-ordered groups. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 37(2):219–221, June 1994. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Harding:1994:MCU

[2782] John Harding. The macneille completion of a uniquely complemented lattice. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 37(2):222–227, June 1994. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Jespers:1994:UIG

[2783] Eric Jespers, Guilherme Leal, and C. Polcino Milies. Units of integral group rings of some metacyclic groups. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 37(2):228–237, June 1994. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Ki:1994:CRH

[2784] U-Hang Ki and Young Jin Suh. On a characterization of real hypersurfaces of type A in a complex space form. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 37(2):238–244, June 1994. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Maeda:1994:IIR

[2785] Sadahiro Maeda and Kazumi Tsukada. Isotropic immersions into a real space form. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 37(2):245–253, June 1994. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Vervaat:1994:SAP

[2786] Wim Vervaat. Self-affine processes and the ergodic theorem. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 37(2): 254–262, June 1994. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Villamor:1994:GPS

[2787] Enrique Villamor. Geometric proofs of some classical results on boundary values for analytic functions. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 37(2):263–269, June 1994. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Watson:1994:MBO

[2788] Stephen Watson. Measures below outer measures. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 37(2):270–271, June 1994. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic). See acknowledgement of priority and addendum [2850].

Xiao:1994:OSS

[2789] Yu Fei Xiao. One sided SF-rings with certain chain conditions. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 37(2): 272–277, June 1994. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Yildirim:1994:TES

[2790] C. Yalçin Yildirim. On the tails of the exponential series. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 37(2):278– 286, June 1994. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Yousif:1994:SFR

[2791] Mohamed Yousif. On semiperfect FPF-rings. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 37(2):287–288, June 1994. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Beauregard:1994:FLD

[2792] Raymond A. Beauregard. Factorization in LCM domains with conjugation. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 37(3):289–293, September 1994. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Beer:1994:NEC

[2793] Gerald Beer. A note on epi-convergence. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 37(3):294–300, September 1994. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Bergamasco:1994:GHC

[2794] Adalberto P. Bergamasco and Sérgio Luís Zani. Global hypoellipticity of a class of second order operators. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 37(3):301–305, September 1994. CO-DEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Bresar:1994:NPL

[2795] Matej Bresar and Peter Semrl. Normal-preserving linear mappings. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 37(3):306–309, September 1994. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Britten:1994:LAL

[2796] D. J. Britten and F. W. Lemire. On level 0 affine Lie modules. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 37(3): 310–314, September 1994. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Centore:1994:CAI

[2797] P. M. Centore and E. R. Vrscay. Continuity of attractors and invariant measures for iterated function systems. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 37(3):315–329, September 1994. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Dajczer:1994:SNF

[2798] Marcos Dajczer and Ruy Tojeiro. Submanifolds with nonparallel first normal bundle. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 37(3):330–337, September 1994. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Dokovic:1994:CSR

[2799] D. Z. Doković, P. Check, and J.-Y. Hée. On closed subsets of root systems. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 37(3):338–345, September 1994. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Drnovsek:1994:POB

[2800] Roman Drnovsek. Parareductive operators on Banach spaces. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 37(3): 346–350, September 1994. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Hengartner:1994:DTD

[2801] W. Hengartner and D. Poulin. Distortion theorems for diffeomorphisms. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 37(3):351–360, September 1994. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Hirano:1994:RWS

[2802] Yasuyuki Hirano. On rings whose simple modules are flat. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 37(3): 361–364, September 1994. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Kokologiannaki:1994:UPP

[2803] C. G. Kokologiannaki, M. E. Muldoon, and P. D. Siafarikas. A unimodal property of purely imaginary zeros of Bessel and related functions. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 37(3):365–373, September 1994. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Li:1994:LOP

[2804] Chi-Kwong Li and Stephen Pierce. Linear operators preserving similarity classes and related results. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 37(3):374–383, September 1994. CO-DEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Lvovsky:1994:CSP

[2805] S. L’vovsky. On curves and surfaces with projectively equivalent hyperplane sections. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 37(3):384–392, September 1994. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Parameswaran:1994:GTC

[2806] Mangalam R. Parameswaran. A general Tauberian condition that implies Euler summability. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 37(3):393–398, September 1994. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Rahe:1994:FCF

[2807] M. Rahe. On finite coding factors of a class of random Markov chains. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 37(3):399–407, September 1994. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Sawyer:1994:UBH

[2808] P. Sawyer. On an upper bound for the heat kernel on SU^{∗}(2n)/Sp(n). Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 37(3):408–418, September 1994. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Tsuda:1994:USC

[2809] Koichi Tsuda. Universal spaces for closed images of σ-discrete metric spaces. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 37(3):419–427, September 1994. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Zorboska:1994:ADC

[2810] Nina Zorboska. Angular derivative and compactness of composition operators on large weighted Hardy spaces. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 37(3):428–432, September 1994. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Allenby:1994:RFP

[2811] R. B. J. T. Allenby. The residual finiteness of polygonal products — two counterexamples. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 37(4):433–436, December 1994. CODEN CM-BUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Anderson:1994:OHF

[2812] David F. Anderson, Scott T. Chapman, and William W. Smith. Overrings of half-factorial domains. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 37(4):437–442, December 1994. CO-DEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Bell:1994:CSC

[2813] Howard E. Bell and Mohamad Nagy Daif. On commutativity and strong commutativity-preserving maps. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 37(4):443–447, December 1994. CO-DEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Borwein:1994:MO

[2814] David Borwein and Xiaopeng Gao. Matrix operators on l_{p} to l_{q}. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 37(4): 448–456, December 1994. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Bresar:1994:SCP

[2815] Matej Bresar and C. Robert Miers. Strong commutativity preserving maps of semiprime rings. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 37(4):457–460, December 1994. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Craigen:1994:TSO

[2816] R. Craigen. Trace, symmetry and orthogonality. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 37(4):461–467, December 1994. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Evans:1994:PFM

[2817] Martin J. Evans. Presentations of the free metabelian group of rank 2. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 37(4):468–472, December 1994. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Ginsburg:1994:DRO

[2818] John Ginsburg. Dismantlability revisited for ordered sets and graphs and the fixed-clique property. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 37(4):473–481, December 1994. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Heil:1994:SFS

[2819] Wolfgang Heil and Wilbur Whitten. The Seifert fiber space conjecture and torus theorem for nonorientable 3-manifolds. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 37(4):482–489, December 1994. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Johnson:1994:MHS

[2820] E. W. Johnson, Johnny A. Johnson, and Monty B. Taylor. On modules having small cofinite irreducibles. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 37(4):490–494, December 1994. CO-DEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Katchalski:1994:TCS

[2821] Meir Katchalski and János Pach. Touching convex sets in the plane. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 37(4):495–504, December 1994. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Kawamura:1994:CCT

[2822] Kazuhiro Kawamura. A characterization of mLC^{n} compacta in terms of Gromov–Hausdorff convergence. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 37(4):505–513, December 1994. CO-DEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Kearnes:1994:INB

[2823] Keith Kearnes and Ross Willard. Inherently nonfinitely based solvable algebras. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 37(4):514–521, December 1994. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Manne:1994:CAE

[2824] Per E. Manne. Carleman approximation by entire functions on the union of two totally real subspaces of C^{n}. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 37(4):522–526, December 1994. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Marshall:1994:ATC

[2825] Murray A. Marshall. An approximation theorem for coarse V -topologies on rings. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 37(4):527–533, December 1994. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Maruyama:1994:FCW

[2826] Ken ichi Maruyama. Finite complexes whose self-homotopy equivalence groups realize the infinite cyclic group. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 37(4):534–536, December 1994. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Rhemtulla:1994:ILF

[2827] Akbar Rhemtulla and Howard Smith. On infinite locally finite groups. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 37(4):537–544, December 1994. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Rooney:1994:RIT

[2828] P. G. Rooney. On the range of an integral transformation. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 37(4): 545–548, December 1994. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Schwartzman:1994:TTT

[2829] Sol Schwartzman. Topological transitivity on the torus. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 37(4):549– 551, December 1994. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Steprans:1994:TBF

[2830] Juris Steprans, Stephen Watson, and Winfried Just. A topological Banach fixed point theorem for compact Hausdorff spaces. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 37(4):552– 555, December 1994. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Tzannes:1994:ECR

[2831] V. Tzannes. Every countable regular space without isolated points is connectifiable. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 37(4):556–559, December 1994. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Wang:1994:GE

[2832] Feng Yu Wang. Gradient estimates on R^{d}. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 37(4):560–570, December 1994. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Bae:1995:TPD

[2833] Sunghan Bae and Ja Kyung Koo. Torsion points of Drinfel’d modules. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 38(1):3–10, March 1995. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Bellow:1995:HET

[2834] A. Bellow, A. P. Calderón, and U. Krengel. Hopf’s ergodic theorem for particles with different velocities and the “strong sweeping out property”. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 38(1):11–15, March 1995. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Blair:1995:SVT

[2835] D. E. Blair and D. Perrone. Second variation of the “total scalar curvature” on contact manifolds. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 38(1):16–22, March 1995. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Bonilla:1995:ACS

[2836] A. Bonilla and J. C. Fariña. Lip, α approximation on closed sets with lip, α extension. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 38(1):23–33, March 1995. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Brown:1995:RWL

[2837] R. C. Brown, D. B. Hinton, and M. K. Kwong. A remark on a weighted Landau inequality of Kwong and Zettl. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 38(1):34–41, March 1995. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Choksi:1995:GER

[2838] J. R. Choksi and M. G. Nadkarni. The group of eigenvalues of a rank one transformation. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 38(1):42–54, March 1995. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Ferrero:1995:SFC

[2839] Miguel Ferrero. Semisimplicity of free centred extensions. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 38(1):55–58, March 1995. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Hoshino:1995:VHC

[2840] Mitsuo Hoshino. Vanishing of Hochschild cohomologies for local rings with embedding dimension two. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 38(1):59–65, March 1995. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Hurtubise:1995:CPE

[2841] J. C. Hurtubise. Configurations de particules et espaces de modules. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 38(1):66–79, March 1995. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Jespers:1995:BUS

[2842] E. Jespers. Bicyclic units in some integral group rings. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 38(1):80–86, March 1995. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Kodaka:1995:AFT

[2843] Kazunori Kodaka. Anzai and Furstenberg transformations on the 2-torus and topologically quasi-discrete spectrum. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 38(1):87–92, March 1995. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Koldobsky:1995:ISP

[2844] Alexander Koldobsky. Isometric stability property of certain Banach spaces. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 38(1):93–97, March 1995. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Mitsuhiro:1995:CNF

[2845] T. Mitsuhiro, T. Nakahara, and T. Uehara. The class number formula of a real quadratic field and an estimate of the value of a unit. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 38(1): 98–103, March 1995. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Nagahara:1995:TSS

[2846] Takasi Nagahara and Seiya Yokota. On traces of separable simple subalgebras in matrix rings. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 38(1):104–111, March 1995. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Rao:1995:CMG

[2847] M. Anwar Rao and Robert Sandling. The characterisation of modular group algebras having unit groups of nilpotency class 3. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 38(1):112– 116, March 1995. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Schwartzman:1995:POG

[2848] Sol Schwartzman. Periodic orbits for generalized gradient flows. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 38(1): 117–119, March 1995. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Tang:1995:CSG

[2849] C. Y. Tang. Conjugacy separability of generalized free products of certain conjugacy separable groups. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 38(1):120–127, March 1995. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Watson:1995:APA

[2850] Stephen Watson. Acknowledgement of priority. addendum: “Measures below outer measures”. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 38(1):128, March 1995. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Atkins:1995:IPC

[2851] Richard Atkins and Zhong Ge. An inverse problem in the calculus of variations and the characteristic curves of connections on SO(3)-bundles. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 38(2):129–140, June 1995. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Bley:1995:LTR

[2852] W. Bley. A leopoldt-type result for rings of integers of cyclotomic extensions. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 38(2):141–148, June 1995. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Bokhari:1995:CAC

[2853] M. A. Bokhari. On constrained L^{2}-approximation of complex functions. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 38(2):149–155, June 1995. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Boroczky:1995:RSC

[2854] Károly Böröczky, Jr. and Martin Henk. Radii and the sausage conjecture. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 38(2):156–166, June 1995. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Clark:1995:GAC

[2855] David A. Clark and Masato Kuwata. Generalized Artin’s conjecture for primitive roots and cyclicity mod p of elliptic curves over function fields. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 38(2):167–173, June 1995. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Clase:1995:SGG

[2856] Michael V. Clase. Semiprime Goldie generalized matrix rings. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 38(2):

174–176, June 1995. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Curzio:1995:SFS

[2857] Mario Curzio and C. Kanta Gupta. Second Fox subgroups of arbitrary groups. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 38(2):177–181, June 1995. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Destrempes:1995:CIC

[2858] F. Destrempes and A. Pianzola. The contragredient isotypic component of the regular representation of pseudoreflection groups. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 38(2):182– 186, June 1995. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Dobbs:1995:X

[2859] David E. Dobbs and Evan G. Houston. On t-Spec(R[[X]]). Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 38(2): 187–195, June 1995. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Fu:1995:AEI

[2860] Siqi Fu. Asymptotic expansions of invariant metrics of strictly pseudoconvex domains. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 38(2):196–206, June 1995. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Girardi:1995:CCP

[2861] Maria Girardi and William B. Johnson. The complete continuity property and finite-dimensional decompositions. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 38(2):207–214, June 1995. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Guo:1995:PRR

[2862] Xiu Zhan Guo. On the periodic radical of a ring. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 38(2):215–217, June 1995. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Kalton:1995:EEB

[2863] N. J. Kalton. An elementary example of a Banach space not isomorphic to its complex conjugate. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 38(2):218–222, June 1995. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Orr:1995:SLO

[2864] John Lindsay Orr. Shuffling of linear orders. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 38(2):223–229, June 1995. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Phillips:1995:EIH

[2865] N. Christopher Phillips. Every invertible Hilbert space operator is a product of seven positive operators. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 38(2):230–236, June 1995. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Ransford:1995:CPE

[2866] T. J. Ransford. Characters and point evaluations. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 38(2):237–241, June 1995. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Weaver:1995:INL

[2867] Nik Weaver. Isometries of noncompact Lipschitz spaces. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 38(2):242– 249, June 1995. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Wehrfritz:1995:JHT

[2868] B. A. F. Wehrfritz. A Jordan–Hölder theorem for finitary groups. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 38 (2):250–251, June 1995. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Zhong:1995:PLM

[2869] Yong Zhong. Positive linear mappings between C^{∗}-algebras. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 38(2): 252–256, June 1995. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Aldaz:1995:CNM

[2870] J. M. Aldaz and P. A. Loeb. Counterexamples in nonstandard measure theory. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 38(3):257–261, September 1995. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Amberg:1995:AGR

[2871] Bernhard Amberg and Oliver Dickenschied. On the adjoint group of a radical ring. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 38(3):262–270, September 1995. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Bak:1995:WLM

[2872] Jong-Guk Bak. Weighted lacunary maximal functions on curves. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 38(3): 271–277, September 1995. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Blei:1995:ETC

[2873] Ron C. Blei. An extension theorem concerning Fréchet measures. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 38(3): 278–285, September 1995. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Daher:1995:HUE

[2874] Mohamad Daher. Homéomorphismes uniformes entre les sphères unité des espaces d’interpolation. (French) [Uniform homomorphisms between unit spheres of interpolation spaces]. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 38(3):286–294, September 1995. CO-DEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Edmunds:1995:TWI

[2875] David E. Edmunds and Vakhtang M. Kokilashvili. Two-weighted inequalities for singular integrals. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 38(3):295–303, September 1995. CO-DEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Hanine:1995:PAM

[2876] El Mostafa Hanine. Polynômes anisotropes modulo p^{2}. (French) [Anisotropic polynomials modulo p^{2}]. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 38(3):304–307, September 1995. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Harrison:1995:TPF

[2877] K. J. Harrison. The tensor product formula for reflexive subspace lattices. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 38(3):308–316, September 1995. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Juriaans:1995:TUI

[2878] Stanley Orlando Juriaans. Torsion units in integral group rings. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 38(3): 317–324, September 1995. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Johnson:1995:HCS

[2879] E. W. Johnson and Johnny A. Johnson. The Hausdorff completion of the space of closed subsets of a module. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 38(3):325–329, September 1995. CO-DEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Kingan:1995:TIC

[2880] Robert J. Kingan. Tournaments and ideal class groups. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 38(3):330–333, September 1995. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

LeMerdy:1995:DOS

[2881] Christian Le Merdy. On the duality of operator spaces. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 38(3):334–346, September 1995. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Miller:1995:DSS

[2882] Allen R. Miller. m-dimensional Schlömilch series. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 38(3):347–351, September 1995. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Nahlus:1995:HLA

[2883] Nazih Nahlus. Homomorphisms of Lie algebras of algebraic groups and analytic groups. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 38(3):352–359, September 1995. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Perez-Garcia:1995:FDS

[2884] C. Perez-Garcia and Wim H. Schikhof. Finite-dimensional subspaces of the p-adic space l^{∞}. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 38(3):360–365, September 1995. CODEN CM-BUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Pergher:1995:AFP

[2885] Pedro L. Q. Pergher. (Z_{2})^{k}-actions fixing a product of spheres and a point. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 38(3):366–372, September 1995. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Vershinin:1995:SEO

[2886] Vladimir V. Vershinin and Aleksandr L. Anisimov. A series of elements of order 4 in the symplectic cobordism ring. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 38(3):373–381, September 1995. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Zhou:1995:CLP

[2887] Yi Qiang Zhou. A characterization of left perfect rings. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 38(3):382–384, September 1995. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic). See erratum [3319].

Arias:1995:SN

[2888] Alvaro Arias and Paul Müller. Similarity of nests in L_{p}, 1 ≤ p ≠ 2 < ∞. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 38(4):385–389, December 1995. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Bhatwadekar:1995:NGF

[2889] S. M. Bhatwadekar and K. P. Russell. A note on geometric factoriality. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 38(4):390–395, December 1995. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Boivin:1995:CMC

[2890] André Boivin and Roman Dwilewicz. CR mappings of circular CR manifolds. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 38(4):396–407, December 1995. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Chatters:1995:SFM

[2891] A. W. Chatters and M. M. Parmenter. Stably free modules over rings of generalised integer quaternions. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 38(4):408–411, December 1995. CO-DEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Davidson:1995:CIC

[2892] Kenneth R. Davidson, John Lindsay Orr, and David R. Pitts. Connectedness of the invertibles in certain nest algebras. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 38(4):412–420, December 1995. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Dlab:1995:QHE

[2893] V. Dlab, P. Heath, and F. Marko. Quasi-hereditary endomorphism algebras. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 38(4):421–428, December 1995. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Dobbs:1995:WNP

[2894] David E. Dobbs and Moshe Roitman. Weak normalization of power series rings. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 38(4):429–433, December 1995. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Ferrero:1995:NDG

[2895] Miguel Ferrero, Antonio Giambruno, and César Polcino Milies. A note on derivations of group rings. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 38(4):434–437, December 1995. CO-DEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Flammang:1995:CDM

[2896] V. Flammang. Comparaison de deux mesures de polynômes. (French) [comparison of two polynomial measures]. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 38(4):438–444, December 1995. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Lee:1995:SRH

[2897] Tsiu Kwen Lee. Semiprime rings with hypercentral derivations. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 38(4): 445–449, December 1995. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Levy:1995:NRT

[2898] Jason Levy. A note on the relative trace formula. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 38(4):450–461, December 1995. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Mitchell:1995:GCG

[2899] Jane M. O. Mitchell. The genus of the Coxeter graph. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 38(4):462– 464, December 1995. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Pappalardi:1995:MSG

[2900] Francesco Pappalardi. On minimal sets of generators for primitive roots. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 38(4):465–468, December 1995. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Polster:1995:CPD

[2901] B. Polster and G. F. Steinke. Cut and paste in 2-dimensional projective planes and circle planes. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 38(4):469–480, December 1995. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Redondo:1995:D

[2902] María Julia Redondo and Andréa Solotar. α-derivations. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 38(4):481– 489, December 1995. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Zheng:1995:QEF

[2903] Jian Hua Zheng. A quantitative estimate on fixed-points of composite meromorphic functions. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 38(4):490–495, December 1995. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Zizler:1995:DOK

[2904] Petr Zizler. Definitizable operators on a Krein space. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 38(4):496–506, December 1995. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Baragar:1996:UCM

[2905] Arthur Baragar. On the unicity conjecture for Markoff numbers. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 39(1):3– 9, March 1996. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Bullen:1996:CLP

[2906] P. S. Bullen and D. N. Sarkhel. A continuity-like property of derivatives. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 39(1):10–20, March 1996. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Coghlan:1996:DGA

[2907] Francis Coghlan and Peter Hoffman. Division graded algebras in the Brauer–Wall group. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 39(1):21–24, March 1996. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Dokovic:1996:SLM

[2908] D. Z. Doković and Nguyêñ Q. Tha’ng. Surjective linear maps between root systems with zero. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 39(1):25–34, March 1996. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Everest:1996:DSG

[2909] G. R. Everest and I. E. Shparlinski. Divisor sums of generalised exponential polynomials. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 39(1):35–46, March 1996. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Gilbert:1996:DAC

[2910] William J. Gilbert. The division algorithm in complex bases. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 39(1): 47–54, March 1996. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Herbera:1996:SIP

[2911] Dolors Herbera and Ahmad Shamsuddin. On self-injective perfect rings. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 39(1):55–58, March 1996. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Heggie:1996:IRG

[2912] M. Heggie and G. R. Nicklason. An integral representation for the generalized binomial function. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 39(1):59–67, March 1996. CODEN CM-BUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Jain:1996:MIH

[2913] S. K. Jain, S. R. López-Permouth, and S. Raza Syed. Mutual injective hulls. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 39(1):68–73, March 1996. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Kye:1996:FSP

[2914] Seung-Hyeok Kye. Facial structures for the positive linear maps between matrix algebras. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 39(1):74–82, March 1996. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Li:1996:CSO

[2915] Chi-Kwong Li and Wayne Whitney. Closed symmetric overgroups of S_{n} in O_{n}. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 39(1):83–94, March 1996. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

March:1996:MCR

[2916] Peter March, Maung Min-Oo, and Ernst A. Ruh. Mean curvature of Riemannian foliations. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 39(1):95–105, March 1996. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Mollin:1996:CGR

[2917] R. A. Mollin. A completely general Rabinowitsch criterion for complex quadratic fields. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 39(1):106–110, March 1996. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Okoh:1996:TFD

[2918] F. Okoh. Torsion-free and divisible modules over finite-dimensional algebras. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 39(1):111–114, March 1996. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Ramsay:1996:SAF

[2919] Keith Ramsay. A sieve on analytic functions. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 39(1):115–116, March 1996. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Shi:1996:MCH

[2920] Ying Guang Shi. Mean convergence of Hermite–Fejér interpolation based on the zeros of Lascenov polynomials. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 39(1):117–128, March 1996. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Baars:1996:LIL

[2921] Jan Baars and Helma Gladdines. On the linear invariance of Lindelöf numbers. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 39(2):129–137, June 1996. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Bass:1996:DMD

[2922] Richard F. Bass. The Doob–Meyer decomposition revisited. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 39(2): 138–150, June 1996. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Beaucoup:1996:BTI

[2923] Franck Beaucoup and Catherine Souchon. Bernstein-type inequalities with Bombieri norm. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 39(2):151–163, June 1996. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Bhatia:1996:JEC

[2924] R. Bhatia and L. Elsner. On joint eigenvalues of commuting matrices. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 39(2):164–168, June 1996. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Biswas:1996:SGI

[2925] I. Biswas and K. Guruprasad. On some geometric invariants associated to the space of flat connections on an open space. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 39(2):169–177, June 1996. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Ferrero:1996:RIS

[2926] M. Ferrero, E. R. Puczylowski, and S. Sidki. On the representation of an idempotent as a sum of nilpotent elements. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 39(2):178–185, June 1996. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Grochenig:1996:BCF

[2927] Karlheinz Gröchenig and Andrew Haas. Backward continued fractions and their invariant measures. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 39(2):186–198, June 1996. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Hassan:1996:PIS

[2928] Al-Zaid Hassan, B. Brindza, and Á. Pintér. On positive integer solutions of the equation xy + yz + xz = n. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 39(2):199–202, June 1996. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Jeong:1996:PIS

[2929] Ja A. Jeong, Kazunori Kodaka, and Hiroyuki Osaka. Purely infinite simple C^{∗}-crossed products. II. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 39(2):203–210, June 1996. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Lee:1996:CDI

[2930] P. H. Lee and T. L. Wong. Central ∗-differential identities in prime rings. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 39(2):211–215, June 1996. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Lorch:1996:FPI

[2931] Lee Lorch and Peter Szego. Further on the points of inflection of Bessel functions. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 39(2):216–218, June 1996. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Nakazi:1996:SMM

[2932] Takahiko Nakazi. Slice maps and multipliers of invariant subspaces. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 39 (2):219–226, June 1996. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Pruss:1996:NMP

[2933] Alexander R. Pruss. Nonexistence of maxima for perturbations of some inequalities with critical growth. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 39(2):227–237, June 1996. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Smith:1996:SRI

[2934] Larry Smith. Some rings of invariants that are Cohen–Macaulay. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 39(2): 238–240, June 1996. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Szwarc:1996:NLQ

[2935] Ryszard Szwarc. Nonnegative linearization and quadratic transformation of Askey–Wilson polynomials. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 39(2):241–249, June 1996. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Xue:1996:QDL

[2936] Weimin Xue. Quasi-duality, linear compactness and Morita duality for power series rings. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 39(2):250–256, June 1996. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Clark:1996:DNC

[2937] John Clark and Nguyen Viet Dung. On the decomposition of nonsingular CS-modules. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 39(3):257–265, September 1996. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Dilworth:1996:UKK

[2938] S. J. Dilworth and C. J. Lennard. Uniform Kadec–Klee Lorentz spaces L_{w,1} and uniformly concave functions. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 39(3):266–274, September 1996. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Graef:1996:OGA

[2939] J. R. Graef, C. Qian, and P. W. Spikes. Oscillation and global attractivity in a periodic delay equation. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 39(3):275–283, September 1996. CO-DEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Izuchi:1996:BAS

[2940] Keiji Izuchi, Kazuhiro Kasuga, and Yasuo Matsugu. Bourgain algebras of spaces of n-harmonic functions in the unit polydisk. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 39(3): 284–293, September 1996. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Kim:1996:CSC

[2941] Goansu Kim. Conjugacy separability of certain polygonal products. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 39(3):294–307, September 1996. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Lipparini:1996:CVD

[2942] Paolo Lipparini. A characterization of varieties with a difference term. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 39(3):308–315, September 1996. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Magill:1996:NCF

[2943] K. D. Magill, Jr. Nearrings of continuous functions from topological spaces into topological nearrings. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 39(3):316–329, September 1996. CO-DEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Maier:1996:PHC

[2944] Peter Maier and Markus Stroppel. Pseudo-homogeneous coordinates for Hughes planes. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 39(3):330–345, September 1996. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Marshall:1996:NCC

[2945] Mary K. Marshall. Numbers of conjugacy class sizes and derived lengths for A-groups. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 39(3):346–351, September 1996. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Park:1996:ATM

[2946] Joon-Sik Park and Won Tae Oh. The Abbena–Thurston manifold as a critical point. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 39(3):352–359, September 1996. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Soleev:1996:CBP

[2947] A. Soleev. Complicated bifurcations of periodic solutions in some system of ODE. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 39(3):360–366, September 1996. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Weaver:1996:WCD

[2948] Nik Weaver. Weak^{∗}-closed derivations from C[0, 1] into L^{∞}[0, 1]. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 39(3): 367–375, September 1996. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Ye:1996:DHO

[2949] Youpei Ye and Jiang Luh. Derivations of higher order in prime rings. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 39(3):376–384, September 1996. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Fantham:1996:RPC

[2950] Peter Fantham, Ioan James, and Michael Mather. On the reduced product construction. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 39(4):385–389, December 1996. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Bongiorno:1996:FDD

[2951] B. Bongiorno, W. F. Pfeffer, and B. S. Thomson. A full descriptive definition of the gage integral. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 39(4):390–401, December 1996. CO-DEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Chabert:1996:CPP

[2952] Jean-Luc Chabert. Une caractérisation des polynômes prenant des valeurs entières sur tous les nombres premiers. (French) [A characterization of polynomials taking whole values for all prime numbers]. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 39(4): 402–407, December 1996. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Chen:1996:SNF

[2953] Huaihui Chen and Paul M. Gauthier. On strongly normal functions. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 39(4):408–419, December 1996. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Donsig:1996:FAI

[2954] Allan P. Donsig and S. C. Power. The failure of approximate inner conjugacy for standard diagonals in regular limit algebras. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 39(4):420–428, December 1996. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Goodearl:1996:ARR

[2955] K. R. Goodearl. C^{∗}-algebras of real rank zero whose K_{0}’s are not Riesz groups. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 39(4):429–437, December 1996. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Hurri-Syrjanen:1996:GIB

[2956] Ritva Hurri-Syrjänen. A generalization of an inequality of Bhattacharya and Leonetti. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 39(4):438–447, December 1996. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Makai:1996:BAG

[2957] Endre Makai, Jr. and Horst Martini. On bodies associated with a given convex body. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 39(4):448–459, December 1996. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Phillips:1996:SAF

[2958] John Phillips. Self-adjoint Fredholm operators and spectral flow. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 39(4): 460–467, December 1996. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Sokolov:1996:TVI

[2959] M. V. Sokolov. The Turaev–Viro invariant for 3-manifolds is a sum of three invariants. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 39(4):468–475, December 1996. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Swardson:1996:WS

[2960] Mary Anne Swardson and Paul J. Szeptycki. When X^{∗} is a P′-space. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 39(4):476–485, December 1996. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Tomita:1996:WPC

[2961] Artur H. Tomita. The Wallace problem: a counterexample from MA_{countable} and p-compactness. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 39(4):486–498, December 1996. CO-DEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Zou:1996:DUE

[2962] Yi Ming Zou. Deformation of the universal enveloping algebra of Γ(σ_{1}, σ_{2}, σ_{3}). Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 39(4):499–506, December 1996. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Ayari:1997:FCL

[2963] S. Ayari and S. Dubuc. La formule de Cauchy sur la longueur d’une courbe. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 40(1):3–9, March 1997. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Borwein:1997:CFB

[2964] Jon Borwein and Jon Vanderwerff. Convex functions on Banach spaces not containing ℓ_{1}. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 40(1):10–18, March 1997. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Derbez:1997:LTS

[2965] Eric Derbez and Gordon Slade. Lattice trees and super-Brownian motion. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 40(1):19–38, March 1997. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Zhao:1997:PF

[2966] Dongsheng Zhao. On projective Z-frames. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 40(1):39–46, March 1997. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Hartl:1997:UCD

[2967] Manfred Hartl. A universal coefficient decomposition for subgroups induced by submodules of group algebras. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 40(1):47–53, March 1997. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Kechagias:1997:NMI

[2968] Nondas E. Kechagias. A note on U_{n} × U_{m} modular invariants. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 40(1): 54–59, March 1997. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Khavinson:1997:CPH

[2969] Dmitry Khavinson. Cauchy’s problem for harmonic functions with entire data on a sphere. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 40(1):60–66, March 1997. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Knight:1997:BMP

[2970] Frank B. Knight. On a Brownian motion problem of T. Salisbury. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 40 (1):67–71, March 1997. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Lee:1997:GSM

[2971] Min Ho Lee. Generalized Siegel modular forms and cohomology of locally symmetric varieties. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 40(1):72–80, March 1997. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Movahhedi:1997:CCS

[2972] A. Movahhedi and A. Salinier. Une caractérisation des corps satisfaisant le théorème de l’axe principal. (French) [A characterization of objects satisfying the principal axis theorem]. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 40(1):81–87, March 1997. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Radulescu:1997:MMV

[2973] M. L. Radulescu and F. H. Clarke. The multidirectional mean value theorem in Banach spaces. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 40(1):88–102, March 1997. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Riley:1997:TCL

[2974] David M. Riley and Vladimir Tasić. The transfer of a commutator law from a nil-ring to its adjoint group. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 40(1):103–107, March 1997. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Schaer:1997:CSM

[2975] J. Schaer. Continuous self-maps of the circle. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 40(1):108–116, March 1997. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Vigue:1997:LSP

[2976] Jean-Pierre Vigué. Un lemme de Schwarz pour les boules-unités ouvertes. (French) [A lemma of Schwarz for open unit balls]. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 40(1):117–128, March 1997. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Badea:1997:CAB

[2977] Catalin Badea. Sur les caractères d’une algèbre de Banach. (French) [On the characters of a Banach algebra]. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 40(2):129–132, June 1997. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Blackmore:1997:DTO

[2978] T. D. Blackmore. Derivations from totally ordered semigroup algebras into their duals. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 40(2):133–142, June 1997. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Bremner:1997:QDS

[2979] Murray Bremner. Quantum deformations of simple Lie algebras. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 40(2): 143–148, June 1997. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Brown:1997:MHC

[2980] Tom C. Brown, Bruce M. Landman, and Marni Mishna. Monochromatic homothetic copies of {1, 1 + s, 1 + s + t}. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 40(2):149–157, June 1997. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Coxeter:1997:THT

[2981] H. S. M. Coxeter. The trigonometry of hyperbolic tessellations. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 40(2): 158–168, June 1997. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Cruz-Uribe:1997:COS

[2982] David Cruz-Uribe. The class A_{∞}^{+}(g) and the one-sided reverse Hölder inequality. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 40(2):169–173, June 1997. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Hungerbuhler:1997:NUH

[2983] Norbert Hungerbühler. Non-uniqueness for the p-harmonic flow. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 40(2): 174–182, June 1997. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Kepert:1997:RGA

[2984] Andrew G. Kepert. The range of group algebra homomorphisms. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 40(2): 183–192, June 1997. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Kucerovsky:1997:FRO

[2985] Dan Kucerovsky. Finite rank operators and functional calculus on Hilbert modules over abelian C^{∗}-algebras. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 40(2):193–197, June 1997. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Meldrum:1997:RMN

[2986] J. D. P. Meldrum and J. H. Meyer. The J_{0}-radical of a matrix nearring can be intermediate. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 40(2):198–203, June 1997. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Meyerhoff:1997:ICH

[2987] Robert Meyerhoff and Mingqing Ouyang. The η-invariants of cusped hyperbolic 3-manifolds. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 40(2):204–213, June 1997. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Mollin:1997:PQT

[2988] R. A. Mollin, B. Goddard, and S. Coupland. Polynomials of quadratic type producing strings of primes. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 40(2):214–220, June 1997. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Rada:1997:SRW

[2989] Juan Rada and Manuel Saorín. On semiregular rings whose finitely generated modules embed in free modules. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 40(2):221–230, June 1997. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Reid:1997:ATF

[2990] N. Reid. Asymptotic theory and the foundations of statistics. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 40(2): 231–243, June 1997. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Naito:1997:NRP

[2991] Yuki Naito and Hiroyuki Usami. Nonexistence results of positive entire solutions for quasilinear elliptic inequalities. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 40(2):244–253, June 1997. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Saito:1997:SAS

[2992] Kichi-Suke Saito and Yasuo Watatani. Subdiagonal algebras for subfactors II (finite dimensional case). Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 40(2):254–256, June 1997. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Baikoussis:1997:CRH

[2993] Christos Baikoussis. A characterization of real hypersurfaces in complex space forms in terms of the Ricci tensor. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 40(3):257–265, September 1997. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Bechtell:1997:FGL

[2994] H. Bechtell, M. Deaconescu, and Gh. Silberberg. Finite groups with large automizers for their Abelian subgroups. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 40(3):266–270, September 1997. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Bergweiler:1997:NRP

[2995] Walter Bergweiler. Non-real periodic points of entire functions. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 40(3):

271–275, September 1997. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Chouikha:1997:FEE

[2996] Raouf Chouikha. Fonctions elliptiques et équations différentielles ordinaires. (French) [Elliptic functions and ordinary differential equations]. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 40(3):276–284, September 1997. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Crawford:1997:SHM

[2997] T. Arleigh Crawford. The space of harmonic maps from the 2-sphere to the complex projective plane. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 40(3):285–295, September 1997. CO-DEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Hare:1997:GAL

[2998] Kathryn E. Hare. A general approach to Littlewood–Paley theorems for orthogonal families. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 40(3):296–308, September 1997. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Hillman:1997:HFA

[2999] J. A. Hillman and M. Sakuma. On the homology of finite abelian coverings of links. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 40(3):309–315, September 1997. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Hudzik:1997:GPO

[3000] H. Hudzik, A. Kamińska, and M. Mastylo. On geometric properties of Orlicz–Lorentz spaces. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 40(3):316–329, September 1997. CODEN CM-BUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Kapovich:1997:APH

[3001] Ilya Kapovich. Amalgamated products and the Howson property. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 40(3):330–340, September 1997. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Lee:1997:SUT

[3002] Hyang-Sook Lee. The stable and unstable types of classifying spaces. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 40(3):341–351, September 1997. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Liriano:1997:NPT

[3003] Sal Liriano. A new proof of a theorem of Magnus. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 40(3):352–355, September 1997. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Mazet:1997:PDM

[3004] Pierre Mazet. Principe du maximum et lemme de Schwarz, à valeurs vectorielles. (French) [Schwarz’ vector-valued maximum principle and lemma]. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 40(3):356–363, September 1997. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Narayanan:1997:NVC

[3005] Sridhar Narayanan. On the non-vanishing of a certain class of Dirichlet series. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 40(3):364–369, September 1997. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Rolfsen:1997:WME

[3006] Dale Rolfsen and Li Zhongmou. Which 3-manifolds embed in Triod × I × I? Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 40(3):370–375, September 1997. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Gross:1997:DP

[3007] Benedict H. Gross and Gordan Savin. The dual pair PGL_{3} ×G_{2}. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 40(3): 376–384, September 1997. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Bae:1997:EUC

[3008] Sunghan Bae and Pyung-Lyun Kang. Elliptic units and class fields of global function fields. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 40(4):385–394, December 1997. CODEN CM-BUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Boudhraa:1997:SR

[3009] Zineddine Boudhraa. D-spaces and resolution. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 40(4):395–401, December 1997. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Carpenter:1997:PRN

[3010] Jenna P. Carpenter. On the preservation of root numbers and the behavior of Weil characters under reciprocity equivalence. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 40(4):402–415, December 1997. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Clemence:1997:SBT

[3011] Dominic P. Clemence. On the singular behaviour of the Titchmarsh– Weyl m-function for the perturbed Hill’s equation on the line. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 40(4): 416–421, December 1997. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Dow:1997:CSR

[3012] Alan Dow. On compact separable radial spaces. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 40(4):422–432, December 1997. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Guo:1997:UES

[3013] Kanghui Guo. A uniform L^{∞} estimate of the smoothing operators related to plane curves. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 40(4):433–442, December 1997. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Hadwin:1997:RRB

[3014] Don Hadwin and Mehmet Orhon. Reflective representations and Banach C^{∗}-modules. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 40(4):443–447, December 1997. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Kaczynski:1997:SIP

[3015] Tomasz Kaczynski and Marian Mrozek. Stable index pairs for discrete dynamical systems. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 40(4):448–455, December 1997. CODEN CM-BUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Kucharz:1997:ASM

[3016] Wojciech Kucharz and Kamil Rusek. Approximation of smooth maps by real algebraic morphisms. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 40(4):456–463, December 1997. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Kuo:1997:SNB

[3017] Chung-Cheng Kuo. On the solvability of a Neumann boundary value problem at resonance. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 40(4):464–470, December 1997. CODEN CM-BUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Lawrence:1997:SPE

[3018] Jim Lawrence. A short proof of Euler’s relation for convex polytopes*. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 40(4):471–474, December 1997. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Lou:1997:CMB

[3019] Zengjian Lou. Coefficient multipliers of Bergman spaces A^{p}, II. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 40(4): 475–487, December 1997. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Maouche:1997:CSD

[3020] Abdelaziz Maouche. Caractérisations spectrales du radical et du socle d’une paire de Jordan–Banach. (French) [Spectral characterization of the radical and pedestal of a Jordan–Banach pair]. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 40(4):488–497, December 1997. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Selvaraj:1997:MTB

[3021] Chikkanna Selvaraj and Suguna Selvaraj. Matrix transformations based on Dirichlet convolution. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 40(4):498–507, December 1997. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Anonymous:1997:AII

[3022] Anonymous. Author index — index des auteurs — for 1997 — pour 1997. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 40(4):508–511, December 1997. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Anderson:1998:RCI

[3023] David F. Anderson and David E. Dobbs. Root closure in integral domains, III. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 41(1):3–9, March 1998. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Borwein:1998:SCM

[3024] David Borwein. Simple conditions for matrices to be bounded operators on l_{p}. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 41(1):10–14, March 1998. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Brown:1998:STC

[3025] Tom Brown, Peter Jau-Shyong Shiue, and X. Y. Yu. Sequences with translates containing many primes. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 41(1):15–19, March 1998. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Brunetti:1998:NCC

[3026] Maurizio Brunetti. A new cohomological criterion for the p-nilpotence of groups. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 41(1):20–22, March 1998. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Clemence:1998:SAA

[3027] Dominic P. Clemence. Subordinacy analysis and absolutely continuous spectra for Sturm–Liouville equations with two singular endpoints. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 41 (1):23–27, March 1998. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Felix:1998:GGA

[3028] Yves Félix and Aniceto Murillo. Gorenstein graded algebras and the evaluation map. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 41(1):28–32, March 1998. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Gallant:1998:AET

[3029] Robert Gallant and Charles J. Colbourn. Asymptotic existence of tight orthogonal main effect plans. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 41(1):33–40, March 1998. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Giner:1998:CSI

[3030] E. Giner. On the Clarke subdifferential of an integral functional on L_{p}, 1 ≤ p < ∞. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 41(1):41–48, March 1998. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Harrison:1998:SWD

[3031] K. J. Harrison, J. A. Ward, and L-J. Eaton. Stability of weighted DARMA filters. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 41(1):49–64, March 1998. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Hassanabadi:1998:CCL

[3032] A. Mohammadi Hassanabadi and Akbar Rhemtulla. Criteria for commutativity in large groups. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 41(1):65–70, March 1998. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Hurrelbrink:1998:SPT

[3033] Jurgen Hurrelbrink and Ulf Rehmann. Splitting patterns and trace forms. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 41(1):71–78, March 1998. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Kelarev:1998:AQK

[3034] A. V. Kelarev. An answer to a question of Kegel on sums of rings. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 41(1):79–80, March 1998. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Lanski:1998:CCR

[3035] Charles Lanski. The cardinality of the center of a Π ring. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 41(1): 81–85, March 1998. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Lubinsky:1998:LEF

[3036] D. S. Lubinsky. On q-exponential functions for |q| = 1. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 41(1):86–97, March 1998. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Papistas:1998:AMG

[3037] Athanassios I. Papistas. Automorphisms of metabelian groups. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 41(1): 98–104, March 1998. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

So:1998:ECC

[3038] Wasin So. An explicit criterion for the convexity of quaternionic numerical range. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 41(1):105–108, March 1998. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Tahara:1998:GTD

[3039] Ken-Ichi Tahara, L. R. Vermani, and Atul Razdan. On generalized third dimension subgroups. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 41(1):109–117, March 1998. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Valenti:1998:PIS

[3040] Angela Valenti. On permanental identities of symmetric and skewsymmetric matrices in characteristic p. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 41(1):118–124, March 1998. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Boyd:1998:UAM

[3041] David W. Boyd. Uniform approximation to Mahler’s measure in several variables. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 41(1):125–128, March 1998. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Lee:1998:PSC

[3042] Young Joo Lee. Pluriharmonic symbols of commuting Toeplitz type operators on the weighted Bergman spaces. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 41(2):129–136, June 1998. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Choksi:1998:GCC

[3043] J. R. Choksi and M. G. Nadkarni. Genericity of certain classes of unitary and self-adjoint operators. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 41(2):137–139, June 1998. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Bullock:1998:SH

[3044] Doug Bullock, Charles Frohman, and Joanna Kania-Bartoszyńska. Skein homology. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 41(2):140–144, June 1998. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Fry:1998:SPU

[3045] R. Fry. Smooth partitions of unity on Banach spaces. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 41(2):145– 150, June 1998. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Fuchs:1998:EPM

[3046] Laszlo Fuchs and Sang Bum Lee. Equivalent presentations of modules over Prüfer domains. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 41(2):151–157, June 1998. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Gaal:1998:PIB

[3047] István Gaál. Power integral bases in composits of number fields. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 41(2): 158–165, June 1998. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Hof:1998:PPT

[3048] A. Hof. Percolation on Penrose tilings. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 41(2):166–177, June 1998. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Krupnik:1998:MPR

[3049] Ilya Krupnik and Peter Lancaster. Minimal pencil realizations of rational matrix functions with symmetries. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 41(2):178–186, June 1998. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Loh:1998:ESR

[3050] W. K. A. Loh. Exponential sums on reduced residue systems. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 41(2): 187–195, June 1998. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Nakazi:1998:BHT

[3051] Takahiko Nakazi. Brown–Halmos type theorems of weighted Toeplitz operators. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 41(2):196–206, June 1998. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Philos:1998:OCF

[3052] Ch. G. Philos and Y. G. Sficas. An oscillation criterion for first order linear delay differential equations. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 41(2):207–213, June 1998. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Shackell:1998:PRC

[3053] John Shackell. On a problem of Rubel concerning the set of functions satisfying all the algebraic differential equations satisfied by a given function. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 41(2):214–224, June 1998. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Vanderwerff:1998:MIP

[3054] Jon Vanderwerff. Mazur intersection properties for compact and weakly compact convex sets. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 41(2):225–230, June 1998. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Worthington:1998:GSC

[3055] R. L. Worthington. The growth series of compact hyperbolic Coxeter groups with 4 and 5 generators. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 41(2):231–239, June 1998. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Xia:1998:CGA

[3056] Jingbo Xia. On certain K-groups associated with minimal flows. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 41(2): 240–244, June 1998. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Yang:1998:NPC

[3057] Lecheng Yang. The normality in products with a countably compact factor. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 41(2):245–251, June 1998. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Yang:1998:DGA

[3058] Qingjie Yang. Dihedral groups of automorphisms of compact Riemann surfaces. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 41(2):252–256, June 1998. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Bagby:1998:NSS

[3059] Thomas Bagby and P. M. Gauthier. Note on the support of Sobolev functions. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 41(3):257–260, September 1998. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Barthwal:1998:SRW

[3060] S. Barthwal, S. Jhingan, and P. Kanwar. A simple ring over which proper cyclics are continuous is a PCI-ring. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 41(3):261–266, September 1998. CO-DEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Fukuma:1998:NAL

[3061] Yoshiaki Fukuma. On the nonemptiness of the adjoint linear system of polarized manifold. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 41(3):267–278, September 1998. CODEN CM-BUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Acosta:1998:NCR

[3062] María D. Acosta and Manuel Ruiz Galán. New characterizations of the reflexivity in terms of the set of norm attaining functionals. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 41(3): 279–289, September 1998. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Gratzer:1998:CLF

[3063] G. Grätzer, H. Lakser, and E. T. Schmidt. Congruence lattices of finite semimodular lattices. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 41(3):290–297, September 1998. CODEN CM-BUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Jahandideh:1998:ITS

[3064] M. T. Jahandideh. On the ideal-triangularizability of semigroups of quasinilpotent positive operators on C(K). Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 41(3):298–305, September 1998. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Kolasa:1998:OIN

[3065] Lawrence A. Kolasa. Oscillatory integrals with nonhomogeneous phase functions related to Schrödinger equations. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 41(3):306–317, September 1998. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Lipparini:1998:CVD

[3066] Paolo Lipparini. A characterization of varieties with a difference term, II: neutral = meet semi-distributive. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 41(3):318–327, September 1998. CO-DEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Mollin:1998:CNO

[3067] R. A. Mollin. Class number one and prime-producing quadratic polynomials revisited. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 41(3):328–334, September 1998. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Codeca:1998:EV

[3068] P. Codecà and M. Nair. Extremal values of Δ(x, N) = ∑_{{n<xN \atop (n,N)=1}} 1−xϕ(N).

Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 41(3):335–??, September 1998. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Tymchatyn:1998:CCD

[3069] E. D. Tymchatyn and Chang-Cheng Yang. Characterizing continua by disconnection properties. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 41(3):348–??, September 1998. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

VanOystaeyen:1998:EHA

[3070] Fred Van Oystaeyen and Yinhuo Zhang. Embedding the Hopf automorphism group into the Brauer group. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 41(3):359–??, September 1998. CO-DEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Moskowitz:1998:ECR

[3071] Martin Moskowitz and Michael Wüstner. Exponentiality of certain real solvable Lie groups. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 41(3):368–??, September 1998. CODEN CM-BUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Young:1998:NIL

[3072] Carmen M. Young. Normal invariants of lens spaces. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 41(3):374–??, September 1998. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Burns:1998:IBI

[3073] John Burns and Graham Ellis. Inequalities for Baer invariants of finite groups. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 41(4):385–??, December 1998. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Daly:1998:NML

[3074] James E. Daly and Keith L. Phillips. A note on H^{1} multipliers for locally compact vilenkin groups. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 41(4):392–??, December 1998. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Dziubanski:1998:BLW

[3075] Jacek Dziubański and Eugenio Hernández. Band-limited wavelets with subexponential decay. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 41(4):398–??, December 1998. CODEN CM-BUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Al-Hasan:1998:BSI

[3076] Abdelnaser J. Al-Hasan and Dashan Fan. L^{p}-boundedness of a singular integral operator. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 41(4):404–??, December 1998. CODEN CM-BUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Llorens-Fuster:1998:FPP

[3077] Enrique Llorens-Fuster and Brailey Sims. The fixed point property in c_{0}. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 41(4):413–??, December 1998. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Long:1998:FPA

[3078] D. D. Long and A. W. Reid. Free products with amalgamation and p-adic Lie groups. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 41(4):423–??, December 1998. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Mascioni:1998:LMF

[3079] Vania Mascioni and Lajos Molnár. Linear maps on factors which preserve the extreme points of the unit ball. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 41(4):434–??, December 1998. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Chamberland:1998:MPJ

[3080] Marc Chamberland and Gary Meisters. A mountain pass to the Jacobian conjecture. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 41(4):442–??, December 1998. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Bresar:1998:DAP

[3081] Matej Bresar, W. S. Martindale 3rd, and C. Robert Miers. Dependent automorphisms in prime rings. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 41(4):452–??, December 1998. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Moran:1998:RRR

[3082] Alan Moran. The right regular representation of a compact right topological group. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 41(4):463–??, December 1998. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Muller:1998:SSH

[3083] Jürgen Müller and Jochen Wengenroth. Separating singularities of holomorphic functions. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 41(4):473–??, December 1998. CODEN CM-BUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Oberlin:1998:CMC

[3084] Daniel M. Oberlin. Convolution with measures on curves in R^{3}. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 41(4): 478–??, December 1998. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Parmenter:1998:PRG

[3085] M. M. Parmenter, E. Spiegel, and P. N. Stewart. The periodic radical of group rings and incidence algebras. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 41(4):481–??, December 1998. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Sun:1998:RCM

[3086] Heng Sun. Remarks on certain metaplectic groups. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 41(4):488–??, December 1998. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Borwein:1998:CHP

[3087] J. M. Borwein, R. Girgensohn, and Xianfu Wang. On the construction of Hölder and proximal subderivatives. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 41(4):497–??, December 1998. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Anonymous:1998:AII

[3088] Anonymous. Author index — index des auteurs — for 1998 — pour 1998. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 41(4):508–??, December 1998. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Liu:1999:BMS

[3089] Peide Liu, Eero Saksman, and Hans-Olav Tylli. Boundedness of the qmean-square operator on vector-valued analytic martingales. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 42(??): 221–230, ???? 1999. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Beauzamy:1999:HRP

[3090] Bernard Beauzamy. How the roots of a polynomial vary with its coefficients: a local quantitative result. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 42(1):3–12, March 1999. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Brendle:1999:DPS

[3091] Jörg Brendle. Dow’s principle and Q-sets. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 42(1):13–24, March 1999. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Brown:1999:SCD

[3092] Tom C. Brown, Ronald L. Graham, and Bruce M. Landman. On the set of common differences in van der Waerden’s theorem on arithmetic progressions. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 42(1):25–36, March 1999. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Christensen:1999:OCR

[3093] Ole Christensen. Operators with closed range, pseudo-inverses, and perturbation of frames for a subspace. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 42(1):37–45, March 1999. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Dijkstra:1999:GPT

[3094] Jan J. Dijkstra. Generic partial two-point sets are extendable. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 42(1): 46–51, March 1999. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Edmonds:1999:ECB

[3095] Allan L. Edmonds. Embedding coverings in bundles. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 42(1):52–55, March 1999. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Elbert:1999:SFZ

[3096] Árpád Elbert and Panayiotis D. Siafarikas. On the square of the first zero of the Bessel function J_{ν}(z). Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 42(1):56–77, March 1999. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Gittenberger:1999:MSD

[3097] Bernhard Gittenberger and Jörg M. Thuswaldner. The moments of the sum-of-digits function in number fields. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 42(1):68–77, March 1999. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Gonzalez:1999:FJP

[3098] Josep González. Fermat Jacobians of prime degree over finite fields. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 42(1):78–86, March 1999. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Kittaneh:1999:SNI

[3099] Fuad Kittaneh. Some norm inequalities for operators. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 42(1):87–96, March 1999. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Kwon:1999:AFB

[3100] E. G. Kwon. On analytic functions of Bergman BMO in the ball. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 42(1): 97–103, March 1999. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Nikolskaia:1999:IVP

[3101] Ludmila Nikolskaia. Instabilité de vecteurs propres d’opérateurs linéaires. (French) [Instability of eigenvectors of linear operators]. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 42(1):104– 117, March 1999. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Rao:1999:PWN

[3102] T. S. S. R. K. Rao. Points of weak^{∗}-norm continuity in the unit ball of the space WC(K, X)^{∗}. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 42(1):118–124, March 1999. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Smith:1999:MVI

[3103] Larry Smith. Modular vector invariants of cyclic permutation representations. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 42(1):125–128, March 1999. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Baker:1999:HOA

[3104] Andrew Baker. Hecke operations and the Adams E_{2}-term based on elliptic cohomology. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 42(2):129–138, June 1999. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Bonet:1999:ENW

[3105] José Bonet, Paweł Domański, and Mikael Lindström. Essential norm and weak compactness of composition operators on weighted Banach spaces of analytic functions. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 42(2):139–148, June 1999. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Boyer:1999:NFD

[3106] S. Boyer and X. Zhang. A note on finite Dehn fillings. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 42(2):149– 154, June 1999. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Campbell:1999:NCM

[3107] H. E. A. Campbell, A. V. Geramita, I. P. Hughes, R. J. Shank, and D. L. Wehlau. Non–Cohen-Macaulay vector invariants and a Noether bound for a Gorenstein ring of invariants. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 42(2):155–161, June 1999. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Cobos:1999:LSC

[3108] Fernando Cobos and Thomas Kühn. Lorentz-Schatten classes and pointwise domination of matrices. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 42(2):162–168, June 1999. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Ding:1999:HKL

[3109] Hongming Ding. Heat kernels of Lorentz cones. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 42(2):169–173, June 1999. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Ferrero:1999:RC

[3110] Miguel Ferrero and Alveri Sant’Ana. Rings with comparability. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 42(2): 174–183, June 1999. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Fiorito:1999:AMS

[3111] Giovanni Fiorito. On arithmetic means of sequences generated by a periodic function. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 42(2):184–189, June 1999. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Gilmer:1999:TQF

[3112] Patrick M. Gilmer. Topological quantum field theory and strong shift equivalence. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 42(2):190–197, June 1999. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Guadalupe:1999:CAD

[3113] José J. Guadalupe, Mario Pérez, and Juan L. Varona. Commutators and analytic dependence of Fourier–Bessel series on (0, ∞). Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 42(2): 198–208, June 1999. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Lanteri:1999:AVB

[3114] Antonio Lanteri and Hidetoshi Maeda. Ample vector bundles of curve genus one. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 42(2):209–213, June 1999. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Paeng:1999:CRS

[3115] Seong-Hun Paeng and Jong-Gug Yun. Conjugate radius and sphere theorem. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 42(2):214–220, June 1999. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Peide:1999:BMS

[3116] Liu Peide, Eero Saksman, and Hans-Olav Tylli. Boundedness of the qmean-square operator on vector-valued analytic martingales. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 42(2): 221–230, June 1999. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Rush:1999:GIR

[3117] David E. Rush. Generating ideals in rings of integer-valued polynomials. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 42(2):231–236, June 1999. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Thompson:1999:BDA

[3118] A. C. Thompson. On Benson’s definition of area in Minkowski space. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 42(2):237–247, June 1999. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Weber:1999:CBC

[3119] Christian Weber. The classification of Pin_{4}-bundles over a 4-complex. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 42(2):248–256, June 1999. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Austin:1999:HKA

[3120] David Austin and Dale Rolfsen. Homotopy of knots and the Alexander polynomial. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 42(3):257–262, September 1999. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Choie:1999:MTM

[3121] Youngju Choie and Min Ho Lee. Mellin transforms of mixed cusp forms. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 42(3):263–273, September 1999. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Dadarlat:1999:BMN

[3122] Marius Dădărlat and Søren Eilers. The Bockstein map is necessary. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 42(3):274–284, September 1999. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Deng:1999:KSO

[3123] Peiming Deng. On Kloosterman sums with oscillating coefficients. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 42(3):285–290, September 1999. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Grubb:1999:SQM

[3124] D. J. Grubb and Tim LaBerge. Spaces of quasi-measures. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 42(3): 291–297, September 1999. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Jespers:1999:SAM

[3125] Eric Jespers and Jan Okniński. Semigroup algebras and maximal orders. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 42(3):298–306, September 1999. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Kapovich:1999:MSS

[3126] Michael Kapovich and John J. Millson. On the moduli space of a spherical polygonal linkage. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 42(3):307–320, September 1999. CODEN CM-BUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Kikuchi:1999:AOM

[3127] Masato Kikuchi. Averaging operators and martingale inequalities in rearrangement invariant function spaces. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 42(3):321–334, September 1999. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Kim:1999:CSS

[3128] Goansu Kim and C. Y. Tang. Cyclic subgroup separability of HNNextensions with cyclic associated subgroups. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 42(3):335–343, September 1999. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Koldobsky:1999:PDD

[3129] Alexander Koldobsky. Positive definite distributions and subspaces of L_{p} with applications to stable processes. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 42(3):344–353, September 1999. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Marshall:1999:RHR

[3130] Murray A. Marshall. A real holomorphy ring without the Schmüdgen property. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 42(3):354–358, September 1999. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Martin:1999:GRB

[3131] W. J. Martin and D. R. Stinson. A generalized Rao bound for ordered orthogonal arrays and (t, m, s)-nets. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 42(3):359–370, September 1999. CO-DEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Marubayashi:1999:PPI

[3132] H. Marubayashi and A. Ueda. Prime and primary ideals in a Prüfer order in a simple Artinian ring with finite dimension over its center. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 42(3): 371–379, September 1999. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Nurmela:1999:ABO

[3133] Kari J. Nurmela, Patric R. J. Östergård, and Rainer aus dem Spring. Asymptotic behavior of optimal circle packings in a square. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 42(3): 380–385, September 1999. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Polat:1999:MS

[3134] Norbert Polat. Minimal separators. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 42(3):386–392, September 1999. CO-DEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Savin:1999:CSR

[3135] Gordan Savin. A class of supercuspidal representations of G_{2}(k). Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 42(3): 393–400, September 1999. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Swain:1999:LDP

[3136] Gordon A. Swain and Philip S. Blau. Lie derivations in prime rings with involution. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 42(3):401–411, September 1999. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Tai:1999:PD

[3137] Yung-Sheng Tai. Peirce domains. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 42(3):412–416, September 1999. CO-DEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Aziz-Ul-Auzeem:1999:SPR

[3138] Abdul Aziz-Ul-Auzeem and B. A. Zarger. Some properties of rational functions with prescribed poles. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 42(4):417–426, December 1999. CO-DEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Berndt:1999:RMI

[3139] Bruce C. Berndt and Heng Huat Chan. Ramanujan and the modular j-invariant. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 42(4):427–440, December 1999. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Berrizbeitia:1999:PBR

[3140] P. Berrizbeitia and P. D. T. A. Elliott. Product bases for the rationals. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 42(4):441–444, December 1999. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Bochnak:1999:SMR

[3141] J. Bochnak and W. Kucharz. Smooth maps and real algebraic morphisms. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 42(4):445–451, December 1999. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Bradley:1999:FRO

[3142] Sean Bradley. Finite rank operators in certain algebras. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 42(4):452– 462, December 1999. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Hofmann:1999:GCC

[3143] Steve Hofmann, Xinwei Li, and Dachun Yang. A generalized characterization of commutators of parabolic singular integrals. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 42(4):463–477, December 1999. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Pruss:1999:RMA

[3144] Alexander R. Pruss. A remark on the Moser-Aubin inequality for axially symmetric functions on the sphere. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 42(4):478–485, December 1999. CO-DEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Sawyer:1999:SF

[3145] P. Sawyer. Spherical functions on SO_{0}(p, q)/SO(p) × SO(q). Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 42(4): 486–498, December 1999. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Zaharia:1999:CSI

[3146] Alexandru Zaharia. Characterizations of simple isolated line singularities. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 42(4):499–506, December 1999. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Anonymous:1999:AII

[3147] Anonymous. Author index — index des auteurs — for 1999 — pour 1999. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 42(4):507–??, December 1999. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Adin:2000:RAL

[3148] Ron Adin and David Blanc. Resolutions of associative and Lie algebras. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 43(1):3–16, March 2000. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Bak:2000:MPC

[3149] Jong-Guk Bak. Multilinear proofs for convolution estimates for degenerate plane curves. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 43(1):17–20, March 2000. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Barnes:2000:CAB

[3150] Bruce A. Barnes. The commutant of an abstract backward shift. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 43(1): 21–24, March 2000. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Bounkhel:2000:SRD

[3151] M. Bounkhel and L. Thibault. Subdifferential regularity of directionally Lipschitzian functions. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 43(1):25–36, March 2000. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Bousaidi:2000:MSR

[3152] M. A. Bousaidi. Multiplicative structure of the ring K(S(T^{∗}RP^{2n+1})). Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 43(1):37–46, March 2000. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Bozicevic:2000:PLG

[3153] Mladen Božičević. A property of Lie group orbits. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 43(1):47–50, March 2000. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Edward:2000:EDN

[3154] Julian Edward. Eigenfunction decay for the Neumann Laplacian on hornlike domains. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 43(1):51–59, March 2000. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Farkas:2000:TUG

[3155] Daniel R. Farkas and Peter A. Linnell. Trivial units in group rings. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 43(1):60–62, March 2000. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Iosevich:2000:SRK

[3156] Alex Iosevich and Guozhen Lu. Sharpness results and Knapp’s homogeneity argument. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 43(1):63–68, March 2000. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Kaminker:2000:TIS

[3157] Jerome Kaminker and Vicumpriya Perera. Type II spectral flow and the eta invariant. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 43(1):69–73, March 2000. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Kimura:2000:GMI

[3158] Makoto Kimura and Sadahiro Maeda. Geometric meaning of isoparametric hypersurfaces in a real space form. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 43(1):74–78, March 2000. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Konig:2000:CSA

[3159] Steffen König. Cyclotomic Schur algebras and blocks of cyclic defect. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 43(1):79–86, March 2000. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Lindstrom:2000:LTB

[3160] Mikael Lindström and Georg Schlüchtermann. Lomonosov’s techniques and Burnside’s Theorem. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 43(1):87–89, March 2000. CODEN CM-BUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Muic:2000:CSH

[3161] Goran Muić and Gordan Savin. Complementary series for Hermitian quaternionic groups. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 43(1):90–99, March 2000. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Okon:2000:GRL

[3162] James S. Okon and J. Paul Vicknair. A Gorenstein ring with larger Dilworth number than Sperner number. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 43(1):100–104, March 2000. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Overholt:2000:SUU

[3163] Marius Overholt. Sets of uniqueness for univalent functions. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 43(1): 105–107, March 2000. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Sanders:2000:ECC

[3164] Daniel P. Sanders and Yue Zhao. On the entire coloring conjecture. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 43(1):108–114, March 2000. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Schaller:2000:PNE

[3165] Paul Schmutz Schaller. Perfect non-extremal Riemann surfaces. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 43(1): 115–125, March 2000. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Soto:2000:ACM

[3166] José J. M. Soto. Sur l’annulation de certains modules de cohomologie d’André–Quillen. (French) [On the nullification of certain cohomology modules of d’André–Quillen]. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 43(1):126–128, March 2000. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Ballico:2000:MSR

[3167] E. Ballico. Maximal subbundles of rank 2 vector bundles on projective curves. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 43(2):129–137, June 2000. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Boyd:2000:ELN

[3168] C. Boyd. Exponential laws for the Nachbin ported topology. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 43(2): 138–144, June 2000. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Chang:2000:PVL

[3169] Jae-Ho Chang, Sang Youl Lee, and Chan-Young Park. On the 2-parallel versions of links. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 43(2):145–156, June 2000. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

ElAbdalaoui:2000:LCO

[3170] El Houcein El Abdalaoui. A larger class of Ornstein transformations with mixing property. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 43(2):157–161, June 2000. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Foth:2000:MSP

[3171] Philip Foth. Moduli spaces of polygons and punctured Riemann spheres. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 43(2):162–173, June 2000. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Gantz:2000:SPB

[3172] Christian Gantz and Brian Steer. Stable parabolic bundles over elliptic surfaces and over Riemann surfaces. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 43(2):174–182, June 2000. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Ionesei:2000:GTP

[3173] Gheorghe Ionesei. A gauge theoretic proof of the Abel–Jacobi theorem. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 43(2):183–192, June 2000. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Magajna:2000:CNR

[3174] Bojan Magajna. C^{∗}-convexity and the numerical range. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 43(2):193– 207, June 2000. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Matouskova:2000:ECL

[3175] Eva Matoušková. Extensions of continuous and Lipschitz functions. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 43 (2):208–217, June 2000. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Mollin:2000:CFJ

[3176] R. A. Mollin and A. J. van der Poorten. Continued fractions, Jacobi symbols, and quadratic Diophantine equations. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 43(2):218–225, June 2000. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Neisendorfer:2000:JHI

[3177] Joseph Neisendorfer. James–Hopf invariants, Anick’s spaces, and the double loops on odd primary Moore spaces. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 43(2):226–235, June 2000. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Voloch:2000:QB

[3178] José Felipe Voloch. On a question of Buium. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 43(2):236–238, June 2000. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Yu:2000:NDQ

[3179] Gang Yu. On the number of divisors of the quadratic form m^{2} + n^{2}. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 43(2):239–256, June 2000. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Androulakis:2000:SMW

[3180] George Androulakis, Peter G. Casazza, and Denka N. Kutzarova. Some more weak Hilbert spaces. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 43(3):257–267, September 2000. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Bogley:2000:CPT

[3181] W. A. Bogley, N. D. Gilbert, and James Howie. Cockcroft properties of Thompson’s group. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 43(3):268–281, September 2000. CODEN CM-BUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Boston:2000:CGR

[3182] Nigel Boston and David T. Ose. Characteristic p Galois representations that arise from Drinfeld modules. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 43(3):282–293, September 2000. CO-DEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Bracci:2000:FPC

[3183] Filippo Bracci. Fixed points of commuting holomorphic maps without boundary regularity. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 43(3):294–303, September 2000. CODEN CM-BUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Darmon:2000:CHM

[3184] Henri Darmon and Jean-François Mestre. Courbes hyperelliptiques à multiplications réelles et une construction de Shih. (French) [Hyperelliptic curves with real multiplication and a construction of Shih]. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 43(3): 304–311, September 2000. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Dobbs:2000:PIC

[3185] David E. Dobbs. On the prime ideals in a commutative ring. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 43(3): 312–319, September 2000. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Elliott:2000:CCA

[3186] George Elliott and Igor Fulman. On classification of certain C^{∗}-algebras. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 43(3):320–329, September 2000. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Hare:2000:MOC

[3187] Kathryn E. Hare. Maximal operators and Cantor sets. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 43(3):330–342, September 2000. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Hughes:2000:CHN

[3188] Bruce Hughes, Larry Taylor, and Bruce Williams. Controlled homeomorphisms over nonpositively curved manifolds. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 43(3):343–354, September 2000. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Kelly:2000:DFW

[3189] Brian P. Kelly. A dimension-free weak-type estimate for operators on UMD-valued functions. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 43(3):355–361, September 2000. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Kim:2000:EHF

[3190] Hwankoo Kim. Examples of half-factorial domains. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 43(3):362–367, September 2000. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Litvak:2000:KKI

[3191] A. E. Litvak. Kahane–Khinchin’s Inequality for quasi-norms. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 43(3): 368–379, September 2000. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Shahidi:2000:TGC

[3192] Freydoon Shahidi. Twists of a general class of L-functions by highly ramified characters. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 43(3):380–384, September 2000. CODEN CM-BUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Bluskov:2000:ICC

[3193] I. Bluskov, M. Greig, and K. Heinrich. Infinite classes of covering numbers. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 43(4):385–396, December 2000. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Bonato:2000:TOP

[3194] Anthony Bonato, Peter Cameron, and Dejan Delić. Tournaments and orders with the pigeonhole property. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 43(4):397–405, December 2000. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Borwein:2000:WMO

[3195] David Borwein. Weighted mean operators on l_{p}. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 43(4):406–412, December 2000. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Chatters:2000:NIM

[3196] A. W. Chatters. Non-isomorphic maximal orders with isomorphic matrix rings. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 43(4):413–417, December 2000. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Gong:2000:OSU

[3197] Guihua Gong, Xinhui Jiang, and Hongbing Su. Obstructions to Zstability for unital simple C^{∗}-algebras. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 43(4):418–426, December 2000. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Ivey:2000:HHS

[3198] Thomas A. Ivey. Helices, Hasimoto surfaces and Bäcklund transformations. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 43(4):427–439, December 2000. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Koufogiorgos:2000:ENC

[3199] Themis Koufogiorgos and Charalambos Tsichlias. On the existence of a new class of contact metric manifolds. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 43(4):440–447, December 2000. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Li:2000:NGN

[3200] Chi-Kwong Li and Alexandru Zaharia. Nonconvexity of the generalized numerical range associated with the principal character. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 43(4): 448–458, December 2000. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Ndogmo:2000:PIS

[3201] J. C. Ndogmo. Properties of the invariants of solvable Lie algebras. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 43(4):459–471, December 2000. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Oberlin:2000:ERX

[3202] Daniel M. Oberlin. An estimate for a restricted X-ray transform. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 43(4):

472–476, December 2000. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Ward:2000:DLS

[3203] Michael J. Ward. The dynamics of localized solutions of nonlocal reaction-diffusion equations. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 43(4):477–495, December 2000. CO-DEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Xu:2000:HPA

[3204] Yuan Xu. Harmonic polynomials associated with reflection groups. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 43(4):496–507, December 2000. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Anonymous:2000:AII

[3205] Anonymous. Author index — index des auteurs — for 2000 — pour 2000. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 43(4):508–??, December 2000. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Alexandru:2001:GD

[3206] Victor Alexandru, Nicolae Popescu, and Alexandru Zaharescu. The generating degree of C_{p}. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 44(1):3–11, March 2001. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Anisca:2001:TSS

[3207] Razvan Anisca and Monica Ilie. A technique of studying sums of central Cantor sets. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 44(1):12–18, March 2001. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Brindza:2001:MBR

[3208] B. Brindza, Á. Pintér, and W. M. Schmidt. Multiplicities of binary recurrences. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 44(1):19–21, March 2001. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Evans:2001:GSO

[3209] Ronald Evans. Gauss sums of orders six and twelve. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 44(1):22–26, March 2001. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Goodaire:2001:NSI

[3210] Edgar G. Goodaire and César Polcino Milies. Normal subloops in the integral loop ring of an RA loop. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 44(1):27–35, March 2001. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Kapovich:2001:QBD

[3211] Michael Kapovich and John J. Millson. Quantization of bending deformations of polygons in E^{3}, hypergeometric integrals and the gassner representation. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 44(1):36–60, March 2001. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Kats:2001:IPI

[3212] B. A. Kats. The inequalities for polynomials and integration over fractal arcs. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 44(1):61–69, March 2001. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Lempert:2001:TBA

[3213] László Lempert and Róbert Szőke. The tangent bundle of an almost complex manifold. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 44(1):70–79, March 2001. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Levin:2001:CCD

[3214] Michael Levin. Constructing compacta of different extensional dimensions. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 44(1):80–86, March 2001. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Lieman:2001:NES

[3215] Daniel Lieman and Igor Shparlinski. On a new exponential sum. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 44(1): 87–92, March 2001. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Neumann:2001:SSL

[3216] B. H. Neumann. Some semigroup laws in groups. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 44(1):93–96, March 2001. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Ou:2001:DCQ

[3217] Zhiming M. Ou and Kenneth S. Williams. On the density of cyclic quartic fields. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 44(1):97–104, March 2001. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Pilipovic:2001:CEP

[3218] Stevan Pilipović. Convolution equation in S^{′∗} — propagation of singularities. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 44(1):105–114, March 2001. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Roy:2001:AAS

[3219] Damien Roy. Approximation algébrique simultanée de nombres de Liouville. (French) [Simultaneous algebraic approximation for Liouville numbers]. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 44(1):115–120, March 2001. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Wojciechowski:2001:NCM

[3220] Michal Wojciechowski. A necessary condition for multipliers of weak type (1, 1). Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 44(1):121–125, March 2001. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Zeron:2001:ECR

[3221] E. Santillan Zeron. Each copy of the real line in C^{2} is removable. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 44(1): 126–128, March 2001. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Curras-Bosch:2001:LSD

[3222] Carlos Currás-Bosch. Linéarisation symplectique en dimension 2. (French) [Simplectic linearization in dimension 2]. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 44(2):129–139, June 2001. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Gotay:2001:QNS

[3223] Mark J. Gotay and Janusz Grabowski. On quantizing nilpotent and solvable basic algebras. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 44(2):140–149, June 2001. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Jakobczak:2001:ESS

[3224] Piotr Jakóbczak. Exceptional sets of slices for functions from the Bergman space in the ball. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 44(2):150–159, June 2001. CODEN CM-BUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Langlands:2001:TFA

[3225] Robert P. Langlands. The trace formula and its applications: An introduction to the work of James Arthur. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 44(2):160–209, June 2001. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Leung:2001:GEP

[3226] Man Chun Leung. Growth estimates on positive solutions of the equation Δu+Ku^{\frac{n+2}{n−2}} = 0 in R^{n}. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 44(2):210–222, June 2001. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Marshall:2001:EAP

[3227] M. Marshall. Extending the Archimedean Positivstellensatz to the noncompact case. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 44(2):223–230, June 2001. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Rosenblatt:2001:WCS

[3228] Joseph M. Rosenblatt and George A. Willis. Weak convergence is not strong convergence for amenable groups. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 44(2):231–241, June 2001. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Schueller:2001:ZFP

[3229] Laura Mann Schueller. The zeta function of a pair of quadratic forms. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 44(2):242–256, June 2001. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Abanades:2001:AHR

[3230] Miguel A. Abánades. Algebraic homology for real hyperelliptic and real projective ruled surfaces. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 44(3):257–265, September 2001. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Cencelj:2001:EMN

[3231] M. Cencelj and A. N. Dranishnikov. Extension of maps to nilpotent spaces. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 44(3):266–269, September 2001. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Cheung:2001:LOP

[3232] Wai-Shun Cheung and Chi-Kwong Li. Linear operators preserving generalized numerical ranges and radii on certain triangular algebras of matrices. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 44(3):270–281, September 2001. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Lee:2001:HOJ

[3233] Min Ho Lee and Hyo Chul Myung. Hecke operators on Jacobi-like forms. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 44(3):282–291, September 2001. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

McKay:2001:ANT

[3234] Angela McKay. An analogue of Napoleon’s Theorem in the hyperbolic plane. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 44(3):292–297, September 2001. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Muic:2001:PCS

[3235] Goran Muić. A proof of Casselman–Shahidi’s Conjecture for quasi-split classical groups. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 44(3):298–312, September 2001. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Reverter:2001:IMG

[3236] Amadeu Reverter and Núria Vila. Images of mod p Galois representations associated to elliptic curves. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 44(3):313–322, September 2001. CO-DEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Schuman:2001:CHP

[3237] Bertrand Schuman. Une classe d’hamiltoniens polynomiaux isochrones. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 44(3):323–334, September 2001. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Stacey:2001:ILT

[3238] P. J. Stacey. Inductive limit toral automorphisms of irrational rotation algebras. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 44(3):335–336, September 2001. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Vinet:2001:STL

[3239] Luc Vinet and Alexei Zhedanov. Spectral transformations of the Laurent biorthogonal polynomials, II. Pastro polynomials. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 44(3):337–345, September 2001. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Wang:2001:PSS

[3240] Wei Wang. Positive solution of a subelliptic nonlinear equation on the Heisenberg group. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 44(3):346–354, September 2001. CODEN CM-BUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Weaver:2001:HBI

[3241] Nik Weaver. Hilbert bimodules with involution. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 44(3):355–369, September 2001. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Weston:2001:LI

[3242] Anthony Weston. On locating isometric ℓ_{1}^{(n)}. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 44(3):370–375, September 2001. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Zhang:2001:NHF

[3243] Xi Zhang. A note on p-harmonic 1-forms on complete manifolds. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 44(3): 376–384, September 2001. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Ballantine:2001:HCP

[3244] Cristina M. Ballantine. A hypergraph with commuting partial Laplacians. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 44(4):385–397, December 2001. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Cardon:2001:ECG

[3245] David A. Cardon and M. Ram Murty. Exponents of class groups of quadratic function fields over finite fields. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 44(4):398–407, December 2001. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Falbel:2001:FGG

[3246] E. Falbel. Finite groups generated by involutions on Lagrangian planes of C^{2}. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 44(4):408–419, December 2001. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Gauthier:2001:AMF

[3247] P. M. Gauthier and M. R. Pouryayevali. Approximation by meromorphic functions with Mittag-Leffler type constraints. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 44(4):420–428, December 2001. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Henniger:2001:ERN

[3248] J. P. Henniger. Ergodic rotations of nilmanifolds conjugate to their inverses. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 44(4):429–439, December 2001. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Hironaka:2001:LPP

[3249] Eriko Hironaka. The Lehmer polynomial and pretzel links. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 44(4): 440–451, December 2001. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic). See erratum [3294].

Ishihara:2001:SAP

[3250] Hironobu Ishihara. Some adjunction properties of ample vector bundles. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 44(4):452–458, December 2001. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Kahl:2001:LCA

[3251] Thomas Kahl. LS-catégorie algébrique et attachement de cellules. (French) [Algebraic LS category and cellular attachment]. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 44(4):459– 468, December 2001. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Marcoux:2001:SPW

[3252] Laurent W. Marcoux. Sums and products of weighted shifts. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 44(4): 469–481, December 2001. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Mezo:2001:MWO

[3253] Paul Mezo. Matching of weighted orbital integrals for metaplectic correspondences. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 44(4):482–490, December 2001. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Wang:2001:RSN

[3254] Weiqiang Wang. Resolution of singularities of null cones. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 44(4):491– 503, December 2001. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Zhang:2001:WAC

[3255] Yong Zhang. Weak amenability of a class of Banach algebras. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 44(4): 504–508, December 2001. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Anonymous:2001:AII

[3256] Anonymous. Author index — index des auteurs — for 2001 — pour 2001. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 44(4):509–??, December 2001. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Allison:2002:CTG

[3257] Bruce Allison and Oleg Smirnov. Coordinatization theorems for graded algebras. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 45(4):451–465, December 2002. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Anonymous:2002:AII

[3258] Anonymous. Author index — index des auteurs — for 2002 — pour 2002. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 45(4):732–??, December 2002. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Ardizzone:2002:SFC

[3259] Lucia Ardizzone, Renata Grimaldi, and Pierre Pansu. Sur les singularités de la fonction croissance d’une variété non simplement connexe. (French) [On the singularities of the growth function of a nonsimply-connected variety]. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 45(2):161–167, June 2002. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Arthur:2002:NAL

[3260] James Arthur. A note on the Automorphic Langlands Group. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 45(4): 466–482, December 2002. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Azagra:2002:RAN

[3261] D. Azagra and T. Dobrowolski. Real-analytic negligibility of points and subspaces in Banach spaces, with applications. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 45(1):3–10, March 2002. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Baake:2002:DWL

[3262] Michael Baake. Diffraction of weighted lattice subsets. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 45(4):483– 498, December 2002. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Bahturin:2002:GGM

[3263] Yu. A. Bahturin and M. V. Zaicev. Group gradings on matrix algebras. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 45(4):499–508, December 2002. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Bahturin:2002:PCH

[3264] Yuri Bahturin, Mikhail Kochetov, and Susan Montgomery. Polycharacters of cocommutative Hopf algebras. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 45(1):11–24, March 2002. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Benkart:2002:LSG

[3265] Georgia Benkart and Alberto Elduque. Lie superalgebras graded by the root systems C(n), D(m, n), D(2, 1; α), F (4), G(3). Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 45(4):509–524, December 2002. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Berman:2002:SFU

[3266] Stephen Berman, Jun Morita, and Yoji Yoshii. Some factorizations in universal enveloping algebras of three dimensional Lie algebras and generalizations. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 45(4):525–536, December 2002. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Bloom:2002:EDM

[3267] Steven Bloom and Ron Kerman. Extrapolation of L^{p} data from a modular inequality. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 45(1):25–35, March 2002. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Brenken:2002:AIG

[3268] Berndt Brenken. C^{∗}-algebras of infinite graphs and Cuntz–Krieger algebras. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 45(3):321–336, September 2002. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Byott:2002:BEO

[3269] Nigel P. Byott and G. Griffith Elder. Biquadratic extensions with one break. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 45(2):168–179, June 2002. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Chapoton:2002:PRG

[3270] Frédéric Chapoton, Sergey Fomin, and Andrei Zelevinsky. Polytopal realizations of generalized associahedra. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 45(4):537–566, December 2002. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Chen:2002:SRA

[3271] Imin Chen. Surjectivity of modℓ representations attached to elliptic curves and congruence primes. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 45(3):337–348, September 2002. CO-DEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Connolly:2002:EGV

[3272] Francis X. Connolly and Stratos Prassidis. On the exponent of the NK_{0}- groups of virtually infinite cyclic groups. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 45(2):180–195, June 2002. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Coppens:2002:VAL

[3273] Marc Coppens. Very ample linear systems on blowings-up at general points of projective spaces. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 45(3):349–354, September 2002. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Cresson:2002:OLC

[3274] Jacky Cresson. Obstruction à la linéarisation des champs de vecteurs polynomiaux. (French) [Obstruction in the linearization of vector polynomial fields]. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 45(3):355–363, September 2002. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Cummins:2002:MED

[3275] C. J. Cummins. Modular equations and discrete, genus-zero subgroups of SL(2, R) containing Γ(N). Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 45(1):36–45, March 2002. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Dafni:2002:LWC

[3276] Galia Dafni. Local VMO and weak convergence in h^{1}. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 45(1):46–59, March 2002. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Deitmar:2002:MTW

[3277] Anton Deitmar. Mellin transforms of Whittaker functions. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 45(3):

364–377, September 2002. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

deSalas:2002:RGC

[3278] Fernando Sancho de Salas. Residue: a geometric construction. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 45(??): 284–293, ???? 2002. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

DeSole:2002:SJP

[3279] Alberto De Sole and Victor G. Kac. Subalgebras of gc_{N} and Jacobi polynomials. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 45(4):567–622, December 2002. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic). Dedicated to Robert Moody on his 60th birthday.

Dranishnikov:2002:UEH

[3280] A. N. Dranishnikov, G. Gong, V. Lafforgue, and G. Yu. Uniform embeddings into Hilbert space and a question of Gromov. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 45(1):60–70, March 2002. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Dubickas:2002:MMC

[3281] Arturas Dubickas. Mahler measures close to an integer. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 45(2):196– 203, June 2002. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Fakhruddin:2002:CGS

[3282] Najmuddin Fakhruddin. On the Chow groups of supersingular varieties. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 45(2):204–212, June 2002. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Fernandez-Lopez:2002:LME

[3283] Manuel Fernández-López, Eduardo García-Río, and Demir N. Kupeli. The local Möbius equation and decomposition theorems in Riemannian geometry. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 45(3):378–387, September 2002. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

SanchodeSalas:2002:RGC

[3284] Fernando Sancho de Salas. Residue: a geometric construction. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 45(2): 284–??, June 2002. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Gao:2002:FBR

[3285] Yun Gao. Fermionic and bosonic representations of the extended affine Lie algebra ˜}↕_{N} (C_{q}). Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 45(4):623–633, December 2002. CODEN CM-BUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Gauthier:2002:AAD

[3286] P. M. Gauthier and E. S. Zeron. Approximation on arcs and dendrites going to infinity in C^{n}. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 45(1):80–85, March 2002. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Gerth:2002:CFO

[3287] Frank Gerth III. On cyclic fields of odd prime degree p with infinite Hilbert p-class field towers. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 45(1):86–88, March 2002. CODEN CM-BUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Gille:2002:ASC

[3288] Philippe Gille. Algèbres simples centrales de degré 5 et E_{8}. (French) [Simple central algebras of degree 5 and E_{8}]. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 45(3):388–398, September 2002. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Gordon:2002:GGS

[3289] B. Brent Gordon and Kirti Joshi. Griffiths groups of supersingular Abelian varieties. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 45(2):213–219, June 2002. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Grant:2002:GPF

[3290] David Grant. On Gunning’s prime form in genus 2. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 45(1):89–96, March 2002. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Haas:2002:IMN

[3291] Andrew Haas. Invariant measures and natural extensions. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 45(1):97–108, March 2002. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Hakim:2002:GDS

[3292] Jeffrey Hakim and Fiona Murnaghan. Globalization of distinguished supercuspidal representations of GL(n). Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 45(2):220–230, June 2002. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Hall:2002:DT

[3293] R. R. Hall and P. Shiu. The distribution of totatives. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 45(1):109– 114, March 2002. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Hironaka:2002:ELP

[3294] Eriko Hironaka. Erratum: “The Lehmer Polynomial and Pretzel Links”. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 45(2):231, June 2002. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic). See [3249].

Iakovlev:2002:SBI

[3295] Serguei Iakovlev. On the singular behavior of the inverse Laplace transforms of the functions \frac{I_{n}(s)}{sI′_{n}(s)}. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 45(3):399–416, September 2002. CO-DEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Ji:2002:SCI

[3296] Min Ji and Zhongmin Shen. On strongly convex indicatrices in Minkowski geometry. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 45(2):232–256, June 2002. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Kamiyama:2002:DCS

[3297] Yasuhiko Kamiyama and Shuichi Tsukuda. On deformations of the complex structure on the moduli space of spatial polygons. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 45(3):417–421, September 2002. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Lagarias:2002:LCD

[3298] Jeffrey C. Lagarias and Peter A. B. Pleasants. Local complexity of Delone sets and crystallinity. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 45(4):634–652, December 2002. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Ledet:2002:EDS

[3299] Arne Ledet. On the essential dimension of some semi-direct products. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 45(3):422–427, September 2002. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Lee:2002:MFA

[3300] Min Ho Lee. Modular forms associated to theta functions. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 45(2): 257–264, June 2002. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Luca:2002:NNZ

[3301] Florian Luca. The number of non-zero digits of n!. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 45(1):115–118, March 2002. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Marcelo:2002:GCC

[3302] Agustín Marcelo, Félix Marcelo, and César Rodríguez. The grade conjecture and the S_{2} condition. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 45(1):119–122, March 2002. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Martinez:2002:SJS

[3303] Consuelo Martínez and Efim Zelmanov. Specializations of Jordan superalgebras. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 45(4):653–671, December 2002. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Mollin:2002:CSS

[3304] R. A. Mollin. Criteria for simultaneous solutions of X^{2} − DY^{2} = c and x^{2} − Dy^{2} = −c. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 45(3):428–435, September 2002. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Moody:2002:UDM

[3305] Robert V. Moody. Uniform distribution in model sets. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 45(1):123– 130, March 2002. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Nawrocki:2002:SCD

[3306] Marek Nawrocki. On the Smirnov class defined by the maximal function. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 45(2):265–271, June 2002. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Neusel:2002:TIT

[3307] Mara D. Neusel. The transfer in the invariant theory of modular permutation representations II. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 45(2):272–283, June 2002. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Przytycki:2002:TIC

[3308] Józef H. Przytycki and Witold Rosicki. The topological interpretation of the core group of a surface in S^{4}. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 45(1):131–137, March 2002. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Rao:2002:NCR

[3309] S. Eswara Rao and Punita Batra. A new class of representations of EALA coordinated by quantum tori in two variables. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 45(4):672–685, December 2002. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Rauschning:2002:AIS

[3310] Jan Rauschning and Peter Slodowy. An aspect of icosahedral symmetry. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 45(4):686–696, December 2002. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Sawyer:2002:SFR

[3311] P. Sawyer. The spherical functions related to the root system B_{2}. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 45(3): 436–447, September 2002. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Sebbar:2002:MSF

[3312] Abdellah Sebbar. Modular subgroups, forms, curves and surfaces. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 45 (2):294–308, June 2002. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Sirvent:2002:PDS

[3313] V. F. Sirvent and B. Solomyak. Pure discrete spectrum for onedimensional substitution systems of Pisot type. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 45(4):697–710, December 2002. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Spearman:2002:DDQ

[3314] Blair K. Spearman and Kenneth S. Williams. The discriminant of a dihedral quintic field defined by a trinomial X^{5} +aX +b. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 45(1):138–153, March 2002. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

vandenDries:2002:IAP

[3315] Lou van den Dries and Franz-Viktor Kuhlmann. Images of additive polynomials in F_{q}((t)) have the optimal approximation property. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 45(1): 71–79, March 2002. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Weitsman:2002:PIS

[3316] Allen Weitsman. On the Poisson integral of step functions and minimal surfaces. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 45(1):154–160, March 2002. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Xia:2002:JMO

[3317] Jingbo Xia. Joint mean oscillation and local ideals in the Toeplitz algebra II: Local commutivity and essential commutant. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 45(2):309– 320, June 2002. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Yoshii:2002:CQT

[3318] Yoji Yoshii. Classification of quantum tori with involution. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 45(4): 711–731, December 2002. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Zhou:2002:ECL

[3319] Yiqiang Zhou. Erratum: “A Characterization of Left Perfect Rings”. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 45(3):448–450, September 2002. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic). See [2887].

Anderson:2003:CD

[3320] D. D. Anderson and Tiberiu Dumitrescu. Condensed domains. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 46(1):3– 13, March 2003. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Bahturin:2003:GCG

[3321] Yu. A. Bahturin and M. M. Parmenter. Generalized commutativity in group algebras. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 46(1):14–25, March 2003. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Bernardi:2003:RPR

[3322] D. Bernardi, E. Halberstadt, and A. Kraus. Remarques sur les points rationnels des variétés de Fermat. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 46(1):26–38, March 2003. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Bulow:2003:PRC

[3323] Tommy Bülow. Power residue criteria for quadratic units and the negative Pell equation. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 46(1):39–53, March 2003. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Cheung:2003:LMT

[3324] Wai-Shun Cheung and Chi-Kwong Li. Linear maps transforming the unitary group. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 46(1):54–58, March 2003. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Constantinescu:2003:NNI

[3325] T. Constantinescu and J. L. Johnson. A note on noncommutative interpolation. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 46(1):59–70, March 2003. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Cutter:2003:NFG

[3326] Pamela Cutter, Andrew Granville, and Thomas J. Tucker. The number of fields generated by the square root of values of a given polynomial. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 46(1):71–79, March 2003. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Erlijman:2003:MSB

[3327] Juliana Erlijman. Multi-sided braid type subfactors, II. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 46(1):80–94, March 2003. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Gauthier:2003:CRR

[3328] P. M. Gauthier. Cercles de remplissage for the Riemann zeta function. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 46(1):95–97, March 2003. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Larsen:2003:CPS

[3329] Nadia S. Larsen. Crossed products by semigroups of endomorphisms and groups of partial automorphisms. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 46(1):98–112, March 2003. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Lee:2003:PHC

[3330] Jaesung Lee and Kyung Soo Rim. Properties of the M-harmonic conjugate operator. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 46(1):113–121, March 2003. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Moon:2003:CFG

[3331] Myoungho Moon. On certain finitely generated subgroups of groups which split. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 46(1):122–129, March 2003. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Petersen:2003:FT

[3332] Peter Petersen and Frederick Wilhelm. On Frankel’s Theorem. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 46(1): 130–139, March 2003. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Renner:2003:ECD

[3333] Lex E. Renner. An explicit cell decomposition of the wonderful compactification of a semisimple algebraic group. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 46(1):140–148, March 2003. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Scherk:2003:RPC

[3334] John Scherk. The ramification polygon for curves over a finite field. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 46(1):149–156, March 2003. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Wieczorek:2003:TPC

[3335] Małgorzata Wieczorek. Torsion points on certain families of elliptic curves. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 46(1):157–160, March 2003. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Sanchez:2003:AQR

[3336] Félix Cabello Sánchez and Jesús M. F. Castillo. Answer to a question of S. Rolewicz. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 46(2):161–163, June 2003. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Dean:2003:CF

[3337] Andrew J. Dean. Classification of AF flows. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 46(2):164–177, June 2003. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Jaulent:2003:IIT

[3338] Jean-François Jaulent and Christian Maire. Sur les invariants d’Iwasawa des tours cyclotomiques. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 46(2):178–190, June 2003. CODEN CM-BUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Kim:2003:WTE

[3339] Yong-Cheol Kim. Weak type estimates of the maximal quasiradial Bochner–Riesz operator on certain Hardy spaces. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 46(2):191–203, June 2003. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Levy:2003:ROC

[3340] Jason Levy. Rationality and orbit closures. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 46(2):204–215, June 2003. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Li:2003:LMS

[3341] Chi-Kwong Li, Leiba Rodman, and Peter Semrl. Linear maps on selfadjoint operators preserving invertibility, positive definiteness, numerical range. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 46(2):216–228, June 2003. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Lin:2003:CNI

[3342] Ke-Pao Lin and Stephen S.-T. Yau. Counting the number of integral points in general n-dimensional tetrahedra and Bernoulli polynomials. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 46(2):229–241, June 2003. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Litvak:2003:ESD

[3343] A. E. Litvak and V. D. Milman. Euclidean sections of direct sums of normed spaces. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 46(2):242–251, June 2003. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Miyamoto:2003:BDS

[3344] Ikuko Miyamoto, Minoru Yanagishita, and Hidenobu Yoshida. Beurling– Dahlberg–Sjögren type theorems for minimally thin sets in a cone. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 46 (2):252–264, June 2003. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Oh:2003:RSK

[3345] Seungsang Oh. Reducing spheres and Klein bottles after Dehn fillings. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 46(2):265–267, June 2003. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Puls:2003:GCC

[3346] Michael J. Puls. Group cohomology and L^{p}-cohomology of finitely generated groups. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 46(2):268–276, June 2003. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Rochon:2003:RHA

[3347] Frédéric Rochon. Rigidity of Hamiltonian actions. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 46(2):277–290, June 2003. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Sankaran:2003:CTH

[3348] Parameswaran Sankaran. A coincidence theorem for holomorphic maps to G/P. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 46(2):291–298, June 2003. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Tomaszewski:2003:BBT

[3349] Witold Tomaszewski. A basis of Bachmuth type in the commutator subgroup of a free group. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 46(2):299–303, June 2003. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Traves:2003:LHS

[3350] William N. Traves. Localization of the Hasse–Schmidt algebra. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 46(2): 304–309, June 2003. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Wang:2003:SOD

[3351] Xiaofeng Wang. Second order Dehn functions of asynchronously automatic groups. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 46(2):310–320, June 2003. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Ballico:2003:DSC

[3352] E. Ballico. Discreteness for the set of complex structures on a real variety. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 46(3):321–322, September 2003. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Chamberland:2003:CTD

[3353] Marc Chamberland. Characterizing two-dimensional maps whose Jacobians have constant eigenvalues. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 46(3):323–331, September 2003. CO-DEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Dokovic:2003:SQA

[3354] Dragomir Z. Doković and Tin-Yau Tam. Some questions about semisimple Lie groups originating in matrix theory. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 46(3):332–343, September 2003. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Gurak:2003:GES

[3355] S. Gurak. Gauss and Eisenstein sums of order twelve. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 46(3):344–355, September 2003. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Ishiwata:2003:BCT

[3356] Makiko Ishiwata, Józef H. Przytycki, and Akira Yasuhara. Branched covers of tangles in three-balls. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 46(3):356–364, September 2003. CO-DEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Kishimoto:2003:HPS

[3357] Akitaka Kishimoto, Narutaka Ozawa, and Shôichirô Sakai. Homogeneity of the pure state space of a separable C^{∗}-algebra. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 46(3):365–372, September 2003. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Laugesen:2003:PTF

[3358] Richard S. Laugesen and Igor E. Pritsker. Potential theory of the farthest-point distance function. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 46(3):373–387, September 2003. CO-DEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Lin:2003:TQE

[3359] Huaxin Lin. Tracially quasidiagonal extensions. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 46(3):388–399, September 2003. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Marshall:2003:APP

[3360] M. Marshall. Approximating positive polynomials using sums of squares. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 46(3):400–418, September 2003. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Masuda:2003:NSF

[3361] Toshihiko Masuda. On non-strongly free automorphisms of subfactors of type III_{0}. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 46(3):419–428, September 2003. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Sastry:2003:GTR

[3362] Pramathanath Sastry and Yue Lin L. Tong. The Grothendieck trace and the de Rham integral. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 46(3):429–440, September 2003. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Stacey:2003:ILM

[3363] P. J. Stacey. An inductive limit model for the K-theory of the generatorinterchanging antiautomorphism of an irrational rotation algebra. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 46(3):441–456, September 2003. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Toms:2003:SPG

[3364] Andrew Toms. Strongly perforated K_{0}-groups of simple C^{∗}-algebras. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 46(3):457–472, September 2003. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Yeats:2003:MAS

[3365] Karen Yeats. A multiplicative analogue of Schur’s Tauberian Theorem. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 46(3):473–480, September 2003. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Bachir:2003:CDF

[3366] M. Bachir and G. Lancien. On the composition of differentiable functions. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 46(4):481–494, December 2003. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Baragar:2003:CVH

[3367] Arthur Baragar. Canonical vector heights on algebraic K3 surfaces with Picard Number Two. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 46(4):495–508, December 2003. CODEN CM-BUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Benson:2003:SKA

[3368] David J. Benson, Alex Kumjian, and N. Christopher Phillips. Symmetries of Kirchberg algebras. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 46(4):509–528, December 2003. CO-DEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Billig:2003:RTH

[3369] Yuly Billig. Representations of the twisted Heisenberg–Virasoro algebra at level zero. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 46(4):529–537, December 2003. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Borwein:2003:SWG

[3370] Jonathan Borwein, Simon Fitzpatrick, and Roland Girgensohn. Subdifferentials whose graphs are not norm × weak^{∗} closed. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 46(4):538–545, December 2003. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287

(electronic). URL http://cms.math.ca/10.4153/CMB-2003-051-5;
http://docserver.carma.newcastle.edu.au/149/.

Long:2003:SCE

[3371] Ling Long. L-series of certain elliptic surfaces. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 46(4):546–558, December 2003. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Marco:2003:DCI

[3372] Nicolas Marco and Xavier Massaneda. On density conditions for interpolation in the ball. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 46(4):559–574, December 2003. CODEN CM-BUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Marshall:2003:OPF

[3373] M. Marshall. Optimization of polynomial functions. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 46(4):575–587, December 2003. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Monteiro:2003:WSR

[3374] Martha Salerno Monteiro. Weakly stable relations and inductive limits of C^{∗}-algebras. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 46(4):588–596, December 2003. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Neeb:2003:CS

[3375] Karl-Hermann Neeb and Ivan Penkov. Cartan subalgebras of }↕_{∞}. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 46(4):597–616, December 2003. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic). See erratum [3875].

Pak:2003:HTF

[3376] Hong Kyung Pak. On harmonic theory in flows. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 46(4):617–631, December 2003. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Runde:2003:OAU

[3377] Volker Runde. The operator amenability of uniform algebras. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 46(4): 632–634, December 2003. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Anonymous:2003:AII

[3378] Anonymous. Author index — index des auteurs — for 2003 — pour 2003. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 46(4):635–??, December 2003. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Al-Salman:2004:SIR

[3379] Ahmad Al-Salman and Yibiao Pan. Singular integrals with rough kernels. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 47(1):3–11, March 2004. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Burger:2004:NMR

[3380] Edward B. Burger. On Newton’s method and rational approximations to quadratic irrationals. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 47(1):12–21, March 2004. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Goto:2004:NHF

[3381] Yasuhiro Goto. A note on the height of the formal Brauer group of a K3 surface. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 47(1):22–29, March 2004. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Khalkhali:2004:GCE

[3382] M. Khalkhali and B. Rangipour. On the generalized cyclic Eilenberg– Zilber Theorem. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 47(1):38–48, March 2004. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Lindstrom:2004:ENB

[3383] Mikael Lindström, Shamil Makhmutov, and Jari Taskinen. The essential norm of a Bloch-to-Q_{p} composition operator. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 47(1):49–59, March 2004. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Little:2004:RII

[3384] Robert D. Little. Rational integer invariants of regular cyclic actions. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 47(1):60–72, March 2004. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Li:2004:SHQ

[3385] Ma Li and Chen Dezhong. Systems of Hermitian quadratic forms. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 47 (1):73–81, March 2004. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Moody:2004:PSD

[3386] Robert V. Moody and Nicolae Strungaru. Point sets and dynamical systems in the autocorrelation topology. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 47(1):82–99, March 2004. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Seto:2004:IS

[3387] Michio Seto. Invariant subspaces on T^{N} and R^{N}. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 47(1):100–107, March 2004. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Sliwa:2004:USB

[3388] Wiesław Śliwa. On universal Schauder bases in non-Archimedean Fréchet spaces. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 47(1):108–118, March 2004. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Theriault:2004:PEB

[3389] Stephen D. Theriault. 2-primary exponent bounds for Lie groups of low rank. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 47(1):119–132, March 2004. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Wang:2004:EST

[3390] Wei Wang. Embeddability of some three-dimensional weakly pseudoconvex CR structures. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 47(1):133–143, March 2004. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Xia:2004:UWO

[3391] Jingbo Xia. On the uniqueness of wave operators associated with nontrace class perturbations. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 47(1):144–151, March 2004. CODEN CM-BUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Zheng:2004:UMF

[3392] Jian-Hua Zheng. On uniqueness of meromorphic functions with shared values in some angular domains. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 47(1):152–160, March 2004. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Alspach:2004:SSP

[3393] Brian Alspach and Shaofei Du. Suborbit structure of permutation pgroups and an application to Cayley digraph isomorphism. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 47(2): 161–167, June 2004. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Baake:2004:KDM

[3394] Michael Baake and Bernd Sing. Kolakoski-(3, 1) is a (deformed) model set. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 47(2):168–190, June 2004. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Gratzer:2004:CCS

[3395] G. Grätzer and E. T. Schmidt. Congruence class sizes in finite sectionally complemented lattices. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 47(2):191–205, June 2004. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Hurri-Syrjanen:2004:PIR

[3396] Ritva Hurri-Syrjänen. The Poincaré Inequality and reverse doubling weights. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 47(2):206–214, June 2004. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Jaworski:2004:CAI

[3397] Wojciech Jaworski. Countable amenable identity excluding groups. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 47 (2):215–228, June 2004. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Jensen:2004:SRR

[3398] C. U. Jensen. Solvability by real radicals and Fermat primes. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 47(2): 229–236, June 2004. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Laubie:2004:RSF

[3399] François Laubie. Ramification des séries formelles. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 47(2):237–245, June 2004. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Makai:2004:MSR

[3400] Endre Makai, Jr. and Horst Martini. On maximal k-sections and related common transversals of convex bodies. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 47(2):246–256, June 2004. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Marwaha:2004:GCN

[3401] Alka Marwaha. A geometric characterization of nonnegative bands. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 47 (2):257–263, June 2004. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

McKinnon:2004:CRP

[3402] David McKinnon. Counting rational points on ruled varieties. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 47(2): 264–270, June 2004. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Naumann:2004:LRA

[3403] Niko Naumann. Linear relations among the values of canonical heights from the existence of non-trivial endomorphisms. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 47(2):271–279, June 2004. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Perkins:2004:SBM

[3404] Ed Perkins. Super-Brownian motion and critical spatial stochastic systems. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 47(2):280–297, June 2004. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Yahaghi:2004:NTI

[3405] Bamdad R. Yahaghi. Near triangularizability implies triangularizability. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 47(2):298–313, June 2004. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Yun:2004:MCC

[3406] Jong-Gug Yun. Mean curvature comparison with L^{1}-norms of Ricci curvature. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 47(2):314–320, June 2004. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

He:2004:ELS

[3407] Xinyu He. Existence of Leray’s self-similar solutions of the Navier–Stokes equations in D ⊂ R^{∋}. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 47(1):30–37, March 2004. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Bullejos:2004:CSM

[3408] M. Bullejos and A. M. Cegarra. Classifying spaces for monoidal categories through geometric nerves. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 47(3):321–331, September 2004. CO-DEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Charette:2004:RGF

[3409] Virginie Charette, William M. Goldman, and Catherine A. Jones. Recurrent geodesics in flat Lorentz 3-manifolds. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 47(3):332–342, September 2004. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Drensky:2004:CWS

[3410] Vesselin Drensky and Lakhdar Hammoudi. Combinatorics of words and semigroup algebras which are sums of locally nilpotent subalgebras. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 47(3):343–353, September 2004. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Fawaz:2004:IFS

[3411] Amine Fawaz. An integral formula on Seifert bundles. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 47(3):354–357, September 2004. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Ford:2004:SFP

[3412] Kevin Ford. A strong form of a problem of R. L. Graham. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 47(3): 358–368, September 2004. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Fosner:2004:SBL

[3413] Ajda Fosner and Peter Semrl. Spectrally bounded linear maps on B(X). Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 47(3):369–372, September 2004. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Gyori:2004:DE

[3414] K. Győri, L. Hajdu, and N. Saradha. On the Diophantine equation n(n + d) ⋯ (n + (k − 1)d) = by^{l}. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 47(3):373–388, September 2004. CO-DEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic). See correction [3497].

He:2004:IFR

[3415] Jianxun He. An inversion formula of the Radon transform on the Heisenberg group. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 47(3):389–397, September 2004. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

McKinnon:2004:RBM

[3416] David McKinnon. A reduction of the Batyrev–Manin conjecture for Kummer surfaces. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 47(3):398–406, September 2004. CODEN CM-BUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Nedelec:2004:MRB

[3417] L. Nedelec. Multiplicity of resonances in black box scattering. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 47(3): 407–416, September 2004. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Niu:2004:HTL

[3418] Pengcheng Niu, Yanwu Han, and Junqiang Han. A Hopf type lemma and a CR type inversion for the generalized Greiner operator. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 47(3): 417–430, September 2004. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Osburn:2004:NRD

[3419] Robert Osburn. A note on 4-rank densities. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 47(3):431–438, September 2004. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Parker:2004:SBT

[3420] John R. Parker. On the stable basin theorem. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 47(3):439–444, September 2004. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Pirkovskii:2004:BBC

[3421] A. Yu. Pirkovskii. Biprojectivity and biflatness for convolution algebras of nuclear operators. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 47(3):445–455, September 2004. CODEN CM-BUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Seto:2004:BCL

[3422] Michio Seto. On the Berger–Coburn–Lebow problem for Hardy submodules. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 47(3):456–467, September 2004. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Soundararajan:2004:SMO

[3423] K. Soundararajan. Strong multiplicity one for the Selberg class. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 47(3): 468–474, September 2004. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Wade:2004:UAE

[3424] W. R. Wade. Uniqueness of almost everywhere convergent vilenkin series. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 47(3):475–480, September 2004. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Bekjan:2004:NCH

[3425] Turdebek N. Bekjan. A new characterization of Hardy martingale cotype space. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 47(4):481–491, December 2004. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Boumuki:2004:IIL

[3426] Nobutaka Boumuki. Isotropic immersions with low codimension of complex space forms into real space forms. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 47(4):492–503, December 2004. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Cardoso:2004:HFR

[3427] Fernando Cardoso and Georgi Vodev. High frequency resolvent estimates and energy decay of solutions to the wave equation. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 47(4):504–514, December 2004. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Frigon:2004:RET

[3428] M. Frigon. Remarques sur l’enlacement en théorie des points critiques pour des fonctionnelles continues. (French) [Remarks on the intertwining in critical-point theory for continuous functionals]. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 47(4):515–529, December 2004. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Iranmanesh:2004:C

[3429] A. Iranmanesh and B. Khosravi. A characterization of P SU_{11}(q). Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 47(4):530–539, December 2004. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Jain:2004:CHT

[3430] Pankaj Jain, Pawan K. Jain, and Babita Gupta. Compactness of Hardytype operators over star-shaped regions in R^{N}. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 47(4):540–552, December 2004. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Kerr:2004:GAV

[3431] David Kerr. A geometric approach to Voiculescu–Brown entropy. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 47(4):553–565, December 2004. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Koike:2004:ASW

[3432] Kenji Koike. Algebraicity of some Weil Hodge classes. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 47(4):566– 572, December 2004. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Liu:2004:GTT

[3433] Yu-Ru Liu. A generalization of the Turán Theorem and its applications. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 47(4):573–588, December 2004. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Liu:2004:GEK

[3434] Yu-Ru Liu. A generalization of the Erdős–Kac Theorem and its applications. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 47(4):589–606, December 2004. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Plamenevskaya:2004:RFV

[3435] Olga Plamenevskaya. A residue formula for SU(2)-valued moment maps. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 47(4):607–614, December 2004. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Randrianantoanina:2004:AFT

[3436] Narcisse Randrianantoanina. C^{∗}-algebras and factorization through diagonal operators. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 47(4):615–623, December 2004. CODEN CM-BUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Zhang:2004:CTY

[3437] Xi Zhang. A compactness theorem for Yang–Mills connections. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 47(4): 624–634, December 2004. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Anonymous:2004:AII

[3438] Anonymous. Author index — index des auteurs — for 2004 — pour 2004. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 47(4):635–??, December 2004. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Burq:2005:QEB

[3439] N. Burq. Quantum ergodicity of boundary values of eigenfunctions: a control theory approach. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 48(1):3–15, March 2005. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Cojocaru:2005:SGR

[3440] Alina Carmen Cojocaru. On the surjectivity of the Galois representations associated to non-CM elliptic curves. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 48(1):16–31, March 2005. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Dabkowski:2005:NLO

[3441] Mieczysław K. Dabkowski, Józef H. Przytycki, and Amir A. Togha. Nonleft-orderable 3-manifold groups. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 48(1):32–40, March 2005. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Dixon:2005:DHS

[3442] John D. Dixon and A. Rahnamai Barghi. Degree homogeneous subgroups. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 48(1):41–49, March 2005. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Elliott:2005:ICM

[3443] George A. Elliott, Guihua Gong, and Liangqing Li. Injectivity of the connecting maps in AH inductive limit systems. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 48(1):50–68, March 2005. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Fabian:2005:BSW

[3444] M. Fabian, V. Montesinos, and V. Zizler. Biorthogonal systems in weakly Lindelöf spaces. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 48(1):69–79, March 2005. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Herman:2005:TUG

[3445] Allen Herman, Yuanlin Li, and M. M. Parmenter. Trivial units for group rings with G-adapted coefficient rings. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 48(1):80–89, March 2005. CO-DEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Jeffrey:2005:PCC

[3446] Lisa C. Jeffrey and Augustin-Liviu Mare. Products of conjugacy classes in SU(2). Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 48(1):90–96, March 2005. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Katavolos:2005:RBP

[3447] Aristides Katavolos and Vern I. Paulsen. On the ranges of bimodule projections. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 48(1):97–111, March 2005. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Mo:2005:NCF

[3448] Xiaohuan Mo and Zhongmin Shen. On negatively curved Finsler manifolds of scalar curvature. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 48(1):112–120, March 2005. CODEN CM-BUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Mollin:2005:NSC

[3449] R. A. Mollin. Necessary and sufficient conditions for the central norm to equal 2^{h} in the simple continued fraction expansion of \sqrt{2^{h}c} for any odd c > 1. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 48(1):121–132, March 2005. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Talvila:2005:EHK

[3450] Erik Talvila. Estimates of Henstock–Kurzweil Poisson integrals. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 48(1): 133–146, March 2005. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Vaananen:2005:BTE

[3451] Keijo Väänänen and Wadim Zudilin. Baker-type estimates for linear forms in the values of q-series. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 48(1):147–160, March 2005. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Betancor:2005:HCO

[3452] Jorge J. Betancor. Hankel convolution operators on spaces of entire functions of finite order. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 48(2):161–174, June 2005. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Borwein:2005:WCO

[3453] David Borwein and Werner Kratz. Weighted convolution operators on ℓ_{p}. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 48(2):175–179, June 2005. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Cynk:2005:GAC

[3454] Slawomir Cynk and Christian Meyer. Geometry and arithmetic of certain double octic Calabi–Yau manifolds. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 48(2):180–194, June 2005. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Daniel:2005:SC

[3455] D. Daniel, J. Nikiel, L. B. Treybig, H. M. Tuncali, and E. D. Tymchatyn. On Suslinian continua. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 48(2):195–202, June 2005. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

deQuehen:2005:NCM

[3456] Victoria E. de Quehen and Leslie G. Roberts. Non-Cohen–Macaulay projective monomial curves with positive h-vector. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 48(2):203–210, June 2005. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Germain:2005:DT

[3457] Jam Germain. The distribution of totatives. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 48(2):211–220, June 2005. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Kerr:2005:EPS

[3458] Matt Kerr. An elementary proof of Suslin reciprocity. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 48(2):221– 236, June 2005. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Kimura:2005:IHC

[3459] Kenichiro Kimura. Indecomposable higher chow cycles. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 48(2):237– 243, June 2005. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

McLeod:2005:CMC

[3460] Alice McLeod and William Moser. Counting multiple cyclic choices without adjacencies. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 48(2):244–250, June 2005. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Murphy:2005:ITA

[3461] G. J. Murphy. The index theory associated to a non-finite trace on a C^{∗}- algebra. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 48(2):251–259, June 2005. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Oberlin:2005:RTS

[3462] Daniel M. Oberlin. A restriction theorem for a k-surface in R^{n}. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 48(2): 260–266, June 2005. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Rodman:2005:CAP

[3463] Leiba Rodman, Peter Semrl, and Ahmed R. Sourour. Continuous adjacency preserving maps on real matrices. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 48(2):267–274, June 2005. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Smith:2005:KDI

[3464] Patrick F. Smith. Krull dimension of injective modules over commutative Noetherian rings. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 48(2):275–282, June 2005. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Thibault:2005:EIS

[3465] Lionel Thibault and Dariusz Zagrodny. Enlarged inclusion of subdifferentials. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 48(2):283–301, June 2005. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Yokonuma:2005:DSA

[3466] Takeo Yokonuma. Discrete sets and associated dynamical systems in a non-commutative setting. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 48(2):302–316, June 2005. CODEN CM-BUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Yousif:2005:PFR

[3467] Mohamed F. Yousif, Yiqiang Zhou, and Nasr Zeyada. On pseudo-Frobenius rings. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 48(2):317–320, June 2005. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Akbary:2005:NVC

[3468] Amir Akbary and Shahab Shahabi. On non-vanishing of convolution of Dirichlet series. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 48(3):321–332, September 2005. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Alzer:2005:MPH

[3469] Horst Alzer. Monotonicity properties of the Hurwitz zeta function. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 48(3):333–339, September 2005. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Andruchow:2005:SGU

[3470] Esteban Andruchow. Short geodesics of unitaries in the L^{2} metric. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 48(3):340–354, September 2005. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Chebotar:2005:MPP

[3471] M. A. Chebotar, W.-F. Ke, P.-H. Lee, and L.-S. Shiao. On maps preserving products. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 48(3):355–369, September 2005. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Daly:2005:TM

[3472] J. E. Daly and S. Fridli. Trigonometric multipliers on H_{2π}. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 48(3): 370–381, September 2005. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

DeCarli:2005:UEU

[3473] Laura De Carli. Uniform estimates of ultraspherical polynomials of large order. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 48(3):382–393, September 2005. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Dokovic:2005:DPT

[3474] D. Z. Doković, F. Szechtman, and K. Zhao. Diagonal plus tridiagonal representatives for symplectic congruence classes of symmetric matrices. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 48(3):394–404, September 2005. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Froese:2005:LTR

[3475] Richard Froese. Liouville’s Theorem in the radially symmetric case. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 48(3):405–408, September 2005. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Gauthier:2005:EUI

[3476] P. M. Gauthier and J. Xiao. The existence of universal inner functions on the unit ball of C^{n}. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 48(3):409–413, September 2005. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Kaveh:2005:VFC

[3477] Kiumars Kaveh. Vector fields and the cohomology ring of toric varieties. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 48(3):414–427, September 2005. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Miyamoto:2005:REC

[3478] Roland Miyamoto and Jaap Top. Reduction of elliptic curves in equal characteristic 3 (and 2). Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 48(3):428–444, September 2005. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Patras:2005:GLI

[3479] Frédéric Patras, Christophe Reutenauer, and Manfred Schocker. On the Garsia Lie idempotent. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 48(3):445–454, September 2005. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Rychtar:2005:GDC

[3480] Jan Rychtár. On Gâteaux differentiability of convex functions in WCG spaces. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 48(3):455–459, September 2005. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Sommers:2005:SIN

[3481] Eric N. Sommers. B-stable ideals in the nilradical of a Borel subalgebra. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 48(3):460–472, September 2005. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Zeron:2005:LTC

[3482] E. S. Zeron. Logarithms and the topology of the complement of a hypersurface. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 48(3):473–480, September 2005. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Azagra:2005:EFF

[3483] D. Azagra, M. Fabian, and M. Jiménez-Sevilla. Exact filling of figures with the derivatives of smooth mappings between Banach spaces. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 48(4): 481–499, December 2005. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Baracco:2005:EHF

[3484] Luca Baracco. Extension of holomorphic functions from one side of a hypersurface. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 48(4):500–504, December 2005. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Bouikhalene:2005:GDW

[3485] Belaid Bouikhalene. On the generalized d’Alembert’s and Wilson’s functional equations on a compact group. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 48(4):505–522, December 2005. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Duvelmeyer:2005:AMB

[3486] Nico Düvelmeyer. Angle measures and bisectors in Minkowski planes. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 48(4):523–534, December 2005. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Ellenberg:2005:ETD

[3487] Jordan S. Ellenberg. On the error term in Duke’s estimate for the average special value of L-functions. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 48(4):535–546, December 2005. CO-DEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Feher:2005:DFF

[3488] L. M. Fehér, A. Némethi, and R. Rimányi. Degeneracy of 2-forms and 3-forms. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 48(4):547–560, December 2005. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Foth:2005:NLL

[3489] Philip Foth. A note on Lagrangian loci of quotients. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 48(4):561–575, December 2005. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Ichimura:2005:TKN

[3490] Humio Ichimura. On a theorem of Kawamoto on normal bases of rings of integers, II. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 48(4):576–579, December 2005. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Kot:2005:ESH

[3491] Piotr Kot. Exceptional sets in Hartogs domains. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 48(4):580–586, December 2005. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Lopes:2005:SV

[3492] Samuel A. Lopes. Separation of variables for U_{q}(∫↕_{\+∞})^{+}. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 48(4): 587–600, December 2005. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Mashreghi:2005:RDM

[3493] Javad Mashreghi and Mohamad R. Pouryayevali. On the regularity of the s-differential metric. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 48(4):601–606, December 2005. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Park:2005:TAE

[3494] Efton Park. Toeplitz algebras and extensions of irrational rotation algebras. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 48(4):607–613, December 2005. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Tuncali:2005:FOM

[3495] H. Murat Tuncali and Vesko Valov. On finite-to-one maps. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 48(4): 614–621, December 2005. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Venereau:2005:HFV

[3496] Stéphane Vénéreau. Hyperplanes of the form f_{1}(x, y)z_{1} + ⋯ + f_{k}(x, y)z_{k} + g(x, y) are variables. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 48(4):622–635, December 2005. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Gyori:2005:CDE

[3497] K. Győri, L. Hajdu, and N. Saradha. Correction to: “On the Diophantine Equation n(n + d) ⋯ (n + (k − 1)d) = by^{l}”. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 48(4):636, December 2005. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic). See [3414].

Anonymous:2005:AII

[3498] Anonymous. Author index — index des auteurs — for 2005 — pour 2005. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 48(4):637–??, December 2005. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Al-Salman:2006:CSI

[3499] Ahmad Al-Salman. On a class of singular integral operators with rough kernels. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 49(1):3–10, March 2006. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Bevelacqua:2006:GRF

[3500] Anthony J. Bevelacqua and Mark J. Motley. Going-down results for C_{i}- fields. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 49(1):11–20, March 2006. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Chapman:2006:EDE

[3501] Robin Chapman and William Hart. Evaluation of the Dedekind eta function. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 49(1):21–35, March 2006. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Daskalopoulos:2006:HFW

[3502] Georgios D. Daskalopoulos and Richard A. Wentworth. Holomorphic frames for weakly converging holomorphic vector bundles. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 49(1):36–40, March 2006. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Doeraene:2006:GWV

[3503] Jean-Paul Doeraene and Mohammed El Haouari. The Ganea and Whitehead variants of the Lusternik–Schnirelmann category. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 49(1):41–54, March 2006. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Dubois:2006:NAT

[3504] Jérôme Dubois. Non Abelian twisted Reidemeister torsion for fibered knots. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 49(1):55–71, March 2006. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Dwilewicz:2006:ARH

[3505] Roman J. Dwilewicz. Additive Riemann–Hilbert problem in line bundles over CP^{1}. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 49(1):72–81, March 2006. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Gogatishvili:2006:EDT

[3506] Amiran Gogatishvili and Lubos Pick. Embeddings and duality theorem for weak classical Lorentz spaces. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 49(1):82–95, March 2006. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Kulshammer:2006:RSM

[3507] Burkhard Külshammer. Roots of simple modules. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 49(1):96–107, March 2006. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Kwapisz:2006:DPP

[3508] Jarosław Kwapisz. A dynamical proof of Pisot’s Theorem. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 49(1):108– 112, March 2006. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Ledet:2006:E

[3509] Arne Ledet. PSL(2, 2^{n})-extensions over F_{2^{n}}. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 49(1):113–116, March 2006. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Levene:2006:DTO

[3510] R. H. Levene. A double triangle operator algebra from SL_{2}(R). Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 49(1):

117–126, March 2006. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Lewis:2006:CDG

[3511] Mark L. Lewis. Character degree graphs of solvable groups of fitting height 2. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 49(1):127–133, March 2006. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Suh:2006:RHC

[3512] Young Jin Suh. Real hypersurfaces in complex two-plane Grassmannians with vanishing Lie derivative. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 49(1):134–143, March 2006. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Taylor:2006:SLS

[3513] Michael Taylor. Scattering length and the spectrum of −Δ + V. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 49(1): 144–151, March 2006. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Yun:2006:CGN

[3514] Jong-Gug Yun. Comparison geometry with L^{1}-norms of Ricci curvature. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 49(1):152–160, March 2006. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Agapito:2006:WBG

[3515] José Agapito. Weighted Brianchon–Gram decomposition. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 49(2):161– 169, June 2006. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Atkins:2006:GES

[3516] Richard Atkins. The geometry of d^{2}y^{1}/dt^{2} = f(y, ẏ, t) and d^{2}y^{2}/dt^{2} = g(y, ẏ, t), and Euclidean spaces. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 49(2):170–184, June 2006. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Averkov:2006:IVM

[3517] Gennadiy Averkov. On the inequality for volume and Minkowskian thickness. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 49(2):185–195, June 2006. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Chernousov:2006:APT

[3518] Vladimir Chernousov. Another proof of Totaro’s Theorem on E_{8}-torsors. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 49(2):196–202, June 2006. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Comez:2006:EHT

[3519] Dogan Çömez. The ergodic Hilbert transform for admissible processes. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 49(2):203–212, June 2006. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Dean:2006:ILT

[3520] Andrew J. Dean. On inductive limit type actions of the Euclidean motion group on stable UHF algebras. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 49(2):213–225, June 2006. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Engman:2006:SIE

[3521] Martin Engman. The spectrum and isometric embeddings of surfaces of revolution. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 49(2):226–236, June 2006. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Gauthier:2006:ARM

[3522] P. M. Gauthier and E. S. Zeron. Approximation by rational mappings, via homotopy theory. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 49(2):237–246, June 2006. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Myjak:2006:SMT

[3523] Józef Myjak, Tomasz Szarek, and Maciej Śleczka. A Szpilrajn– marczewski type theorem for concentration dimension on Polish spaces. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 49(2):247–255, June 2006. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Neelon:2006:BWT

[3524] Tejinder Neelon. A Bernstein–Walsh type inequality and applications. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 49(2):256–264, June 2006. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Nicholson:2006:ESU

[3525] W. K. Nicholson and Y. Zhou. Endomorphisms that are the sum of a unit and a root of a fixed polynomial. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 49(2):265–269, June 2006. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Occhetta:2006:CPP

[3526] Gianluca Occhetta. A characterization of products of projective spaces. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 49(2):270–280, June 2006. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Ragnarsson:2006:CTT

[3527] Carl Johan Ragnarsson, Wesley Wai Suen, and David G. Wagner. Correction to a theorem on total positivity. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 49(2):281–284, June 2006. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Riedl:2006:OSS

[3528] Jeffrey M. Riedl. Orbits and stabilizers for solvable linear groups. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 49(2): 285–295, June 2006. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Schutt:2006:MTC

[3529] Matthias Schütt. On the modularity of three Calabi–Yau threefolds with bad reduction at 11. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 49(2):296–312, June 2006. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Wagner:2006:RBG

[3530] Roy Wagner. On the relation between the Gaussian orthogonal ensemble and reflections, or a self-adjoint version of the Marcus–Pisier inequality. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 49(2):313–320, June 2006. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Balser:2006:PPG

[3531] Andreas Balser. Polygons with prescribed Gauss map in Hadamard spaces and Euclidean buildings. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 49(3):321–336, September 2006. CO-DEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Berlanga:2006:HEG

[3532] R. Berlanga. Homotopy equivalence and groups of measure-preserving homeomorphisms. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 49(3):337–346, September 2006. CODEN CM-BUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Ecker:2006:ACG

[3533] Jürgen Ecker. Affine completeness of generalised dihedral groups. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 49(3):347–357, September 2006. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

ElKhalil:2006:PEL

[3534] Abdelouahed El Khalil, Said El Manouni, and Mohammed Ouanan. On the principal eigencurve of the p-Laplacian: Stability phenomena. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 49(3):358–370, September 2006. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Floricel:2006:ISI

[3535] Remus Floricel. Inner E_{0}-semigroups on infinite factors. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 49(3):371– 380, September 2006. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Girela:2006:MBS

[3536] Daniel Girela and José Ángel Peláez. On the membership in Bergman spaces of the derivative of a Blaschke product with zeros in a Stolz domain. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 49(3):381–388, September 2006. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Hiai:2006:FLS

[3537] Fumio Hiai, Dénes Petz, and Yoshimichi Ueda. A free logarithmic Sobolev inequality on the circle. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 49(3):389–406, September 2006. CO-DEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Jardine:2006:IMS

[3538] J. F. Jardine. Intermediate model structures for simplicial presheaves. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 49(3):407–413, September 2006. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Jiang:2006:CET

[3539] Liya Jiang, Houyu Jia, and Han Xu. Commutators estimates on Triebel– Lizorkin spaces. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 49(3):414–427, September 2006. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Lee:2006:VVM

[3540] Min Ho Lee. Vector-valued modular forms of weight two associated with Jacobi-like forms. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 49(3):428–437, September 2006. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Mercer:2006:URS

[3541] Idris David Mercer. Unimodular roots of special Littlewood polynomials. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 49(3):438–447, September 2006. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Pacelli:2006:LBN

[3542] Allison M. Pacelli. A lower bound on the number of cyclic function fields with class number divisible by n. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 49(3):448–463, September 2006. CO-DEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Ravindra:2006:NDA

[3543] G. V. Ravindra. A note on detecting algebraic cycles. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 49(3):464–471, September 2006. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Silvester:2006:CCF

[3544] Alan K. Silvester, Blair K. Spearman, and Kenneth S. Williams. Cyclic cubic fields of given conductor and given index. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 49(3):472–480, September 2006. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Browkin:2006:SSC

[3545] J. Browkin and J. Brzeziński. On sequences of squares with constant second differences. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 49(4):481–491, December 2006. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Chan:2006:CCS

[3546] Kai-Cheong Chan and Dragomir Z. Doković. Conjugacy classes of subalgebras of the real sedenions. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 49(4):492–525, December 2006. CO-DEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Choi:2006:VMF

[3547] So Young Choi. The values of modular functions and modular forms. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 49(4):526–535, December 2006. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Dostal:2006:MCM

[3548] Petr Dostál, Jaroslav Lukes, and Jirí Spurný. Measure convex and measure extremal sets. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 49(4):536–548, December 2006. CODEN CM-BUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Fuhr:2006:HYI

[3549] Hartmut Führ. Hausdorff–Young inequalities for group extensions. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 49(4):549–559, December 2006. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

vanLuijk:2006:SAC

[3550] Ronald van Luijk. A K3 surface associated with certain integral matrices having integral eigenvalues. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 49(4):560–577, December 2006. CO-DEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Muic:2006:SFL

[3551] Goran Muić. On the structure of the full lift for the Howe correspondence of (Sp(n), O(V)) for rank-one reducibilities. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 49(4):578–591, December 2006. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Sarti:2006:GAC

[3552] Alessandra Sarti. Group actions, cyclic coverings and families of K3surfaces. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 49(4):592–608, December 2006. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Stoimenow:2006:PVV

[3553] A. Stoimenow. On the Polyak–Viro Vassiliev invariant of degree 4. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 49(4):609–623, December 2006. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Teragaito:2006:NID

[3554] Masakazu Teragaito. On non-integral Dehn surgeries creating nonorientable surfaces. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 49(4):624–627, December 2006. CODEN CM-BUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Zeron:2006:ATR

[3555] E. S. Zeron. Approximation and the topology of rationally convex sets. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 49(4):628–636, December 2006. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Anonymous:2006:AII

[3556] Anonymous. Author index — index des auteurs — for 2006 — pour 2006. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 49(4):637–??, December 2006. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Basener:2007:HDS

[3557] Richard F. Basener. Higher dimensional spaces of functions on the spectrum of a uniform algebra. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 50(1):3–10, March 2007. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Borwein:2007:VPE

[3558] David Borwein and Jonathan Borwein. van der Pol expansions of Lseries. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 50(1):11–23, March 2007. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic). URL http://docserver.carma.

newcastle.edu.au/283/.

Brown:2007:IMN

[3559] Nathan Brown, Rachel Finck, Matthew Spencer, Kristopher Tapp, and Zhongtao Wu. Invariant metrics with nonnegative curvature on compact Lie groups. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 50(1):24–34, March 2007. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Duyckaerts:2007:SCP

[3560] Thomas Duyckaerts. A singular critical potential for the Schrödinger operator. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 50(1):35–47, March 2007. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Dvorsky:2007:TSM

[3561] Alexander Dvorsky. Tensor square of the minimal representation of O(p, q). Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 50(1):48–55, March 2007. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Gourdeau:2007:SCS

[3562] F. Gourdeau, A. Pourabbas, and M. C. White. Simplicial cohomology of some semigroup algebras. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 50(1):56–70, March 2007. CODEN CM-BUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Gurak:2007:PKS

[3563] S. Gurak. Polynomials for Kloosterman sums. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 50(1):71–84, March 2007. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Han:2007:CFG

[3564] Deguang Han. Classification of finite group-frames and super-frames. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 50(1):85–96, March 2007. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Kim:2007:CRH

[3565] In-Bae Kim, Ki Hyun Kim, and Woon Ha Sohn. Characterizations of real hypersurfaces in a complex space form. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 50(1):97–104, March 2007. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Klep:2007:VPG

[3566] Igor Klep. On valuations, places and graded rings associated to ∗- orderings. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 50(1):105–112, March 2007. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Li:2007:HHM

[3567] ZhenYang Li and Xi Zhang. Hermitian harmonic maps into convex balls. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 50(1):113–122, March 2007. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Nikolov:2007:SAI

[3568] Nikolai Nikolov and Peter Pflug. Simultaneous approximation and interpolation on arakelian sets. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 50(1):123–125, March 2007. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Ongay:2007:DCM

[3569] Fausto Ongay. φ-dialgebras and a class of matrix “Coquecigrues”. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 50 (1):126–137, March 2007. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Sari:2007:SSS

[3570] Bünyamin Sari. On the structure of the set of symmetric sequences in Orlicz sequence spaces. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 50(1):138–148, March 2007. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Sliwa:2007:QNA

[3571] Wiesław Śliwa. On quotients of non-Archimedean Köthe spaces. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 50(1): 149–157, March 2007. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Tipu:2007:NGC

[3572] Vicentiu Tipu. A note on Giuga’s Conjecture. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 50(1):158–160, March 2007. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Arapura:2007:FCF

[3573] Donu Arapura and Su-Jeong Kang. Functoriality of the Coniveau filtration. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 50(2):161–171, June 2007. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Aron:2007:IDV

[3574] Richard Aron and Pamela Gorkin. An infinite dimensional vector space of universal functions for H^{∞} of the ball. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 50(2):172–181, June 2007. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Chapoton:2007:CTT

[3575] Frédéric Chapoton. On the Coxeter transformations for Tamari posets. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 50(2):182–190, June 2007. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Drungilas:2007:ERA

[3576] Paulius Drungilas and Arturas Dubickas. Every real algebraic integer is a difference of two Mahler measures. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 50(2):191–195, June 2007. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Fernandez:2007:PQT

[3577] Julio Fernández, Josep González, and Joan-C. Lario. Plane quartic twists of X(5, 3). Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 50(2):196–205, June 2007. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Golasinski:2007:SSF

[3578] Marek Golasiński and Daciberg Lima Gonçalves. Spherical space forms: Homotopy types and self-equivalences for the group (Z/a[rtimes]Z/b) × SL_{2} (F_{p}). Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 50(2):206–214, June 2007. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Kloosterman:2007:ESG

[3579] Remke Kloosterman. Elliptic K3 surfaces with geometric Mordell–Weil rank 15. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 50(2):215–226, June 2007. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Kucerovsky:2007:ASR

[3580] D. Kucerovsky and P. W. Ng. AF-skeletons and real rank zero algebras with the corona factorization property. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 50(2):227–233, June 2007. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Kuo:2007:RMA

[3581] Wentang Kuo. A remark on a modular analogue of the Sato–Tate conjecture. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 50(2):234–242, June 2007. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Langlands:2007:NPR

[3582] Robert P. Langlands. Un nouveau point de repère dans la théorie des formes automorphes. (French) [A new landmark in the theory of automorphic forms]. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 50(2):243–267, June 2007. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Manuilov:2007:LIE

[3583] V. Manuilov and K. Thomsen. On the lack of inverses to C^{∗}-extensions related to Property T groups. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 50(2):268–283, June 2007. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

McIntosh:2007:SOM

[3584] Richard J. McIntosh. Second order mock theta functions. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 50(2):284– 290, June 2007. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Sarkar:2007:BTC

[3585] Rudra P. Sarkar and Jyoti Sengupta. Beurling’s Theorem and characterization of heat kernel for Riemannian symmetric spaces of noncompact type. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 50(2):291–312, June 2007. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Tzermias:2007:CLM

[3586] Pavlos Tzermias. On Cauchy–Liouville–Mirimanoff polynomials. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 50 (2):313–320, June 2007. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Blair:2007:LC

[3587] David E. Blair. On Lagrangian catenoids. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 50(3):321–333, September 2007. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Chiang-Hsieh:2007:DHC

[3588] Hung-Jen Chiang-Hsieh and Yifan Yang. Determination of Hauptmoduls and construction of Abelian extensions of quadratic number fields. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 50(3):334–346, September 2007. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Perez:2007:RHC

[3589] Juan de Dios Pérez, Florentino G. Santos, and Young Jin Suh. Real hypersurfaces in complex projective space whose structure Jacobi operator is of Codazzi type. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 50(3):347–355, September 2007. CODEN CM-BUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Filippakis:2007:EPS

[3590] Michael E. Filippakis and Nikolaos S. Papageorgiou. Existence of positive solutions for nonlinear noncoercive hemivariational inequalities. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 50(3): 356–364, September 2007. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Godinho:2007:ECA

[3591] Leonor Godinho. Equivariant cohomology of S^{1}-actions on 4-manifolds. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 50(3):365–376, September 2007. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Gutierrez:2007:GIM

[3592] C. Gutierrez, X. Jarque, J. Llibre, and M. A. Teixeira. Global injectivity of C^{1} maps of the real plane, inseparable leaves and the Palais–Smale condition. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 50(3):377–389, September 2007. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Hebda:2007:LNS

[3593] James J. Hebda, Chun-Chung Hsieh, and Chichen M. Tsau. Linking number of singular links and the Seifert matrix. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 50(3):390–398, September 2007. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Komornik:2007:ECB

[3594] Vilmos Komornik and Paola Loreti. Expansions in complex bases. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 50(3):399–408, September 2007. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Luca:2007:DCM

[3595] Florian Luca and Igor E. Shparlinski. Discriminants of complex multiplication fields of elliptic curves over finite fields. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 50(3):409–417, September 2007. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Matui:2007:SPA

[3596] Hiroki Matui. A short proof of affability for certain Cantor minimal Z^{2}- systems. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 50(3):418–426, September 2007. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Mejia:2007:ICE

[3597] Israel Moreno Mejía. On the image of certain extension maps. I. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 50(3): 427–433, September 2007. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Ozarslan:2007:MTO

[3598] M. Ali Õzarslan and Oktay Duman. MKZ type operators providing a better estimation on [1/2, 1). Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 50(3):434–439, September 2007. CO-DEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Raghuram:2007:KTA

[3599] A. Raghuram. A Künneth Theorem for p-adic groups. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 50(3):440–446, September 2007. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Sniatycki:2007:GFT

[3600] Jedrzej Śniatycki. Generalizations of Frobenius’ Theorem on manifolds and subcartesian spaces. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 50(3):447–459, September 2007. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Spielberg:2007:WSP

[3601] Jack Spielberg. Weak semiprojectivity for purely infinite C^{∗}-algebras. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 50(3):460–468, September 2007. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Tvalavadze:2007:SDE

[3602] M. V. Tvalavadze. Simple decompositions of the exceptional Jordan algebra. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 50(3):469–473, September 2007. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Zhou:2007:WIS

[3603] Jiazu Zhou. On Willmore’s Inequality for submanifolds. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 50(3):474– 480, September 2007. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Blanloeil:2007:CND

[3604] Vincent Blanloeil and Osamu Saeki. Concordance des noeuds de dimension 4. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 50(4):481–485, December 2007. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Cynk:2007:HDM

[3605] S. Cynk and K. Hulek. Higher-dimensional modular Calabi–Yau manifolds. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 50(4):486–503, December 2007. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Dukes:2007:AER

[3606] Peter Dukes and Alan C. H. Ling. Asymptotic existence of resolvable graph designs. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 50(4):504–518, December 2007. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Henson:2007:ANC

[3607] C. Ward Henson, Yves Raynaud, and Andrew Rizzo. On axiomatizability of non-commutative L_{p}-spaces. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 50(4):519–534, December 2007. CO-DEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Hohlweg:2007:GDA

[3608] Christophe Hohlweg. Generalized descent algebras. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 50(4):535–546, December 2007. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Iakovlev:2007:ILT

[3609] Serguei Iakovlev. Inverse Laplace transforms encountered in hyperbolic problems of non-stationary fluid-structure interaction. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 50(4):547–566, December 2007. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Joshi:2007:ETF

[3610] Kirti Joshi. Exotic torsion, Frobenius splitting and the slope spectral sequence. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 50(4):567–587, December 2007. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Labute:2007:CDS

[3611] John Labute, Nicole Lemire, Ján Minác, and John Swallow. Cohomological dimension and Schreier’s formula in Galois cohomology. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 50(4): 588–593, December 2007. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Laubie:2007:RGA

[3612] François Laubie. Ramification des groupes abéliens d’automorphismes des corps F_{q}((X)). (French) [Branching Abelian groups for automorphism of the object F_{q}((X))]. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 50(4):594–597, December 2007. CODEN CM-BUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Lorestani:2007:ALC

[3613] Keivan Borna Lorestani, Parviz Sahandi, and Siamak Yassemi. Artinian local cohomology modules. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 50(4):598–602, December 2007. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Penkov:2007:CGH

[3614] Ivan Penkov and Gregg Zuckerman. Construction of generalized Harish-Chandra modules with arbitrary minimal ∥-type. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 50(4):603–609, December 2007. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Rychtar:2007:WKK

[3615] Jan Rychtár and Jirí Spurný. On weak^{∗} Kadec–Klee norms. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 50(4): 610–618, December 2007. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Tcaciuc:2007:EAS

[3616] Adi Tcaciuc. On the existence of asymptotic-l_{p} structures in Banach spaces. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 50(4):619–631, December 2007. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Zelenyuk:2007:TCG

[3617] Yevhen Zelenyuk and Yuliya Zelenyuk. Transformations and colorings of groups. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 50(4):632–636, December 2007. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Anonymous:2007:AII

[3618] Anonymous. Author index — index des auteurs — for 2007 — pour 2007. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 50(4):637, December 2007. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Alaca:2008:CS

[3619] Ayşe Alaca, Şaban Alaca, and Kenneth S. Williams. The convolution sum ∑_{m<n/16} σ(m)σ(n − 16m). Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 51(1):3–14, March 2008. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Aqzzouz:2008:DPC

[3620] Belmesnaoui Aqzzouz, Redouane Nouira, and Larbi Zraoula. The duality problem for the class of AM-compact operators on Banach lattices. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 51(1):15–20, March 2008. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Baracco:2008:REC

[3621] Luca Baracco. A remark on extensions of CR functions from hyperplanes. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 51(1):21–25, March 2008. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Belinschi:2008:HTM

[3622] S. T. Belinschi and H. Bercovici. Hinčin’s Theorem for multiplicative free convolution. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 51(1):26–31, March 2008. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Choi:2008:LIC

[3623] Stephen Choi and Ping Zhou. On linear independence of a certain multivariate infinite product. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 51(1):32–46, March 2008. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Croot:2008:MNT

[3624] Ernie Croot. The minimal number of three-term arithmetic progressions modulo a prime converges to a limit. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 51(1):47–56, March 2008. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Dobrowolski:2008:NIS

[3625] Edward Dobrowolski. A note on integer symmetric matrices and Mahler’s measure. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 51(1):57–59, March 2008. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Janzen:2008:FNS

[3626] David Janzen. Følner nets for semidirect products of amenable groups. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 51(1):60–66, March 2008. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Kalton:2008:RIF

[3627] Nigel Kalton and Fyodor Sukochev. Rearrangement-invariant functionals with applications to traces on symmetrically normed ideals. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 51(1): 67–80, March 2008. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Kassel:2008:HFC

[3628] Christian Kassel. Homotopy formulas for cyclic groups acting on rings. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 51(1):81–85, March 2008. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Nakazato:2008:NRD

[3629] Hiroshi Nakazato, Natália Bebiano, and João da Providência. The numerical range of 2-dimensional Krein space operators. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 51(1):86–99, March 2008. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Petkov:2008:DZF

[3630] Vesselin Petkov. Dynamical zeta function for several strictly convex obstacles. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 51(1):100–113, March 2008. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Petrov:2008:ZCT

[3631] V. Petrov, N. Semenov, and K. Zainoulline. Zero cycles on a twisted Cayley plane. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 51(1):114–124, March 2008. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Polo-Blanco:2008:ERC

[3632] Irene Polo-Blanco and Jaap Top. Explicit real cubic surfaces. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 51(1): 125–133, March 2008. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Rosales:2008:NSH

[3633] J. C. Rosales and P. A. García-Sánchez. Numerical semigroups having a Toms decomposition. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 51(1):134–139, March 2008. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Rossi:2008:FVB

[3634] Julio D. Rossi. First variations of the best Sobolev trace constant with respect to the domain. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 51(1):140–145, March 2008. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Zhou:2008:SSM

[3635] Xiaowen Zhou. Stepping-stone model with circular Brownian migration. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 51(1):146–160, March 2008. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Agarwal:2008:WIT

[3636] Ravi P. Agarwal, Victoria Otero-Espinar, Kanishka Perera, and Dolores R. Vivero. Wirtinger’s inequalities on time scales. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 51(2):161– 171, June 2008. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Alkan:2008:CLG

[3637] Emre Alkan and Alexandru Zaharescu. Consecutive large gaps in sequences defined by multiplicative constraints. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 51(2):172–181, June 2008. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic). See corrigendum [3764].

Bahturin:2008:GGA

[3638] Y. A. Bahturin and A. Giambruno. Group gradings on associative algebras with involution. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 51(2):182–194, June 2008. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Chen:2008:BBC

[3639] Huaihui Chen and Paul Gauthier. Boundedness from below of composition operators on α-Bloch spaces. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 51(2):195–204, June 2008. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Duda:2008:GDP

[3640] Jakub Duda. On Gâteaux differentiability of pointwise Lipschitz mappings. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 51(2):205–216, June 2008. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Filippakis:2008:MNS

[3641] Michael E. Filippakis and Nikolaos S. Papageorgiou. A multivalued nonlinear system with the vector p-Laplacian on the semi-infinity interval. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 51(2):217–228, June 2008. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Hanley:2008:ESP

[3642] Mary Hanley. Existence of solutions to Poisson’s equation. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 51(2): 229–235, June 2008. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Konovalov:2008:KLP

[3643] Victor N. Konovalov and Kirill A. Kopotun. Kolmogorov, linear and pseudo-dimensional widths of classes of s-monotone functions in L_{p}, 0 < p < 1. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 51(2):236–248, June 2008. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Mangoubi:2008:IRN

[3644] Dan Mangoubi. On the inner radius of a nodal domain. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 51(2):249– 260, June 2008. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Neeb:2008:CRQ

[3645] Karl-Hermann Neeb. On the classification of rational quantum tori and the structure of their automorphism groups. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 51(2):261–282, June 2008. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Ravindra:2008:NLT

[3646] G. V. Ravindra. The Noether–Lefschetz theorem via vanishing of coherent cohomology. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 51(2):283–290, June 2008. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Spinelli:2008:GAM

[3647] Ernesto Spinelli. Group algebras with minimal strong Lie derived length. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 51(2):291–297, June 2008. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Tocon:2008:KRI

[3648] Maribel Tocón. The Kostrikin radical and the invariance of the core of reduced extended affine Lie algebras. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 51(2):298–309, June 2008. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Witbooi:2008:RHR

[3649] P. J. Witbooi. Relative homotopy in relational structures. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 51(2):310– 320, June 2008. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Asaeda:2008:QMC

[3650] Marta Asaeda. Quantum multiple construction of subfactors. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 51(3): 321–333, September 2008. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Ascah-Coallier:2008:VDR

[3651] I. Ascah-Coallier and P. M. Gauthier. Value distribution of the Riemann zeta function. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 51(3):334–336, September 2008. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Bennett:2008:DBP

[3652] Michael A. Bennett. Differences between perfect powers. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 51(3):337– 347, September 2008. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Casazza:2008:RPB

[3653] Peter G. Casazza and Ole Christensen. The reconstruction property in Banach spaces and a perturbation theorem. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 51(3):348–358, September 2008. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Cho:2008:RHC

[3654] Jong Taek Cho and U-Hang Ki. Real hypersurfaces in complex space forms with Reeb flow symmetric structure Jacobi operator. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 51(3): 359–371, September 2008. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Ezquerro:2008:PII

[3655] J. A. Ezquerro and M. A. Hernández. Picard’s iterations for integral equations of mixed Hammerstein type. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 51(3):372–377, September 2008. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Izuchi:2008:CVS

[3656] Kou Hei Izuchi. Cyclic vectors in some weighted L^{p} spaces of entire functions. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 51(3):378–385, September 2008. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Lan:2008:PSF

[3657] K. Q. Lan and G. C. Yang. Positive solutions of the Falkner–Skan equation arising in the boundary layer theory. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 51(3):386–405, September 2008. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Mimouni:2008:CSC

[3658] Abdeslam Mimouni. Condensed and strongly condensed domains. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 51(3):406–412, September 2008. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

The:2008:BRD

[3659] L. Nguyen Van Thé. Big Ramsey degrees and divisibility in classes of ultrametric spaces. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 51(3):413–423, September 2008. CODEN CM-BUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Novelli:2008:NSB

[3660] Jean-Christophe Novelli and Jean-Yves Thibon. Noncommutative symmetric Bessel functions. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 51(3):424–438, September 2008. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Samei:2008:MSS

[3661] Karim Samei. On the maximal spectrum of semiprimitive multiplication modules. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 51(3):439–447, September 2008. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Sasahara:2008:SBL

[3662] Toru Sasahara. Stability of biharmonic Legendrian submanifolds in Sasakian space forms. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 51(3):448–459, September 2008. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Smoktunowicz:2008:PIG

[3663] Agata Smoktunowicz. On primitive ideals in graded rings. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 51(3): 460–466, September 2008. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Wang:2008:CVE

[3664] Yue Wang. Coupled vortex equations on complete Kähler manifolds. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 51(3):467–480, September 2008. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Bayart:2008:UIF

[3665] Frédéric Bayart. Universal inner functions on the ball. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 51(4):481– 486, December 2008. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Betancor:2008:LTT

[3666] Jorge J. Betancor, Teresa Martínez, and Lourdes Rodríguez-Mesa. Laplace transform type multipliers for Hankel transforms. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 51(4):487– 496, December 2008. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Borwein:2008:ENZ

[3667] Peter Borwein, Kwok-Kwong Stephen Choi, and Idris Mercer. Expected norms of zero-one polynomials. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 51(4):497–507, December 2008. CO-DEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Cavicchioli:2008:RST

[3668] Alberto Cavicchioli and Fulvia Spaggiari. A result in surgery theory. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 51(4):508–518, December 2008. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Coskun:2008:ECK

[3669] Izzet Coskun, Joe Harris, and Jason Starr. The effective cone of the Kontsevich moduli space. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 51(4):519–534, December 2008. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Csorba:2008:SHT

[3670] Péter Csorba. On the simple Z_{2}-homotopy types of graph complexes and their simple Z_{2}-universality. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 51(4):535–544, December 2008. CO-DEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Ionescu:2008:AAM

[3671] Marius Ionescu and Yasuo Watatani. C^{∗}-algebras associated with Mauldin–Williams graphs. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 51(4):545–560, December 2008. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Kuznetsov:2008:ERF

[3672] Alexey Kuznetsov. Expansion of the Riemann Ξ function in Meixner– Pollaczek polynomials. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 51(4):561–569, December 2008. CODEN CM-BUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Lutzer:2008:API

[3673] D. J. Lutzer, J. van Mill, and V. V. Tkachuk. Amsterdam properties of C_{p}(X) imply discreteness of X. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 51(4):570–578, December 2008. CO-DEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Matet:2008:GMS

[3674] Pierre Matet. Guessing with mutually stationary sets. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 51(4):579– 583, December 2008. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Purbhoo:2008:TPP

[3675] Kevin Purbhoo and Stephanie van Willigenburg. On tensor products of polynomial representations. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 51(4):584–592, December 2008. CO-DEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Roslanowski:2008:CS

[3676] Andrzej Rosłanowski and Juris Steprāns. Chasing silver. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 51(4):593– 603, December 2008. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Sliwa:2008:ISP

[3677] Wiesław Śliwa. The invariant subspace problem for non-Archimedean Banach spaces. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 51(4):604–617, December 2008. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Valmorin:2008:VTC

[3678] V. Valmorin. Vanishing theorems in Colombeau algebras of generalized functions. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 51(4):618–626, December 2008. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Vidanovic:2008:SSB

[3679] Mirjana V. Vidanović, Slobodan B. Tričković, and Miomir S. Stanković. Summation of series over Bourget functions. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 51(4):627–636, December 2008. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Anonymous:2008:AII

[3680] Anonymous. Author index — index des auteurs. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 51(4):637–??, December 2008. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287

(electronic). URL http://cms.math.ca/cmb/v51/p637.

Anonymous:2009:P

[3681] Anonymous. Photo. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 52(??):321–322, ???? 2009. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic). URL http://cms.math.

ca/cmb/v52/p321.

Anonymous:2009:AII

[3682] Anonymous. Author index — index des auteurs — for 2009 — pour 2009. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 52(??):637–??, ???? 2009. CODEN CMBUA3. ISSN 0008-4395 (print),

1496-4287 (electronic). URL http://cms.math.ca/cmb/v52/p637.

Banks:2009:CNS

[3683] W. D. Banks. Carmichael numbers with a square totient. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 52(1):3–8, March 2009. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Chasse:2009:SMO

[3684] Dominique Chassé and Yvan Saint-Aubin. On the spectrum of an n!×n! matrix originating from statistical mechanics. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 52(1):9–17, March 2009. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Chinea:2009:HHM

[3685] Domingo Chinea. Harmonicity of holomorphic maps between almost Hermitian manifolds. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 52(1):18–27, March 2009. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Choi:2009:RLW

[3686] Changsun Choi, Ju Myung Kim, and Keun Young Lee. Right and left weak approximation properties in Banach spaces. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 52(1):28–38, March 2009. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Cimpric:2009:RTA

[3687] Jakob Cimprič. A representation theorem for Archimedean quadratic modules on ∗-rings. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 52(1):39–52, March 2009. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Cummins:2009:CFL

[3688] C. J. Cummins. Cusp forms like Δ. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 52(1):53–62, March 2009. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Dietmann:2009:SZQ

[3689] Rainer Dietmann. Small zeros of quadratic forms avoiding a finite number of prescribed hyperplanes. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 52(1):63–65, March 2009. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Dryden:2009:HTH

[3690] Emily B. Dryden and Alexander Strohmaier. Huber’s Theorem for hyperbolic orbisurfaces. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 52(1):66–71, March 2009. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Duncan:2009:SB

[3691] Alexander Duncan, Michael LeBlanc, and David L. Wehlau. A SAGBI basis for F[V_{2} ⊕V_{2} ⊕V_{3}]^{C_{p}}. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 52(1):72–83, March 2009. CODEN CM-BUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Gauthier:2009:HTS

[3692] P. M. Gauthier and E. S. Zeron. Hartogs’ Theorem on separate holomorphicity for projective spaces. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 52(1):84–86, March 2009. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Lee:2009:HFV

[3693] Junho Lee. Holomorphic 2-forms and vanishing theorems for Gromov– Witten invariants. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 52(1):87–94, March 2009. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Miranian:2009:MVO

[3694] L. Miranian. Matrix valued orthogonal polynomials on the unit circle: Some extensions of the classical theory. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 52(1):95–104, March 2009. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Okoudjou:2009:GEB

[3695] Kasso A. Okoudjou, Luke G. Rogers, and Robert S. Strichartz. Generalized eigenfunctions and a Borel Theorem on the Sierpiński gasket. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 52(1):105–116, March 2009. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Poulakis:2009:RPC

[3696] Dimitrios Poulakis. On the rational points of the curve f(X, Y)^{q} = h(X)g(X, Y). Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 52(1):117–126, March 2009. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Shelah:2009:ERA

[3697] Saharon Shelah. The Erdős–Rado arrow for singular cardinals. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 52(1): 127–131, March 2009. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Shen:2009:PFM

[3698] Zhongmin Shen. On projectively flat (α, β)-metrics. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 52(1):132–144, March 2009. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Wang:2009:CR

[3699] Z. Wang and J. L. Chen. 2-clean rings. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 52(1):145–153, March 2009. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Ye:2009:BPT

[3700] Yasheng Ye and Min Ru. A Big Picard Theorem for holomorphic maps into complex projective space. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 52(1):154–160, March 2009. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Arcara:2009:NTR

[3701] D. Arcara and Y.-P. Lee. A new tautological relation in \overline{M}_{3,1} via the invariance constraint. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 52(2):161–174, June 2009. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Biswas:2009:CPP

[3702] Indranil Biswas. Connections on a parabolic principal bundle, II. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 52(2): 175–185, June 2009. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Broughan:2009:ERF

[3703] Kevin A. Broughan. Extension of the Riemann ξ-function’s logarithmic derivative positivity region to near the critical strip. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 52(2):186– 194, June 2009. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Garaev:2009:WPR

[3704] M. Z. Garaev, V. C. Garcia, and S. V. Konyagin. The Waring problem with the Ramanujan τ-function, II. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 52(2):195–199, June 2009. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Gatto:2009:SCG

[3705] Letterio Gatto and Taíse Santiago. Schubert calculus on a Grassmann algebra. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 52(2):200–212, June 2009. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Ghenciu:2009:DPP

[3706] Ioana Ghenciu and Paul Lewis. Dunford–Pettis properties and spaces of operators. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 52(2):213–223, June 2009. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Ghiloni:2009:ECD

[3707] Riccardo Ghiloni. Equations and complexity for the Dubois–Efroymson dimension theorem. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 52(2):224–236, June 2009. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Ghioca:2009:PSH

[3708] Dragos Ghioca. Points of small height on varieties defined over a function field. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 52(2):237–244, June 2009. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Goodaire:2009:IRL

[3709] Edgar G. Goodaire and César Polcino Milies. Involutions of RA loops. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 52(2):245–256, June 2009. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Ikeda:2009:ESG

[3710] Toru Ikeda. Essential surfaces in graph link exteriors. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 52(2):257– 266, June 2009. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Kosan:2009:ERH

[3711] Muhammet Tamer Koşan. Extensions of rings having McCoy condition. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 52(2):267–272, June 2009. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

MacDonald:2009:AC

[3712] John MacDonald and Laura Scull. Amalgamations of categories. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 52(2): 273–284, June 2009. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Parker:2009:GGC

[3713] John R. Parker and Ioannis D. Platis. Global geometrical coordinates on Falbel’s cross-ratio variety. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 52(2):285–294, June 2009. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Plotka:2009:FWG

[3714] Krzysztof Płotka. On functions whose graph is a Hamel basis, II. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 52(2): 295–302, June 2009. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Shelah:2009:C

[3715] Saharon Shelah. A comment on “√ < ⊔”. Bulletin canadien de mathé-

matiques = Canadian Mathematical Bulletin, 52(2):303–314, June 2009. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Yi:2009:GQC

[3716] Taishan Yi and Xingfu Zou. Generic quasi-convergence for essentially strongly order-preserving semiflows. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 52(2):315–320, June 2009. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Boroczky:2009:DTB

[3717] K. Böröczky, K. J. Böröczky, F. Fodor, H. Harborth, and W. Kuperberg. Dedication: Ted bisztriczky. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 52(3):323–326, September 2009. CO-DEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Berman:2009:GFC

[3718] Leah Wrenn Berman, Jürgen Bokowski, Branko Grünbaum, and Tomaž Pisanski. Geometric “floral” configurations. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 52(3):327–341, September 2009. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Bezdek:2009:XRN

[3719] K. Bezdek and Gy. Kiss. On the X-ray number of almost smooth convex bodies and of convex bodies of constant width. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 52(3):342–348, September 2009. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Campi:2009:PBO

[3720] Stefano Campi and Paolo Gronchi. On projection bodies of order one. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 52(3):349–360, September 2009. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Toth:2009:NCC

[3721] Gábor Fejes Tóth. A note on covering by convex bodies. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 52(3): 361–365, September 2009. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Gevay:2009:CCS

[3722] Gábor Gévay. A class of cellulated spheres with non-polytopal symmetries. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 52(3):366–379, September 2009. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Henk:2009:SMR

[3723] Martin Henk and María A. Hernández Cifre. Successive minima and radii. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 52(3):380–387, September 2009. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Heppes:2009:TRM

[3724] Aladár Heppes. Transversals with residue in moderately overlapping T (k)-families of translates. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 52(3):388–402, September 2009. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Jeronimo-Castro:2009:SRT

[3725] J. Jerónimo-Castro, L. Montejano, and E. Morales-Amaya. Shaken Rogers’s Theorem for homothetic sections. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 52(3):403–406, September 2009. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Langi:2009:BPC

[3726] Zsolt Lángi and Márton Naszódi. On the Bezdek–Pach conjecture for centrally symmetric convex bodies. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 52(3):407–415, September 2009. CO-DEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Malik:2009:HPG

[3727] Shabnam Malik, Ahmad Mahmood Qureshi, and Tudor Zamfirescu. Hamiltonian properties of generalized Halin graphs. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 52(3):416–423, September 2009. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Martini:2009:CDM

[3728] Horst Martini and Margarita Spirova. Covering discs in Minkowski planes. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 52(3):424–434, September 2009. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Monson:2009:MRA

[3729] B. Monson and Egon Schulte. Modular reduction in abstract polytopes. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 52(3):435–450, September 2009. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Pach:2009:IC

[3730] János Pach, Gábor Tardos, and Géza Tóth. Indecomposable coverings. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 52(3):451–463, September 2009. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Stancu:2009:TVP

[3731] Alina Stancu. Two volume product inequalities and their applications. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 52(3):464–480, September 2009. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Alaca:2009:SIP

[3732] Ayşe Alaca, Şaban Alaca, and Kenneth S. Williams. Some infinite products of Ramanujan type. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 52(4):481–492, December 2009. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Artebani:2009:ODF

[3733] Michela Artebani. A one-dimensional family of K3 surfaces with a Z_{4} action. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 52(4):493–510, December 2009. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Bonciocat:2009:IPO

[3734] Anca Iuliana Bonciocat and Nicolae Ciprian Bonciocat. The irreducibility of polynomials that have one large coefficient and take a prime value. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 52(4):511–520, December 2009. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Chen:2009:PLP

[3735] Yanping Chen and Yong Ding. The parabolic Littlewood–Paley operator with Hardy space kernels. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 52(4):521–534, December 2009. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Daigle:2009:NLN

[3736] Daniel Daigle and Shulim Kaliman. A note on locally nilpotent derivations and variables of k[X, Y, Z]. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 52(4):535–543, December 2009. CO-DEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Hanafy:2009:IFC

[3737] I. M. Hanafy. Intuitionistic fuzzy γ-continuity. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 52(4):544–554, December 2009. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Hirata:2009:BBS

[3738] Kentaro Hirata. Boundary behavior of solutions of the Helmholtz equation. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 52(4):555–563, December 2009. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Jin:2009:GAQ

[3739] Hai Lan Jin, Jaekyung Doh, and Jae Keol Park. Group actions on quasi-Baer rings. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 52(4):564–582, December 2009. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Konstantinou:2009:CPR

[3740] Elisavet Konstantinou and Aristides Kontogeorgis. Computing polynomials of the Ramanujan t_{n} class invariants. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 52(4):583–597, December 2009. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Moreno:2009:NSI

[3741] M. A. Moreno, J. Nicola, E. Pardo, and H. Thomas. Numerical semigroups that are not intersections of d-squashed semigroups. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 52(4): 598–612, December 2009. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Wulan:2009:LTC

[3742] Hasi Wulan and Kehe Zhu. Lipschitz type characterizations for Bergman spaces. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 52(4):613–626, December 2009. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Yu:2009:CFS

[3743] Dan Sheng Yu, Ping Zhou, and Song Ping Zhou. On L^{1}-convergence of Fourier series under the MVBV condition. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 52(4):627–636, December 2009. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Athanasiadis:2010:CRT

[3744] Christos A. Athanasiadis. A combinatorial reciprocity theorem for hyperplane arrangements. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 53(1):3–10, March 2010. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Burke:2010:AIE

[3745] Maxim R. Burke. Approximation and interpolation by entire functions of several variables. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 53(1):11–22, March 2010. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Chen:2010:BBM

[3746] Huaihui Chen and Minzhu Zhang. Boundedness from below of multiplication operators between α-Bloch spaces. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 53(1):23–36, March 2010. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Choi:2010:CPO

[3747] Man-Duen Choi and Frédéric Latrémolière. C^{∗}-crossed-products by an order-two automorphism. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 53(1):37–50, March 2010. CODEN CM-BUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Cobos:2010:RBI

[3748] Fernando Cobos and Luz M. Fernández-Cabrera. On the relationship between interpolation of Banach algebras and interpolation of bilinear operators. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 53(1):51–57, March 2010. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Dabrowski:2010:RFJ

[3749] Andrzej Dabrowski and Tomasz Jedrzejak. Ranks in families of Jacobian varieties of twisted Fermat curves. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 53(1):58–63, March 2010. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Dodos:2010:ASM

[3750] Pandelis Dodos. On antichains of spreading models of Banach spaces. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 53(1):64–76, March 2010. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Finston:2010:CAR

[3751] David Finston and Stefan Maubach. Constructing (almost) rigid rings and a UFD having infinitely generated Derksen and Makar–Limanov invariants. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 53(1):77–86, March 2010. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Ghioca:2010:ECP

[3752] Dragos Ghioca. Elliptic curves over the perfect closure of a function field. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 53(1):87–94, March 2010. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Ghioca:2010:TFM

[3753] Dragos Ghioca. Towards the full Mordell–Lang conjecture for Drinfeld modules. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 53(1):95–101, March 2010. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Khan:2010:SBI

[3754] Rizwanur Khan. Spacings between integers having typically many prime factors. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 53(1):102–117, March 2010. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Lewis:2010:US

[3755] Paul Lewis. The uncomplemented spaces W (X, Y) and K(X, Y). Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 53 (1):118–121, March 2010. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Mo:2010:CFM

[3756] Xiaohuan Mo and Linfeng Zhou. A class of Finsler metrics with bounded Cartan torsion. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 53(1):122–132, March 2010. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Moritoh:2010:FDE

[3757] Shinya Moritoh and Kyoko Tomoeda. A further decay estimate for the Dziubański–Hernández wavelets. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 53(1):133–139, March 2010. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Mukunda:2010:PNP

[3758] Keshav Mukunda. Pisot numbers from {0, 1}-polynomials. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 53(1):140– 152, March 2010. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Niu:2010:SHT

[3759] Pengcheng Niu, Yafei Ou, and Junqiang Han. Several Hardy type inequalities with weights related to generalized Greiner operator. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 53 (1):153–162, March 2010. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Taylor:2010:VAS

[3760] Michael Taylor. Variants of Arnold’s stability results for 2D Euler equations. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 53(1):163–170, March 2010. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Thomas:2010:MFS

[3761] Hugh Thomas and Alexander Yong. Multiplicity-free Schubert calculus. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 53(1):171–186, March 2010. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Unver:2010:LUF

[3762] Sinan Ünver. On the local unipotent fundamental group scheme. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 53(1): 187–192, March 2010. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Agarwal:2010:OSO

[3763] Ravi P. Agarwal, Cezar Avramescu, and Octavian G. Mustafa. On the oscillation of a second order strictly sublinear differential equation. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 53 (2):193–203, June 2010. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Alkan:2010:CCL

[3764] Emre Alkan and Alexandru Zaharescu. Corrigendum: “Consecutive large gaps in sequences defined by multiplicative constraints” [MR2414205]. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 53(2):204–205, June 2010. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic). See [3637].

Atceken:2010:SSS

[3765] Mehmet Atc.eken. Semi-slant submanifolds of an almost paracontact metric manifold. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 53(2):206–217, June 2010. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Biswas:2010:RTB

[3766] Indranil Biswas. Restriction of the tangent bundle of G/P to a hypersurface. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 53(2):218–222, June 2010. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Chuang:2010:DPM

[3767] Chen-Lian Chuang and Tsiu-Kwen Lee. Density of polynomial maps. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 53(2):223–229, June 2010. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Dogruoz:2010:MUC

[3768] S. Doğruöz, A. Harmanci, and P. F. Smith. Modules with unique closure relative to a torsion theory. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 53(2):230–238, June 2010. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Dong:2010:NEO

[3769] Z. Dong. A note on the exactness of operator spaces. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 53(2):239– 246, June 2010. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Etingof:2010:REF

[3770] P. Etingof, P. Malcolmson, and F. Okoh. Root extensions and factorization in affine domains. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 53(2):247–255, June 2010. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Fang:2010:EDI

[3771] Xiaochun Fang and Lin Wang. Equivalent definitions of infinite positive elements in simple C^{∗}-algebras. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 53(2):256–262, June 2010. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Feuto:2010:WNI

[3772] Justin Feuto, Ibrahim Fofana, and Konin Koua. Weighted norm inequalities for a maximal operator in some subspace of amalgams. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 53(2): 263–277, June 2010. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Galego:2010:CBS

[3773] Elói M. Galego. Cantor–Bernstein sextuples for Banach spaces. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 53(2): 278–285, June 2010. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Gorelic:2010:OB

[3774] Isaac Gorelic. Orders of π-bases. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 53(2):286–294, June 2010. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Guo:2010:GAD

[3775] Boling Guo and Zhaohui Huo. The global attractor of a damped, forced Hirota equation in H^{1}. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 53(2):295–310, June 2010. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Jasiczak:2010:RZS

[3776] Michał Jasiczak. Remark on zero sets of holomorphic functions in convex domains of finite type. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 53(2):311–320, June 2010. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Lee:2010:TUR

[3777] Tsiu-Kwen Lee and Yiqiang Zhou. A theorem on unit-regular rings. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 53(2):321–326, June 2010. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Luor:2010:MEI

[3778] Dah-Chin Luor. Multidimensional exponential inequalities with weights. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 53(2):327–339, June 2010. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Lusala:2010:RPS

[3779] Tsasa Lusala, Jedrzej Śniatycki, and Jordan Watts. Regular points of a subcartesian space. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 53(2):340–346, June 2010. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Papageorgiou:2010:MNS

[3780] Nikolaos S. Papageorgiou and Vasile Staicu. Multiple nontrivial solutions for doubly resonant periodic problems. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 53(2):347–359, June 2010. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Porter:2010:SHS

[3781] Jack Porter and Mohan Tikoo. Separating H-sets by open sets. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 53(2): 360–366, June 2010. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Porter:2010:SSO

[3782] Jack Porter and Mohan Tikoo. Separating H-sets by open sets. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 53(2): 360–366, June 2010. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Stamov:2010:APL

[3783] Gani Tr. Stamov. Almost periodicity and Lyapunov’s functions for impulsive functional differential equations with infinite delays. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 53(2): 367–377, June 2010. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Zhou:2010:NSC

[3784] Sizhong Zhou. A new sufficient condition for a graph to be (g, f, n)- critical. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 53(2):378–384, June 2010. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Achter:2010:ECS

[3785] Jeffrey D. Achter. Exceptional covers of surfaces. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 53(3):385–393, September 2010. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Averkov:2010:NES

[3786] Gennadiy Averkov. On nearly equilateral simplices and nearly l_{∞} spaces. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 53(3):394–397, September 2010. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Botelho:2010:PCH

[3787] Fernanda Botelho and James Jamison. Projections in the convex hull of surjective isometries. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 53(3):398–403, September 2010. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Broer:2010:ITA

[3788] Abraham Broer. Invariant theory of Abelian transvection groups. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 53(3):404–411, September 2010. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Calvaruso:2010:ELL

[3789] G. Calvaruso. Einstein-like Lorentz metrics and three-dimensional curvature homogeneity of order one. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 53(3):412–424, September 2010. CO-DEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Chapoton:2010:FPL

[3790] Frédéric Chapoton. Free pre-Lie algebras are free as Lie algebras. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 53(3):

425–437, September 2010. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Chigogidze:2010:NHN

[3791] A. Chigogidze and A. Nagórko. Near-homeomorphisms of Nöbeling manifolds. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 53(3):438–446, September 2010. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Choi:2010:ICO

[3792] Yemon Choi. Injective convolution operators on ℓ^{∞}(Γ) are surjective. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 53(3):447–452, September 2010. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Desgroseilliers:2010:SRT

[3793] Marc Desgroseilliers, Benoit Larose, Claudia Malvenuto, and Christelle Vincent. Some results on two conjectures of Schützenberger. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 53(3): 453–465, September 2010. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Dubarbie:2010:SMB

[3794] Luis Dubarbie. Separating maps between spaces of vector-valued absolutely continuous functions. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 53(3):466–474, September 2010. CO-DEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Jankowski:2010:NMB

[3795] Tadeusz Jankowski. Nonlinear multipoint boundary value problems for second order differential equations. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 53(3):475–490, September 2010. CO-DEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Huang:2010:WTE

[3796] Jizheng Huang and Heping Liu. The weak type (1, 1) estimates of maximal functions on the Laguerre hypergroup. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 53(3):491–502, September 2010. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Jizheng:2010:WTE

[3797] Huang Jizheng and Heping Liu. The weak type (1, 1) estimates of maximal functions on the Laguerre hypergroup. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 53(3):491–515, September 2010. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Kurenok:2010:TCM

[3798] V. P. Kurenok. The time change method and SDEs with nonnegative drift. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 53(3):503–515, September 2010. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Maurmann:2010:APM

[3799] Quinn Maurmann, Max Engelstein, Anthony Marcuccio, and Taryn Pritchard. Asymptotics of perimeter-minimizing partitions. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 53(3): 516–525, September 2010. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Milian:2010:SSP

[3800] Anna Milian. On some stochastic perturbations of semilinear evolution equations. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 53(3):526–533, September 2010. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Pambuccian:2010:ATT

[3801] Victor Pambuccian. Acute triangulation of a triangle in a general setting. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 53(3):534–541, September 2010. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Pintea:2010:SMH

[3802] Cornel Pintea. Smooth mappings with higher dimensional critical sets. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 53(3):542–549, September 2010. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Shalit:2010:RPS

[3803] Orr Moshe Shalit. Representing a product system representation as a contractive semigroup and applications to regular isometric dilations. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 53(3):550–563, September 2010. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Watanabe:2010:DNK

[3804] Yoshiyuki Watanabe and Young Jin Suh. On 6-dimensional nearly Kähler manifolds. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 53(3):564–570, September 2010. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Trifkovic:2010:PMF

[3805] Mak Trifković. Periods of modular forms and imaginary quadratic base change. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 53(3):571–576, September 2010. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Asgharzadeh:2010:UAL

[3806] Mohsen Asgharzadeh and Massoud Tousi. A unified approach to local cohomology modules using Serre classes. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 53(4):577–586, December 2010. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Birkenmeier:2010:HRE

[3807] Gary F. Birkenmeier, Jae Keol Park, and S. Tariq Rizvi. Hulls of ring extensions. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 53(4):587–601, December 2010. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Boij:2010:NDC

[3808] Mats Boij and Anthony Geramita. Notes on diagonal coinvariants of the dihedral group. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 53(4):602–613, December 2010. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Boroczky:2010:MWC

[3809] Károly J. Böröczky and Rolf Schneider. The mean width of circumscribed random polytopes. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 53(4):614–628, December 2010. CODEN CM-BUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Chinen:2010:ADP

[3810] Naotsugu Chinen and Tetsuya Hosaka. Asymptotic dimension of proper CAT(0) spaces that are homeomorphic to the plane. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 53(4):629–638, December 2010. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Coykendall:2010:GI

[3811] Jim Coykendall and Tridib Dutta. A generalization of integrality. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 53(4):639–653, December 2010. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Elliott:2010:VPE

[3812] P. D. T. A. Elliott. Variations on a paper of Erdős and Heilbronn. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 53(4):654–660, December 2010. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Johnstone:2010:CNE

[3813] Jennifer A. Johnstone and Blair K. Spearman. Congruent number of elliptic curves with rank at least three. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 53(4):661–666, December 2010. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Khashyarmanesh:2010:ERL

[3814] Kazem Khashyarmanesh. On the endomorphism rings of local cohomology modules. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 53(4):667–673, December 2010. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Kristaly:2010:MSC

[3815] Alexandru Kristály, Nikolaos S. Papageorgiou, and Csaba Varga. Multiple solutions for a class of Neumann elliptic problems on compact Riemannian manifolds with boundary. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 53(4):674–683, December 2010. CO-DEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Proctor:2010:IDI

[3816] Emily Proctor and Elizabeth Stanhope. An isospectral deformation on an infranil-orbifold. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 53(4):684–689, December 2010. CODEN CM-BUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Puerta:2010:MOI

[3817] M. E. Puerta and G. Loaiza. On the maximal operator ideal associated with a tensor norm defined by interpolation spaces. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 53(4):690–705, December 2010. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Roberts:2010:NRO

[3818] R. Roberts and J. Shareshian. Non-right-orderable 3-manifold groups. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 53(4):706–718, December 2010. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Stasyuk:2010:CEO

[3819] I. Stasyuk and E. D. Tymchatyn. A continuous extension operator for convex metrics. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 53(4):719–729, December 2010. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Theriault:2010:CWF

[3820] Stephen D. Theriault. A case when the fiber of the double suspension is the double loops on Anick’s space. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 53(4):730–736, December 2010. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Vougalter:2010:NIT

[3821] Vitali Vougalter. On the negative index theorem for the linearized non-linear Schrödinger problem. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 53(4):737–745, December 2010. CO-DEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Werner:2010:S

[3822] Caryn Werner. On surfaces with p_{g} = 0 and K^{2} = 5. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 53(4):746–756, December 2010. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Woo:2010:IPA

[3823] Alexander Woo. Interval pattern avoidance for arbitrary root systems. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 53(4):757–762, December 2010. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Bakonyi:2011:EPD

[3824] M. Bakonyi and D. Timotin. Extensions of positive definite functions on amenable groups. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 54(1):3–11, March 2011. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Bingham:2011:HKB

[3825] N. H. Bingham and A. J. Ostaszewski. Homotopy and the Kestelman– Borwein–Ditor theorem. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 54(1):12–20, March 2011. CODEN CM-BUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Bouali:2011:GDS

[3826] S. Bouali and M. Ech-chad. Generalized D-symmetric operators II. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 54 (1):21–27, March 2011. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Chang:2011:GSP

[3827] Yu-Hsien Chang and Cheng-Hong Hong. Generalized solution of the photon transport problem. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 54(1):28–38, March 2011. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Chapman:2011:ENS

[3828] S. T. Chapman, P. A. García-Sánchez, D. Llena, and J. Marshall. Elements in a numerical semigroup with factorizations of the same length. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 54(1):39–43, March 2011. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Cheung:2011:SSO

[3829] Wai-Shun Cheung and Tin-Yau Tam. Star-shapedness and K-orbits in complex semisimple Lie algebras. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 54(1):44–55, March 2011. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Dinh:2011:CVC

[3830] Thi Anh Thu Dinh. Characteristic varieties for a class of line arrangements. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 54(1):56–67, March 2011. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Eilers:2011:NSK

[3831] Søren Eilers, Gunnar Restorff, and Efren Ruiz. Non-splitting in Kirchberg’s ideal-related KK-theory. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 54(1):68–81, March 2011. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Emerson:2011:LNA

[3832] Heath Emerson. Lefschetz numbers for C^{∗}-algebras. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 54(1):82–99, March 2011. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Fan:2011:GMI

[3833] Dashan Fan and Huoxiong Wu. On the generalized Marcinkiewicz integral operators with rough kernels. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 54(1):100–112, March 2011. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Hytonen:2011:NBA

[3834] Tuomas P. Hytönen. On the norm of the Beurling–Ahlfors operator in several dimensions. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 54(1):113–125, March 2011. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Jin:2011:FSK

[3835] Yongyang Jin and Genkai Zhang. Fundamental solutions of Kohn sub-Laplacians on anisotropic Heisenberg groups and H-type groups. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 54(1): 126–140, March 2011. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Kim:2011:LMA

[3836] Sang Og Kim and Choonkil Park. Linear maps on C^{∗}-algebras preserving the set of operators that are invertible in A/I. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 54(1):141–146, March 2011. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Nelson:2011:GQP

[3837] Sam Nelson. Generalized Quandle polynomials. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 54(1):147–158, March 2011. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Sababheh:2011:HIR

[3838] Mohammad Sababheh. Hardy inequalities on the real line. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 54(1): 159–171, March 2011. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Shayya:2011:MFT

[3839] Bassam Shayya. Measures with Fourier transforms in L^{2} of a half-space. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 54(1):172–179, March 2011. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Spurny:2011:AFL

[3840] J. Spurný and M. Zelený. Additive families of low Borel classes and Borel measurable selectors. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 54(1):180–192, March 2011. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Bennett:2011:MSD

[3841] Harold Bennett and David Lutzer. Measurements and G_{δ}-subsets of domains. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 54(2):193–206, June 2011. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Chen:2011:BFI

[3842] Jiecheng Chen and Dashan Fan. A bilinear fractional integral on compact Lie groups. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 54(2):207–216, June 2011. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Chen:2011:RRS

[3843] William Y. C. Chen, Larry X. W. Wang, and Arthur L. B. Yang. Recurrence relations for strongly q-log-convex polynomials. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 54(2):217– 229, June 2011. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Clouatre:2011:UPS

[3844] Raphaël Clouâtre. Universal power series in C^{N}. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 54(2):230–236, June 2011. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Creedon:2011:SUG

[3845] Leo Creedon and Joe Gildea. The structure of the unit group of the group algebra F_{2^{k}}D_{8}. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 54(2):237–243, June 2011. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Daniel:2011:HSC

[3846] D. Daniel, J. Nikiel, L. B. Treybig, H. M. Tuncali, and E. D. Tymchatyn. Homogeneous Suslinian continua. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 54(2):244–248, June 2011. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

daSilva:2011:NAA

[3847] Paulo L. Dattori da Silva. A note about analytic solvability of complex planar vector fields with degeneracies. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 54(2):249–254, June 2011. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Dehaye:2011:IDB

[3848] Paul-Olivier Dehaye. On an identity due to Bump and Diaconis, and Tracy and Widom. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 54(2):255–269, June 2011. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Dow:2011:SOU

[3849] Alan Dow. Sequential order under PFA. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 54(2):270–276, June 2011. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Farley:2011:MSF

[3850] Jonathan David Farley. Maximal sublattices of finite distributive lattices. III: a conjecture from the 1984 Banff Conference on Graphs and Order. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 54(2):277–282, June 2011. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Hillman:2011:SM

[3851] J. A. Hillman and S. K. Roushon. Surgery on S̃L×E^{n}-manifolds. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 54(2):

283–287, June 2011. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Jacobs:2011:RCP

[3852] David P. Jacobs, Mohamed O. Rayes, and Vilmar Trevisan. The resultant of Chebyshev polynomials. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 54(2):288–296, June 2011. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Johnson:2011:LPP

[3853] Marianne Johnson and Ralph Stöhr. Lie powers and pseudo-idempotents. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 54(2):297–301, June 2011. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Kurka:2011:SSN

[3854] Ondrej Kůrka. Structure of the set of norm-attaining functionals on strictly convex spaces. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 54(2):302–310, June 2011. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Marzougui:2011:SRC

[3855] Habib Marzougui. Some remarks concerning the topological characterization of limit sets for surface flows. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 54(2):311–315, June 2011. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Mazhouda:2011:SPM

[3856] Kamel Mazhouda. The saddle-point method and the Li coefficients. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 54 (2):316–329, June 2011. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Mouhib:2011:BID

[3857] A. Mouhib. Sur la borne inférieure du rang du 2-groupe de classes de certains corps multiquadratiques. (French) [on the lower limit of the rank of the 2-group of classes of certain multiquadratic objects]. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 54(2): 330–337, June 2011. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Nakazi:2011:STU

[3858] Takahiko Nakazi. Szegő’s Theorem and uniform algebras. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 54(2):338– 346, June 2011. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Potapov:2011:HSP

[3859] D. Potapov and F. Sukochev. The Haar system in the preduals of hyperfinite factors. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 54(2):347–363, June 2011. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Preda:2011:LTA

[3860] Ciprian Preda and Petre Preda. Lyapunov Theorems for the asymptotic behavior of evolution families on the half-line. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 54(2):364–369, June 2011. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Stout:2011:MVH

[3861] Edgar Lee Stout. Manifold-valued holomorphic approximation. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 54(2): 370–380, June 2011. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Veluscek:2011:SNH

[3862] Dejan Veluscek. A short note on the higher level version of the Krull– Baer Theorem. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 54(2):381–384, June 2011. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Blackadar:2011:IRI

[3863] Bruce Blackadar and Eberhard Kirchberg. Irreducible representations of inner quasidiagonal C^{∗}-algebras. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 54(3):385–395, September 2011. CO-DEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Cho:2011:PGS

[3864] Jong Taek Cho, Jun ichi Inoguchi, and Ji-Eun Lee. Parabolic geodesics in Sasakian 3-manifolds. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 54(3):396–410, September 2011. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Davidson:2011:OAU

[3865] Kenneth R. Davidson and Alex Wright. Operator algebras with unique preduals. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 54(3):411–421, September 2011. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Perez:2011:TCS

[3866] Juan de Dios Pérez and Young Jin Suh. Two conditions on the structure Jacobi operator for real hypersurfaces in complex projective space. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 54(3):422–429, September 2011. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

DeLand:2011:CFL

[3867] Matthew DeLand. Complete families of linearly non-degenerate rational curves. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 54(3):430–441, September 2011. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Garcia:2011:NLT

[3868] Esther García, Miguel Gómez Lozano, and Erhard Neher. Nondegeneracy for Lie triple systems and Kantor pairs. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 54(3):442–455, September 2011. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Gustafson:2011:OSP

[3869] Karl Gustafson. On operator sum and product adjoints and closures. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 54(3):456–463, September 2011. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Hwang:2011:CCE

[3870] Tea-Yuan Hwang and Chin-Yuan Hu. A characterization of the compound-exponential type distributions. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 54(3):464–471, September 2011. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Iacono:2011:SMA

[3871] Donatella Iacono. A semiregularity map annihilating obstructions to deforming holomorphic maps. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 54(3):472–486, September 2011. CO-DEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Kong:2011:SPA

[3872] Xiangjun Kong. Some properties associated with adequate transversals. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 54(3):487–497, September 2011. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Mortad:2011:ACS

[3873] Mohammed Hichem Mortad. On the adjoint and the closure of the sum of two unbounded operators. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 54(3):498–505, September 2011. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Neamaty:2011:CSS

[3874] A. Neamaty and S. Mosazadeh. On the canonical solution of the Sturm– Liouville problem with singularity and turning point of even order. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 54(3):506–518, September 2011. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Neeb:2011:ECS

[3875] K. H. Neeb and I. Penkov. Erratum: “Cartan Subalgebras of gl_{∞}. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 54 (3):519, September 2011. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic). See [3375].

Polishchuk:2011:SHF

[3876] A. Polishchuk. Simple helices on Fano threefolds. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 54(3):520–526, September 2011. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Preda:2011:DEF

[3877] Ciprian Preda and Ciprian Sipos. On the dichotomy of the evolution families: a discrete-argument approach. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 54(3):527–537, September 2011. CO-DEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Srinivasan:2011:HM

[3878] Gopala Krishna Srinivasan and P. Zvengrowski. On the horizontal monotonicity of |Γ(s)|. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 54(3):538–543, September 2011. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Strungaru:2011:PDM

[3879] Nicolae Strungaru. Positive definite measures with discrete Fourier transform and pure point diffraction. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 54(3):544–555, September 2011. CO-DEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Teragaito:2011:CSB

[3880] Masakazu Teragaito. Cyclic surgery between toroidal surgeries. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 54(3): 556–560, September 2011. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Uren:2011:NTV

[3881] James J. Uren. A note on toric varieties associated with moduli spaces. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 54(3):561–565, September 2011. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Zhou:2011:NUR

[3882] Xiang-Jun Zhou, Lei Shi, and Ding-Xuan Zhou. Non-uniform randomized sampling for multivariate approximation by high order Parzen windows. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 54(3):566–576, September 2011. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Aqzzouz:2011:EDP

[3883] Belmesnaoui Aqzzouz. Erratum: The Duality Problem For The Class of AM-Compact Operators On Banach Lattices. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 54(4):577–??, December 2011. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Baoguo:2011:KTO

[3884] Jia Baoguo, Lynn Erbe, and Allan Peterson. Kiguradze-type oscillation theorems for second order superlinear dynamic equations on time scales. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 54(4):580–??, December 2011. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Camargo:2011:LIM

[3885] Javier Camargo. Lightness of induced maps and homeomorphisms. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 54 (4):607–??, December 2011. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Dibaei:2011:ANA

[3886] Mohammad T. Dibaei and Alireza Vahidi. Artinian and non-Artinian local cohomology modules. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 54(4):619–??, December 2011. CODEN CM-BUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Fiorenza:2011:MNT

[3887] Alberto Fiorenza, Babita Gupta, and Pankaj Jain. Mixed norm type Hardy inequalities. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 54(4):630–??, December 2011. CODEN CM-BUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Flores:2011:ECF

[3888] André Luiz Flores, J. Carmelo Interlando, and Trajano Pires da Nóbrega Neto. An extension of Craig’s family of lattices. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 54(4):645–??, December 2011. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Forrest:2011:NOI

[3889] Brian E. Forrest and Volker Runde. Norm one idempotent cb-multipliers with applications to the Fourier algebra in the cb-multiplier norm. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 54 (4):654–??, December 2011. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Haas:2011:AST

[3890] Ruth Haas and Aloysius G. Helminck. Admissible sequences for twisted involutions in Weyl groups. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 54(4):663–??, December 2011. CODEN CM-BUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Hammerlindl:2011:QIP

[3891] Andy Hammerlindl. Quasi-isometry and plaque expansiveness. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 54(4):

676–??, December 2011. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Jimenez-Vargas:2011:LIS

[3892] A. Jiménez-Vargas and Moisés Villegas-Vallecillos. 2-local isometries on spaces of Lipschitz functions. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 54(4):680–??, December 2011. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Lusala:2011:SSS

[3893] Tsasa Lusala and Jedrzej Śniatycki. Stratified subcartesian spaces. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 54 (4):693–??, December 2011. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Moonens:2011:NCF

[3894] Laurent Moonens. Nonconstant continuous functions whose tangential derivative vanishes along a smooth curve. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 54(4):706–??, December 2011. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Okassa:2011:SLR

[3895] Eugène Okassa. Symplectic Lie–Rinehart–Jacobi algebras and contact manifolds. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 54(4):716–??, December 2011. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Ostrovskii:2011:ABM

[3896] M. I. Ostrovskii. Auerbach bases and minimal volume sufficient enlargements. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 54(4):726–??, December 2011. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Samuels:2011:IMM

[3897] Charles L. Samuels. The infimum in the metric Mahler measure. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 54(4): 739–??, December 2011. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Shparlinski:2011:DIT

[3898] Igor E. Shparlinski. On the distribution of irreducible trinomials. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 54(4):

748–??, December 2011. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Sun:2011:CCF

[3899] Qingfeng Sun. Cancellation of cusp forms coefficients over Beatty sequences on GL(m). Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 54(4):757–??, December 2011. CODEN CM-BUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Agarwal:2012:LTA

[3900] Ravi P. Agarwal and Octavian G. Mustafa. On a local theory of asymptotic integration for nonlinear differential equations. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 55(1):3–??, March 2012. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Akiyama:2012:BCO

[3901] Shigeki Akiyama and Tomonari Suzuki. Browder’s convergence for oneparameter nonexpansive semigroups. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 55(1):15–??, March 2012. CO-DEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Bertin:2012:MMS

[3902] Marie José Bertin. A Mahler measure of a K3 surface expressed as a Dirichlet L-series. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 55(1):26–??, March 2012. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Butske:2012:ETD

[3903] William Butske. Endomorphisms of two dimensional Jacobians and related finite algebras. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 55(1):38–??, March 2012. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Chebolu:2012:FGH

[3904] Sunil K. Chebolu, J. Daniel Christensen, and Ján Minác. Freyd’s generating hypothesis for groups with periodic cohomology. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 55(1):48–??, March 2012. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Coons:2012:EST

[3905] Michael Coons. Extension of some theorems of W. Schwarz. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 55(1): 60–??, March 2012. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Dean:2012:CIL

[3906] Andrew J. Dean. Classification of inductive limits of outer actions of R on approximate circle algebras. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 55(1):73–??, March 2012. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Divaani-Aazar:2012:CGL

[3907] Kamran Divaani-Aazar and Alireza Hajikarimi. Cofiniteness of generalized local cohomology modules for one-dimensional ideals. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 55(1):81–??, March 2012. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Ghanbari:2012:IEG

[3908] K. Ghanbari and B. Shekarbeigi. Inequalities for eigenvalues of a general clamped plate problem. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 55(1):88–??, March 2012. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Glied:2012:SCI

[3909] Svenja Glied. Similarity and coincidence isometries for modules. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 55(1): 98–??, March 2012. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Guler:2012:SFP

[3910] Dincer Guler. On Segrè forms of positive vector bundles. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 55(1):108– ??, March 2012. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Kon:2012:CRH

[3911] S. H. Kon and Tee-How Loo. On characterizations of real hypersurfaces in a complex space form with η-parallel shape operator. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 55(1):

114–??, March 2012. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

LaGrange:2012:CTC

[3912] John D. LaGrange. Characterizations of three classes of zero-divisor graphs. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 55(1):127–??, March 2012. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Li:2012:PFF

[3913] Benling Li and Zhongmin Shen. Projectively flat fourth root Finsler metrics. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 55(1):138–??, March 2012. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Li:2012:CBS

[3914] Songxiao Li, Hasi Wulan, and Kehe Zhu. A characterization of Bergman spaces on the unit ball of C^{n}. II. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 55(1):146–??, March 2012. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Mafi:2012:ACG

[3915] Amir Mafi and Hero Saremi. Artinianness of certain graded local cohomology modules. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 55(1):153–??, March 2012. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Mine:2012:SSC

[3916] Kotaro Mine and Katsuro Sakai. Subdivisions of simplicial complexes preserving the metric topology. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 55(1):157–??, March 2012. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Pergher:2012:IF

[3917] Pedro L. Q. Pergher. Involutions fixing F^{n} ∪{Indecomposable}. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 55(1): 164–??, March 2012. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Rhoades:2012:HPM

[3918] B. E. Rhoades. Hausdorff prime matrices. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 55(1):172–??, March 2012. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Spirn:2012:LDD

[3919] Daniel Spirn and J. Douglas Wright. Linear dispersive decay estimates for the 3 + 1 dimensional water wave equation with surface tension. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 55(1): 176–??, March 2012. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Steinberg:2012:YAS

[3920] Benjamin Steinberg. Yet another solution to the Burnside problem for matrix semigroups. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 55(1):188–??, March 2012. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Ulas:2012:RPA

[3921] Maciej Ulas. Rational points in arithmetic progressions on y^{2} = x^{n} + k. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 55(1):193–??, March 2012. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Valenti:2012:AGU

[3922] Angela Valenti and Mikhail Zaicev. Abelian gradings on upper block triangular matrices. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 55(1):208–??, March 2012. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Wang:2012:PSI

[3923] Da-Bin Wang. Positive solutions of impulsive dynamic system on time scales. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 55(1):214–??, March 2012. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Bernardes:2012:LST

[3924] Nilson C. Bernardes. Limit sets of typical homeomorphisms. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 55(2): 225–??, June 2012. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Bishnoi:2012:AMV

[3925] Anuj Bishnoi and Sudesh K. Khanduja. On algebraically maximal valued fields and defectless extensions. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 55(2):233–??, June 2012. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Cegrell:2012:CC

[3926] Urban Cegrell. Convergence in capacity. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 55(2):242–??, June 2012. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Chang:2012:DES

[3927] Der-Chen Chang and Bao Qin Li. Description of entire solutions of eiconal type equations. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 55(2):249–??, June 2012. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Delvaux:2012:NAA

[3928] L. Delvaux, A. Van Daele, and Shuanhong Wang. A note on the antipode for algebraic quantum groups. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 55(2):260–??, June 2012. CO-DEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

DiVincenzo:2012:EGE

[3929] M. Onofrio Di Vincenzo and Vincenzo Nardozza. On the existence of the graded exponent for finite dimensional Z_{p}-graded algebras. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 55(2): 271–??, June 2012. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Eloe:2012:UIE

[3930] Paul W. Eloe, Johnny Henderson, and Rahmat Ali Khan. Uniqueness implies existence and uniqueness conditions for a class of (k + j)-point boundary value problems for n-th order differential equations. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 55(2): 285–??, June 2012. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Glasner:2012:GRP

[3931] Eli Glasner. The group Aut(μ) is Roelcke precompact. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 55(2):297–??, June 2012. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Han:2012:ADB

[3932] Yongsheng Han, Ming-Yi Lee, and Chin-Cheng Lin. Atomic decomposition and boundedness of operators on weighted Hardy spaces. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 55(2):

303–??, June 2012. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Hellus:2012:NVC

[3933] M. Hellus. A note on the vanishing of certain local cohomology modules. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 55(2):315–??, June 2012. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Jardine:2012:VHT

[3934] J. F. Jardine. The Verdier hypercovering theorem. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 55(2):319–??, June 2012. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Kamiya:2012:NDC

[3935] Shigeyasu Kamiya, John R. Parker, and James M. Thompson. Nondiscrete complex hyperbolic triangle groups of type (n, n, ∞; k). Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 55(2): 329–??, June 2012. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Loring:2012:MOI

[3936] Terry A. Loring. From matrix to operator inequalities. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 55(2):339–??, June 2012. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

MacDougall:2012:RHA

[3937] J. A. MacDougall and L. G. Sweet. Rational homogeneous algebras. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 55(2):351–??, June 2012. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Nhan:2012:CIW

[3938] Nguyen Du Vi Nhan and Dinh Thanh Duc. Convolution inequalities in l_{p} weighted spaces. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 55(2):355–??, June 2012. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Nie:2012:SCE

[3939] Zhaohu Nie. The secondary Chern–Euler class for a general submanifold. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 55(2):368–??, June 2012. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Oman:2012:MWP

[3940] Greg Oman and Adam Salminen. On modules whose proper homomorphic images are of smaller cardinality. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 55(2):378–??, June 2012. CO-DEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Riedl:2012:AIW

[3941] Jeffrey M. Riedl. Automorphisms of iterated wreath product p-groups. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 55(2):390–??, June 2012. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Sebbar:2012:ESM

[3942] Abdellah Sebbar and Ahmed Sebbar. Eisenstein series and modular differential equations. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 55(2):400–??, June 2012. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Service:2012:RTA

[3943] Robert Service. A Ramsey theorem with an application to sequences in Banach spaces. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 55(2):410–??, June 2012. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Vinh:2012:MSP

[3944] Le Anh Vinh. Maximal sets of pairwise orthogonal vectors in finite fields. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 55(2):418–??, June 2012. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Yang:2012:CRC

[3945] Jianbin Yang and Song Li. Convergence rates of cascade algorithms with infinitely supported masks. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 55(2):424–??, June 2012. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Zelator:2012:NDE

[3946] Konstantine Zelator. A note on the Diophantine equation x^{2} + y^{6} = z^{e}, e ≥ 4. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 55(2):435–??, June 2012. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Zorboska:2012:UIC

[3947] Nina Zorboska. Univalently induced, closed range, composition operators on the Bloch-type spaces. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 55(2):441–??, June 2012. CODEN CM-BUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Bahreini:2012:CSL

[3948] Manijeh Bahreini, Elizabeth Bator, and Ioana Ghenciu. Complemented subspaces of linear bounded operators. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 55(3):449–??, September 2012. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Campbell:2012:HCF

[3949] Peter S. Campbell and Anna Stokke. Hook-content formulae for symplectic and orthogonal tableaux. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 55(3):462–??, September 2012. CO-DEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Chen:2012:NRM

[3950] Bin Chen and Lili Zhao. A note on Randers metrics of scalar flag curvature. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 55(3):474–??, September 2012. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Deng:2012:WMS

[3951] Xinghua Deng and Robert V. Moody. Weighted model sets and their higher point-correlations. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 55(3):487–??, September 2012. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Fradelizi:2012:SEB

[3952] Matthieu Fradelizi, Grigoris Paouris, and Carsten Schütt. Simplices in the Euclidean Ball. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 55(3):498–??, September 2012. CODEN CM-BUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Gauthier:2012:DIH

[3953] P. M. Gauthier and V. Nestoridis. Domains of injective holomorphy. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 55(3):509–??, September 2012. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Iwase:2012:MSF

[3954] Norio Iwase, Mamoru Mimura, Nobuyuki Oda, and Yeon Soo Yoon. The Milnor–Stasheff filtration on spaces and generalized cyclic maps. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 55(3): 523–??, September 2012. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Kang:2012:APS

[3955] Dongsheng Kang. Asymptotic properties of solutions to semilinear equations involving multiple critical exponents. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 55(3):537–??, September 2012. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Lewis:2012:NCS

[3956] Paul Lewis and Polly Schulle. Non-complemented spaces of operators, vector measures, and c_{o}. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 55(3):548–??, September 2012. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Michalowski:2012:WBP

[3957] Nicholas Michalowski, David J. Rule, and Wolfgang Staubach. Weighted L^{p} boundedness of pseudodifferential operators and applications. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 55 (3):555–??, September 2012. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Miller:2012:GKT

[3958] A. R. Miller and R. B. Paris. A generalised Kummer-type transformation for the _{p}F_{p}(x) hypergeometric function. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 55(3):571–??, September 2012. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Ndogmo:2012:CON

[3959] J. C. Ndogmo. Casimir operators and nilpotent radicals. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 55(3):579–??, September 2012. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Nie:2012:SSC

[3960] Zhaohu Nie. On Sha’s secondary Chern–Euler class. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 55(3):586–??, September 2012. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Osekowski:2012:SID

[3961] Adam Osekowski. Sharp inequalities for differentially subordinate harmonic functions and martingales. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 55(3):597–??, September 2012. CO-DEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Ozgur:2012:CIS

[3962] Cihan Özgür and Adela Mihai. Chen inequalities for submanifolds of real space forms with a semi-symmetric non-metric connection. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 55(3): 611–??, September 2012. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic). See correction [4193].

Pan:2012:CDN

[3963] Jiaqing Pan. The continuous dependence on the nonlinearities of solutions of fast diffusion equations. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 55(3):623–??, September 2012. CO-DEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Pigola:2012:CMM

[3964] S. Pigola and M. Rimoldi. Characterizations of model manifolds by means of certain differential systems. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 55(3):632–??, September 2012. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Zhou:2012:MCL

[3965] Jiang Zhou and Bolin Ma. Marcinkiewicz commutators with Lipschitz functions in non-homogeneous spaces. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 55(3):646–??, September 2012. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Zhou:2012:OTI

[3966] Chunqin Zhou. An onofri-type inequality on the sphere with two conical singularities. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 55(3):663–??, September 2012. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Aizenbud:2012:MFJ

[3967] Avraham Aizenbud and Dmitry Gourevitch. Multiplicity free Jacquet modules. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 55(4):673–??, December 2012. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Berndt:2012:PEF

[3968] Ryan Berndt. A pointwise estimate for the Fourier transform and maxima of a function. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 55(4):689–??, December 2012. CODEN CM-BUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Borwein:2012:CUC

[3969] Jonathan M. Borwein and Jon Vanderwerff. Constructions of uniformly convex functions. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 55(4):697–707, December 2012. CODEN CM-BUA3. ISSN 0008-4395 (print), 1496-4287 (electronic). URL http://

docserver.carma.newcastle.edu.au/778/.

Demeter:2012:IRR

[3970] Ciprian Demeter. Improved range in the return times theorem. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 55(4): 708–??, December 2012. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Gigli:2012:FVF

[3971] Nicola Gigli and Shin-Ichi Ohta. First variation formula in Wasserstein spaces over compact Alexandrov spaces. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 55(4):723–??, December 2012. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Hernandez:2012:ESA

[3972] Eduardo Hernández and Donal O’Regan. Existence of solutions for abstract non-autonomous neutral differential equations. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 55(4):736–??, December 2012. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Hickel:2012:AHS

[3973] M. Hickel and G. Rond. Approximation of holomorphic solutions of a system of real analytic equations. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 55(4):752–??, December 2012. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Li:2012:SAL

[3974] Hanfeng Li. Smooth approximation of Lipschitz projections. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 55(4): 762–??, December 2012. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Martini:2012:ZCN

[3975] Horst Martini and Senlin Wu. On Zindler curves in normed planes. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 55 (4):767–??, December 2012. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Mollin:2012:PEN

[3976] R. A. Mollin and A. Srinivasan. Pell equations: Non-principal Lagrange criteria and central norms. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 55(4):774–??, December 2012. CODEN CM-BUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Motallebi:2012:PDS

[3977] M. R. Motallebi and H. Saiflu. Products and direct sums in locally convex cones. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 55(4):783–??, December 2012. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Novelli:2012:MCL

[3978] Carla Novelli and Gianluca Occhetta. Manifolds covered by lines and extremal rays. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 55(4):799–??, December 2012. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Oberlin:2012:RRT

[3979] Daniel M. Oberlin. Restricted Radon transforms and projections of planar sets. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 55(4):815–??, December 2012. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Perez-Garcia:2012:NEN

[3980] C. Perez-Garcia and W. H. Schikhof. New examples of non-Archimedean Banach spaces and applications. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 55(4):821–??, December 2012. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Reinhold:2012:AEC

[3981] Karin Reinhold, Anna K. Savvopoulou, and Christopher M. Wedrychowicz. Almost everywhere convergence of convolution measures. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 55(4): 830–??, December 2012. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Sairaiji:2012:RJV

[3982] Fumio Sairaiji and Takuya Yamauchi. The rank of Jacobian varieties over the maximal Abelian extensions of number fields: Towards the Frey–Jarden Conjecture. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 55(4):842–??, December 2012. CODEN CM-BUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Shparlinski:2012:CSD

[3983] Igor E. Shparlinski and Katherine E. Stange. Character sums with division polynomials. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 55(4):850–??, December 2012. CODEN CM-BUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

vonRenesse:2012:OTV

[3984] Max-K. von Renesse. An optimal transport view of Schrödinger’s equation. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 55(4):858–??, December 2012. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Wang:2012:LIE

[3985] Hui Wang and Shaoqiang Deng. Left invariant Einstein–Randers metrics on compact Lie groups. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 55(4):870–??, December 2012. CODEN CM-BUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Xueli:2012:ESE

[3986] Song Xueli and Peng Jigen. Equivalence of L_{p} stability and exponential stability of nonlinear Lipschitzian semigroups. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 55(4):882–??, December 2012. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Aissiou:2013:SLE

[3987] Tayeb Aïssiou. Semiclassical limits of eigenfunctions on flat ndimensional tori. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 56(1):3–??, March 2013. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Alon:2013:ORU

[3988] Gil Alon and Gady Kozma. Ordering the representations of S_{n} using the interchange process. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 56(1):13–??, March 2013. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Ayuso:2013:DVR

[3989] Fortuny P. Ayuso. Derivations and valuation rings. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 56(1):31–??, March 2013. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Amara:2013:CTC

[3990] Jamel Ben Amara. Comparison theorem for conjugate points of a fourthorder linear differential equation. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 56(1):39–??, March 2013. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Biswas:2013:PPP

[3991] Indranil Biswas and Arijit Dey. Polystable parabolic principal G-bundles and hermitian-einstein connections. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 56(1):44–??, March 2013. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Bouziad:2013:CQT

[3992] A. Bouziad. Cliquishness and quasicontinuity of two-variable maps. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 56 (1):55–??, March 2013. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Ghenciu:2013:US

[3993] Ioana Ghenciu. The uncomplemented subspace K(X, Y). Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 56(1):65–??, March 2013. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Hrubes:2013:ABC

[3994] P. Hrubes, A. Wigderson, and A. Yehudayoff. An asymptotic bound on the composition number of integer sums of squares formulas. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 56(1): 70–??, March 2013. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Islam:2013:TFP

[3995] Muhammad N. Islam. Three fixed point theorems: Periodic solutions of a Volterra type integral equation with infinite heredity. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 56(1):80–??, March 2013. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Jacob:2013:PCM

[3996] Benoît Jacob. On perturbations of continuous maps. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 56(1):92–??, March 2013. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Kong:2013:EAE

[3997] Qingkai Kong and Min Wang. Eigenvalue approach to even order system periodic boundary value problems. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 56(1):102–??, March 2013. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Krepski:2013:CEL

[3998] Derek Krepski. Central extensions of loop groups and obstruction to prequantization. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 56(1):116–??, March 2013. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Li:2013:EEA

[3999] Junfang Li. Evolution of eigenvalues along rescaled Ricci flow. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 56(1): 127–??, March 2013. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Munteanu:2013:CEH

[4000] Radu-Bogdan Munteanu. On constructing ergodic hyperfinite equivalence relations of non-product type. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 56(1):136–??, March 2013. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Oukhaba:2013:GCI

[4001] Hassan Oukhaba and Stéphane Viguié. On the Gras conjecture for imaginary quadratic fields. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 56(1):148–??, March 2013. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Rego:2013:EDD

[4002] L. C. Rêgo and R. J. Cintra. An extension of the Dirichlet density for sets of Gaussian integers. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 56(1):161–??, March 2013. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Sahin:2013:SIS

[4003] Bayram Sahin. Semi-invariant submersions from almost Hermitian manifolds. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 56(1):173–??, March 2013. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Shen:2013:SNR

[4004] Zhongmin Shen. On some non-Riemannian quantities in Finsler geometry. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 56(1):184–??, March 2013. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Stefansson:2013:SLZ

[4005] Úlfar F. Stefánsson. On the smallest and largest zeros of Müntz–Legendre polynomials. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 56(1):194–??, March 2013. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Tall:2013:PLS

[4006] Franklin D. Tall. Productively Lindelöf spaces may all be D. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 56(1): 203–??, March 2013. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Tausk:2013:LCN

[4007] Daniel V. Tausk. A locally compact non divisible Abelian group whose character group is torsion free and divisible. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 56(1):213–??, March 2013. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Yang:2013:FEF

[4008] Dilian Yang. Functional equations and Fourier analysis. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 56(1):218– ??, March 2013. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Agashe:2013:NVT

[4009] Amod Agashe. On the notion of visibility of torsors. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 56(2):225– ??, June 2013. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Arvanitidis:2013:COH

[4010] Athanasios G. Arvanitidis and Aristomenis G. Siskakis. Cesàro operators on the Hardy spaces of the half-plane. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 56(2):229–??, June 2013. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Betsakos:2013:VSL

[4011] Dimitrios Betsakos and Stamatis Pouliasis. Versions of Schwarz’s lemma for condenser capacity and inner radius. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 56(2):241–??, June 2013. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Borwein:2013:SCL

[4012] Peter Borwein, Stephen K. K. Choi, and Himadri Ganguli. Sign changes of the Liouville function on quadratics. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 56(2):251–??, June 2013. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Chandoul:2013:SPE

[4013] A. Chandoul, M. Jellali, and M. Mkaouar. The smallest Pisot element in the field of formal power series over a finite field. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 56(2):258– ??, June 2013. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Chen:2013:EDG

[4014] Yichao Chen, Toufik Mansour, and Qian Zou. Embedding distributions of generalized fan graphs. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 56(2):265–??, June 2013. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Cheng:2013:SWC

[4015] Lixin Cheng, Zhenghua Luo, and Yu Zhou. On super weakly compact convex sets and representation of the dual of the normed semigroup they generate. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 56(2):272–??, June 2013. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Coons:2013:TSC

[4016] Michael Coons. Transcendental solutions of a class of minimal functional equations. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 56(2):283–??, June 2013. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Dai:2013:QMM

[4017] Mei-Feng Dai. Quasisymmetrically minimal Moran sets. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 56(2):292–??, June 2013. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Perez:2013:RHC

[4018] Juan de Dios Pérez and Young Jin Suh. Real hypersurfaces in complex projective space whose structure Jacobi operator is Lie D-parallel. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 56(2):306–??, June 2013. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Dorais:2013:NCF

[4019] François G. Dorais. A note on conjectures of F. Galvin and R. Rado. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 56(2):317–??, June 2013. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Erdogan:2013:RFT

[4020] M. Burak Erdogan and Daniel M. Oberlin. Restricting Fourier transforms of measures to curves in R^{2}. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 56(2):326–??, June 2013. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Fan:2013:CPM

[4021] Qingzhai Fan. Certain properties of K_{0}-monoids preserved by tracial approximation. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 56(2):337–??, June 2013. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Goodaire:2013:IAG

[4022] Edgar G. Goodaire and César Polcino Milies. Involutions and anticommutativity in group rings. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 56(2):344–??, June 2013. CODEN CM-BUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Hare:2013:SRC

[4023] Kathryn E. Hare, Franklin Mendivil, and Leandro Zuberman. The sizes of rearrangements of Cantor sets. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 56(2):354–??, June 2013. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Kyritsi:2013:MSN

[4024] Sophia Th. Kyritsi and Nikolaos S. Papageorgiou. Multiple solutions for nonlinear periodic problems. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 56(2):366–??, June 2013. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Ma:2013:STG

[4025] Li Ma and Jing Wang. Sharp threshold of the Gross–Pitaevskii equation with trapped dipolar quantum gases. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 56(2):378–??, June 2013. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Mursaleen:2013:AMN

[4026] M. Mursaleen. Application of measure of noncompactness to infinite systems of differential equations. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 56(2):388–??, June 2013. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Oancea:2013:CAM

[4027] D. Oancea. Coessential abelianization morphisms in the category of groups. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 56(2):395–??, June 2013. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Prunaru:2013:FTM

[4028] Bebe Prunaru. A factorization theorem for multiplier algebras of reproducing kernel Hilbert spaces. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 56(2):400–??, June 2013. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Rad:2013:DZD

[4029] Nader Jafari Rad, Sayyed Heidar Jafari, and Doost Ali Mojdeh. On domination in zero-divisor graphs. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 56(2):407–??, June 2013. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Sanders:2013:SSL

[4030] T. Sanders. Structure in sets with logarithmic doubling. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 56(2):412–??, June 2013. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Thom:2013:CSD

[4031] Andreas Thom. Convergent sequences in discrete groups. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 56(2):424–??, June 2013. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Wnuk:2013:SRA

[4032] Witold Wnuk. Some remarks on the algebraic sum of ideals and Riesz subspaces. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 56(2):434–??, June 2013. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Zelenyuk:2013:CLI

[4033] Yevhen Zelenyuk. Closed left ideal decompositions of U(G). Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 56(2): 442–??, June 2013. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Akbari:2013:CIG

[4034] S. Akbari, M. Chavooshi, M. Ghanbari, and S. Zare. The f-chromatic index of a graph whose f-core has maximum degree 2. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 56(3):449–??, September 2013. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Athavale:2013:CMS

[4035] Ameer Athavale and Pramod Patil. On certain multivariable subnormal weighted shifts and their duals. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 56(3):459–??, September 2013. CO-DEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Aulaskari:2013:IRN

[4036] Rauno Aulaskari and Jouni Rättyä. Inclusion relations for new function spaces on Riemann surfaces. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 56(3):466–??, September 2013. CO-DEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Ayadi:2013:HAG

[4037] Adlene Ayadi. Hypercyclic Abelian groups of affine maps on C^{n}. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 56(3): 477–??, September 2013. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Bahmanpour:2013:NHD

[4038] Kamal Bahmanpour. A note on homological dimensions of Artinian local cohomology modules. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 56(3):491–??, September 2013. CODEN CM-BUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Browning:2013:LWE

[4039] T. D. Browning. The Lang–Weil estimate for cubic hypersurfaces. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 56 (3):500–??, September 2013. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Bu:2013:WSC

[4040] Qingying Bu. Weak sequential completeness of K(X , Y). Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 56(3):503–??, September 2013. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Dubickas:2013:LFM

[4041] Arturas Dubickas. Linear forms in monic integer polynomials. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 56(3): 510–??, September 2013. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Elbasraoui:2013:EFS

[4042] Abdelkrim Elbasraoui and Abdellah Sebbar. Equivariant forms: Structure and geometry. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 56(3):520–??, September 2013. CODEN CM-BUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Filali:2013:CPI

[4043] M. Filali and M. Sangani Monfared. A cohomological property of πinvariant elements. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 56(3):534–??, September 2013. CODEN CM-BUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Gauthier:2013:UOP

[4044] P. M. Gauthier. Universally overconvergent power series via the Riemann zeta-function. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 56(3):544–??, September 2013. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Handelman:2013:RDG

[4045] David Handelman. Real dimension groups. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 56(3):551–??, September 2013. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Herzog:2013:ZIC

[4046] Ivo Herzog. Ziegler’s indecomposability criterion. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 56(3):564–??, September 2013. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Hoang:2013:CCB

[4047] Giabao Hoang and Wendell Ressler. Conjugacy classes and binary quadratic forms for the Hecke groups. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 56(3):570–??, September 2013. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Liau:2013:ACS

[4048] Pao-Kuei Liau and Cheng-Kai Liu. On automorphisms and commutativity in semiprime rings. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 56(3):584–??, September 2013. CODEN CM-BUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Liu:2013:NIO

[4049] Congwen Liu and Lifang Zhou. On the p-norm of an integral operator in the half plane. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 56(3):593–??, September 2013. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Louboutin:2013:RCP

[4050] Stéphane R. Louboutin. Resultants of Chebyshev polynomials: A short proof. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 56(3):602–??, September 2013. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Mazorchuk:2013:CSH

[4051] Volodymyr Mazorchuk and Kaiming Zhao. Characterization of simple highest weight modules. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 56(3):606–??, September 2013. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Sevim:2013:RMC

[4052] Esra Sengelen Sevim and Zhongmin Shen. Randers metrics of constant scalar curvature. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 56(3):615–??, September 2013. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Shang:2013:OCS

[4053] Yilun Shang. Optimal control strategies for virus spreading in inhomogeneous epidemic dynamics. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 56(3):621–??, September 2013. CO-DEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Sundar:2013:ISS

[4054] S. Sundar. Inverse semigroups and Sheu’s groupoid for the odd dimensional quantum spheres. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 56(3):630–??, September 2013. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Turkmen:2013:RIC

[4055] Inan Utku Türkmen. Regulator indecomposable cycles on a product of elliptic curves. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 56(3):640–??, September 2013. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Valverde:2013:IRD

[4056] Cesar Valverde. On induced representations distinguished by orthogonal groups. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 56(3):647–??, September 2013. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Yu:2013:AUT

[4057] Zhi-Xian Yu and Ming Mei. Asymptotics and uniqueness of travelling waves for non-monotone delayed systems on 2D lattices. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 56(3):659–??, September 2013. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Ayadi:2013:DAC

[4058] K. Ayadi, M. Hbaib, and F. Mahjoub. Diophantine approximation for certain algebraic formal power series in positive characteristic. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 56(4): 673–??, December 2013. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Nikseresht:2013:EDM

[4059] A. Nikseresht and A. Azizi. Envelope dimension of modules and the simplified radical formula. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 56(4):683–??, December 2013. CODEN CM-BUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Banks:2013:CMC

[4060] William D. Banks, Ahmet M. Güloğlu, and Aaron M. Yeager. Carmichael meets Chebotarev. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 56(4):695–??, December 2013. CODEN CM-BUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Bartosova:2013:UMF

[4061] Dana Bartosová. Universal minimal flows of groups of automorphisms of uncountable structures. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 56(4):709–??, December 2013. CODEN CM-BUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Berczes:2013:SDN

[4062] Attila Bérczes and Florian Luca. On the sum of digits of numerators of Bernoulli numbers. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 56(4):723–??, December 2013. CODEN CM-BUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Currey:2013:ODP

[4063] B. Currey and A. Mayeli. The orthonormal dilation property for abstract Parseval wavelet frames. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 56(4):729–??, December 2013. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Elliott:2013:RCC

[4064] George A. Elliott and Zhuang Niu. On the radius of comparison of a commutative C∗-algebra. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 56(4):737–??, December 2013. CODEN CM-BUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Fu:2013:DFS

[4065] Xiaoye Fu and Jean-Pierre Gabardo. Dimension functions of self-affine scaling sets. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 56(4):745–??, December 2013. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Issa:2013:GTB

[4066] Zahraa Issa and Matilde Lalín. A generalization of a theorem of Boyd and Lawton. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 56(4):759–??, December 2013. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Lahiri:2013:NZV

[4067] Indrajit Lahiri and Imrul Kaish. A non-zero value shared by an entire function and its linear differential polynomials. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 56(4):769–??, December 2013. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Liu:2013:SPS

[4068] Zhixin Liu. Small prime solutions to cubic Diophantine equations. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 56 (4):785–??, December 2013. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

MacDonald:2013:UBE

[4069] Mark L. MacDonald. Upper bounds for the essential dimension of E_{7}. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 56(4):795–??, December 2013. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Oberlin:2013:ECM

[4070] Richard Oberlin. Estimates for compositions of maximal operators with singular integrals. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 56(4):801–??, December 2013. CODEN CM-BUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Petridis:2013:QLE

[4071] Yiannis N. Petridis, Nicole Raulf, and Morten S. Risager. Quantum limits of Eisenstein series and scattering states. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 56(4):814–??, December 2013. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic). See erratum [4072].

Petridis:2013:EQL

[4072] Yiannis N. Petridis, Nicole Raulf, and Morten S. Risager. Erratum to “Quantum Limits of Eisenstein Series and Scattering States”. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 56(4): 827–??, December 2013. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic). See [4071].

Pollack:2013:MTB

[4073] Paul Pollack. On Mertens’ Theorem for Beurling primes. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 56(4):829–??, December 2013. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Shparlinski:2013:ANS

[4074] Igor E. Shparlinski. On the average number of Square-free values of polynomials. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 56(4):844–??, December 2013. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Teragaito:2013:LOE

[4075] Masakazu Teragaito. Left-orderability and exceptional Dehn surgery on twist knots. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 56(4):850–??, December 2013. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

vanMill:2013:CDH

[4076] Jan van Mill. On countable dense and n-homogeneity. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 56(4):860–??, December 2013. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Wei:2013:NKP

[4077] Changguo Wei. Note on Kasparov product of C^{∗}-algebra extensions. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 56 (4):870–??, December 2013. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Xie:2013:FGG

[4078] BaoHua Xie, JieYan Wang, and YuePing Jiang. Free groups generated by two Heisenberg translations. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 56(4):881–??, December 2013. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Adamczak:2014:SPP

[4079] Radosław Adamczak, Rafal Latala, Alexander E. Litvak, Krzysztof Oleszkiewicz, Alain Pajor, and Nicole Tomczak-Jaegermann. A short proof of Paouris’ Inequality. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 57(1):3–??, March 2014. CODEN CM-BUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Alperin:2014:ISC

[4080] Roger C. Alperin and Brian L. Peterson. Integral sets and the center of a finite group. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 57(1):9–??, March 2014. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Aribi:2014:CES

[4081] Amine Aribi, Sorin Dragomir, and Ahmad El Soufi. On the continuity of the eigenvalues of a sublaplacian. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 57(1):12–??, March 2014. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Bourin:2014:SIC

[4082] Jean-Christophe Bourin, Tetsuo Harada, and Eun-Young Lee. Subadditivity inequalities for compact operators. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 57(1):25–??, March 2014. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Dashti:2014:CAL

[4083] Mahshid Dashti, Rasoul Nasr-Isfahani, and Sima Soltani Renani. Character amenability of Lipschitz algebras. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 57(1):37–??, March 2014. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Fonf:2014:CUS

[4084] Vladimir P. Fonf and Clemente Zanco. Covering the unit sphere of certain Banach spaces by sequences of slices and balls. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 57(1):42–??, March 2014. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Fosner:2014:JDF

[4085] Ajda Fosner and Tsiu-Kwen Lee. Jordan ∗-derivations of finitedimensional semiprime algebras. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 57(1):51–??, March 2014. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Geschke:2014:DCN

[4086] Stefan Geschke. 2-dimensional convexity numbers and P_{4}-free graphs. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 57(1):61–??, March 2014. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Grari:2014:ACA

[4087] A. Grari. Un anneau commutatif associé à un design symétrique. (French) [A commutative ring associated with a symmetric design]. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 57(1):72–??, March 2014. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Khemphet:2014:SPD

[4088] Anchalee Khemphet and Justin R. Peters. Semicrossed products of the disk algebra and the Jacobson radical. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 57(1):80–??, March 2014. CO-DEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Lazar:2014:CSG

[4089] Aldo J. Lazar. Compact subsets of the Glimm space of a C^{∗}-algebra. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 57(1):90–??, March 2014. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Levy:2014:RJG

[4090] Jason Levy. Rationality and the Jordan–Gatti–Viniberghi decomposition. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 57(1):97–??, March 2014. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Luca:2014:CFE

[4091] Florian Luca and Igor E. Shparlinski. On the counting function of elliptic Carmichael numbers. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 57(1):105–??, March 2014. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Madras:2014:LBE

[4092] Neal Madras. A lower bound for the end-to-end distance of self-avoiding walk. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 57(1):113–??, March 2014. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Mildenberger:2014:SFC

[4093] Heike Mildenberger, Dilip Raghavan, and Juris Steprans. Splitting families and complete separability. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 57(1):119–??, March 2014. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Mlaiki:2014:CT

[4094] Nabil M. Mlaiki. Camina triples. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 57(1):125–??, March 2014. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Mubeena:2014:TCC

[4095] T. Mubeena and P. Sankaran. Twisted conjugacy classes in Abelian extensions of certain linear groups. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 57(1):132–??, March 2014. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Mukwembi:2014:SOC

[4096] Simon Mukwembi. Size, order, and connected domination. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 57(1):141–??, March 2014. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Mustafayev:2014:ESE

[4097] H. S. Mustafayev. The essential spectrum of the essentially isometric operator. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 57(1):145–??, March 2014. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Oral:2014:SDM

[4098] Kürsat Hakan Oral, Neslihan Aysen Özkirisci, and Ünsal Tekir. Strongly 0-dimensional modules. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 57(1):159–??, March 2014. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Oztop:2014:MMO

[4099] Serap Öztop and Nico Spronk. On minimal and maximal p-operator space structures. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 57(1):166–??, March 2014. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Rabier:2014:QDT

[4100] Patrick J. Rabier. Quasiconvexity and density topology. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 57(1):178– ??, March 2014. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Rad:2014:CBZ

[4101] Nader Jafari Rad and Sayyed Heidar Jafari. A characterization of bipartite zero-divisor graphs. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 57(1):188–??, March 2014. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Zhao:2014:LBL

[4102] Wei Zhao. A lower bound for the length of closed geodesics on a Finsler manifold. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 57(1):194–??, March 2014. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic). See erratum [4103].

Zhao:2014:EPL

[4103] Wei Zhao. Erratum to the paper “A Lower Bound for the Length of Closed Geodesics on a Finsler Manifold”. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 57(1):209–??, March 2014. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic). See

Zhang:2014:EFG

[4104] Jiao Zhang and Qing-Wen Wang. An explicit formula for the generalized cyclic shuffle map. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 57(1):210–??, March 2014. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Adamaszek:2014:SFC

[4105] Michal Adamaszek. Small flag complexes with torsion. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 57(2):225–??, June 2014. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Bagherian:2014:MCT

[4106] J. Bagherian. On the multiplicities of characters in table algebras. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 57 (2):231–??, June 2014. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Bernardes:2014:ALS

[4107] Nilson C. Bernardes, Jr. Addendum to “Limit Sets of Typical Homeomorphisms”. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 57(2):240–??, June 2014. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Brodskiy:2014:AND

[4108] N. Brodskiy, J. Dydak, and U. Lang. Assouad–Nagata dimension of wreath products of groups. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 57(2):245–??, June 2014. CODEN CM-BUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Christensen:2014:PWF

[4109] Ole Christensen, Hong Oh Kim, and Rae Young Kim. On Parseval wavelet frames with two or three generators via the unitary extension principle. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 57(2):254–??, June 2014. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Dai:2014:SHA

[4110] Li Dai and Jingcheng Dong. On semisimple Hopf algebras of dimension pq^{n}. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 57(2):264–??, June 2014. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Didas:2014:DTA

[4111] Michael Didas and Jörg Eschmeier. Derivations on Toeplitz algebras. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 57(2):270–??, June 2014. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Elkholy:2014:MPP

[4112] A. M. Elkholy and M. H. Abd El-Latif. On mutually m-permutable product of smooth groups. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 57(2):277–??, June 2014. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Fulp:2014:IDD

[4113] Ronald Fulp. Infinite dimensional DeWitt supergroups and their bodies. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 57(2):283–??, June 2014. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic). See correction [?].

Ghasemi:2014:CCS

[4114] Mehdi Ghasemi, Murray Marshall, and Sven Wagner. Closure of the cone of sums of 2d-powers in certain weighted ℓ_{1}-seminorm topologies. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 57(2):289–??, June 2014. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Gille:2014:OAR

[4115] Philippe Gille. Octonion algebras over rings are not determined by their norms. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 57(2):303–??, June 2014. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Hakamata:2014:LOF

[4116] Ryoto Hakamata and Masakazu Teragaito. Left-orderable fundamental group and Dehn surgery on the knot 5_{2}. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 57(2):310–??, June 2014. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Huang:2014:DPP

[4117] Zhaoyong Huang. Duality of preenvelopes and pure injective modules. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 57(2):318–??, June 2014. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Ivanov:2014:ZDG

[4118] S. V. Ivanov and Roman Mikhailov. On zero-divisors in group rings of groups with torsion. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 57(2):326–??, June 2014. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Karassev:2014:AMH

[4119] A. Karassev, V. Todorov, and V. Valov. Alexandroff manifolds and homogeneous continua. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 57(2):335–??, June 2014. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Kishimoto:2014:LUG

[4120] Daisuke Kishimoto, Akira Kono, and Mitsunobu Tsutaya. On localized unstable K^{1}-groups and applications to self-homotopy groups. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 57(2): 344–??, June 2014. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Lauret:2014:REB

[4121] Emilio A. Lauret. Representation equivalent Bieberbach groups and strongly isospectral flat manifolds. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 57(2):357–??, June 2014. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Li:2014:HLF

[4122] Lei Li and Ya-Shu Wang. How Lipschitz functions characterize the underlying metric spaces. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 57(2):364–??, June 2014. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Lopez:2014:PEM

[4123] S. C. López, Muntaner-Batle, and Rius-Font. A problem on edge-magic labelings of cycles. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 57(2):375–??, June 2014. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Lydka:2014:CEF

[4124] Adrian Lydka. On complex explicit formulae connected with the Möbius function of an elliptic curve. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 57(2):381–??, June 2014. CODEN CM-BUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Morita:2014:SST

[4125] Jun Morita and Bertrand Rémy. Simplicity of some twin tree automorphism groups with trivial commutation relations. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 57(2):390–??, June 2014. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Perrone:2014:CCS

[4126] Domenico Perrone. Curvature of K-contact semi-Riemannian manifolds. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 57(2):401–??, June 2014. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Samei:2014:CGC

[4127] Karim Samei. On the comaximal graph of a commutative ring. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 57(2): 413–??, June 2014. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Soltan:2014:NAL

[4128] Piotr M. Soltan and Ami Viselter. A note on amenability of locally compact quantum groups. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 57(2):424–??, June 2014. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Tagami:2014:RIF

[4129] Keiji Tagami. The Rasmussen invariant, four-genus and three-genus of an almost positive knot are equal. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 57(2):431–??, June 2014. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Yang:2014:FPL

[4130] YanHong Yang. The fixed point locus of the Verschiebung on M_{X} (∈, ′) for genus-2 curves X in characteristic 2. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 57(2):439–??, June 2014. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Alaghmandan:2014:ZAC

[4131] Mahmood Alaghmandan, Yemon Choi, and Ebrahim Samei. ZLamenability constants of finite groups with two character degrees. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 57 (3):449–??, September 2014. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Bownik:2014:CPC

[4132] Marcin Bownik and John Jasper. Constructive proof of Carpenter’s theorem. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 57(3):463–??, September 2014. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Eghbali:2014:STC

[4133] Majid Eghbali. On set theoretically and cohomologically complete intersection ideals. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 57(3):477–??, September 2014. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Franc:2014:FCV

[4134] Cameron Franc and Geoffrey Mason. Fourier coefficients of vector-valued modular forms of dimension 2. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 57(3):485–??, September 2014. CO-DEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Fujita:2014:JCC

[4135] Yasutsugu Fujita and Takafumi Miyazaki. Jeśmanowicz’ Conjecture with congruence relations. II. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 57(3):495–??, September 2014. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Galindo:2014:BRU

[4136] César Galindo. On braided and ribbon unitary fusion categories. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 57(3): 506–??, September 2014. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Goncalves:2014:SPS

[4137] Daniel Gonçalves. Simplicity of partial skew group rings of Abelian groups. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 57(3):511–??, September 2014. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Guo:2014:MIT

[4138] Guangquan Guo and Guoping Wang. Maximizing the index of trees with given domination number. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 57(3):520–??, September 2014. CO-DEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Heil:2014:MTK

[4139] Wolfgang Heil and Dongxu Wang. On 3-manifolds with torus or Klein bottle category two. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 57(3):526–??, September 2014. CODEN CM-BUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Ide:2014:IFS

[4140] Joshua Ide and Lenny Jones. Infinite families of A_{4}-sextic polynomials. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 57(3):538–??, September 2014. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Kalantar:2014:COR

[4141] Mehrdad Kalantar. Compact operators in regular LCQ groups. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 57(3): 546–??, September 2014. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Kane:2014:AIL

[4142] Daniel M. Kane and Scott Duke Kominers. Asymptotic improvements of lower bounds for the least common multiples of arithmetic progressions. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 57(3):551–??, September 2014. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Kaveh:2014:NGG

[4143] Kiumars Kaveh and A. G. Khovanskii. Note on the Grothendieck group of subspaces of rational functions and Shokurov’s Cartier b-divisors. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 57 (3):562–??, September 2014. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Kiani:2014:SRD

[4144] Sima Kiani, Hamid Reza Maimani, and Reza Nikandish. Some results on the domination number of a zero-divisor graph. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 57(3):573–??, September 2014. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Larson:2014:HPL

[4145] Paul Larson and Franklin D. Tall. On the hereditary paracompactness of locally compact, hereditarily normal spaces. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 57(3):579–??, September 2014. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Lehec:2014:SPP

[4146] Joseph Lehec. Short probabilistic proof of the Brascamp-Lieb and barthe theorems. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 57(3):585–??, September 2014. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Lu:2014:IMS

[4147] Yufeng Lu, Dachun Yang, and Wen Yuan. Interpolation of Morrey spaces on metric measure spaces. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 57(3):598–??, September 2014. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Nasr-Isfahani:2014:JRS

[4148] Alireza Nasr-Isfahani. Jacobson radicals of skew polynomial rings of derivation type. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 57(3):609–??, September 2014. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Parusinski:2014:NWP

[4149] Adam Parusiński and Jean-Philippe Rolin. A note on the Weierstrass preparation theorem in quasianalytic local rings. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 57(3):614–??, September 2014. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Petrich:2014:CFC

[4150] Mario Petrich. Combinatorially factorizable cryptic inverse semigroups. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 57(3):621–??, September 2014. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Sokic:2014:ICA

[4151] Miodrag Sokić. Indicators, chains, antichains, Ramsey property. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 57(3):

631–??, September 2014. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Swanepoel:2014:ESS

[4152] Konrad J. Swanepoel. Equilateral sets and a Schütte theorem for the 4-norm. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 57(3):640–??, September 2014. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Tang:2014:HFG

[4153] Juping Tang and Long Miao. On the F⊕-hypercentre of finite groups. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 57(3):648–??, September 2014. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Thang:2014:ALS

[4154] Nguyen Tat Thang. Admissibility of local systems for some classes of line arrangements. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 57(3):658–??, September 2014. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Ahmadi:2014:CLG

[4155] S. Ruhallah Ahmadi and Bruce Gilligan. Complexifying Lie group actions on homogeneous manifolds of non-compact dimension two. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 57 (4):673–??, December 2014. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Aurichi:2014:TGA

[4156] Leandro F. Aurichi and Rodrigo R. Dias. Topological games and Alster spaces. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 57(4):683–??, December 2014. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Bailet:2014:MMF

[4157] Pauline Bailet. On the monodromy of Milnor fibers of hyperplane arrangements. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 57(4):697–??, December 2014. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Brannan:2014:SAF

[4158] Michael Brannan. Strong asymptotic freeness for free orthogonal quantum groups. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 57(4):708–??, December 2014. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Bruillard:2014:CIM

[4159] Paul Bruillard, César Galindo, Seung-Moon Hong, Yevgenia Kashina, Deepak Naidu, Sonia Natale, Julia Yael Plavnik, and Eric C. Rowell. Classification of integral modular categories of Frobenius–Perron dimension pq^{4} and p^{2}q^{2}. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 57(4):721–??, December 2014. CODEN CM-BUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Cagliero:2014:TPE

[4160] Leandro Cagliero and Fernando Szechtman. On the theorem of the primitive element with applications to the representation theory of associative and Lie algebras. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 57(4):735–??, December 2014. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Cavalieri:2014:GPP

[4161] Renzo Cavalieri and Steffen Marcus. Geometric perspective on piecewise polynomiality of double Hurwitz numbers. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 57(4):749–??, December 2014. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

daSilva:2014:HMS

[4162] Rosângela Maria da Silva and Keti Tenenblat. Helicoidal minimal surfaces in a Finsler space of Randers type. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 57(4):765–??, December 2014. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Erzakova:2014:MNR

[4163] Nina A. Erzakova. Measures of noncompactness in regular spaces. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 57 (4):780–??, December 2014. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Fang:2014:NCW

[4164] Zhong-Shan Fang and Ze-Hua Zhou. New characterizations of the weighted composition operators between Bloch type spaces in the polydisk. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 57(4):794–??, December 2014. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Gabriyelyan:2014:FLC

[4165] S. S. Gabriyelyan. Free locally convex spaces and the k-space property. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 57(4):803–??, December 2014. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Godefroy:2014:UPS

[4166] G. Godefroy. Uniqueness of preduals in spaces of operators. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 57(4): 810–??, December 2014. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Hou:2014:GDT

[4167] Ruchen Hou. On global dimensions of tree type finite dimensional algebras. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 57(4):814–??, December 2014. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Jeong:2014:RHC

[4168] Imsoon Jeong, Seonhui Kim, and Young Jin Suh. Real hypersurfaces in complex two-plane Grassmannians with Reeb parallel structure Jacobi operator. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 57(4):821–??, December 2014. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Koh:2014:ROA

[4169] Doowon Koh. Restriction operators acting on radial functions on vector spaces over finite fields. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 57(4):834–??, December 2014. CODEN CM-BUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Lei:2014:FTV

[4170] Antonio Lei. Factorisation of two-variable p-adic L-functions. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 57(4): 845–??, December 2014. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Pan:2014:BEL

[4171] Qingfei Pan and Kun Wang. On the bound of the C^{∗} exponential length. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 57(4):853–??, December 2014. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Parlier:2014:SNS

[4172] Hugo Parlier. A short note on short pants. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 57(4):870–??, December 2014. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Schoen:2014:CCS

[4173] Tomasz Schoen. On convolutions of convex sets and related problems. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 57(4):877–??, December 2014. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Xu:2014:ESN

[4174] Yong Xu and Xinjian Zhang. m-embedded subgroups and p-nilpotency of finite groups. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 57(4):884–??, December 2014. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Alaghmandan:2015:AAS

[4175] Mahmood Alaghmandan. Approximate amenability of Segal algebras II. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 58(1):3–??, March 2015. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Boulabiar:2015:C

[4176] Karim Boulabiar. Characters on C(X). Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 58(1):7–??, March 2015. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Chavan:2015:ITB

[4177] Sameer Chavan. Irreducible tuples without the boundary property. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 58 (1):9–??, March 2015. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Chen:2015:CCR

[4178] Jiecheng Chen and Guoen Hu. Compact commutators of rough singular integral operators. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 58(1):19–??, March 2015. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Chung:2015:EFI

[4179] Jaeyoung Chung. On an exponential functional inequality and its distributional version. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 58(1):30–??, March 2015. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Daniilidis:2015:OGD

[4180] A. Daniilidis, D. Drusvyatskiy, and A. S. Lewis. Orbits of geometric descent. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 58(1):44–??, March 2015. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

DeNitties:2015:SFD

[4181] Giuseppe De Nitties and Hermann Schulz-Baldes. Spectral flows of dilations of Fredholm operators. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 58(1):51–??, March 2015. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Fulp:2015:CID

[4182] Ronald Owen Fulp. Correction to “Infinite Dimensional DeWitt Supergroups and Their Bodies”. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 58(1):69–??, March 2015. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic). See [4113].

Ghenciu:2015:LSB

[4183] Ioana Ghenciu. Limited sets and bibasic sequences. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 58(1):71–??, March 2015. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Harada:2015:ECR

[4184] Megumi Harada, Tatsuya Horiguchi, and Mikiya Masuda. The equivariant cohomology rings of Peterson varieties in all Lie types. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 58(1): 80–??, March 2015. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Hasegawa:2015:ECS

[4185] Kei Hasegawa. Essential commutants of semicrossed products. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 58(1): 91–??, March 2015. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Hossein-Zadeh:2015:GAC

[4186] Samaneh Hossein-Zadeh, Ali Iranmanesh, Mohammad Ali Hosseinzadeh, and Mark L. Lewis. On graphs associated with character degrees and conjugacy class sizes of direct products of finite groups. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 58(1):105– ??, March 2015. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Kamalov:2015:PAT

[4187] F. Kamalov. Property T and amenable transformation group C^{∗}- algebras. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 58(1):110–??, March 2015. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Mantilla-Soler:2015:WAE

[4188] Guillermo Mantilla-Soler. Weak arithmetic equivalence. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 58(1):115– ??, March 2015. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Markovic:2015:SCB

[4189] Marijan Marković. A sharp constant for the Bergman projection. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 58(1): 128–??, March 2015. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Nasseh:2015:GAR

[4190] Saeed Nasseh. On the generalized Auslander–Reiten conjecture under certain ring extensions. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 58(1):134–??, March 2015. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Olevskii:2015:LC

[4191] Victor Olevskii. Localization and completeness in L_{2}(R). Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 58(1):144– ??, March 2015. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Ostrovskii:2015:CBM

[4192] Mikhail I. Ostrovskii. Connections between metric characterizations of superreflexivity and the Radon–Nikodýn property for dual Banach spaces. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 58(1):150–??, March 2015. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Ozgur:2015:CCI

[4193] Cihan Özgür and Adela Mihai. Corrigendum to “Chen Inequalities for Submanifolds of Real Space Forms with a Semi-symmetric Non-metric Connection”. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 58(1):158–??, March 2015. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic). See [3962].

Pollack:2015:SNN

[4194] Paul Pollack and Joseph Vandehey. Some normal numbers generated by arithmetic functions. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 58(1):160–??, March 2015. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Raffoul:2015:PSA

[4195] Youssef N. Raffoul. Periodic solutions of almost linear Volterra integrodynamic equation on periodic time scales. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 58(1):174–??, March 2015. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Tarnauceanu:2015:FGD

[4196] Marius Tarnauceanu. On finite groups with dismantlable subgroup lattices. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 58(1):182–??, March 2015. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Wirths:2015:TES

[4197] Karl Joachim Wirths. Telescoping estimates for smooth series. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 58(1): 188–??, March 2015. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Yang:2015:DGO

[4198] Qingjie Yang and Weiting Zhong. Dihedral groups of order 2p of automorphisms of compact Riemann surfaces of genus p − 1. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 58(1):196– ??, March 2015. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Moslehian:2015:EAO

[4199] Mohammad Sal Moslehian and Ali Zamani. Exact and approximate operator parallelism. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 58(1):207–??, March 2015. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Aghigh:2015:CDP

[4200] Kamal Aghigh and Azadeh Nikseresht. Characterizing distinguished pairs by using liftings of irreducible polynomials. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 58(2):225–??, June 2015. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Bergen:2015:AAP

[4201] Jeffrey Bergen. Affine actions of U_{q}(sl(2)) on polynomial rings. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 58(2): 233–??, June 2015. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Botelho:2015:IHO

[4202] Fernanda Botelho. Isometries and Hermitian operators on Zygmund spaces. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 58(2):241–??, June 2015. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Cartwright:2015:LDG

[4203] Dustin Cartwright, David Jensen, and Sam Payne. Lifting divisors on a generic chain of loops. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 58(2):250–??, June 2015. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

DeFilippis:2015:GJS

[4204] Vincenzo De Filippis, Abdellah Mamouni, and Lahcen Oukhtite. Generalized Jordan semiderivations in prime rings. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 58(2):263–??, June 2015. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Jafari:2015:DZD

[4205] Sayyed Heidar Jafari and Nader Jafari Rad. On domination of zerodivisor graphs of matrix rings. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 58(2):271–??, June 2015. CODEN CM-BUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Johnson:2015:ITO

[4206] William Johnson, Amir Bahman Nasseri, Gideon Schechtman, and Tomasz Tkocz. Injective Tauberian operators on L_{1} and operators with dense range on ℓ_{∞}. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 58(2):276–??, June 2015. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Kalus:2015:RRC

[4207] Matthias Kalus. On the relation of real and complex Lie supergroups. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 58(2):281–??, June 2015. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Karpukhin:2015:SPF

[4208] Mikhail Karpukhin. Spectral properties of a family of minimal tori of revolution in five-dimensional sphere. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 58(2):285–??, June 2015. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Khamsi:2015:AFP

[4209] M. A. Khamsi. Approximate fixed point sequences of nonlinear semigroup in metric spaces. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 58(2):297–??, June 2015. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Khoshkhah:2015:LDM

[4210] Kaveh Khoshkhah and Manouchehr Zaker. On the largest dynamic monopolies of graphs with a given average threshold. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 58(2):306–??, June 2015. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Lewkowicz:2015:CFU

[4211] Marek Kazimierz Lewkowicz. Coloring four-uniform hypergraphs on nine vertices. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 58(2):317–??, June 2015. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Llamas:2015:CPB

[4212] Aurora Llamas and José Martínez-Bernal. Cover product and Betti polynomial of graphs. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 58(2):320–??, June 2015. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Medini:2015:CDH

[4213] Andrea Medini. Countable dense homogeneity in powers of zerodimensional definable spaces. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 58(2):334–??, June 2015. CODEN CM-BUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Merino-Cruz:2015:CIC

[4214] Héctor Merino-Cruz and Antoni Wawrzynczyk. On closed ideals in a certain class of algebras of holomorphic functions. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 58(2):350–??, June 2015. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Sebag:2015:HPG

[4215] Julien Sebag. Homological planes in the Grothendieck ring of varieties. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 58(2):356–??, June 2015. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Sharma:2015:FSL

[4216] R. K. Sharma and Swati Sidana. Finite semisimple loop algebras of indecomposable RA loops. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 58(2):363–??, June 2015. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Szabo:2015:SNC

[4217] Gábor Szabó. A short note on the continuous Rokhlin property and the universal coefficient theorem in E-theory. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 58(2):374–??, June 2015. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Tang:2015:SLB

[4218] Xiaomin Tang and Taishun Liu. The Schwarz lemma at the boundary of the egg domain B_{p_{1},p_{2}} in C^{n}. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 58(2):381–??, June 2015. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Tang:2015:SDC

[4219] Zhongming Tang. On Stanley depths of certain monomial factor algebras. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 58(2):393–??, June 2015. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Tikuisis:2015:SCS

[4220] Aaron Peter Tikuisis and Andrew Toms. On the structure of Cuntz semigroups in (possibly) nonunital C∗-algebras. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 58(2):402–??, June 2015. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Willson:2015:FPT

[4221] Benjamin Willson. A fixed point theorem and the existence of a Haar measure for hypergroups satisfying conditions related to amenability. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 58(2):415–??, June 2015. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Yamagishi:2015:RCP

[4222] Masakazu Yamagishi. Resultants of Chebyshev polynomials: The first, second, third, and fourth kinds. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 58(2):423–??, June 2015. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Yang:2015:SOR

[4223] Dachun Yang and Sibei Yang. Second-order Riesz transforms and maximal inequalities associated with magnetic Schrödinger operators. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 58 (2):432–??, June 2015. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Boynton:2015:GDI

[4224] Jason Greene Boynton and Jim Coykendall. On the graph of divisibility of an integral domain. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 58(3):449–??, September 2015. CODEN CM-BUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Casini:2015:HSC

[4225] Emanuele Casini, Enrico Miglierina, and Lukasz Piasecki. Hyperplanes in the space of convergent sequences and preduals of ℓ_{1}. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 58(3):459–??, September 2015. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Demirbas:2015:ASG

[4226] Seckin Demirbas. Almost sure global well-posedness for the fractional cubic Schrödinger equation on torus. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 58(3):471–??, September 2015. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Duc:2015:IPD

[4227] Dinh Thanh Duc, Nguyen Du Vi Nhan, and Nguyen Tong Xuan. Inequalities for partial derivatives and their applications. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 58(3):486–??, September 2015. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Edmunds:2015:CDM

[4228] Charles C. Edmunds. Constructing double magma on groups using commutation operations. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 58(3):497–??, September 2015. CODEN CM-BUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Hsu:2015:VSA

[4229] Ming-Hsiu Hsu and Ming-Yi Lee. VMO space associated with parabolic sections and its application. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 58(3):507–??, September 2015. CO-DEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Kang:2015:RMD

[4230] Su-Jeong Kang. Refined Motivic dimension. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 58(3):519–??, September 2015. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Li:2015:RCT

[4231] Benling Li and Zhongmin Shen. Ricci curvature tensor and non-Riemannian quantities. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 58(3):530–??, September 2015. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Li:2015:MNS

[4232] Lili Li and Guiyun Chen. Minimal non self dual groups. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 58(3):538–??, September 2015. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Lu:2015:HMF

[4233] Guangshi Lü and Ayyadurai Sankaranarayanan. Higher moments of Fourier coefficients of cusp forms. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 58(3):548–??, September 2015. CO-DEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Martinez-Maure:2015:PLF

[4234] Yves Martinez-Maure. Plane Lorentzian and Fuchsian hedgehogs. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 58 (3):561–??, September 2015. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Martinez-Torres:2015:DTC

[4235] David Martinez-Torres. The diffeomorphism type of canonical integrations of Poisson tensors on surfaces. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 58(3):575–??, September 2015. CO-DEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Matringe:2015:SBF

[4236] Nadir Matringe. A specialisation of the Bump–Friedberg L-function. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 58(3):580–??, September 2015. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Ongaro:2015:NPS

[4237] Jared Ongaro and Boris Shapiro. A note on planarity stratification of Hurwitz spaces. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 58(3):596–??, September 2015. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Rodger:2015:PDK

[4238] C. A. Rodger and Thomas Richard Whitt. Path decompositions of Kneser and generalized Kneser graphs. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 58(3):610–??, September 2015. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Sands:2015:FQC

[4239] Jonathan W. Sands. L-functions for quadratic characters and annihilation of Motivić cohomology groups. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 58(3):620–??, September 2015. CO-DEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Silberman:2015:QUE

[4240] Lior Silberman. Quantum unique ergodicity on locally symmetric spaces: the degenerate lift. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 58(3):632–??, September 2015. CODEN CM-BUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Tang:2015:GSS

[4241] Xianhua Tang. Ground state solutions of Nehari–Pankov type for a superlinear Hamiltonian elliptic system on R^{N}. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 58(3):651–??, September 2015. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Vahidi:2015:BNF

[4242] Alireza Vahidi. Betti numbers and flat dimensions of local cohomology modules. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 58(3):664–??, September 2015. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Achter:2015:LHE

[4243] Jeffrey Achter and Cassandra Williams. Local heuristics and an exact formula for Abelian surfaces over finite fields. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 58(4):673–??, December 2015. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Anona:2015:ALA

[4244] F. M. Anona, Princy Randriambololondrantomalala, and H. S. G. Ravelonirina. Sur les algèbres de Lie associées à une connexion. (French) [On Lie algebras associated with a connection]. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 58(4):692–??, December 2015. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Benamar:2015:CFE

[4245] H. Benamar, A. Chandoul, and M. Mkaouar. On the continued fraction expansion of fixed period in finite fields. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 58(4):704–??, December 2015. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Brendle:2015:MCC

[4246] Simon Brendle and Otis Chodosh. On the maximum curvature of closed curves in negatively curved manifolds. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 58(4):713–??, December 2015. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Castro:2015:IMR

[4247] Alfonso Castro and Emily Fischer. Infinitely many rotationally symmetric solutions to a class of semilinear Laplace–Beltrami equations on spheres. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 58(4):723–??, December 2015. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Efrat:2015:VMP

[4248] Ido Efrat and Eliyahu Matzri. Vanishing of Massey products and Brauer groups. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 58(4):730–??, December 2015. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Gao:2015:HPR

[4249] Zenghui Gao. Homological properties relative to injectively resolving subcategories. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 58(4):741–??, December 2015. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Han:2015:ETI

[4250] Yanchang Han. Embedding theorem for inhomogeneous Besov and Triebel–Lizorkin spaces on RD-spaces. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 58(4):757–??, December 2015. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Hanson:2015:CSB

[4251] Brandon Hanson. Character sums over Bohr sets. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 58(4):774–??, December 2015. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Kitabeppu:2015:NBR

[4252] Yu Kitabeppu and Sajjad Lakzian. Non-branching RCD(0, N) geodesic spaces with small linear diameter growth have finitely generated fundamental groups. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 58(4):787–??, December 2015. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Kong:2015:SQE

[4253] Qingjun Kong and Xiuyun Guo. On s-semipermutable or squasinormally embedded subgroups of finite groups. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 58(4):799–??, December 2015. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Liu:2015:RMM

[4254] Feng Liu and Huoxiong Wu. On the regularity of the multisublinear maximal functions. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 58(4):808–??, December 2015. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Llibre:2015:LCL

[4255] Jaume Llibre and Xiang Zhang. On the limit cycles of linear differential systems with homogeneous nonlinearities. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 58(4):818–??, December 2015. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Luo:2015:EMC

[4256] Xiu-Hua Luo. Exact morphism category and Gorenstein-projective representations. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 58(4):824–??, December 2015. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Perez:2015:RHC

[4257] Juan de Dios Pérez, Young Jin Suh, and Changhwa Woo. Real hypersurfaces in complex two-plane Grassmannians with GTW harmonic curvature. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 58(4):835–??, December 2015. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Sundar:2015:CCT

[4258] S. Sundar. A computation with the Connes–Thom isomorphism. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 58(4): 846–??, December 2015. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Williams:2015:TQF

[4259] Kenneth S. Williams. Ternary quadratic forms and eta quotients. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 58 (4):858–??, December 2015. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Wright:2015:VKC

[4260] Thomas Wright. Variants of Korselt’s criterion. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 58(4):869–??, December 2015. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Zaatra:2015:GSS

[4261] Mohamed Zaatra. Generating some symmetric semi-classical orthogonal polynomials. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 58(4):877–??, December 2015. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Alfuraidan:2016:CPM

[4262] Monther Rashed Alfuraidan. The contraction principle for multivalued mappings on a modular metric space with a graph. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 59(1):3–??, March 2016. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Aulaskari:2016:CHR

[4263] Rauno Aulaskari and Huaihui Chen. On classes Q_{p}^{#} for hyperbolic Riemann surfaces. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 59(1):13–??, March 2016. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Cleanthous:2016:GES

[4264] Galatia Cleanthous. A geometric extension of Schwarz’s lemma and applications. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 59(1):30–??, March 2016. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Donovan:2016:DAD

[4265] Diane M. Donovan, Terry S. Griggs, Thomas A. McCourt, Jakub Oprsal, and David Stanovský. Distributive and anti-distributive Mendelsohn triple systems. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 59(1):36–??, March 2016. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Dorfmeister:2016:BPT

[4266] Josef F. Dorfmeister, Jun ichi Inoguchi, and Shimpei Kobayashi. On the Bernstein problem in the three-dimensional Heisenberg group. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 59(1): 50–??, March 2016. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Feng:2016:UPW

[4267] Han Feng. Uncertainty principles on weighted spheres, balls and simplexes. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 59(1):62–??, March 2016. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Gasinski:2016:PSG

[4268] Leszek Gasiński and Nikolaos S. Papageorgiou. Positive solutions for the generalized nonlinear logistic equations. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 59(1):73–??, March 2016. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Gauthier:2016:AFD

[4269] Paul M. Gauthier and Julie Kienzle. Approximation of a function and its derivatives by entire functions. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 59(1):87–??, March 2016. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Goncalves:2016:FRG

[4270] Daniel Gonçalves, Hui Li, and Danilo Royer. Faithful representations of graph algebras via branching systems. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 59(1):95–??, March 2016. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

He:2016:LPC

[4271] Ziyi He, Dachun Yang, and Wen Yuan. Littlewood–Paley characterizations of second-order Sobolev spaces via averages on balls. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 59(1): 104–??, March 2016. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Hu:2016:SPS

[4272] Pei-Chu Hu and Bao Qin Li. A simple proof and strengthening of a uniqueness theorem for L-functions. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 59(1):119–??, March 2016. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Jensen:2016:DST

[4273] Gerd Jensen and Christian Pommerenke. Discrete space-time and Lorentz transformations. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 59(1):123–??, March 2016. CODEN CM-BUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Kajihara:2016:TFB

[4274] Yasushi Kajihara. Transformation formulas for bilinear sums of basic hypergeometric series. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 59(1):136–??, March 2016. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Laterveer:2016:BNC

[4275] Robert Laterveer. A brief note concerning hard Lefschetz for Chow groups. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 59(1):144–??, March 2016. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

MacColl:2016:RKH

[4276] Joseph MacColl. Rotors in Khovanov homology. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 59(1):159–??, March 2016. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Martinez-Pedroza:2016:NFG

[4277] Eduardo Martínez-Pedroza. A note on fine graphs and homological isoperimetric inequalities. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 59(1):170–??, March 2016. CODEN CM-BUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Naylor:2016:GTK

[4278] Geoff Naylor and Dale Rolfsen. Generalized torsion in knot groups. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 59 (1):182–??, March 2016. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Raeisi:2016:RNW

[4279] Ghaffar Raeisi and Ali Zaghian. Ramsey number of wheels versus cycles and trees. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 59(1):190–??, March 2016. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Rajaee:2016:QCS

[4280] Saeed Rajaee. Quasi-copure submodules. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 59(1):197–??, March 2016. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Spektor:2016:RKI

[4281] Susanna Spektor. Restricted Khinchine inequality. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 59(1):204–??, March 2016. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Totik:2016:UUS

[4282] Vilmos Totik. Universality under Szegő’s condition. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 59(1):211–??, March 2016. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Atici:2016:CFD

[4283] Ferhan M. Atici and Hatice Yaldiz. Convex functions on discrete time domains. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 59(2):225–??, June 2016. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Beardon:2016:NDF

[4284] Alan F. Beardon. Non-discrete Frieze groups. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 59(2):234–??, June 2016. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Cao:2016:NQH

[4285] Wensheng Cao and Xiaolin Huang. A note on quaternionic hyperbolic ideal triangle groups. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 59(2):244–??, June 2016. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

DeFilippis:2016:APV

[4286] Vincenzo De Filippis. Annihilators and power values of generalized skew derivations on Lie ideals. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 59(2):258–??, June 2016. CODEN CM-BUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Dehghani-Zadeh:2016:ACG

[4287] Fatemeh Dehghani-Zadeh. Artinianness of composed graded local cohomology modules. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 59(2):271–??, June 2016. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Dimca:2016:PDP

[4288] Alexandru Dimca. The Poincaré–Deligne polynomial of Milnor fibers of triple point line arrangements is combinatorially determined. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 59(2): 279–??, June 2016. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Dukes:2016:ETI

[4289] Peter Dukes, E. R. Lamken, and Alan C. H. Ling. An existence theory for incomplete designs. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 59(2):287–??, June 2016. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Gauthier:2016:NEZ

[4290] P. M. Gauthier. Non-extendable zero sets of harmonic and holomorphic functions. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 59(2):303–??, June 2016. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Ilten:2016:PRD

[4291] Nathan Ilten and Zach Teitler. Product ranks of the 3 × 3 determinant and permanent. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 59(2):311–??, June 2016. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Ino:2016:PNS

[4292] Shoji Ino. Perturbations of von Neumann subalgebras with finite index. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 59(2):320–??, June 2016. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Jiang:2016:UJC

[4293] Chunlan Jiang and Rui Shi. On the uniqueness of Jordan canonical form decompositions of operators by K-theoretical data. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 59(2):326–??, June 2016. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Kepczyk:2016:NAS

[4294] Marek Kepczyk. A note on algebras that are sums of two subalgebras. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 59(2):340–??, June 2016. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Krantz:2016:TBA

[4295] Steven Krantz. On a theorem of Bers, with applications to the study of automorphism groups of domains. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 59(2):346–??, June 2016. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Li:2016:FQO

[4296] Chi-Kwong Li and Ming-Cheng Tsai. Factoring a quadratic operator as a product of two positive contractions. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 59(2):354–??, June 2016. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Li:2016:DAS

[4297] Dan Li and Wanbiao Ma. Dynamical analysis of a stage-structured model for Lyme disease with two delays. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 59(2):363–??, June 2016. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Moameni:2016:SED

[4298] Abbas Moameni. Supports of extremal doubly stochastic measures. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 59 (2):381–??, June 2016. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Prajapati:2016:TCG

[4299] S. K. Prajapati and R. Sarma. Total character of a group G with (G, Z(G)) as a generalized Camina pair. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 59(2):392–??, June 2016. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Zargar:2016:FGF

[4300] Majid Rahro Zargar and Hossein Zakeri. On flat and Gorenstein flat dimensions of local cohomology modules. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 59(2):403–??, June 2016. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Song:2016:EMS

[4301] Hongxue Song, Caisheng Chen, and Qinglun Yan. Existence of multiple solutions for a p-Laplacian system in R^{N} with sign-changing weight functions. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 59(2):417–??, June 2016. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Yao:2016:ESF

[4302] Hongliang Yao. On extensions of stably finite C∗-algebras (II). Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 59(2): 435–??, June 2016. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Zhang:2016:NCP

[4303] Haihui Zhang. A note on 3-choosability of planar graphs related to Montassier’s conjecture. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 59(2):440–??, June 2016. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Abdallah:2016:HTS

[4304] Nancy Abdallah. On Hodge theory of singular plane curves. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 59(3): 449–??, September 2016. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Ara:2016:NRE

[4305] Pere Ara and Kevin C. O’Meara. The nilpotent regular element problem. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 59(3):461–??, September 2016. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Clay:2016:TBO

[4306] Adam Clay, Colin Desmarais, and Patrick Naylor. Testing bi-orderability of knot groups. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 59(3):472–??, September 2016. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Crooks:2016:GEC

[4307] Peter Crooks and Tyler Holden. Generalized equivariant cohomology and stratifications. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 59(3):483–??, September 2016. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

DeCarli:2016:OPG

[4308] Laura De Carli and Gohin Shaikh Samad. One-parameter groups of operators and discrete Hilbert transforms. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 59(3):497–??, September 2016. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

DeNicola:2016:GGF

[4309] Antonio De Nicola and Ivan Yudin. Generalized Goldberg formula. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 59 (3):508–??, September 2016. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Hare:2016:RBK

[4310] Kathryn Hare and L. Thomas Ramsey. The relationship between ϵkronecker sets and Sidon sets. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 59(3):521–??, September 2016. CO-DEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Jahan:2016:CLP

[4311] Qaiser Jahan. Characterization of low-pass filters on local fields of positive characteristic. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 59(3):528–??, September 2016. CODEN CM-BUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Jiang:2016:SHS

[4312] Yongxin Jiang, Wei Wang, and Zhaosheng Feng. Spatial homogenization of stochastic wave equation with large interaction. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 59(3):542–??, September 2016. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Kachmar:2016:NFE

[4313] Ayman Kachmar. A new formula for the energy of bulk superconductivity. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 59(3):553–??, September 2016. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Li:2016:NER

[4314] Boyu Li. Normal extensions of representations of Abelian semigroups. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 59(3):564–??, September 2016. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Li:2016:COR

[4315] Jifu Li, Zhiguang Hu, and Shaoqiang Deng. Cohomogeneity one Randers metrics. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 59(3):575–??, September 2016. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Lin:2016:DII

[4316] Minghua Lin. A determinantal inequality involving partial traces. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 59 (3):585–??, September 2016. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Liu:2016:DDP

[4317] H. Q. Liu. The Dirichlet divisor problem of arithmetic progressions. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 59(3):592–??, September 2016. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Liu:2016:SPS

[4318] Zhixin Liu. Small prime solutions to cubic Diophantine equations II. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 59(3):599–??, September 2016. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Mihailescu:2016:EUN

[4319] Mihai Mihailescu and Gheorghe Morosanu. Eigenvalues of −Δ_{p} − Δ_{q} under Neumann boundary condition. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 59(3):606–??, September 2016. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Nakashima:2016:CSB

[4320] Norihiro Nakashima, Hiroaki Terao, and Shuhei Tsujie. Canonical systems of basic invariants for unitary reflection groups. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 59(3):617–??, September 2016. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Otsubo:2016:HFT

[4321] Noriyuki Otsubo. Homology of the Fermat tower and universal measures for Jacobi sums. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 59(3):624–??, September 2016. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Shaveisi:2016:SRA

[4322] Farzad Shaveisi. Some results on the annihilating-ideal graphs. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 59(3): 641–??, September 2016. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Su:2016:DUC

[4323] Huadong Su. On the diameter of unitary Cayley graphs of rings. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 59(3): 652–??, September 2016. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Ying:2016:RWE

[4324] Zhiling Ying, Tamer Kosan, and Yiqiang Zhou. Rings in which every element is a sum of two tripotents. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 59(3):661–??, September 2016. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Bacak:2016:LRH

[4325] Miroslav Bacák and Leonid V. Kovalev. Lipschitz retractions in Hadamard spaces via gradient flow semigroups. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 59(4):673–??, December 2016. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Carlson:2016:GSG

[4326] Jon F. Carlson, Sunil K. Chebolu, and Ján Minác. Ghosts and strong ghosts in the stable category. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 59(4):682–??, December 2016. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Chen:2016:RCO

[4327] Chung-Chuan Chen. Recurrence of cosine operator functions on groups. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 59(4):693–??, December 2016. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Chen:2016:TCP

[4328] Yichao Chen and Xuluo Yin. The thickness of the Cartesian product of two graphs. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 59(4):705–??, December 2016. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Perez:2016:RHC

[4329] Juan de Dios Pérez, Hyunjin Lee, Young Jin Suh, and Changhwa Woo. Real hypersurfaces in complex two-plane Grassmannians with Reeb parallel Ricci tensor in the GTW connection. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 59(4):721–??, December 2016. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Dimassi:2016:SCA

[4330] Mouez Dimassi. Semi-classical asymptotics for Schrödinger operator with oscillating decaying potential. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 59(4):734–??, December 2016. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Dolzan:2016:MDT

[4331] David Dolzan. The metric dimension of the total graph of a finite commutative ring. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 59(4):748–??, December 2016. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Fichou:2016:AAC

[4332] Goulwen Fichou, Ronan Quarez, and Masahiro Shiota. Artin approximation compatible with a change of variables. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 59(4):760–??, December 2016. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Garcia-Pacheco:2016:GCH

[4333] Francisco Javier García-Pacheco and Justin R. Hill. Geometric characterizations of Hilbert spaces. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 59(4):769–??, December 2016. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Gauthier:2016:CRP

[4334] Paul M. Gauthier and Fatemeh Sharifi. The Carathéodory reflection principle and Osgood–Carathéodory theorem on Riemann surfaces. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 59 (4):776–??, December 2016. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Hashemi:2016:ZDG

[4335] Ebrahim Hashemi and R. Amirjan. Zero-divisor graphs of Ore extensions over reversible rings. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 59(4):794–??, December 2016. CODEN CM-BUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Izumiya:2016:GIL

[4336] Shyuichi Izumiya. Geometric interpretation of Lagrangian equivalence. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 59(4):806–??, December 2016. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Kaimakamis:2016:CTD

[4337] George Kaimakamis, Konstantina Panagiotidou, and Juan de Dios Pérez. A classification of three-dimensional real hypersurfaces in non-flat complex space forms in terms of their generalized Tanaka–Webster Lie derivative. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 59(4):813–??, December 2016. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Karpenko:2016:IPP

[4338] Nikita A. Karpenko. Incompressibility of products of pseudohomogeneous varieties. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 59(4):824–??, December 2016. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Liao:2016:SPT

[4339] Fanghui Liao and Zongguang Liu. Some properties of Triebel–Lizorkin and Besov spaces associated with Zygmund dilations. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 59(4):834–??, December 2016. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Nah:2016:STS

[4340] Kyeongah Nah and Gergely Röst. Stability threshold for scalar linear periodic delay differential equations. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 59(4):849–??, December 2016. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Osserman:2016:SVB

[4341] Brian Osserman. Stability of vector bundles on curves and degenerations. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 59(4):858–??, December 2016. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Pal:2016:MRS

[4342] Sarbeswar Pal. Moduli of rank 2 stable bundles and Hecke curves. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 59 (4):865–??, December 2016. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Wang:2016:CMP

[4343] Jianfei Wang. The Carleson measure problem between analytic Morrey spaces. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 59(4):878–??, December 2016. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Akbari:2017:CAI

[4344] Saeeid Akbari, Abbas Alilou, Jafar Amjadi, and Seyed Mahmoud Sheikholeslami. The co-annihilating-ideal graphs of commutative rings. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 60 (1):3–??, March 2017. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Akbari:2017:CMG

[4345] Saieed Akbari, Babak Miraftab, and Reza Nikandish. Co-maximal graphs of subgroups of groups. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 60(1):12–??, March 2017. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Azimi:2017:SDG

[4346] Ali Azimi and Mohammad Farrokhi Derakhshandeh Ghouchan. Self 2distance graphs. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 60(1):26–??, March 2017. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Bouchemakh:2017:DKP

[4347] Isma Bouchemakh and Kaci Fatma. On the dual König property of the order-interval hypergraph of two classes of N-free posets. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 60(1):

43–??, March 2017. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Button:2017:TFC

[4348] Jack Button. Tubular free by cyclic groups act freely on CAT(0) cube complexes. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 60(1):54–??, March 2017. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Chang:2017:PSR

[4349] Gyu Whan Chang. Power series rings over Prüfer v-multiplication domains, II. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 60(1):63–??, March 2017. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Christ:2017:NGC

[4350] Michael Christ and Marc A. Rieffel. Nilpotent group C^{∗}-algebras as compact quantum metric spaces. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 60(1):77–??, March 2017. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Choi:2017:CFE

[4351] Chang-Kwon Choi, Jaeyoung Chung, Yumin Ju, and John Rassias. Cubic functional equations on restricted domains of Lebesgue measure zero. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 60(1):95–??, March 2017. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Diestel:2017:ENF

[4352] Geoff Diestel. An extension of Nikishin’s factorization theorem. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 60(1): 104–??, March 2017. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Farashahi:2017:APT

[4353] Arash Ghaani Farashahi. Abstract Plancherel (trace) formulas over homogeneous spaces of compact groups. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 60(1):111–??, March 2017. CO-DEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Ghanei:2017:HPA

[4354] Mohammad Reza Ghanei, Rasoul Nasr-Isfahani, and Mehdi Nemati. A homological property and Arens regularity of locally compact quantum groups. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 60(1):122–??, March 2017. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Gurbuz:2017:SEG

[4355] Ferit Gürbüz. Some estimates for generalized commutators of rough fractional maximal and integral operators on generalized weighted Morrey spaces. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 60(1):131–??, March 2017. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Khavinson:2017:DPS

[4356] Dmitry Khavinson, Erik Lundberg, and Hermann Render. The Dirichlet problem for the slab with entire data and a difference equation for harmonic functions. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 60(1):146–??, March 2017. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Liu:2017:CFS

[4357] Ye Liu. On chromatic functors and stable partitions of graphs. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 60(1): 154–??, March 2017. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Morimoto:2017:CHB

[4358] Masaharu Morimoto. Cokernels of homomorphisms from Burnside rings to inverse limits. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 60(1):165–??, March 2017. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Oubbi:2017:USF

[4359] Lahbib Oubbi. On Ulam stability of a functional equation in Banach modules. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 60(1):173–??, March 2017. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Pathak:2017:CL

[4360] Siddhi Pathak. On a conjecture of Livingston. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 60(1):184–??, March 2017. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Rhaly:2017:CT

[4361] H. C. Rhaly, Jr. Corrigendum to “Generalized Cesàro Matrices”. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 60(1): 196–??, March 2017. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic). See [2061].

Tang:2017:DKI

[4362] Zikai Tang and Hanyuan Deng. Degree Kirchhoff index of bicyclic graphs. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 60(1):197–??, March 2017. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Vetrik:2017:MDC

[4363] Tomás Vetrik. The metric dimension of circulant graphs. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 60(1):206– ??, March 2017. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Wang:2017:COS

[4364] Yuanyi Wang. Condition C′_{∧} of operator spaces. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 60(1):217–??, March 2017. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Bahmanpour:2017:FFD

[4365] Kamal Bahmanpour and Reza Naghipour. Faltings’ finiteness dimension of local cohomology modules over local Cohen–Macaulay rings. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 60(2): 225–??, June 2017. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Basu:2017:TCQ

[4366] Samik Basu and B. Subhash. Topology of certain quotient spaces of Stiefel manifolds. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 60(2):235–??, June 2017. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Bhuniya:2017:RGR

[4367] Anjan Kumar Bhuniya and Kalyan Hansda. On radicals of Green’s relations in ordered semigroups. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 60(2):246–??, June 2017. CODEN CM-BUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Chen:2017:YTP

[4368] Bin Chen and Lili Zhao. On a Yamabe type problem in Finsler geometry. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 60(2):253–??, June 2017. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Chen:2017:CRC

[4369] Jianlong Chen, Huihui Zhu, Pedro Patricio, and Yulin Zhang. Characterizations and representations of core and dual core inverses. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 60(2): 269–??, June 2017. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Friedl:2017:TAI

[4370] Stefan Friedl and Stefano Vidussi. Twisted Alexander invariants detect trivial links. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 60(2):283–??, June 2017. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Gauthier:2017:LTH

[4371] Paul M. Gauthier and Fatemeh Sharifi. Luzin-type holomorphic approximation on closed subsets of open Riemann surfaces. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 60(2):300–??, June 2017. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Hein:2017:CMF

[4372] Nickolas Hein, Frank Sottile, and Igor Zelenko. A congruence modulo four for real Schubert calculus with isotropic flags. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 60(2):309–??, June 2017. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Khojasteh:2017:WNG

[4373] Sohiela Khojasteh and Mohammad Javad Nikmehr. The weakly nilpotent graph of a commutative ring. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 60(2):319–??, June 2017. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Fourn:2017:NCV

[4374] Samuel Le Fourn. Nonvanishing of central values of L-functions of newforms in S_{2}(Γ_{0}(dp^{2})) twisted by quadratic characters. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 60(2):329– ??, June 2017. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Ma:2017:ILQ

[4375] Yumei Ma. Isometry on linear n-G-quasi normed spaces. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 60(2):350–??, June 2017. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Preda:2017:RQP

[4376] Ciprian Preda. On the roughness of quasinilpotency property of oneparameter semigroups. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 60(2):364–??, June 2017. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Rao:2017:CL

[4377] M. Sambasiva Rao. Coaxer lattices. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 60(2):372–??, June 2017. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Kumar:2017:IMC

[4378] N. Shravan Kumar. Invariant means on a class of von Neumann algebras related to ultraspherical hypergroups II. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 60(2):402–??, June 2017. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Stoyanov:2017:GMS

[4379] Luchezar Stoyanov. On Gibbs measures and spectra of Ruelle transfer operators. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 60(2):411–??, June 2017. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Tang:2017:NSQ

[4380] Xianhua Tang. New super-quadratic conditions for asymptotically periodic Schrödinger equations. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 60(2):422–??, June 2017. CODEN CM-BUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Weng:2017:GAS

[4381] Peixuan Weng and Li Liu. Globally asymptotic stability of a delayed integro-differential equation with nonlocal diffusion. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 60(2):436–??, June 2017. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Rousseau:2017:BDC

[4382] C. Rousseau. The bifurcation diagram of cubic polynomial vector fields on CP^{1}. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 60(2):??, June 2017. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Alaghmandan:2017:CDC

[4383] Mahmood Alaghmandan and Jason Crann. Character density in central subalgebras of compact quantum groups. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 60(3):449–??, September 2017. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Bayart:2017:FUA

[4384] Frédéric Bayart and Paul M. Gauthier. Functions universal for all translation operators in several complex variables. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 60(3):462–??, September 2017. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Buijs:2017:MCE

[4385] Urtzi Buijs, Yves Félix, Aniceto Murillo, and Daniel Tanré. Maurer– Cartan elements in the Lie models of finite simplicial complexes. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 60(3): 470–??, September 2017. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Carrell:2017:SWG

[4386] Jim Carrell and Kiumars Kaveh. Springer’s Weyl group representation via localization. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 60(3):478–??, September 2017. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Dobrowolski:2017:NLT

[4387] Edward Dobrowolski. A note on Lawton’s Theorem. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 60(3):484–??, September 2017. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Fiori:2017:RHT

[4388] Andrew Fiori. A Riemann–Hurwitz theorem for the algebraic Euler characteristic. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 60(3):490–??, September 2017. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Haase:2017:CNP

[4389] Christian Haase and Jan Hofmann. Convex-normal (pairs of) polytopes. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 60(3):510–??, September 2017. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Iena:2017:SSF

[4390] Oleksandr Iena and Alain Leytem. On the singular sheaves in the fine Simpson moduli spaces of 1-dimensional sheaves. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 60(3):522–??, September 2017. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Kalaj:2017:GSP

[4391] David Kalaj and Djordjije Vujadinović. The gradient of a solution of the Poisson equation in the unit ball and related operators. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 60(3): 536–??, September 2017. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Karzhemanov:2017:PKS

[4392] Ilya Karzhemanov. On polarized K3 surfaces of genus 33. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 60(3):546–??, September 2017. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Kurdyka:2017:NTP

[4393] Krzysztof Kurdyka and Laurentiu Paunescu. Nuij type pencils of hyperbolic polynomials. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 60(3):561–??, September 2017. CODEN CM-BUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Li:2017:WFH

[4394] Ji Li and Brett D. Wick. Weak factorizations of the Hardy space H^{1}(R^{n}) in terms of multilinear Riesz transforms. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 60(3):571–??, September 2017. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Liu:2017:ERM

[4395] Feng Liu and Huoxiong Wu. Endpoint regularity of multisublinear fractional maximal functions. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 60(3):586–??, September 2017. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Louder:2017:SCC

[4396] Larsen Louder and Henry Wilton. Stackings and the W -cycles conjecture. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 60(3):604–??, September 2017. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Reichstein:2017:DLD

[4397] Zinovy Reichstein and Angelo Vistoli. On the dimension of the locus of determinantal hypersurfaces. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 60(3):613–??, September 2017. CO-DEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Shahrokhi-Dehkordi:2017:TMS

[4398] M. S. Shahrokhi-Dehkordi. Traceless maps as the singular minimizers in the multi-dimensional calculus of variations. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 60(3):631–??, September 2017. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Werner:2017:MDM

[4399] Elisabeth Werner and Deping Ye. Mixed f-divergence for multiple pairs of measures. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 60(3):641–??, September 2017. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Zhuo:2017:CBT

[4400] Ciqiang Zhuo, Winfried Sickel, Dachun Yang, and Wen Yuan. Characterizations of Besov-type and Triebel–Lizorkin-type spaces via averages on balls. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 60(3):655–??, September 2017. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Abtahi:2017:CAI

[4401] Fatemeh Abtahi, Mohsen Azizi, and Ali Rejali. Character amenability of the intersection of Lipschitz algebras. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 60(4):673–??, December 2017. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Bao:2017:SDT

[4402] Guanlong Bao, Nihat Gökhan Gögüs, and Stamatis Pouliasis. Q_{√} spaces

and Dirichlet type spaces. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 60(4):690–??, December 2017. CODEN CM-BUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Benelkourchi:2017:EAD

[4403] Slimane Benelkourchi. Envelope approach to degenerate complex Monge– Ampère equations on compact Kähler manifolds. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 60(4):705–??, December 2017. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Chen:2017:DHW

[4404] Chung-Chuan Chen. Disjoint hypercyclicity and weighted translations on discrete groups. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 60(4):712–??, December 2017. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Eroglu:2017:ICG

[4405] Münevver Pinar Eroğlu and Nurcan Argaç. On identities with composition of generalized derivations. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 60(4):721–??, December 2017. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Huang:2017:IGW

[4406] Yanhe Huang, Frank Sottile, and Igor Zelenko. Injectivity of generalized Wronski maps. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 60(4):747–??, December 2017. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Jantzen:2017:MWC

[4407] Jens Carsten Jantzen. Maximal weight composition factors for Weyl modules. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 60(4):762–??, December 2017. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Jensen:2017:SSG

[4408] Gerd Jensen and Christian Pommerenke. On the structure of the Schild group in relativity theory. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 60(4):774–??, December 2017. CODEN CM-BUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Jiang:2017:RDT

[4409] Chunlan Jiang. Reduction to dimension two of local spectrum for AH algebra with ideal property. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 60(4):791–??, December 2017. CODEN CM-BUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Liu:2017:SCC

[4410] Zhongyun Liu, Xiaorong Qin, Nianci Wu, and Yulin Zhang. The shifted classical circulant and skew circulant splitting iterative methods for Toeplitz matrices. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 60(4):807–??, December 2017. CODEN CM-BUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Moslehian:2017:COB

[4411] Mohammad Sal Moslehian and Ali Zamani. Characterizations of operator Birkhoff–James orthogonality. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 60(4):816–??, December 2017. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Motegi:2017:GTE

[4412] Kimihiko Motegi and Masakazu Teragaito. Generalized torsion elements and bi-orderability of 3-manifold groups. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 60(4):830–??, December 2017. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Pachl:2017:CCS

[4413] Jan Pachl and Juris Steprans. Continuity of convolution and SIN groups. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 60(4):845–??, December 2017. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

delaFuente:2017:KCH

[4414] Jesús Suárez de la Fuente. The Kottman constant for α-Hölder maps. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 60(4):855–??, December 2017. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Wang:2017:COG

[4415] Long Wang, Nieves Castro-Gonzalez, and Jianlong Chen. Characterizations of outer generalized inverses. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 60(4):861–??, December 2017. CO-DEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Xu:2017:RFO

[4416] Xu Xu and Laiyi Zhu. Rational function operators from Poisson integrals. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 60(4):872–??, December 2017. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Zheng:2017:TEI

[4417] Yuefei Zheng and Zhaoyong Huang. Triangulated equivalences involving Gorenstein projective modules. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 60(4):879–??, December 2017. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Gilligan:2017:LPP

[4418] Bruce Gilligan. Levi’s problem for pseudoconvex homogeneous manifolds. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 60(4):??, December 2017. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Alhasanat:2018:PSS

[4419] Ahmad Alhasanat and Chunhua Ou. Periodic steady-state solutions of a liquid film model via a classical method. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 61(1):3–15, March 01, 2018. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-

4287 (electronic). URL https://www.cambridge.org/core/journals/
canadian-mathematical-bulletin/article/periodic-steadystate-
solutions-of-a-liquid-film-model-via-a-classical-method/1B78B795B290B1B9FCE65C19EE20AB23.

Bavula:2018:CSW

[4420] V. V. Bavula and T. Lu. Classification of simple weight modules over the Schrödinger algebra. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 61(1):16–39, March 01, 2018. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic). URL https://www.

cambridge.org/core/journals/canadian-mathematical-bulletin/
article/classification-of-simple-weight-modules-over-the-schrodinger-
algebra/599671DEF67720E3D343813E769D7DED.

Chen:2018:SBR

[4421] Wengu Chen and Huanmin Ge. A sharp bound on RIC in generalized orthogonal matching pursuit. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 61(1):40–54, March 01, 2018. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-

4287 (electronic). URL https://www.cambridge.org/core/journals/
canadian-mathematical-bulletin/article/sharp-bound-on-ric-in-
generalized-orthogonal-matching-pursuit/F343FAE65C5D94B3E0713023BC74D2F5.

Chen:2018:EUE

[4422] Yichao Chen, Xiaojian Gao, and Yuanqiu Huang. Enumerating unlabelled embeddings of digraphs. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 61(1):55–69, March 01, 2018. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-

4287 (electronic). URL https://www.cambridge.org/core/journals/
canadian-mathematical-bulletin/article/enumerating-unlabelled-
embeddings-of-digraphs/890E123C93C7C0B367D621960FB69356.

Dang:2018:HTR

[4423] Pei Dang, Hua Liu, and Tao Qian. Hilbert transformation and representation of the ax + b group. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 61(1):70–84, March 01, 2018. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-

4287 (electronic). URL https://www.cambridge.org/core/journals/
canadian-mathematical-bulletin/article/hilbert-transformation-
and-representation-of-the-ax-b-group/2D4443D87E37B1DF31E95469F36AC8FC.

Ding:2018:SSF

[4424] Fan Ding, Hansjörg Geiges, and Guangjian Zhang. On subcritically Stein fillable 5-manifolds. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 61(1):85–96, March 01, 2018. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic). URL https://www.

cambridge.org/core/journals/canadian-mathematical-bulletin/
article/on-subcritically-stein-fillable-5manifolds/5B3354418F50B87BA5893FF75EED8DED.

Ding:2018:SIC

[4425] Yong Ding and Xudong Lai. On a singular integral of Christ– Journé type with homogeneous kernel. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 61(1):97–113, March 01, 2018. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-

4287 (electronic). URL https://www.cambridge.org/core/journals/
canadian-mathematical-bulletin/article/on-a-singular-integral-
of-christjourne-type-with-homogeneous-kernel/9BBD2808D4C2E2E9ACAEEC958AA9033F.

Haralampidou:2018:CCN

[4426] Marina Haralampidou, Mohamed Oudadess, Lourdes Palacios, and Carlos Signoret. A characterization of C^{∗}-normed algebras via positive functionals. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 61(1):114–123, March 01, 2018. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic). URL https://www.

cambridge.org/core/journals/canadian-mathematical-bulletin/
article/characterization-of-cnormed-algebras-via-positive-functionals/
C8330988197C7DABB7E0DD239DF1EBC6.

Hemasundar:2018:JCT

[4427] Gollakota V. V. Hemasundar and R. R. Simha. The Jordan curve theorem via complex analysis. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 61(1):124–129, March 01, 2018. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-

4287 (electronic). URL https://www.cambridge.org/core/journals/
canadian-mathematical-bulletin/article/jordan-curve-theorem-
via-complex-analysis/C35BD0E6EC72FE1A1AD8A1FE56430A12.

Kosan:2018:AMU

[4428] Tamer Kosan, Serap Sahinkaya, and Yiqiang Zhou. Additive maps on units of rings. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 61(1):130–141, March 01, 2018. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic). URL https://www.

cambridge.org/core/journals/canadian-mathematical-bulletin/
article/additive-maps-on-units-of-rings/E19846CA1BC2E78F6FD3D3C29EA8362D.

Li:2018:EFP

[4429] Bao Qin Li. An equivalent form of Picard’s theorem and beyond. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 61(1):142–148, March 01, 2018. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic). URL https://www.

cambridge.org/core/journals/canadian-mathematical-bulletin/
article/an-equivalent-form-of-picards-theorem-and-beyond/F6BD7A04544ACD2016D59BD93F97CC40.

Llibre:2018:GPP

[4430] Jaume Llibre and Claudia Valls. Global phase portraits for the Abel quadratic polynomial differential equations of the second kind with Z^{2}- symmetries. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 61(1):149–165, March 01, 2018. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic). URL https://www.

cambridge.org/core/journals/canadian-mathematical-bulletin/

article/global-phase-portraits-for-the-abel-quadratic-polynomial-
differential-equations-of-the-second-kind-with-z2symmetries/
6721B99A744F2795F07E69D837F73FC1.

Miranda-Neto:2018:MTC

[4431] Cleto B. Miranda-Neto. A module-theoretic characterization of algebraic hypersurfaces. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 61(1):166–173, March 01, 2018. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic). URL https://www.

cambridge.org/core/journals/canadian-mathematical-bulletin/
article/moduletheoretic-characterization-of-algebraic-hypersurfaces/
39944945CD8D8F377E68E7A3DD9C2479.

Roche:2018:FRC

[4432] Alan Roche and C. Ryan Vinroot. A factorization result for classical and similitude groups. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 61(1):174–190, March 01, 2018. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287

(electronic).  URL https://www.cambridge.org/core/journals/
canadian-mathematical-bulletin/article/factorization-result-
for-classical-and-similitude-groups/FE5948800F09B8FBFCC474E15CD6BE0A.

Saito:2018:FST

[4433] Hiroki Saito and Hitoshi Tanaka. The Fefferman-Stein type inequalities for strong and directional maximal operators in the plane. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 61(1):191– 200, March 01, 2018. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-

4287 (electronic). URL https://www.cambridge.org/core/journals/
canadian-mathematical-bulletin/article/feffermanstein-type-
inequalities-for-strong-and-directional-maximal-operators-in-
the-plane/2D1CA6DF1631B44B13713BDF6AA453CB.

Takahashi:2018:PPB

[4434] Tomokuni Takahashi. Projective plane bundles over an elliptic curve. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 61(1):201–210, March 01, 2018. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic). URL https://www.

cambridge.org/core/journals/canadian-mathematical-bulletin/
article/projective-plane-bundles-over-an-elliptic-curve/C5350FC76DDA5FFEDBD77AAE4E373189.

Tran:2018:AHD

[4435] Anh T. Tran and Yoshikazu Yamaguchi. The asymptotics of the higher dimensional Reidemeister torsion for exceptional surgeries along twist knots. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 61(1):211–224, March 01, 2018. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic). URL https://www.

cambridge.org/core/journals/canadian-mathematical-bulletin/
article/asymptotics-of-the-higher-dimensional-reidemeister-
torsion-for-exceptional-surgeries-along-twist-knots/02D7C643341A7092B40C45EF59E9A9EC.

Bichon:2018:HBN

[4436] Julien Bichon, David Kyed, and Sven Raum. Higher ℓ^{2}-Betti numbers of universal quantum groups. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 61(2):225–235, June 01, 2018. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-

4287 (electronic). URL https://www.cambridge.org/core/journals/
canadian-mathematical-bulletin/article/higher-2betti-numbers-
of-universal-quantum-groups/DCA8CDEAB810857211E9DCE8D864A023.

Boutonnet:2018:NUB

[4437] Remi Boutonnet and Jean Roydor. A note on uniformly bounded cocycles into finite von Neumann algebras. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 61(2):236–239, June 01, 2018. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-

4287 (electronic). URL https://www.cambridge.org/core/journals/
canadian-mathematical-bulletin/article/note-on-uniformly-bounded-
cocycles-into-finite-von-neumann-algebras/25E5721CDB75874CB4499B39635F2C6C.

Bu:2018:HCS

[4438] Shangquan Bu and Gang Cai. Hölder continuous solutions of degenerate differential equations with finite delay. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 61(2):240–251, June 01, 2018. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-

4287 (electronic). URL https://www.cambridge.org/core/journals/
canadian-mathematical-bulletin/article/holder-continuous-solutions-
of-degenerate-differential-equations-with-finite-delay/01740E419DF063C09373417D26B859E3.

Dewar:2018:CH

[4439] Megan Dewar, David Pike, and John Proos. Connectivity in hypergraphs. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 61(2):252–271, June 01, 2018. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic). URL https://www.

cambridge.org/core/journals/canadian-mathematical-bulletin/
article/connectivity-in-hypergraphs/99EE7E7BDDD26CA3365AED071948707D.

Franz:2018:SPE

[4440] Matthias Franz. Symmetric products of equivariantly formal spaces. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 61 (2):272–281, June 01, 2018. CODEN CMBUA3. ISSN 0008-4395 (print),

1496-4287 (electronic). URL https://www.cambridge.org/core/
journals/canadian-mathematical-bulletin/article/symmetric-products-
of-equivariantly-formal-spaces/DF163F05DB0616E04B2E1077001CBC82.

Gauthier:2018:ITC

[4441] Paul M. Gauthier. The initial and terminal cluster sets of an analytic curve. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 61(2):282–288, June 01, 2018. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic). URL https://www.

cambridge.org/core/journals/canadian-mathematical-bulletin/
article/initial-and-terminal-cluster-sets-of-an-analytic-curve/
A402B8F838CE94B91E14615786B07799.

Gupta:2018:RAN

[4442] Purvi Gupta. A real-analytic nonpolynomially convex isotropic torus with no attached discs. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 61(2):289–291, June 01, 2018. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287

(electronic).  URL https://www.cambridge.org/core/journals/
canadian-mathematical-bulletin/article/realanalytic-nonpolynomially-
convex-isotropic-torus-with-no-attached-discs/230E49ED9D4B74D29A7DFB2FA29D248B.

Ha:2018:NUT

[4443] Pham Hoang Ha and Yu Kawakami. A note on a unicity theorem for the Gauss maps of complete minimal surfaces in Euclidean four-space. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 61(2):292–300, June 01, 2018. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic). URL https://www.

cambridge.org/core/journals/canadian-mathematical-bulletin/
article/note-on-a-unicity-theorem-for-the-gauss-maps-of-complete-
minimal-surfaces-in-euclidean-fourspace/80A7A4C3F25948AF6C4510FD8576807B.

Joziak:2018:RHI

[4444] Paweł Józiak. Remarks on Hopf images and quantum permutation groups S_{n}^{+}. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 61(2):301–317, June 01, 2018. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic). URL https://www.

cambridge.org/core/journals/canadian-mathematical-bulletin/

article/remarks-on-hopf-images-and-quantum-permutation-groups-
sn/C2E37A2F3B335F234707F5CCF5736D7C.

Lee:2018:ANE

[4445] Tsiu-Kwen Lee. Ad-nilpotent elements of semiprime rings with involution. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 61(2):318–327, June 01, 2018. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic). URL https://www.

cambridge.org/core/journals/canadian-mathematical-bulletin/
article/adnilpotent-elements-of-semiprime-rings-with-involution/
F9F50D95ABCFD31BDE926845BEFDD209.

Maican:2018:MSS

[4446] Mario Maican. Moduli of space sheaves with Hilbert polynomial 4m + 1. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 61(2):328–345, June 01, 2018. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic). URL https://www.

cambridge.org/core/journals/canadian-mathematical-bulletin/
article/moduli-of-space-sheaves-with-hilbert-polynomial-4m-
1/56EE108C4B4C330AA522EA474236CCBF.

Polak:2018:CSP

[4447] Jason K. C. Polak. Counting separable polynomials in Z/n[x]. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 61(2): 346–352, June 01, 2018. CODEN CMBUA3. ISSN 0008-4395 (print),

1496-4287 (electronic). URL https://www.cambridge.org/core/
journals/canadian-mathematical-bulletin/article/counting-separable-
polynomials-in-nx/77D5E3E32DF411544DF02370A2DAE26Bo.

Qin:2018:GSM

[4448] Dongdong Qin, Yubo He, and Xianhua Tang. Ground state and multiple solutions for Kirchhoff type equations with critical exponent. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 61(2):353–369, June 01, 2018. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic). URL https://www.

cambridge.org/core/journals/canadian-mathematical-bulletin/
article/ground-state-and-multiple-solutions-for-kirchhoff-type-
equations-with-critical-exponent/F746AC742EEA60FABC7CEFC4621CC09A.

Rocha:2018:RCI

[4449] Pablo Alejandro Rocha. A remark on certain integral operators of fractional type. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 61(2):370–375, June 01, 2018. CODEN CMBUA3.

ISSN 0008-4395 (print), 1496-4287 (electronic). URL https://www.

cambridge.org/core/journals/canadian-mathematical-bulletin/
article/remark-on-certain-integral-operators-of-fractional-
type/B8DE846A3F74613D34A415342091AA8B.

Sebbar:2018:EZF

[4450] Abdellah Sebbar and Isra Al-Shbeil. Elliptic zeta functions and equivariant functions. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 61(2):376–389, June 01, 2018. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic). URL https://www.

cambridge.org/core/journals/canadian-mathematical-bulletin/
article/elliptic-zeta-functions-and-equivariant-functions/CFF0E4F61D4F5523C34C854826FD52CD.

Wang:2018:MTA

[4451] Li an Daniel Wang. A multiplier theorem on anisotropic Hardy spaces. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 61(2):390–404, June 01, 2018. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic). URL https://www.

cambridge.org/core/journals/canadian-mathematical-bulletin/
article/multiplier-theorem-on-anisotropic-hardy-spaces/DDA9EF209F0FD5AB8750C56FF3872976.

Yukita:2018:GRD

[4452] Tomoshige Yukita. Growth rates of 3-dimensional hyperbolic Coxeter groups are Perron numbers. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 61(2):405–422, June 01, 2018. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-

4287 (electronic). URL https://www.cambridge.org/core/journals/
canadian-mathematical-bulletin/article/growth-rates-of-3dimensional-
hyperbolic-coxeter-groups-are-perron-numbers/CC4992B9431214E0F9766111E72E7F63.

Zhang:2018:STW

[4453] Guo-Bao Zhang and Ge Tian. Stability of traveling wavefronts for a two-component lattice dynamical system arising in competition models. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 61(2):423–437, June 01, 2018. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic). URL https://www.

cambridge.org/core/journals/canadian-mathematical-bulletin/
article/stability-of-traveling-wavefronts-for-a-twocomponent-
lattice-dynamical-system-arising-in-competition-models/434D8D574722460F93362EB8D2EC42DF.

Zhang:2018:CSH

[4454] Tao Zhang and Chunqin Zhou. Classification of solutions for harmonic functions with Neumann boundary value. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 61(2):438–448, June 01, 2018. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-

4287 (electronic). URL https://www.cambridge.org/core/journals/
canadian-mathematical-bulletin/article/classification-of-solutions-
for-harmonic-functions-with-neumann-boundary-value/85F6664895C2D014E17126B89030CC0D.

Abrahamsen:2018:SEP

[4455] Trond A. Abrahamsen, Petr Hájek, Olav Nygaard, and Stanimir L. Troyanski. Strongly extreme points and approximation properties. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 61(3):449–457, September 01, 2018. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic). URL https://www.

cambridge.org/core/journals/canadian-mathematical-bulletin/
article/strongly-extreme-points-and-approximation-properties/
884FB3B41394A0F96FF357B0B465AD0A.

Agler:2018:GHF

[4456] Jim Agler and John McCarthy. Global holomorphic functions in several non-commuting variables II. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 61(3):458–463, September 01, 2018. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-

4287 (electronic). URL https://www.cambridge.org/core/journals/
canadian-mathematical-bulletin/article/global-holomorphic-functions-
in-several-noncommuting-variables-ii/0D223D105BF892DC867397FD59F1D205.

Aujogue:2018:SPC

[4457] Jean-Baptiste Aujogue. A short proof of the characterization of model sets by almost automorphy. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 61(3):464–472, September 01, 2018. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-

4287 (electronic). URL https://www.cambridge.org/core/journals/
canadian-mathematical-bulletin/article/short-proof-of-the-characterization-
of-model-sets-by-almost-automorphy/28B8C8607DC77372D04CAA3F9A4C1E7C.

Awonusika:2018:SIJ

[4458] Richard Awonusika and Ali Taheri. A spectral identity on Jacobi polynomials and its analytic implications. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 61(3):473–482, September 01, 2018. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-

4287 (electronic). URL https://www.cambridge.org/core/journals/
canadian-mathematical-bulletin/article/spectral-identity-on-
jacobi-polynomials-and-its-analytic-implications/2DBF7BFEA25AA355A4A8A025732EDAA8.

Banica:2018:TDA

[4459] Teodor Banica. Tannakian duality for affine homogeneous spaces. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 61(3):483–494, September 01, 2018. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic). URL https://www.

cambridge.org/core/journals/canadian-mathematical-bulletin/
article/tannakian-duality-for-affine-homogeneous-spaces/E339E010741F5F2266BD9E4B2903896A.

Chang:2018:PLQ

[4460] Der-Chen Chang, Nanping Yang, and Hsi-Chun Wu. Poincaré lemma on quaternion-like Heisenberg groups. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 61(3):495–508, September 01, 2018. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-

4287 (electronic). URL https://www.cambridge.org/core/journals/
canadian-mathematical-bulletin/article/poincare-lemma-on-quaternionlike-
heisenberg-groups/A994A1A0EC2FCFE38B571C9373899B8D.

Cui:2018:AIM

[4461] Xiaohui Cui, Chunjie Wang, and Kehe Zhu. Area integral means of analytic functions in the unit disk. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 61(3):509–517, September 01, 2018. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-

4287 (electronic). URL https://www.cambridge.org/core/journals/
canadian-mathematical-bulletin/article/area-integral-means-
of-analytic-functions-in-the-unit-disk/DE2066BE7C2DE2CA5C322B6EB0C871DC.

Hong:2018:BMS

[4462] Kyungpyo Hong and Seungsang Oh. Bounds on multiple self-avoiding polygons. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 61(3):518–530, September 01, 2018. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic). URL https://www.

cambridge.org/core/journals/canadian-mathematical-bulletin/
article/bounds-on-multiple-selfavoiding-polygons/908A99F47209C96267873F62F3023971.

Ingram:2018:AUA

[4463] Patrick Ingram. p-adic uniformization and the action of Galois on certain affine correspondences. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 61(3):531–542, September 01, 2018. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-

4287 (electronic). URL https://www.cambridge.org/core/journals/
canadian-mathematical-bulletin/article/padic-uniformization-
and-the-action-of-galois-on-certain-affine-correspondences/
FB5257BE4BD1611910CB9E76EA871876.

Jeong:2018:LDR

[4464] Imsoon Jeong, Juan de Dios Pérez, Young Jin Suh, and Changhwa Woo. Lie derivatives and Ricci tensor on real hypersurfaces in complex two-plane Grassmannians. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 61(3):543–552, September 01, 2018. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-

4287 (electronic). URL https://www.cambridge.org/core/journals/
canadian-mathematical-bulletin/article/lie-derivatives-and-
ricci-tensor-on-real-hypersurfaces-in-complex-twoplane-grassmannians/
5DD84F0AC4916EAF6ECC6A84361C9077.

Karimianpour:2018:BRF

[4465] Camelia Karimianpour. Branching rules for n-fold covering groups of SL_{2} over a non-Archimedean local field. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 61(3):553–571, September 01, 2018. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-

4287 (electronic). URL https://www.cambridge.org/core/journals/
canadian-mathematical-bulletin/article/branching-rules-for-
nfold-covering-groups-of-sl2-over-a-nonarchimedean-local-field/
F58F464840072635AEBDA8CC7A2919D1.

Koskivirta:2018:NCE

[4466] Jean-Stefan Koskivirta. Normalization of closed Ekedahl–Oort strata. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 61(3):572–587, September 01, 2018. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic). URL https://www.

cambridge.org/core/journals/canadian-mathematical-bulletin/
article/normalization-of-closed-ekedahloort-strata/3D55E0E4EA18FBA0BD60269B0F618C4A.

Lang:2018:VCA

[4467] Honglei Lang, Yunhe Sheng, and Aïssa Wade. VB-Courant algebroids, e-Courant algebroids and generalized geometry. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 61(3):588–607, September 01, 2018. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic). URL https://www.

cambridge.org/core/journals/canadian-mathematical-bulletin/
article/vbcourant-algebroids-ecourant-algebroids-and-generalized-
geometry/2D575448B26A24C7E9E05FC0370217FD.

Loeffler:2018:NAR

[4468] David Loeffler. A note on p-adic Rankin–Selberg L-functions. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 61(3):608–621, September 01, 2018. CODEN CMBUA3.

ISSN 0008-4395 (print), 1496-4287 (electronic). URL https://www.

cambridge.org/core/journals/canadian-mathematical-bulletin/
article/note-on-padic-rankinselberg-lfunctions/C8EEBF440001F2C6ADDA46136AAB8A15.

Maier:2018:SEN

[4469] Helmut Maier and Michael Th. Rassias. On the size of an expression in the Nyman–Beurling–Báez–Duarte criterion for the Riemann hypothesis. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 61(3):622–627, September 01, 2018. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic). URL https://www.

cambridge.org/core/journals/canadian-mathematical-bulletin/
article/on-the-size-of-an-expression-in-the-nymanbeurlingbaezduarte-
criterion-for-the-riemann-hypothesis/77DAE25DB88A44714D7A7B271A072900.

Markovic:2018:DFC

[4470] Marijan Marković. Differential-free characterisation of smooth mappings with controlled growth. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 61(3):628–636, September 01, 2018. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-

4287 (electronic). URL https://www.cambridge.org/core/journals/
canadian-mathematical-bulletin/article/differentialfree-characterisation-
of-smooth-mappings-with-controlled-growth/C42398A0B457B2AA0A4B4368A0B15C09.

Nemirovski:2018:US

[4471] Stefan Nemirovski and Rasul Gazimovich Shafikov. Uniformization and steinness. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 61(3):637–639, September 01, 2018. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic). URL https://www.

cambridge.org/core/journals/canadian-mathematical-bulletin/
article/uniformization-and-steinness/B2F94F1EFCEC22F1E87E0BBF65D617D9.

Reijonen:2018:RIM

[4472] Atte Reijonen. Remark on integral means of derivatives of Blaschke products. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 61(3):640–649, September 01, 2018. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic). URL https://www.

cambridge.org/core/journals/canadian-mathematical-bulletin/
article/remark-on-integral-means-of-derivatives-of-blaschke-
products/B6C7B9604889934EE1346FE4DA946B95.

Shirane:2018:CNE

[4473] Taketo Shirane. Connected numbers and the embedded topology of plane curves. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 61(3):650–658, September 01, 2018. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic). URL https://www.

cambridge.org/core/journals/canadian-mathematical-bulletin/
article/connected-numbers-and-the-embedded-topology-of-plane-
curves/7C0014A475AC1453C7A71A63B7C664BA.

Wang:2018:DNH

[4474] Zhenjian Wang. On deformations of nodal hypersurfaces. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 61(3):659–672, September 01, 2018. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic). URL https://www.

cambridge.org/core/journals/canadian-mathematical-bulletin/
article/on-deformations-of-nodal-hypersurfaces/7E675DE4655D4D81AB2EDA3E10147471.

Alhowaity:2018:REC

[4475] Sawsan Alhowaity, Florin Diacu, and Ernesto Pérez-Chavela. Relative equilibria in curved restricted 4-body problems. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 61(4):673–687, December 01, 2018. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-

4287 (electronic). URL https://www.cambridge.org/core/journals/
canadian-mathematical-bulletin/article/relative-equilibria-
in-curved-restricted-4body-problems/72A5D665B27D7AA6AAB80386961F6D41.

Bavula:2018:UEA

[4476] V. V. Bavula and T. Lu. The universal enveloping algebra of the Schrödinger algebra and its prime spectrum. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 61(4):688–703, December 01, 2018. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-

4287 (electronic). URL https://www.cambridge.org/core/journals/
canadian-mathematical-bulletin/article/universal-enveloping-
algebra-of-the-schrodinger-algebra-and-its-prime-spectrum/98F2A260E1F37041E14625194C9FEF65.

Beneteau:2018:RIF

[4477] Catherine Bénéteau, Matthew C. Fleeman, Dmitry S. Khavinson, Daniel Seco, and Alan A. Sola. Remarks on inner functions and optimal approximants. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 61(4):704–716, December 01, 2018. CODEN CMBUA3. ISSN 0008-4395 (print),

1496-4287 (electronic). URL https://www.cambridge.org/core/
journals/canadian-mathematical-bulletin/article/remarks-on-
inner-functions-and-optimal-approximants/346102AD9B0AEB6F058EC80BE9C4017E.

Bu:2018:PSS

[4478] Shangquan Bu and Gang Cai. Periodic solutions of second order degenerate differential equations with delay in Banach spaces. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 61(4):717–737, December 01, 2018. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic). URL https://www.

cambridge.org/core/journals/canadian-mathematical-bulletin/
article/periodic-solutions-of-second-order-degenerate-differential-
equations-with-delay-in-banach-spaces/1C3B43166778D5EADC01AA996AFA6D50.

Cruz-Uribe:2018:PIN

[4479] David Cruz-Uribe, Scott Rodney, and Emily Rosta. Poincaré inequalities and Neumann problems for the p-Laplacian. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 61(4):738–753, December 01, 2018. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic). URL https://www.

cambridge.org/core/journals/canadian-mathematical-bulletin/
article/poincare-inequalities-and-neumann-problems-for-the-
plaplacian/11AF022C9FE45D0D105DC2E01D1ABFE1.

Lidman:2018:NCP

[4480] Tye Lidman and Eamonn Tweedy. A note on concordance properties of fibers in Seifert homology spheres. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 61(4):754–767, December 01, 2018. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-

4287 (electronic). URL https://www.cambridge.org/core/journals/
canadian-mathematical-bulletin/article/note-on-concordance-
properties-of-fibers-in-seifert-homology-spheres/FFC5455615E2DCE6FA45348F646DA731.

Li:2018:CVT

[4481] Liangliang Li, Jing Tian, and Goong Chen. Chaotic vibration of a two-dimensional non-strictly hyperbolic equation. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 61(4):768–786, December 01, 2018. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-

4287 (electronic). URL https://www.cambridge.org/core/journals/
canadian-mathematical-bulletin/article/chaotic-vibration-of-
a-twodimensional-nonstrictly-hyperbolic-equation/93957C8C62CCE0019A9F1B1F655B3C92.

Liu:2018:EER

[4482] Yu Liu and Shuai Qi. Endpoint estimates of Riesz transforms associated with generalized Schrödinger operators. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 61(4):787–801, December 01, 2018. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-

4287 (electronic). URL https://www.cambridge.org/core/journals/
canadian-mathematical-bulletin/article/endpoint-estimates-of-
riesz-transforms-associated-with-generalized-schrodinger-operators/
CE2137311104386DC8338107D4A144F0.

Li:2018:OHH

[4483] Junfeng Li and Haixia Yu. The oscillatory hyper-Hilbert transform associated with plane curves. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 61(4):802–811, December 01, 2018. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-

4287 (electronic). URL https://www.cambridge.org/core/journals/
canadian-mathematical-bulletin/article/oscillatory-hyperhilbert-
transform-associated-with-plane-curves/AA21DCA9A792782D851E2B3C29FA3B0E.

Medini:2018:IPC

[4484] Andrea Medini, Jan van Mill, and Lyubomyr Zdomskyy. Infinite powers and Cohen reals. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 61(4):812–821, December 01, 2018. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-

4287 (electronic). URL https://www.cambridge.org/core/journals/
canadian-mathematical-bulletin/article/infinite-powers-and-
cohen-reals/845A4959060347A78C0F97436479C8A5.

Pollack:2018:MRS

[4485] Aaron Pollack and Shrenik Shah. Multivariate Rankin–Selberg integrals on GL_{4} and GU(2, 2). Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 61(4):822–835, December 01, 2018. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-

4287 (electronic). URL https://www.cambridge.org/core/journals/
canadian-mathematical-bulletin/article/multivariate-rankinselberg-
integrals-on-gl4-and-gu2-2/C576E78DEB4FC0538DF6D82EEC828768.

Purbhoo:2018:TNS

[4486] Kevin Purbhoo. Total nonnegativity and stable polynomials. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 61(4): 836–847, December 01, 2018. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic). URL https://www.cambridge.org/

core/journals/canadian-mathematical-bulletin/article/total-
nonnegativity-and-stable-polynomials/4796CABDD1859B72928234BC7FB676E0.

Schmidt:2018:QSG

[4487] Simon Schmidt and Moritz Weber. Quantum symmetries of graph C^{∗}- algebras. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 61(4):848–864, December 01, 2018. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic). URL https://www.

cambridge.org/core/journals/canadian-mathematical-bulletin/
article/quantum-symmetries-of-graph-c-algebras/8C58371D0DB1F70F8BCD0FC3A4C88C24.

Shaul:2018:HDL

[4488] Liran Shaul. Homological dimensions of local (co)homology over commutative DG-rings. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 61(4):865–877, December 01, 2018. CODEN CM-BUA3. ISSN 0008-4395 (print), 1496-4287 (electronic). URL https://

www.cambridge.org/core/journals/canadian-mathematical-bulletin/
article/homological-dimensions-of-local-cohomology-over-commutative-
dgrings/FAE2D1E3D0C9EE06236A63BBFB25EFBE.

Sun:2018:WAP

[4489] Chia-Liang Sun. Weak approximation for points with coordinates in rank-one subgroups of global function fields. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 61(4):878–890, December 01, 2018. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-

4287 (electronic). URL https://www.cambridge.org/core/journals/
canadian-mathematical-bulletin/article/weak-approximation-for-
points-with-coordinates-in-rankone-subgroups-of-global-function-
fields/E78A64BFD1DDCC49228D978475DD8A97.

AlAlam:2019:EMS

[4490] Ihab Al Alam and Pascal Lefèvre. Embeddings of Müntz spaces in L^{∞}(μ). Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 62(1):1–9, March 2019. CODEN CMBUA3. ISSN 0008-4395 (print),

1496-4287 (electronic). URL https://www.cambridge.org/core/
journals/canadian-mathematical-bulletin/article/embeddings-
of-muntz-spaces-in-linfty-unicodestixx1d707/BD75940FE51DB19961B9C4F3693FD90A.

Bertapelle:2019:DM

[4491] A. Bertapelle and N. Mazzari. On deformations of 1-motives. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 62(1): 11–22, March 2019. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-

4287 (electronic). URL https://www.cambridge.org/core/journals/
canadian-mathematical-bulletin/article/on-deformations-of-1motives/
38AA635C8C32C15270824D0885AA848F.

vanBatenburg:2019:SCN

[4492] Wouter Cames van Batenburg and Ross J. Kang. Squared chromatic number without claws or large cliques. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 62(1):23–35, March 2019. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-

4287 (electronic). URL https://www.cambridge.org/core/journals/
canadian-mathematical-bulletin/article/squared-chromatic-number-
without-claws-or-large-cliques/C70B18104C1B3D5F6AD0300847F9BA88.

Criswell:2019:PBM

[4493] Jackson A. Criswell, Ben Salisbury, and Peter Tingley. PBW bases and marginally large tableaux in types B and C. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 62(1):37– 54, March 2019. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-

4287 (electronic). URL https://www.cambridge.org/core/journals/
canadian-mathematical-bulletin/article/pbw-bases-and-marginally-
large-tableaux-in-types-b-and-c/0C6CDF3D5ACAB8A90DCF351A2567DB01.

DeCarli:2019:TPE

[4494] Laura De Carli, Alberto Mizrahi, and Alexander Tepper. Three problems on exponential bases. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 62(1):55–70, March 2019. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic). URL https://www.

cambridge.org/core/journals/canadian-mathematical-bulletin/
article/three-problems-on-exponential-bases/9296D5C7E9713B624A3E71816BA73D06.

Figiel:2019:QEE

[4495] Tadeusz Figiel and William Johnson. Quotients of essentially Euclidean spaces. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 62(1):71–74, March 2019. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic). URL https://www.

cambridge.org/core/journals/canadian-mathematical-bulletin/
article/quotients-of-essentially-euclidean-spaces/7F4D213443FC17D2CFCA995C37F4B92C.

Gaal:2019:MQS

[4496] Marcell Gaál. Maps on quantum states in C^{∗}-algebras preserving von Neumann entropy or Schatten p-norm of convex combinations. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 62(1):75–80, March 2019. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic). URL https://www.

cambridge.org/core/journals/canadian-mathematical-bulletin/
article/maps-on-quantum-states-in-cast-algebras-preserving-

von-neumann-entropy-or-schatten-p-norm-of-convex-combinations/
AEF6F13CE2B7FE13E409E703FFE100FE.

Gu:2019:FTA

[4497] Miao Gu and Greg Martin. Factorization tests and algorithms arising from counting modular forms and automorphic representations. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 62(1): 81–97, March 2019. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-

4287 (electronic). URL https://www.cambridge.org/core/journals/
canadian-mathematical-bulletin/article/factorization-tests-
and-algorithms-arising-from-counting-modular-forms-and-automorphic-
representations/A369C374D63481F0DC91AF446E5C76C3.

Kocel-Cynk:2019:PMS

[4498] Beata Kocel-Cynk, Wiesław Pawłucki, and Anna Valette. C^{√}- parametrization in O-minimal structures. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 62(1):99–108, March 2019. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287

(electronic).  URL https://www.cambridge.org/core/journals/
canadian-mathematical-bulletin/article/mathscrcp-parametrization-
in-ominimal-structures/54B2E072C3EA5BABC79B2CA52E6BB132.

Lee:2019:IWO

[4499] Hao Lee. Irregular weight one points with D_{4} image. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 62(1):109– 118, March 2019. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-

4287 (electronic). URL https://www.cambridge.org/core/journals/
canadian-mathematical-bulletin/article/irregular-weight-one-
points-with-d4-image/33657787CAA43CF63C9AC947BE5F163F.

Liu:2019:FWP

[4500] Huaifu Liu and Xiaohuan Mo. Finsler warped product metrics of Douglas type. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 62(1):119–130, March 2019. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic). URL https://www.

cambridge.org/core/journals/canadian-mathematical-bulletin/
article/finsler-warped-product-metrics-of-douglas-type/EB29E76344C4613D7AD71B584AE99559.

Liu:2019:ICH

[4501] Zhichao Liu. Injectivity of the connecting homomorphisms in inductive limits of Elliott–Thomsen algebras. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 62(1):131–148, March 2019. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-

4287 (electronic). URL https://www.cambridge.org/core/journals/
canadian-mathematical-bulletin/article/injectivity-of-the-connecting-
homomorphisms-in-inductive-limits-of-elliottthomsen-algebras/
D9DEE36FAC7B77E2294364400BB1DA30.

Meisner:2019:OLD

[4502] Patrick Meisner. One level density for cubic Galois number fields. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 62 (1):149–167, March 2019. CODEN CMBUA3. ISSN 0008-4395 (print),

1496-4287 (electronic). URL https://www.cambridge.org/core/
journals/canadian-mathematical-bulletin/article/one-level-density-
for-cubic-galois-number-fields/F6E3B717ACE76A3562E2EF157A646789.

Nguyen:2019:HJP

[4503] Khoa Dang Nguyen. The Hermite–Joubert problem and a conjecture of Brassil and Reichstein. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 62(1):169–177, March 2019. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic). URL https://www.

cambridge.org/core/journals/canadian-mathematical-bulletin/
article/hermitejoubert-problem-and-a-conjecture-of-brassil-
and-reichstein/8C414A6AB8ABDED8B00D940DF2FDE9E3.

Reichstein:2019:PRP

[4504] Zinovy B. Reichstein. On a property of real plane curves of even degree. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 62(1):179–182, March 2019. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic). URL https://www.

cambridge.org/core/journals/canadian-mathematical-bulletin/
article/on-a-property-of-real-plane-curves-of-even-degree/4C221D17B3568E10AF675DFBF1AFB0D3.

Rosales:2019:GHB

[4505] Leobardo Rosales. Generalizing Hopf’s boundary point lemma. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 62(1): 183–197, March 2019. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-

4287 (electronic). URL https://www.cambridge.org/core/journals/
canadian-mathematical-bulletin/article/generalizing-hopfs-boundary-
point-lemma/1873D3B9D8B51A24FBFD2B62A81551E9.

vanGool:2019:MDT

[4506] Samuel J. van Gool and Benjamin Steinberg. Merge decompositions, two-sided Krohn–Rhodes, and aperiodic pointlikes. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 62 (1):199–208, March 2019. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic). URL https://www.cambridge.org/

core/journals/canadian-mathematical-bulletin/article/merge-
decompositions-twosided-krohnrhodes-and-aperiodic-pointlikes/
FA9203781E0AF185215A7F418448689D.

Zhang:2019:MRC

[4507] Zheng Zhang. On Motivic realizations of the canonical Hermitian variations of Hodge structure of Calabi–Yau type over type D^{H} domains. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 62(1):209–221, March 2019. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic). URL https://www.

cambridge.org/core/journals/canadian-mathematical-bulletin/
article/on-motivic-realizations-of-the-canonical-hermitian-
variations-of-hodge-structure-of-calabiyau-type-over-type-dmathbbh-
domains/9EDE810FA415290612E512B55B0A3C18.

Anonymous:2019:BVIa

[4508] Anonymous. BCM volume 62 issue 1 cover and front matter. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 62(1): f1–f2, March 2019. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-

4287 (electronic). URL https://www.cambridge.org/core/journals/
canadian-mathematical-bulletin/article/bcm-volume-62-issue-
1-cover-and-front-matter/7BF842EFBC413EBD6CFC4BE57055041D.

Anonymous:2019:BVIb

[4509] Anonymous. BCM volume 62 issue 1 cover and back matter. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 62(1): b1–b2, March 2019. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-

4287 (electronic). URL https://www.cambridge.org/core/journals/
canadian-mathematical-bulletin/article/bcm-volume-62-issue-
1-cover-and-back-matter/3EBDC4D137F13790757015F02F8ED4A7.

Asgarli:2019:SBT

[4510] Shamil Asgarli. Sharp Bertini theorem for plane curves over finite fields. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 62(2):223–230, June 2019. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic). URL https://www.

cambridge.org/core/journals/canadian-mathematical-bulletin/
article/sharp-bertini-theorem-for-plane-curves-over-finite-
fields/EF8EE748FE1806D0079AAFE3845A5833.

Budd:2019:NCD

[4511] Samuel Budd and Adam Van Tuyl. Newton complementary duals of f-ideals. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 62(2):231–241, June 2019. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic). URL https://www.

cambridge.org/core/journals/canadian-mathematical-bulletin/
article/newton-complementary-duals-of-f-ideals/6B5B236B1096519B3C69E39C3E66CA73.

Cook:2019:DMS

[4512] Brian Cook. Discrete multilinear spherical averages. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 62(2):243– 246, June 2019. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-

4287 (electronic). URL https://www.cambridge.org/core/journals/
canadian-mathematical-bulletin/article/discrete-multilinear-
spherical-averages/E1C3F16CD684E674E33DCF67F37EB119.

Elmadani:2019:CDS

[4513] Y. Elmadani and I. Labghail. Cyclicity in Dirichlet spaces. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 62(2): 247–257, June 2019. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-

4287 (electronic). URL https://www.cambridge.org/core/journals/
canadian-mathematical-bulletin/article/cyclicity-in-dirichlet-
spaces/E2A1B5DC06E742E1FA8B7C41C0105282.

Gigli:2019:NAS

[4514] Nicola Gigli and Chiara Rigoni. A note about the strong maximum principle on RCD spaces. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 62(2):259–266, June 2019. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic). URL https://www.

cambridge.org/core/journals/canadian-mathematical-bulletin/
article/note-about-the-strong-maximum-principle-on-rcd-spaces/
A6BBDD6FD4EA98897F8AEAB03649D89A.

Gonzalez:2019:ENF

[4515] José Luis González and Kalle Karu. Examples of non-finitely generated Cox rings. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 62(2):267–285, June 2019. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic). URL https://www.

cambridge.org/core/journals/canadian-mathematical-bulletin/
article/examples-of-nonfinitely-generated-cox-rings/B016185FADCBC51F1DF1459C4F18C785.

Haslhofer:2019:LCS

[4516] Robert Haslhofer and Mohammad N. Ivaki. Low complexity solutions of the Allen–Cahn equation on three-spheres. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 62(2):287– 291, June 2019. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-

4287 (electronic). URL https://www.cambridge.org/core/journals/
canadian-mathematical-bulletin/article/low-complexity-solutions-
of-the-allencahn-equation-on-threespheres/E2F7AB64649FC10E30A01717325E5437.

Hassanzadeh:2019:MCT

[4517] Mohammad Hassanzadeh, Masoud Khalkhali, and Ilya Shapiro. Monoidal categories, 2-traces, and cyclic cohomology. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 62(2):293– 312, June 2019. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-

4287 (electronic). URL https://www.cambridge.org/core/journals/
canadian-mathematical-bulletin/article/monoidal-categories-
2traces-and-cyclic-cohomology/7AAA3B9C18DA63798A339C59B9D8D5B7.

Hou:2019:LWT

[4518] Xianming Hou and Huoxiong Wu. On the limiting weak-type behaviors for maximal operators associated with power weighted measure. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 62(2):313–326, June 2019. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic). URL https://www.

cambridge.org/core/journals/canadian-mathematical-bulletin/
article/on-the-limiting-weaktype-behaviors-for-maximal-operators-
associated-with-power-weighted-measure/20D686BCC3E11B167A717EE542DACE91.

Jiang:2019:CSC

[4519] Cao Jiang, Xing-Tang Dong, and Ze-Hua Zhou. Commuting and semicommuting monomial-type Toeplitz operators on some weakly pseudoconvex domains. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 62(2):327–340, June 2019. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic). URL https://www.

cambridge.org/core/journals/canadian-mathematical-bulletin/
article/commuting-and-semicommuting-monomialtype-toeplitz-operators-
on-some-weakly-pseudoconvex-domains/569A7B792AFE5A9C544AA99787B34A84.

Kalvin:2019:DLT

[4520] Victor Kalvin and Alexey Kokotov. Determinant of the Laplacian on tori of constant positive curvature with one conical point. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 62(2):341– 347, June 2019. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-

4287 (electronic). URL https://www.cambridge.org/core/journals/
canadian-mathematical-bulletin/article/determinant-of-the-laplacian-
on-tori-of-constant-positive-curvature-with-one-conical-point/
234370997526606FAFA4CDE3D688E037.

Cagliero:2019:JCD

[4521] Leandro Cagliero and Fernando Szechtman. Jordan-Chevalley decomposition in Lie algebras. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 62(2):349–354, June 2019. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic). URL https://www.

cambridge.org/core/journals/canadian-mathematical-bulletin/
article/jordanchevalley-decomposition-in-lie-algebras/CCF61625B61266406372E0D9376EBB5D.

Liu:2019:RBO

[4522] Ling Liu, Abdenacer Makhlouf, Claudia Menini, and Florin Panaite. {σ, τ} Rota–Baxter operators, infinitesimal Hom-bialgebras and the associative (bi)Hom–Yang–Baxter equation. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 62(2):355–372, June 2019. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-

4287 (electronic). URL https://www.cambridge.org/core/journals/
canadian-mathematical-bulletin/article/unicodestixx1d70eunicodestixx1d70f-
rotabaxter-operators-infinitesimal-hombialgebras-and-the-associative-
bihomyangbaxter-equation/21818E4A2BB997EE76379AB5317270F1.

Loring:2019:SFA

[4523] Terry A. Loring and Hermann Schulz-Baldes. Spectral flow argument localizing an odd index pairing. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 62(2):373–381, June 2019. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287

(electronic).  URL https://www.cambridge.org/core/journals/
canadian-mathematical-bulletin/article/spectral-flow-argument-
localizing-an-odd-index-pairing/B40129BE59FE9C1E2209C83AD7449EA6.

Maeda:2019:GCM

[4524] Sadahiro Maeda, Hiromasa Tanabe, and Seiichi Udagawa. Generating curves of minimal ruled real hypersurfaces in a nonflat complex space form. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 62(2):383–392, June 2019. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic). URL https://www.

cambridge.org/core/journals/canadian-mathematical-bulletin/
article/generating-curves-of-minimal-ruled-real-hypersurfaces-
in-a-nonflat-complex-space-form/D167239E3918D892FC4B4099C8C07AE1.

Nillsen:2019:VFT

[4525] Rodney Nillsen. Vanishing Fourier transforms and generalized differences in L^{2}(R). Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 62(2):393–404, June 2019. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic). URL https://www.

cambridge.org/core/journals/canadian-mathematical-bulletin/
article/vanishing-fourier-transforms-and-generalized-differences-
in-l2mathbbr/7C09A7F24B8261E9043446A870A2BAAA.

Oberlin:2019:SBP

[4526] Richard Oberlin. Sparse bounds for a prototypical singular Radon transform. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 62(2):405–415, June 2019. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic). URL https://www.

cambridge.org/core/journals/canadian-mathematical-bulletin/
article/sparse-bounds-for-a-prototypical-singular-radon-transform/
90244569C82FD03F37985923C73F14F1.

Pascoe:2019:WET

[4527] J. E. Pascoe. The wedge-of-the-edge theorem: Edge-of-the-wedge type phenomenon within the common real boundary. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 62(2):417– 427, June 2019. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-

4287 (electronic). URL https://www.cambridge.org/core/journals/
canadian-mathematical-bulletin/article/wedgeoftheedge-theorem-
edgeofthewedge-type-phenomenon-within-the-common-real-boundary/
5B2DE1CA809F4C33A660504DA9BBF9E1.

Porritt:2019:IPF

[4528] Sam Porritt. Irreducible polynomials over a finite field with restricted coefficients. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 62(2):429–439, June 2019. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic). URL https://www.

cambridge.org/core/journals/canadian-mathematical-bulletin/
article/irreducible-polynomials-over-a-finite-field-with-restricted-
coefficients/6148551FD7FFC9804C6AC0DB711D30C4.

vanMill:2019:ASG

[4529] Jan van Mill and Vesko M. Valov. Actions of semitopological groups. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 62 (2):441–450, June 2019. CODEN CMBUA3. ISSN 0008-4395 (print),

1496-4287 (electronic). URL https://www.cambridge.org/core/

journals/canadian-mathematical-bulletin/article/actions-of-
semitopological-groups/F5D31E23D7C08AE1453B300418276FED.

Collari:2019:SUD

[4530] Carlo Collari and Paolo Lisca. Symmetric union diagrams and refined spin models. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 62(3):451–468, September 2019. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic). URL https://www.

cambridge.org/core/journals/canadian-mathematical-bulletin/
article/symmetric-union-diagrams-and-refined-spin-models/62FED6AC02BE78DDEF6F8DA17959463D.

DeKoninck:2019:CIC

[4531] Jean-Marie De Koninck and Florian Luca. Consecutive integers with close kernels. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 62(3):469–473, September 2019. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic). URL https://www.

cambridge.org/core/journals/canadian-mathematical-bulletin/
article/consecutive-integers-with-close-kernels/95143DE6D827240FAF167677B6205C6C.

Gauthier:2019:ARZ

[4532] P. M. Gauthier. Approximating the Riemann zeta-function by polynomials with restricted zeros. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 62(3):475–478, September 2019. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-

4287 (electronic). URL https://www.cambridge.org/core/journals/
canadian-mathematical-bulletin/article/approximating-the-riemann-
zetafunction-by-polynomials-with-restricted-zeros/2A57330CEBA251AEAD8CF8405F83A961.

Guo:2019:QPE

[4533] Ji Guo. The quotient problem for entire functions. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 62(3):479–489, September 2019. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-

4287 (electronic). URL https://www.cambridge.org/core/journals/
canadian-mathematical-bulletin/article/quotient-problem-for-
entire-functions/E3A00172A87EA931699895A3AF4B2E3D.

Gutierrez:2019:MCI

[4534] Armando W. Gutiérrez. On the metric compactification of infinitedimensional ℓ_{p} spaces. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 62(3):491–507, September 2019. CODEN CM-BUA3. ISSN 0008-4395 (print), 1496-4287 (electronic). URL https://

www.cambridge.org/core/journals/canadian-mathematical-bulletin/

article/on-the-metric-compactification-of-infinitedimensional-
ell-p-spaces/BA73E83FEC3E2E7BD7215211C5005EC3.

Huang:2019:HEF

[4535] Libing Huang and Xiaohuan Mo. Homogeneous Einstein Finsler metrics on (4n + 3)-dimensional spheres. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 62(3):509–523, September 2019. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-

4287 (electronic). URL https://www.cambridge.org/core/journals/
canadian-mathematical-bulletin/article/homogeneous-einstein-
finsler-metrics-on-4n3-dimensional-spheres/6F6147A576EAA6B6751A1555D3B767D4.

Huang:2019:HEM

[4536] Libing Huang and Zhongmin Shen. Homogeneous Einstein manifolds with vanishing S curvature. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 62(3):525–537, September 2019. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-

4287 (electronic). URL https://www.cambridge.org/core/journals/
canadian-mathematical-bulletin/article/homogeneous-einstein-
manifolds-with-vanishing-s-curvature/187BB486D9DD60A3AC43B4475C748EEF.

Ismailov:2019:DLA

[4537] Nurlan Ismailov, Ivan Kaygorodov, and Yury Volkov. Degenerations of Leibniz and anticommutative algebras. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 62(3):539–549, September 2019. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-

4287 (electronic). URL https://www.cambridge.org/core/journals/
canadian-mathematical-bulletin/article/degenerations-of-leibniz-
and-anticommutative-algebras/460DC8581F4222FC46B24AEF3C6452D7.

Jahnel:2019:FAB

[4538] Jörg Jahnel and Damaris Schindler. On the frequency of algebraic Brauer classes on certain log K3 surfaces. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 62(3):551–563, September 2019. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-

4287 (electronic). URL https://www.cambridge.org/core/journals/
canadian-mathematical-bulletin/article/on-the-frequency-of-
algebraic-brauer-classes-on-certain-log-k3-surfaces/BA9934965B4546F035E5264B0E1AE7FF.

Mackall:2019:CFK

[4539] Eoin Mackall. The Coniveau filtration on K_{1} for some Severi–Brauer varieties. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 62(3):565–576, September 2019. CODEN CMBUA3.

ISSN 0008-4395 (print), 1496-4287 (electronic). URL https://www.

cambridge.org/core/journals/canadian-mathematical-bulletin/
article/coniveau-filtration-on-textk1-for-some-severibrauer-
varieties/587B7E0A05677E887E2E9F6AD5EE5519.

Moore:2019:NIM

[4540] Justin Tatch Moore. Nonexistence of idempotent means on free binary systems. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 62(3):577–581, September 2019. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic). URL https://www.

cambridge.org/core/journals/canadian-mathematical-bulletin/
article/nonexistence-of-idempotent-means-on-free-binary-systems/
E73B01B1EFAB52DED76C7EACC2BBAD8B.

Muhlherr:2019:TT

[4541] Bernhard Mühlherr and Richard M. Weiss. Tits triangles. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 62(3):583– 601, September 2019. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-

4287 (electronic). URL https://www.cambridge.org/core/journals/
canadian-mathematical-bulletin/article/tits-triangles/FD121DD3A40CF22D5A0434E6441E123A.

Nicholson:2019:TBS

[4542] W. K. Nicholson. On a theorem of Burgess and Stephenson. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 62(3): 603–605, September 2019. CODEN CMBUA3. ISSN 0008-4395 (print),

1496-4287 (electronic). URL https://www.cambridge.org/core/
journals/canadian-mathematical-bulletin/article/on-a-theorem-
of-burgess-and-stephenson/8588219E307E87C3E116B0B143E5858F.

Rosales:2019:HTB

[4543] Leobardo Rosales. A Hopf-type boundary point lemma for pairs of solutions to quasilinear equations. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 62(3):607–621, September 2019. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-

4287 (electronic). URL https://www.cambridge.org/core/journals/
canadian-mathematical-bulletin/article/hopftype-boundary-point-
lemma-for-pairs-of-solutions-to-quasilinear-equations/58DF9CA7DCF313C7E4B4EE8E22632936.

Shen:2019:HVT

[4544] Bin Shen. Holomorphic vanishing theorems on Finsler holomorphic vector bundles and complex Finsler manifolds. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 62(3):623–641, September 2019. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-

4287 (electronic). URL https://www.cambridge.org/core/journals/
canadian-mathematical-bulletin/article/holomorphic-vanishing-
theorems-on-finsler-holomorphic-vector-bundles-and-complex-
finsler-manifolds/E7F7C8AC15AEBE4D174CB54F2FFD25CD.

Steinerberger:2019:EAB

[4545] Stefan Steinerberger. An endpoint Alexandrov Bakelman Pucci estimate in the plane. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 62(3):643–651, September 2019. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic). URL https://www.

cambridge.org/core/journals/canadian-mathematical-bulletin/
article/an-endpoint-alexandrov-bakelman-pucci-estimate-in-the-
plane/1FC755A2D95C63836736F94BEC234415.

Suh:2019:YSR

[4546] Young Jin Suh and Uday Chand De. Yamabe solitons and Ricci solitons on almost co-Kähler manifolds. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 62(3):653–661, September 2019. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-

4287 (electronic). URL https://www.cambridge.org/core/journals/
canadian-mathematical-bulletin/article/yamabe-solitons-and-
ricci-solitons-on-almost-cokahler-manifolds/BFDB36B5512FC7BE41BD36C6DEF37F8C.

Troitsky:2019:KFC

[4547] V. G. Troitsky and M. S. Türer. Krivine’s function calculus and Bochner integration. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 62(3):663–669, September 2019. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic). URL https://www.

cambridge.org/core/journals/canadian-mathematical-bulletin/
article/krivines-function-calculus-and-bochner-integration/
F109A463BDA23FA91D5B4E7C59893BF6.

Wang:2019:EOB

[4548] Jianfei Wang and Danli Zhang. Extension operators for biholomorphic mappings. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 62(3):671–679, September 2019. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic). URL https://www.

cambridge.org/core/journals/canadian-mathematical-bulletin/
article/extension-operators-for-biholomorphic-mappings/72C5A5D3B527B30AC623E32CD12E2E18.

Xie:2019:PCE

[4549] Guangheng Xie, Dachun Yang, and Wen Yuan. Pointwise characterizations of even order Sobolev spaces via derivatives of ball averages. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 62(3):681–699, September 2019. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic). URL https://www.

cambridge.org/core/journals/canadian-mathematical-bulletin/
article/pointwise-characterizations-of-even-order-sobolev-spaces-
via-derivatives-of-ball-averages/CCEB3E16AE660B9D6A236051CFAF7D3B.

Banerjee:2019:DTT

[4550] Abhishek Banerjee. On differential torsion theories and rings with several objects. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 62(4):703–714, December 2019. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic). URL https://www.

cambridge.org/core/journals/canadian-mathematical-bulletin/
article/on-differential-torsion-theories-and-rings-with-several-
objects/574D81145FAAAD6B7C5648253FA41833.

Bu:2019:WPT

[4551] Shangquan Bu and Gang Cai. Well-posedness of third order differential equations in Hölder continuous function spaces. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 62(4):715–726, December 2019. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-

4287 (electronic). URL https://www.cambridge.org/core/journals/
canadian-mathematical-bulletin/article/wellposedness-of-third-
order-differential-equations-in-holder-continuous-function-
spaces/1581A259CE2551CDE96C55BC825E6672.

Deng:2019:TPC

[4552] Guotai Deng, Chuntai Liu, and Sze-Man Ngai. Topological properties of a class of higher-dimensional self-affine tiles. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 62(4):727–740, December 2019. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-

4287 (electronic). URL https://www.cambridge.org/core/journals/
canadian-mathematical-bulletin/article/topological-properties-
of-a-class-of-higherdimensional-selfaffine-tiles/33177BC0B27017278A5509D7D7E3576F.

Deza:2019:DDM

[4553] Antoine Deza and Lionel Pournin. Diameter, decomposability, and Minkowski sums of polytopes. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 62(4):741–755, December 2019. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-

4287 (electronic). URL https://www.cambridge.org/core/journals/
canadian-mathematical-bulletin/article/diameter-decomposability-
and-minkowski-sums-of-polytopes/2FAF87EA12CE35ABC5294CDAC9CB7999.

Farhat:2019:SCH

[4554] Yasser Farhat and Frédéric Gourdeau. Simplicial (co)-homology of ℓ^{1}(Z_{+}). Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 62(4):756–766, December 2019. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic). URL https://www.

cambridge.org/core/journals/canadian-mathematical-bulletin/
article/simplicial-cohomology-of-ell-1mathbbz/2DCC062BA7496742C3EE072C5B87EA95.

Gauthier:2019:UEF

[4555] P. M. Gauthier. Universal entire functions that define order isomorphisms of countable real sets. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 62(4):767–779, December 2019. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-

4287 (electronic). URL https://www.cambridge.org/core/journals/
canadian-mathematical-bulletin/article/universal-entire-functions-
that-define-order-isomorphisms-of-countable-real-sets/62EA366DE28DAECB30DCB76D0B231A61.

Geng:2019:PAO

[4556] Pengbo Geng, Wengu Chen, and Huanmin Ge. Perturbation analysis of orthogonal least squares. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 62(4):780–797, December 2019. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-

4287 (electronic). URL https://www.cambridge.org/core/journals/
canadian-mathematical-bulletin/article/perturbation-analysis-
of-orthogonal-least-squares/083C47B31275DB909B3039859592C26F.

Hare:2019:SSP

[4557] Kathryn E. Hare and Robert (Xu) Yang. Sidon sets are proportionally Sidon with small Sidon constants. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 62(4):798– 809, December 2019. CODEN CMBUA3. ISSN 0008-4395 (print),

1496-4287 (electronic). URL https://www.cambridge.org/core/
journals/canadian-mathematical-bulletin/article/sidon-sets-
are-proportionally-sidon-with-small-sidon-constants/775CC970CB0CA7A2B98B1A007D2C8700.

Kosan:2019:RWE

[4558] M. Tamer Kosan, Tülay Yildirim, and Y. Zhou. Rings whose elements are the sum of a tripotent and an element from the Jacobson radical. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 62(4):810–821, December 2019. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic). URL https://www.cambridge.org/

core/journals/canadian-mathematical-bulletin/article/rings-

whose-elements-are-the-sum-of-a-tripotent-and-an-element-from-
the-jacobson-radical/CC48C39759F43BE94D68C51EF26DA11B.

Koca:2019:EME

[4559] Caner Koca and Mehdi Lejmi. Einstein–Maxwell equations on four-dimensional Lie algebras. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 62(4):822–840, December 2019. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-

4287 (electronic). URL https://www.cambridge.org/core/journals/
canadian-mathematical-bulletin/article/einsteinmaxwell-equations-
on-fourdimensional-lie-algebras/7CC8E42A01F8CD35F4D251B73F77DFB4.

Kuo:2019:STS

[4560] Wentang Kuo, Yu-Ru Liu, Sávio Ribas, and Kevin Zhou. The shifted Turán sieve method on tournaments. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 62(4):841–855, December 2019. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-

4287 (electronic). URL https://www.cambridge.org/core/journals/
canadian-mathematical-bulletin/article/shifted-turan-sieve-
method-on-tournaments/A87787C3261474455B665969E71E73CC.

Raghavan:2019:CPP

[4561] Dilip Raghavan and Jonathan L. Verner. Chains of P -points. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 62(4): 856–868, December 2019. CODEN CMBUA3. ISSN 0008-4395 (print),

1496-4287 (electronic). URL https://www.cambridge.org/core/
journals/canadian-mathematical-bulletin/article/chains-of-ppoints/
54AA30D01C39E5BA0A7A3201EEA79CD6.

Roy:2019:NAA

[4562] Damien Roy and Johannes Schleischitz. Numbers with almost all convergents in a Cantor set. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 62(4):869–875, December 2019. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-

4287 (electronic). URL https://www.cambridge.org/core/journals/
canadian-mathematical-bulletin/article/numbers-with-almost-
all-convergents-in-a-cantor-set/13651E945614063A1065FFBC4B3EE709.

Saez:2019:BPM

[4563] Pablo Sáez, Xavier Vidaux, and Maxim Vsemirnov. Büchi’s problem in modular arithmetic for arbitrary quadratic polynomials. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 62(4):876– 885, December 2019. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-

4287 (electronic). URL https://www.cambridge.org/core/journals/
canadian-mathematical-bulletin/article/buchis-problem-in-modular-
arithmetic-for-arbitrary-quadratic-polynomials/D8BAF1011C13AAFD71D1F51673099808.

Steinberg:2019:LC

[4564] Benjamin Steinberg. Linear conjugacy. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 62(4):886–895, December 2019. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-

4287 (electronic). URL https://www.cambridge.org/core/journals/
canadian-mathematical-bulletin/article/linear-conjugacy/D3758252819C82CA0AEC99C44A5B0A27.

Ueyama:2019:KPQ

[4565] Kenta Ueyama. On Knörrer periodicity for quadric hypersurfaces in skew projective spaces. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 62(4):896–911, December 2019. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-

4287 (electronic). URL https://www.cambridge.org/core/journals/
canadian-mathematical-bulletin/article/on-knorrer-periodicity-
for-quadric-hypersurfaces-in-skew-projective-spaces/90BDF8E9B0B6A8AE71247A49F9A97334.

Wang:2019:RSA

[4566] Yaning Wang. Ricci solitons on almost co-Kähler manifolds. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 62(4):912–922, December 2019. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic). URL https://www.cambridge.org/

core/journals/canadian-mathematical-bulletin/article/ricci-
solitons-on-almost-cokahler-manifolds/F5E33F1D175AC694751FAD5D14E8A8FB.

Yang:2019:MLT

[4567] Qi Yang and Chuanming Zong. Multiple lattice tilings in Euclidean spaces. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 62(4):923–929, December 2019. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic). URL https://www.

cambridge.org/core/journals/canadian-mathematical-bulletin/
article/multiple-lattice-tilings-in-euclidean-spaces/46E220E2E0B77E6BDAA0B788CB74619A.

Anonymous:2019:BVIc

[4568] Anonymous. BCM volume 62 issue 4 cover and front matter. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 62(4): f1–f2, December 2019. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-

4287 (electronic). URL https://www.cambridge.org/core/journals/
canadian-mathematical-bulletin/article/bcm-volume-62-issue-
4-cover-and-front-matter/18489AEF8424E3647D031429F91BB69A.

Anonymous:2019:BVId

[4569] Anonymous. BCM volume 62 issue 4 cover and back matter. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 62(4): b1–b2, December 2019. CODEN CMBUA3. ISSN 0008-4395 (print),

1496-4287 (electronic). URL https://www.cambridge.org/core/
journals/canadian-mathematical-bulletin/article/bcm-volume-
62-issue-4-cover-and-back-matter/6678347E22C7DD1FC8B1402A090DC7F8.

Ballico:2020:VTT

[4570] E. Ballico. A vanishing theorem for the twisted normal bundle of curves in Pn, n ≥ 8. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 63(1):1–5, March 2020. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic). URL https://www.

cambridge.org/core/journals/canadian-mathematical-bulletin/
article/vanishing-theorem-for-the-twisted-normal-bundle-of-
curves-in-mathbbpn-ngeqslant-8/0E3FD4D05C7D193C56AB1F7D8AC03559.

Bavula:2020:RDC

[4571] V. V. Bavula and V. Levandovskyy. A remark on the Dixmier Conjecture. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 63(1):6–12, March 2020. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic). URL https://www.

cambridge.org/core/journals/canadian-mathematical-bulletin/
article/remark-on-the-dixmier-conjecture/9B13849B2F981EF776E9D1F5098829A5.

Bramburger:2020:UPD

[4572] Jason J. Bramburger. Ultracontractive properties for directed graph semigroups with applications to coupled oscillators. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 63(1):13– 30, March 2020. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-

4287 (electronic). URL https://www.cambridge.org/core/journals/
canadian-mathematical-bulletin/article/ultracontractive-properties-
for-directed-graph-semigroups-with-applications-to-coupled-
oscillators/ACE2FD7B5A34558F0F987D555753B9C8.

Chatterjee:2020:LIL

[4573] Tapas Chatterjee and Sonika Dhillon. Linear independence of logarithms of cyclotomic numbers and a conjecture of Livingston. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 63(1): 31–45, March 2020. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-

4287 (electronic). URL https://www.cambridge.org/core/journals/
canadian-mathematical-bulletin/article/linear-independence-

of-logarithms-of-cyclotomic-numbers-and-a-conjecture-of-livingston/
0FF737B76ED7E0FAB2705D778E7BAAC8.

Colbois:2020:HPB

[4574] Bruno Colbois, Alexandre Girouard, and Antoine Métras. Hypersurfaces with prescribed boundary and small Steklov eigenvalues. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 63(1):46–57, March 2020. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic). URL https://www.

cambridge.org/core/journals/canadian-mathematical-bulletin/
article/hypersurfaces-with-prescribed-boundary-and-small-steklov-
eigenvalues/E72D408A4ADD20B4BF2C26C2680D70BC.

Dillery:2020:SBG

[4575] D. Scott Dillery and John D. LaGrange. Spectra of Boolean graphs over finite fields of characteristic two. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 63(1): 58–65, March 2020. CODEN CMBUA3. ISSN 0008-4395 (print),

1496-4287 (electronic). URL https://www.cambridge.org/core/
journals/canadian-mathematical-bulletin/article/spectra-of-
boolean-graphs-over-finite-fields-of-characteristic-two/904BAE1A61AA22E66F5A529716D49FC0.

Healy:2020:RPH

[4576] Brendan Burns Healy. Rigidity properties for hyperbolic generalizations. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 63(1):66–76, March 2020. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic). URL https://www.

cambridge.org/core/journals/canadian-mathematical-bulletin/
article/rigidity-properties-for-hyperbolic-generalizations/
79A88BACA5E5F5B9495A1AAE175D395D.

Lancien:2020:CGJ

[4577] Gilles Lancien, Colin Petitjean, and Antonin Procházka. On the coarse geometry of James spaces. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 63(1):77–93, March 2020. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287

(electronic).  URL https://www.cambridge.org/core/journals/
canadian-mathematical-bulletin/article/on-the-coarse-geometry-
of-james-spaces/DD7520659B74E6E2F1F2C644FD242702.

Li:2020:TMP

[4578] Yueyue Li, Yan Tian, and Xiankun Du. Triangularization of matrices and polynomial maps. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 63(1):94–105, March 2020. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic). URL https://www.

cambridge.org/core/journals/canadian-mathematical-bulletin/
article/triangularization-of-matrices-and-polynomial-maps/E99253D6BB66F6F0083028905F960FE5.

Li:2020:ETD

[4579] Songxiao Li, Junming Liu, and Cheng Yuan. Embedding theorems for Dirichlet type spaces. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 63(1):106–117, March 2020. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic). URL https://www.

cambridge.org/core/journals/canadian-mathematical-bulletin/
article/embedding-theorems-for-dirichlet-type-spaces/CA0E16B123056EF0C91DA2EC3A1A8212.

Lucic:2020:IHW

[4580] Danka Lucić and Enrico Pasqualetto. Infinitesimal hilbertianity of weighted Riemannian manifolds. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 63(1):118–140, March 2020. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287

(electronic).  URL https://www.cambridge.org/core/journals/
canadian-mathematical-bulletin/article/infinitesimal-hilbertianity-
of-weighted-riemannian-manifolds/B241174C2193831ACA6A974CD54D6DF0.

Saito:2020:BDW

[4581] Hiroki Saito, Hitoshi Tanaka, and Toshikazu Watanabe. Block decomposition and weighted Hausdorff content. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 63(1):141–156, March 2020. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-

4287 (electronic). URL https://www.cambridge.org/core/journals/
canadian-mathematical-bulletin/article/block-decomposition-
and-weighted-hausdorff-content/709C8E39124CCC7F513E6F68EA6CA215.

SanAntolin:2020:PWF

[4582] A. San Antolín. On Parseval wavelet frames via multiresolution analyses in H_{G}^{2}. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 63(1):157–172, March 2020. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic). URL https://www.

cambridge.org/core/journals/canadian-mathematical-bulletin/
article/on-parseval-wavelet-frames-via-multiresolution-analyses-
in-hg2/16680ADE970C7627B55DCE600B230C62.

Scavia:2020:RRA

[4583] Federico Scavia. Retract rationality and algebraic tori. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 63(1):173– 186, March 2020. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-

4287 (electronic). URL https://www.cambridge.org/core/journals/
canadian-mathematical-bulletin/article/retract-rationality-
and-algebraic-tori/34898D6DF9E60DDC07A095970F72D8E5.

Shparlinski:2020:VSS

[4584] Igor E. Shparlinski and José Felipe Voloch. Value sets of sparse polynomials. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 63(1):187–196, March 2020. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic). URL https://www.cambridge.org/

core/journals/canadian-mathematical-bulletin/article/value-
sets-of-sparse-polynomials/7476A220D33C8D618FCB285F1B880A5F.

Bella:2020:CEA

[4585] Angelo Bella and Santi Spadaro. A common extension of Arhangel’ski’s theorem and the Hajnal–Juhász inequality. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 63(1):197–203, March 2020. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-

4287 (electronic). URL https://www.cambridge.org/core/journals/
canadian-mathematical-bulletin/article/common-extension-of-
arhangelskis-theorem-and-the-hajnaljuhasz-inequality/3F10A549A20675EBB2A62CDE06015FBE.

Suh:2020:RHC

[4586] Young Jin Suh and Gyu Jong Kim. Real hypersurfaces in the complex quadric with Lie invariant structure Jacobi operator. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 63(1):204– 221, March 2020. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-

4287 (electronic). URL https://www.cambridge.org/core/journals/
canadian-mathematical-bulletin/article/real-hypersurfaces-in-
the-complex-quadric-with-lie-invariant-structure-jacobi-operator/
4F15D164AE5B0473925DB6A101456324.

Sun:2020:OBE

[4587] Hongyan Sun. Orlicz–Besov extension and imbedding. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 63(1):222– 241, March 2020. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-

4287 (electronic). URL https://www.cambridge.org/core/journals/
canadian-mathematical-bulletin/article/orliczbesov-extension-
and-imbedding/77EE3159A9D4BF4876BBB885C00B3D2F.

Zubelevic:2020:PSC

[4588] Oleg Zubelevic. On periodic solutions to constrained Lagrangian system. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 63(1):242–255, March 2020. CODEN CMBUA3.

ISSN 0008-4395 (print), 1496-4287 (electronic). URL https://www.

cambridge.org/core/journals/canadian-mathematical-bulletin/
article/on-periodic-solutions-to-constrained-lagrangian-system/
A87592A9B33DAA4FE2474593DAFC3F28.

Anonymous:2020:BVIa

[4589] Anonymous. BCM volume 63 issue 1 cover and front matter. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 63(1): f1–f2, March 2020. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-

4287 (electronic). URL https://www.cambridge.org/core/journals/
canadian-mathematical-bulletin/article/bcm-volume-63-issue-
1-cover-and-front-matter/B495F57CF56C49DED6E1D73293487041.

Frej:2020:CET

[4590] Bartosz Frej and Dawid Huczek. A comment on ergodic theorem for amenable groups. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 63(2):257–261, June 2020. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic). URL https://www.

cambridge.org/core/journals/canadian-mathematical-bulletin/
article/comment-on-ergodic-theorem-for-amenable-groups/EF02EA6F75010AE046137A9DC833713A.

Mlaiki:2020:NFA

[4591] Nabil Mlaiki. New facts about the vanishing off subgroup V (G). Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 63(2): 262–268, June 2020. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-

4287 (electronic). URL https://www.cambridge.org/core/journals/
canadian-mathematical-bulletin/article/new-facts-about-the-
vanishing-off-subgroup-vg/D1D852ED2609FA43600DA497103133F1.

Ohrysko:2020:SRF

[4592] Przemysław Ohrysko and Maria Roginskaya. The spectral radius formula for Fourier–Stieltjes algebras. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 63(2):269–275, June 2020. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287

(electronic).  URL https://www.cambridge.org/core/journals/
canadian-mathematical-bulletin/article/spectral-radius-formula-
for-fourierstieltjes-algebras/D58332A72F3C44A307F9774F34B59420.

Naito:2020:OCS

[4593] Manabu Naito. Oscillation criteria for second order ordinary differential equations. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 63(2):276–286, June 2020. CODEN CMBUA3.

ISSN 0008-4395 (print), 1496-4287 (electronic). URL https://www.

cambridge.org/core/journals/canadian-mathematical-bulletin/
article/oscillation-criteria-for-second-order-ordinary-differential-
equations/3063E6EC5C65D09DF92B7CB2347BC3C9.

Ohno:2020:SIR

[4594] Takao Ohno and Tetsu Shimomura. Sobolev’s inequality for Riesz potentials of functions in Musielak–Orlicz–Morrey spaces over nondoubling metric measure spaces. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 63(2):287–303, June 2020. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287

(electronic).  URL https://www.cambridge.org/core/journals/
canadian-mathematical-bulletin/article/sobolevs-inequality-
for-riesz-potentials-of-functions-in-musielakorliczmorrey-spaces-
over-nondoubling-metric-measure-spaces/E5E411CD4D6F48A87CAA022F15D08D72.

Tan:2020:DPO

[4595] Jian Tan. Discrete paraproduct operators on variable Hardy spaces. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 63 (2):304–317, June 2020. CODEN CMBUA3. ISSN 0008-4395 (print),

1496-4287 (electronic). URL https://www.cambridge.org/core/
journals/canadian-mathematical-bulletin/article/discrete-paraproduct-
operators-on-variable-hardy-spaces/5994D83367C9BC64AA0F04B90690E23A.

Wang:2020:NSP

[4596] Ye Wang, Xin-Han Dong, and Yue-Ping Jiang. Non-spectral problem for some self-similar measures. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 63(2):318–327, June 2020. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287

(electronic).  URL https://www.cambridge.org/core/journals/
canadian-mathematical-bulletin/article/nonspectral-problem-
for-some-selfsimilar-measures/31BA989B435DA922DC65C6136964CE69.

Croome:2020:CCM

[4597] Sarah Croome and Mark L. Lewis. Character codegrees of maximal class p-groups. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 63(2):328–334, June 2020. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic). URL https://www.

cambridge.org/core/journals/canadian-mathematical-bulletin/
article/character-codegrees-of-maximal-class-p-groups/34294AD7A4C118B649A3F2348DEE4BFE.

Dicks:2020:OTF

[4598] Warren Dicks and Zoran Sunić. Orders on trees and free products of left-ordered groups. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 63(2):335–347, June 2020. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic). URL https://www.

cambridge.org/core/journals/canadian-mathematical-bulletin/
article/orders-on-trees-and-free-products-of-leftordered-groups/
6B495DF5B15873C2A1C73EDA28A514F8.

Cushman:2020:SSL

[4599] Richard Cushman and Jedrzej Śniatycki. On subcartesian spaces Leibniz’ rule implies the chain rule. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 63(2):348–357, June 2020. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287

(electronic).  URL https://www.cambridge.org/core/journals/
canadian-mathematical-bulletin/article/on-subcartesian-spaces-
leibniz-rule-implies-the-chain-rule/EF03598F02B950BF1132E1198EC38545.

Wang:2020:HPDa

[4600] Zhenjian Wang. On homogeneous polynomials determined by their partial derivatives. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 63(2):358–365, June 2020. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic). URL https://www.

cambridge.org/core/journals/canadian-mathematical-bulletin/
article/on-homogeneous-polynomials-determined-by-their-partial-
derivatives/8FF1A09FC92B9F51E6726742307C86BA. See correction

[4609].

Chen:2020:SCM

[4601] Ming-Liang Chen, Jing-Cheng Liu, Juan Su, and Xiang-Yang Wang. Spectrality of a class of Moran measures. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 63(2):366–381, June 2020. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-

4287 (electronic). URL https://www.cambridge.org/core/journals/
canadian-mathematical-bulletin/article/spectrality-of-a-class-
of-moran-measures/55C79F123EC665C60F5535476F0AFDDF.

Huang:2020:GPS

[4602] Keping Huang. A gap principle for subvarieties with finitely many periodic points. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 63(2):382–392, June 2020. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic). URL https://www.

cambridge.org/core/journals/canadian-mathematical-bulletin/
article/gap-principle-for-subvarieties-with-finitely-many-periodic-
points/261E826CA85075D13425C7E1C520A85D.

Clay:2020:DOS

[4603] Adam Clay and Tessa Reimer. Dense orderings in the space of leftorderings of a group. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 63(2):393–404, June 2020. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic). URL https://www.

cambridge.org/core/journals/canadian-mathematical-bulletin/
article/dense-orderings-in-the-space-of-leftorderings-of-a-
group/8D64A2CA505B6E946299CB27483972A2.

Huang:2020:SAP

[4604] Chuangxia Huang, Xin Long, Lihong Huang, and Si Fu. Stability of almost periodic Nicholson’s blowflies model involving patch structure and mortality terms. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 63(2):405–422, June 2020. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic). URL https://www.

cambridge.org/core/journals/canadian-mathematical-bulletin/
article/stability-of-almost-periodic-nicholsons-blowflies-model-
involving-patch-structure-and-mortality-terms/98933E35B7CC1780266AFD1EB5BA67DA.

Belkale:2020:TPP

[4605] P. Belkale and N. Fakhruddin. Triviality properties of principal bundles on singular curves. II. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 63(2):423–433, June 2020. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic). URL https://www.

cambridge.org/core/journals/canadian-mathematical-bulletin/
article/triviality-properties-of-principal-bundles-on-singular-
curves-ii/3CE2ACCBACB11AB0BF68D910CD30E0BA.

Troscheit:2020:AST

[4606] Sascha Troscheit. Assouad spectrum thresholds for some random constructions. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 63(2):434–453, June 2020. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic). URL https://www.

cambridge.org/core/journals/canadian-mathematical-bulletin/
article/assouad-spectrum-thresholds-for-some-random-constructions/
FA2A8D9A1EE3BC6FFE10337257524508.

Ito:2020:STP

[4607] Yohei Ito and Kiyoshi Takeuchi. On some topological properties of Fourier transforms of regular holonomic D-modules. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 63(2):454– 468, June 2020. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-

4287 (electronic). URL https://www.cambridge.org/core/journals/
canadian-mathematical-bulletin/article/on-some-topological-
properties-of-fourier-transforms-of-regular-holonomic-mathcald-
modules/99AC9C4AB14E8D32696B46308EA30457.

Murugan:2020:MCU

[4608] S. P. Murugan and S. Sundar. Measurable and continuous units of an E_{0}-semigroup. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 63(2):469–478, June 2020. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic). URL https://www.

cambridge.org/core/journals/canadian-mathematical-bulletin/
article/measurable-and-continuous-units-of-an-e0-semigroup/
C174365AC67BF348C65D09E679FE9E10.

Wang:2020:HPDb

[4609] Zhenjian Wang. On homogeneous polynomials determined by their partial derivatives — Corrigendum. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 63(2):479, June 2020. CODEN CM-BUA3. ISSN 0008-4395 (print), 1496-4287 (electronic). URL https://

www.cambridge.org/core/journals/canadian-mathematical-bulletin/
article/on-homogeneous-polynomials-determined-by-their-partial-
derivatives-corrigendum/F191EFD55F544787D7673D89CFBEC6E7. See

[4600].

Louboutin:2020:DPA

[4610] Stéphane R. Louboutin. On the discriminants of the powers of an algebraic integer. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 63(3):481–483, September 2020. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic). URL https://www.

cambridge.org/core/journals/canadian-mathematical-bulletin/
article/on-the-discriminants-of-the-powers-of-an-algebraic-
integer/4809D5BCA8DB45D374B8DB7CB78A8027.

Badea:2020:ENA

[4611] Catalin Badea, Vincent Devinck, and Sophie Grivaux. Escaping a neighborhood along a prescribed sequence in Lie groups and Banach algebras. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 63(3):484–505, September 2020. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic). URL https://www.

cambridge.org/core/journals/canadian-mathematical-bulletin/
article/escaping-a-neighborhood-along-a-prescribed-sequence-
in-lie-groups-and-banach-algebras/40FD7184A2A96B1C9DC5A68EAA16565C.

Miller:2020:ERF

[4612] Chris Miller and Patrick Speissegger. Expansions of the real field by canonical products. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 63(3):506–521, September 2020. CODEN CM-BUA3. ISSN 0008-4395 (print), 1496-4287 (electronic). URL https://

www.cambridge.org/core/journals/canadian-mathematical-bulletin/
article/expansions-of-the-real-field-by-canonical-products/
80244AEA650C9D4E2EB7A2E0BDAC5BC3.

Maier:2020:CSR

[4613] Helmut Maier and Michael Th. Rassias. Cotangent sums related to the Riemann hypothesis for various shifts of the argument. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 63(3):522–535, September 2020. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic). URL https://www.

cambridge.org/core/journals/canadian-mathematical-bulletin/
article/cotangent-sums-related-to-the-riemann-hypothesis-for-
various-shifts-of-the-argument/8FCB735178F5389D7AB6385677B7711E.

Chalebgwa:2020:AVE

[4614] Taboka Prince Chalebgwa. Algebraic values of entire functions with extremal growth orders: An extension of a theorem of Boxall and Jones. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 63(3):536–546, September 2020. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic). URL https://www.

cambridge.org/core/journals/canadian-mathematical-bulletin/
article/algebraic-values-of-entire-functions-with-extremal-
growth-orders-an-extension-of-a-theorem-of-boxall-and-jones/
E95C5585E805719046D8780AC53BED3F.

Llibre:2020:LCP

[4615] Jaume Llibre, Paulina Martínez, and Claudio Vidal. Limit cycles of a perturbation of a polynomial Hamiltonian systems of degree 4 symmetric with respect to the origin. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 63(3): 547–561, September 2020. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic). URL https://www.cambridge.org/

core/journals/canadian-mathematical-bulletin/article/limit-
cycles-of-a-perturbation-of-a-polynomial-hamiltonian-systems-
of-degree-4-symmetric-with-respect-to-the-origin/A674A609A183CA3504C3F75C8CFA5B60.

Jardine:2020:SCL

[4616] J. F. Jardine. Stable components and layers. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 63(3):562–576, September 2020. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-

4287 (electronic). URL https://www.cambridge.org/core/journals/
canadian-mathematical-bulletin/article/stable-components-and-
layers/115EEF4BDE091E9C93D9473D0B4EEFA0.

Kazda:2020:DEM

[4617] Alexandr Kazda, Jakub Oprsal, Matt Valeriote, and Dmitriy Zhuk. Deciding the existence of minority terms. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 63(3):577–591, September 2020. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-

4287 (electronic). URL https://www.cambridge.org/core/journals/
canadian-mathematical-bulletin/article/deciding-the-existence-
of-minority-terms/9974D0C36388FF4BF835D3A8E0289B85.

Ito:2020:PRA

[4618] Atsushi Ito, Makoto Miura, and Kazushi Ueda. Projective reconstruction in algebraic vision. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 63(3):592–609, September 2020. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-

4287 (electronic). URL https://www.cambridge.org/core/journals/
canadian-mathematical-bulletin/article/projective-reconstruction-
in-algebraic-vision/26FB870678ABDDCBC7120A248C162A6E.

Leiderman:2020:SQF

[4619] Arkady Leiderman and Mikhail Tkachenko. Separable quotients of free topological groups. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 63(3):610–623, September 2020. CODEN CM-BUA3. ISSN 0008-4395 (print), 1496-4287 (electronic). URL https://

www.cambridge.org/core/journals/canadian-mathematical-bulletin/
article/separable-quotients-of-free-topological-groups/C593B0295D5A6BA51CECF62593299BAA.

Alavinejad:2020:CSR

[4620] Mahnaz Alavinejad and Jianhong Wu. Coupled systems of renewal equations for forces of infection through a contact network. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 63(3):624–632, September 2020. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic). URL https://www.

cambridge.org/core/journals/canadian-mathematical-bulletin/
article/coupled-systems-of-renewal-equations-for-forces-of-
infection-through-a-contact-network/5014FFDAFF4E4351B71FA7EC8061A390.

Bao:2020:RCO

[4621] Guanlong Bao, Hasi Wulan, and Fangqin Ye. The range of the Cesàro operator acting on H^{∞}. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 63(3):633–642, September 2020. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-

4287 (electronic). URL https://www.cambridge.org/core/journals/
canadian-mathematical-bulletin/article/range-of-the-cesaro-
operator-acting-on-hinfty/BA558B6663F7622766CF46DD16F39C61.

Lin:2020:JNI

[4622] Haibo Lin, Zhen Liu, and Chenyan Wang. The John–Nirenberg inequality for the regularized BLO space on non-homogeneous metric measure spaces. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 63(3):643–654, September 2020. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic). URL https://www.

cambridge.org/core/journals/canadian-mathematical-bulletin/
article/johnnirenberg-inequality-for-the-regularized-blo-space-
on-nonhomogeneous-metric-measure-spaces/EF14F264D7D9D55537CC0EDA0703EAB5.

Pepin:2020:NMH

[4623] A. Pepin, S. S. Beauchemin, S. Léger, and N. Beaudoin. A new method for high-degree spline interpolation: Proof of continuity for piecewise polynomials. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 63(3):655–669, September 2020. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic). URL https://www.

cambridge.org/core/journals/canadian-mathematical-bulletin/
article/new-method-for-highdegree-spline-interpolation-proof-
of-continuity-for-piecewise-polynomials/0037D4376FC0FB10F80C74804776A121.

Awtrey:2020:GGE

[4624] Chad Awtrey and Peter Jakes. Galois groups of even sextic polynomials. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 63(3):670–676, September 2020. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic). URL https://www.

cambridge.org/core/journals/canadian-mathematical-bulletin/
article/galois-groups-of-even-sextic-polynomials/1D8CABFE7E25607425AADE8731B8EFB2.

Li:2020:EUS

[4625] Qingwei Li. Existence and uniqueness of solutions to singular p-Laplace equations of Kirchhoff type. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 63(3):677–691, September 2020. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-

4287 (electronic). URL https://www.cambridge.org/core/journals/
canadian-mathematical-bulletin/article/existence-and-uniqueness-
of-solutions-to-singular-plaplace-equations-of-kirchhoff-type/
902CAE6DEC1015975E0F8CBAB5C51754.

Salame:2020:AFP

[4626] Khadime Salame. Amenability and fixed point properties of semitopological semigroups in modular vector spaces. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 63(3):692–704, September 2020. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-

4287 (electronic). URL https://www.cambridge.org/core/journals/
canadian-mathematical-bulletin/article/amenability-and-fixed-
point-properties-of-semitopological-semigroups-in-modular-vector-
spaces/7DCB1403B2D348F9D5F59075B26CDFD3.

Anonymous:2020:BVIb

[4627] Anonymous. BCM volume 63 issue 3 cover and front matter. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 63(3): f1–f2, September 2020. CODEN CMBUA3. ISSN 0008-4395 (print),

1496-4287 (electronic). URL https://www.cambridge.org/core/
journals/canadian-mathematical-bulletin/article/bcm-volume-
63-issue-3-cover-and-front-matter/79E3F5C4958276A9387BEF2CC22AFFBE.

Anonymous:2020:BVIc

[4628] Anonymous. BCM volume 63 issue 3 cover and back matter. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 63(3): b1–b2, September 2020. CODEN CMBUA3. ISSN 0008-4395 (print),

1496-4287 (electronic). URL https://www.cambridge.org/core/
journals/canadian-mathematical-bulletin/article/bcm-volume-
63-issue-3-cover-and-back-matter/8C1FBD7C416AAC32ECA5BC739413DCDC.

Menegatto:2020:PDP

[4629] V. A. Menegatto and C. P. Oliveira. Positive definiteness on products of compact two-point homogeneous spaces and locally compact Abelian groups. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 63(4):705–715, December 2020. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic). URL https://www.

cambridge.org/core/journals/canadian-mathematical-bulletin/
article/positive-definiteness-on-products-of-compact-twopoint-
homogeneous-spaces-and-locally-compact-abelian-groups/434A6963A2D01EC0AEEAF4EF678F72BD.

Protas:2020:DBP

[4630] David Protas. Derivatives of Blaschke products and model space functions. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 63(4):716–725, December 2020. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic). URL https://www.

cambridge.org/core/journals/canadian-mathematical-bulletin/
article/derivatives-of-blaschke-products-and-model-space-functions/
A13BC7319F6C4485D89DE4AEAD3239D6.

Mayeli:2020:NCO

[4631] Azita Mayeli. Non-convex optimization via strongly convex majorizationminimization. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 63(4):726–737, December 2020. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic). URL https://www.

cambridge.org/core/journals/canadian-mathematical-bulletin/
article/nonconvex-optimization-via-strongly-convex-majorizationminimization/
6032A3F0EC23573D905B6FE4DDE86A03.

Tongsomporn:2020:IDP

[4632] Janyarak Tongsomporn and Jörn Steuding. Irregularities in the distribution of prime numbers in a Beatty sequence. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 63(4):738–743, December 2020. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-

4287 (electronic). URL https://www.cambridge.org/core/journals/
canadian-mathematical-bulletin/article/irregularities-in-the-
distribution-of-prime-numbers-in-a-beatty-sequence/67078D583A1C98EEDD680B37DB34B024.

Davis:2020:CCC

[4633] Christopher W. Davis. Concordance, crossing changes, and knots in homology spheres. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 63(4):744–754, December 2020. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic). URL https://www.

cambridge.org/core/journals/canadian-mathematical-bulletin/
article/concordance-crossing-changes-and-knots-in-homology-
spheres/06770C1D0E68D598F9381D565DB44294.

Saha:2020:OBE

[4634] Kuldeep Saha. On open book embedding of contact manifolds in the standard contact sphere. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 63(4):755–770, December 2020. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-

4287 (electronic). URL https://www.cambridge.org/core/journals/
canadian-mathematical-bulletin/article/on-open-book-embedding-
of-contact-manifolds-in-the-standard-contact-sphere/393261E1C3FC466D510A56B10D47D2D2.

He:2020:SBO

[4635] Danqing He and Zuoshunhua Shi. Sharp bounds for oscillatory integral operators with homogeneous polynomial phases. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 63(4): 771–786, December 2020. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic). URL https://www.cambridge.org/

core/journals/canadian-mathematical-bulletin/article/sharp-
bounds-for-oscillatory-integral-operators-with-homogeneous-
polynomial-phases/2BE8A9E21DBE29E4E3A0503B31A5D704.

Linden:2020:SCF

[4636] Christopher Linden. Slow continued fractions and permutative representations of O_{N}. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 63(4):787–801, December 2020. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic). URL https://www.

cambridge.org/core/journals/canadian-mathematical-bulletin/
article/slow-continued-fractions-and-permutative-representations-
of-mathcalon/873CBD0CB7E8D8F505498EA540A72B1D.

Bannai:2020:TGE

[4637] Shinzo Bannai and Momoko Ohno. Two-graphs and the embedded topology of smooth quartics and its bitangent lines. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 63(4):802–812, December 2020. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-

4287 (electronic). URL https://www.cambridge.org/core/journals/
canadian-mathematical-bulletin/article/twographs-and-the-embedded-
topology-of-smooth-quartics-and-its-bitangent-lines/A693C7E47A2583161765C6EFEE5D924F.

Tarcsay:2020:OAD

[4638] Zsigmond Tarcsay and Tamás Titkos. Operators on anti-dual pairs: Selfadjoint extensions and the strong Parrott theorem. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 63(4):813–824, December 2020. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-

4287 (electronic). URL https://www.cambridge.org/core/journals/
canadian-mathematical-bulletin/article/operators-on-antidual-
pairs-selfadjoint-extensions-and-the-strong-parrott-theorem/
5E2CF1E3E3357B528073B8FEFF3A98C4.

Nemati:2020:IQG

[4639] Mehdi Nemati and Maryam Rajaei Rizi. Ideals of the quantum group algebra, Arens regularity and weakly compact multipliers. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 63(4):825–836, December 2020. CODEN CMBUA3.

ISSN 0008-4395 (print), 1496-4287 (electronic). URL https://www.

cambridge.org/core/journals/canadian-mathematical-bulletin/
article/ideals-of-the-quantum-group-algebra-arens-regularity-
and-weakly-compact-multipliers/3CE6884EF4E3C406C56C57E4557A8445.

Devin:2020:LPD

[4640] Lucile Devin. Limiting properties of the distribution of primes in an arbitrarily large number of residue classes. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 63(4):837–849, December 2020. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-

4287 (electronic). URL https://www.cambridge.org/core/journals/
canadian-mathematical-bulletin/article/limiting-properties-
of-the-distribution-of-primes-in-an-arbitrarily-large-number-
of-residue-classes/B4921D5D0857F8C5BFAC8FA437D9A051.

Barany:2020:PPR

[4641] Imre Bárány, Greg Martin, Eric Naslund, and Sinai Robins. Primitive points in rational polygons. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 63(4):850–870, December 2020. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-

4287 (electronic). URL https://www.cambridge.org/core/journals/
canadian-mathematical-bulletin/article/primitive-points-in-
rational-polygons/4676795EFB835B1B1CF0960CF6582F4C.

Currey:2020:TFH

[4642] B. Currey and V. Oussa. Translates of functions on the Heisenberg group and the HRT conjecture. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 63(4):871–881, December 2020. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-

4287 (electronic). URL https://www.cambridge.org/core/journals/
canadian-mathematical-bulletin/article/translates-of-functions-
on-the-heisenberg-group-and-the-hrt-conjecture/45DB381E236EA64B81A609ABCDE7C03B.

Chapman:2020:EDS

[4643] Adam Chapman and Kelly McKinnie. Essential dimension, symbol length and p-rank. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 63(4):882–890, December 2020. CODEN CM-BUA3. ISSN 0008-4395 (print), 1496-4287 (electronic). URL https://

www.cambridge.org/core/journals/canadian-mathematical-bulletin/
article/essential-dimension-symbol-length-and-p-rank/E05F2878980F585A47897FFD831FBD14.

Okumura:2020:SLS

[4644] Katsuhiko Okumura. SNC log symplectic structures on Fano products. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 63(4):891–900, December 2020. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic). URL https://www.

cambridge.org/core/journals/canadian-mathematical-bulletin/
article/snc-log-symplectic-structures-on-fano-products/C63844DFDB98B80EC154ECF7F059B60B.

Reiser:2020:MSM

[4645] Philipp Reiser. Moduli spaces of metrics of positive scalar curvature on topological spherical space forms. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 63(4):901–908, December 2020. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-

4287 (electronic). URL https://www.cambridge.org/core/journals/
canadian-mathematical-bulletin/article/moduli-spaces-of-metrics-
of-positive-scalar-curvature-on-topological-spherical-space-
forms/63C8193520B59387FBC6D4F935DC1A50.

Wang:2020:GER

[4646] Yaning Wang. Generalized D–Einstein real hypersurfaces in CP^{2} and CH^{2}. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 63(4):909–920, December 2020. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic). URL https://www.

cambridge.org/core/journals/canadian-mathematical-bulletin/
article/generalized-mathcald-einstein-real-hypersurfaces-in-
mathbbcp2-and-mathbbch2/D2211B882497DF0DE730449AAF6AF8F8.

Kundu:2020:GFS

[4647] Debanjana Kundu. Growth of fine Selmer groups in infinite towers. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 63(4):921–936, December 2020. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic). URL https://www.

cambridge.org/core/journals/canadian-mathematical-bulletin/
article/growth-of-fine-selmer-groups-in-infinite-towers/CD4EFE6D0F0F40113E30B21426E49762.

Anonymous:2020:BVId

[4648] Anonymous. BCM volume 63 issue 4 cover and front matter. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 63(4): f1–f2, December 2020. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-

4287 (electronic). URL https://www.cambridge.org/core/journals/
canadian-mathematical-bulletin/article/bcm-volume-63-issue-
4-cover-and-front-matter/7100529999084DEF116F34005EFF8D4A.

Hare:2021:NAN

[4649] Kathryn E. Hare and Parasar Mohanty. A non-abelian, non-Sidon, completely bounded Λ(p) set. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 64(1):1–7, March 2021. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287

(electronic).  URL https://www.cambridge.org/core/journals/
canadian-mathematical-bulletin/article/nonabelian-nonsidon-
completely-bounded-lambda-p-set/683B3CB7E78FF5C4AB118159FC7B9047.

Bonet:2021:ESC

[4650] José Bonet. Every separable complex Fréchet space with a continuous norm is isomorphic to a space of holomorphic functions. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 64(1):8–12, March 2021. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic). URL https://www.cambridge.org/

core/journals/canadian-mathematical-bulletin/article/every-
separable-complex-frechet-space-with-a-continuous-norm-is-isomorphic-
to-a-space-of-holomorphic-functions/2085241BD02F54A263FD2E307C8B1DDC.

Bau:2021:SDG

[4651] Sheng Bau, Peter Johnson, and Matt Noble. On single-distance graphs on the rational points in Euclidean spaces. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 64(1):13– 24, March 2021. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-

4287 (electronic). URL https://www.cambridge.org/core/journals/
canadian-mathematical-bulletin/article/on-singledistance-graphs-
on-the-rational-points-in-euclidean-spaces/2DE2703398C1D4C409326EA97858360A.

Tan:2021:LAA

[4652] Youjun Tan and Senrong Xu. On a Lie algebraic approach to Abelian extensions of associative algebras. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 64(1):25–38, March 2021. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287

(electronic).  URL https://www.cambridge.org/core/journals/
canadian-mathematical-bulletin/article/on-a-lie-algebraic-approach-
to-abelian-extensions-of-associative-algebras/A3D4F0E45657D6809E51DF43AF9D1BBA.

Maeda:2021:MSC

[4653] Yota Maeda. The modularity of special cycles on orthogonal Shimura varieties over totally real fields under the Beilinson–Bloch conjecture. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 64(1):39–53, March 2021. CODEN CMBUA3.

ISSN 0008-4395 (print), 1496-4287 (electronic). URL https://www.

cambridge.org/core/journals/canadian-mathematical-bulletin/
article/modularity-of-special-cycles-on-orthogonal-shimura-
varieties-over-totally-real-fields-under-the-beilinsonbloch-
conjecture/69CFD43AD60EB976221A2FB94DDD01E6.

Cramer:2021:DRP

[4654] Zachary Cramer. The distance from a rank n−1 projection to the nilpotent operators on C^{n}. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 64(1):54–74, March 2021. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic). URL https://www.

cambridge.org/core/journals/canadian-mathematical-bulletin/
article/distance-from-a-rank-n1-projection-to-the-nilpotent-
operators-on-mathbb-cn/179470305E3279373A9AAB4F5B31C427.

Hurri-Syrjanen:2021:TTI

[4655] Ritva Hurri-Syrjänen, Takao Ohno, and Tetsu Shimomura. On Trudinger-type inequalities in Orlicz–Morrey spaces of an integral form. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 64(1):75–90, March 2021. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic). URL https://www.

cambridge.org/core/journals/canadian-mathematical-bulletin/
article/on-trudingertype-inequalities-in-orliczmorrey-spaces-
of-an-integral-form/4F1307FFD6AE83D24D7F2E96A75FF946.

Tout:2021:GPI

[4656] Omar Tout. Gelfand pairs involving the wreath product of finite Abelian groups with symmetric groups. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 64(1):91–97, March 2021. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-

4287 (electronic). URL https://www.cambridge.org/core/journals/
canadian-mathematical-bulletin/article/gelfand-pairs-involving-
the-wreath-product-of-finite-abelian-groups-with-symmetric-
groups/86F2FAB69F9BAD75AAFD079B507E2EDB.

Camara:2021:DCS

[4657] M. C. Câmara and W. T. Ross. The dual of the compressed shift. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 64(1):98–111, March 2021. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic). URL https://www.

cambridge.org/core/journals/canadian-mathematical-bulletin/
article/dual-of-the-compressed-shift/4D3DCD47D084D361066810909E95AC0D.

Seis:2021:NVV

[4658] Christian Seis. A note on the vanishing viscosity limit in the Yudovich class. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 64(1):112–122, March 2021. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic). URL https://www.

cambridge.org/core/journals/canadian-mathematical-bulletin/
article/note-on-the-vanishing-viscosity-limit-in-the-yudovich-
class/5C553111AC7317439D93762A61BFBB89.

Matsuno:2021:CCD

[4659] Masaki Matsuno. A complete classification of 3-dimensional quadratic AS-regular algebras of type EC. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 64(1):123–141, March 2021. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287

(electronic).  URL https://www.cambridge.org/core/journals/
canadian-mathematical-bulletin/article/complete-classification-
of-3dimensional-quadratic-asregular-algebras-of-type-ec/AAD58B46D3AA865D430F24A2852DF91F.

Azagra:2021:KTE

[4660] Daniel Azagra, Erwan Le Gruyer, and Carlos Mudarra. Kirszbraun’s theorem via an explicit formula. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 64(1):142–153, March 2021. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287

(electronic).  URL https://www.cambridge.org/core/journals/
canadian-mathematical-bulletin/article/kirszbrauns-theorem-
via-an-explicit-formula/15797B44C630B0E2A4BB12547759929D.

Yelton:2021:SME

[4661] Jeffrey Yelton. Semistable models of elliptic curves over residue characteristic 2. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 64(1):154–162, March 2021. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic). URL https://www.

cambridge.org/core/journals/canadian-mathematical-bulletin/
article/semistable-models-of-elliptic-curves-over-residue-characteristic-
2/2AD334268D191312C0FA5E9FF67ED799.

Ito:2021:NCC

[4662] Tetsuya Ito. A note on chirally cosmetic surgery on cable knots. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 64(1): 163–173, March 2021. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-

4287 (electronic). URL https://www.cambridge.org/core/journals/
canadian-mathematical-bulletin/article/note-on-chirally-cosmetic-
surgery-on-cable-knots/690DE16138573D0051FF223F9269C77C.

Betsakos:2021:HMM

[4663] Dimitrios Betsakos, Christina Karafyllia, and Nikolaos Karamanlis. Hyperbolic metric and membership of conformal maps in the Bergman space. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 64(1):174–181, March 2021. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic). URL https://www.

cambridge.org/core/journals/canadian-mathematical-bulletin/
article/hyperbolic-metric-and-membership-of-conformal-maps-
in-the-bergman-space/4D2DF3341268F8A270F83604AF03CECA.

Yang:2021:FWP

[4664] Zhao Yang and Xiaoling Zhang. Finsler warped product metrics with relatively isotropic Landsberg curvature. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 64(1):182–191, March 2021. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-

4287 (electronic). URL https://www.cambridge.org/core/journals/
canadian-mathematical-bulletin/article/finsler-warped-product-
metrics-with-relatively-isotropic-landsberg-curvature/CA893B7BA2B686B9D002C80F8BFBCB44.

Sha:2021:MRL

[4665] Min Sha and Igor E. Shparlinski. Möbius randomness law for Frobenius traces of ordinary curves. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 64(1):192–203, March 2021. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287

(electronic).  URL https://www.cambridge.org/core/journals/
canadian-mathematical-bulletin/article/mobius-randomness-law-
for-frobenius-traces-of-ordinary-curves/49D580862ED6B46C77E600F349AB1F6B.

Pollack:2021:NNC

[4666] Paul Pollack. The number of non-cyclic Sylow subgroups of the multiplicative group modulo n. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 64(1):204–215, March 2021. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287

(electronic).  URL https://www.cambridge.org/core/journals/
canadian-mathematical-bulletin/article/number-of-noncyclic-
sylow-subgroups-of-the-multiplicative-group-modulo-n/EDD3D00D8016B42AC281BEC0189FD805.

Kopper:2021:NCH

[4667] John Kopper. The Nef cone of the Hilbert scheme of points on rational elliptic surfaces and the cone conjecture. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 64(1):216–227, March 2021. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-

4287 (electronic). URL https://www.cambridge.org/core/journals/
canadian-mathematical-bulletin/article/nef-cone-of-the-hilbert-
scheme-of-points-on-rational-elliptic-surfaces-and-the-cone-
conjecture/6F7AE92B80FF9C79BF6A79699C19BFD8.

Ray:2021:ECT

[4668] Anwesh Ray and R. Sujatha. Euler characteristics and their congruences in the positive rank setting. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 64(1):228–245, March 2021. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287

(electronic).  URL https://www.cambridge.org/core/journals/
canadian-mathematical-bulletin/article/euler-characteristics-
and-their-congruences-in-the-positive-rank-setting/679D596990DFDE426146DEFE6AADDBC3.

He:2021:FEH

[4669] Tongmu He. Faltings extension and Hodge–Tate filtration for abelian varieties over p -adic local fields with imperfect residue fields. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 64(2):247–263, June 2021. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic). URL https://www.

cambridge.org/core/journals/canadian-mathematical-bulletin/
article/faltings-extension-and-hodgetate-filtration-for-abelian-
varieties-over-padic-local-fields-with-imperfect-residue-fields/
897E763FC7FFD679BCF243BB586748DA.

Mouze:2021:GFH

[4670] Augustin Mouze and Vincent Munnier. Growth of frequently hypercyclic functions for some weighted Taylor shifts on the unit disc. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 64(2):264–281, June 2021. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic). URL https://www.

cambridge.org/core/journals/canadian-mathematical-bulletin/
article/growth-of-frequently-hypercyclic-functions-for-some-
weighted-taylor-shifts-on-the-unit-disc/A06AE156CA4BE8887FDFAC2AEAFF1ACB.

Eskenazis:2021:PIU

[4671] Alexandros Eskenazis. On Pisier’s inequality for UMD targets. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 64 (2):282–291, June 2021. CODEN CMBUA3. ISSN 0008-4395 (print),

1496-4287 (electronic). URL https://www.cambridge.org/core/
journals/canadian-mathematical-bulletin/article/on-pisiers-
inequality-for-umd-targets/AF381BEA786F8D88DF5CB304BB681599.

Harlander:2021:RVA

[4672] Jens Harlander and Stephan Rosebrock. Relative vertex asphericity. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 64 (2):292–305, June 2021. CODEN CMBUA3. ISSN 0008-4395 (print),

1496-4287 (electronic). URL https://www.cambridge.org/core/
journals/canadian-mathematical-bulletin/article/relative-vertex-
asphericity/31C45B969D6D6DA3EF94C62E69241CDC.

Freslon:2021:PDF

[4673] Amaury Freslon. Positive definite functions and cut-off for discrete groups. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 64(2):306–322, June 2021. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic). URL https://www.

cambridge.org/core/journals/canadian-mathematical-bulletin/
article/positive-definite-functions-and-cutoff-for-discrete-
groups/AD3136C1E51BA45004D444B792C82061.

Abrams:2021:IMJ

[4674] Gene Abrams, Francesca Mantese, and Alberto Tonolo. Injective modules over the Jacobson algebra K⟨X, Y |XY = 1⟩. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 64(2):323– 339, June 2021. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-

4287 (electronic). URL https://www.cambridge.org/core/journals/
canadian-mathematical-bulletin/article/injective-modules-over-
the-jacobson-algebra-klangle-x-y-xy1rangle/828F508A34ADEC7E9790729374A657A4.

Pouliasis:2021:CCC

[4675] Stamatis Pouliasis. Continuity of condenser capacity under holomorphic motions. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 64(2):340–348, June 2021. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic). URL https://www.

cambridge.org/core/journals/canadian-mathematical-bulletin/
article/continuity-of-condenser-capacity-under-holomorphic-
motions/AA7359CC723A21E8E760FB4EEE450798.

Wang:2021:DLV

[4676] Haining Wang. Deligne–Lusztig varieties and basic EKOR strata. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 64 (2):349–367, June 2021. CODEN CMBUA3. ISSN 0008-4395 (print),

1496-4287 (electronic). URL https://www.cambridge.org/core/
journals/canadian-mathematical-bulletin/article/delignelusztig-
varieties-and-basic-ekor-strata/ECFA05927B3BC4493CE644EA962985D0.

Griniasty:2021:ESD

[4677] Itay Griniasty and Ruth Lawrence. Explicit symmetric DGLA models of 3-cells. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 64(2):368–382, June 2021. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic). URL https://www.

cambridge.org/core/journals/canadian-mathematical-bulletin/
article/explicit-symmetric-dgla-models-of-3cells/1EB9B2150A10C31DF58C15D1F8C21DE5.

Endo:2021:AGR

[4678] Naoki Endo, Shiro Goto, and Ryotaro Isobe. Almost Gorenstein rings arising from fiber products. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 64(2):383–400, June 2021. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287

(electronic).  URL https://www.cambridge.org/core/journals/
canadian-mathematical-bulletin/article/almost-gorenstein-rings-
arising-from-fiber-products/9FD56A6CB3B224EE8A97630C6D3A8C26.

Bridy:2021:QIG

[4679] Andrew Bridy, John R. Doyle, Dragos Ghioca, Liang-Chung Hsia, and Thomas J. Tucker. A question for iterated Galois groups in arithmetic dynamics. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 64(2):401–417, June 2021. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic). URL https://www.

cambridge.org/core/journals/canadian-mathematical-bulletin/
article/question-for-iterated-galois-groups-in-arithmetic-dynamics/
BB3B19931BA92F5E03C1B870FE0A0E62.

Akhmedov:2021:GSC

[4680] Anar Akhmedov and B. Doug Park. Geography of simply connected nonspin symplectic 4-manifolds with positive signature. II. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 64(2): 418–428, June 2021. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-

4287 (electronic). URL https://www.cambridge.org/core/journals/
canadian-mathematical-bulletin/article/geography-of-simply-
connected-nonspin-symplectic-4manifolds-with-positive-signature-
ii/285C4BAD31A725E7E11D441C881A6C9A.

Dijkstra:2021:CAZ

[4681] Jan J. Dijkstra and David S. Lipham. On cohesive almost zerodimensional spaces. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 64(2):429–441, June 2021. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic). URL https://www.

cambridge.org/core/journals/canadian-mathematical-bulletin/
article/on-cohesive-almost-zerodimensional-spaces/3DC9FADB729C00A92E74EAEBF95808AE.

Mello:2021:IPS

[4682] Jorge Mello. On intersections of polynomial semigroups orbits with plane lines. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 64(2):442–451, June 2021. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic). URL https://www.

cambridge.org/core/journals/canadian-mathematical-bulletin/
article/on-intersections-of-polynomial-semigroups-orbits-with-
plane-lines/EF37043EF7CBE2BED935C0485CCEEE91.

Baldi:2021:FDO

[4683] Gregorio Baldi and Giada Grossi. Finite descent obstruction for Hilbert modular varieties. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 64(2):452–473, June 2021. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic). URL https://www.

cambridge.org/core/journals/canadian-mathematical-bulletin/
article/finite-descent-obstruction-for-hilbert-modular-varieties/
B21E7CDF8F5A95AEBE13A457A2A8C199.

Coine:2021:CBM

[4684] Clément Coine. Complete boundedness of multiple operator integrals. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 64(2):474–490, June 2021. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic). URL https://www.

cambridge.org/core/journals/canadian-mathematical-bulletin/
article/complete-boundedness-of-multiple-operator-integrals/
82925AE3EF31DCAA45CFD3CDAD89D71E.

Klyachko:2021:KBG

[4685] Anton A. Klyachko. The Klein bottle group is not strongly verbally closed, though awfully close to being so. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 64(2):491–497, June 2021. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-

4287 (electronic). URL https://www.cambridge.org/core/journals/
canadian-mathematical-bulletin/article/klein-bottle-group-is-
not-strongly-verbally-closed-though-awfully-close-to-being-
so/406D0D56F8F0982AFDBBD51FDB1F8ED9.

Anonymous:2021:BVIa

[4686] Anonymous. BCM volume 64 issue 2 cover and front matter. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 64(2): f1–f2, June 2021. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-

4287 (electronic). URL https://www.cambridge.org/core/journals/
canadian-mathematical-bulletin/article/bcm-volume-64-issue-
2-cover-and-front-matter/F3DF666D8298E7C3BE2F11EA4249B675.

Anonymous:2021:BVIb

[4687] Anonymous. BCM volume 64 issue 2 cover and back matter. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 64(2): b1–b2, June 2021. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-

4287 (electronic). URL https://www.cambridge.org/core/journals/
canadian-mathematical-bulletin/article/bcm-volume-64-issue-
2-cover-and-back-matter/58ACDCFA0757CA1F384B94DD831A38FD.

Vandervelde:2021:BBS

[4688] Sam Vandervelde. Beyond Beatty sequences: Complementary lattices. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 64(3):499–511, September 2021. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic). URL https://www.

cambridge.org/core/journals/canadian-mathematical-bulletin/
article/beyond-beatty-sequences-complementary-lattices/9B065104110E7B0A1F9D1A1C3CEDCB65.

Debernardi:2021:AHO

[4689] Alberto Debernardi and Elijah Liflyand. Approximation via Hausdorff operators. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 64(3):512–529, September 2021. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic). URL https://www.

cambridge.org/core/journals/canadian-mathematical-bulletin/
article/approximation-via-hausdorff-operators/B8BA7994B2259A53FC54C07A831474EA.

deShalit:2021:CPA

[4690] Ehud de Shalit. Criteria for periodicity and an application to elliptic functions. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 64(3):530–540, September 2021. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic). URL https://www.

cambridge.org/core/journals/canadian-mathematical-bulletin/
article/criteria-for-periodicity-and-an-application-to-elliptic-
functions/54D6A39DA999CD043F60D03BBD999ADE.

Anderson:2021:MAM

[4691] Robert M. Anderson, Haosui Duanmu, and Aaron Smith. Mixing and average mixing times for general Markov processes. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 64(3):541– 552, September 2021. CODEN CMBUA3. ISSN 0008-4395 (print),

1496-4287 (electronic). URL https://www.cambridge.org/core/
journals/canadian-mathematical-bulletin/article/mixing-and-
average-mixing-times-for-general-markov-processes/84EA62734FA0297C1630C7C2886F5AA0.

Bhatia:2021:VPS

[4692] Rajendra Bhatia and Tanvi Jain. Variational principles for symplectic eigenvalues. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 64(3):553–559, September 2021. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic). URL https://www.

cambridge.org/core/journals/canadian-mathematical-bulletin/
article/variational-principles-for-symplectic-eigenvalues/B560F355B04672E5348F59AAF8DCB09E.

Burelle:2021:RDE

[4693] Jean-Philippe Burelle. Rigidity of diagonally embedded triangle groups. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 64(3):560–579, September 2021. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic). URL https://www.

cambridge.org/core/journals/canadian-mathematical-bulletin/
article/rigidity-of-diagonally-embedded-triangle-groups/CC7F88AE8F0132A6C3262520A962F27D.

Chan:2021:ECC

[4694] William Chan. Equivalence of codes for countable sets of reals. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 64(3):580–590, September 2021. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic). URL https://www.

cambridge.org/core/journals/canadian-mathematical-bulletin/
article/equivalence-of-codes-for-countable-sets-of-reals/1A2479F5626E06EDBFB9B5B8E807A015.

Dwivedi:2021:SRR

[4695] Shubham Dwivedi. Some results on Ricci–Bourguignon solitons and almost solitons. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 64(3):591–604, September 2021. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic). URL https://www.

cambridge.org/core/journals/canadian-mathematical-bulletin/
article/some-results-on-riccibourguignon-solitons-and-almost-
solitons/FDBE26EA8B4FBC7A3A9640F151734828.

Lei:2021:STS

[4696] Tao Lei. Small trees in supercritical random forests. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 64(3): 605–623, September 2021. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic). URL https://www.cambridge.org/

core/journals/canadian-mathematical-bulletin/article/small-
trees-in-supercritical-random-forests/B1EC0593B6CF63840D689955218402CE.

theKhoi:2021:LOS

[4697] Vu the Khoi, Masakazu Teragaito, and Anh T. Tran. Left orderable surgeries of double twist knots II. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 64(3):624–637, September 2021. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-

4287 (electronic). URL https://www.cambridge.org/core/journals/
canadian-mathematical-bulletin/article/left-orderable-surgeries-
of-double-twist-knots-ii/D23D5A76C7A71FA5356104604F16E1B9.

Khodabandehlou:2021:LCA

[4698] J. Khodabandehlou, S. Maghsoudi, and J. B. Seoane-Sepúlveda. Lineability, continuity, and antiderivatives in the non-Archimedean setting. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 64(3):638–650, September 2021. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic). URL https://www.

cambridge.org/core/journals/canadian-mathematical-bulletin/
article/lineability-continuity-and-antiderivatives-in-the-nonarchimedean-
setting/DDFB17CD19DCA259259E21E299F86EC1.

Ingalls:2021:CCG

[4699] Colin Ingalls, Bruce W. Jordan, Allan Keeton, Adam Logan, and Yevgeny Zaytman. The Clifford-cyclotomic group and Euler–Poincaré characteristics. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 64(3):651–666, September 2021. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic). URL https://www.

cambridge.org/core/journals/canadian-mathematical-bulletin/
article/cliffordcyclotomic-group-and-eulerpoincare-characteristics/
548CDB6A8B0BB40FE2D219160C9BAFE7.

Nghiem:2021:RPF

[4700] Tran-Trung Nghiem and Zinovy Reichstein. On a rationality problem for fields of cross-ratios II. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 64(3):667–677, September 2021. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-

4287 (electronic). URL https://www.cambridge.org/core/journals/
canadian-mathematical-bulletin/article/on-a-rationality-problem-
for-fields-of-crossratios-ii/A4B7925271DB5B24E17ED2D0A8FA0288.

Abtahi:2021:ASC

[4701] Fatemeh Abtahi. C^{∗}-algebra structure on certain Banach algebra products. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 64(3):678–686, September 2021. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic). URL https://www.

cambridge.org/core/journals/canadian-mathematical-bulletin/
article/c-algebra-structure-on-certain-banach-algebra-products/
D12CAAC8668629FA1AC337B1B2F8CEE5.

Zhang:2021:DRO

[4702] Helen W. J. Zhang. Dyson’s rank, overpartitions, and universal mock theta functions. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 64(3):687–696, September 2021. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic). URL https://www.

cambridge.org/core/journals/canadian-mathematical-bulletin/
article/dysons-rank-overpartitions-and-universal-mock-theta-
functions/5F8441A848F7BFBBFD3718AEA56B7B88.

Qian:2021:EDT

[4703] Ruishen Qian and Xiangling Zhu. Embedding of Dirichlet type spaces D_{p−1}^{p} into tent spaces and Volterra operators. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 64(3):697–708, September 2021. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-

4287 (electronic). URL https://www.cambridge.org/core/journals/
canadian-mathematical-bulletin/article/embedding-of-dirichlet-
type-spaces-mathcal-dpp1-into-tent-spaces-and-volterra-operators/
7B8ED910B37276493DF0DFD2AE9F821C.

Garcia:2021:HSL

[4704] Daniel López Garcia. Homology supported in Lagrangian submanifolds in mirror quintic threefolds. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 64(3):709–724, September 2021. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-

4287 (electronic). URL https://www.cambridge.org/core/journals/
canadian-mathematical-bulletin/article/homology-supported-in-
lagrangian-submanifolds-in-mirror-quintic-threefolds/9AABC8B4F270BEC9908FC2D77F24AAF1.

Anonymous:2021:BVIc

[4705] Anonymous. BCM volume 64 issue 3 cover and front matter. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 64(3): f1–f2, September 2021. CODEN CMBUA3. ISSN 0008-4395 (print),

1496-4287 (electronic). URL https://www.cambridge.org/core/
journals/canadian-mathematical-bulletin/article/bcm-volume-
64-issue-3-cover-and-front-matter/5065BF6D28B3708CBCD42BD4A2C8B3AF.

Anonymous:2021:BVId

[4706] Anonymous. BCM volume 64 issue 3 cover and back matter. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 64(3): b1–b2, September 2021. CODEN CMBUA3. ISSN 0008-4395 (print),

1496-4287 (electronic). URL https://www.cambridge.org/core/
journals/canadian-mathematical-bulletin/article/bcm-volume-
64-issue-3-cover-and-back-matter/4C5BA311D9B2EBE66F02EB3A531EF74B.

Kinzebulatov:2021:RTK

[4707] Damir Kinzebulatov. Regularity theory of Kolmogorov operator revisited. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 64(4):725–736, December 2021. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic). URL https://www.

cambridge.org/core/journals/canadian-mathematical-bulletin/
article/regularity-theory-of-kolmogorov-operator-revisited/
86211F0086CC5E34B11DD42B326A910B.

Dweik:2021:OFE

[4708] Samer Dweik. Optimal free export/import regions. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 64(4):737–751, December 2021. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-

4287 (electronic). URL https://www.cambridge.org/core/journals/
canadian-mathematical-bulletin/article/optimal-free-exportimport-
regions/7FFDCFC69F2AC71263C07B3C986F7C52.

Bakshi:2021:FLM

[4709] Rhea Palak Bakshi, Dionne Ibarra, Gabriel Montoya-Vega, Józef H. Przytycki, and Deborah Weeks. On framings of links in 3-manifolds. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 64(4):752–764, December 2021. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic). URL https://www.cambridge.org/

core/journals/canadian-mathematical-bulletin/article/on-framings-
of-links-in-3manifolds/FA7F58F8346F9534DF7E7C3B31B3BD5B.

Duy:2021:SAT

[4710] Nguyen Tuan Duy, Nguyen Lam, and Phi Long Le. Sharp affine Trudinger–Moser inequalities: a new argument. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 64(4): 765–778, December 2021. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic). URL https://www.cambridge.org/

core/journals/canadian-mathematical-bulletin/article/sharp-
affine-trudingermoser-inequalities-a-new-argument/C74B4F70483183B53BEDC8A004F78251.

Perez:2021:NPR

[4711] Rolando Perez III. A note on the phase retrieval of holomorphic functions. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 64(4):779–786, December 2021. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic). URL https://www.

cambridge.org/core/journals/canadian-mathematical-bulletin/
article/note-on-the-phase-retrieval-of-holomorphic-functions/
BF3DB2146B5D94DF75271B1082A99403.

He:2021:BNG

[4712] Xiang He. Brill–Noether generality of binary curves. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 64(4):787–807, December 2021. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-

4287 (electronic). URL https://www.cambridge.org/core/journals/
canadian-mathematical-bulletin/article/brillnoether-generality-
of-binary-curves/AECDD55EC351D6F98B139C0B4ECD3A85.

Lucchini:2021:VGG

[4713] Andrea Lucchini. The virtually generating graph of a profinite group. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 64(4):808–819, December 2021. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic). URL https://www.

cambridge.org/core/journals/canadian-mathematical-bulletin/
article/virtually-generating-graph-of-a-profinite-group/3D5F4D549424F841C35199C95D3D70A9.

Chen:2021:RBM

[4714] Yin Chen. Relations between modular invariants of a vector and a covector in dimension two. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 64(4):820–827, December 2021. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-

4287 (electronic). URL https://www.cambridge.org/core/journals/
canadian-mathematical-bulletin/article/relations-between-modular-
invariants-of-a-vector-and-a-covector-in-dimension-two/DD00C28A3B9FF3C140288123A41CE8D1.

Beardon:2021:SIM

[4715] A. F. Beardon and D. Minda. Surjective isometries of metric geometries. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 64(4):828–839, December 2021. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic). URL https://www.

cambridge.org/core/journals/canadian-mathematical-bulletin/
article/surjective-isometries-of-metric-geometries/FCAFCD10121D7FFBF26DF628DE6FF93B.

Cellarosi:2021:AFQ

[4716] Francesco Cellarosi. Autocorrelation functions for quantum particles in supersymmetric Pöschl–Teller potentials. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 64(4):840–852, December 2021. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-

4287 (electronic). URL https://www.cambridge.org/core/journals/
canadian-mathematical-bulletin/article/autocorrelation-functions-
for-quantum-particles-in-supersymmetric-poschlteller-potentials/
7BDFC0A9CDF5D2731F90B23F8113E7BB.

Kuznetsova:2021:IBW

[4717] Yulia Kuznetsova and Safoura Zadeh. On isomorphisms between weighted L^{p}-algebras. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 64(4):853–866, December 2021. CODEN CM-BUA3. ISSN 0008-4395 (print), 1496-4287 (electronic). URL https://

www.cambridge.org/core/journals/canadian-mathematical-bulletin/
article/on-isomorphisms-between-weighted-lp-algebras/637DB43AA5634EB3D70C456D3B1F1BCC.

Ingram:2021:DGM

[4718] Patrick Ingram. Degree gaps for multipliers and the dynamical André-Oort conjecture. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 64(4):867–876, December 2021. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic). URL https://www.

cambridge.org/core/journals/canadian-mathematical-bulletin/
article/degree-gaps-for-multipliers-and-the-dynamical-andreoort-
conjecture/5E1BCA92B13123EE277C8D60F1D0DC9C.

Szechtman:2021:DRR

[4719] Fernando Szechtman. On the degree of repeated radical extensions. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 64(4):877–885, December 2021. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic). URL https://www.

cambridge.org/core/journals/canadian-mathematical-bulletin/
article/on-the-degree-of-repeated-radical-extensions/0B6DA506113644F80E89BE756F3672C7.

Ballesteros-Chavez:2021:PSC

[4720] Daniel Ballesteros-Chávez, Wilhelm Klingenberg, and Ben Lambert. Prescribed k-symmetric curvature hypersurfaces in de Sitter space. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 64(4): 886–901, December 2021. CODEN CMBUA3. ISSN 0008-4395 (print),

1496-4287 (electronic). URL https://www.cambridge.org/core/
journals/canadian-mathematical-bulletin/article/prescribed-
ksymmetric-curvature-hypersurfaces-in-de-sitter-space/1A86ADBFA486070E676B2E37FFCDE244.

Ransford:2021:DSI

[4721] Thomas Ransford. On the decay of singular inner functions. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 64(4): 902–905, December 2021. CODEN CMBUA3. ISSN 0008-4395 (print),

1496-4287 (electronic). URL https://www.cambridge.org/core/
journals/canadian-mathematical-bulletin/article/on-the-decay-
of-singular-inner-functions/E8F3D10B6A303C3086BFB23B7C967D33.

Bordignon:2021:PVI

[4722] Matteo Bordignon. A Pólya–Vinogradov inequality for short character sums. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 64(4):906–910, December 2021. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic). URL https://www.

cambridge.org/core/journals/canadian-mathematical-bulletin/
article/polyavinogradov-inequality-for-short-character-sums/
C27802DED380B7C1547CC60FDBF77AE4.

Tao:2021:ECR

[4723] Xiangxing Tao and Guoen Hu. An estimate for the composition of rough singular integral operators. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 64(4):911–922, December 2021. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-

4287 (electronic). URL https://www.cambridge.org/core/journals/
canadian-mathematical-bulletin/article/an-estimate-for-the-
composition-of-rough-singular-integral-operators/9779989CC5A14BAB2886F5F570FDE8FC.

Scandone:2021:CSS

[4724] Raffaele Scandone, Lorenzo Luperi Baglini, and Kyrylo Simonov. A characterization of singular Schrödinger operators on the halfline. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 64(4):923–941, December 2021. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic). URL https://www.

cambridge.org/core/journals/canadian-mathematical-bulletin/
article/characterization-of-singular-schrodinger-operators-
on-the-halfline/C8CDC2483059758E20EE081B0AE61A8A.

Ivanov:2021:TFR

[4725] Grigory Ivanov. Tight frames and related geometric problems. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 64(4):942–963, December 2021. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic). URL https://www.cambridge.org/

core/journals/canadian-mathematical-bulletin/article/tight-
frames-and-related-geometric-problems/7ADB1EED216CF79C8F836BCACFE42B17.

Fu:2021:IPT

[4726] Xiaohui Fu, Pan-Shun Lau, and Tin-Yau Tam. Inequalities on partial traces of positive semidefinite block matrices. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 64(4):964–969, December 2021. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-

4287 (electronic). URL https://www.cambridge.org/core/journals/
canadian-mathematical-bulletin/article/inequalities-on-partial-
traces-of-positive-semidefinite-block-matrices/4A9A4F15205EB8570FD2CD24A24DE8E9.

Martin:2022:SSR

[4727] Greg Martin and Amir Parvardi. Subproducts of small residue classes. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 65(1):1–8, March 2022. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic). URL https://www.

cambridge.org/core/journals/canadian-mathematical-bulletin/
article/subproducts-of-small-residue-classes/C0F5EC7105A4C010903EC7F9EA1909B4.

Shi:2022:VMV

[4728] Shuhui Shi. Vanishing of multizeta values over F_{q}[t] at negative integers. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 65(1):9–29, March 2022. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic). URL https://www.

cambridge.org/core/journals/canadian-mathematical-bulletin/
article/vanishing-of-multizeta-values-over-mathbb-fqt-at-negative-
integers/BFE650CC896BC2F4232721B695EDC7FC.

Azad:2022:TCA

[4729] Morteza Baniasad Azad and Behrooz Khosravi. On two conjectures about the sum of element orders. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 65(1):30–38, March 2022. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287

(electronic).  URL https://www.cambridge.org/core/journals/
canadian-mathematical-bulletin/article/on-two-conjectures-about-
the-sum-of-element-orders/2480BEAD6C8B35EEAF98E499DF99F6D0.

Zhou:2022:MDN

[4730] Yiqiang Zhou. A multiplicative dual of nil-clean rings. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 65(1):39– 43, March 2022. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-

4287 (electronic). URL https://www.cambridge.org/core/journals/
canadian-mathematical-bulletin/article/multiplicative-dual-
of-nilclean-rings/BD63782127A74F71C0299433C00573D9.

Yan:2022:CHO

[4731] Yi Yan. Characterizations of Hankel operators in the essential commutant of quasicontinuous Toeplitz operators. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 65(1):44–51, March 2022. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-

4287 (electronic). URL https://www.cambridge.org/core/journals/
canadian-mathematical-bulletin/article/characterizations-of-
hankel-operators-in-the-essential-commutant-of-quasicontinuous-
toeplitz-operators/58B6D7B029FDE7FEAEB550A2AE4CAB30.

Ingram:2022:EFS

[4732] Patrick Ingram. Effective finiteness of solutions to certain differential and difference equations. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 65(1):52–67, March 2022. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic). URL https://www.

cambridge.org/core/journals/canadian-mathematical-bulletin/
article/effective-finiteness-of-solutions-to-certain-differential-
and-difference-equations/6B4F5B553ACFF649EF3CB9E27CEC7BC9.

Yao:2022:NEC

[4733] Zeke Yao, Bangchao Yin, and Zejun Hu. Non-existence of conformally flat real hypersurfaces in both the complex quadric and the complex hyperbolic quadric. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 65(1):68–83, March 2022. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic). URL https://www.

cambridge.org/core/journals/canadian-mathematical-bulletin/
article/nonexistence-of-conformally-flat-real-hypersurfaces-
in-both-the-complex-quadric-and-the-complex-hyperbolic-quadric/
3D5705BE187B078532C67512794DFD51.

Ruzsa:2022:SSS

[4734] Imre Ruzsa and Jozsef Solymosi. Sumsets of semiconvex sets. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 65 (1):84–94, March 2022. CODEN CMBUA3. ISSN 0008-4395 (print),

1496-4287 (electronic). URL https://www.cambridge.org/core/
journals/canadian-mathematical-bulletin/article/sumsets-of-
semiconvex-sets/7386409D38B966DFD6E2822519E28900.

Xi:2022:SMI

[4735] Bo-Yan Xi and Fuzhen Zhang. Some matrix inequalities of logmajorization type. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 65(1):95–105, March 2022. CODEN CMBUA3.

ISSN 0008-4395 (print), 1496-4287 (electronic). URL https://www.

cambridge.org/core/journals/canadian-mathematical-bulletin/
article/some-matrix-inequalities-of-logmajorization-type/CA48A5CBB694D4C7C759B310519B1C61.

Goncharov:2022:QEB

[4736] Alexander Goncharov and Yasemin Sengül. Quasi-equivalence of bases in some Whitney spaces. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 65(1):106–115, March 2022. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287

(electronic).  URL https://www.cambridge.org/core/journals/
canadian-mathematical-bulletin/article/quasiequivalence-of-
bases-in-some-whitney-spaces/1B477C8F0C2A5733CBA7DFB3FC0F5041.

Ghioca:2022:ZDO

[4737] Dragos Ghioca and Sina Saleh. Zariski dense orbits for regular self-maps on split semiabelian varieties. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 65(1):116–122, March 2022. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-

4287 (electronic). URL https://www.cambridge.org/core/journals/
canadian-mathematical-bulletin/article/zariski-dense-orbits-
for-regular-selfmaps-on-split-semiabelian-varieties/737B46F1B9D1B7B24991852C4D2DDE53.

Huang:2022:UEM

[4738] Hansong Huang and Dechao Zheng. Unitary equivalence of multiplication operators on the Bergman spaces of polygons. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 65(1):123– 133, March 2022. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-

4287 (electronic). URL https://www.cambridge.org/core/journals/
canadian-mathematical-bulletin/article/unitary-equivalence-
of-multiplication-operators-on-the-bergman-spaces-of-polygons/
6832DAE98594F0C527CDF8A3126E8594.

Bosch:2022:TAW

[4739] Henry Bosch, Tyler Gonzales, Kamryn Spinelli, Gabe Udell, and Yunus E. Zeytuncu. A Tauberian approach to Weyl’s law for the Kohn Laplacian on spheres. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 65(1):134–154, March 2022. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic). URL https://www.

cambridge.org/core/journals/canadian-mathematical-bulletin/
article/tauberian-approach-to-weyls-law-for-the-kohn-laplacian-
on-spheres/6416669B09A504A1572091F7304C9E95.

Rumynin:2022:CAH

[4740] Dmitriy Rumynin and Matthew Westaway. Centrification of algebras and Hopf algebras. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 65(1):155–169, March 2022. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic). URL https://www.

cambridge.org/core/journals/canadian-mathematical-bulletin/
article/centrification-of-algebras-and-hopf-algebras/8259A4B7AD2AB857916437071B2B6C63.

Cuckovic:2022:ZPT

[4741] Zeljko Cucković, Zhenghui Huo, and Sönmez Sahutoglu. Zero products of Toeplitz operators on Reinhardt domains. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 65(1):170–179, March 2022. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-

4287 (electronic). URL https://www.cambridge.org/core/journals/
canadian-mathematical-bulletin/article/zero-products-of-toeplitz-
operators-on-reinhardt-domains/32FF02AE2F070F18C70768DEBC8A0BEA.

Karapetrovic:2022:VIM

[4742] Boban Karapetrović. Volume integral means over spherical shell. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 65(1): 180–197, March 2022. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-

4287 (electronic). URL https://www.cambridge.org/core/journals/
canadian-mathematical-bulletin/article/volume-integral-means-
over-spherical-shell/EB919C7C2646CD10AE55D877D78B65EA.

Moshchevitin:2022:DNL

[4743] Nikolay G. Moshchevitin. On the distribution of nonprimitive lattice points in the plane. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 65(1):198–207, March 2022. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic). URL https://www.

cambridge.org/core/journals/canadian-mathematical-bulletin/
article/on-the-distribution-of-nonprimitive-lattice-points-
in-the-plane/D8E4CFA4E060C761C1EC72E9D0CC89A7.

Benyamine:2022:FSM

[4744] Charif Abdallah Benyamine. On finite sections of the multiplicative Hilbert inequalities. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 65(1):208–216, March 2022. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic). URL https://www.

cambridge.org/core/journals/canadian-mathematical-bulletin/
article/on-finite-sections-of-the-multiplicative-hilbert-inequalities/
337B14B5FC144940D1BFB495D5C09356.

Corson:2022:BOD

[4745] Samuel M. Corson. Bi-orders do not arise from total orders. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 65(1): 217–224, March 2022. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-

4287 (electronic). URL https://www.cambridge.org/core/journals/
canadian-mathematical-bulletin/article/biorders-do-not-arise-
from-total-orders/5E653DA97E087707B0001202F98978AE.

Rump:2022:USY

[4746] Wolfgang Rump. Uniconnected solutions to the Yang–Baxter equation arising from self-maps of groups. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 65(1):225–233, March 2022. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287

(electronic).  URL https://www.cambridge.org/core/journals/
canadian-mathematical-bulletin/article/uniconnected-solutions-
to-the-yangbaxter-equation-arising-from-selfmaps-of-groups/
3BC160E9761F5723CEADD8DB97AEE9E8.

Lamkin:2022:LCL

[4747] Philip Lamkin and Tomasz Tkocz. Log-concavity and log-convexity of moments of averages of i.i.d. random variables. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 65(2):271– 278, June 2022. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-

4287 (electronic). URL https://www.cambridge.org/core/journals/
canadian-mathematical-bulletin/article/logconcavity-and-logconvexity-
of-moments-of-averages-of-iid-random-variables/E2D62AA1B9C4A6324457BC6DBE47711F.

Gambini:2022:DDQ

[4748] Alessandro Gambini, Remis Tonon, and Alessandro Zaccagnini. On the distribution of the digits of quotients of integers and primes. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 65(2): 279–295, June 2022. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-

4287 (electronic). URL https://www.cambridge.org/core/journals/
canadian-mathematical-bulletin/article/on-the-distribution-
of-the-digits-of-quotients-of-integers-and-primes/0189F8DD681306D81113F7EF8BF173D0.

Golasinski:2022:HNL

[4749] Marek Golasiński. On homotopy nilpotency of loop spaces of Moore spaces. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 65(2):296–307, June 2022. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic). URL https://www.

cambridge.org/core/journals/canadian-mathematical-bulletin/

article/on-homotopy-nilpotency-of-loop-spaces-of-moore-spaces/
417DB3D52A872057B6E5565253DD703D.

Laterveer:2022:CRS

[4750] Robert Laterveer. On the Chow ring of some special Calabi–Yau varieties. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 65(2):308–327, June 2022. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic). URL https://www.

cambridge.org/core/journals/canadian-mathematical-bulletin/
article/on-the-chow-ring-of-some-special-calabiyau-varieties/
B7A8AE006B9C985027AAA3A20EA3B607.

Sun:2022:CMT

[4751] Fangmei Sun and Hasi Wulan. Characterizations of Morrey type spaces. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 65(2):328–344, June 2022. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic). URL https://www.

cambridge.org/core/journals/canadian-mathematical-bulletin/
article/characterizations-of-morrey-type-spaces/2A4197A0DB194AE8D8C0AF5132648557.

Donato:2022:ESM

[4752] Paul Donato and Patrick Iglesias-Zemmour. Every symplectic manifold is a (linear) coadjoint orbit. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 65(2):345–360, June 2022. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287

(electronic).  URL https://www.cambridge.org/core/journals/
canadian-mathematical-bulletin/article/every-symplectic-manifold-
is-a-linear-coadjoint-orbit/60EEA4D0397E72FAE554ADEE9EC800B0.

Gauthier:2022:AFB

[4753] Paul M. Gauthier and Mohammad Shirazi. Asymptotic first boundary value problem for holomorphic functions of several complex variables. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 65(2):361–380, June 2022. CODEN CMBUA3. ISSN 0008-4395 (print),

1496-4287 (electronic). URL https://www.cambridge.org/core/
journals/canadian-mathematical-bulletin/article/asymptotic-
first-boundary-value-problem-for-holomorphic-functions-of-several-
complex-variables/5555DECBF7D497CA90693A3825234EA2.

McKee:2022:AIA

[4754] Andrew McKee and Reyhaneh Pourshahami. Amenable and inner amenable actions and approximation properties for crossed products by locally compact groups. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 65(2):381–399, June 2022. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic). URL https://www.

cambridge.org/core/journals/canadian-mathematical-bulletin/
article/amenable-and-inner-amenable-actions-and-approximation-
properties-for-crossed-products-by-locally-compact-groups/2704BBB5F9ADFC32CA2CEED8725D3E25.

Billig:2022:BSF

[4755] Yuly Billig. Bihomogeneous symmetric functions. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 65(2):400– 408, June 2022. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-

4287 (electronic). URL https://www.cambridge.org/core/journals/
canadian-mathematical-bulletin/article/bihomogeneous-symmetric-
functions/3F9113DC906FB7DE68B3726391B2885E.

Dubon:2022:NDZ

[4756] Eric Dubon. Une note sur la densité des zéros des sommes partielles de la fonction zeta de Dedekind sur un corps quadratique. (French) [A note on the density of the zeros of the partial sums of the Dedekind zeta function over a quadratic field]. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 65(2):409–415, June 2022. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic). URL https://www.

cambridge.org/core/journals/canadian-mathematical-bulletin/
article/une-note-sur-la-densite-des-zeros-des-sommes-partielles-
de-la-fonction-zeta-de-dedekind-sur-un-corps-quadratique/F2AE5F0A3D8CFF5E4E797B28EEE6A544.

Leonetti:2022:TID

[4757] Paolo Leonetti and Michele Caprio. Turnpike in infinite dimension. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 65(2):416–430, June 2022. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic). URL https://www.

cambridge.org/core/journals/canadian-mathematical-bulletin/
article/turnpike-in-infinite-dimension/3B2E5D2C11A41B513DDD7C89B69556BA.

Bourget:2022:SDS

[4758] Alain Bourget. Spectral distribution of symmetrized circulant matrices. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 65(2):431–446, June 2022. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic). URL https://www.

cambridge.org/core/journals/canadian-mathematical-bulletin/
article/spectral-distribution-of-symmetrized-circulant-matrices/
507B07BEEB4D9684437E6799AB0938EC.

Kuwahara:2022:EBR

[4759] Shuhei Kuwahara and Michio Seto. Exponentials of de Branges– Rovnyak kernels. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 65(2):447–455, June 2022. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic). URL https://www.

cambridge.org/core/journals/canadian-mathematical-bulletin/
article/exponentials-of-de-brangesrovnyak-kernels/4E7484E0A9D5ABFD2DCBBF3AFAE8F7F6.

Christensen:2022:CVR

[4760] Ole Christensen and Marzieh Hasannasab. Completion versus removal of redundancy by perturbation. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 65(2):456–465, June 2022. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287

(electronic).  URL https://www.cambridge.org/core/journals/
canadian-mathematical-bulletin/article/completion-versus-removal-
of-redundancy-by-perturbation/4769303430C768AA828EDC74665AF0E1.

Vijiitu:2022:RTC

[4761] Viorel Vîjîitu. A Radó theorem for complex spaces. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 65(2):466– 472, June 2022. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-

4287 (electronic). URL https://www.cambridge.org/core/journals/
canadian-mathematical-bulletin/article/rado-theorem-for-complex-
spaces/849BD32CC21ECE434C13D8026D778540.

Su:2022:NPT

[4762] Yu Su and Senli Liu. Nehari–Pohozaev-type ground state solutions of Kirchhoff-type equation with singular potential and critical exponent. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 65(2):473–495, June 2022. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic). URL https://www.

cambridge.org/core/journals/canadian-mathematical-bulletin/
article/neharipohozaevtype-ground-state-solutions-of-kirchhofftype-
equation-with-singular-potential-and-critical-exponent/6CA121F2DA12FC5115C0A568E0622361.

Jin:2022:ECA

[4763] Steven Jin and Lawrence C. Washington. An elliptic curve analogue of Pillai’s lower bound on primitive roots. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 65(2):496– 505, June 2022. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-

4287 (electronic). URL https://www.cambridge.org/core/journals/
canadian-mathematical-bulletin/article/an-elliptic-curve-analogue-
of-pillais-lower-bound-on-primitive-roots/436AACEA84E5E62EE1CF715A51128270.

Arcila-Maya:2022:DTA

[4764] Niny Arcila-Maya. Decomposition of topological Azumaya algebras. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 65(2):506–524, June 2022. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic). URL https://www.

cambridge.org/core/journals/canadian-mathematical-bulletin/
article/decomposition-of-topological-azumaya-algebras/0B3A4AB673A78F281B97686CBD45623A.

Quadrelli:2022:GTF

[4765] Claudio Quadrelli. Galois-theoretic features for 1-smooth pro-p groups. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 65(2):525–541, June 2022. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic). URL https://www.

cambridge.org/core/journals/canadian-mathematical-bulletin/
article/galoistheoretic-features-for-1smooth-prop-groups/696DC8538C88AF73841B3551D572AD56.

Anonymous:2022:BVI

[4766] Anonymous. BCM volume 65 issue 2 cover and front matter. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 65(2): f1–f3, June 2022. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-

4287 (electronic). URL https://www.cambridge.org/core/journals/
canadian-mathematical-bulletin/article/bcm-volume-65-issue-
2-cover-and-front-matter/D6F8B53369C53D8457C5F526E20DD8F2.

Hui:2022:EHE

[4767] Kin Ming Hui. Existence of hypercylinder expanders of the inverse mean curvature flow. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 65(3):543–551, September 2022. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic). URL https://www.

cambridge.org/core/journals/canadian-mathematical-bulletin/
article/existence-of-hypercylinder-expanders-of-the-inverse-
mean-curvature-flow/2ED15BE5721E4BD616925720912FD974.

Martinez-Maure:2022:NAC

[4768] Yves Martinez-Maure. Noncircular algebraic curves of constant width: an answer to Rabinowitz. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 65(3):552–556, September 2022. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-

4287 (electronic). URL https://www.cambridge.org/core/journals/
canadian-mathematical-bulletin/article/noncircular-algebraic-
curves-of-constant-width-an-answer-to-rabinowitz/525D3FDBC6F2FDFC96A73088AD69DE32.

Tsukamoto:2022:CRR

[4769] Mayu Tsukamoto. Characterizations of right rejective chains. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 65(3): 557–570, September 2022. CODEN CMBUA3. ISSN 0008-4395 (print),

1496-4287 (electronic). URL https://www.cambridge.org/core/
journals/canadian-mathematical-bulletin/article/characterizations-
of-right-rejective-chains/DB8DB1D75B8F51CD0BB058E815C041A5.

Falco:2022:AFB

[4770] Javier Falcó and Paul M. Gauthier. Asymptotic first boundary value problem for elliptic operators. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 65(3):571–581, September 2022. CODEN CMBUA3. ISSN 0008-4395 (print),

1496-4287 (electronic). URL https://www.cambridge.org/core/
journals/canadian-mathematical-bulletin/article/asymptotic-
first-boundary-value-problem-for-elliptic-operators/8022DFD1A994B28C07C271427D1B8A2C.

Lazar:2022:PQM

[4771] Aldo J. Lazar and Douglas W. B. Somerset. Pure quotients and Morita’s theorem for k_{ω}-spaces. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 65(3):582–597, September 2022. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-

4287 (electronic). URL https://www.cambridge.org/core/journals/
canadian-mathematical-bulletin/article/pure-quotients-and-moritas-
theorem-for-komega-spaces/BEFA5B925481B43EC99EA2D772A5366B.

Eliahou:2022:SRF

[4772] Shalom Eliahou and Youssef Fares. Some results on the Flynn–Poonen– Schaefer conjecture. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 65(3):598–611, September 2022. CODEN CM-BUA3. ISSN 0008-4395 (print), 1496-4287 (electronic). URL https://

www.cambridge.org/core/journals/canadian-mathematical-bulletin/
article/some-results-on-the-flynnpoonenschaefer-conjecture/
45A7226AE913B666669D8B6CC05E304B.

Kadourek:2022:SCP

[4773] Jirí Kad’ourek. On semidirectly closed pseudovarieties of finite semigroups and monoids. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 65(3):612–627, September 2022. CODEN CM-BUA3. ISSN 0008-4395 (print), 1496-4287 (electronic). URL https://

www.cambridge.org/core/journals/canadian-mathematical-bulletin/
article/on-semidirectly-closed-pseudovarieties-of-finite-semigroups-
and-monoids/C97088F30E561DB1FDD6F5A20DC414D5.

Robert:2022:CAW

[4774] Leonel Robert. C^{∗}-algebras where every element is a limit of products of positive elements. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 65(3):628–632, September 2022. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-

4287 (electronic). URL https://www.cambridge.org/core/journals/
canadian-mathematical-bulletin/article/calgebras-where-every-
element-is-a-limit-of-products-of-positive-elements/688C6D7A19194D270DD864AEB0E8E35E.

Camps-Moreno:2022:PPQ

[4775] Eduardo Camps-Moreno, Craig Kohne, Eliseo Sarmiento, and Adam Van Tuyl. Powers of principal Q-Borel ideals. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 65(3):633–652, September 2022. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-

4287 (electronic). URL https://www.cambridge.org/core/journals/
canadian-mathematical-bulletin/article/powers-of-principal-
qborel-ideals/FE3F77B10E95F51A56522EC6A632FF8E.

Burke:2022:AOI

[4776] Maxim R. Burke. Analytic order-isomorphisms of countable dense subsets of the unit circle. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 65(3):653–664, September 2022. CODEN CM-BUA3. ISSN 0008-4395 (print), 1496-4287 (electronic). URL https://

www.cambridge.org/core/journals/canadian-mathematical-bulletin/
article/analytic-orderisomorphisms-of-countable-dense-subsets-
of-the-unit-circle/560EA525EED3876B548BD79C4DF4081C.

Azam:2022:UQH

[4777] Mahmud Azam and Ebrahim Samei. Uniformly quasi-Hermitian groups are supramenable. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 65(3):665–673, September 2022. CODEN CM-BUA3. ISSN 0008-4395 (print), 1496-4287 (electronic). URL https://

www.cambridge.org/core/journals/canadian-mathematical-bulletin/
article/uniformly-quasihermitian-groups-are-supramenable/CCA4A5E8A76830B33B5F9ED3827E39CC.

Sane:2022:DUB

[4778] Abdoul Karim Sane. On dual unit balls of Thurston norms. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 65(3): 674–685, September 2022. CODEN CMBUA3. ISSN 0008-4395 (print),

1496-4287 (electronic). URL https://www.cambridge.org/core/
journals/canadian-mathematical-bulletin/article/on-dual-unit-
balls-of-thurston-norms/092C22C67E8A31E42C4B655F73216147.

Petrov:2022:AFC

[4779] Victor Petrov and Simon W. Rigby. The Allison–Faulkner construction of E_{8}. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 65(3):686–701, September 2022. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic). URL https://www.

cambridge.org/core/journals/canadian-mathematical-bulletin/
article/allisonfaulkner-construction-of-e8/9326CC94236F22D285A091C8941D01C7.

Kleine:2022:FSG

[4780] Sören Kleine and Katharina Müller. Fine Selmer groups of congruent p-adic Galois representations. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 65(3):702–722, September 2022. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-

4287 (electronic). URL https://www.cambridge.org/core/journals/
canadian-mathematical-bulletin/article/fine-selmer-groups-of-
congruent-padic-galois-representations/41D8079C650AADE7852612A5E46C4028.

Dayan:2022:WSB

[4781] Alberto Dayan. Weakly separated Bessel systems of model spaces. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 65(3):723–742, September 2022. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic). URL https://www.cambridge.org/

core/journals/canadian-mathematical-bulletin/article/weakly-
separated-bessel-systems-of-model-spaces/91D6DACABA0413026E52CE6D3681C940.

Kodaka:2022:PGI

[4782] Kazunori Kodaka. The Picard groups of inclusions of C^{∗}-algebras induced by equivalence bimodules. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 65(3):743–758, September 2022. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-

4287 (electronic). URL https://www.cambridge.org/core/journals/
canadian-mathematical-bulletin/article/picard-groups-of-inclusions-
of-c-algebras-induced-by-equivalence-bimodules/48810D839B5C53A7EFF40CDF1E5D4E57.

Wallis:2022:NCS

[4783] Ben Wallis. A new complemented subspace for the Lorentz sequence spaces, with an application to its lattice of closed ideals. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 65(3):759–769, September 2022. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic). URL https://www.

cambridge.org/core/journals/canadian-mathematical-bulletin/
article/new-complemented-subspace-for-the-lorentz-sequence-

spaces-with-an-application-to-its-lattice-of-closed-ideals/
8B5A8400A397836E7EDDB69BE352596D.

Ceniceros:2022:SSS

[4784] Jose Ceniceros, Indu R. Churchill, and Mohamed Elhamdadi. Singquandle shadows and singular knot invariants. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 65(3):770–787, September 2022. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-

4287 (electronic). URL https://www.cambridge.org/core/journals/
canadian-mathematical-bulletin/article/singquandle-shadows-
and-singular-knot-invariants/9734175886AF943E8C83498053DEC49D.

Jakhar:2022:NNF

[4785] Anuj Jakhar and Surender Kumar. On nonmonogenic number fields defined by x^{6} + ax + b. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 65(3):788–794, September 2022. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-

4287 (electronic). URL https://www.cambridge.org/core/journals/
canadian-mathematical-bulletin/article/on-nonmonogenic-number-
fields-defined-by-x6axb/B624FA0AC34604E88005A4B2B83784A4.

Mizusawa:2022:MCF

[4786] Yasushi Mizusawa. On metabelian 2-class field towers over Z_{2}- extensions of real quadratic fields. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 65(3):795–805, September 2022. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-

4287 (electronic). URL https://www.cambridge.org/core/journals/
canadian-mathematical-bulletin/article/on-metabelian-2class-
field-towers-over-mathbb-z2-extensions-of-real-quadratic-fields/
BF99E228190328E2ED86CDBDD067058F.

Brits:2022:CSE

[4787] Rudi Brits, François Schulz, and Cheick Touré. Commutativity via spectra of exponentials. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 65(4):815–824, December 2022. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-

4287 (electronic). URL https://www.cambridge.org/core/journals/
canadian-mathematical-bulletin/article/commutativity-via-spectra-
of-exponentials/7B53DE6D50E53768702A5413DEAAF9A3.

Francoeur:2022:MSN

[4788] Dominik Francoeur and Anitha Thillaisundaram. Maximal subgroups of nontorsion Grigorchuk–Gupta–Sidki groups. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 65(4):825–844, December 2022. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-

4287 (electronic). URL https://www.cambridge.org/core/journals/
canadian-mathematical-bulletin/article/maximal-subgroups-of-
nontorsion-grigorchukguptasidki-groups/7000310A88BFA9AFAB5E5DEC9F37BAEE.

Tapp:2022:PBU

[4789] Kristopher Tapp. On pull-backs of the universal connection. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 65(4): 845–859, December 2022. CODEN CMBUA3. ISSN 0008-4395 (print),

1496-4287 (electronic). URL https://www.cambridge.org/core/
journals/canadian-mathematical-bulletin/article/on-pullbacks-
of-the-universal-connection/F25AE0A2C0FB8350AB227C79DBD0BB74.

Bondarenko:2022:SCX

[4790] Andriy Bondarenko, Andriy Prymak, and Danylo Radchenko. Spherical coverings and x-raying convex bodies of constant width. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 65(4):860– 866, December 2022. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-

4287 (electronic). URL https://www.cambridge.org/core/journals/
canadian-mathematical-bulletin/article/spherical-coverings-
and-xraying-convex-bodies-of-constant-width/B7E86F17BB41846A9E239AC829DA0AE4.

Motegi:2022:GTK

[4791] Kimihiko Motegi and Masakazu Teragaito. Generalized torsion for knots with arbitrarily high genus. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 65(4):867–881, December 2022. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-

4287 (electronic). URL https://www.cambridge.org/core/journals/
canadian-mathematical-bulletin/article/generalized-torsion-
for-knots-with-arbitrarily-high-genus/C653CF0F705F56BE07AFA891CABFD12D.

Cambie:2022:ITB

[4792] Stijn Cambie and Ross J. Kang. Independent transversals in bipartite correspondence-covers. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 65(4):882–894, December 2022. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-

4287 (electronic). URL https://www.cambridge.org/core/journals/
canadian-mathematical-bulletin/article/independent-transversals-
in-bipartite-correspondencecovers/D3C2B75086C881FFA69E994A34BD8DC6.

Chen:2022:SFA

[4793] Lu Chen, Guozhen Lu, and Maocun Zhu. A sharpened form of Adamstype inequalities on higher-order Sobolev spaces W^{m, \frac{n}{m}} (R^{n}): a simple approach. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 65(4):895–905, December 2022. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic). URL https://www.

cambridge.org/core/journals/canadian-mathematical-bulletin/
article/sharpened-form-of-adamstype-inequalities-on-higherorder-
sobolev-spaces-wmfrac-nmmathbb-rn-a-simple-approach/54994BF5FD7F6B80D709603F4AC1A6C5.

Pasten:2022:ADT

[4794] Hector Pasten. Arithmetic derivatives through geometry of numbers. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 65(4):906–923, December 2022. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic). URL https://www.

cambridge.org/core/journals/canadian-mathematical-bulletin/
article/arithmetic-derivatives-through-geometry-of-numbers/
773C65B65935B90A6F5E60D8570AD04C.

Mizuta:2022:TIR

[4795] Yoshihiro Mizuta and Tetsu Shimomura. Trudinger’s inequalities for Riesz potentials in Morrey spaces of double phase functionals on half spaces. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 65(4):924–935, December 2022. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic). URL https://www.

cambridge.org/core/journals/canadian-mathematical-bulletin/
article/trudingers-inequalities-for-riesz-potentials-in-morrey-
spaces-of-double-phase-functionals-on-half-spaces/A1D917297FC360DD56C09953D84B3FD3.

Bienvenu:2022:KTS

[4796] Pierre-Yves Bienvenu and François Hennecart. Kneser’s theorem in σ-finite abelian groups. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 65(4):936–942, December 2022. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-

4287 (electronic). URL https://www.cambridge.org/core/journals/
canadian-mathematical-bulletin/article/knesers-theorem-in-sigma-
finite-abelian-groups/C37CC828089B939B7ABD2EBD17F18462.

Bourin:2022:MIM

[4797] Jean-Christophe Bourin and Eun-Young Lee. Matrix inequalities and majorizations around Hermite–Hadamard’s inequality. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 65(4):943– 952, December 2022. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-

4287 (electronic). URL https://www.cambridge.org/core/journals/
canadian-mathematical-bulletin/article/matrix-inequalities-
and-majorizations-around-hermitehadamards-inequality/F55B26A0FBF16BAED086B317EE2947F9.

Sepulcre:2022:NSA

[4798] Juan Matías Sepulcre and Tomás Vidal. A note on spaces of almost periodic functions with values in Banach spaces. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 65(4):953–962, December 2022. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-

4287 (electronic). URL https://www.cambridge.org/core/journals/
canadian-mathematical-bulletin/article/note-on-spaces-of-almost-
periodic-functions-with-values-in-banach-spaces/F6C346F0C1FC021C724247DFD19AE2E8.

Hu:2022:CWT

[4799] Lian Hu, Songxiao Li, and Rong Yang. Corona and Wolff theorems for the multiplier algebra of Dirichlet–Morrey spaces. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 65(4):963– 975, December 2022. CODEN CMBUA3. ISSN 0008-4395 (print),

1496-4287 (electronic). URL https://www.cambridge.org/core/
journals/canadian-mathematical-bulletin/article/corona-and-
wolff-theorems-for-the-multiplier-algebra-of-dirichletmorrey-
spaces/B6B5E2E34EF8F0AA427CB6E301CABB02.

Frendreiss:2022:SIZ

[4800] William Frendreiss, Jennifer Gao, Austin Lei, Amy Woodall, Hui Xue, and Daozhou Zhu. Stieltjes interlacing of the zeros of j_{n}. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 65(4):976–993, December 2022. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic). URL https://www.

cambridge.org/core/journals/canadian-mathematical-bulletin/
article/stieltjes-interlacing-of-the-zeros-of-jn/D4A323146FF6FF677A129BDEBF291E66.

Berdellima:2022:NMC

[4801] A. Bërdëllima. A note on Mosco convergence in CAT(0) spaces. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 65(4): 994–1003, December 2022. CODEN CMBUA3. ISSN 0008-4395 (print),

1496-4287 (electronic). URL https://www.cambridge.org/core/
journals/canadian-mathematical-bulletin/article/note-on-mosco-
convergence-in-operatorname-mathrm-cat0-spaces/F4F7A3DB7AD9899846E5CC482C44F8B3.

Bao:2022:NSC

[4802] Jiguang Bao and Qiaoli Feng. Necessary and sufficient conditions on global solvability for the p-k-Hessian inequalities. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 65(4):1004–1019, December 2022. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-

4287 (electronic). URL https://www.cambridge.org/core/journals/
canadian-mathematical-bulletin/article/necessary-and-sufficient-
conditions-on-global-solvability-for-the-pkhessian-inequalities/
C5762CD8490BED48E8BEC89B8BC59D9E.

Allu:2022:OVA

[4803] Vasudevarao Allu and Himadri Halder. Operator valued analogues of multidimensional Bohr’s inequality. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 65(4):1020–1035, December 2022. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-

4287 (electronic). URL https://www.cambridge.org/core/journals/
canadian-mathematical-bulletin/article/operator-valued-analogues-
of-multidimensional-bohrs-inequality/4B3016C1558B9E95933A00616044E686.

Patra:2022:SCR

[4804] Dhriti Sundar Patra. Some characterizations of ρ-Einstein solitons on Sasakian manifolds. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 65(4):1036–1049, December 2022. CODEN CM-BUA3. ISSN 0008-4395 (print), 1496-4287 (electronic). URL https://

www.cambridge.org/core/journals/canadian-mathematical-bulletin/
article/some-characterizations-of-rho-einstein-solitons-on-
sasakian-manifolds/BCA61BA921E23DC4A88476C6CD1F364D.

Ray:2022:AGM

[4805] Anwesh Ray. Asymptotic growth of Mordell–Weil ranks of elliptic curves in noncommutative towers. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 65(4):1050–1062, December 2022. CODEN CMBUA3. ISSN 0008-4395 (print),

1496-4287 (electronic). URL https://www.cambridge.org/core/
journals/canadian-mathematical-bulletin/article/asymptotic-
growth-of-mordellweil-ranks-of-elliptic-curves-in-noncommutative-
towers/C462E9B117BBEECC3CD39999DFD00E57.

Fan:2022:UPC

[4806] Qingzhai Fan and Shan Zhang. Uniform property Γ for certain C^{∗}- algebras. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 65(4):1063–1070, December 2022. CODEN CMBUA3.

ISSN 0008-4395 (print), 1496-4287 (electronic). URL https://www.

cambridge.org/core/journals/canadian-mathematical-bulletin/
article/uniform-property-gamma-for-certain-mathrm-c-algebras/
731727B1E5BA04B242775BEABBD22E1A.

Tamiozzo:2023:SCM

[4807] Matteo Tamiozzo. Special curves in modular surfaces. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 66(1):1– 18, March 2023. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-

4287 (electronic). URL https://www.cambridge.org/core/journals/
canadian-mathematical-bulletin/article/special-curves-in-modular-
surfaces/E327F297E517F193F470C7D305DA41A6.

Donovan:2023:HSH

[4808] Connor Donovan, Maxwell Lin, and Nicholas A. Scoville. On the homotopy and strong homotopy type of complexes of discrete Morse functions. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 66(1):19–37, March 2023. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic). URL https://www.

cambridge.org/core/journals/canadian-mathematical-bulletin/
article/on-the-homotopy-and-strong-homotopy-type-of-complexes-
of-discrete-morse-functions/8AC564478FFC7FE18AA721021D03B998.

Zafrullah:2023:RDD

[4809] M. Zafrullah. Revisiting G-Dedekind domains. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 66(1): 38–52, March 2023. CODEN CMBUA3. ISSN 0008-4395 (print),

1496-4287 (electronic). URL https://www.cambridge.org/core/
journals/canadian-mathematical-bulletin/article/revisiting-
gdedekind-domains/18655B4198F69FA3CE3CEFFDD6F50471.

Itaba:2023:QPP

[4810] Ayako Itaba and Izuru Mori. Quantum projective planes finite over their centers. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 66(1):53–67, March 2023. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic). URL https://www.

cambridge.org/core/journals/canadian-mathematical-bulletin/
article/quantum-projective-planes-finite-over-their-centers/
0616A02CF005FE80823B0CF3BA9B516F.

Lilienfeldt:2023:TPM

[4811] David T.-B. G. Lilienfeldt. Torsion properties of modified diagonal classes on triple products of modular curves. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 66(1):68– 86, March 2023. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-

4287 (electronic). URL https://www.cambridge.org/core/journals/
canadian-mathematical-bulletin/article/torsion-properties-of-
modified-diagonal-classes-on-triple-products-of-modular-curves/
EC1562E54BF8774BA33E229F4FED23CF.

Hernandez-Gutierrez:2023:CPR

[4812] Rodrigo Hernández-Gutiérrez and Alfredo Zaragoza. A characterization of the product of the rational numbers and complete Erdős space. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 66(1):87–102, March 2023. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic). URL https://www.

cambridge.org/core/journals/canadian-mathematical-bulletin/
article/characterization-of-the-product-of-the-rational-numbers-
and-complete-erdos-space/70B7DE324FCE062EA0E09F0C868CB849.

Dostert:2023:PHS

[4813] Maria Dostert and Alexander Kolpakov. Packable hyperbolic surfaces with symmetries. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 66(1):103–113, March 2023. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic). URL https://www.

cambridge.org/core/journals/canadian-mathematical-bulletin/
article/packable-hyperbolic-surfaces-with-symmetries/DA6F6D3379BD57AD029A68B6846D6C35.

Meyrath:2023:UFM

[4814] Thierry Meyrath. On unavoidable families of meromorphic functions. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 66(1):114–123, March 2023. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic). URL https://www.

cambridge.org/core/journals/canadian-mathematical-bulletin/
article/on-unavoidable-families-of-meromorphic-functions/6D4D3B92AFA7515EF8D89F2E10855D44.

Ulivelli:2023:GKT

[4815] Jacopo Ulivelli. Generalization of Klain’s theorem to Minkowski symmetrization of compact sets and related topics. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 66(1):124–141, March 2023. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-

4287 (electronic). URL https://www.cambridge.org/core/journals/
canadian-mathematical-bulletin/article/generalization-of-klains-
theorem-to-minkowski-symmetrization-of-compact-sets-and-related-
topics/5E7BE2F4B93FC0F54C29DD8ABA6E0060.

Yattselev:2023:SPC

[4816] Maxim L. Yattselev. On smooth perturbations of Chebyshëv polynomials and ¯∂-Riemann–Hilbert method. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 66(1):142–155, March 2023. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287

(electronic).  URL https://www.cambridge.org/core/journals/
canadian-mathematical-bulletin/article/on-smooth-perturbations-
of-chebyshev-polynomials-and-bar-partial-riemannhilbert-method/
3DBB296E42C6916DCE7C6DC162C89066.

Kubis:2023:TRT

[4817] Wiesław Kubi’s and Paul Szeptycki. On a topological Ramsey theorem. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 66(1):156–165, March 2023. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic). URL https://www.

cambridge.org/core/journals/canadian-mathematical-bulletin/
article/on-a-topological-ramsey-theorem/F9BA28C498082E1DFF0C925BE4EF5210.

Newton:2023:IST

[4818] Alexis Newton and Jeremy Rouse. Integers that are sums of two rational sixth powers. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 66(1):166–177, March 2023. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic). URL https://www.

cambridge.org/core/journals/canadian-mathematical-bulletin/
article/integers-that-are-sums-of-two-rational-sixth-powers/
36425CBE1A7009E8FCFCE0D383B1BE3F.

Monavari:2023:LDS

[4819] Sergej Monavari and Andrea T. Ricolfi. Sur la lissité du schéma quot ponctuel emboîté. (French) [On the smoothness of the embedded point quot schema]. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 66(1):178–184, March 2023. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic). URL https://www.

cambridge.org/core/journals/canadian-mathematical-bulletin/
article/sur-la-lissite-du-schema-quot-ponctuel-emboite/7DEE062EC8DCE202FFC47CEB35CDFD6B.

Johnston:2023:AVP

[4820] Daniel R. Johnston. On the average value of π(t) − li(t). Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 66(1):185– 195, March 2023. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-

4287 (electronic). URL https://www.cambridge.org/core/journals/
canadian-mathematical-bulletin/article/on-the-average-value-
of-pi-toperatorname-textrm-lit/EF10526C74AF1817EB3B798E2FA9EBD2.

Azimi:2023:HRM

[4821] Y. Azimi. Hilbert rings with maximal ideals of different heights and unruly Hilbert rings. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 66(1):196–203, March 2023. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic). URL https://www.

cambridge.org/core/journals/canadian-mathematical-bulletin/
article/hilbert-rings-with-maximal-ideals-of-different-heights-
and-unruly-hilbert-rings/1C2907278C1F20FBB7DE15489049B11C.

Muthukumar:2023:MSI

[4822] P. Muthukumar and Jaydeb Sarkar. Model spaces invariant under composition operators. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 66(1):204–217, March 2023. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic). URL https://www.

cambridge.org/core/journals/canadian-mathematical-bulletin/
article/model-spaces-invariant-under-composition-operators/
AE8FC2883C59A4138225666AF8B48D80.

Kojin:2023:RNO

[4823] Kenta Kojin. A refined nc Oka–Weil theorem. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 66(1): 218–224, March 2023. CODEN CMBUA3. ISSN 0008-4395 (print),

1496-4287 (electronic). URL https://www.cambridge.org/core/
journals/canadian-mathematical-bulletin/article/refined-nc-
okaweil-theorem/CD83B72C4AF3BFC70629F10604DB5B63.

Balaji:2023:TLS

[4824] Nikhil Balaji and Florian Luca. Terms of Lucas sequences having a large smooth divisor. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 66(1):225–231, March 2023. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic). URL https://www.

cambridge.org/core/journals/canadian-mathematical-bulletin/
article/terms-of-lucas-sequences-having-a-large-smooth-divisor/
128ECA19406A6BD8438863DB92C5838D.

Bredon:2023:HCG

[4825] Naomi Bredon. Hyperbolic Coxeter groups of minimal growth rates in higher dimensions. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 66(1):232–242, March 2023. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic). URL https://www.

cambridge.org/core/journals/canadian-mathematical-bulletin/
article/hyperbolic-coxeter-groups-of-minimal-growth-rates-in-
higher-dimensions/358207F12A49F9C2D97547D04E53C74D.

Griffin:2023:LBD

[4826] Michael Griffin, Ken Ono, Larry Rolen, and Wei-Lun Tsai. Limiting Betti distributions of Hilbert schemes on n points. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 66(1):243– 258, March 2023. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-

4287 (electronic). URL https://www.cambridge.org/core/journals/
canadian-mathematical-bulletin/article/limiting-betti-distributions-
of-hilbert-schemes-on-n-points/DC9492926095DD32BBCC5D3FE6D2B4F0.